A Complete Bibliography of Publications in the Journal of Computer and System Sciences

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

30 September 2025 Version 1.30

Title word cross-reference

#09111 [1814]. #AM01053M [1862, 1877, 1904]. #better 5X8 els [1856]. #better 5X8 els.pdf [1856]. #BIS [3184]. #BIS-hardness [3184]. #CSP [2696, 2908, 2500, 3508]. #econdirectM [1893]. #HA03022M [1860, 1875]. #HA05062N [1876]. #HA05062O [1849, 1861]. #HA06043M [1892, 1871, 1889]. #HA08111M [1846]. #P [1195, 3598, 1261, 1264]. #praise 5X8 els.pdf [1832]. #sciencejobs N [1802].

#P [1373]. (1 + 1/ dim ϵ) [1590]. (1 ± ϵ) [3581]. (2 − 7/(3p + 1))· [2414]. (2^{m}) [1306]. (s, t) [1475]. 0(n) [1160]. 1 [270]. 1 − L [1371, 924]. 12n^{2} [450]. 2

[3525, 3184, 3798, 2191, 1048, 3402, 1500, ; 3364, 2034, 2993, 834, 2473, 3101]. {2, 3} ; [2843]. 2 − ϵ [2353]. 2^{0.802n} [3693]. 2^{2^{2^{pn}}} ; [445]. 2^{O(n^{1/3})} [1977]. 3 [1920, 3773, 2242, ; 2827, 2374, 1961, 2825, 3240, 1071, 1133]. ; 3 + 1 [3895]. 4 + ϵ [3670]. 43 [3620]. ; 4 · Log_{2}N [655]. 5/4 [1612]. 8 [2998]. =? ; [1433]. ^{∗} [858]. ^{0} [1625, 3418, 1085]. ^{1} [1523]. ; ^{∗} [3194]. ^{NP} [2235]. ^{NP [O(log n)]} [995]. _{2}^{p} ⊆ ; [2235]. ^{P H} [1501]. _{2} [3418]. _{H} [1565]. _{p} ; [3693]. a [426, 289]. A[P (x), 2x, x + 1] [500]. ; Ax = λBx [377]. b [3253]. β [3444]. · [3418]. ; d [2527, 3843, 3854, 1362, 3256, 3563]. δ ; [3553]. ℓ_{p} [2154]. ϵ [1551, 2617, 1864, 1693]. ; g [3858, 3937, 3312]. H ; [3920, 3727, 3701, 1019]. K [1320, 2756, 191, ; 3494, 1300, 1546, 3286, 2357, 3085, 3869, ; 3590, 1788, 2535, 1940, 199, 2491, 3924, 3803, ; 3256, 1188, 3603, 3754, 961, 3252, 3394, 1681,

1398, 2805, 3506, 3670, 630, 2690, 1339, 585]. L [541, 433, 314]. L = {a^{2i}|; i ≥ 0} [772]. L_{∞} [3358, 1736, 2474]. L_{p} [3344]. λ [2997]. log^{∗} n + O(1) [1735]. log n [473]. M [708, 1294]. N^{n} [174]. Z_{m} [2157, 704]. Zm [1196]. Z_{m} [1486]. F [3629, 3878]. μ [3127]. N [1293, 655, 1433, 2780, 3717, 2135, 1792, 229, 895]. n^{2} [2135]. N^{m} [1597]. n log n [751]. O(EV log^{2} V) [566]. O(log^{2} n) [847]. O(log^{3/2} n) [1422]. O(LogLogN) [960]. O(log n) [2370]. O(M(n^{2})) [1488]. O(n) [1359, 3350]. O(n^{1.5}\sqrt{log n}) [938]. O(n^{log log n}) [1268]. O(n log log n) [1311]. O(n log n) [1864, 1359, 1430]. O(|V | log |V |) [1475]. O_{n} [3717]. ω [2678, 3196, 242, 416, 417, 443, 909, 722]. ω^{2} [298]. ω^{n} [460]. p [3367, 3758, 1141, 1294, 1359]. P F · #P [1373]. Q [3470, 2918]. q = 7 [38]. r [3343]. s [1427, 585]. S^{3/2} [1565]. Σ_{3}^{p} [1552]. Σ_{2} [895]. st [3350]. ⊆ [1565]. t [1427, 3205]. T^{ω} [469]. Θ(n log^{2} n) [1426]. V log V [213]. W [2] [1940]. X + Y [641]. Z^{2} [3101].

* [159].

-acyclic [3444]. -anonymous [3085]. -Approval [3343]. -approximation [2191, 3581, 3620]. -arc [3252]. -armed [2756]. -ary [1792]. -automata [909]. -Biased [1551]. -bit [2998]. -bound [2135]. -cheatable [1141]. -chromatic [3253]. -closeness [1141]. -cluster [3563]. -Clustering [3603]. -CNFs [2357]. -colorable [1920]. -coloring [3869]. -Complete [1371, 1940, 1552]. -complex [1019]. -Computations [443]. -connected [834, 3670]. -connectivity [1133]. -construction [708]. -context-free [416]. -cores [3754]. -D [3525]. -Depth [1320]. -Dimensional [1339, 2780, 1362]. -distinct [3394]. -Edge [3758]. -Edge/vertex-connected [3758].

-extendible [2997]. -extra [3858, 3312]. -Extraconnectivities [2843]. -Feature [1940]. -Flows [1475]. -Free [1693, 2374, 1864, 3920, 3727, 3878, 3701]. -good-neighbor [3937]. -Group [1359]. -hard [1085]. -Hitting [2527]. -hub [3367]. -hyperbolic [3553]. -internal [2535]. -languages [416, 417, 722]. -Leaf [2491]. -letter [2135]. -Limit [3127]. -Majority [3494]. -MCFL [3101, 3717]. -Means [2805]. -Median [1788]. -min-wise [3256]. -Minor-Free [3629]. -MST [1546]. -Nilpotent [1048]. -norm [2474, 3344]. -numbering [3350]. -page [961]. -parallel [229]. -party [3798]. -path [3843]. -Process [655]. -query [2034]. -random [2617]. -regular [3773, 2678, 3196, 3854]. -Resolution [3470]. -SAT [2242, 2473, 2357, 1681]. -secluded [3803]. -server [3286, 1188]. -space [1961]. -spin [3184]. -subcolorable [3364]. -subsequence [3924]. -systems [298]. -tapes [242, 460]. -Term [1300, 3506, 2690]. -Testability [1500]. -transducers [426, 289]. -trees [858]. -tuple [2993]. -type [417]. -Uniform [3205]. -valued [655]. -vertex [2535]. -Way [270].

/= [1461].

0 [535].

1 [535]. 1-center [3190]. 1-convexes [174]. 1-median [3248, 3190]. 1.375-approximation [3564]. 1.5-approximation [2080]. 1.75-approximation [2305]. 1/2 [2236]. 1971 [85]. 1994 [1378]. 1Heads [942]. 1st [3829].

2 [3171, 2671, 2522]. 2-CSP [3171, 2671]. 2-Sat [2671]. 2013 [3301]. 2015 [3465]. 2017 [3576]. 2020 [3542, 3567, 3537, 3520, 3528, 3556, 3548].

2021 [3622, 3658, 3578, 3613, 3594, 3599, 3651, 3637]. 2022 [3713, 3738, 3666, 3703, 3680, 3695, 3730, 3722]. 2023 [3808, 3766, 3743, 3760, 3749, 3755, 3792, 3777]. 2024 [3842, 3868, 3816, 3837, 3825, 3831, 3861, 3857]. 2025 [3907, 3927, 3875, 3888, 3894, 3917, 3914]. 2026 [3942]. 20th [3301]. 21st [3576]. 22nd [3647]. 2CNF [3781]. 2D [1230, 3412]. 2nd [2525].

324-332 [3859]. 35th [2029]. 3rd [2525].

46th [3610].

56 [3817]. 571 [1605].

65 [2926]. 66 [2185].

73 [2359]. 74 [2469]. 75 [2501]. 76 [3859]. 78 [3109].

81 [3187]. 82 [3199]. 88 [3933].

9th [2971, 3465, 1378].

= [1461, 978].

A- [466]. A-transducers [558]. A.E. [1019]. abduction [2534]. ABE [3322]. abelian [371, 2297, 3258, 1359]. Ability [1547, 50]. abortion [3115]. above [2552, 2585, 2659, 2442, 2997]. Absolute [3734, 3451]. Absolutely [179]. absorbing [2046]. Abstract [126, 1742, 3425, 2713, 151, 1961, 306, 92, 357, 617, 264, 1060, 347, 2510, 845, 885, 583, 862, 31]. Abstraction [1698, 881, 700]. AC [1625, 3418, 1085]. ACC [1161]. accelerated [325]. acceptance [397]. accepted [1203]. accepting [514, 1095]. acceptors [115, 114, 104, 249, 168, 69]. Access [1366, 3280, 3432, 217, 3116, 3805, 2103, 743, 3005, 3746, 762, 3007]. Accessibility [634]. account [2975]. accrual [2823]. accumulations [3425]. accuracy [1965, 3862, 2482, 2210, 2111, 2807]. Accurate [2662, 1808]. achieve [2881]. achieving [832]. Achilles [1535]. ACM [274, 345, 2029, 214]. acoustic [2809]. across [2966]. Action [2602, 2657, 544, 31]. action-based [2657]. actions [908]. activation [3250, 2051]. Active [1531, 3804, 2436, 3008, 3734, 3237]. ACTL [1688]. actor [1132]. actualization [893]. Acyclic [3920, 717, 1494, 3363, 3444, 3880, 2911, 3390, 2638, 2540]. acyclicity [972]. Ad [1893, 1892, 1904, 3021, 2209, 2875, 2870, 2537, 3143]. ad-hoc [3143]. Adaps [2574]. adapt [3664]. Adapting [3871, 3391]. Adaptive [1631, 1744, 2576, 2404, 1275, 2332, 2963, 3019, 2338, 3546, 3757, 2772, 3734, 2818, 2574, 3233, 2767, 3152, 1835, 2337, 2400, 2407, 2816, 3052]. Adaptivity [2572]. adapts [2754]. adder [2998]. adder/subtractor [2998]. adding [3504]. Addition [3783, 2788, 546, 3203, 3197, 826]. Additional [1750, 3254]. Additive [3607]. additively [3798]. address [2464]. Addressable [456]. Addressing [573, 129]. Adequacy [574]. Adherences [547]. adjacencies [3707]. adjoint [25]. adjunct [235, 291]. administration [3007]. Admissible [1361]. admission [2252, 2740]. admission-control [2252]. admit [3656]. ADMs [2441]. advance [2801]. Advanced [3236, 2880, 2868, 2827]. Advances [3230, 3454, 2600]. Advantage [1292]. Adversarial [1631, 3432, 3805, 2755]. Adversary [2509]. Advertisement [1817, 1802, 1814, 1829, 1936, 1862, 1877, 1907, 1908, 1846, 1849, 1860, 1861, 1875, 1876, 1832, 1952, 1856, 1871, 1889, 1955, 1956, 1957, 1958, 1970]. Advice [1262, 1495, 3849, 3286, 2497]. Aerial [3461]. AFA [196]. affect [514]. Affine [23, 1822]. AFL [264, 99, 139, 196, 292, 289, 309]. AFLs [115, 114, 781, 359, 133]. AFPTAS [3250]. after [825]. again [3476]. against [3231, 2563, 2285, 1557, 813, 2929, 3431]. Agent [3049, 3082, 3610, 2346, 2345]. Agent-based [3049]. agents [2448, 3456, 3580, 3934, 3663, 3504, 3055, 2674, 3678]. Aggregate [1769, 1463, 2641, 1840]. aggregating [2423]. Aggregation [1410, 3551, 2258, 2750, 2855, 1879, 2993, 2768, 2796, 2662]. Agnostic [2436, 1670, 2753]. Agreement [3417, 3092, 2658, 836, 2439, 2371, 2615, 2906]. agreements [1866]. Aided [1617]. airports [2341]. alert [2935, 2930]. Algebra [1573, 1685, 1466, 1354, 1771, 118, 863, 833, 905, 2402, 1110, 337, 553, 2270, 977, 699, 510, 899]. Algebraic [3790, 906, 884, 111, 1150, 9, 1443, 148, 1988, 602, 420, 678, 709, 653, 1066, 378, 1915, 716, 3735, 617, 448, 609, 5, 668, 953, 1012, 1227, 879, 862]. Algebras [1307, 1528, 715, 1356, 635, 376, 1912, 1082, 3154, 3039, 866, 729, 1101, 634, 710, 900, 695]. algébriques [378]. Algorithm [1293, 1556, 1311, 1612, 1546, 1634, 1788, 1583, 1576, 1320, 2535, 1367, 1392, 1473, 1789, 1504, 1268, 259, 1386, 1726, 1359, 1795, 2527, 3943, 2424, 3130, 3796, 2480, 3869, 3019, 3669, 2242, 779, 956, 2275, 3336, 1913, 2305, 749, 2772, 3157, 3922, 3171, 3423, 2209, 2062, 3702, 545, 557, 789, 566, 2553, 2142, 2883, 2901, 2671, 2938, 1943, 1063, 2080, 213, 2752, 2105, 605, 444, 3606, 3736, 2423, 1213, 187, 2695, 2330, 3027, 944, 2378, 3505, 2371, 2805, 3620, 2703, 2723, 2160, 2086, 531, 3563, 1220, 2668, 3564, 2475, 3417, 1184, 751, 3908, 2005, 1136, 2079, 3066, 3570, 3113, 3421]. algorithm [2807, 823, 1511]. Algorithmic [3319, 3829, 3395, 3677, 1362, 122, 859, 1486, 1451, 3124, 3698, 2145, 2886]. Algorithms [1497, 1744, 1249, 1521, 1317, 1298, 1299, 1751, 1560, 3643, 1598, 1644, 1600, 446, 2986, 48, 3434, 1290, 200, 809, 1368, 1393, 1446, 1547, 1245, 1429, 2934, 3881, 1685, 1306, 2869, 3169, 495, 1040, 1221, 2452, 385, 3735, 2462, 1947, 2560, 3668, 1949, 2168, 2418, 3409, 3311, 1898, 3478, 2491, 2020, 3742, 2394, 2496, 3889, 2132, 3758, 1879, 2669, 1232, 1188, 2191, 2701, 2433, 2497, 950, 1192, 807, 2149, 1983, 2069, 792, 2563, 3380, 2165, 849, 2228, 1926, 2044, 3393, 3205, 265, 758, 3789, 3878, 276, 251, 2121, 1053, 3364, 827, 1164, 113, 2564, 2297, 3607, 3289, 3258, 2509]. algorithms [2134, 2309, 1811, 2474, 2768, 935, 718, 343, 488, 2388, 1222, 205, 2586, 3536, 2166, 941, 3879, 3215, 338, 766, 3330, 492, 1812, 3190, 3056, 1052, 2540]. Aligner [2779]. Alignment [2309, 2298]. All-Pairs-Shortest-Path [1303]. allocation [956, 3801, 2767, 3156, 3015, 618, 2628, 2406]. allowing [863]. Almost [1920, 3796, 2637, 2610, 1713, 1093, 2473, 3901, 3493, 990, 3702, 3899]. Almost-natural [2610]. alpha [2505]. alpha-equivalence [2505]. alphabet [3903, 2135, 310, 1945, 3288]. Alternating [1352, 3226, 974, 909, 1152]. alternation [925, 802, 567]. Alternative [1560]. alternator [2133]. AM [2201]. AM01053LMS.zip [1877, 1904]. Amalgamation [886, 409]. Ambient [2713, 3048]. Ambients [2601]. Ambiguity [235, 323, 1993, 1014]. ambiguous [3718, 237]. among [561, 1042, 311, 421, 2767, 760, 41]. Amoroso [468]. Amortization [1478]. amount [3768]. ample [3715]. amplification [2563, 2014]. Amsterdam [1378]. Analog [1418, 3448]. Analysing [2402]. Analysis [3119, 868, 1598, 1558, 442, 275, 1704, 1835, 1650, 1669, 2936, 1306, 485, 2916, 2657, 2343, 1040, 3104, 2894, 3943, 3032, 2299, 3401, 3033, 2467, 1960, 76, 2623, 2174, 2209, 3060, 3061, 2062, 3067, 2207, 3236, 902, 2856, 874, 2319, 444, 2330, 9, 3822, 1811, 569, 801, 3545, 2265, 106, 2316, 2871, 373, 2900, 2397, 2250, 1140, 835, 826, 2314, 3328, 2427, 2819, 2893, 2605, 2515, 2345, 3582, 3653].

Analytic [2401, 3648]. analytical [2828, 2207, 2622, 3025]. Analytics [3175, 3321, 3179]. analyze [2462]. anaphora [2920]. anarchy [1916]. ancestors [1178]. ancestral [1804]. Ancient [1299]. AND-OR-NOT [2984]. and/or [2874, 3884, 3381]. annealing [2160]. annotated [2593]. annotation [2306, 2728, 2732]. annotations [1805]. Announcement [85, 135, 636, 784]. announcements [2974]. Annual [274, 162, 345, 2029, 214, 1378, 250]. anomaly [3348, 3233]. anonymity [2615]. anonymization [3012, 2967]. anonymous [3085, 3351, 3676, 3911, 3009]. answer [905, 2655]. Answering [1543, 3135, 3136, 3310, 3933, 2653, 2654, 2597, 3137]. answers [2326, 1064, 2367, 2860]. ant [2883, 2768, 3066]. anti [3166]. anti-piracy [3166]. antimonotone [2797]. Antler [3844]. Any [1578, 762]. AOCD [2345]. AOCD-based [2345]. aperiodic [2252, 3262]. Apex [1695]. Appear [1629, 1640, 1646, 1654, 1664, 1673, 1683, 1691, 1699, 1708, 1718, 1728, 1739, 1749, 1762, 1773, 1781, 1790, 1801, 1813, 1827, 1845, 1855, 1870, 1887, 1903, 1921, 1932, 1950, 1969, 1984, 1995, 2006, 2015, 2026, 2039, 2052, 2065, 2075, 2087, 2099, 2107, 2115, 2125, 2138, 2150, 2161, 2170, 2186, 2195, 2217, 2230, 2244, 2253, 2266]. Application [2525, 1478, 1614, 1455, 1564, 3866, 3051, 3467, 2306, 2222, 2535, 2272, 2355, 2583, 1807, 2707, 3109, 2796, 1835, 2826, 852, 181, 769, 2454, 112, 2608, 2882, 3908, 3449]. application-layer [3467]. Applications [2525, 1317, 1296, 3677, 1643, 1723, 1474, 3465, 1348, 809, 1248, 1499, 1724, 1326, 3004, 940, 3832, 2193, 1221, 3124, 1166, 2713, 3398, 2518, 3350, 82, 2263, 104, 2704, 320, 2751, 10, 2069, 3671, 3331, 2175, 3392, 2937, 866, 2212, 459, 3107, 2691, 3011, 2767, 2980, 1811, 3403, 2197, 2822, 2400, 1127, 3093, 2730, 2605, 2885, 2600, 2321]. Applied [1005, 118, 2286, 3114, 580]. apportionment [3939]. Approach [1441, 2633, 1247, 1467, 1696, 1407, 2916, 2313, 2709, 2657, 128, 2000, 2495, 3728, 2404, 3385, 3085, 2660, 242, 2944, 2790, 2517, 351, 3655, 2866, 1117, 2733, 3293, 2901, 2728, 304, 3231, 2485, 3202, 2585, 2735, 2285, 2405, 864, 3088, 3604, 3324, 2707, 3109, 256, 2661, 181, 1003, 949, 612, 2832, 396, 2074, 3168, 2896, 1071, 2626, 3391, 3010, 2341, 2834, 2967]. approaches [3931, 3065]. Approval [3343]. Approximability [1519, 1440, 3367, 2700, 2110, 1931, 3683]. approximable [3416]. Approximate [1428, 3876, 212, 3693, 3938, 1736, 1853, 2692, 3535, 1064, 3609, 3495, 2304, 2432, 3339, 3351, 2353, 2197, 2358]. Approximated [1468]. Approximately [3211, 1866, 2758, 1913, 3678]. Approximating [1549, 1534, 2045, 1590, 3429, 3699, 1632, 3588, 2791, 3035, 2290, 941, 2458, 2021, 3034, 1912, 2561, 2038, 2154, 1823, 3193, 2638]. Approximation [1554, 1748, 1751, 2390, 1546, 2379, 1634, 1788, 2168, 2020, 3742, 2496, 1983, 1457, 3205, 276, 1256, 2134, 1245, 2166, 1731, 3169, 1910, 3370, 2402, 3409, 3247, 2305, 2683, 3410, 2500, 3758, 2178, 3702, 2669, 3665, 3375, 2191, 3712, 2553, 2142, 1926, 2080, 1109, 3606, 3736, 3607, 1852, 3892, 2378, 2391, 3620, 2703, 2723, 2307, 3581, 3675, 3670, 1075, 2200, 26, 109, 3564, 141, 3269, 177, 95, 497]. Approximations [1459, 1644, 1514, 268, 1243, 3788, 3130, 3473, 122, 2857, 456]. arbitrary [143, 1028, 2038, 3909, 2279, 3911, 761, 102, 153]. Arborescences [1247, 3771]. Arc [1506, 3840, 3360, 2640, 3252, 1805]. arc-disjoint [3840]. Architecture [1367, 1234, 2822, 2772, 2827, 2824, 2906]. Architecture-Independent [1367]. Architectures [1367, 3014]. arcs [3724]. Area [796, 588, 3510]. Area-period [796].

Area-time [588]. argument [3639, 2034, 1191]. Arguments [1775, 940, 923]. arising [3689]. Arithmetic [1625, 1589, 1491, 1377, 1675, 3697, 3820, 265, 500, 2690, 445]. Arithmetical [1535, 3763, 719]. Arity [1461, 1529]. armed [2756]. arms [827]. Armstrong [852]. arrangement [963]. arrangements [1911, 1163]. Array [1234, 1386, 537, 915]. Arrays [1387, 846, 975, 743, 395]. arrival [3909, 2216]. Arrow [1725]. art [2560]. Arthur [2494, 919, 1596]. Article [1605]. artifacts [2506, 3282]. Artificial [2317]. ary [1792]. AS004M [1908]. AS004M.pdf [1908]. AS007M [1907]. AS007M.pdf [1907]. AS008M [1922]. AS008M.pdf [1922]. Asking [1335]. Aspects [1486, 3463, 2814, 859, 46, 3400, 3052]. assays [2085]. assembler [58]. assemblers [369]. assemblies [3791]. assembly [3804, 324, 3791, 2912]. assertion [326]. assertions [290, 2991]. Assigning [3587]. assignment [3376, 3122, 227, 3118, 3054]. assignments [3496, 2758, 3506]. assisted [3323, 3044]. associated [225]. Association [1539, 3323, 2902, 2484]. Associative [424, 1614, 1578, 790, 1993, 1158]. assuming [2105]. Assumption [1738, 2597]. Assumptions [1462, 1714]. assurance [2769]. Asymmetric [1733, 1512, 2809, 2213]. asymmetry [3788]. Asymptotic [399, 3223, 3410, 3254]. Asymptotically [1663, 692]. Asynchronous [1758, 1967, 1051, 3511, 43, 127, 2072, 390, 606, 373, 3387, 2250, 470, 2227, 2684, 2516]. asynchronously [3787]. Asynchrony [1292]. AT-free [3688]. ATL [3308]. atom [3631]. atomic [3063]. Atomicity [1467, 2105]. atoms [3195]. Attack [3581, 3012]. attackers [3391]. attacking [3516]. attacks [3234, 2402, 2940, 2285, 2936, 2939, 3329, 3431]. attainability [999]. attained [19]. Attention [1503].

Attraction [1614, 1690]. attractor [3220]. Attractors [1574]. Attribute [1618, 1642, 1505, 2073, 1613, 1896, 646, 2993, 2909]. Attribute-Efficient [1505, 1613, 1896]. Attributed [277]. Attributes [1237, 1539, 2299, 890, 2392, 1843, 2584, 508]. auction [2050]. auctions [2652]. Auditing [1843, 2392, 2889]. augmentation [2378, 581, 894, 1133]. Augmented [1698, 192]. August [85, 3460, 3542, 3622, 3713, 3766, 3842, 3907]. authenticated [2198, 2658, 2694]. authenticating [2931]. Authentication [1656, 592, 3114, 2795, 2182, 2199, 2377, 2906, 2416]. Author [28, 53, 78, 116, 155, 195, 233, 257, 282, 307, 333, 360, 386, 407, 429, 455, 481, 507, 529, 552, 575, 603, 627, 651, 671, 696, 723, 748, 775, 800, 820, 843, 869, 889, 907, 931, 954, 979, 1000, 1020, 1038, 1057, 1080, 1106, 1128, 1145, 1167, 1193, 1233, 1272, 1314, 1358, 1404, 1444, 1482, 1508, 1537, 1579, 1606, 1641, 1665, 1700, 1740, 1782, 1828, 1888, 1951, 2007, 2053, 2100, 2231]. authorisation [2745, 2962]. authoritative [3431]. authorization [2657, 2963]. Authorship [3114]. Automata [2234, 162, 1575, 3465, 3610, 1516, 1693, 1698, 1721, 1486, 1574, 1799, 1494, 2259, 250, 830, 679, 401, 89, 3226, 1900, 2780, 794, 3467, 3468, 75, 3782, 908, 3125, 3352, 301, 3660, 3718, 3440, 3589, 207, 1114, 3862, 2947, 3127, 2903, 354, 3530, 3800, 3382, 2798, 868, 1122, 119, 3745, 946, 242, 460, 2533, 3608, 929, 3903, 1009, 1135, 3625, 50, 3662, 329, 537, 1230, 1994, 3574, 351, 3763, 1203, 3067, 3407, 33, 388, 822, 2724, 355, 1864, 258, 10, 640, 2131, 3355, 64, 2413, 144, 484, 317, 515, 121, 204, 644, 321, 12, 30, 802, 125, 197]. automata [245, 2665, 459, 704, 3481, 3583, 166, 1102, 2686, 1174, 910, 2998, 210, 2923, 3359, 892, 2776, 782, 2072, 623, 2389, 3095, 2707, 3109, 909, 490, 2762, 2271, 260, 486, 211, 366, 461, 509, 3437, 2393, 2872, 3527, 534, 703, 74, 295, 3295, 2706, 2410, 2469, 339, 2454, 483, 3290, 153, 26, 2837, 1200, 470, 480, 415, 467, 20, 228, 952, 1215, 414, 1196, 219, 777, 13, 371, 565, 149, 171, 699, 286, 2227, 1087, 2546, 3547, 363, 687, 21, 2684, 419, 302, 505, 409, 2725, 578, 3288, 3918, 335, 90, 3826]. automata-based [3467]. Automata-Theoretic [1698, 830]. Automated [3467]. Automatic [2578, 2952, 3621, 1927, 966, 2657, 3185, 2782, 2950, 3783]. automatically [2866]. Automaticity [1360]. automating [2341]. Automaton [374, 1755, 3935, 394, 337, 270, 3042, 224]. Automorphism [910]. Automorphisms [509]. Autonomic [2971]. Autonomous [2767, 144]. Autoreducibility [2282]. Auxiliary [1799, 515, 125, 614, 3322]. available [2322]. Average [1962, 2614, 1056, 2176, 2887, 1620, 1338, 1920, 3493, 1092, 661, 2659]. Average-Case [1620, 2887]. AVL [1661]. avoidable [2702]. avoidance [3457]. aware [2575, 3087, 2733, 2818, 3156, 3021, 3152, 2342, 2826, 2742, 2832, 3048, 2601, 2824]. awareness [2674, 2992]. axiomatic [1071, 1851]. axiomatization [711]. Axiomatizations [3332, 3306]. Axiomatizing [818]. axioms [911].

B [3199, 2003, 1148, 1669, 858, 3646, 1103]. B-fairness [1103]. B-splines [3646]. B-Trees [1669, 2003, 1148]. B2B [3093]. back [580, 3476, 3487]. back-references [3487]. Backdoor [3851]. Backdoors [3267, 3835]. Backoff [1558, 3909]. backtrack [2887]. backtracking [1826]. Backwards [3882]. Backwards-reachability [3882]. Bags [1356, 1463]. Balance [1661, 2551]. Balanced [3801, 2628, 1339]. Balancing [1389, 1702, 2688, 1894, 2317, 2824]. Ball [3661]. balls [2197]. Banach [82]. Band [3449, 3720, 2298]. Band-restricted [3449]. bandits [2749, 2756]. Bandwidth [1632, 1607, 2561, 2213, 3373]. Bankrupting [3769]. Barrier [1270, 1426]. Bartky [373]. Base [1397, 809, 1326, 3650, 717, 2896, 3077]. Based [1604, 1748, 2404, 1598, 1316, 1355, 1619, 2656, 3710, 2828, 2657, 3467, 2343, 2811, 3551, 2426, 2548, 3304, 2518, 2653, 868, 2841, 3185, 3644, 3175, 3178, 3865, 2827, 2729, 2209, 2771, 2866, 1025, 3407, 2594, 2574, 3231, 3139, 2228, 2715, 2658, 3554, 2609, 2145, 2735, 2285, 3736, 864, 3118, 2854, 2103, 3505, 2754, 3179, 2389, 2993, 2239, 3759, 2808, 2941, 2826, 3115, 3403, 2864, 3161, 2935, 3232, 2825, 3049, 3848, 2312, 3168, 2896, 3017, 2203, 2427, 2199, 2868, 2770, 3937, 3137, 2740, 3143, 2961, 2603, 2906, 2416, 2932, 3010, 2802, 2345, 2746, 2834, 3653, 3007, 3065]. Bases [3160, 1267, 1353, 3091, 563, 1147, 2411, 1059]. Basic [1663, 876, 884, 1005, 3412, 3653]. basis [888, 3028]. Batch [1502, 353]. Bayesian [3455]. BBC [2989]. Be [1468, 1055, 3422, 2353, 842, 654, 338, 825]. Beach [2068]. beacons [713]. beats [3835]. Behavior [1253, 2771, 803, 864, 2961]. Behavior-based [2771]. behaviors [818, 3177]. Behaviour [3175, 2574, 3277]. behavioural [3086]. behaviours [747]. Belief [2922, 1668, 2369]. below [3130, 2585, 2442]. benchmarking [3087]. Benchmarks [2399]. Bendix [605]. Benefit [1671]. Benefits [1284]. Berman [659]. Bernoulli [1453]. Best [1309, 3381]. best-first [3381]. Better [2873, 2441, 1421, 942, 2996, 598, 1148, 2061]. Betti [1911, 1448]. betting [3108]. between [3870, 1111, 2201, 2093, 1677, 2024, 173, 816, 388, 2671, 2995, 2012, 1036, 1540, 108, 665, 3359, 1119, 3055, 1084, 3389, 231, 1610, 414, 91, 32, 1791, 2953]. Betweenness [2552, 2941]. Beyond [2997, 3559, 3294]. bi [1141, 3288, 3020]. bi-immunity [1141]. bi-reversible [3288]. Biased [1551]. Bichromatic [1348, 3326]. Biconnectivity [3350]. Bidding [3632]. bidirectional [3929]. Big [3319, 3199, 3146, 3235, 3232, 3260, 3328, 3320, 3322, 3326, 3010, 2967, 2966]. bigram [2923]. bilanguages [435]. bilinear [535, 480, 3187, 3023]. Bin [1612, 2789, 3451, 3552, 3764, 3410, 3250, 251]. binarization [2951]. Binary [1342, 1499, 821, 1792, 1339, 1881, 2479, 3094, 1022, 3099, 661, 3510, 1864, 2791, 1063, 3765, 3753, 1100, 3155, 440, 260, 551, 687, 1136, 3485]. bins [2789]. biochemical [9]. bioinformatics [2110, 2306, 2069]. biological [2790, 322]. biology [320, 200, 2302, 2564]. biometric [2735]. biomolecular [1811]. bipancyclicity [3189]. Bipartite [3773, 3184, 1949, 2705]. bipartization [2228, 3532]. Bisection [3627, 3834, 2390, 3854, 3521]. Bisimilarity [1615, 3547]. Bisimulation [2947, 1608, 3531]. Bisimulations [881]. bispecial [2946]. Bistable [1726, 1602]. Bit [1261, 2998]. Bits [1160, 2122, 3350, 2143]. Bivariate [3646]. black [2159]. Blackout [3827]. Blackout-tolerant [3827]. Blank [1859]. Block [3601, 1656, 3790, 2672, 2485]. blocked [3369]. Blocking [1867, 1429, 464, 2528]. Blum [3379, 1416, 3378]. board [1816, 1831, 8, 29, 54, 86, 123, 161, 201, 241, 266, 283, 308, 334, 387, 408, 1168, 1194, 1857, 1872, 1890, 1905, 1847, 1924, 1933, 1954, 1971, 1985, 1996, 2008, 2016, 2028, 2040, 2054, 2066, 2076, 2088, 2101, 2108, 2116, 2126, 2139, 2151, 2162, 2171, 2187, 2196, 2205, 2219, 2233, 2245, 2254, 2268, 2278, 2288, 2301, 2311, 2324, 2335, 2348, 2361, 2376, 2396, 2408, 2409, 2421, 2429, 2438, 2444, 2450, 2456, 2461, 2466, 2471, 2477, 2487, 2493, 2503, 2513, 2524, 2539, 2556, 2571, 2589, 2599, 2620, 2630]. Board [2649, 2676, 2699, 2718, 2738, 2765, 2785, 2800, 2813, 2821, 2830, 2848, 2862, 2879, 2892, 2914, 2928, 2943, 2958, 2970, 2983, 3003, 3013, 3037, 3047, 3058, 3070, 3079, 3090, 3111, 3134, 3141, 3149, 3164, 3174, 3181, 3201, 3213, 3218, 3229, 3244, 3264, 3275, 3292, 3300, 3318, 3345, 3354, 3361, 3371, 3372, 3377, 3388, 3408, 3414, 3426, 3430, 3436, 3442, 3450, 3458, 3464, 3486, 3498, 3507, 3518, 3529, 3538, 3543, 3549, 3557, 3568, 3579, 3595, 3600, 3614, 3623, 3638, 3652, 3659, 3739, 3667, 3681, 3696, 3704, 3714, 3723, 3731]. Board [3744, 3750, 3756, 3761, 3767, 3778, 3793, 3807, 3815, 3824, 3830, 3836, 3841, 3856, 3860, 3867, 3874, 3887, 3893, 3902, 3906, 3913, 3916, 3926, 3941]. bodies [2160]. bond [3892]. boolean [877, 462, 454, 3787, 3535, 3352, 3130, 1310, 3711, 1267, 2566, 2908, 3418, 2637, 2411, 1179, 2500, 3307, 3588, 3618, 3508, 3765, 1286, 3737, 1271, 1843, 1471, 3833, 87, 1596, 687, 1418, 3781]. Boosting [1455, 2120, 1547, 1746, 2371]. Bootstrap [2900]. bootstrapping [2935]. bordered [2440]. bornée [237]. Borsuk [3611]. bottleneck [1164]. Bottom [1507]. Bottom-Up [1507]. Bound [1680, 1501, 1505, 3019, 1576, 1563, 1284, 1715, 1439, 2061, 1067, 896, 2023, 1043, 3248, 2637, 923, 365, 450, 2038, 1820, 1162, 849, 2552, 1946, 2135, 2034, 215, 2474, 2997, 445]. boundaries [2046]. Boundary [1498, 1015, 107, 39, 212]. Bounded [1461, 3308, 2999, 962, 1254, 2888, 1265, 3497, 2143, 359, 3821, 1202, 1437, 3491, 1418, 31, 1594, 1524, 2559, 3104, 252, 3533, 3706, 2201, 2672, 115, 114, 140, 831, 925, 3184, 868, 2760, 946, 3669, 2637, 217, 237, 2705, 3818, 3729, 2468, 3700, 3627, 2173, 1171, 3479, 2583, 3588, 42, 157, 3751, 2283, 1134, 35, 838, 874, 865, 311, 421, 3737, 2948, 2373, 654, 582, 164, 655, 309, 567, 736, 2488, 79, 286, 1161, 2049, 234, 585, 3373, 3315, 2540, 2989]. bounded-context [585]. bounded-degree [3588]. bounded-state [3818]. Bounded-Width [1254, 2999, 962]. boundedness [2788, 826]. Bounding [3477, 1262, 3912, 14]. Bounds [1587, 1775, 1780, 1254, 1748, 1785, 1026, 1663, 1520, 3192, 1529, 1489, 1745, 1697, 1415, 1271, 1322, 1721, 1689, 1386, 1799, 1421, 1364, 1761, 1333, 2136, 3716, 3169, 1974, 940, 1981, 2159, 3223, 3533, 2382, 221, 1115, 354, 180, 1975, 2226, 3418, 212, 749, 1179, 2412, 1051, 2441, 2616, 694, 3000, 2433, 871, 313, 2905, 3921, 3663, 2585, 964, 1993, 974, 745, 656, 810, 2036, 3737, 3129, 187, 2113, 210, 2509, 1989, 628, 1044, 618, 3419, 164, 612, 2643, 2011, 3215, 141, 2434, 2501, 3491, 2356, 565, 681, 2094, 1052]. box [2159]. BP [1565]. BPP [1032]. Bracketed [1]. Braess [2194]. Branch [3019, 2061]. Branch-and-Bound [3019]. Branching [1254, 1732, 1589, 2903, 1992, 1608, 1713, 2919, 1746, 1364, 1470, 2645, 872, 896, 962, 2535, 2788, 776, 1209, 824]. Branching-time [2903, 2919]. branchings [3877, 3394]. Brandenburger [3072]. breadth [3193]. Breaking [1270, 1426, 1408]. bribery [3719]. bridging [2567]. brief [3379, 2165]. broadband [2212]. Broadcast [1635, 2932, 1067, 1111, 2770]. Broadcasting [2451, 3432]. Brownian [1747]. browsing [2338]. Bruijn [3475]. brute [2702]. brute-force [2702]. Büchi [892]. Bucklin [3060, 3061]. Budget [2989]. buffer [2216, 1069]. Building [3804, 3754, 2935, 3320, 3005, 3088]. buildings [3116]. Burks [2494]. Bursty [1631, 2627]. bus [2827]. Buses [1441]. Business [2689, 3020, 2657, 3080, 3282]. bynondeterministic [335]. bytecode [2074]. Byzantine [3511, 2439].

C [705, 413, 3610, 1372]. C-grammars [413]. c.e [2284]. C1 [1857]. C2 [1857, 1872, 1890, 1905, 1971, 1985, 1996, 2008, 2016, 2028, 2040, 2054, 2066, 2076, 2088, 2101, 2108, 2116, 2126, 2139, 2151, 2162, 2171]. C3 [1814, 1829, 1952, 1904, 1856, 1922, 1871, 1889, 1970]. C4 [1815, 1830, 1858, 1873, 1891, 1906, 1923, 1953, 1972, 1986, 1997, 2009, 2017, 2041, 2055, 2067, 2077, 2089, 2102, 2109, 2117, 2127, 2140, 2152, 2163, 2172]. C90 [1388]. Cônes [506]. CA [1603]. CAC [2212]. Cache [1706, 2802]. caching [1898]. calculation [223]. Calculi [3490, 845]. Calculus [1604, 1677, 1666, 1669, 2601, 1795, 626, 969, 536, 1109, 1150, 750]. calibration [2333]. Call [1942]. Calvin [705]. campaigns [3939]. Can [1430, 1468, 1055, 3575, 3422, 2167, 654]. cancellation [759]. Candidate [472]. Cannot [1398, 1169]. canonical [890, 771, 3028]. canonicalizing [2536]. canonization [2526]. capabilities [3448, 390]. capability [3453, 322]. capacities [2142, 3886]. capacity [3705, 857, 1869, 1069, 3690, 3025]. capping [3616]. capture [1112]. captured [3101]. Capturing [2854]. Carathéodory [3242]. carbon [3616]. cardinal [2467]. cardinality [3552, 3410]. cards [2416]. cares [2489]. Carl [2362, 2363]. Carlo [1249, 2750]. carpooling [3054]. Cartesian [3189, 3039, 3884]. Cascade [1755, 401]. cascading [1975]. Case [3802, 1280, 3055, 2331, 1620, 1338, 1092, 789, 2374, 1056, 2854, 3504, 2373, 2887, 3222, 2475, 2896, 897, 3288, 3343]. Case-factor [2331]. cases [3618, 3269]. Catalan [2759]. categorial [3690]. Categorical [3400, 3784]. categories [284, 3039]. category [998, 3219, 2975, 3344, 570]. Causal [2531, 3178, 298]. causality [2985, 2582]. Cayley [2321]. CCS [2229]. CD [3901, 2706]. CD-systems [2706]. CDPS [3325]. ceiling [2543]. celebration [3378]. Cell [1780, 2036, 466, 2356, 3046]. Cell-probe [2036]. cells [598]. Cellular [1609, 1486, 1574, 1900, 2780, 794, 3127, 868, 3662, 329, 1230, 3763, 822, 915, 2724, 3355, 317, 321, 975, 704, 1174, 2998, 2072, 486, 2393, 2872, 2410, 2469, 316, 2837, 1200, 470, 1196, 171, 2227, 1087, 419, 3042, 578, 224, 335]. centdian [3421]. center [3367, 3190, 3570]. centers [3616, 3151]. centre [3337]. centres [2966]. centric [3521, 3282, 2342]. Certain [414, 2291, 97, 2135, 210, 2156, 797, 4]. certificate [2351]. certification [3919]. Cervonenkis [1552]. CFG [2951]. CFS [3044]. Chaining [1656]. chains [3782, 2876, 3392, 3822, 3527]. Chaitin [3223]. change [3694, 3077]. change-making [3694]. changes [2368]. changing [3288]. channel [3630, 3757, 3573, 3296]. Channels [1733, 1671, 2313, 2403, 3432, 3805, 2825, 2809]. Chaos [864, 3316]. character [522, 2445]. characterisation [3470]. Characterisations [2721, 1077]. Characterising [3353]. Characteristic [402, 403]. Characteristics [1767, 794]. Characterization [1433, 1376, 3810, 110, 52, 2352, 2208, 3416, 3662, 1203, 2751, 138, 394, 204, 530, 2980, 668, 98, 366, 349, 229, 972, 709, 2243, 418]. Characterizations [1296, 416, 125, 1719, 220, 3790, 3154, 1885, 983, 3795, 1195]. characterize [1072]. characterized [211]. characterizes [2291]. Characterizing [1533, 1201, 3707, 1548, 1456, 21]. charges [868]. charts [131]. chase [3241]. Chasing [1679]. chatting [2835]. cheatable [1141]. Chebyshev [3]. Check [2149, 2505]. checkable [1990]. Checkers [1639]. Checking [677, 1313, 64, 1668, 3308, 2903, 550, 2174, 2876, 3339, 3001, 2905, 2177, 484, 2744, 3359, 2682, 2074, 2679, 149, 3596]. checkpointing [2516]. chess [701]. child [3850, 3572]. Chinese [3252]. Chip [2824, 2828, 2956, 2320, 2827]. choice [1867, 839, 2528]. choosing [3934]. Chordality [855]. Chromatic [1478, 1522, 1228, 2373, 3253]. chunking [2495]. churn [3769]. Cipher [1656].

Circuit [1518, 1717, 2435, 2433, 2113, 2058, 1097, 462, 597, 2314, 808]. Circuits [1625, 1684, 1260, 1626, 1294, 1563, 1471, 1548, 1624, 1418, 495, 831, 790, 2956, 3191, 2355, 3285, 1134, 1825, 2926, 2636, 3737, 43, 127, 2047, 373, 3491, 539, 3283, 3254]. circulant [56]. Circular [3590, 1506, 2723]. circumscription [1180]. circumventing [2386]. clarifications [287]. Class [1460, 662, 2398, 1048, 3139, 2063, 2051, 803, 864, 474, 2627, 368, 747, 903, 314, 492, 756, 585, 57]. Classes [1793, 1255, 494, 1555, 1628, 240, 1357, 1256, 1553, 1791, 3645, 3363, 110, 2327, 919, 1042, 3440, 273, 799, 3715, 479, 3524, 2024, 2507, 3859, 2020, 1173, 3259, 3267, 2907, 504, 1036, 285, 838, 874, 125, 1078, 3789, 3802, 2782, 2950, 3266, 3878, 3919, 2757, 3930, 1098, 132, 572, 2445, 192, 293, 347, 1075, 263, 984, 392, 391, 1185, 286, 3294]. Classical [3779, 2037, 1682, 2977, 3918]. Classification [120, 1340, 3626, 2897, 2123, 2794, 2808, 1108, 2130, 2836]. classified [3554]. classifier [3086]. Classifiers [1800, 2281, 3219, 3344]. Classifying [3482, 667, 2607]. clause [2049]. clause-variable [2049]. claw [3560]. claw-free [3560]. Clearing [3456]. Clifford [1912]. Climbing [1535]. Clique [1384, 3524, 1290, 3477, 3736, 3558, 1931]. Clique-width [3524, 3477]. cliques [3656]. clock [832, 3198]. Clocked [3657]. clocking [2956]. Clone [2940]. clones [2411, 2308, 3249]. close [38]. close-packed [38]. Closed [677, 1242, 1180, 609, 284, 3039, 2596, 895]. closeness [1141]. closest [2660, 396, 2722]. Closure [3645, 1638, 1474, 3246, 1257, 1789, 1373, 480, 1120, 3160, 3641, 1164, 3437, 1139, 3400]. Closures [37]. Cloud [2890, 3185, 2739, 3112, 3144, 2741, 2743, 3348, 3710, 3349, 2745, 2883, 3323, 3156, 3118, 3604, 3150, 2742, 2880, 2933, 3009, 2881, 3347, 2746, 2834, 2967, 3007, 3020, 2747, 2740, 3010, 2966]. cloud-assisted [3323]. Cloud-based [3185]. Cloud-integrated [3144]. clouds [2744, 3321, 3152, 3151, 2882]. Cluster [2827, 3000, 3901, 3855, 1982, 2664, 2804, 3899, 3563, 2397, 3017, 2890]. Cluster-based [2827]. Clustered [1792]. Clustering [2379, 1982, 2124, 2209, 3603, 2896, 2576, 3832, 2097, 3742, 3762, 3665, 3784, 2634, 3505, 2968]. clusters [2322, 3000, 3265]. clutter [2818]. clutter-aware [2818]. CNFs [2357]. Co [1361, 2338, 2411, 2320]. co-browsing [2338]. co-clones [2411]. co-design [2320]. Co-Finite [1361]. coal [2719]. coalgebra [3076]. coalgebras [3078]. coarse [1941]. coarse-grained [1941]. Code [1656, 356]. Codes [1428, 38, 2001, 3331, 1078, 2034, 2189]. Codeterministic [521]. codifications [413]. Coding [424, 2013, 3393]. coefficients [2637]. Cohen [769]. coherent [3247]. coimplications [2921]. coin [1981]. coincidence [999]. coins [1881]. Collaboration [3145, 2521]. Collaborative [2634, 2425, 3085, 2517, 3118, 2329, 3455, 2935, 2992]. Collages [1362]. collapses [996]. Collapsing [1142, 945]. colleague [591]. colleagues [3641]. Collection [1351, 3811]. collective [2407]. Collision [3281, 3457]. Collision-free [3281]. Colloquium [3610]. colony [2883, 2768, 3066]. color [3393]. color-coding [3393]. colorable [1920]. colored [2616]. Coloring [3674, 3687, 3901, 3869, 2587, 3435, 3356, 2887, 3642, 3484]. colorings [2394]. colors [3641]. colour [2735]. colourable [2279]. Colouring [3342, 3503, 3920, 3928]. Column [2955]. columns [641]. combinations [2281]. Combinatorial [1742, 56, 2749, 172, 1685, 3269, 1183, 3169, 2352, 2652, 3433, 2791, 276, 311, 421, 1157, 3337, 1076, 3678].

Combinatorics [1706]. combinators [3499]. combinatory [3690]. Combined [3165, 1800]. Combined-semantics [3165]. Combining [1544, 1705, 1806, 2871]. Comment [248, 3187]. Commerce [1651, 1837]. commit [2215]. common [556, 3850, 200, 3258, 1805, 3591, 1945]. Communicating [3593, 564]. Communication [1733, 1293, 1710, 2726, 1489, 2497, 1294, 3040, 1724, 1157, 1512, 739, 891, 1679, 2314, 1536, 2850, 2000, 2157, 2296, 2024, 2169, 1820, 527, 1036, 1210, 2870, 1098, 1119, 3206, 850, 1161]. Communication-efficient [2726]. Communication-optimal [3040]. communications [3720, 3117, 2407, 2831, 2809]. communities [2811]. community [2790, 2899, 3118]. commutatif [174]. commutatifs [506]. Commutation [1794]. Commutative [1578, 1492, 753, 904, 174, 405, 1046, 1161, 2684]. Commutatively [677]. Commutativity [1025]. Commutativity-based [1025]. commute [1007]. Compact [2383, 3919, 2380, 402, 403, 1886]. compaction [3931, 1987, 2130]. comparabilities [3866]. Comparable [1593]. comparative [3937]. Comparing [273, 2532, 2406, 1810]. Comparison [1406, 1642, 1306, 2401, 2900, 2459]. comparisons [3172, 3938]. compatibility [893]. compatible [2545, 2787]. competency [2122]. competing [3604]. Competitive [1684, 2425, 1298, 1246, 1188, 1753]. compiler [304]. compilers [68, 231]. complement [895]. complementary [2994, 2703]. complementation [2924]. Complementing [892]. complements [1170, 433]. Complete [1371, 1521, 3530, 3800, 1760, 1570, 1374, 3155, 440, 750, 169, 1438, 1717, 3525, 998, 2999, 3104, 990, 1940, 2751, 3661, 861, 451, 605, 995, 2529, 543, 659, 747, 1552, 1195, 3654, 305, 2130, 2722]. completed [971]. Completely [2797, 25]. Completeness [1655, 1341, 1400, 348, 1353, 951, 1868, 2185, 260, 604, 1443, 872, 669, 664, 3586, 2593, 1056, 660, 3222, 339, 90, 326]. completion [3209, 605, 3753]. Complex [1283, 1930, 1466, 3327, 969, 1064, 2908, 3497, 1019, 1983, 3508, 1150, 1231, 1149, 398, 2577, 2337]. Complex-valued [1930]. complexes [3708, 3847]. Complexities [1294, 3420, 1084, 91]. Complexity [1518, 3466, 1549, 2646, 1710, 1793, 2225, 1711, 3551, 1510, 1743, 1796, 1712, 3925, 1255, 3711, 3334, 1573, 2781, 3791, 1384, 1765, 3239, 1628, 1489, 346, 617, 1397, 1591, 3672, 1262, 1236, 504, 1627, 1578, 384, 1304, 1681, 1730, 1692, 1344, 1719, 1658, 1577, 3602, 1696, 1391, 572, 1689, 1668, 3774, 1724, 368, 1512, 3721, 797, 1372, 1567, 1460, 1725, 1799, 602, 1679, 678, 1452, 263, 999, 1619, 1264, 1378, 1338, 1613, 1360, 1466, 3484, 1594, 1448, 1412, 1017, 2351, 2973, 1974, 3687, 960, 3804, 2327, 2453, 2559, 2155, 126, 919, 3104, 1111, 2000]. complexity [3310, 3933, 2023, 2990, 669, 1092, 833, 3923, 3360, 3809, 3286, 3920, 273, 799, 3043, 3127, 2446, 354, 3204, 2566, 2296, 2696, 1180, 1226, 2908, 2357, 2653, 1939, 657, 2758, 1949, 2050, 2226, 2291, 3034, 580, 76, 216, 3697, 3221, 269, 2024, 2790, 2705, 1914, 3132, 3729, 3434, 489, 2412, 3441, 2911, 3598, 3928, 3762, 183, 3060, 3061, 3763, 3885, 2542, 3293, 3628, 3457, 3501, 2904, 3000, 3784, 1820, 641, 902, 3390, 2905, 2417, 3351, 3386, 3355, 535, 3899, 3508, 587, 2037, 3252, 964, 1036, 2176, 1029, 2945, 3719, 2534, 3845, 2018, 838, 523, 745, 2145, 125]. complexity [3639, 3365, 642, 3890, 1928, 3447, 3064, 2392, 3249, 3633, 2686, 1210, 918, 530, 934, 3863, 3753, 3621, 930, 564, 1098, 3206, 2058, 2270, 623, 983, 3353, 132, 1099, 1100, 3770, 1157, 1093, 569, 634, 1810, 2303, 1091, 648, 1944, 2887, 2584, 2859, 1097, 2354, 3873, 3898, 471, 1823, 3515, 765, 1139, 3290, 445, 3253, 739, 738, 933, 1075, 1191, 3795, 850, 1945, 953, 3489, 927, 1118, 767, 1854, 3227, 3541, 597, 3733, 3747, 219, 441, 218, 984, 392, 391, 1185, 686, 2243, 2874, 701, 286, 1161, 3449, 381, 319, 1987, 2130, 2096, 1199, 742, 342, 808]. complexity [3918, 3343, 3254, 3826, 3512, 3294]. Complexity-class-encoding [368]. Component [2249, 746]. Components [1422, 3852, 3399, 3238, 3315]. composite [1160, 2215]. Composition [2851, 1007, 2389, 2742]. composition-based [2389]. Compositions [1759, 5]. Compressed [1337, 1796, 3287, 3617, 2762, 3748]. compressible [3645]. Compression [1707, 1290, 2228, 2495, 3715, 3847, 2562, 3445, 2184, 973, 2434, 2501, 3491, 2459, 3010]. Compression-based [2228]. Compteur [673]. Comput [3817, 2359, 3859, 2926, 2185, 3199, 3109, 2469, 2501, 3187]. Computability [1323, 3550, 2510, 695]. Computable [1686, 563, 516, 2617, 580, 3662, 1091, 192]. Computation [324, 1260, 2845, 1523, 2633, 1591, 3647, 156, 1481, 1132, 3576, 361, 687, 419, 1418, 3849, 3425, 1172, 25, 783, 925, 1051, 822, 2907, 1142, 457, 665, 2673, 2072, 1044, 3301, 1155, 648, 577, 921, 612, 3215, 79, 768]. Computational [1743, 1573, 3448, 1592, 2392, 602, 953, 1675, 1613, 3449, 1987, 118, 89, 2559, 126, 128, 3127, 2226, 2181, 1005, 3157, 183, 3779, 642, 2302, 2564, 983, 344, 3544, 2318, 2047, 765, 2091, 2317, 381, 2130, 342]. Computations [676, 443, 1479, 1758, 1234, 1611, 1437, 126, 3223, 3362, 2201, 489, 35, 245, 691, 1095, 2948, 3151, 736]. compute [2641, 2403, 272, 3265]. compute-intensive [3265]. Computer [3319, 3933, 1348, 3112, 3379, 3135, 3378, 2206, 2621, 2801, 616, 94, 883, 73, 2992, 31, 762]. computers [1165, 846, 231]. Computing [1250, 3383, 3327, 3167, 1911, 3727, 2890, 2739, 2960, 2651, 3611, 1348, 3112, 1367, 1976, 1533, 3014, 628, 3794, 1126, 1645, 991, 3004, 34, 2806, 2614, 3877, 2798, 2275, 3852, 1179, 3889, 2553, 2883, 2069, 535, 3863, 3676, 1206, 1079, 390, 531, 2397, 2071, 2832, 2312, 2822, 2880, 3240, 2243, 2407, 3320, 539, 2567, 2831, 2933, 3142, 3009, 2740, 2881, 2600, 3369, 3347, 2746, 2966, 274, 2971, 345, 2029, 214]. concatenable [395]. concealment [2344]. Concept [1462, 118, 2327, 287, 2299, 397]. Concepts [1238, 1333, 1928, 1190]. conceptual [855, 3179]. Concerning [1794, 484, 210, 84, 133]. concise [3038]. CONCUR [3636]. Concurrency [1277, 527, 859, 1060, 2635, 764, 1151, 917, 3636]. Concurrent [888, 1363, 1669, 858, 2841, 976, 2978, 891, 1103, 699]. Condition [2602, 1965, 296]. Conditional [837, 428, 2480, 3389, 3412]. conditions [2468, 16, 697, 80, 343]. cone [65]. cones [668]. Conference [2971, 3465, 1378]. Confident [1744]. configuration [287, 3084, 3186, 224]. configurations [615, 3646, 422]. Conflict [1366, 951, 3161]. conflict-free [951]. conflicting [3005]. Conflicts [1540]. Confluence [837, 903]. confluent [2168]. congested [3558]. congestion [3609, 2824, 3143]. congestion-aware [2824]. Congress [85]. Congruence [1608, 265, 903]. congruences [710]. congruential [815, 797]. congruity [751]. Conjecture [1382, 1561, 1436, 2558, 2068, 1114, 3862, 3561, 3618, 2636, 659, 3675]. conjugacy [3128]. conjugate [325, 2921]. Conjunctive [1571, 2591, 2654, 1653, 3226, 3309, 3626, 3165, 2911, 2856, 733, 2860, 2597, 3485, 3302]. Conjunctive-Query [1653]. Connected [2541, 1422, 3519, 3929, 2415, 3852, 2132, 3495, 3758, 2616, 2904, 3562, 3399, 459, 3064, 910, 834, 3670, 1902, 3582]. connected-hereditary [2415]. Connection [1081, 1898, 2989, 3686, 3545].

Connection-trap-free [1081]. connections [855, 771]. Connectives [1542, 3307]. Connectivity [1311, 1427, 1247, 1392, 1778, 1682, 1841, 3691, 2954, 2191, 3443, 1053, 3863, 1158, 2378, 3389, 2156, 539, 3858, 1199, 894, 1133, 3312]. connectors [3467]. Conquer [1243]. Consecutive [2309, 385, 2496]. Consensus [2379, 3511, 2222, 3611, 1895, 2554, 3099, 2386, 2573, 1810, 3406]. consequences [985]. Conservative [3159, 1351, 3034]. conservativity [3535, 1899]. considerations [3138, 74]. Consistency [1694, 1581, 3001, 1758, 1692, 757, 3360, 1002]. consistent [3310, 3933]. Consistently [1349]. consolidation [3152]. conspiracy [622]. Constant [1431, 1546, 1788, 1897, 2558, 3494, 3493, 3352, 1913, 3416, 1094, 3712, 2385, 1825, 2926, 3606, 2439, 3892, 2031]. constant-depth [1825, 2926]. Constant-Factor [1546, 1788, 3416, 3712, 3892]. constant-round [2439]. constants [501, 535]. Constrained [1277, 1949, 325, 2181, 2414, 2084, 2700, 113, 1811, 3811, 2770]. Constraint [1584, 1769, 2550, 3588, 1580, 3197, 1276, 1653, 2488, 1841, 3360, 2792, 3159, 2381, 3094, 3380, 2175, 2659, 3365, 2279, 3602]. Constraint-Generating [1584]. Constraints [1568, 1649, 1648, 1313, 1286, 1357, 1607, 2657, 2343, 3552, 3474, 3194, 2463, 2520, 2260, 2359, 3410, 1844, 3001, 3784, 1231, 2666, 2846, 2401, 3753, 2474, 746, 2661, 548, 3333, 3100, 3932]. constructibility [224]. Constructing [1311, 3285, 2223]. Construction [1325, 3475, 1506, 1449, 1472, 1564, 2227, 1390, 352, 708, 988, 3573, 2901, 941, 419]. Constructions [1550, 635, 2114, 599, 3237, 2158, 1163]. constructors [1007, 663]. Constructs [1667, 885]. Consumer [2342, 3236].

Consumer-centric [2342]. consumption [3082, 3308, 3728]. Contagion [3940]. Contain [1471]. Containment [1580, 1653, 3366, 2591, 3625, 350, 733, 1836, 309]. Content [2495, 2735, 2579, 1840, 2043, 2734, 2730]. Content-based [2735]. Content-dependent [2495]. Contention [1393, 1558, 3757]. contentless [313]. Contents [2027, 2220, 2204, 2218, 2232, 2267, 2277, 2287, 2300, 2310, 2323, 2334, 2347, 2360, 2375, 2395, 2420, 2428, 2437, 2443, 2449, 2455, 2460, 2465, 2470, 2476, 2486, 2492, 2502, 2512, 2523, 2538, 2555, 2570, 2588, 2598, 2619, 2629, 2647, 2648, 2675, 2697, 2698, 2717, 2736, 2737, 2763, 2764, 2783, 2784, 2799, 2812, 2820, 2829, 2847, 2861, 2877, 2878, 2891, 2913, 2927, 2942, 2957, 2969, 2982, 3002, 3036, 3057, 3089, 3110, 3133, 3140, 3148, 3163, 3173, 3180, 3200, 3212, 3217, 3228, 3243, 3263, 3274, 3291, 3299, 3317]. Context [1476, 1756, 1617, 312, 674, 1401, 1477, 1235, 1492, 560, 1373, 288, 1799, 367, 2601, 3078, 2917, 904, 222, 3526, 2781, 416, 423, 3105, 867, 2670, 1070, 2733, 404, 1, 792, 67, 411, 138, 384, 350, 3481, 108, 2549, 604, 690, 51, 137, 327, 3048, 170, 21, 300, 319, 1199, 77, 585]. Context-aware [2601, 2733, 3048]. Context-Free [1756, 1235, 1492, 1799, 1476, 312, 1401, 1477, 560, 367, 3078, 2917, 904, 222, 423, 3105, 2670, 1070, 404, 1, 792, 411, 138, 384, 350, 3481, 108, 2549, 604, 690, 51, 327, 170, 21, 300, 1199]. context-freeness [2781]. context-independent [319]. Context-Sensitive [288, 867, 108, 77]. contexts [3179]. contextual [3485]. Contextuality [3517]. contingency [1913]. continuation [3943]. Continuous [1307, 1690, 1753, 336, 635, 3074, 7, 708, 25, 2402, 3114, 3368, 1107, 2603]. Continuous-Media [1753]. continuous-time [1107]. Continuously [184]. Contraction [1082, 683, 712, 3733]. contractions [3309, 3443, 3640, 3780]. Control [3060, 3061, 411, 2216, 1284, 1611, 2248, 1942, 2252, 3551, 3720, 82, 2772, 859, 1060, 3805, 113, 2103, 2635, 769, 764, 1018, 583, 2964, 3005, 1151, 2740, 3143, 3343, 57, 3007]. controllability [3241, 2597]. controlled [773, 2735, 584, 1011, 367]. Controlling [3785, 2627]. controls [81]. Convergence [1496, 62, 1614, 1690, 970, 3817, 3740, 2209, 3734, 1806, 377, 343, 55]. Converging [3241]. Conversion [231]. Convex [1457, 2388, 561, 22, 3823, 3895, 82, 989, 2419, 1213, 2160, 2307, 3646, 1217]. convexes [174]. convolution [3735, 2483]. convolutions [3627]. Cook [1035]. cooperating [3882]. Cooperation [1575]. cooperative [2242, 3565]. coordinate [2122]. coordinated [3277]. coordinates [3585]. coordination [3374, 3833]. cop [3846]. Coping [551]. Coproducts [209]. Copying [1369, 1484, 670]. Copyless [3437]. Core [1429, 2541, 2567, 2802]. cores [3754]. coresets [3832]. corner [3715]. corners [3726]. Correct [280, 608, 304]. correcting [38, 1166, 1200]. Correction [953]. Corrections [3933]. Correctness [714, 508, 1487, 253, 605, 61, 130]. corrector [40]. correlated [3024, 2627]. Correlation [2379, 2012, 1084, 2964, 3328, 2953]. correlations [2251]. correspondence [639]. correspondences [2995]. corresponding [2930]. corridor [3897]. Corrigendum [1049, 1605, 3817, 3859, 2926, 2185, 3199, 650, 3109, 2501, 610]. corrupted [3234]. Cost [793, 1703, 1355, 353, 2191, 3616, 3437, 2825, 600, 57]. cost-register [3437]. costs [3250]. count [1913]. countable [3790]. Counter [673, 2947, 3479, 249, 296, 517, 3649, 753, 534, 625, 302, 1120]. counterexample [3561]. Counterexamples [1688, 1331, 1068, 2365]. Counterfactuals [1397]. countering [2938]. countermeasures [3653]. counters [946, 166]. Counting [1587, 3444, 3572, 2295, 1383, 1435, 809, 2534, 1490, 3605, 1363, 3535, 430, 1042, 3711, 3773, 829, 2758, 3034, 2024, 2394, 3157, 2911, 1173, 2701, 737, 3588, 3211, 1063, 3064, 2787, 2859, 3881, 2860, 2662]. counts [3566, 2923]. cover [3514, 2645, 2045, 1221, 3843, 1949, 2761, 3192, 3922, 3758, 2925, 2142, 2665, 2353, 3027, 3892, 3504, 3780, 765]. Coverability [3480, 2149]. Covering [3741, 1499, 1003, 2458, 3764, 3905, 2788, 3456, 350, 2134, 1806, 412]. covering/packing [2134]. coverings [989]. Covers [1429, 3333]. CPU [3018, 3017]. CPU-GPU [3017]. Crane [2068]. crash [2447]. crash-recovery [2447]. Cray [1388]. creates [1088]. Creating [1578]. Creation [1341, 1407]. CREW [1179, 1422]. criteria [2806, 2867]. Critical [3230, 3454, 3231, 3427, 3721, 2060, 3431]. Criticisms [2624]. cross [3051, 3284, 2742]. cross-platform [2742]. crossing [2020, 2532, 1976]. CRT [1776]. Cryptanalysis [2377]. Cryptographic [1271, 2431, 2158, 2113, 3325, 3007]. cryptography [593, 1096]. CSL [3339]. CSP [3171, 2671, 3267]. CSPs [1910, 3535, 3034, 3416, 3297]. CTMCs [3339]. Cubature [65]. cube [3708]. Cubes [1499, 3155]. Cubic [1577, 3103, 2777, 545, 300]. curve [62, 3403]. Curves [1562, 1456, 1915]. Cut [3205, 1569, 2954, 3671, 3289, 2723, 1634, 1731, 1983, 2997]. cut-and-paste [2723]. cutback [821]. cuts [3706, 3560]. cutset [3879]. Cutting [113, 3726]. cutwidth [2452]. CVP [3693]. Cyber [3230, 3427, 3581, 2819, 3428]. cyber-physical [2819]. cyber-security [3581]. cybersecurity [3454, 2965]. CYBEX [3428]. Cycle [3727, 1577, 1657, 3376, 1135, 694, 834]. Cycles [1583, 3840, 3266, 3884, 1987].

Cyclic [3461, 262, 771, 1063, 3770, 1088]. Cyclic-routing [3461]. cycling [126]. cylindric [729]. Cylindrical [172].

D

[3525, 2827, 3402, 2825, 3240, 1071, 711, 3406]. D-scheme [711]. D0L [1927, 2135, 2793, 555]. D0L-sequences [1927]. DAGs [3771]. d’ambiguïté [237]. Data [3169, 3319, 1298, 1686, 2633, 809, 737, 1767, 1636, 1402, 1353, 1274, 1357, 1770, 1635, 3321, 2529, 1365, 1611, 2596, 3135, 1512, 967, 129, 1326, 1412, 863, 3332, 3384, 1882, 2344, 3234, 1915, 855, 2000, 3310, 3933, 2299, 2479, 3689, 716, 3474, 2852, 2760, 3475, 563, 3185, 2224, 2986, 2814, 2263, 959, 3574, 617, 3762, 2272, 1074, 145, 3236, 2276, 2003, 3144, 2818, 3012, 3331, 2481, 2744, 3616, 1060, 717, 729, 456, 559, 1928, 2365, 3199, 3146, 2769, 2274, 3366, 3282, 1229, 3151, 3325, 3324, 3102, 2768, 3235, 3759, 966, 2747, 2265, 1836, 2380, 2635, 1808, 568, 847, 973]. data [1018, 880, 3210, 2340, 885, 3232, 3062, 1184, 693, 3328, 3811, 3337, 3828, 3320, 2546, 2819, 410, 2893, 2770, 3278, 510, 2810, 3066, 3322, 3009, 744, 3326, 3010, 3347, 2834, 2967, 2966, 2519, 3016]. data-driven [2263, 3282]. data-independent [3574]. data-intensive [3066]. data-streaming [2519]. Database [1881, 1341, 2256, 1529, 1483, 1530, 1567, 1279, 1407, 732, 1027, 1061, 2590, 890, 1081, 731, 1110, 1170, 3138, 1065, 2859, 764, 3210, 2340, 2269, 2048, 854, 1125]. Database-friendly [1881]. Databases [1542, 1277, 1605, 1648, 1527, 1572, 1289, 1667, 1585, 1283, 1464, 1304, 1692, 1504, 1541, 1531, 1410, 1652, 3310, 3933, 1841, 817, 968, 2592, 3001, 2273, 2593, 2175, 2595, 3323, 2285, 1231, 2392, 2529, 2289, 574, 3279, 967, 1018, 2597, 2865, 3120, 853, 3136, 2603, 1214]. datacenter [3521]. DATALOG [1439, 1061, 1285, 1224, 2792, 1526, 1468, 1409, 1280, 1282]. DBMS [3056]. Dbsan [2896]. Dbsan-Gmeans [2896]. DDM [2404]. DDoS [3431]. De-anonymization [3012]. deadline [2250, 3320]. Deadlock [854, 2229, 3848]. deadlock-free [3848]. Deadlock-freedom [854]. decaying [2258]. December [3500, 3567, 3658, 3738, 3808, 3868, 3927]. Decentralized [3848, 2576, 1043, 3144, 2938, 2330, 3088]. Decidability [1605, 3304, 3165, 1572, 1580, 3335, 1978, 511, 465, 330, 660, 490, 3515, 579, 294, 328]. Decidable [3839, 3440, 3499, 1516, 3196, 3625, 865, 2546, 189]. Deciding [1615, 3305, 2457, 3531, 3261, 2923, 3584, 1552, 3547, 3339, 903]. Decision [186, 1310, 870, 776, 1455, 1794, 1375, 2727, 584, 1239, 1547, 2499, 1596, 2427, 1448, 3028, 3255, 3196, 3441, 1232, 514, 484, 587, 235, 240, 372, 631, 3633, 2121, 239, 1079, 2059, 2993, 2239, 2796, 440, 797, 2358, 649, 45, 2909, 2183]. decision-making [2239]. Decision-Theoretic [1455]. declarative [3866]. Declustered [1387, 3062]. decodability [2923]. Decodable [3331, 2034]. Decoding [1776, 2001]. decomposability [41]. Decomposable [482, 209, 2063, 570]. Decomposing [1022, 718]. Decomposition [3392, 1755, 2479, 890, 2672, 3539, 647, 71, 262, 2405, 366, 2871, 46, 17, 3137, 5]. decomposition-based [3137]. decomposition-parameters [3539]. Decompositions [1768, 3790, 3132, 1170, 574, 2197, 2264]. decrease [2031]. Dedalus [2985]. Dedicated [591]. Deductive [1289, 1667, 1410, 968, 967, 1214]. deduplication [3121]. Deep [3834]. defenders [3391]. Defending [2929]. defense [3008, 3581, 3431]. Definability [1383, 619, 3305, 2980, 449]. definable [460, 1173, 2179]. Defined [1642, 887, 480]. defining [2714]. Definition [1529, 3412, 478, 2933]. definitional [13].

Definitions [1321, 2388, 754, 2714].

degenerative [3511]. Degree

[1433, 1520, 397, 1920, 2155, 3876, 3800, 3184, 1171, 3209, 3588, 3751, 865, 2846, 2661, 1202, 3619, 2909, 3932]. Degree-languages [397]. Degrees [1311, 1677, 183, 1719, 1332, 3108, 105, 3526, 3273, 2284, 1142, 225, 945, 1033, 649]. Delay [2072, 2216, 43, 127, 2627, 2314, 3908, 2605, 2961]. Delay-insensitive [2072]. Delayed [1609, 2051]. Delaying [3853]. delegation [3059, 3280]. deletes [1148]. Deleting [3624, 971, 890, 3878, 3853]. Deletion [3629, 3789, 3802, 3363, 3423, 3495, 3707, 3396, 543, 3279, 3563, 847]. deletions [3443]. DELFIN [2229]. delivery [3592, 3278]. demand [2406]. Denials [2889]. denotational [1039]. Dense [1554, 3896, 2045]. Densest [1743, 3904]. Density [1458, 269, 2069, 293]. Denumerable [1722]. Dependable [2971, 2960]. Dependence [3306, 2624, 3307]. Dependencies [1584, 727, 2977, 816, 733, 3271, 3605, 3866, 726, 2722, 3106, 653]. Dependency [2274, 3412, 3224, 717, 1206, 2721]. Dependency-preserving [2274]. dependent [2495, 1016, 2796]. deployed [2623]. deployment [2804]. Depth [1296, 1320, 3835, 1258, 1724, 79, 2558, 3923, 831, 432, 3257, 117, 425, 2236, 2995, 1134, 1825, 2926, 2636, 462, 632, 3491, 922, 3373, 3254]. Depth-bounded [79]. depth-first [3923]. depths [1176]. deques [528]. Derandomization [1738]. Derivation [2452, 42, 341, 420, 21, 234]. Derivation-bounded [42]. derivations [476, 173, 200]. Derivatives [3483, 497]. deriving [2506]. Description [2653, 1401, 1477, 2654, 3890]. Descriptional [1799, 1360]. descriptions [508]. Descriptive [3598, 3628, 1696, 983, 1264, 2840]. Design [890, 1367, 1304, 1471, 852, 1125, 2632, 2424, 882, 1850, 3157, 2604, 3885, 2320, 2846, 2998, 2387, 2180, 3337, 1991]. designated [3235]. Designing [2085, 2082, 3393, 2194, 3668]. designs [2712]. desirable [1125]. despite [3769]. Destroying [3904]. destructing [2810]. Detecting [1583, 2953, 3881]. Detection [1743, 2939, 2896, 1244, 3348, 3084, 2940, 2899, 694, 2229, 3233, 2064, 3040, 3403, 2211, 248, 2964, 3277, 103, 3086, 2961]. detector [2447]. detectors [1965]. Determinacy [3268, 2978]. determinant [634]. determination [265]. determined [3220]. Determining [2210, 751]. Determinism [1784, 1111]. Deterministic [669, 1165, 1135, 3757, 1759, 3502, 3487, 3296, 30, 245, 2333, 2569, 1623, 721, 3585, 952, 302, 1524, 110, 3226, 3432, 3589, 207, 511, 815, 3745, 443, 362, 3305, 3889, 786, 476, 3479, 3067, 3628, 404, 2553, 1287, 402, 403, 466, 2563, 3340, 3393, 208, 513, 3673, 1210, 892, 490, 3527, 576, 1222, 2706, 483, 941, 91, 897, 363, 681, 452, 633]. determinization [3075]. detouring [2541]. Detours [3797]. develop [2992]. Developing [3144, 2825, 2882]. Development [2580, 3086, 3052]. developmental [200, 323, 319]. deviation [2964]. device [2338, 372]. devices [3117, 3280]. DF0L [1798]. DFA [2022]. DFS [3350]. DG [2896]. diagnosability [3389]. diagnosis [3237, 3937]. Diagonalization [601, 3072]. diagonally [3449]. diagram [253, 1961]. diagrams [2331, 1163, 3100, 1071]. Diameter [3336, 1514, 3915, 3753]. diameter-reducing [3915]. diameters [3869, 1982]. diametral [3928]. diamond [3342]. diamond-free [3342]. dichotomous [1192]. Dichotomy [1312, 2238, 2548, 3204, 3629, 3751, 3602, 3701, 2859]. dictionaries [3247]. dictionary [1208, 2003]. Dietze [3862]. Difference [1660, 107, 3504, 2049]. Differences [3338, 1767, 936]. different [3119, 3502, 2714, 3812]. differentiability [80]. differential [2495, 76, 3828]. Difficult [601]. difficulty [1866]. diffusion [3026]. Difunctionally [261]. digital [2607, 2732, 3405, 3597, 3406]. digitization [3402, 3404]. digraph [3800, 3289]. digraphs [3363, 3494, 2112, 3394, 1164, 2321]. DIL [555]. Dilworth [3814]. Dimension [1742, 1462, 2184, 1415, 1417, 1372, 1552, 2293, 2462, 1975, 3239, 3665, 2098, 2018, 3621, 2754, 761, 1823, 153]. Dimensional [1796, 1582, 1339, 2780, 2046, 27, 22, 2867, 2181, 1362, 1979, 1994, 1884, 915, 3355, 866, 974, 260, 2410, 2469, 339, 3515, 2668, 2837, 171, 1217]. Dimensions [1496, 3817, 1226, 3534, 3170]. Direct [401, 366]. Directed [2491, 1136, 70, 58, 2032, 147, 3456, 3797, 3712, 3910, 3390, 1102, 508, 2687, 2638, 2540, 2669]. direction [325]. Directions [3230, 3454]. Directory [1725]. Dirk [1605]. disassembly [2294]. disclosure [943, 2265]. Disconnected [3560]. Discovering [2902, 2607, 2484]. Discovery [2479, 3177, 1451, 3178, 2733, 3021]. Discrepancy [1348, 2568]. Discrete [3398, 284, 177, 1107, 2225, 3401, 3374, 3740, 2419, 1160, 463, 240, 803, 522, 3403, 1126, 1216, 318, 203, 762]. Discrete-time [284, 1107, 463, 803]. Discretized [1582]. Discriminantal [3735]. Disjoint [3688, 1514, 1607, 3333, 3840, 789, 2417, 1926, 66, 2695, 492, 2638, 2540]. disjunctions [2275]. Disk [1387, 3921, 1806, 2803, 3881]. disk-covering [1806]. disks [2868]. Dispatch [2248]. dispersed [2744, 2769]. dispersion [3911]. Displacement [2920]. displacements [55]. displaying [3912]. dissemination [2579, 3225, 2380]. dissimilarity [1964]. Distance [1723, 3685, 2305, 3924, 3818, 3423, 644, 3606, 3834, 2289, 2993, 2474, 2200, 3679, 3547, 2611]. distance-hereditary [3423]. distances [2464, 3883, 531, 3838]. Distinct [1677, 3394]. distinguish [3170]. distortion [3493]. Distributed [2632, 1556, 3092, 2249, 1298, 1581, 3555, 2940, 2631, 2633, 950, 1479, 1367, 2744, 1758, 1530, 1101, 1336, 1725, 1657, 2577, 3004, 2850, 2884, 2344, 3705, 18, 3183, 2726, 2404, 2402, 882, 3720, 2817, 2224, 2181, 3221, 888, 2132, 1051, 2223, 737, 3122, 3616, 3919, 730, 3676, 618, 2941, 2934, 2635, 2962, 2211, 2346, 854, 57, 774, 2966]. Distributed-Memory [1367]. Distribution [1431, 1473, 1271, 19, 1557, 1268, 1968, 1190, 821, 2318, 2823, 2199, 3619, 2932]. distribution-free [1190]. Distributional [3485]. Distributions [1458, 1644, 226, 3119]. distributively [638]. Divide [1243]. Divide-and-Conquer [1243]. divisibility [2022]. Division [600, 2062, 1825, 2926, 2270, 3046, 2809]. divisor [3258]. DL [2854]. DL-Lite [2854]. DNA [1521, 2082, 2912]. DNF [1300, 1968, 2119, 2092, 1473, 1977, 3506, 1268, 1734]. DNFs [3851]. DNLC [1017]. DNS [3431]. Do [1511, 1398, 1169]. documents [2634, 2073, 2722]. Does [3504, 2925, 2355, 514]. DOL [579]. DOL-DTOL [579]. Domain [1614, 1690, 998, 2575, 2733, 1007, 3024, 469]. Domains [1792, 1412, 1006, 267, 487, 431]. Dominant [3403, 3449]. Dominating [2129, 3940, 2132, 3899]. domination [3725, 3495]. Dominator [3901]. Domino [840]. Domino-tiling [840]. don’t [3641, 2489]. Dot [1296, 117, 432, 2236, 2995, 2998]. Dot-Depth [1296, 117, 432, 2236, 2995]. Double [442, 38, 3284]. double-cross [3284]. doubling [2462]. Down [1507, 1547, 3268, 882, 1211, 2457]. downward [3332]. dpda [625]. DPDAs [514]. Drawing [2683]. drawings [3510]. drift [1136]. Drifting [1747]. Driven [2580, 2263, 2517, 3282, 2318, 3290, 3052]. Driver [3277]. DSPACE [1565, 1433].

DTOL [579]. Dual

[3376, 3725, 2543, 3521, 2921]. dual-port [3521]. Duality [405, 463, 570]. dueling [2756]. Dumortier [1605]. duplication [378, 2988]. duplications [1804]. Duration [1684]. during [1706]. During’cannot [825]. Duris [780]. Dyn [1460]. Dyn-FO [1460]. Dynamic [1208, 3829, 1325, 3677, 2404, 1527, 1389, 3139, 3053, 1789, 1460, 900, 3083, 3302, 3092, 3183, 2257, 669, 3593, 3691, 2804, 2451, 956, 2190, 3742, 2759, 3655, 1015, 735, 496, 3663, 685, 3676, 3664, 2747, 619, 1163, 891, 3232, 1894, 3006, 693, 3022, 2906]. Dynamical [794, 1687, 1486, 2225, 2845, 23, 267, 487]. dynamically [3050]. Dynamics [3008, 318, 3220, 2837].

e-business [3080]. e-mail [3114]. E. [898]. each [466]. ear [2875]. earliest [2250]. earth [3022]. easier [3198]. Easiness [1714]. easy [3618, 1824, 937]. eavesdropped [3147]. EC3 [2890]. eccentricities [3889]. Eccentricity [3553]. econdirectLMS.zip [1893]. ecosystems [2607]. ECTL [3194]. Ed [1872, 1890, 1905, 1971, 1985, 1996, 2008, 2016, 2028, 2040, 2054, 2066, 2076, 2088, 2101, 2108, 2116, 2126, 2139, 2151, 2162, 2171]. Ed. [1857, 1847, 1924, 1933, 1954]. Eden [287]. Edge [683, 3189, 1247, 1011, 894, 3582, 2370, 3641, 3669, 3519, 2190, 1015, 3521, 3846, 2228, 1926, 3443, 3640, 3780, 3701, 3563]. edge-connected [3519, 3582]. Edge-connectivity [894]. Edge-contraction [683]. edge-disjoint [1926]. Edge-fault-tolerant [3189]. Edge-label [1011]. edge-length [3669]. edge-periodic [3846]. Edge/vertex [3758]. Edges [1423, 3189, 3624, 2192, 3501, 3238, 3883, 3838]. edit [3924, 531]. Editing [3153, 3273, 3000, 2661, 3855, 2664]. Editor [724, 315, 2972, 436, 1999, 828, 688, 2247, 1241, 2504, 1973, 2078, 538, 980, 993, 491, 379, 785, 2494, 1525, 1432, 1566, 1621, 2849, 1350, 2478, 332, 2708, 1878, 1146, 2188, 2010, 1783, 2915, 878, 1425, 1405, 1674, 1159, 1509, 1219, 1741, 1834, 1630, 1701, 955, 1370, 1538, 1252, 1774, 590, 1050, 1154, 1030, 932, 734, 1266, 1259, 1297, 1330, 1414, 920, 1281, 1729, 913, 2650, 1113, 1647, 1709, 1616, 1909, 1273, 611, 1378, 652, 658, 1493, 965, 1419, 562, 806, 844, 851, 1073, 1001, 1021, 1058, 1186, 986, 1090]. Editorial [2572, 1816, 1831, 3276, 3636, 3004, 2590, 8, 29, 54, 86, 123, 161, 201, 241, 266, 283, 308, 334, 387, 408, 1168, 1194, 2187, 2196, 2205, 2219, 2233, 2245, 2254, 2268, 2278, 2288, 2301, 2311, 2324, 2335, 2348, 2361, 2376, 2396, 2408, 2409, 2421, 2429, 2438, 2444, 2450, 2456, 2461, 2466, 2471, 2477, 2487, 2493, 2503, 2513, 2524, 2539, 2556, 2571, 2589, 2599, 2620, 2630, 2649, 2676, 2699, 2718, 2738, 2765, 2785, 2800, 2813, 2821, 2830, 2848, 2862, 2879, 2892, 2914, 2928, 2943, 2958, 2970, 2983, 3003, 3013, 3037]. Editorial [3047, 3058, 3070, 3079, 3090, 3111, 3134, 3141, 3149, 3164, 3174, 3181, 3201, 3213, 3218, 3229, 3244, 3264, 3275, 3292, 3300, 3318, 3345, 3354, 3361, 3371, 3372, 3377, 3388, 3408, 3414, 3426, 3430, 3436, 3442, 3450, 3458, 3464, 3486, 3498, 3507, 3518, 3529, 3538, 3543, 3549, 3557, 3568, 3579, 3595, 3600, 3614, 3623, 3638, 3652, 3659, 3739, 3667, 3681, 3696, 3704, 3714, 3723, 3731, 3744, 3750, 3756, 3761, 3767, 3778, 3793, 3807, 3815, 3824, 3830, 3836, 3841, 3856, 3860, 3867, 3874, 3887, 3893, 3902, 3906, 3913, 3916, 3926]. Editorial [3941]. editors [1938, 1559, 1385, 2422, 1445, 1517, 1586, 2774, 1343, 1469, 1315, 1545, 1763, 939, 1450, 3071]. Edmonds [3107]. education [3082]. Edward [3378, 3379]. EFADS [3009]. Effect [1158, 501]. Effective [2907, 285, 1248, 159, 1323, 839, 118, 2485, 2086, 2864, 218, 3093]. Effects [672, 2628]. Efficiency [1497, 1279, 1232, 215, 973]. Efficient [1707, 1649, 2828, 1751, 1221, 3472, 1228, 1505, 1643, 2394, 1369, 2433, 1368, 1393, 3340, 1473, 2121, 1164, 1190, 1212, 1429, 1811, 2474, 2307, 193, 835, 2106, 1613, 1470, 1812, 3190, 3009, 3326, 2280, 626, 2726, 1968, 3117, 1960, 956, 1896, 3577, 3236, 557, 2547, 2229, 2084, 223, 3224, 2315, 1852, 3021, 944, 3152, 2768, 3225, 3279, 1220, 3150, 977, 2668, 3232, 2825, 3208, 3083, 2770, 2615, 3137, 3010, 2834]. Efficiently [3689, 1342, 2063, 2521, 1795, 3662, 2938]. eigenvalue [247, 47, 181]. eigenvalues [25]. eigenvector [25]. Eilenberg [2685]. elastic [3431, 2890]. Elasticity [3321]. Election [1735, 2875]. elections [2548, 3939, 3900, 2945, 3719]. Electronic [1651, 3147, 1837, 2810]. element [801]. Elementary [73, 833, 870, 3541]. Eleventh [162]. Elgot [705]. eligible [3845]. Elimination [1710, 1391, 3094, 3413]. elliptic [107, 212, 95]. Embedded [2249, 2248, 2320, 2823]. Embedding [1121, 1797, 3884, 957, 779, 3128, 2586, 3492]. Embeddings [1632, 3155]. emergency [2606]. Emerging [2766, 2811, 2965, 3150]. empirical [1897, 2625]. Emptiness [1308, 3625, 2546]. Empty [3136, 1148]. emulate [1054]. enabled [3116, 2340]. Enabling [2338, 2674]. enacted [2741]. eNCE [1308]. encoded [196]. Encoding [1380, 2485, 2445, 368]. encounters [3897]. encrypted [3347]. Encryption [740, 3348, 3710, 3573, 3798, 3341, 3329, 3009, 2932]. End [1490, 3476, 3569, 2936]. End-to-End [1490]. Endomorphisms [461]. ends [690]. Energy [2315, 2768, 3082, 3728, 3117, 3616, 3156, 3152, 3015, 2826, 2628, 3150, 3811]. energy- [3811]. energy-aware [3156, 2826]. Energy-efficient [2315, 3117]. enforcement [457, 3225, 3848]. Enforcing [2741]. engineering [2341]. engineers [2731]. enhanced [2962, 3347]. enhancement [2807]. enhancements [2605]. Enlarge [2149]. enough [1147]. entailment [3043]. entanglement [2037, 1863]. entity [3439, 3785]. entity-linking [3439]. entropies [232]. Entropy [1627, 1302, 2184, 3890]. entry [355, 258, 505]. enumerability [132]. enumerable [3045, 3031]. Enumerating [2693, 3689, 3572, 3878]. Enumeration [1710, 3463, 3472, 3671, 3224, 3908]. Environment [1581, 2993, 3015, 2742]. Environments [2572, 2525, 1755, 3664, 2747, 3006, 3320, 2770, 2740, 2183, 2525]. EPDF [2468]. epidemic [3624]. epsilon [3540, 112]. epsilon-nets [3540]. Epstein [468]. equal [2130]. Equality [1751, 592, 222, 571, 682, 3341, 1041]. Equals [1623]. equation [24, 180, 3272]. equational [2975, 911, 663, 2285, 1041]. Equations [1428, 3203, 1396, 998, 107, 712, 2672, 2533, 76, 2944, 176, 2924, 3673, 2064, 88, 1899, 102, 2499, 32, 131, 95]. equifair [861]. Equilibria [2646, 3634, 3609, 1914, 2419, 3351]. equilibrium [2446, 2333]. Equivalence [165, 1571, 1604, 1409, 857, 1507, 3131, 404, 1798, 1580, 3583, 483, 3596, 2850, 207, 511, 2947, 3043, 56, 2505, 3165, 189, 348, 465, 550, 537, 2457, 388, 249, 253, 176, 2413, 2135, 2793, 350, 160, 559, 3531, 72, 2707, 3109, 637, 625, 51, 327, 579, 879, 32, 814, 363, 711, 756, 633]. Equivalences [760, 93, 3359]. Equivalent [1734, 301, 2714]. era [3150]. erasable [292]. erasure [3331]. EREW [1392]. ERfair [1967]. ergodic [3042]. Erratum [142, 331, 2359, 406, 2469]. Error [2201, 1697, 1714, 1800, 1787, 248, 103, 38, 2281, 212, 2062, 3734, 1820, 3365, 2064, 1853, 141, 2964]. Error-bounded [2201]. error-detection [2064]. Errors [1335, 2586, 2464, 2326, 2489, 3649, 551]. essays [2866]. essential [3724]. Establishing [210, 3428]. Estimating [2930, 3728]. Estimation [1748, 1614, 1690, 1474, 1355, 1275, 2427, 2578, 27, 1839, 2207, 2941, 2864, 3025]. estimators [2986]. ET0L [476, 524, 447]. Ethereum [3775]. Etherless [3775]. ETOL [674, 519]. Euclidean [3832, 3402, 2289]. Euler [770]. Eulerian [3153]. EV [3147]. EV-Linker [3147]. evacuation [3921, 3934]. evaluated [3527]. Evaluating [3199, 3146]. Evaluation [1529, 437, 1466, 2869, 1717, 970, 2280, 3234, 2206, 2262, 2604, 2621, 2212, 2762, 1083, 193, 205, 2397, 2211, 2475, 2643, 2399, 2719, 2931]. evaluations [2467]. evasion [112]. even [2904, 3415]. even/odd [2904]. Event [2602, 2833, 3561, 2978, 1101]. Event-Condition-Action [2602]. events [120, 96, 97, 117, 3231, 356, 667, 3912]. eventually [3040]. Every [2659, 146, 66, 203]. Everything [1500, 2671]. everywhere [1093]. evolution [320, 2854, 3419]. evolutionary [2902, 2895, 3428]. evolvability [2751]. evolving [3448]. Exact [2252, 1742, 1618, 1346, 1179, 1340, 2092, 3879, 2093, 3420, 3409, 3311, 3611, 3768, 3227]. exactly [962]. example [852]. Examples [1333, 883]. Exception [3563]. exchange [2852, 2198, 3296, 3011, 2694, 692, 2596, 2265, 2581, 3059, 3428]. excluding [2020]. exclusion [655]. Excursion [1434]. execution [2517]. Executions [1657, 1487]. Exist [1436]. existence [985, 352, 1947, 1191, 487]. Existential [836, 3043, 3699]. Expand [2149]. Expanders [926]. expansion [3497]. expansions [3197]. Expected [1292, 1745, 1282, 2030, 3732, 2439]. experimental [2467, 3061, 2345]. experiments [2900]. Expert [1495]. explaining [3068]. Explicit [1542, 599]. Exploiting [2276, 3170, 3156]. Exploration [3663, 3281, 3635, 3776, 2385]. explorer [3650]. Exploring [3935, 2093, 2622]. explosion [2174]. Exponent [1424, 3167]. Exponential [1609, 749, 1563, 1730, 2034, 1456, 1438, 1111, 1756, 3423, 996, 745, 1824, 3878, 3027]. Exponentially [601, 3038, 2021]. Exponentiation [1491, 2688]. exponentiations [3346, 3653]. expressed [825]. expressibility [2068, 656]. Expressing [1076]. expression [3697, 523, 3544]. Expressions [1764, 1599, 1693, 1624, 1466, 460, 3305, 346, 3499, 3487, 1864, 2292, 3340, 2775, 72, 623, 3102, 154, 1083, 3483]. Expressive [872, 3439, 3220, 1528, 1667, 1280, 1443, 1439, 1278, 3440, 1138, 2651, 1065, 2048]. Expressively [1760]. Expressiveness [664, 2856, 660, 1531, 776, 3154, 3437]. Extended [773, 870, 2624, 2856, 3272, 2060]. extendible [2997]. Extending [589]. Extension [1352, 1351, 2083, 199, 3341, 206, 3686, 393]. Extensions [43, 127, 710, 1061, 3790, 2842, 3774, 3076, 510]. extent [269]. exterior [3735]. exterior-algebraic [3735]. External [1543]. externalities [2651]. extra [3519, 3389, 3858, 3312]. Extraconnectivities [2843]. Extracting [2057, 1550, 1787, 2060]. extraction [1198, 2734]. Extractors [1787, 2189, 2569, 2568]. extracts [2294]. Extrema [1289, 1043]. extrema-finding [1043]. Extremal [2350]. extremals [59]. eyes [2761].

F [554]. Facebook [2898]. faces [2148]. facets [738, 933]. Facilitating [3066]. facility [2541]. Factor [1546, 1590, 1788, 3416, 3712, 3606, 910, 3892, 2331, 2416]. Factoring [791, 812, 3162]. Factorization [88, 2299, 3130, 3128, 2106]. factors [3167, 2479, 144]. fading [2809]. failure [872, 1965, 2137, 3040, 2447]. failures [3511]. Fair [2581, 3059, 3832, 2604]. Fairness [1705, 1702, 1156, 1103]. faithful [922]. fallback [3060, 3061]. false [3234, 1197, 3233]. families [75, 230, 97, 423, 151, 647, 682, 303, 306, 37, 92, 3168]. family [908, 3256, 264, 502]. Fan [1418, 790]. Fan-In [1418, 790]. far [675]. farms [2806]. Fast [495, 1249, 2299, 1300, 254, 3820, 3889, 2132, 279, 2385, 251, 3258, 936, 1245, 2483, 1504, 3210, 977, 3575, 3592, 1806, 568, 744, 2483]. Faster [3311, 2701, 1446, 3532, 2445, 3748, 3803, 3171, 2223, 2304, 3415, 1035, 531]. fastest [2671]. Fat [1347]. Fat-Shattering [1347]. Fault [1659, 317, 2393, 2872, 316, 3492, 1244, 3189, 2956, 2211, 2314, 3320, 3582]. Fault-Tolerance [1659, 3582]. Fault-tolerant [3492, 2956, 2314]. Faults [1426, 1390, 2681]. faulty [3189]. FBDDs [1761]. FCFS [1127]. FCT [3647, 3576]. fd [972]. fd-acyclicity [972]. fDatalog [1461]. Feasibility [2833, 1313]. feasible [3453, 1139]. Feature [1940, 2729, 2901, 3485]. Feature-based [2729]. features [3626, 2735]. February [3509, 2246, 3578, 3666, 3743, 3816, 3875, 3942]. federated [2963]. Federation [85, 2743]. Feedback [1690, 2640, 2952, 2418, 3630, 3844, 2228, 947, 3069]. Feedforward [1108]. ferromagnetic [3535]. few [3741, 3923, 3068, 2489, 3910, 1210, 1095, 2544]. Fewer [3396, 905]. Fi [3147]. fiber [2146]. Fibonacci [2979]. fidelity [3375]. Field [1331]. fields [628, 791, 3891]. FIFO [3313]. Fifth [274]. file [1862, 1877, 1907, 1908, 1893, 1849, 1860, 1861, 1875, 1876, 1892, 1904, 1856, 1922, 1871, 1889, 1934, 1925, 2771]. file-sharing [2771]. Files [1337]. fill [2996]. fill-in [2996]. filling [62]. filter [3269]. filtered [3341]. filtered-equality-test [3341]. filtering [3085, 2329, 3455, 2941, 2484]. FIN [1270]. FIN-Type [1270]. Final [510, 294]. Find [1556, 1595, 1322, 3771]. Finding [556, 770, 1178, 3915, 1776, 3803, 425, 1247, 2374, 3562, 2951, 3415, 2681, 1577, 1786, 834, 752, 638, 3405, 375, 3870, 1043, 2637, 2535, 3478, 2790, 3768, 2616, 2701, 533, 2044, 3266, 875, 3772, 396, 2540, 3512]. Fine [3687, 3236]. Fine-grained [3687, 3236]. finger [1918]. fingerprinting [2936]. Fingerprints [3287]. finitary [3073]. Finite [3459, 1361, 669, 460, 1327, 1693, 1102, 2389, 1494, 1484, 149, 722, 1309, 679, 1244, 2225, 3660, 3589, 511, 3593, 354, 119, 3745, 3732, 3903, 3416, 3577, 1994, 3067, 3407, 355, 258, 3241, 2795, 2137, 849, 144, 587, 3195, 644, 146, 3126, 262, 2665, 66, 3802, 2392, 210, 2923, 3621, 2948, 3359, 628, 623, 791, 3095, 2707, 1927, 554, 801, 148, 2454, 483, 949, 903, 625, 502, 179, 415, 467, 2597, 20, 972, 286, 992, 21, 726, 505, 3596, 2725, 3918, 3826]. finite-precision [992]. Finite-state [3459, 722, 3593, 3621, 3359, 949, 179, 3596]. Finite-turn [149, 587, 625]. finite-valued [3416]. Finitely [1237, 1464, 196]. finiteness [3531, 37]. firefighting [2990]. Firing [680, 474, 577]. First [863, 1250, 2068, 1529, 3410, 2273, 1008, 16, 841, 3790, 1224, 3923, 3593, 3220, 1218, 536, 656, 3531, 3583, 2757, 43, 3381, 239, 500, 3193, 750, 3435]. First-Order [1250, 1529, 2068, 2273, 1008, 16, 841, 1224, 3593, 3220, 1218, 536, 3531, 3583, 2757, 239, 500, 750]. fission [3353]. Fit [3410, 3435]. Fitting [1915]. five [3172]. Fixed [3309, 1440, 2704, 3380, 518, 2281, 2984, 3711, 2496, 3423, 1117, 2143, 1943, 2228, 1946, 2789, 3584, 1177, 3242, 1945, 2473, 3269, 2049]. Fixed-Parameter [1440, 3380, 2496, 3423, 2143, 1943, 2228, 3242, 2473, 3269, 2049]. Fixed-point [2704]. fixing [2681]. Fixpoint [1352, 1062, 1152]. Fixpoints [1321, 754]. flair [3872]. Flake [3406]. flattening [2977]. flexible [3323, 3009]. Flipping [3123, 1981]. flooding [3029]. Flow [1533, 1245, 382, 3357, 425, 566, 253, 714, 559, 2405, 2290, 3115, 2964, 131, 281, 2515, 2836]. flow-decomposition [2405]. flow-time [3357]. flowchart [430, 721, 897]. flowchartable [220]. flowcharts [271, 709]. Flows [1533, 3321, 1475, 3462, 2168, 1006, 3242]. Floyd [3783]. FO [1460]. FOCS [2153]. Focus [1503]. folder [1856]. folding [1140]. forêts [378]. Forbidden [3897, 3925, 3871, 3094]. force [2702]. forced [3871]. forecasting [3083]. Foreground [1122]. forehead [3206]. Forest [1607, 3417, 3925, 2669]. forests [378, 3429]. Foreword [1525, 1559, 1432, 1385, 1566, 1621, 1350, 1445, 2349, 1878, 1517, 2557, 1938, 1783, 1425, 1405, 1586, 1674, 1159, 1509, 1343, 3647, 2247, 1741, 1469, 1315, 1834, 1630, 1701, 2164, 1241, 1545, 1370, 1538, 1252, 1763, 1774, 1266, 1450, 1803, 1259, 1297, 1330, 1414, 1281, 1729, 1647, 1709, 1616, 2748, 1909, 1273, 1378, 1493, 1419, 1819, 2893, 724, 315, 2849, 2972, 2478, 2708, 436, 1146, 2188, 1999, 2010, 2915, 828, 688, 878, 1219, 955, 2504, 939, 1973, 2078, 590, 1050, 1154, 1030, 932, 734, 538, 920, 980, 993, 491, 913, 2650, 1113, 379, 785, 611, 652, 658, 965, 562, 806, 844]. foreword [851, 1073, 1001, 1021, 2600, 1058, 1186, 986, 1090]. forgetful [321]. Form [1797, 719, 838, 863]. Formal [1752, 801, 3316, 322, 2657, 2299, 200, 2508, 2840, 202, 256, 2265, 248, 449, 103, 2522]. formalised [94]. format [3369]. formation [3053]. formed [2060]. Forms [1366, 1248, 1351, 3072, 222, 2122, 312, 535, 411, 447, 560, 604, 1899, 193, 194]. Formula [1604, 1713, 2566, 3697, 3781]. Formula-Based [1604]. formulae [2757]. Formulas [1300, 1267, 1491, 1688, 1421, 3339, 3285, 2092, 572, 3506, 883, 2568, 2049]. formulating [542]. formulation [2725]. Forthcoming [1629, 1640, 1646, 1654, 1664, 1673, 1683, 1691, 1699, 1708, 1718, 1728, 1739, 1749, 1762, 1773, 1781, 1790, 1801, 2186, 2195, 2217, 2230, 2244, 2253, 2266]. forward [2199]. Foundation [3319, 2426]. Foundations [549, 3829, 7, 3610, 2595]. four [2458, 1053, 2130, 957]. four- [1053]. Fourier [2483, 762, 2637, 148]. Fourth [214]. FPT [1941, 2491, 3922, 3789, 3878, 3027, 3330]. Fractal [1462]. fractional [1975]. fragment [3813]. fragmentation [2283]. fragments [2922]. Framework [1519, 1474, 1767, 1323, 3412, 3349, 2806, 741, 3925, 2817, 3185, 2607, 3144, 402, 403, 3116, 2625, 2212, 3359, 3179, 3018, 3049, 3330, 2581, 2882, 2407, 3320, 2966]. Frameworks [3668, 3439]. Freddy [1605]. Fredholm [88, 32]. Free [1325, 1756, 3629, 674, 1693, 1235, 1492, 795, 1799, 635, 3167, 2917, 904, 1476, 222, 352, 3920, 3593, 3727, 432, 3334, 1081, 117, 416, 312, 423, 3105, 3281, 3342, 3524, 2670, 2759, 3573, 1070, 1401, 1477, 404, 2374, 3503, 1867, 1864, 499, 1, 10, 3688, 792, 411, 138, 951, 384, 350, 3481, 3878, 1148, 2133, 227, 108, 1190, 127, 2808, 3560, 3701, 2549, 560, 604, 690, 1046, 710, 51, 841, 327, 367, 3848, 3208, 170, 21, 300, 1199, 2528, 3078]. free-at-empty [1148]. free-choice [2528]. freedom [854]. freeness [3526, 2781]. French [378, 174, 237, 673]. Frequency [1549, 41, 2826]. frequent [2485]. friend [591]. friendly [1881, 2977]. friends [2989]. From-below [3130]. frontier [2663]. Frontiers [2801, 2261, 3186]. FRP [2833]. FTP [1849, 1876, 1862, 1877, 1907, 1908, 1893, 1860, 1861, 1875, 1892, 1904, 1856, 1922, 1871, 1889, 1934, 1925]. Full [1666, 2275, 158, 2787]. Fully [2190, 1789, 1420, 845, 152, 1852]. Function [1261, 1578, 1553, 272, 3035, 106]. Functional [1406, 3268, 887, 727, 1045, 1231, 733, 3271, 3605, 3866, 2722, 3106]. Functionals [1437, 6]. Functions [1347, 494, 1459, 1383, 1262, 1578, 1757, 1271, 1463, 1564, 602, 1436, 1264, 1470, 1161, 1418, 592, 1454, 3466, 2023, 3124, 831, 887, 790, 580, 2637, 1179, 3598, 2147, 3256, 364, 1019, 2419, 3588, 516, 3386, 504, 144, 644, 1993, 2051, 642, 3783, 2113, 1044, 192, 462, 259, 2035, 2350, 839, 921, 26, 953, 87, 1118, 2921, 263, 396, 1185, 687, 649, 539, 57]. functions-type [259]. functor [518]. functors [3073, 793, 3076]. Fundamental [1583]. Fundamentals [3647, 3576]. Further [40, 3930]. furthest [991]. fusion [1162]. Future [3020, 3230, 3454]. Fuzzy [2533, 1544, 26, 2183, 2299, 2467, 2057, 2735, 1964, 2059, 2239, 2483, 74, 2921, 2484, 2909]. FVS [3683].

G [2966]. G-Hadoop [2966]. GA [3119]. Galactic [3786]. Galil [780]. GALS [2826, 2825]. GALS-based [2826, 2825]. Game [3293, 1495, 3470, 929, 3846, 1885, 3356, 3604, 999, 3428]. Games [1747, 1345, 3675, 813, 1596, 2677, 2646, 3517, 3632, 919, 44, 3634, 3609, 2678, 2841, 3314, 840, 2419, 3351, 2978, 2680, 3565, 2989, 2025, 3515, 767, 441, 3260, 2679, 3391, 928, 1980]. Gamma [3455]. Gap [1518, 1173, 1111, 2386, 2179]. Gap-definable [1173, 2179]. Gaps [1677, 159]. Garden [287]. Garden-of-Eden [287]. gate [3413, 2622, 3254]. gate-limited [2622]. Gates [1471, 1624, 3491]. gathering [3387]. Gaussian [1391, 3455]. Gaussian-Gamma [3455]. Gelder [1352]. gene [2902, 1808]. gene-order [1808]. General [2366, 2114, 96, 1415, 1333, 300, 1454, 1066, 2403, 3349, 2293, 662, 482, 870, 1085, 2147, 3250, 2398, 3195, 838, 2051, 2308, 181, 493, 59, 1854, 2130, 3343, 762]. generalisations [2840]. generalised [2581]. Generalization [1455, 1800, 2281, 2462, 2644, 3131, 185, 21, 3251]. generalizations [1114, 666]. Generalized [3496, 1399, 1267, 1839, 1312, 14, 2530, 3107, 100, 190, 3823, 2612, 3905, 40, 3306, 16, 15, 2405, 2980, 755, 5, 2796, 256, 232, 294]. Generalizing [3788]. generate [2280]. generated [3261, 3288]. Generating [1584, 2415, 848, 1070]. Generation [3276, 87, 1187, 3185, 417, 83, 2175, 966, 90, 2746]. generative [3690]. generator [77]. Generators [1676, 985, 1116, 196, 3031, 1919]. Generic [1329, 2673, 3222, 1856, 3385, 3573, 3763, 3359, 883]. Genetic [2869, 2828, 2901]. genome [1804, 1807, 1812]. genomic [2082]. Genomorphisms [1010]. genotyping [2085]. Gentzen [1851]. Genus [2291, 2705]. Geodesic [3099, 3401, 991]. Geographic [3328]. Geographical [3240]. geographically [2744]. geographically-dispersed [2744]. geolocated [3012]. Geometric [1441, 1670, 1570, 1243, 1172, 3029, 2583, 3919, 243, 2537, 992, 1002]. geometrical [3425, 1928]. Geometry [3130, 1675, 833, 1221, 44, 3398, 1005]. geospatial [3176]. GF [2944, 394, 1306]. gift [1217]. girth [3838]. given [3903, 3273, 903, 3619, 3838]. Gladkij [206]. Global [24, 1609, 1479, 2051, 3937, 2899, 2316, 607, 842]. Gmeans [2896]. Goldschmidt [2062]. Good [1520, 1187, 990, 3937, 43]. gossip [2811]. gossiping [2222]. gotos [877]. GPU [3018, 3017, 3369, 3814]. GPUs [2805]. Gradient [1748, 1598, 247]. Gradient-Based [1748, 1598]. gradients [377]. grading [2866]. grained [3687, 1941, 3236]. grammaires [237]. Grammar [3617, 1791, 312, 411, 341, 384, 2762, 560, 604]. Grammar-compressed [3617, 2762]. Grammars [1642, 1799, 198, 3226, 413, 3615, 222, 140, 2781, 1207, 1137, 3105, 237, 199, 867, 2670, 1992, 320, 173, 646, 773, 1015, 1070, 200, 3407, 3285, 428, 792, 67, 208, 3531, 3583, 584, 865, 1132, 291, 3690, 3544, 898, 1011, 2584, 2073, 412, 637, 51, 498, 150, 179, 3559, 137, 327, 367, 170, 21, 322, 190, 3485]. Grammatical [3390, 423, 647, 682]. grant [622]. Graph [1948, 1556, 1311, 3732, 1583, 1600, 1754, 320, 1401, 1477, 1507, 2587, 1290, 1737, 1990, 1752, 1695, 2103, 3532, 3135, 493, 3558, 1140, 1678, 948, 1308, 1791, 381, 1682, 1017, 2370, 3517, 3309, 3566, 3632, 3310, 3933, 3896, 741, 3401, 2452, 886, 385, 3668, 2291, 990, 2415, 3273, 3524, 1015, 1211, 3635, 2976, 2056, 425, 557, 3259, 3390, 2385, 2907, 3751, 3205, 666, 3138, 3789, 3802, 3266, 3707, 3878, 584, 3919, 1868, 2185, 3064, 1053, 3930, 692, 842, 1084, 3640, 3235, 3770, 1011, 3642, 3873, 1003, 2537, 2586, 581, 3210, 1988, 150, 2832, 1202, 367, 2874]. graph [281, 3136, 2839, 3858, 1199, 1133, 3315, 3838, 1052, 2540]. Graph-based [2103]. graph-controlled [367]. graph-realization [557]. Graph-Structured [3135]. Graph-theoretic [381]. graphic [3018]. graphical [2480]. Graphics [1348]. Graphs [1461, 1295, 3550, 1327, 1423, 1446, 1577, 1303, 198, 1920, 382, 3575, 2129, 2032, 855, 3840, 3522, 3469, 3901, 3533, 3896, 3708, 3706, 3920, 3823, 385, 3727, 3800, 3184, 3475, 3087, 3691, 1949, 2531, 3189, 3519, 990, 779, 3929, 147, 3725, 3897, 3153, 3342, 3477, 1085, 2705, 2020, 3497, 3700, 3239, 3854, 2759, 2642, 3553, 3889, 3928, 3627, 2169, 3587, 3521, 3846, 2192, 3665, 2616, 545, 2536, 3754, 3797, 3510, 961, 3503, 1218, 3688, 3899, 2221, 2283, 213, 2063, 3810, 456, 1132, 1102, 3364, 2695, 2401, 3366, 3911, 3940, 3389, 3397, 654, 1927, 3772, 2887, 3674, 3560]. graphs [2661, 3686, 3794, 3898, 834, 577, 3881, 3536, 752, 3210, 3333, 3492, 3489, 1902, 129, 1917, 1978, 3400, 410, 2130, 3137, 2687, 2638, 957, 2692, 3582, 3512]. greatest [3258]. Greedy [3698, 1386, 2961]. green [3616]. grey [3083]. Grid [3081, 2525, 2404, 1362, 3779, 3234, 3935, 2181, 2397, 2804, 3733]. Grid-Based [2404]. Grids [2575, 1121, 2318, 2316, 2317]. ground [2869]. Group [908, 204, 522, 1603, 1196, 1359, 3546, 3921, 2297, 1880, 2573, 2796, 2909]. grouping [3081]. Groups [901, 1249, 1711, 1560, 690, 3577, 1048, 3126, 3447, 910, 935, 211, 3288]. growing [867, 3873]. Growth [1756, 578]. guarantee [3435, 3427, 3320]. guaranteed [2442]. guaranteeing [338]. guarantees [2097]. guards [1885]. Gucht [1605]. Guess [1865]. Guess-and-verify [1865]. Guest [2590, 1999, 2247, 1973, 2078, 1938, 3280, 1525, 1559, 2422, 1445, 1878, 2188, 2010, 1783, 1586, 2774, 1674, 1509, 1741, 1834, 1630, 1763, 1774, 1729, 1647, 1709, 1909, 1493, 986]. GUHA [2481]. GYO [771]. Gyssens [1605].

H [3187, 2394, 2759, 2279]. H-colorings [2394]. H-colourable [2279]. H-minor-free [2759]. H.264 [2931]. H.264/SVC [2931]. HA01053LMS.zip [1862]. HA03022LMS.sea [1860, 1875]. HA05062ABCNOP.pdf.sea [1849, 1861, 1876]. HA06043LMSU.zip [1892, 1871, 1889]. HA11093LMS.zip [1934]. HA11093M [1934]. Hadoop [2966]. Hairy [3661]. Half [1499, 2944, 1148]. Halfspaces [1431, 2565, 2004, 2328, 2431, 3828]. halting [369, 3338, 268]. halving [3611]. Hamiltonian [3376, 495, 1577, 1112]. Hamming [38, 2474]. Hamming-error-correcting [38]. HAMS [2214]. Handle [1137, 898, 1110]. Handle-rewriting [1137]. handlers [2833]. Handling [1274, 1795]. Hankel [549]. Hard [1554, 1381, 1595, 1590, 1561, 3904, 3395, 3123, 2949, 2955, 3247, 1085, 3768, 3067, 2142, 3122, 2951, 2353, 2194, 1127, 2358, 2251]. harder [730]. Hardness [1428, 2561, 2432, 2563, 3128, 1714, 2154, 2391, 2387, 1204, 2014, 3494, 3184, 1926, 2565, 2431, 2787, 2388, 3675]. hardnesses [1919]. hardware [1850]. hardware/software [1850]. Hartmanis [1561, 659]. Hash [424, 494, 592, 3710, 3256]. hash-based [3710]. Hash-Coding [424]. Hashing [2536, 573, 442, 1183, 2136]. Hausdorff [1226]. Having [1688, 3869]. Hazard [373]. head [3574]. Heads [1398, 1169]. healing [2211]. health [3349, 2607]. healthcare [2733]. heap [902]. heaps [2710]. heaviest [1811]. heavy [3771]. hedge [3481]. height [3352, 96, 97, 661, 769]. Hellinger [3206]. Helly [3889]. Help [1251, 1176, 2122, 2167, 3771, 2925, 914]. helps [2875]. hereditary [2415, 3423, 543]. Heterogeneity [3152]. Heterogeneous [2572, 3016, 2208, 3019, 3580, 3267, 2621, 3053, 2214, 2319, 2767, 2318, 3017, 2407]. Heuristic [2564, 2405, 3324]. Heuristics [1737, 2575, 2735, 3152, 3381, 1140]. HEX [3198]. Hidden [1004, 2752, 2370, 3170, 1810, 1220]. Hide [3446]. hides [1160]. hiding [982, 1894]. Hierarchical [2897, 1754, 3643, 2097, 3214, 2794, 1084, 2682, 3455, 1988, 2166, 2407, 2516]. Hierarchies [781, 310, 988, 269, 2995, 206, 576, 520, 736, 3101]. Hierarchy [1535, 1501, 1296, 1592, 1675, 1678, 1364, 3685, 919, 432, 958, 216, 3574, 2976, 3139, 513, 996, 108, 243, 763, 698, 948, 633]. High [3014, 1514, 2322, 3534, 1979, 2177, 3021, 3265, 2154, 1093, 2622, 2393, 698, 1184]. High-Diameter [1514]. high-dimensional [1979]. high-level [2177]. high-performance [2622]. Higher [1443, 3082, 1975, 3293, 1899, 1163]. Higher-Order [1443, 3293, 1899]. Highly [1796, 2322]. highway [3665]. Hilbert [6, 247, 62, 916, 3328]. Hilbert-Huang [3328]. historical [2746]. histories [888, 1016, 1959]. history [3379, 1807]. Hitting [2527, 3533, 3814, 3796, 3689, 3495]. HMMs [2306]. HNN [3774]. HNN-extensions [3774]. Hoare [664, 1007]. hoc [2209, 3021, 2875, 2870, 2537, 3143]. HOL4 [2917]. Holds [1436]. holes [989]. Holistic [2250]. Holographic [2560]. holonomic [2918]. HOM [3891]. HOM-problem [3891]. homeomorphism [540]. homogeneous [3068]. Homomorphic [11, 3348, 3798, 2933]. Homomorphism [2507, 3859, 465, 550, 3289, 3536, 170]. Homomorphisms [665, 2693, 881, 253, 3751, 356]. Homotopic [3402]. Honest [994, 293]. honeycombs [3198]. hop [1111, 3571, 2209, 2941, 3025]. Hopfield [1614, 1690]. horizon [3732]. horizontally [3323]. horn [883, 1660]. hosting [2882]. hosts [2930]. HPCC [2399]. Huang [3328]. hub [3367]. hubs [3747]. Huge [3242]. hull [1217]. human [1807]. HV [3434]. HV-planarity [3434]. Hybrid [2733, 1516, 1060, 2734, 2711, 2804, 3655, 2976, 2728, 2104, 3095, 2885, 2967]. hybridization [2544, 3207]. Hyper [1534, 781]. hyper-AFLs [781]. Hyper-Rectangles [1534]. hyperbolic [3553, 3889]. hypercube [2843, 1165, 3333, 3492, 1894, 3312]. hypercube-like [2843, 3333, 3492, 3312]. hypercubes [1121, 2401, 3884]. hyperedge [3615]. Hypergrammars [393]. Hypergraph [1171, 1137, 1070]. hypergraphs [3469, 3444, 3689, 3801]. Hyperplane [3314]. Hypersimple [1361]. hyperspheres [3406]. hypertournaments [2526]. Hypertree [1768, 1885, 2264]. Hypotheses [1655, 2092]. hypothesis [1197, 383, 3781].

I. [468]. I/O [3382, 3224, 2803]. I/O-efficient [3224]. IBE [3187, 3023]. ICALP [3610]. idealized [600]. Identification [1600, 1344, 34, 3395, 2411, 2308, 1201, 2615]. identifier [2834]. identify [2674]. Identifying [3176, 1960, 3618]. identity [2658, 1124, 2933, 2906]. identity-based [2658]. IDS [2934]. IEEE [2971, 250]. if [2057, 2508]. IF-like [2508]. if-then [2057]. IFIP [85]. Ignorant [1349]. II [1144, 417, 134, 189, 434, 403, 3789, 127, 88, 1220, 594]. III [3533]. illegal [3115]. image [868, 3399, 2735, 2808, 3227]. images [11]. imaging [3398]. IMB [2399]. imbedding [107, 545]. immersion [2907]. Immunity [1438, 2282, 1141]. impact [3862, 3231, 3313, 1151]. impersonation [2936]. implement [2320]. Implementation [716, 3573, 2401, 885]. implementations [617, 280]. implemented [3814]. Implementing [3183, 2447]. Implicants [1734]. Implication [1231, 3195, 2721, 726, 2722, 3028]. Implications [1605, 1572, 3535, 3880]. Implicit [2645, 2003, 568, 1074, 2924, 847]. implies [2582]. impossibility [832, 2386]. Improved [2382, 1634, 1644, 2418, 2468, 2669, 810, 1053, 1689, 3759, 3679, 1387, 2540, 3678, 3169, 3669, 2142, 3921, 2658, 3789, 2703, 582, 3322, 3570, 2111, 3421]. Improvement [3336, 2769, 2841]. Improving [3084, 1946, 2803]. in- [3394]. in-band [3720]. in-degree [3800]. in-place [3668]. Inaccessible [923]. Inaccuracies [1403]. Inapproximability [1607, 3367, 2013, 3129]. inapproximable [3370, 2797]. Incentive [2545, 3118]. Incentive-based [3118]. incentives [3821]. including [2769]. Inclusion [727, 3271, 1345, 2775, 733, 2434, 2501, 814]. incomparable [264]. Incomplete [1304, 1770, 3762, 767]. incompletely [185]. incompleteness [1109]. Incompressibility [3701]. inconsistency [1943, 3161]. inconsistent [2273]. Increasing [1277, 1530, 2095]. Incremental [1529, 1365, 1502, 2275, 2834, 3065]. incrementing [3649]. Indecomposable [370, 761, 468, 152]. Indefinite [1412, 73]. Independence [911, 1462, 2977, 616, 1611, 3459, 2480, 3195, 583]. Independent [2249, 1637, 1367, 731, 1484, 3870, 3309, 2383, 3555, 2338, 3574, 399, 3256, 3771, 303, 2289, 967, 3879, 1916, 319]. Index [53, 1272, 1314, 1358, 1404, 1444, 1482, 1508, 1537, 1579, 1606, 1641, 1665, 1700, 1740, 1782, 1828, 1888, 1951, 2007, 2053, 2100, 2231, 2613, 28, 78, 116, 155, 195, 233, 257, 282, 307, 333, 360, 386, 407, 429, 455, 481, 507, 529, 552, 575, 603, 627, 651, 671, 696, 723, 748, 775, 800, 820, 843, 869, 889, 907, 931, 954, 979, 1000, 1020, 1038, 1057, 1080, 1106, 1128, 1145, 1167, 1193, 1233, 3554, 839, 502, 2834]. Indexed [1635]. indexes [3617]. Indexing [1842, 1357, 1650, 3710, 3475, 1884, 2668, 3137]. indices [2551, 2816]. Indistinguishability [1592]. Individual [2296]. individualized [2773]. individuals [3147]. indivisible [3026]. indoor [2864]. Induced [3905, 2415, 3503, 2639, 2695, 2993, 261, 2719, 335, 3688]. inducing [3400]. Induction [1520, 663]. Inductive [1750, 1099, 449, 2095, 326]. Inductively [887]. inefficient [1191, 338]. Inequalities [1627, 916, 169, 3678]. inequality [3367]. inequivalence [675]. Infeasibility [2562]. Inferability [1332]. Inference [435, 1750, 1778, 1771, 3850, 3390, 2285, 1212, 747, 1018, 1202, 2511]. Inferring [2840, 2367]. infinitary [3790]. Infinite [49, 1597, 2177, 1374, 1403, 1274, 2948, 27, 22, 712, 2792, 2991, 432, 3448, 3644, 1082, 3480, 2978, 866, 3109]. infinite-dimensional [27, 22]. Infinite-state [2177]. infinite-use [3644]. Infinitely [1237]. infix [1078]. Influence [1551, 2989]. Information [85, 1291, 1777, 1766, 1544, 1750, 1636, 1304, 2727, 1770, 1735, 3240, 2515, 982, 3685, 2000, 2114, 2124, 2773, 2899, 3768, 2863, 1162, 2731, 3039, 3303, 3296, 2057, 2315, 2801, 2889, 1989, 3301, 2265, 3115, 2043, 548, 612, 3387, 767, 441, 3093, 2581, 2674, 2611, 3428, 2096, 2968, 2336]. information-theoretic [2114, 612]. informational [2547]. Infrastructure [3230, 3236, 3454, 3088, 3152, 3427, 3431]. infrastructures [2963, 3231, 2741]. Inherent [1765]. inhibitor [2308, 2490]. Initial [883, 176]. injecting [2435]. injection [3234]. injective [3920]. Injectivity [486, 186, 3660]. ink [582]. ink-bounded [582]. inner [3418, 47]. inner-product [47]. innovative [2863, 2866]. Inoculation [2202]. Input [1711, 1016, 3448, 227, 1853, 3290, 3322, 203]. Input-dependent-only [1016]. input-driven [3290]. input-free [227]. inputs [3172, 313, 2755]. insecure [2182, 2377]. insensitive [2072]. insensitivity [720]. Insertion [3479, 3279, 847]. insertion/deletion [3279]. inserts [1148]. insightful [2122]. inspection [2341]. Instability [3909, 3662]. Instance [1341, 1023, 2141, 2562, 2854]. Instance-Completeness [1341]. Instances [1554, 1381]. instant [2609]. Instantaneous [1745]. instantiation [3573]. institutions [3082, 862]. instructions [20]. Integer [1519, 3108, 1423, 3162, 3017, 2031, 1717, 3091, 3068, 279, 2134, 3188, 2584, 3242, 941, 2136, 2061]. Integers [1776, 156, 3203, 178]. integral [176, 267, 487]. integrated [2252, 3144]. integration [2272, 1836, 73]. Integrity [1313, 1844, 3001, 3785, 548]. intelligence [3020]. Intelligent [2893, 2863, 2212, 3088]. intensive [3265, 3324, 3066]. Inter [2506, 2575, 2827]. Inter-deriving [2506]. inter-domain [2575]. inter-layer [2827]. Interaction [3800, 727, 3271, 548, 2770]. Interactions [816, 1336, 3055, 2898, 314]. Interactive [3072, 1265, 340, 2818, 1442, 1094, 1863, 3055, 2454, 3828, 2932]. Intercalation [380]. interconnection [2313, 1130, 2321]. interconnections [884]. Interdependent [586]. interface [2991]. Interfacing [2339]. intermediate [1065]. Internal [3620, 2896, 2535, 2786]. International [2525, 2971, 3465, 3610, 3647, 3576, 85]. Internet [2402, 1723, 2156, 2968]. interoperability [2580]. Interpolating [1491]. interpolation [109, 1108]. Interpretation [2511, 1023, 3855, 178]. interpretation-oriented [178]. interpretations [479, 255, 721, 560]. interpreter [626]. Interpreters [231]. Intersection [1242, 1431, 1368, 1666, 3684, 1975]. intersections [829, 2565, 2004, 2328, 2431]. Interval [1293, 2467, 1758, 2921, 2427, 2833, 3706, 385, 3929, 3845, 3396, 3584, 2483, 581, 2900, 2909]. Interval-valued [2467, 2921, 2900, 2909]. Intervals [1776, 1336]. intractability [981, 3682]. intractable [842]. Intrinsic [1344, 1719, 1928, 2754]. Intrinsically [2780]. Introduction [3014, 2422]. introspection [374]. intrusion [3084, 2772, 2929, 2930]. intuitionistic [1964, 2059, 2239, 2796, 2909]. invariance [3566, 132]. Invariant [107, 1205, 3482, 3024, 1100]. Invariants [1652]. inverse [3602, 2474, 452]. inverses [836]. inversion [2851]. Invertible [1603, 1486]. invertors [227]. investigations [1127]. investment [2316]. invocation [2742]. involving [1341]. IO [427, 434, 3101]. IoT [3116, 3179, 3811]. IoT-enabled [3116]. IP [2208, 3554, 3759, 1557, 2836]. Irreducibility [1248, 811]. irreflexive [1987]. Irrelevant [1237]. IRS [296]. island [3603, 2826]. isn’t [444]. Isolation [1587, 2357, 124, 1189]. isomorphic [3800]. Isomorphism [1382, 1371, 1518, 2526, 1600, 654, 1436, 1359, 2558, 1048, 2283, 666, 1096, 213, 1868, 2185, 2297, 493, 948]. Isomorphisms [924, 2122, 451]. isoperimetric [3521]. isotonic [3358]. Issue [3319, 3829, 2153, 2739, 2960, 3175, 3230, 3112, 2118, 2430, 2766, 3135, 1378, 2677, 3142, 3004, 274, 2572, 2590, 2422, 3677, 2514, 3454, 345, 2325, 3465, 2863, 188, 2363, 3014, 2801, 136, 2302, 2029, 3150, 214, 3636, 2091, 2269, 2831, 2600, 2336]. Issues [1629, 1640, 1646, 1654, 1664, 1673, 1683, 1691, 1699, 1708, 1718, 1728, 1739, 1749, 1762, 1773, 1781, 1790, 1801, 3276, 1531, 2186, 2195, 2217, 2230, 2244, 2253, 2266, 3054]. item [3093]. Iterated [1397, 1825, 2926, 742]. iteration [400, 708, 1143, 1144, 230, 715, 911, 1014, 215]. Iterative [2191, 352, 2901, 499, 975, 181, 259, 710]. IV [3925].

J [468, 3817, 2359, 3859, 2926, 2185, 3199, 3109, 2469, 2501, 3187]. Java [2074]. JCSS [2504, 2831]. job [2806, 2406]. jobs [353, 3357, 2574, 3265]. Johnson [1881, 1576]. Join [1772, 3932, 1326, 3224, 717, 3056]. Joins [1355, 905, 2270, 3684]. Journal [3933, 3135, 3319, 3112, 3379, 3378]. Jr. [860]. judgment [3551]. July [1378]. jumps [270]. June [3452, 1378, 3537, 3613, 3703, 3760, 3837]. juntas [2002]. justification [2973, 968]. justifications [2974].

K-rational [1124]. K4 [3027]. K4-minor [3027]. Kannan [1114]. Karp [1035]. keen [2875]. Keisler [3072]. Kernel [1045, 2664, 3818, 2786, 2659, 2994, 3606, 3161]. Kernel-LEAF [1045]. Kernelization [3259, 2527, 2614, 3629, 3209, 3158, 3266, 3834, 3814]. Kernelizations [3207]. Kernels [3363, 2640, 2112, 2472, 3843, 2491, 3245, 3671, 3188, 3881, 2839, 3315]. key [3032, 3644, 2198, 3785, 2658, 1096, 3341, 3011, 2694, 3088, 2199, 3329, 2031, 2615, 2906, 2932, 3653]. Keyed [1571]. Keys [573, 3195, 472]. keyword [3121]. keywords [3573]. kind [44, 43, 127]. kinds [295]. kleene [1203]. kleene-like [1203]. knapsack [3709, 2030, 3936, 365, 3702, 450]. knot [1944]. knots [2654]. Know [1511, 3641, 3516]. Knowledge [1397, 1522, 1610, 1055, 2159, 943, 1147, 964, 3585, 2655]. Knowledgebase [1411]. Knuth [605, 3783]. Kolmogorov [2559, 1712, 1434, 1226, 3729, 1627, 3621, 1099].

L [2842, 1524]. L-morphism [2842]. label [2897, 2542, 2482, 584, 1011]. Labeled [3847]. labelings [3863]. labelled [3489]. lags [102]. lambda [626, 1109, 845, 1666]. lambda-calculus [626]. Langages [673]. Langevin [2750]. Language [2249, 1291, 1476, 3465, 1362, 1344, 1719, 1396, 2064, 1620, 1465, 2851, 2917, 3352, 1850, 2689, 3763, 200, 1045, 341, 101, 3138, 866, 974, 2924, 1201, 836, 3261, 187, 2289, 3301, 3591, 2073, 2499, 885, 449, 171, 478, 1851, 2048, 280]. Languages [677, 1341, 1361, 1769, 1756, 532, 1320, 3903, 1327, 1401, 1477, 1759, 3610, 1597, 2236, 1283, 1570, 1328, 1798, 1276, 1235, 1695, 1492, 673, 1324, 1463, 2410, 2469, 1308, 1351, 1017, 1027, 110, 1055, 901, 962, 1104, 2068, 904, 3718, 2888, 3472, 68, 230, 547, 815, 140, 432, 3334, 887, 416, 417, 134, 362, 244, 465, 3246, 1171, 340, 1203, 2056, 2840, 404, 3390, 202, 2593, 2594, 2857, 1, 42, 264, 546, 157, 303, 3202, 64, 296, 138, 1068, 513, 235, 702, 1929, 1150, 350, 206, 1078, 2240, 2365, 2950, 3447, 108, 243, 473, 433]. languages [3031, 2529, 753, 192, 256, 475, 898, 690, 797, 502, 380, 150, 3038, 480, 3559, 759, 248, 37, 229, 314, 814, 103, 722, 2546, 620, 319, 234, 1199, 77, 1120, 397, 633, 167]. Laplace [3505]. Laplace-like [3505]. Laplacian [2426]. Laplacian-based [2426]. Large [1390, 2806, 2404, 1941, 3852, 2899, 3754, 2374, 41, 3210, 3, 3062, 2885, 3007]. large-scale [2404, 2885, 3007]. Largest [3591, 3405]. LATA [3465]. latencies [3609]. Latency [3705, 3432, 3712, 2419, 2213]. Latent [1650]. Lattice [3644, 1633, 3653, 2295, 2013, 2144, 1157]. Lattice-based [3644, 3653]. Lattices [1428, 3160, 1010]. Lawvere [3072, 2975]. Lax [3076]. Layer [1340, 3051, 3467, 2827, 746]. layered [3131, 3581, 2743]. layout [741, 2452, 846, 692]. Layouts [1387, 3062]. LC [697]. LC-regular [697]. LCP [2987]. LCS [758, 2700, 2309]. LDAC [2938]. Leader [2875, 1735, 3676]. Leaf [2491, 1045]. leak [2889]. leakage [3303, 3329]. learn [1225]. Learnability [1361, 1347, 1582, 2782, 1271, 1553, 1494, 1502, 1181, 2093, 2241]. Learnable [1555]. Learner [1334, 1620]. Learners [1686, 2952, 2095]. Learning [2370, 2435, 2128, 1534, 1744, 1742, 1291, 1496, 1748, 2369, 1458, 1503, 1618, 2326, 1300, 1237, 1431, 1498, 1267, 1346, 1505, 2119, 2760, 1459, 1617, 1270, 3534, 1348, 1342, 1349, 1455, 1670, 2118, 1457, 1068, 2430, 1269, 2665, 1473, 1532, 1719, 1672, 2240, 2365, 1547, 1238, 1977, 2004, 2328, 1345, 1989, 1365, 2755, 1177, 1689, 1268, 2035, 1792, 1216, 1452, 1331, 1456, 1613, 1333, 1453, 1755, 2327, 2366, 2422, 2425, 2424, 2436, 2293, 908, 3817, 1221, 2480, 1968, 2462, 2364, 1207, 2019, 2122, 2750, 1896, 2325, 2545, 2141, 2432, 2751, 2433, 2901, 2092, 2752, 1928, 2498, 2950, 2753, 1190]. learning [2144, 2565, 2431, 3506, 1966, 1184, 2091, 3828, 3391, 2005, 1136, 2094, 3485]. learning-rich [2019]. Least [518]. leaves [3923, 3771]. Left [234, 2617, 137]. left-computable [2617]. Left-derivation [234]. left-to-right [137]. Leman [3566]. Lemma [2357, 1676, 3344]. lemmas [475, 380]. Length [3706, 1423, 3126, 2382, 1115, 3669, 3697, 3132, 533, 3554, 1213, 2058, 1811, 569, 306, 3333, 2540]. length-based [3554]. Length-bounded [3706, 2540]. length-constrained [1811]. length-preserving [306]. length-weighted [2382]. lengths [226]. less [300]. Let [1337]. letter [2781, 303, 2135]. letters [2068]. level [1178, 2432, 3910, 2177, 2854]. level-1 [3910]. level-ancestors [1178]. Levels [1458, 1176, 2369, 1163]. leveraging [3811]. license [2807]. Lie [1337]. life [2317]. lifetime [2916]. Lifting [3826]. lightweight [3348, 2906]. like [3470, 2843, 1203, 2001, 2508, 3505, 3333, 3492, 3312]. Likelihood [1454]. Limit [1374, 3763, 3127]. limitation [3046]. Limitations [1591]. Limited [1772, 1543, 1372, 1965, 3764, 3041, 3630, 1863, 1069, 2622]. Limitedness [644]. Limiting [422]. Limits [1366, 1633, 3413]. Lindenmayer [521, 323]. Lindenstrauss [1881]. Line [1744, 1598, 1455, 1454, 3850, 1226, 3569, 36, 279, 1753, 3618, 569, 648, 612, 1136]. lineal [3923]. Linear [1649, 1784, 1515, 1428, 1694, 1612, 1687, 1688, 2, 1573, 1677, 3798, 1368, 1713, 2069, 3331, 2044, 1580, 704, 1234, 187, 2297, 1548, 1429, 1486, 1574, 488, 1318, 439, 3161, 1685, 415, 267, 3315, 1076, 118, 794, 2332, 369, 2685, 3708, 2618, 2234, 868, 3869, 779, 405, 3105, 2944, 3818, 3662, 60, 365, 450, 2132, 2192, 2001, 2786, 1820, 3510, 599, 789, 3513, 49, 2728, 111, 124, 1990, 71, 11, 204, 463, 66, 3673, 2994, 3606, 1213, 113, 2695, 3188, 390, 461, 509, 88, 102, 194, 2106, 467, 1854, 229, 3, 487, 1196, 169, 3908, 1087, 2048, 1136]. linear [2839, 3781, 823]. Linear-Consistency [1694]. linear-size [2132]. linear-sized [599]. Linear-time [3798, 2069, 3331, 789, 1213]. linear-weighted [2728]. linearization [3069]. Linearly [1761, 1412, 325, 394, 146, 203]. linguistic [2993, 2796]. link [3736]. linked [2710]. Linker [3147]. Linking [3475, 3216, 3439]. links [2213]. LINMAP [2183]. Lipski [860]. List [1388, 3182, 2394, 3289]. Listing [1218]. listings [382]. lists [121]. Lite [2854]. lived [1895, 3099, 2213]. Ljubljana [85]. LL [697, 637, 498, 686]. LL- [697, 637]. LL-regular [498]. LLP [630]. LMN [1727]. Load [1389, 1702, 1240, 3357, 2688, 2318, 1894, 2317, 2824]. loads [3026, 1894]. Lob [1662]. Lob-Pass [1662]. Local [2702, 2495, 2897, 2899, 2644, 3233, 235, 1198, 937, 2673, 3215, 182, 2074, 3828, 607, 3251]. Localising [2520]. Locality [1246, 2070, 3182, 2224, 3202]. localizability [2582]. localized [2938]. Locally [167, 3422, 3211, 2283, 2034]. locally-optimal [3211]. locating [1811, 947]. location [2541, 2864, 2166, 2197]. locations [3176]. Locking [1527, 1394, 1025, 730, 1013]. Locks [1277, 3032]. log [2283, 2636]. log-bounded [2283]. log-depth [2636]. log2n [2143]. logarithm [1160]. Logarithmic [1435, 1258, 1392, 3617, 1165, 1074]. LOGCFL [1696, 1072]. Logic [1250, 1722, 1642, 1024, 1279, 1278, 2973, 3790, 1224, 1026, 669, 664, 3593, 3043, 1138, 2922, 3307, 2654, 776, 3306, 735, 2432, 2976, 2977, 496, 1045, 660, 685, 1198, 2285, 1209, 1149, 755, 3301, 3076, 619, 2510, 2920, 891, 841, 184, 778, 1854, 667, 1152, 1041, 1214]. logic-logic [1198]. Logical [1296, 1383, 1477, 1153, 2714, 1885, 3202, 227, 853, 3028]. Logics [3474, 2653, 2158, 3468, 3304, 2508, 2919, 760, 830]. logspace [1116]. long [2624, 1895, 3099, 3503, 3266]. long-lived [1895, 3099]. long-range [2624]. longest [1805, 1945]. look [1227]. lookahead [589, 1028]. looking [580]. lookup [3759]. lookups [3554]. Loomis [3224]. Loop [675, 530, 226, 192]. loosely [1902]. Loss [1745, 1454, 2207]. Lotka [3740]. Low [1628, 1393, 1725, 3523, 3665, 3707, 1095, 2872, 3515, 2043, 2825, 2936, 698, 948]. Low-Contention [1393]. low-cost [2825]. low-dimensional [3515]. low-end [2936]. low-rank [3707]. Lower [1680, 1775, 2159, 1780, 1254, 1501, 1115, 1975, 923, 2412, 1489, 694, 2905, 1563, 974, 3737, 1271, 1322, 1721, 1715, 1799, 1421, 2356, 1364, 1761, 565, 681, 1052, 3716, 1974, 940, 1067, 1981, 896, 2023, 1043, 3248, 2226, 3418, 1179, 365, 450, 1051, 2616, 1820, 2433, 3921, 849, 2552, 1993, 656, 2036, 2034, 2113, 2509, 628, 2643, 2011, 3491, 2094, 3533, 964]. Lowness [1793, 984]. LR [191, 3226, 589, 1028, 199, 3510, 697, 637].

LR- [697]. LR-drawings [3510]. LR-regular [199, 637]. LTL [1760]. Luc [1605]. luck [593]. LWPP [2179]. Lyndon [3272].

M [3187]. MA [2201, 3402]. MA-digitization [3402]. Machine [1555, 1348, 1437, 1485, 2424, 1918, 3262, 438, 2883, 35, 146, 3845, 625, 46, 3752, 100, 203]. Machines [2713, 1366, 1322, 1716, 1244, 1066, 209, 252, 3425, 3593, 525, 596, 571, 2991, 2, 1941, 443, 217, 60, 786, 541, 773, 3479, 36, 246, 3513, 961, 49, 3812, 284, 849, 238, 517, 587, 3649, 11, 874, 974, 125, 310, 3783, 1095, 5, 3152, 2290, 582, 41, 185, 261, 164, 349, 294, 1865, 389, 3654, 31, 3596, 3746]. Machtey [591]. Macro [805, 1759, 773]. macrogrammars [393]. Magic [2272, 1352]. mail [3114]. maintain [492]. Maintaining [1789, 2674]. Maintenance [615, 3812, 2896]. Majority [3494, 3355]. Make [1381, 3504, 2989]. makes [3768, 101]. Making [3929, 959, 2727, 3694, 2059, 2993, 2239, 2796, 2909, 2687, 2183]. malicious [3427]. malleable [2666]. Malware [3086]. manage [2579]. Management [2525, 1298, 2631, 2937, 2606, 2771, 3139, 3321, 3151, 2340, 3006, 2953, 3083, 2933]. manager [3017]. Managing [1540, 3055]. MANET [2608]. MANETs [3050, 2934, 2932]. manifold [2426]. manipulating [3821]. manipulation [2945]. Many [1237, 225, 3692, 1786, 3136, 1795, 3923, 2021, 3771, 2707, 3109, 2148, 2802]. many-core [2802]. Many-one [225]. Many-visits [3692]. map [3187, 3023]. Mapping [3147, 2828, 2826, 1809, 294]. mappings [2853, 2596]. MapReduce [3383, 2747, 2967, 2519]. maps [468, 186, 370, 482, 3502, 3584, 486, 761, 414, 100, 335]. Marc [1605]. March [2255, 3520, 3594, 3680, 3749, 3825, 3888]. margin [3344, 2123, 2328]. marginals [3383]. Marked [147]. Markov [2656, 3782, 3255, 3196, 3441, 2876, 3392, 2752, 3633, 187, 3822, 1810, 3527, 2427]. Marriages [1602, 1117]. marshals [1885]. masquerading [3283]. mass [3427]. match [2036]. matchgates [3283]. Matching [1587, 1337, 1317, 1796, 2505, 1398, 2656, 1208, 2464, 804, 3590, 2489, 3221, 1085, 2705, 3147, 2304, 689, 1169, 3671, 3340, 2639, 936, 2445, 2808, 2011, 3227, 2961]. matching-based [2961]. Matchings [1389, 495, 3871, 3880, 2391, 3794]. materialized [2592, 1840, 2720]. mates [3641]. Mathematical [130, 343]. Matrices [1260, 1600, 56, 152, 641, 3765, 3045, 3284, 796, 1959, 801, 3449]. matricial [1143, 1144]. Matrix [1560, 1143, 1144, 3618, 1706, 1724, 2842, 549, 2479, 3130, 2553, 2057, 3753, 502, 588, 32, 762]. Matrix-vector [3618]. Matroid [1247, 1368, 2791]. matroids [3847]. matter [883, 3453]. matters [3655]. max [3384, 2618, 3289, 2638, 1910, 1731, 3171, 2671, 1983, 2997]. Max-3-Cut [1983]. MAX-CSPs [1910]. MAX-CUT [1731, 2997]. max-plus [2618]. Maximal [1555, 2384, 2123, 1396, 3309, 3167, 3555, 2415, 2777, 566, 2994, 2703, 3833, 2568, 2511, 607]. maximal-exponent [3167]. maximal-partition [2568]. maximality [2064]. maximally [3582]. maximization [3607, 4, 3811, 1931, 3678]. maximize [973]. Maximizing [2224, 1766, 1866]. Maximum [2984, 1311, 1384, 1576, 1312, 1348, 3606, 2279, 3620, 3417, 1475, 1454, 2146, 128, 2495, 3715, 3357, 3478, 3683, 2786, 2069, 3736, 629, 3794, 1006]. maximum-density [2069]. May [3528, 3599, 3695, 3755, 3831, 3894]. Maze [219]. Mazurkiewicz [1760]. MCFL [3717, 3101]. MCMG [3018]. Mealy [1485, 2991, 388]. Mealy-type [388]. mean [3314, 2627]. mean-payoff [3314]. Meanders [940]. meaning [2481]. meanings [2714]. Means [2805]. Measure [1360, 3231, 2058, 2707, 3109, 922, 2111]. measure-many [2707, 3109]. measurements [3234]. Measures [1215, 430, 346, 1964, 2993, 3024, 3795, 478]. Measuring [3655, 1767, 2131]. mechanisation [2917]. mechanism [2424, 886, 2635, 2935, 3232]. mechanisms [3632, 2652, 2319, 3453, 2211]. Media [1753]. Median [1556, 1788, 3708, 3248, 638, 375, 3190, 3096]. Medians [3708, 2614]. medical [3179]. medium [3805, 3210]. medium-sized [3210]. Meet [2168]. meeting [2346]. meets [2554, 3356]. Membership [1251, 867, 1473, 1593, 2326, 1968, 3929, 2432, 3511]. Membership-Query [1473]. Membrane [2243]. Membranes [1645, 3208]. Memorial [2363]. memories [272]. Memory [1784, 1614, 1367, 2643, 3063, 2262, 2760, 2385, 1158, 3015, 3387, 1054, 3378, 3254]. memoryless [3099]. Mengerian [3810]. merge [2168, 3798, 1148]. merge-at-half [1148]. merging [3862, 783]. Merlin [919, 1596]. Mesh [1441, 1390, 3119, 1187, 2604, 3113, 2321]. meshes [1869]. Message [1656, 2961, 3063, 3225, 3848]. message-based [3848]. message-passing [3063]. messages [3114]. Meta [3158, 2575]. Meta-kernelization [3158]. meta-scheduling [2575]. metadata [3156]. metaregular [75]. metatheorems [384]. Metering [3236]. Method [1305, 118, 3081, 2403, 39, 2479, 325, 247, 868, 2467, 1960, 514, 71, 2481, 80, 2870, 1806, 2993, 2808, 554, 377, 88, 1222, 544, 2742, 17, 4, 396, 700, 2909, 2183, 3369, 3069, 2061, 326]. Methodology [1304, 2992]. Methods [1807, 5, 1724, 3032, 2426, 716, 3735, 40, 176, 2912, 2059, 2239, 181, 3150, 3648, 3240]. Metric [1731, 998, 3832, 2193, 3741, 3248, 3239, 2123, 917]. Metrics [1480, 2237, 2173, 2038, 2626]. Metropolis [1511]. Michael [591]. microblogging [2609]. Middle [1261]. middleware [1879]. might [2353]. migration [3098, 1962, 3764]. mild [80]. mim [3823]. mim-width [3823]. Min [1637, 2636, 3256, 3289, 2638]. min-max [3289, 2638]. Min-rank [2636]. Min-Wise [1637]. Mind [3539, 2368]. minded [2050]. mine [2719]. Minimal [2022, 1476, 1396, 2049, 855, 3444, 1207, 3262, 446, 3478, 269, 3683, 533, 581, 487]. minimal-program [269]. Minimalism [968]. Minimality [1755]. minimax [48]. minimise [3587]. Minimization [1256, 2358, 325, 2663, 3043, 2566, 351, 2432, 2532, 2482, 2271, 3853]. Minimize [1335, 1982, 3357]. Minimizing [2146, 2292, 3616, 743, 2441, 3884]. Minimum [943, 272, 1316, 3864, 1607, 1118, 141, 1734, 3705, 3493, 3800, 1949, 3429, 3132, 2683, 3854, 2954, 1804, 2223, 2542, 2191, 3501, 1192, 1943, 947, 1133]. Minimum-complexity [1118]. minimum-cost [2191]. minimum-length [3132]. Mining [1539, 2814, 2729, 2818, 2485, 2481, 3323, 2898, 2895]. minmax [3570, 3096, 3421, 2540]. Minor [3629, 2759, 2925, 3027]. minors [3533, 2020]. mismatches [3590]. miss [399]. Misses [1706]. missing [3768]. Mistake [1505, 2434, 2501]. Mistake-Bound [1505]. mitigation [2939]. mitoticity [2282]. MIX [3101]. mixability [2423]. Mixed [1967, 3773, 2568]. mixed-sources [2568]. Mixing [2899]. mixture [2329, 3505, 2005]. Mobile [3142, 3117, 2940, 3456, 2209, 2818, 2938, 2609, 3822, 2864, 2730, 3278, 2933, 2603, 2336, 2931]. mobility [2578, 2209, 2625, 3021]. mobility-based [2209]. Möbius [3155]. MOD [3418, 1470, 2012, 1294]. modal [3304, 2530, 830]. modalities [778]. mode [3117, 2772]. Model [3308, 2580, 1780, 1399, 1659, 2876, 2177, 1416, 1668, 2682, 1442, 1619, 1453, 3052, 2313, 1023, 1172, 2903, 2804, 3008, 1898, 2174, 3346, 3339, 2207, 3131, 2905, 3934, 3341, 2308, 729, 2214, 3736, 2854, 1158, 2210, 3055, 1229, 3389, 3455, 2447, 259, 3585, 3581, 2071, 2734, 1088, 2074, 2356, 2679, 622, 2522, 2427, 2838, 3828, 2516, 2567, 2743, 2770, 3005, 3086, 3937, 503, 1131]. Model-Based [1619, 2854, 2427]. Model-checking [3308, 2177, 2682, 2903, 2905]. Model-driven [3052]. modelers [2711]. Modeling [3234, 2249, 3117, 2837, 3051, 3728, 2208, 2956, 2621, 2625, 3822, 2331, 2885, 3056]. modelling [2206, 2624, 2602]. Models [1521, 1505, 1599, 2752, 1357, 1124, 1437, 1439, 2248, 2828, 855, 669, 2480, 783, 2402, 417, 3497, 2625, 1109, 691, 665, 2329, 2529, 2059, 3095, 1810, 2622, 1018, 588, 1103, 3428, 2005, 1107]. modification [3701]. Modifications [503]. Modifying [1413]. mods [1042]. Modular [3865, 254, 1850, 3895, 3346, 1010, 3162, 227, 2688, 3653]. modularity [3607]. module [893]. Modules [1752, 884]. Modulo [1677, 1160, 1041]. molecular [2302, 2564]. Moments [1549]. Monadic [1638, 1642, 501, 404, 745, 952]. monitor [3427]. Monitoring [3883, 3838, 3612]. monochromatic [3326]. monoid [174]. Monoidal [3106, 284]. monoïde [174]. monoids [532, 10, 1046, 1161]. monopolar [3364]. Monotone [1821, 1258, 1563, 1421, 831, 3702, 2797]. Monotonic [2104, 1410, 343]. monotonicity [968, 558]. Monte [1249, 2750]. Monte-Carlo [2750]. Moore [388]. Moore-reducibility [388]. Moore-type [388]. Moran [3686]. morphism [2842]. morphisms [836]. most [2306, 1135, 2130]. motifs [2616, 3064]. Motion [1747]. motivated [2537]. Moving [1842, 3177]. MPI [2885]. MPI/OpenMP [2885]. MPSoC [2823]. MS [2086]. MS/MS [2086]. MSCs [2177]. MSNP [3142]. MSO [3468, 3268]. MSO1 [2905]. MST [1546]. Mu [1604]. Mu-Calculus [1604]. much [2167]. Muller [373, 2189]. Multi [3082, 3219, 2867, 1094, 3565, 1420, 2573, 764, 2815, 2679, 1380, 3081, 3098, 2710, 1111, 2806, 3117, 2897, 3571, 1883, 3574, 646, 2209, 3067, 3610, 3480, 2883, 3139, 3344, 3296, 1863, 3882, 1880, 2627, 2941, 2902, 1126, 3747, 2346, 3277, 2730, 2567, 2909, 2251, 3066, 2816, 2809, 3025, 2906]. Multi-agent [3082, 3610, 2346]. multi-attribute [2909]. Multi-category [3219, 3344]. multi-channel [3296]. multi-class [2627]. multi-core [2567]. multi-criteria [2806]. Multi-dimensional [2867]. Multi-group [2573]. multi-head [3574]. multi-hop [1111, 3571, 2209, 2941, 3025]. multi-label [2897]. multi-linked [2710]. multi-mode [3117]. multi-node [3747]. multi-objective [2883, 2902, 3066]. Multi-oracle [1094]. multi-parameter [3067]. multi-path [2809]. multi-processor [3098, 2251]. Multi-prover [1420, 1380, 1863]. multi-pushdown [3882]. multi-query [1883]. Multi-round [3565]. Multi-route [2815, 2816]. multi-scale [3480]. multi-scheduling [3081]. multi-server [2906]. multi-UAV [3277]. multi-user [2730]. Multi-valued [2679, 1126]. Multi-version [764, 3139]. multi-visit [646]. multi-way [1880]. Multiattribute [2059, 2183]. Multicast [1703, 2178]. Multicasting [1707]. multicasts [1869]. multichromosomal [1812]. Multicommodity [1245, 3242]. Multicore [2822, 2885]. multicounter [587, 874]. Multicriteria [2239]. Multicut [2761, 3204]. Multidimensional [3214, 3314, 3729, 2413, 3128, 141]. Multidisciplinary [2766]. multifurcating [3417]. multihead [974, 125, 197, 703, 286]. multikey [1074]. multilective [766]. Multilevel [3554, 912, 3721, 870].

Multilinear [2568]. multilink [827]. multimedia [2343, 2212, 2214, 3011]. multiobjective [2895]. multiparameter [826]. Multiparametric [47]. multipartite [3517]. Multiparty [1116, 2633, 1489]. Multiple [571, 1544, 258, 1705, 1558, 1238, 3016, 3432, 2181, 2141, 3573, 3702, 355, 2744, 2767, 3565, 2993, 2298, 2900, 3417, 3811, 505, 3022, 3912, 3207, 3814]. Multiple-entry [258, 355, 505]. multiple-instance [2141]. multiplexed [2085]. Multiplication [1711, 3618, 1706, 1306, 279, 1825, 2926, 796, 588, 2136, 762]. Multiplicative [678, 215]. multiplying [50]. multiprocess [778]. Multiprocessor [1659, 2468]. multiprocessors [1894, 2203]. multipushdown [252]. Multiset [2979, 3270]. Multistack [1575]. Multistage [3709]. Multiswapped [2886]. Multitape [33, 20, 438, 2413, 245]. Multiterminal [1533]. multivalued [2721, 1216]. Multivariate [1331, 2944, 3233, 3393, 156, 791, 3545, 177, 3330, 812, 811]. multiversion [859]. multiway [3289, 1634]. music [2732]. musketeers [3895]. must [338]. mutual [2795, 655, 2416].

n [2374, 3406]. n2 [2079]. n4 [2160]. names [2383, 2975]. Narrow [3298]. Nash [2446, 3609, 2419, 2333]. native [3775]. Natural [299, 1447, 2610, 1110, 3541, 430]. nature [364, 813]. navigational [3154]. NC [962, 1031, 1104, 1221, 1523, 1472, 1189, 1465]. NCE [1695]. Near [1926, 3911, 2298, 3818, 2192, 3510, 1213, 544, 87]. near-linear [3818, 3510]. Near-optimal [1926, 3911, 544, 87]. near-optimum-length [1213]. Nearest [1780, 1736, 2197, 2865, 3137]. Nearly [1675, 1207, 1165]. necessary [1965, 16]. need [3641]. Needed [1333, 272]. Needs [2727]. negation [2977, 3737, 1062, 1152]. Negative [1291, 2192, 2365]. Neighbor [1780, 3937]. Neighborhood [338]. Neighbors [1736, 2197, 991, 3137]. neighbour [2865]. Nemhauser [2644, 3251]. Nested [1397, 301, 1992, 1025, 816, 1110, 1229, 1805, 912]. Net [1597, 1867, 3541, 1103]. Nets [1340, 1601, 1129, 3540, 1079, 763, 1108, 700, 485, 620, 2490, 2528]. Network [3030, 2575, 2606, 680, 2025, 2312, 3240, 2824, 2656, 3084, 3850, 2258, 3281, 3885, 2207, 3131, 807, 2165, 3233, 2319, 2846, 2210, 3280, 1930, 3664, 2622, 1097, 2628, 2929, 778, 1916, 3337, 3278, 3912]. Network-aware [2575]. Network-based [2312]. Network-on-Chip [2824]. network-wide [2628]. Networking [3112, 1557, 3142]. Networks [3829, 3677, 3270, 1609, 1690, 1720, 1533, 2937, 2766, 1415, 1778, 1499, 1514, 1417, 1386, 3276, 1475, 3113, 3327, 3462, 2313, 2403, 2146, 2578, 2984, 3309, 3787, 2828, 3092, 2576, 1111, 3119, 2257, 2208, 3530, 3711, 3800, 3257, 3220, 3448, 3572, 2897, 2843, 2817, 3008, 3571, 3178, 2451, 3476, 2623, 2790, 2899, 3497, 3624, 2209, 3521, 2604, 2771, 1117, 3610, 3910, 3144, 2938, 3663, 2320, 2621, 2084, 2051, 3819, 2212, 2214, 2767, 227, 3021, 3118, 474, 2769, 2875, 2870, 1177, 3883, 2182, 2768, 3724, 2941, 2902, 3419, 2537, 528, 2602, 464, 3833, 2194, 2939, 3277, 3059, 2886, 687]. networks [1107, 3145, 3937, 3143, 2377, 2321, 2961, 2809, 3025, 2345, 2931, 3312, 2827]. Networks-on-Chip [2827]. Neural [1609, 1690, 1340, 1720, 1415, 1499, 1417, 3220, 3448, 2897, 2051, 1079, 1107]. neurons [2047, 1126, 1216]. neutral [2068]. newcomers [2935]. Newman [2790]. Newton [39, 80]. NEXP [1420]. NEXP-Time [1420]. Next [3276]. nexttime [755]. nfa [2292, 2663]. NFAs [2412]. Nilpotent [1048]. nine [3172]. Nivat [3468]. NL [1524]. NLC [865, 1695, 898]. No [1605, 1657, 159].

No-Z-Cycle [1657]. NoC [2826, 2825, 2802]. Node [382, 3119, 1797, 2383, 584, 3156, 543, 3721, 3747, 2722]. node-deletion [543]. nodes [2767, 752, 3113]. Noether [1248]. Noetherian [487]. Noise [1497, 2120, 2746]. Noisy [1686, 1915]. Nominal [2975, 3631, 2505]. Non [2711, 2364, 1287, 1210, 3724, 3527, 3091, 2985, 1821, 2159, 968, 3526, 596, 3184, 3745, 3546, 3757, 3734, 124, 466, 3673, 2980, 483, 464, 3038, 1894, 3828, 2881, 2932, 863]. non- [466]. non-active [3734]. non-adaptive [3546, 3757]. non-black-box [2159]. non-blocking [464]. non-causality [2985]. non-context-freeness [3526]. non-definability [2980]. Non-deterministic [1287, 1210, 3527, 3673, 483]. Non-essential [3724]. non-integer [3091]. non-interactive [3828, 2932]. non-linear [124]. non-monotone [1821]. non-monotonicity [968]. non-oblivious [596]. non-regularity [3526]. non-repudiation [2881]. non-RMM [3038]. Non-standard [2711]. Non-U-shaped [2364]. non-uniformity [1894]. non-uniqueness [3184]. nonadaptive [2509]. nonaddressable [456]. Nonapproximability [1633, 2611]. Nonassociative [1260, 1624]. nonbinary [3207]. nonclosure [3645]. nonconcave [4]. nonconvex [2061]. noncounting [120]. Nondeterminism [1784, 1372, 1407, 454, 1994, 527, 2131, 795, 1215, 1865]. Nondeterministic [1523, 1321, 1599, 1667, 1693, 1309, 3352, 512, 754, 981, 216, 2143, 961, 1864, 714, 721, 576, 91, 219, 757, 335]. Nondiamond [1123]. nonemptiness [874]. Nonerasing [12]. Nonexistence [357, 38]. Nonexpressibility [1156]. nonhierarchical [1084]. nonisolate [396]. Nonlinear [81, 365, 71, 803, 864, 88, 55, 492, 95, 3069]. nonnegative [3751]. nonprobabilistic [26]. Nonrecursive [1409]. nonreducing [720]. nonregular [52, 685]. Nonserializable [1487]. nontrivial [961]. Nonuniform [1587, 1181, 1716, 2018, 1093]. nonwriting [33]. Norm [1736, 3344, 2474]. Normal [863, 674, 1129, 1797, 1351, 3459, 2122, 719, 838, 1899]. Normality [3091, 3125, 3621]. normalization [2274]. Normalized [3685, 2611]. norms [2154, 57]. Notation [2536]. Note [1833, 1551, 1502, 468, 453, 332, 2959, 580, 1019, 355, 1141, 2221, 875, 643, 470, 45, 2638]. Notice [175]. Notions [1655, 3493, 3671]. novel [2479, 2817, 846, 2795, 2998, 3015]. November [3488, 3556, 3651, 3730, 3792, 3861, 3917]. novo [2086]. NP [1638, 994, 1655, 1554, 1793, 1521, 1510, 2949, 1440, 1590, 2562, 2951, 2563, 1757, 1374, 719, 1182, 3516, 995, 1379, 629, 1256, 543, 643, 1035, 659, 440, 2014, 2043, 978, 1917, 698, 621, 1112, 305, 3294]. NP-Complete [1521, 1374, 440, 543, 305]. NP-Completeness [1655]. NP-Hard [1554, 1590, 2949, 2951]. NP-sets [719]. NSPACE [1433]. NTS [901, 815, 814]. null [1023, 732]. Nullstellensatz [1520]. Number [1605, 2246, 2255, 1431, 1522, 1757, 598, 1403, 1333, 1390, 2984, 3494, 3429, 2777, 1913, 3000, 3663, 1990, 517, 2659, 3238, 3802, 2789, 2373, 3206, 3780, 2445, 41, 3253, 3232, 2930, 3912, 3814, 3207]. number-on-the-forehead [3206]. number-theoretic [517]. numbering [3350]. Numberings [2498, 2613]. Numbers [1448, 1911, 2612, 923, 1976, 2979, 629, 2796, 621]. numeration [2612]. Numeric [1539]. numerical [1166, 88, 106, 607, 2779].

O [2079, 3382, 2803]. O* [2160]. O-efficient [3224]. OBDD [1663, 2136]. OBDD-size [2136]. OBDDs [1663, 1865, 1761]. Obituary [2362, 2494, 725]. Object [1341, 1283, 1692, 3404, 1407, 2709, 2506, 1016, 2320, 1059, 2285, 2870, 2823, 3405]. Object-Oriented [1692, 2709, 2506, 2320, 2285, 2870]. objective [2883, 2902, 2895, 3066]. Objectives [1705, 3196]. objects [3183, 969, 3401, 1064, 3177, 1150, 1231, 1149, 992]. Oblivious [676, 1069, 1470, 2033, 2726, 596, 3166]. observable [3196]. observation [278, 3022]. observational [879]. Observations [523, 621]. observing [888]. obstacle [3729]. obstructions [2907]. obtain [581]. occupancy [2403]. Odd [3727, 1176, 2904]. off [278, 36, 3618]. off-line [36, 3618]. offs [1116, 3370, 1427, 871, 624, 1119, 1724, 3795, 824]. OI [427, 434, 3101]. OLAP [2633]. OLSR [2941]. OLSR-based [2941]. omega [3511, 1980, 3223, 2447]. omega-regular [1980]. on-chip [2828, 2320]. On-Line [1744, 1598, 1455, 1454, 1753, 279, 648, 612, 1136]. Once [1267, 1491, 2716, 896, 2023]. One [2046, 296, 1578, 1929, 3481, 1398, 3011, 753, 746, 534, 1047, 1620, 1436, 2434, 2501, 2448, 2947, 958, 1175, 3592, 1135, 670, 33, 36, 3513, 961, 1169, 303, 2795, 1008, 64, 1993, 3516, 225, 938, 2707, 3109, 260, 181, 2706, 2410, 2469, 339, 625, 2668, 1865, 171, 3547, 649, 302, 1105, 3918, 673]. One-counter [753, 2947, 625, 302, 673]. One-dimensional [2046, 260, 2410, 2469, 339, 2668, 171]. One-inclusion [2434, 2501]. One-layer [746]. one-letter [303]. one-one [649]. one-parameter [2448]. One-sided [1047]. one-state [670]. one-step [181]. one-tape [36, 3513, 961]. One-time [3011, 2795]. one-truth-table [1105]. One-Variable [1620, 3481]. One-Way [1578, 1398, 1436, 1929, 534, 33, 1169, 64, 1993, 2707, 3109, 1865, 3918]. ones [385, 2496]. Online [3849, 2332, 3552, 3764, 3936, 2237, 3618, 1705, 1792, 3799, 3451, 2193, 3550, 3144, 2945, 3719, 2121, 3265]. only [896, 1016, 3203, 2694]. ontological [2341]. ontology [2656, 3139, 2854]. ontology-based [3139]. OODBs [1288]. Open [1225, 573, 3783, 2596, 2597]. Open-Addressing [573]. open-world [2597]. OpenMP [2885]. Operation [144, 3016, 3616]. Operation-preserving [144]. operational [3074]. Operations [1287, 1483, 1669, 1326, 3352, 3479, 902, 1201, 3279, 3290, 858, 744, 454]. Operator [2778, 3496, 25, 870, 2817, 124, 30, 755, 3664, 3559, 1112, 2719]. operator-precedence [3559]. Operators [1769, 3368, 158, 1209, 159, 2993, 2796, 88, 3400]. opinion [2729]. optical [2146, 3316]. Optima [1428]. Optimal [960, 1293, 2033, 3223, 3362, 1785, 1441, 1663, 2667, 1918, 3257, 3420, 1506, 3630, 1979, 2147, 1879, 2304, 987, 2084, 3237, 1392, 1449, 3505, 1243, 1394, 205, 1620, 2203, 497, 57, 3406, 3804, 18, 793, 3451, 2479, 3796, 44, 3877, 2817, 1960, 2651, 3105, 3702, 689, 2414, 3211, 81, 1926, 2498, 113, 2133, 692, 3054, 3911, 3040, 2298, 2387, 544, 87, 2475, 338, 2264, 613, 280, 3369]. Optimal-depth [3257]. Optimality [672, 3409, 2025]. optimally [2085]. Optimisation [3080, 3082]. Optimization [191, 1519, 1440, 1528, 3112, 1286, 1504, 3276, 1075, 774, 3051, 561, 2332, 3385, 2181, 1883, 2644, 3433, 3380, 16, 15, 240, 2279, 930, 2858, 2815, 2832, 1153, 3251, 1076]. Optimized [1539, 3025, 2882]. optimizers [357]. Optimizing [3062, 2816, 2817]. optimum [82, 1213, 3863]. Or-Sets [1324]. Oracle [1501, 1379, 1436, 982, 3223, 958, 1197, 1094]. Oracles [1255, 1346, 1096, 3362, 3523, 643, 3679]. orbital [864]. orchard [2372]. Order [1250, 1642, 1529, 1443, 3790, 1224, 2068, 3593, 2286, 3220, 3304, 3293, 2977, 2273, 1218, 1008, 16, 536, 656, 3531, 3583, 2757, 3313, 2980, 553, 239, 1899, 471, 1808, 1163, 500, 841, 750, 1012, 1227, 497, 607]. order-based [3304]. order-sorted [1012, 1227]. Ordered [1293, 997, 1412, 376, 3307, 2056, 3510, 2691, 3619, 578]. ordering [3524, 3381]. Orderings [1309, 2685, 2234, 3848]. orders [3770]. ordinal [2760, 2467]. ordinal-bounded [2760]. Ordinals [1722]. ordinary [76]. organised [2578]. organization [2802]. organizing [916]. Orientation [1480]. Oriented [1692, 2709, 3840, 3935, 3847, 2506, 2320, 2285, 178, 2870, 2838]. Orienting [3819, 3898]. origin [2881]. Orthogonal [3700, 989, 3931, 1003, 2868]. Orthogonally [989, 3895]. oscillation [3685]. osculatory [109]. Other [1793, 1643, 712, 3350, 661, 1983, 2791, 466, 3205, 1191, 3161]. our [2494, 591]. out-branching [2535]. out-branchings [3394]. Out-of-Core [1429]. out-trees [2535]. outcomes [2651]. Outerplanar [3336, 3510]. outerplanaric [3876]. Outfix [1078]. outlier [1979]. Outliers [2896]. Output [1759, 1163, 1685]. Output-Sensitive [1685]. outputs [2755]. outsourcing [3346, 3653]. overcoming [1164]. overhead [2209]. overheads [1894]. overlap [3167, 190]. overlap-free [3167]. Overlaps [1046]. overlays [2811, 3214]. overloaded [1901]. overtaking [858]. overview [2165]. OWA [2719].

p [394, 2833, 1376, 1561, 1141, 1375, 3516, 995, 684, 3353, 978, 1552, 621, 3208, 3046, 1195, 1438, 1524]. P-FRP [2833]. P-Immunity [1438]. P-Selective [1376, 1375, 621]. P-selectivity [684]. P/poly [1141]. P2P [2771, 2747, 2602, 2770, 3278]. P2P-MapReduce [2747]. P= [994]. PAC [1968, 2093, 2432, 1989, 3828]. PAC-learning [1968]. packages [3592]. packed [38]. Packet [1631, 3432, 2624, 3011, 2794, 3143, 2836]. packets [3147]. Packing [3840, 1612, 1247, 2458, 3451, 3552, 3410, 3250, 2789, 251, 2134, 3161, 941]. packings [3298]. page [961]. Pages [3460, 3500, 3452, 3488, 3542, 3567, 3509, 3537, 3520, 3528, 3556, 3548, 3622, 3658, 3578, 3613, 3594, 3599, 3651, 3637, 3713, 3738, 3666, 3703, 3680, 3695, 3730, 3722, 3766, 3743, 3760, 3749, 3755, 3792, 3777, 1605, 957]. Paging [1246, 1557, 2070, 2286]. Pair [3412, 150]. pairing [1118]. pairings [2658]. Pairs [1424, 1423, 1303, 3871, 2190, 3789]. pairwise [2482]. Palindrome [438]. Pancake [3123]. pancyclicity [3189]. paper [248]. Papers [1629, 1640, 1646, 1654, 1664, 1673, 1683, 1691, 1699, 1708, 1718, 1728, 1739, 1749, 1762, 1773, 1781, 1790, 1801, 1813, 1827, 1845, 1855, 1870, 1887, 1903, 1921, 1932, 1950, 1969, 1984, 1995, 2006, 2015, 2026, 2039, 2052, 2065, 2075, 2087, 2099, 2107, 2115, 2125, 2138, 2186, 2195, 2217, 2230, 2150, 2161, 2170, 2244, 2253, 2266, 3465, 3647, 3576]. para [3294]. para-NP [3294]. Paradigm [1458]. Paradigms [1406]. Paradox [2194]. Parallel [1172, 988, 1366, 1367, 1393, 1301, 1449, 1422, 63, 1238, 1345, 2688, 1391, 935, 2747, 1083, 1564, 921, 1240, 1460, 1685, 1484, 3004, 3081, 468, 960, 186, 370, 290, 770, 3643, 783, 3084, 3019, 1941, 3478, 1896, 3192, 786, 2398, 1897, 714, 3380, 1164, 918, 665, 944, 5, 2290, 1044, 3151, 1100, 1155, 486, 1222, 205, 2397, 179, 1223, 2106, 229, 2400, 335, 3369, 3814]. Parallelism [1484, 542]. parallelizable [2063]. parallelization [1147]. Parallelized [1420]. Parallelizing [2519]. Parameter [512, 1440, 2762, 880, 1453, 2448, 18, 2496, 3423, 3067, 2143, 3380, 1943, 2228, 1946, 569, 3242, 2473, 3269, 2049, 57]. Parameter-preserving [880]. Parameterised [3776, 3064]. parameterization [2614, 1227]. parameterizations [3629]. Parameterized [1317, 2990, 3463, 3204, 3880, 3931, 3758, 3375, 2904, 3603, 3784, 3433, 3899, 3252, 3364, 3289, 3770, 3747, 3294, 1948, 2645, 3104, 3923, 3298, 3735, 1947, 1939, 1949, 2226, 3367, 2394, 3762, 3885, 2542, 3501, 3000, 3671, 3779, 2552, 2585, 2659, 3238, 3845, 2797, 2661, 2997, 3253, 1945, 3879, 3733, 3417, 3814]. Parameterizing [2442]. parameters [3119, 3706, 2452, 3539, 3259, 3158, 3541]. parametric [2174, 2062, 3633, 4]. Parametrization [1012, 3788]. parametrized [1944]. Parikh [2842, 1328, 1959]. Parity [1387, 1042, 2012, 2680, 3639, 1191, 3260, 3062]. Parity-Declustered [1387, 3062]. Parliament [3939]. parsable [412]. parser [187]. parser-Linear [187]. parsers [191, 589, 402, 403, 630, 681, 190, 585]. Parsimonious [3029]. parsimony [3818, 3606]. Parsing [1320, 1028, 3615, 3105, 402, 403, 792, 3544, 137]. Part [3199, 1143, 1144, 127, 3412, 897]. partial [2645, 2614, 2036, 3249, 1212, 2787, 130, 263, 3483]. Partially [1369, 2056, 1471, 3196, 3307, 3040, 1046]. partially-ordered [3307]. Partition [856, 1402, 2202, 3588, 3035, 2568, 2839]. partitioned [3643, 3323, 2816]. partitioning [3896, 3122, 3205, 2826, 3369]. Partitions [1290]. partner [3186]. parts [467]. Party [1536, 3798, 2658]. Parysian [3872]. Pass [1662]. passenger [3199, 3146]. passing [3063, 512]. Password [2199, 2198, 2795, 2694, 2182, 2377]. password-authenticated [2198]. Password-based [2199]. password-only [2694]. passwords [2795, 2416]. past [1209]. Paste [1569, 2723]. Paterson [823]. Path [1568, 1424, 1648, 1764, 1369, 3443, 1446, 1513, 1715, 1303, 3522, 3843, 3132, 3928, 2856, 987, 2105, 2919, 765, 977, 3333, 2060, 1112, 3570, 3137, 2809, 3096, 3421]. Path-contractions [3443]. Paths [1295, 1423, 1514, 1607, 3327, 3298, 3869, 3691, 2190, 3456, 2192, 1192, 3503, 1926, 3266, 2695, 3884, 2031, 2638, 2540, 2692, 3688]. patrolling [3391]. Pattern [1337, 2464, 1317, 1796, 2489, 792, 153, 1620, 358, 453, 3590, 3246, 2594, 875, 2445, 318, 90]. pattern-based [2594]. pattern-finding [875]. patterned [2912]. Patterns [1362, 1670, 1239, 143, 556, 3094, 3897, 537, 2625, 1177, 3103, 3881, 2668, 578]. payload [2836]. payoff [3314]. PCPs [2562]. PCTL [3813]. pda [454]. PDAs [942]. PDL [772, 787]. Peak [3056]. peaked [3343]. pebble [3356, 2546]. pebbles [3935, 914]. Pebbling [3356, 614]. peelings [3715]. peer [3092, 2811, 2214]. peer-to-peer [3092, 2811, 2214]. penalty [1155, 259]. people [2989]. peptide [2086]. perceptron [1184]. Perceptrons [1501, 1342]. Perfect [1085, 2995, 1735, 3477, 2705, 3040, 1003, 2011, 441, 2199]. perfect-information [441]. Perfectly [1342]. Performance [2206, 2343, 2097, 2621, 3014, 2212, 2319, 3822, 2397, 2602, 2211, 2399, 2885, 3234, 3728, 2208, 2322, 3255, 3084, 2623, 3435, 3138, 3265, 2401, 3313, 2803, 3151, 2318, 2622, 2628, 766, 2964, 2626, 2605, 2931]. performance-driven [2318]. Performing [1234]. period [3050, 796]. Periodic [1609, 3051, 1967, 2252, 3846, 2385, 3316]. periodicity [238]. periods [3258]. permanent [1912, 2553]. Permutation [1249, 3850, 2467, 2659, 204, 935, 3772]. Permutations [1637, 1929, 2081, 2723, 3103]. permutive [1900]. Perpetual [3812]. Persistent [1369, 959]. person [44, 441]. personal [3349, 2773]. personality [2898]. personalization [3139, 3049]. Personalized [2730]. perspective [3127, 3598, 2337, 3028]. perspectives [3779]. perturbation [1002]. perturbed [2845]. pervasive [2559, 2600, 2835]. petri [1129, 485, 1597, 1601, 1103, 700, 620, 2528]. Pfaffian [1415, 3283]. Pfair [1967, 2468]. Pfair/ERfair [1967]. PFIN [2366]. PFIN-type [2366]. phase [2574]. phenomenon [3698, 163]. phrase [173, 341]. phrase-structure [173]. phylogenetic [2084, 3819, 3606, 3724, 3912]. phylogenetics [3834]. phylogenies [1808]. physical [2819]. physically [3752]. Picking [3032]. Picture [1014, 3468, 3920, 3407]. pictures [3480, 3202]. Pigeonhole [1779]. pipelined [2314]. Pipelining [1883]. piracy [3166]. Pivoting [1391]. pixel [2808]. pixel-wise [2808]. PL [1626]. PLA [1140]. place [3668]. placement [3119, 2817, 2883, 3664, 3113]. plain [2851, 2411]. Planar [2192, 1301, 1446, 1514, 2839, 1475, 2129, 3691, 990, 779, 3273, 3239, 3297, 3035, 3508, 2112, 213, 1164, 2695, 2289, 834, 957, 2540]. Planarity [1223, 385, 3700, 3434, 944, 3770, 3283, 1316]. plane [804, 3720, 113, 2373, 3284, 615]. Planning [1513, 3240, 3104, 3732, 2531]. plans [2280, 2264]. plate [2807]. Platform [2249, 2742, 3083]. Platform-Independent [2249]. player [767]. players [3565]. plus [2582, 2618, 1045]. PMC [3389]. point [518, 3711, 2704, 1117, 3934, 3584, 1177, 3403, 2197]. point-to-fixed-point [3584]. Pointer [1322, 1679, 1918]. Pointers [1377, 743]. Points [1842, 2984, 3309, 6, 2338, 3429, 3742]. polarized [3270]. Policies [1527, 457, 2103, 3313, 2674, 2992]. policy [2181, 3225, 3005]. Poljak [2997]. poly [1141]. polyautomata [474, 606]. Polygon [1301, 987]. polygons [989, 1213, 1003, 991]. polyhedra [751]. polyhedral [2618, 934, 1006]. polylines [3284]. polylog [2236]. polylog-time [2236]. polylogarithmic [3628]. Polymorphic [846, 1771, 2981]. polymorphism [2320, 477]. Polynomial [1285, 1371, 2155, 1554, 1677, 1320, 1529, 3245, 3665, 2976, 691, 1235, 1415, 3676, 347, 1675, 1437, 1678, 1726, 3091, 994, 2033, 962, 2030, 2472, 958, 3609, 3336, 362, 1913, 867, 1123, 3171, 2457, 1232, 3712, 3390, 2995, 2791, 3393, 451, 265, 1824, 798, 1852, 3188, 892, 1806, 788, 654, 2298, 193, 205, 1189, 2350, 398, 1105]. polynomial-size [962]. polynomial-space [3171, 3393]. Polynomial-Time [1371, 1320, 1437, 1726, 958, 3336, 362, 1913, 1232, 2995]. Polynomially [3718, 1265]. Polynomials [1331, 2157, 2791, 2012, 3393, 156, 223, 3126, 791, 437, 1944, 398, 1822, 812, 811]. polytope [3231]. polytope-based [3231]. poor [2019]. popularity [3143]. popularity-based [3143]. population [3657, 3820, 3740]. Port [699, 1791, 3521]. portal [2773]. Portfolio [2316]. portions [2294]. position [2551]. Positional [1114]. positioning [2894]. Positive [2985, 1365, 978, 3043, 2, 2760, 2308, 1928, 2365]. positiveness [3654]. positives [3233]. positivity [3471, 194]. possession [2744]. possibility [832, 2909]. possible [3806, 3768, 2548]. post [571]. Postman [3252, 3238]. potential [3569, 3639]. pour [378]. Power [1528, 1366, 1667, 1261, 1328, 1798, 1234, 1280, 3151, 1443, 431, 3376, 89, 2448, 3439, 3220, 1138, 1994, 670, 1070, 1211, 1007, 3934, 802, 1065, 1013, 3544, 2047, 349, 2825, 759, 2551, 3056, 2095]. power-efficient [2825]. powered [2628]. Powers [1439, 1278, 3869, 156, 223]. powerset [1110]. PP [1242, 1626, 1501, 1095]. PPA [3525, 3586]. PPA-completeness [3586]. PPAD [3661]. PPAD-complete [3661]. PQ [385, 779, 557]. PQ-graph [557]. PQ-tree [385]. PQ-trees [779]. Practical [1028, 330, 1368, 1558, 2814, 2626, 2835]. Practically [3063, 3411]. Practically-self-stabilizing [3411]. Practice [2631]. PRAM [1392, 1422, 1611, 1086]. PRAMs [1179, 1377]. pre [6]. pre-Hilbert [6]. prebuilt [3791]. Precedence [2414, 2666, 2778, 3559]. precise [3473, 2662]. Precisely [3339]. precision [1189, 1184, 992, 1131]. precomputation [2688]. preconditioning [437]. preconditions [418]. Predecessor [1785, 246, 2356]. Predicate [1413, 1394, 354, 750]. Predicates [1289]. Predicative [1406, 1039]. Predicting [2898]. Prediction [1496, 3817, 2098, 1697, 1037, 2213, 1495, 1454, 2574]. Prediction-preserving [1037]. predictions [2145]. Predictive [2096, 3862, 2482]. predictor [40]. predictor-corrector [40]. preemption [2543]. Preemptive [1901, 2147]. Preface [3677]. preferences [3374]. Prefix [2988, 3539, 3591]. Prefix-suffix [2988]. Pregroup [2781]. Preimage [3589]. premise [3551]. premise-based [3551]. Preprocessing [3844, 2925, 3707, 3495, 2013, 3679]. Presburger [3820, 3499, 445]. Prescribed [1311, 3484]. Presence [1237, 1403, 3234, 3385, 3303, 2120, 2858]. presentations [3073]. preservation [2507, 3859, 3443, 2834]. preserve [37]. preserved [5]. Preserving [2633, 3348, 561, 3085, 2745, 2954, 3375, 144, 2274, 306, 880, 1037, 2930, 3005]. presheaves [3077]. prevent [3115]. price [2257, 1914, 1916]. Priced [1777]. Primality [1680, 383]. prime [647, 2485, 2445, 3232]. prime-block [2485]. Primes [1677, 52]. primitive [76, 2793]. primitives [489]. Principal [99, 115, 668, 133]. Principle [1520, 1779, 128, 3833]. principles [9, 2011]. prior [1863]. Priority [3081, 3872, 1478, 1240, 3122, 3921, 2319, 2031]. Priority-grouping [3081]. Prism [2485]. Privacy [3348, 2633, 1636, 1548, 3085, 2745, 3236, 3080, 3828, 2930, 3005, 2881, 3347, 2746, 2834]. privacy-enhanced [3347]. Privacy-preserving [3348]. Private [1636, 2114, 3644]. Proactive [3006, 3050]. Probabilistic [2366, 1615, 2798, 929, 809, 740, 1704, 1650, 850, 46, 941, 1494, 984, 1333, 928, 495, 3718, 3440, 3862, 2201, 2903, 3033, 3382, 2592, 3625, 735, 3001, 1820, 849, 2585, 3821, 1824, 1190, 3021, 595, 787, 2331, 1222, 26, 2071, 736, 3547, 2725, 3326]. probabilities [2403, 3338, 3395]. Probability [1644, 1270, 573, 1451, 958, 2746]. probable [2306]. Probe [1780, 2176, 2036, 2356]. Problem [1424, 1299, 1751, 1785, 1546, 1788, 2955, 1798, 675, 1852, 2154, 1514, 1322, 3620, 1715, 1536, 1303, 1726, 1662, 1734, 1480, 994, 819, 2202, 804, 18, 369, 2045, 3806, 2548, 3471, 3809, 207, 222, 3286, 3925, 2306, 3850, 3420, 3592, 2664, 2660, 3367, 3669, 405, 3813, 2535, 82, 1940, 3221, 1085, 2491, 2174, 2855, 450, 2954, 1230, 3885, 200, 694, 3501, 2355, 2003, 557, 566, 2149, 3661, 680, 2413, 2659, 1946, 1096, 2135, 2793, 2775, 874, 758, 2105, 2285, 1130, 865, 2036, 2994, 2639, 836, 2695, 3027, 43, 127, 629, 875, 474, 2372, 3584, 3602, 540, 623]. problem [239, 543, 3324, 1805, 2474, 2703, 1100, 268, 1810, 260, 3545, 3721, 3891, 637, 232, 483, 3536, 625, 1809, 3879, 978, 826, 579, 294, 3101, 3417, 3208, 2546, 363, 3186, 2031, 613, 3654, 2722, 2839, 3570, 2869, 3113, 3166, 90, 3781, 3096, 3421, 2756, 2061, 3207, 3252]. Problems [1554, 1521, 1780, 1711, 1510, 1785, 1663, 1573, 1440, 1723, 1312, 1737, 1368, 3588, 1633, 1794, 1374, 1730, 1239, 1658, 1778, 1345, 1245, 601, 1685, 1308, 1602, 118, 1066, 639, 2453, 2646, 2645, 107, 39, 683, 2458, 3496, 561, 2225, 2193, 2390, 3839, 3589, 2110, 325, 741, 1172, 247, 1166, 2472, 2672, 3370, 3823, 3711, 3773, 3694, 3880, 1941, 2291, 2418, 3409, 212, 3311, 172, 2381, 47, 189, 3725, 48, 2788, 365, 1005, 2496, 2237, 3762, 3521, 3776, 3245, 2013, 2542, 3067, 2191, 2904, 2925, 404, 514, 3209, 1225, 807, 1983, 3380, 997, 16, 15, 517, 484, 587, 1926]. problems [2585, 3649, 3195, 235, 101, 121, 3205, 2063, 384, 350, 523, 240, 372, 631, 1231, 3107, 3365, 3783, 3707, 584, 276, 311, 421, 3249, 2121, 113, 1174, 2297, 2846, 2378, 930, 564, 3602, 842, 1084, 3774, 440, 2797, 3701, 181, 3675, 797, 765, 2499, 977, 3161, 1945, 750, 2166, 3438, 3648, 17, 4, 112, 588, 814, 1931, 649, 1195, 305, 726, 3005, 3190, 1199, 894, 607, 1076]. Procedural [1027]. procedure [554, 2900]. procedures [186, 870, 776, 285, 259, 551]. Process [660, 2727, 3282, 1669, 655, 218, 2402, 1198, 3686]. Process-Calculus [1669]. Process-centric [3282]. Processes [1615, 2689, 1601, 645, 2427, 2657, 3255, 3196, 3441, 2229, 3531, 3633, 564, 699, 899, 3052]. Processing [85, 2631, 1513, 3407, 313, 728, 3313, 2747, 2815, 508, 3210, 2816, 3326, 2603, 3010]. Processor [1086, 3098, 2405, 1155, 2251]. Processor-time [1086]. Processors [1293, 1543, 1043, 3270, 3412, 306, 92, 2802]. Produce [1500]. producing [200, 402, 403]. product [3790, 401, 3189, 3418, 47, 3883, 366, 3884]. production [3308]. Products [3045, 3858, 1120]. Professor [3378]. profiling [3689]. profits [3604]. Program [1077, 202, 1713, 1403, 239, 327, 228, 357, 134, 189, 269, 285, 160, 559, 63, 569, 93, 756, 3746, 342, 503]. programmable [3453]. programme [3346]. Programming [1519, 3610, 1278, 1354, 68, 887, 134, 82, 2506, 2759, 3655, 2001, 1983, 1198, 1149, 1024, 343, 477, 3016, 2655, 885, 583, 17, 2822, 2895, 3017, 280, 1041, 2061]. Programs [1254, 1732, 1589, 1409, 675, 1746, 1352, 1279, 1364, 1470, 1076, 165, 2985, 2709, 290, 896, 608, 962, 1026, 22, 512, 236, 3033, 365, 450, 496, 976, 516, 253, 714, 71, 685, 631, 72, 271, 530, 2134, 275, 595, 3188, 94, 61, 760, 231, 84, 941, 835, 1161, 1152, 830, 45, 824, 1214]. progression [2690]. projection [728, 3675]. projections [1881, 771, 3380, 717, 1822]. projective [3227]. promise [2237]. Promotion [3872]. promotions [2781]. Proof [1114, 1265, 1779, 1899, 1380, 1974, 358, 919, 536, 605, 1863, 2454, 3795, 1153, 2528]. proof-tree [1153]. Proofs [663, 1610, 1447, 1821, 943, 2610, 84, 508, 1191, 169]. Propagating [2260, 2359]. propagation [1026, 3360, 2837]. Proper [3706, 3506, 285]. properly [2327]. Properties [70, 1321, 1506, 1362, 1257, 356, 1256, 5, 1373, 1360, 1072, 1351, 3645, 2366, 819, 2851, 3566, 855, 430, 96, 2415, 3740, 3246, 2976, 10, 1990, 3402, 1096, 3707, 2440, 543, 3437, 2156, 2997, 2537, 2350, 1139, 480, 759, 420, 314, 414, 2886, 539, 410, 381, 2321, 1052, 2836]. Property [1657, 385, 3905, 139, 339, 2200, 2778, 3937]. prophet [3678]. Propositional [496, 685, 669, 3593, 1180, 2922, 2534, 3795]. Protected [1766]. Protecting [1636]. protection [3236, 713, 1088, 622, 2881, 2746]. proteins [3208]. protocol [3050, 1981, 2576, 3087, 2223, 2658, 2543, 2608, 2615, 3009, 2906, 2215]. Protocols [1733, 1643, 1558, 1284, 1490, 1420, 3657, 1116, 2157, 3820, 3740, 1094, 3909, 2439, 3585, 3115, 2199, 3653]. provability [927]. provable [1175, 504, 2061]. Provably [1187, 2198, 2615, 3247, 956, 3187, 3023]. prover [178, 1863, 1420, 1380]. provers [1175]. provider [2740]. providers [3604]. provides [3729]. Proving [290, 2355, 304, 1580, 194, 309, 2916, 981]. provision [3066]. Proximity [3142]. proxy [3644, 2626, 3009]. Prune [3041]. Prune-and-search [3041]. pruned [2321]. Pseudo [1564, 1919, 3517, 2637]. pseudo-Boolean [2637]. Pseudo-Random [1564, 1919]. pseudo-telepathy [3517]. pseudoknot [2440]. pseudoknot-bordered [2440]. pseudonym [2906]. pseudoperipheral [752]. Pseudorandom [1534, 1032, 1676, 985, 1116]. PSPACE [2999, 958, 1175, 2243, 3046, 1717]. PSPACE-Complete [1717, 2999]. PTIME [1033]. Public [3329, 2974, 1096, 3341, 3088, 3828]. public-key [1096]. Publicly [3279].

Publisher [1833, 2959]. publishing [3325]. Pumping [475]. pure [2446, 3609]. purely [2227]. pursuit [18, 112]. Pushdown [946, 1575, 166, 1799, 401, 58, 89, 3226, 3352, 2903, 3608, 443, 3424, 2131, 515, 3531, 3583, 3882, 3359, 2776, 938, 490, 69, 703, 3295, 3290, 614, 2778, 952, 1215, 777, 363, 452]. Pushing [3618].

QBF [2999, 3539]. QD [3044]. QD-CFS [3044]. QoS [2343, 2181, 3024, 2342, 2573, 2025, 2742, 2741, 2605]. QoS-aware [2342, 2742]. Quadratic [2873, 640, 1417, 2012, 1976, 2659, 642, 17, 2061]. quadratics [440]. quadtrees [3748]. Qualitative [799, 2124]. quality [3084]. quantificational [572]. quantifier [3790]. Quantifiers [1435, 3306, 598, 2980, 928]. Quantifying [3303]. Quantitative [1980, 3032, 3813, 2902]. quantized [1894]. Quantum [2351, 1775, 1712, 2242, 1591, 1720, 1863, 3044, 3746, 1594, 1682, 1981, 2155, 2780, 2046, 2023, 3660, 2876, 3392, 2037, 2034, 2998, 2509, 3095, 2707, 3109, 2454, 3038, 3596, 3918]. Quantum-assisted [3044]. quantum-dot [2998]. quarantine [2341]. quarnets [3910]. quartet [1943]. Quasi [848, 862, 3524, 2507, 3859, 3712, 249, 2834]. quasi-identifier [2834]. quasi-polynomial [3712]. Quasi-Random [848]. quasi-realtime [249]. Quasi-varieties [862]. quasi-wide [2507, 3859]. quasilinear [212]. quasiminimal [3335]. Queries [1568, 1461, 1571, 1251, 1742, 1310, 1300, 1498, 1346, 1588, 1764, 1765, 2592, 1529, 1483, 1768, 1286, 1580, 1543, 1567, 1541, 1263, 1439, 1331, 1282, 1407, 1795, 1027, 1061, 2850, 2280, 2641, 3384, 2370, 3522, 3030, 1841, 905, 2326, 1968, 2952, 563, 657, 3165, 2141, 2911, 2432, 3214, 2056, 2355, 624, 2856, 2367, 987, 3490, 1065, 2665, 733, 3258, 3366, 1100, 966, 3210, 2860, 2720, 2865, 3120, 2048, 3136, 774, 3302]. Query [1542, 1341, 1497, 1406, 1769, 1505, 1384, 1777, 624, 3351, 1283, 1570, 1286, 1473, 1276, 1513, 1653, 1543, 1324, 1463, 1275, 1504, 3135, 1836, 1465, 1326, 1351, 970, 2155, 2293, 3310, 3933, 2023, 3796, 2591, 3420, 2653, 2817, 857, 1883, 2689, 2654, 2751, 2273, 2593, 728, 1063, 1150, 2034, 2144, 2529, 2858, 2815, 2073, 1153, 2597, 2264, 3137, 2816, 3326, 2603, 2094, 3347, 3826]. Querying [1528, 3138, 1770, 1886, 1652, 853, 1412, 2814, 3001]. question [2973]. Questions [1335, 328]. Queue [1287, 226]. queueing [2216]. Queues [1478, 1240, 3024, 2031, 768]. queuing [2627]. quorum [2176]. quota [3551]. quotient [30].

R [532, 2706]. R-automata [2706].

R-trivial [532]. R. [1114, 898]. R.E. [559]. RA [3115]. radicals [788]. Radio [2169, 1067, 1111, 3869, 3571, 2451, 2875]. RAID [2868]. rail [3199, 3146]. rainbow [3687, 3928]. RAM [631]. RAMs [1784, 600]. Ramsey [2193, 1678]. Ramsey-type [2193]. Random [1910, 2030, 1551, 1381, 1434, 1400, 1366, 1389, 1355, 1379, 1499, 1564, 848, 1682, 1881, 3362, 2119, 958, 2617, 2122, 1197, 3029, 217, 3497, 3854, 2541, 2169, 2795, 2887, 2537, 218, 1919, 3746, 762]. random-access [762]. random-walks [3029]. Randomised [1392]. randomization [1111, 1994, 3387]. Randomized [3026, 1590, 1253, 1513, 1731, 1524, 919, 1040, 2652, 1472, 2948, 2779, 1052]. randomly [2623, 3822]. Randomness [1963, 1738, 1318, 1550, 1787, 1263, 3091, 3124, 1155, 1204, 949, 926]. Range [3384, 2987, 1286, 1326, 2624, 3214, 624, 2603]. Range-max [3384]. rank [524, 3765, 3707, 2636, 628]. rankable [3645]. ranked [2275]. Ranking [1388, 2729, 641, 1029, 3445, 2482]. rarely [1176]. Rate [1614, 1690, 2216, 2203]. rate-based [2203]. rates [3909, 2184, 2393, 2872]. rating [3277]. ratio [3451, 2286, 3410]. Rational [1456, 1124, 668, 109, 759]. rationality [3425, 328]. rationals [3197]. rationnels [506]. ratios [399]. RDBMS [3327]. re [2576, 3159, 3644, 3009]. re-clustering [2576]. re-encryption [3009]. re-revisited [3159]. re-signature [3644]. re-signatures [3644]. reach [3575, 2559]. Reachability [1841, 3608, 1474, 1295, 3940, 3295, 2225, 3634, 2841, 3587, 3633, 3882, 3853, 3438, 3648, 3489]. Reactive [2712]. Read [1267, 1491, 1611, 3115, 896, 2023, 2105]. Read-abortion [3115]. Read-Once [1267, 1491, 2023]. read-once-only [896]. read/write [2105]. readjustable [3050]. ready [917]. Real [2249, 1347, 1659, 2141, 1563, 2870, 171, 2947, 1226, 2612, 3497, 2190, 915, 438, 3122, 265, 3011, 2543, 629, 2823, 2410, 2469, 2250, 2668, 3100, 1127, 621, 2730, 768, 2251, 3022, 633]. Real-Time [1659, 2249, 171, 2947, 915, 3122, 3011, 2543, 2823, 2250, 3100, 2730, 768, 2251, 3022, 633]. Real-Valued [1347, 2141]. real-world [3497]. realisability [3284]. Realistic [1040]. realizability [2177]. realizable [394, 146]. Realization [1723, 467, 793, 557, 526, 415]. Realizations [119, 3876, 2618, 487]. Realizing [2974, 227]. reals [2617, 3699, 2284, 3197]. realtime [249]. rearrangeable [464]. Rearrangement [2464]. rearrangements [1812]. Rearranging [973]. Reasoning [2463, 1585, 1180, 3865, 964, 2158, 1099, 544]. rebalancing [1228]. receivers [2936]. Recognition [1506, 537, 1488, 354, 417, 438, 3779, 974, 3810, 3038, 171, 300, 2807]. Recognizable [3469, 362, 2410, 2469]. recognize [962]. recognized [1055]. Recognizing [1917, 3364, 934, 219]. recolorings [2307, 2388]. recombination [2084]. Recommendation [1704, 2728]. recommender [2811, 3455, 3093, 3065]. Reconfiguration [3397, 3373, 3870, 3555, 3864]. Reconnection [3834]. reconstructibility [2389]. Reconstructing [1804, 3910, 1807]. reconstruction [2084, 3558]. reconstructions [1808]. records [3349]. recourse [3936]. recovery [2994, 2703, 2447, 1151]. rectangle [3405]. Rectangles [1534, 1582, 1004]. rectangular [796]. rectangularization [3405]. Recurrence [3471, 102]. recurrent [3220, 1177]. Recursion [708, 609, 3445, 601, 906, 712, 3033, 404, 501, 202, 745, 967, 519, 131, 280]. Recursion-closed [609]. Recursion-theoretic [3445]. recursions [1047, 220]. Recursive [1459, 1409, 1321, 1765, 1353, 106, 1282, 512, 458, 754, 76, 348, 714, 504, 488, 263, 1033]. recursively [3045, 1150, 3031]. reduce [3844, 585]. Reducibilities [1382, 994]. Reducibility [1375, 1791, 1034, 1123, 1963, 388, 311, 421, 684, 344, 1035, 2043, 1037, 398, 281, 1105]. Reducible [1376, 382, 2236, 709]. Reducing [2424, 3233, 1375, 1757, 3640, 3780, 1787, 3915]. Reduction [1402, 1257, 2490, 981, 2533, 351, 36, 2680, 3015, 2762, 2871, 1894, 3269]. Reductions [1518, 1590, 1253, 1400, 1524, 3169, 2558, 924, 561, 771, 1182]. redundancy [3362, 624, 2547, 2577]. Reeb [3401]. Reed [2189]. Reference [1246, 3072, 2070, 3182]. Referenced [75]. references [399, 3487]. referential [2511]. Refined [902, 3671, 2129, 392]. refinement [608, 3100, 700]. Refinements [1752, 1692, 485]. Refining [2453, 3735, 2625]. Reflective [1354]. reflexive [998, 720]. Refutation [1532]. Refutations [1520]. regeneration [322]. Region [1743, 3184, 3120]. Regional [2802]. Regions [1528, 1786, 3177, 1079]. register [3437, 3295]. registers [3069]. registration [2625]. Regression [1598, 2750, 2545, 3358]. regressor [2754]. regret [3570, 3096, 3421]. Regular [1568, 1104, 1764, 1599, 3697, 1401, 499, 1483, 1693, 1601, 72, 3102, 1017, 877, 3466, 3522, 435, 3334, 3773, 3561, 2678, 3196, 460, 96, 97, 199, 3305, 1085, 3854, 346, 3407, 496, 2056, 2856, 3487, 1864, 2857, 546, 2292, 3340, 1068, 356, 702, 3765, 697, 2775, 350, 523, 836, 3261, 753, 623, 3389, 3591, 2584, 747, 637, 498, 3038, 3483, 667, 620, 1980]. regular-grammars [2584]. Regularity [577, 2947, 3526]. regularized [3178]. Regularizing [3626]. regulated [668]. regulation [3118]. regulator [405]. regulatory [3833]. Reinforcement [1748, 3391]. Rejecting [3357]. rejection [2587]. rejections [1942]. relabeling [1015]. Relabelings [1500]. Related [1780, 1785, 3566, 2452, 3694, 3880, 1165, 2791, 1926, 1078, 2297, 3675, 2921, 588, 1931]. Relating [3521, 3890, 392]. Relation [1504, 2533, 1022, 173, 388, 3154, 1229]. Relational [1651, 1571, 2593, 3077, 410, 1354, 1771, 1326, 1795, 890, 563, 657, 816, 728, 536, 729, 2274, 2270, 574, 966, 1837, 912, 1125]. Relations [1042, 1435, 1036, 1792, 863, 3471, 2550, 3608, 856, 2411, 151, 2024, 2260, 2359, 1022, 1110, 2924, 1206, 472, 852, 414, 732]. relationship [3389]. Relationships [1610, 91]. Relative [1510, 1745, 1379, 1467, 1436, 2286, 580, 1853, 634, 3541]. relativisation [2919]. relativizations [799, 978]. Relativized [1042, 925, 808, 593, 576, 922]. Relaxation [1429]. Relaxed [1661, 1880]. Relaxing [1462, 1467, 3283]. relevance [2674]. relevant [3634, 3396, 2035]. Reliability [3112, 2916, 2884]. Reliable [822, 2753, 1471, 1490, 2826, 915, 2521, 3937]. Remarks [1900, 1492, 601, 703, 95, 1089, 493, 505]. Reminiscences [707, 706]. remote [2339]. removal [1004, 3936, 1979, 997, 1220]. Removing [1155]. Renaissance [3720]. rendezvous [3502]. renewed [2795]. repairability [2888]. repairs [3605, 2859]. repeat [3177]. repeats [2044, 1807]. repetitions [2384, 3446, 2900]. replacement [3615, 951]. Replacements [1797]. Replica [1284]. replicas [2744]. Replicated [1530, 817]. Replication [1581, 2221, 3144, 2910]. Reporting [829]. Representable [1464]. Representation [2617, 1327, 702, 649, 3619, 877, 3745, 2651, 719, 866, 459, 2691, 2918, 2289]. Representations [1506, 1324, 2684, 1309, 3577, 795, 488, 2655, 992]. Representative [3168]. Representing [2612, 2044]. Reproduction [297, 143, 358, 453, 153, 318, 336]. repudiation [2881]. reputation [2771, 2935, 3145, 2835]. reputation-based [2935]. require [492]. required [3238]. requirement [2146]. requirements [2262, 362, 3812]. requires [3172]. rescue [3053]. Reset [525, 2490, 3484]. reset/inhibitor [2490]. residuated [3160]. Residues [2873]. Resilience [1530]. resiliency [3431]. resilient [3329]. Resolution [3470, 1561, 1366, 1558, 2011, 2352, 3757]. resolvable [190]. resolved [3552, 933]. resolvents [32]. Resource [3308, 3729, 2728, 1044, 2559, 662, 2181, 2767, 3118, 618, 2510, 3017, 2770]. Resource-Bounded [3308, 3729, 2559]. resource-constrained [2770]. resources [3308, 2953]. Respect [1473, 1454, 922]. Respecting [1632]. Restrained [3805]. restrict [3624]. Restricted [1503, 1364, 3804, 3718, 2531, 2670, 83, 310, 3266, 255, 623, 3770, 3333, 3492, 3449]. restriction [2549, 1809]. Restrictions [1784, 3283]. restructuring [863, 1083]. result [515, 582, 769, 363]. resulting [3897]. Results [1793, 1618, 1478, 1628, 1353, 1692, 1695, 1098, 1364, 3119, 2110, 1043, 3880, 2291, 3165, 348, 40, 2561, 1926, 951, 1868, 2185, 616, 572, 490, 377, 576, 2388, 84, 893, 2130, 2345]. reticulation [3912]. retraction [2130]. Retrieval [424, 1636, 2727, 2114, 2731, 2735, 2315, 2730]. returns [3650]. Reusing [1413]. Reversal [252, 946, 3479, 35, 874]. Reversal-bounded [252, 946, 3479, 874]. reversals [2382, 1804, 36, 2081]. reverse [3326]. Reversibility [1174, 329]. Reversible [3356, 2776, 1623, 3262, 419, 3654, 3288, 3254]. Revision [1668, 2922]. Revisions [1397]. Revisited [1307, 1882, 3862, 3159, 2221, 2789, 2081, 3692, 3621, 1809, 2488]. Revisiting [2874]. reward [3811]. rewind [20]. rewrite [837, 2294, 3412]. Rewriting [1764, 2720, 1484, 3268, 3473, 1137, 2273, 2979, 2549, 1899, 903, 952]. RF [2936]. RFID [3087, 2609, 3055]. RFID-based [2609]. Riccati [24, 180, 32]. Rich [677, 1672, 1553, 2019]. Riemann [383]. right [137, 229]. Rights [3280]. Rigidity [1724]. Rigorously [2956]. rings [1043, 405]. risk [2482, 2316]. River [1399, 873]. RMM [3038]. rNA [2779]. road [3484]. robber [3846]. Robbers [1885]. Robot [1513, 827, 3515]. robotic [3053]. robots [3895]. Robust [3385, 1617, 2019, 2222, 1672, 2950, 1217, 1005, 2956, 2051, 2838]. robustness [2577]. role [2924, 3007]. role-based [3007]. roommate [2758]. Roommates [1726, 1602]. root [2084, 1189]. root-unknown [2084]. Rooted [1608, 448, 3510, 3417, 3161]. roots [124, 3162]. rotating [2686]. rotation [3895]. rotor [3192, 3504]. rotor-router [3504]. rough [2111, 1131]. round [1175, 2439, 3565, 3190]. round-trip [3190]. rounding [2191]. Rounds [1735, 1055, 2544]. roundtrip [2383]. route [3021, 2815, 2816]. router [3504, 2824, 3113]. Routers [1671]. Routing [1631, 1684, 1441, 1399, 783, 1705, 1702, 1386, 3050, 2383, 2332, 2033, 3030, 3698, 2817, 2167, 3461, 3457, 2419, 2212, 1069, 746, 873, 1319, 613, 3025]. rows [1913]. rubric [2866]. rubric-based [2866]. rugby [3641]. rule [2901, 861, 3323, 3821, 3563]. Rulebases [1413]. Rules [1539, 1540, 3551, 837, 2548, 3471, 2057, 3353, 2902, 2602, 2484, 2490, 45, 3052]. run [3513, 2574]. run-time [2574]. running [749, 1806]. runs [2777]. Runtime [3033]. Rural [3238].

S [1565, 2235]. S. [468]. Safe [1527, 1013]. Safety [1323, 2841, 1088, 2719, 854]. SAFIR [3116]. Sakarovitch [1089]. salesman [338]. Sample [1592, 1472, 1658, 1689, 1619, 1613, 3219, 3715, 3847, 200, 2434, 2501]. samples [2380]. Sampling [1355, 1275, 1910, 2667, 2541, 3734, 1835, 3932]. SAND [3829]. SARP [3050]. SAT [2357, 2242, 3835, 3865, 2355, 3267, 2680, 1681, 3249, 3675, 2473, 3208, 1716, 2671]. SAT-based [3865]. satellites [3022]. Satisfaction [3588, 1653, 2792, 3159, 2550, 2381, 3094, 2659, 3365, 3197, 2488]. satisfactory [2390]. Satisfiability [3194, 1253, 1384, 1576, 1312, 1622, 3491, 1974, 2453, 3496, 3813, 2530, 3602, 3781]. satisfy [1218, 2661]. Satisfying [2944, 3506]. Saturated [1687]. Sauer [3344]. Saving [1263]. Scalability [3054, 2626]. Scalable [2214, 2824, 3032, 2322, 2772, 2967]. Scalar [400, 3438]. Scale [1496, 3817, 2404, 3480, 2885, 3007]. Scale-Sensitive [1496, 3817]. Scaled [2018, 2668]. Scaling [2518, 3198, 807, 3098]. Scan [1388, 1172]. Scattered [67, 3789, 3802, 3878]. scenario [2715]. scenario-based [2715]. scenes [934]. Schaefer [2453]. Scheduler [1394, 2346]. schedulers [3017]. Scheduling [3716, 1293, 1299, 1753, 1869, 2666, 3265, 1902, 3081, 1967, 2806, 3643, 353, 2575, 3409, 1901, 2181, 2468, 2517, 2147, 2604, 2414, 3845, 2319, 3736, 3886, 2543, 3324, 2250, 1127, 2203, 305, 3022, 2869, 3143]. Schema [2857, 2853, 2594, 2596, 2060]. Schemas [1305, 1571, 2271, 348, 2276, 771, 731, 255]. schemata [189, 285, 1170, 160, 63, 239, 228, 342]. schematic [1088]. Scheme [1731, 3710, 2956, 2795, 3011, 1852, 2182, 3015, 3235, 764, 2628, 711, 3187, 2810, 503, 2377, 1002, 3347, 3023]. Schemes [1554, 1636, 1380, 906, 890, 818, 3715, 1081, 3847, 3665, 404, 501, 202, 586, 521, 558, 745, 559, 3021, 93, 3581, 129, 555, 2406, 756, 1125]. Schützenberger [2685, 3272]. Sci [3817, 2359, 3859, 2926, 2185, 3199, 3109, 2469, 2501, 3187]. science [2560, 616, 883, 3378]. Sciences [3933, 3135, 3398, 3319, 3112, 3379]. Scientific [2514, 2520, 2517, 2521, 2516, 2741, 2885, 2515]. scope [260]. scope-3 [260]. scoring [2548, 2866]. SD [3908]. sdmathLMS.zip [1925]. sdmathM [1925]. SDN [3720]. Search [1780, 1510, 1478, 2912, 1375, 1619, 3113, 1974, 2129, 3799, 1228, 1918, 2082, 2242, 916, 3617, 3041, 365, 450, 3844, 1232, 2702, 3921, 3053, 1824, 937, 918, 3565, 3381, 3121, 2887, 568, 847, 3193, 551, 2197, 338, 744, 3347]. searchable [3710, 3573, 3341]. Searching [1649, 3580, 2996, 1975, 1074, 2356]. SEAT [2832]. secluded [3803, 3562, 3878]. Second [1642, 1577, 3376, 2977, 127, 2980]. second-order [2977, 2980]. secrecy [2199]. Secure [2971, 1643, 2960, 2633, 3346, 3573, 3236, 3116, 3323, 3121, 3225, 3007, 3710, 3349, 3117, 2198, 3798, 3341, 3011, 3088, 2521, 3187, 2615, 2810, 3347, 3023]. Secure-channel [3573]. Secure-Multiparty-Computation [2633]. securing [2744]. Security [1656, 3230, 3276, 2832, 2180, 3087, 2735, 2285, 3080, 457, 2769, 3427, 3581, 3232, 3428, 3391, 3322, 2966]. Security-aware [2832]. seed [2387]. seeds [2083, 2082, 1960, 2873, 2304, 2303]. seek [3446]. segment [829]. segmentation [2808]. segmentation-free [2808]. segments [3068, 1804, 2069, 1811]. Selected [3465, 3647, 3576]. Selecting [2641, 3170]. Selection [1710, 2955, 1326, 960, 1026, 221, 3248, 749, 641, 871, 950, 3138, 3938, 2342, 3145]. Selective [1604, 1376, 3612, 1375, 2844, 621]. Selectivity [1355, 1839, 684]. selector [792]. Self [1772, 1556, 1744, 2963, 1034, 1166, 1376, 1583, 3178, 916, 1375, 1319, 1087, 3072, 3804, 2578, 2811, 25, 3791, 3720, 2554, 2956, 3411, 2818, 2912, 1929, 2105, 2405, 2133, 684, 2871, 2211, 1200, 336, 2810, 3937]. Self-adaptive [2963, 2818]. self-adjoint [25]. self-assembly [3804, 3791, 2912]. self-comparative [3937]. Self-Confident [1744]. self-correcting [1200]. self-destructing [2810]. self-emerging [2811]. self-healing [2211]. Self-Join [1772]. self-organised [2578]. Self-organizing [916]. Self-Reducibility [1375, 1034, 684]. Self-Reducible [1376]. self-reference [3072]. Self-regularized [3178]. self-reproduction [336]. Self-routing [1319]. self-similar [2405]. Self-similarity [1087]. self-stabilization [2554]. Self-Stabilizing [1556, 1583, 3720, 2956, 2105, 2133, 2871]. Self-testing [1166]. Self-testing/correcting [1166]. self-witnessing [1929]. selfish [2167, 2194]. Semantic [1324, 2858, 1650, 2732, 855, 2506, 2714, 3341, 3531, 3359, 1018, 2595]. Semantics [2653, 527, 1667, 595, 1408, 1531, 1437, 1278, 3074, 2711, 68, 1039, 1064, 3165, 856, 3865, 3293, 2508, 516, 3075, 2510, 2522, 1851, 510, 1214, 917]. SEMD [3225]. Semi [848, 2618, 3643, 3221, 3910, 1010, 225, 3313]. semi-directed [3910]. semi-FIFO [3313]. semi-matching [3221]. semi-modular [1010]. semi-partitioned [3643]. semi-polyhedral [2618]. Semi-Random [848]. semi-Thue [225]. semiAFLs [158].

Semialgebraic [1605, 1572]. semidefinite [1983]. Semideterministic [1407]. semigroup [394, 337]. Semigroups [679]. semilinear [139, 66, 3197, 98]. Semilinearity [1605, 1572, 1794]. Semirandom [1737]. semiring [3472, 795]. semiring-representations [795]. Semistructured [1648, 1770]. sense [373]. sensing [3232]. Sensitive [1496, 288, 1685, 3817, 867, 108, 1163, 137, 2605, 77]. Sensitivity [1960, 3679]. Sensor [2937, 2576, 2817, 3008, 2814, 2769, 2768, 2537, 2628, 3049, 2939, 2819]. sensor-to-sensor [3049]. sensors [3811]. sentences [1218, 1675]. sentential [222]. Separated [1695]. separates [958]. Separating [1655, 3191, 1294, 391, 3315]. Separation [1501, 1258, 1695, 3420, 3314]. Separations [1329]. Separator [1316, 3864]. Separators [1301, 1948, 2583, 961, 3772, 834, 3512]. September [3548, 3637, 3722, 3777, 3857, 3914]. Sequence [1483, 2618, 83, 3209, 2069, 2485, 2793, 1811, 361, 3100]. Sequences [1526, 1786, 848, 3870, 2248, 3459, 3876, 1911, 1115, 269, 3399, 3107, 2918, 1826, 2389, 1805, 1927, 2354, 577]. Sequencing [2544, 2086]. Sequential [1600, 2145, 1066, 916, 60, 364, 49, 238, 484, 11, 2945, 3719, 146, 227, 564, 294, 100, 762]. Sequentially [3930]. Serial [3814, 3821, 1206]. Serializability [817, 1467]. Serializable [1481]. Serialization [976]. Series [1328, 1798, 3469, 5, 759]. server [3286, 353, 3521, 1188, 2694, 2626, 2953, 2906, 3653]. server-centric [3521]. Servers [1558, 2622]. service [353, 3084, 2517, 2733, 2398, 19, 3604, 2742, 2838, 2406, 2743, 3066, 3145, 3052, 2740]. service-driven [2517]. service-oriented [2838]. services [2867, 2855, 2607, 2339, 3116, 3118, 2342, 2573, 2962, 2211, 3048, 2215]. Set [2527, 2129, 1476, 1321, 1940, 288, 592, 52, 556, 3183, 2045, 1221, 2110, 2640, 3796, 2618, 870, 2418, 3311, 3844, 789, 3899, 2228, 66, 642, 3505, 2270, 2655, 999, 3908, 1105, 1131, 3406, 3814]. set-cover [2045]. Sets [1371, 677, 1605, 1534, 1376, 1595, 1561, 1369, 1572, 2272, 1402, 1375, 1235, 3940, 1324, 1593, 1438, 1524, 3645, 877, 3870, 3309, 3689, 3127, 571, 458, 3555, 460, 550, 489, 2132, 2704, 3495, 3763, 3771, 950, 2374, 2417, 1141, 2912, 411, 2978, 3785, 451, 3845, 719, 1182, 3045, 66, 2613, 3203, 995, 1095, 1024, 2059, 2239, 98, 947, 368, 1091, 659, 1205, 2043, 839, 841, 361, 621, 2690, 1033, 3400, 492, 2111, 2834, 917]. settings [357, 2651]. Several [2842, 198, 357, 176, 792]. severity [2194]. SH [2654]. shadow [1946]. Shannon [1627]. shape [2735]. shaped [2364]. shapes [3895, 422]. Shared [1277, 1479, 3757, 3501, 3805, 19, 718, 655, 1054]. Sharing [1643, 1766, 1703, 1240, 3349, 2771, 2795, 2405, 3886, 3009]. sharp [1974]. Sharpening [1727]. Shattering [1347]. Shelah [3344]. shellsort [810]. shift [585, 3069]. shift-reduce [585]. shifted [3433]. Shifting [2434, 2501]. shifts [3839, 3729]. Short [1776, 3870, 2213]. short-lived [2213]. shortcuts [3915]. Shortest [1424, 1295, 1423, 1446, 2154, 1715, 1303, 1734, 3327, 3850, 3691, 2190, 2192, 1192, 987, 2105, 1945, 2031, 3137, 2692]. Shortest-Path [1446]. showing [372]. Shub [1416]. shuffle [692, 3483, 742]. shuffle-exchange [692]. side [1989, 2968]. sided [1047]. sieves [3298]. SIGACT [274, 345, 188, 136, 214]. Sigmoidal [1415]. signal [3147, 2864]. signaling [1156]. Signature [1643, 3644, 864, 3235, 3059, 2933]. signature-based [864]. signatures [3644, 3044, 2180, 1991, 3597]. signed [3355, 2081, 2079]. signs [3773]. Similar [1786, 3900, 2405]. Similarity [1615, 2299, 2082, 1087]. Simon [3420]. Simple [646, 3407, 517, 2549, 1195, 2327, 3522, 661, 915, 304, 987, 836, 530, 3031, 2261, 834, 1894, 991, 280]. simplification [2853]. Simplified [2528, 242]. Simplifying [2857]. Simulatable [2889, 203]. Simulated [2160, 555]. Simulating [3775, 938, 466]. Simulation [1497, 3119, 1114, 2404, 3530, 515, 3055, 3018, 582, 3077]. Simulations [1682, 1821, 961, 1086]. simulator [3018, 2400]. simulators [231]. simultaneity [542]. Simultaneous [2391, 2157, 55, 194]. Single [426, 3691, 2857, 3878, 1485, 353, 2050, 1147, 2020, 3346, 3423, 3027, 718, 2031, 3343, 3653, 3016]. single-crossing [2020]. Single-exponential [3878, 3423, 3027]. single-minded [2050]. single-peaked [3343]. single-server [3653]. Single-source [3691]. Single-type [2857]. Single-Valued [1485, 426]. sinkhole [2939]. sinkless [1129]. SINR [3736, 3585]. SIP [2626]. situations [2606]. Sixth [345]. Size [676, 1384, 1474, 1713, 1563, 2686, 1235, 1548, 1721, 1275, 1724, 1282, 3050, 962, 3470, 1207, 134, 3835, 2132, 3624, 3784, 310, 1853, 2948, 554, 2706, 3795, 567, 565, 681, 2136, 3315]. size-bounded [2948]. Size-Depth [1724]. Size-Estimation [1474]. Size-Space [676, 3795]. sized [3250, 599, 3210]. Sizes [1772]. skeletons [2132, 1886]. skew [3584]. SLA [2740]. SLA-based [2740]. Sleeping [1337]. Slender [1328]. Sliding [3642, 3514, 2894, 3786, 2667]. Slight [1727]. SLURM [3017]. Smale [1416]. Small [3425, 1423, 1533, 1693, 1472, 1379, 3698, 3262, 3869, 2925, 3000, 2355, 3285, 2374, 2569, 3540, 834, 3881, 3210, 3512]. small-space [2569]. small-world [3698]. smallest [2951]. smart [3234, 3116, 2825, 2416]. Smarter [1334]. smartphones [2837]. Smith [2362, 2363]. Smooth [1776]. smoothing [3233]. Smyth [1007]. SNP [2085]. SO-tgds [2851]. Social [3142, 2790, 3144, 2728, 2898, 2730, 3145, 3143, 2835]. Socialized [2773]. soficness [3729]. Soft [3240, 3100]. Software [1752, 2964, 1850, 2731, 2730, 2740]. Software-as-a-Service [2740]. solar [2628]. solar-powered [2628]. Solitaire [782]. Soliton [1009, 1135]. Solomonoff [1697]. Solovay [3124, 2284]. Solution [247, 1663, 659, 18, 3521, 88, 977, 2106, 3, 112, 3208, 2874, 613, 1980]. Solutions [2852, 1609, 15, 1757, 1396, 3166, 906, 1172, 712, 212, 76, 3611, 1005]. Solvability [788, 372]. solvable [935]. solve [325]. Solving [998, 3823, 1941, 2680, 3260, 3113, 741, 365, 557, 176, 4]. Some [985, 287, 1291, 1711, 1296, 1043, 1663, 354, 1573, 2291, 1077, 303, 484, 1794, 1089, 1964, 2796, 377, 74, 576, 84, 893, 314, 1127, 13, 621, 505, 1040, 2917, 230, 3311, 479, 48, 104, 2542, 10, 101, 451, 3205, 125, 372, 631, 2846, 2372, 668, 738, 441, 328, 95, 410, 2130]. Somewhat [3341]. SONET [2441]. Sort [1430]. Sorted [1369, 641, 1012, 1227]. Sorting [1515, 3172, 3476, 1302, 2081, 1426, 528, 2382, 596, 783, 3257, 1165, 871, 163, 1897, 2080, 2474, 2723, 232, 3564]. SOS [3368]. Sound [2633]. Source [1551, 3327, 3691, 3427, 2031]. source-to-target [3327]. Sources [848, 3303, 2569, 1032, 2568]. SPACE [1565, 895, 1549, 1732, 676, 1582, 3577, 2705, 1369, 1435, 1427, 1622, 313, 2547, 1258, 311, 421, 2794, 1623, 648, 1318, 3795, 1799, 3489, 736, 1716, 1594, 1066, 1116, 7, 596, 2402, 925, 62, 3844, 3171, 1094, 3628, 871, 689, 3393, 3673, 2569, 113, 473, 1119, 1961, 2871, 594, 392, 391, 926, 478, 539, 824, 2725, 922, 3781, 762, 3826]. Space-Bounded [1594, 311, 421, 736, 925]. Space-efficient [2547]. spaced [2083, 1960, 2873, 2304, 2303, 2387]. Spaces [1472, 998, 3832, 2193, 3741, 2579, 6, 247, 47, 82, 1979, 466, 3402, 2241, 2123, 316, 917]. Spanners [2642, 3827, 3656, 3193]. Spanning [1316, 3620, 3705, 3493, 3923, 3877, 446, 2683, 2223, 2786, 1192]. Sparse [1255, 1561, 3247, 2750, 659, 398, 1331, 1524, 3896, 3656, 2642, 3754, 3259, 1182, 995, 3188, 3397, 643, 2887, 1105, 812]. Sparsest [3205]. Sparsification [1316]. sparsity [3497]. Spatial [1605, 1572, 1286, 1541, 1652, 2579, 2175, 2289, 778, 2865, 3120, 2048]. spatial-temporal [2579]. Spatially [2258]. Spatially-decaying [2258]. spatiotemporal [3328, 3010]. Special [274, 3319, 2590, 3829, 2422, 3677, 2153, 2514, 2739, 2960, 3175, 3230, 3454, 345, 2325, 3465, 2863, 3112, 188, 2118, 2430, 2766, 2801, 136, 2302, 2029, 3135, 3150, 214, 3269, 3636, 1378, 2091, 2269, 2677, 2831, 3142, 2600, 3004, 2336, 2572, 789, 3014]. specialization [320]. specific [2731, 3686]. Specification [1497, 882, 2263, 1752, 2657, 884, 2103, 1012, 879]. Specifications [1443, 3467, 3368, 2715, 880, 893, 1227]. Specified [1189, 2672, 2550, 185]. spectra [1008, 2086]. Spectral [1724, 1911, 2752, 2330, 2005]. Spectrum [1427]. speculative [2910]. speed [3098, 3360, 3523, 2794, 3654]. Speeding [1290, 2805, 3597]. Speeding-Up [1290, 3597]. Speedups [786]. spelling [2866]. Spike [127]. Spike-free [127]. spiking [2047]. spin [3535, 3184, 3035]. splicing [3261]. spline [141]. splines [3646, 177]. Split [1322, 3899, 3772]. Split-Find [1322]. Splitting [288, 458]. SpMV [3369]. Spot [1639]. Spot-Checkers [1639]. SQL [3271]. Squad [680, 474]. square [2949, 3503, 3162]. square-free [3503]. squares [3804, 2202]. SR [585]. Stability [1687, 1609, 2623, 1895, 3255, 3008, 3571, 3662, 3099, 3672, 2051, 2025, 1097, 102, 3316]. stabilization [2554]. Stabilizing [1556, 1583, 3063, 3720, 2956, 3411, 2105, 2133, 2871]. Stable [1439, 3871, 2758, 1117]. Stack [1575, 473, 301, 773, 516, 249, 168, 64, 12, 30, 125, 270, 534, 703]. stack-counter [249]. Stackelberg [3391]. stage [3324]. stamped [2180, 1991]. standard [2711, 3341]. standardizing [1201]. Star [97, 3920, 3593, 432, 3367, 96, 117, 2401, 769, 1003, 26, 841]. star-free [3593, 432, 117, 841]. starlike [3492]. starting [3934]. State [27, 2713, 227, 3290, 3918, 1244, 299, 3804, 3459, 3074, 7, 3862, 3593, 2402, 2991, 3818, 2174, 670, 2177, 3035, 849, 113, 3621, 3359, 2871, 261, 949, 179, 722, 3596, 2725, 3288]. state-explosion [2174]. state-space [7, 2402]. Stateless [2344, 2706]. statements [84]. States [1479, 41, 294, 3918]. static [2856, 618]. Stationary [6]. stations [2869]. Statistical [1055, 1497, 1610, 2751, 2392, 1018, 2094, 2836]. statistics [819, 2000]. status [2578]. stealthy [3234]. Steiner [2669, 3925, 3204, 3669, 3429, 1607, 1480]. step [430, 3482, 554, 181]. step-counting [430]. step-size [554]. Steps [1808, 3260]. Stepwise [3100, 700, 485]. stiffness [801]. Stochastic [1452, 3575, 3074, 353, 119, 2678, 2841, 262, 2755, 2354, 45]. stop [3177]. Storage [1772, 362, 2579, 3728, 956, 278, 3331, 3616, 3156, 3121, 2622, 757, 3811, 3007]. storage-constrained [3811]. stored [3746]. stored-program [3746]. stores [938]. storing [395]. storyplan [3809]. Straight [569]. Straight-line [569]. strarvation [2133]. strarvation-free [2133]. Strategic [3939]. Strategies [1777, 1265, 2202, 3108, 44, 3105, 999, 3391]. Strategy [2841, 2894, 1147, 2869, 2746]. stream [2000, 3320]. Streaming [3321, 2814, 2214, 3324, 2608, 3025, 2519]. Streams [1753, 2262, 3448, 2986, 3232, 2475, 2643]. strength [2864]. strength-based [2864]. stretch [1962]. Strict [208, 560, 3545]. String [2894, 1759, 1398, 819, 2660, 1070, 2304, 689, 1169, 2044, 3505, 936, 1100, 1810, 531, 3536, 903, 150, 1917, 1978]. String-Matching [1398]. string-rewriting [903]. string-to-graph [150]. Strings [1381, 1434, 1585, 556, 3287, 3850, 2384, 3505, 2388]. strip [2994, 2703]. Strong [1655, 2226, 981, 1578, 3249, 3419, 3691, 996, 1142, 1993, 159, 711, 3187, 3023]. Strongly [1730, 459, 910]. Structural [819, 3074, 1618, 1321, 3497, 3154, 1408, 51, 1338, 2321, 2851, 3706, 2381, 2689, 3629, 3259, 3158, 3649, 1096, 1103]. structural/temporal [2689]. structurally [2531]. Structure [561, 657, 2411, 237, 680, 1332, 1603, 3437, 678, 1378, 105, 580, 3178, 423, 2284, 476, 173, 448, 1074, 2946, 2003, 3170, 341, 211, 847, 3403, 371, 693, 31, 237]. Structured [1367, 1234, 3135, 165, 3382, 71, 2084, 275, 2331, 2073, 2106, 17, 1196, 756]. Structures [1611, 1512, 872, 143, 186, 370, 358, 453, 353, 482, 3475, 3561, 3844, 2759, 959, 173, 1077, 816, 145, 737, 111, 3211, 2978, 297, 3482, 1964, 3759, 1100, 390, 2682, 645, 883, 761, 471, 769, 568, 583, 912, 744]. struggle [2869]. Study [1280, 2773, 417, 121, 3504, 3055, 2626]. studying [542]. Sturmian [2946]. style [792]. Sub [1756, 2795, 2967]. Sub-exponential [1756]. sub-passwords [2795]. sub-tree [2967]. subadditive [3678]. subclasses [3572]. subcolorable [3364]. Subcomplete [666]. subcubic [3679]. subdigraphs [2279]. subdivision [988]. Subexponential [3563, 3536, 1947]. Subgraph [3566, 2283, 2904, 2925, 997, 540, 2895, 3908, 3932]. subgraphs [2032, 3904, 3925, 3877, 2415, 3456, 3495, 2701, 3562, 3415, 3878, 3064, 3670]. sublanguages [870]. Sublearning [1658]. sublinear [3938]. Sublogarithmic [2178]. submatrix [2496]. submodels [845]. submodular [3702]. Subproblem [3381]. Subquadratic [3546]. Subrecursive [134, 189, 105, 293, 347]. subsequence [3924, 1805, 1945]. subsequences [3478].

Subset [1751, 3727, 2955, 1852, 3716, 3796, 3735, 3605]. Subset-Sum [1852]. Subset-Sums [1751]. Subsets [1597]. subshifts [3128, 3335]. Substitution [1413, 157, 154, 3839, 1205]. substitutions [3631]. substring [1839, 1884]. subsumption [2144]. subtractor [2998]. Subtyping [1288]. Subword [1959, 2551]. succession [3471]. Successor [1435]. Succinct [877, 3745, 3820, 3097, 2562, 2691]. succinctly [2550]. Succinctness [2594, 2646]. Sufficient [1346, 343, 1965]. Suffix [1449, 2309, 3334, 2988]. suffix-free [3334]. suggestions [13]. Sum [1852, 3716, 2202, 1982, 3435, 3191]. sum-of-squares [2202]. summarization [2986]. Sums [1751, 2021, 2551]. super [3519, 437, 3675, 3582]. super-preconditioning [437]. super-SAT [3675]. supercomputers [2399, 2885]. Superconcentrators [1176, 632, 599, 1319]. Superdeterministic [514]. Superlinear [1254]. Superpositions [1457]. supersequence [3850, 1945]. superstrings [533]. support [2804, 3086, 744]. supported [2992]. Supporting [1671, 3236, 1231, 3121, 2521, 847, 2742]. supports [759]. surface [1004, 1220]. Surfaces [1562]. Surjectivity [335, 186, 1230, 1174, 486]. Surpassing [1162]. Survey [1550, 2884, 2678, 2937, 2965, 2459, 2259]. survivable [3885]. SVC [2931]. SVD [3065]. SVD-based [3065]. SVP [1590]. swapping [3930]. Sweeping [1721, 963, 2686, 565]. Switching [162, 250, 464, 2314]. SWMR [3063]. Sybil [3769]. syllogistic [870]. Sylow [798]. symbolic [835, 667, 1002]. symbols [38]. Symmetric [2157, 3490, 1044, 462, 999, 3491, 1822, 1161]. Symmetries [2446]. Symmetry [1200, 3788, 2724]. Symposium [274, 3829, 345, 3647, 2029, 3576, 214, 250, 162]. synchronisation [1059]. Synchronization [680, 390, 886, 832, 1992, 2276, 2543, 474, 542, 3115, 2730]. Synchronized [1484, 3226, 3265]. Synchronizing [3787, 3441]. Synchronous [3105, 3157, 606, 786, 466, 3040, 1851]. synchrony [3411]. Synonymous [2655]. Syntactic [158]. syntactical [173]. Syntax [58, 1437, 70, 3631, 508]. syntax-directed [508]. Synthesis [2712, 3097, 2715, 1403, 3467, 623]. synthesized [2855]. Synthesizers [1564]. Synthesizing [1686, 3541]. Syst [3817, 3859, 3199, 3109, 3187]. System [3319, 3933, 2359, 3112, 3379, 2926, 2185, 3135, 2469, 2501, 3378, 1755, 3082, 549, 3349, 919, 34, 2811, 3119, 712, 3084, 2745, 2612, 2944, 351, 2398, 2883, 2574, 536, 2609, 3156, 3261, 1124, 9, 2803, 3325, 618, 801, 747, 2871, 764, 2864, 2628, 903, 548, 2964, 3093, 3908, 649, 2953, 2835]. Systematic [2208, 351]. Systems [1428, 2249, 1604, 1687, 1659, 1265, 524, 674, 1544, 1752, 3014, 2766, 1530, 1336, 1704, 1484, 3240, 1380, 3004, 2850, 2884, 3063, 2344, 2206, 18, 3074, 7, 27, 3183, 2225, 3160, 2711, 2672, 3728, 2404, 2845, 882, 3184, 1850, 405, 1901, 2224, 172, 888, 2788, 2956, 23, 476, 541, 2863, 3097, 2207, 3610, 1225, 3122, 2728, 546, 111, 3035, 2320, 2731, 2176, 521, 463, 2238, 2135, 2793, 951, 225, 262, 3080, 3511, 803, 864, 19, 1863, 3882, 564, 3040, 2104, 3353, 3121, 3412, 3455, 1836, 2823, 2706, 298, 570, 2454, 2397, 3150, 3795, 2935, 2832, 318]. systems [2577, 323, 2936, 2106, 583, 2910, 3, 3541, 826, 267, 487, 519, 2337, 952, 3848, 2822, 3316, 1103, 3208, 3046, 169, 2407, 2838, 2819, 2741, 322, 2251, 2600, 57, 2336, 3007, 3065, 917]. systems-on-chip [2956]. systolic [1203, 846, 801].

Table [1524, 3030, 1074, 3554, 1105]. Tables [1711, 1913, 1992, 522, 3242, 367, 2906]. tabletop [3055]. Tabu [3113]. Tailored [1304]. take [622]. take-grant [622]. Taking [1374]. TAL [1488]. Taming [2368]. tamper [3939]. tampering [3329]. tandem [2044, 1807]. tangible [3055]. Tape [115, 35, 164, 207, 114, 36, 3513, 961, 125, 310, 598, 938, 91, 219, 286, 319, 768]. Tape-bounded [115, 114, 286]. tape-complexity [319]. Tape-reversal [35]. tapes [242, 460]. TARD [3280]. tardiness [2468]. target [3327, 3575, 2147]. Task [3382, 3122, 2731, 1902]. Task-specific [2731]. Task-structured [3382]. Tasks [1293, 1967, 2252, 3865, 2666]. Taverna [2522]. taxes [2167]. taxonomy [3313, 1185, 2865, 3120]. TBR [3834]. TC [1625]. TCAM [2794]. TCEP [3664]. Tchakaloff [65]. Tchebycheff [177]. TCP [2213]. Teacher [1458, 1349]. Teaching [1236, 1334, 1442]. team [2366, 2364, 3865, 3053]. teams [2368]. Technical [2221]. Technique [1580, 2931, 2129, 34, 3314, 780, 981, 2057, 718, 3161, 112]. Techniques [391, 1183, 14, 3865, 3215, 2459, 2317, 830]. technology [3084, 2801, 3759]. telepathy [3517]. telephone [2178]. tells [1867]. template [726]. templates [2792]. Temporal [1542, 3462, 3514, 3656, 1313, 1274, 1778, 1336, 3940, 3853, 3575, 3650, 2579, 3827, 3855, 3877, 2520, 3929, 1138, 3852, 3897, 2689, 776, 3587, 3624, 3635, 3776, 3846, 1198, 3863, 755, 3772, 3674, 3642, 3794, 3898, 778, 1854, 2074, 2484, 3512]. Temporary [3280]. tend [842]. tensor [628]. Tensors [3283]. tent [3584]. tentative [2215]. Term [1300, 3506, 475, 1899, 2690, 695]. terminal [3311, 3545]. termination [837, 253, 861, 3649]. Terms [1325, 1548, 2505, 2021, 1082, 125, 2511, 2981]. ternary [2659]. terrain [3553]. Tessellation [143, 186, 370, 358, 453, 482, 297, 260, 761, 339, 153, 414, 90].

Tessellations [182]. Test [677, 550, 1235, 2110, 3185, 3341, 966, 900]. Testability [1500]. Testable [2757, 167]. tested [654]. Testing [2032, 1694, 385, 1316, 2002, 733, 1658, 1323, 281, 1244, 2022, 3360, 2867, 3546, 3700, 1048, 3340, 3224, 1853, 2389, 2200, 1223, 3489, 686, 3005]. testing/correcting [1166]. Tests [697, 1714, 383, 218]. Text [1528]. Texts [1796, 1719]. texture [2735]. tgds [2851]. Theft [622]. Their [1458, 1483, 1719, 1564, 592, 3073, 940, 3832, 3708, 3124, 818, 884, 140, 3398, 3220, 3572, 727, 76, 104, 2704, 2272, 240, 3107, 1807, 2440, 3271, 2796, 3076, 3581, 26, 2886, 539, 190, 633]. them [410]. Theorem [1518, 1312, 1443, 2916, 2453, 358, 378, 720, 3468, 3611, 2644, 647, 728, 644, 1089, 1109, 866, 798, 178, 2918, 5, 229, 170, 2528, 3251, 1002, 206]. théorème [378]. theorems [2193, 2685, 230, 1123, 3216, 702, 133]. Theoretic [1433, 1362, 1455, 1368, 1698, 2114, 2975, 1162, 1885, 517, 204, 3445, 3604, 612, 1140, 381, 830]. Theoretical [3319, 2814, 2426, 3060]. Theoretically [2633]. Theoretically-Sound [2633]. Theories [242, 2975, 1660, 352, 712, 708, 1143, 1144, 715, 911, 609, 663, 2655, 897]. Theory [274, 2128, 2256, 3677, 162, 416, 417, 345, 2631, 3465, 2583, 3647, 2118, 2430, 1578, 2029, 3576, 214, 1678, 3636, 1378, 1326, 3004, 2559, 549, 2590, 2422, 7, 2917, 1092, 817, 68, 24, 870, 2381, 82, 2325, 3699, 351, 200, 145, 163, 499, 10, 111, 3386, 101, 121, 463, 866, 802, 864, 522, 5, 1100, 1157, 130, 211, 1899, 106, 477, 73, 769, 185, 690, 570, 59, 148, 1139, 46, 927, 449, 13, 1033, 2091, 2243, 2269, 21, 31, 1041, 250]. There [990]. thermal [3152]. Thiagarajan [3561]. Thirteenth [250]. those [962]. threats [2965]. Three [1169, 1536, 2895, 1380, 133, 2660, 915, 2574, 1217, 1214]. three-dimensional [915, 1217]. Three-objective [2895]. Three-Party [1536]. three-phase [2574]. Three-prover [1380]. three-string [2660]. three-valued [1214]. Threshold [1310, 831, 1253, 1395, 1134, 2113, 1421, 1470, 1974, 3740, 3546, 2198, 3682, 1825, 2926, 2543, 1044, 2350, 921, 1136]. Thresholding [2537]. thresholds [1042, 3734, 2004]. Throughput [1705, 1867]. Thue [225]. Tie [1408]. Tie-Breaking [1408]. Tight [2110, 1576, 3574, 3000, 3663, 1993, 1284, 2378, 1721, 2291, 2637, 3410, 2038, 2552, 2585]. Tighter [1780, 3215]. tile [2704, 2912]. tiling [840]. Time [1066, 1285, 1371, 1784, 1515, 1554, 1311, 1732, 1612, 1430, 221, 3370, 114, 140, 1659, 1320, 217, 60, 1529, 1628, 1427, 83, 1622, 689, 1392, 1198, 1738, 1422, 1714, 3886, 1420, 1488, 1437, 539, 1716, 1438, 824, 1475, 762, 3514, 3091, 994, 872, 2033, 1116, 2249, 1111, 2030, 3708, 3360, 2947, 596, 593, 2903, 958, 3357, 3617, 3336, 216, 278, 362, 1913, 1165, 3192, 749, 1179, 1123, 3693, 776, 2457, 2209, 2192, 3798, 3665, 1232, 3628, 1074, 871, 624, 3712, 789, 915, 2724, 3513, 438, 3390, 3122, 1897, 2574, 2236, 2995, 284, 2069, 2795, 3331, 1990, 1976]. time [2216, 2044, 964, 451, 463, 1946, 838, 3938, 745, 125, 1824, 3249, 691, 3011, 803, 1209, 1213, 798, 530, 2297, 2695, 2543, 19, 1977, 3504, 892, 1806, 788, 2919, 2290, 938, 3024, 654, 2298, 2823, 102, 576, 847, 2410, 2469, 164, 520, 612, 594, 2250, 1086, 3100, 3679, 1854, 588, 1127, 1726, 2180, 2730, 926, 171, 3208, 1991, 492, 2031, 300, 768, 2251, 1107, 3022, 1105, 2816, 633, 3826]. Time- [114]. Time-approximation [3370]. Time-bounded [140, 838, 164]. time-complexity [1111]. Time-extraction [1198]. time-free [3208]. time-luck [593]. time-partitioned [2816]. Time-restricted [83]. time-shared [19]. Time-sharing [3886]. Time-Space [1732, 1427, 1622, 1716, 1066, 539, 824, 762, 1116, 596, 871, 3826]. Time-space-optimal [689]. time-stamped [2180, 1991]. time-storage [278]. time-symmetry [2724]. Time-varying [60, 2216]. timed [3608]. timeline [3922]. times [353, 272, 3587, 2398, 226]. Timestamps [1542]. timetable [3199, 3146]. tissue [3353, 3046]. Tits [1560]. TOC [1815, 1830, 1818, 1858, 1873, 1891, 1906, 1923, 1953, 1848, 1935, 1874, 1937, 1972, 1986, 1997, 2009, 2017, 2041, 2055, 1998, 2042, 2067, 2077, 2089, 2102, 2109, 2117, 2127, 2140, 2090, 2152, 2163, 2172]. together [3918]. token [3786]. tokens [3775, 3930]. Tolerance [1497, 1659, 3886, 2393, 2872, 3582]. Tolerant [2200, 3827, 3189, 2772, 2956, 317, 316, 3492, 2314, 3320, 2961]. tolerate [2985]. Tolerating [1686, 1390]. Tomography [1723]. Tool [1402, 3241, 2229, 1110]. Tools [1280, 2822]. Top [1507, 1547, 3268, 882, 1211, 2457, 3326]. top- [3326]. Top-Down [1507, 1547, 3268, 882, 1211, 2457]. TopCluster [2804]. topics [2801, 2880]. Topological [3401, 3399, 1541, 1652, 2345, 2828, 1226, 537, 1109, 389, 2886]. Topologically [963]. Topologies [1666, 3923, 2827, 3337]. topology [2383, 3398, 2156, 927, 1916]. topology-independent [2383]. Tortoise [1535]. Torus [1390, 545, 2314, 3884, 2321]. Total [1578, 3386, 1993, 2290, 471]. Totality [1408]. tour [2609]. tournaments [2526, 2640]. tours [770, 338]. Trace [3075, 3427]. Traces [1760, 3482, 2210]. Track [3610]. Tracking [1772, 3399, 2864]. Tractability [1440, 3603, 3186, 2663, 3165, 2381, 2143, 3380, 1946, 2261, 3269]. Tractable [1768, 1283, 1356, 2860, 3242, 2473, 2049]. Trade [1427, 1119, 1724, 1116, 3370, 278, 871, 624, 3795, 824]. Trade-offs [1427, 1119, 1724, 1116, 3370, 871, 624, 3795, 824]. tradeoff [596, 593, 2025, 3168]. tradeoff-based [3168]. Tradeoffs [1732, 676, 1622, 1716, 1066, 662, 2794, 796, 1086, 588, 539, 762]. Trading [3255, 1714, 1484]. Traffic [1631, 2343, 2624, 2207, 2319, 2405, 226, 3055, 821, 3024, 2627, 2968, 2836]. trajectories [2484, 3176]. Trajectory [3175, 2529, 3178, 3179]. Trajectory-based [3175, 3178, 3179]. Trans [1192]. Trans-dichotomous [1192]. Transaction [1059, 713, 2339]. transactional [2833, 2910]. Transactions [1467, 2867, 1025, 971, 2250, 854]. Transducer [1485, 380]. Transducers [1651, 1526, 1507, 1759, 426, 511, 541, 670, 805, 1211, 3424, 3216, 484, 558, 289, 179, 1837, 876]. Transductions [1642, 452]. transfer [2726, 2213, 3166]. transfers [3062]. transform [522, 2483, 762]. Transformation [3895, 1754, 1752, 679, 3453, 606, 1988, 1153, 3328]. transformational [2709]. Transformations [1411, 1603, 299, 3268, 236, 886, 2457, 3375, 2261, 182]. Transformed [1761]. transformers [1838]. Transforming [1562, 1279, 1485, 1965, 645]. transforms [254, 148]. transit [3199, 3146]. Transition [1604, 917, 3074, 2412, 1867]. Transitions [3664, 1864]. Transitive [1474, 1789, 1164, 3235, 2687]. transitive-closure [1164]. transitively [3898]. Translatability [271, 255]. Translating [1693, 131]. Translation [1864, 1211, 586, 3404, 2148]. translations [70, 58, 277, 150]. translocation [2305, 3564, 2079]. Transmission [1490, 2769]. Transmissions [1703]. transpositions [2080]. transreversals [2080]. Transversals [3727, 3444, 3640]. trap [1081]. travel [3177]. traveling [338]. traversal [1115, 1826]. Tree [635, 639, 1556, 1642, 104, 541, 1507, 1759, 3380, 1449, 291, 1234, 1547, 3620, 567, 777, 3137, 1480, 1448, 2129, 413, 3493, 3268, 3470, 3472, 3862, 385, 3850, 3184, 3572, 3669, 3905, 1022, 2683, 365, 2223, 670, 805, 1211, 1203, 2038, 3154, 2786, 3216, 3285, 2857, 771, 728, 975, 2754, 3690, 3591, 2271, 3193, 380, 1153, 952, 876, 3373, 774, 2967, 3834]. tree-based [2754]. tree-child [3850, 3572]. tree-codifications [413]. tree-depth [3373]. tree-generative [3690]. tree-like [3470]. Tree-size [567]. Tree-Structured [1234]. Tree-to-Graph [1507, 1211]. Trees [1310, 1478, 1316, 1661, 1966, 1669, 1596, 1339, 3332, 3705, 3741, 3601, 3923, 1178, 3915, 2591, 712, 3474, 1228, 1918, 3850, 2817, 2761, 779, 446, 2535, 3429, 3198, 3803, 448, 2532, 661, 2003, 3510, 1162, 1192, 499, 1869, 2691, 1148, 3606, 3356, 1880, 821, 3155, 2762, 743, 2271, 488, 2388, 2073, 3193, 3492, 858, 3417, 2358, 693, 21, 3619, 3570, 3912, 3096, 3421, 3207]. Treewidth [1533, 3533, 3700, 3627, 3655, 2283, 2488]. Trevisan [1787]. Triadic [3641]. trial [3365]. Triangle [3934, 3463, 3367, 3524, 3224]. triangle-free [3524]. Triangulation [1301, 1213]. Tribonacci [2979]. trichotomy [3522, 2500]. triconnected [213]. Trie [424, 3759, 744]. trie-based [3759]. Tries [672]. trip [3190]. Triples [1402]. Triplet [3161]. trivial [3943, 532, 444]. tropical [3472]. Trotter [2644, 3251]. Trust [2962, 2606, 2937, 3080, 3093, 2838]. Trustworthiness [2819]. Trustworthy [2770, 3278, 3145]. Trustworthy-based [2770]. Truth [1524, 3534, 1105]. Truth-Table [1524]. Truthful [2652]. TSP [3788, 2173, 3129, 3692]. TTL [1850]. Tucker [3525]. tuple [2579, 2993]. tuple-spaces [2579]. turbo [2001]. turbo-like [2001]. Turing [369, 3108, 115, 114, 3262, 1939, 981, 3629, 36, 3513, 438, 961, 35, 125, 310, 3266, 995, 1095, 582, 3881, 164, 3752, 1865, 3654]. Turing-complete [995]. turn [2841, 587, 625, 149]. turn-based [2841]. Turzík [2997]. Tutte [2791]. tweet [3114]. twenty [3172]. twenty-five [3172]. Two [301, 1796, 230, 1588, 1598, 1582, 3478, 1051, 1884, 1340, 3039, 2694, 289, 2309, 490, 232, 2416, 110, 1055, 3535, 207, 1114, 44, 1175, 3592, 50, 1994, 541, 3900, 2432, 2398, 2355, 3035, 3355, 2658, 974, 197, 836, 2565, 875, 226, 938, 3324, 625, 2837, 179, 767, 441, 768, 3288, 2540, 3826]. two-class [2398]. Two-Dimensional [1796, 1582, 1884, 1994, 3355, 974, 2837]. Two-factor [2416]. Two-Layer [1340]. two-level [2432]. two-party [2658]. two-person [44, 441]. two-player [767]. Two-server [2694]. two-spin [3535]. two-stage [3324]. two-state [3035, 3288]. two-tape [207, 768]. two-variable [875]. Two-way [301, 289, 110, 1114, 50, 541, 197, 226, 179, 3826]. Type [1270, 1692, 1437, 1771, 2366, 2193, 417, 388, 2714, 2857, 3686, 259, 477, 880, 510]. Type-2 [1437]. type-logical [2714]. type-specific [3686]. Typechecking [1838, 1882, 2261]. Typed [1752, 1150, 845, 726]. Types [1666, 1351, 720, 716, 617, 1060, 1065, 1212, 1091, 2858, 885, 2981].

U [2364]. UAV [3277]. UAVs [3811]. ubiquitous [2773, 2607, 2832, 3059]. UbiSafe [2831]. UG [2955]. UG-hard [2955]. Ulam [3611]. ultimate [2135]. Ultra [2691, 2956]. ultra-robust [2956]. Ultra-succinct [2691]. unaggregated [2986]. unambiguous [3782]. Unary [1799, 3903, 433]. Unbound [1439]. Unbounded [790, 3033, 1820]. unboundedness [2844]. uncertain [3384, 3806, 3326]. uncertainty [3385, 2529, 2483]. uncoordinated [2516]. Undecidability [1628, 2284, 772, 500, 1088, 3654]. Undecidable [1687, 101, 121, 836, 189]. under-specified [2672]. Understanding [3586, 3177]. underwater [2809]. undetermined [2721]. Undirected [1427, 1303, 1682, 2480, 3819]. Unequal [573]. Unequal-Probability [573]. unfoldments [253]. unidirectional [3644]. Unification [1041, 1082, 439, 823]. unified [2381, 3386, 3018, 3168]. Uniform [1587, 1496, 3615, 1431, 2175, 3205, 1825, 2926, 838, 1738, 1473, 1271, 1268, 109, 3817, 3551, 1968, 821, 597, 2095, 917]. Uniformity [1462, 1031, 1894]. Uniformly [292, 3368, 2179]. unifying [3093]. Unilateral [137, 3852]. unimodular [3242]. Union [1322, 893, 870, 789, 66, 2240]. Union-Find [1322]. Unions [1582, 2907, 2417]. Unique [2357, 1253, 3569, 1170, 2923, 3676]. uniquely [3042]. uniqueness [3184]. Unit [3396, 1869, 3881, 600]. unit-capacity [1869]. Unitarity [2582]. units [3186]. Universal [494, 3571, 1826, 1504, 526, 1183, 2780, 1115, 3550, 285, 2613, 1229, 3235, 87, 3752, 2136, 469]. Universal-Relation [1504]. universality [3924, 3335, 2227, 419]. Universally [1481, 2671]. universes [2721]. Unknown [1684, 2084, 3511, 2406]. Unlabeled [3715, 3828]. Unmanned [3461]. unordered [3474]. unpredictability [949]. unranked [2271, 2073]. Unreliable [1498, 1471, 3303, 3938]. unrestricted [1865]. Unsatisfiability [1395]. unsatisfiable [2049]. Unshuffling [2949]. unsigned [2305, 2723, 3564]. unsolvability [222]. Unspecified [1618]. unstructured [2106]. Unsupervised [2836]. untestable [2757]. until [1854]. untrusted [3849, 3346]. untyped [2981]. Unweighted [1303, 2696]. Update [1668, 3182, 2463, 3475, 568, 2340]. Updates [672, 1027, 1061, 3554, 1880, 3684]. updating [446]. UPnP [2339]. Upper [1587, 2616, 871, 656, 1284, 618, 2637, 14, 810, 445]. urban [3053]. urgency [3812]. Usage [1494]. Use [1483, 3387, 592, 3223, 3644, 121, 718]. useful [372]. User [2727, 3143, 2811, 2574, 2625, 3156, 2898, 3093, 2730, 2615]. users [3822, 2864, 2194]. Using [1707, 2280, 2579, 1776, 2462, 3539, 1462, 2320, 2329, 1429, 1543, 1746, 1112, 2916, 877, 2641, 3383, 3710, 2370, 2828, 3796, 385, 2402, 3720, 2897, 3350, 779, 2624, 2827, 2873, 2339, 2654, 1804, 2866, 3910, 3259, 2092, 2658, 2057, 2735, 3199, 3146, 3511, 2769, 2688, 2059, 3883, 2445, 3686, 528, 2900, 2837, 2825, 1886, 2720, 2936, 2197, 2399, 3277, 2895, 2674, 2992, 539, 3086, 3025, 3369, 3838, 2836, 2967]. Utility [2823]. utilization [3630].

V [870, 1372]. vaccine [3086]. Vacillatory [1207, 2364]. valence [654]. validity [2257]. valuable [2581]. value [107, 39, 212, 3094, 176, 629, 2371, 1097]. Valued [1347, 1485, 3108, 426, 2467, 3416, 2141, 3297, 1930, 1126, 2900, 184, 655, 2921, 2679, 2909, 1214]. Values [1618, 1882, 2435, 1023, 272, 634, 2442, 732]. Vandeurzen [1605]. VANET [3088, 3225]. Vapnik [1552]. Variable [3094, 1620, 1131, 3250, 1008, 3481, 875, 718, 655, 2049]. variables [877, 165, 2659, 2035]. Variant [1369]. Variants [1374, 1725, 3687, 3899, 2372, 3536]. Variations [780, 3456]. Varieties [376, 862, 1120]. Various [295, 417]. varying [60, 489, 2216]. VASS [2844]. VC [1415, 1417, 1823]. Vector [2083, 546, 2154, 3438, 2725, 400, 244, 2788, 3534, 3618, 2241, 951, 349, 826, 3086]. vectors [3068, 2295, 3170]. Vegas [1994]. vehicle [3277]. Vehicles [3461]. vehicular [3277]. velocity [3021]. velocity-aware [3021]. verbal [3447]. Verifiable [1643, 3279]. Verification [2709, 2710, 1698, 1837, 2677, 2448, 3030, 3114, 3791, 2263, 2137, 3316, 3597, 2906]. verified [3422]. verifier [3235]. verifiers [1094]. verify [1865]. Verifying [2991, 3513]. Version [1416, 2624, 3139, 764]. Versions [1602, 1084]. verso [905]. versus [1542, 1784, 2646, 1039, 1377, 624, 601, 594, 1865, 1456, 926, 1107]. Vertex [2353, 3575, 3514, 3843, 1949, 2418, 2761, 2535, 3854, 3844, 3928, 3423, 3758, 2616, 2191, 2786, 2925, 2142, 997, 2228, 3707, 3396, 3780, 947, 2839, 2638, 2540]. vertex-colored [2616]. vertex-connected [3758]. vertex-deletion [3707]. vertex-disjoint [2540]. vertical [2825]. vertices [1990, 3396, 2130]. very [744]. via [3716, 3376, 2580, 3496, 2424, 1742, 3030, 1911, 2479, 3823, 1644, 2817, 2226, 3094, 3697, 3477, 3147, 2541, 1632, 2001, 3375, 2532, 2977, 1983, 2485, 1142, 3075, 3583, 2034, 2923, 3640, 3780, 3242, 1652, 3059, 3328, 2136, 2816, 2519]. victims [2202]. Video [2608, 3025]. View [2653, 2294, 1698, 1948, 2338, 1897, 2730, 695]. View-based [2653]. viewed [2761]. views [2280, 2641, 857, 2592, 2294, 1840, 3282, 2720]. Virtual [1684, 1671, 2313, 2403, 3411, 2883, 3265, 3152, 3055, 2868]. virtualization [3015]. virtualized [2803, 3006]. viruses [2202]. visibility [2372]. Visibly [3424, 2778]. visit [646]. visits [3692]. visual [3147]. visualisation [2818]. Visualizing [2484]. VLSI [741, 846, 873, 588, 766]. voltage [2826]. voltage-frequency [2826]. Volterra [3740, 88]. Volume [1272, 1314, 1358, 1404, 1444, 1482, 1508, 1537, 1579, 1606, 1641, 1665, 1700, 1740, 1782, 1828, 1818, 1888, 1951, 1874, 1937, 2007, 2053, 1998, 2042, 2100, 2090, 2231, 2220, 2246, 2255, 1632, 333, 360, 386, 407, 429, 455, 481, 507, 529, 552, 575, 603, 627, 651, 671, 696, 723, 748, 775, 800, 820, 843, 869, 889, 907, 931, 954, 979, 1000, 1020, 1038, 1057, 1080, 1106, 1128, 1145, 1167, 1193, 1233, 3206, 2160, 1605]. Voronoi [3214, 1961, 1163, 2197, 1071, 2603]. Voronoi-based [2603]. Voronoi-diagrams [1071]. voters [3494]. voting [3494, 3060, 3061, 2238]. voxelized [3401]. vs [1285, 1224, 2155, 1375, 3516, 1738, 1824, 2948, 2025, 3853, 1204, 1437]. Vulnerability [3087, 3519].

W [2494, 554]. w.r.t [3531]. waiting [226, 655]. walking [3822]. Walks [1682, 2046, 2119, 3029, 3192, 1202, 3400]. warehousing [2814]. wars [2940]. wavelet [1930]. Way [1578, 1398, 270, 1436, 110, 301, 1114, 1939, 50, 541, 33, 1169, 64, 1929, 1993, 197, 289, 226, 1880, 2707, 3109, 534, 179, 1865, 3918, 3826]. WCOID [2896]. WCOID-DG [2896]. WEAC [2608]. Weak [1382, 1366, 1269, 1416, 1779, 756, 1023, 3535, 2977, 2423, 684, 534, 3077, 3653]. weak-key [3653]. weak-stack-counter [534]. weakest [418]. weakly [2132]. web [2343, 2867, 2342, 2573, 2962, 2337, 2215, 3052, 198, 2855, 2263, 2595, 3145, 2336]. websites [1886]. wedges [3935]. Wegman [823]. weight [2192]. Weighted [3799, 3727, 3921, 3205, 2264, 3358, 3468, 2382, 3472, 2696, 2908, 3609, 2021, 2911, 3245, 3499, 2728, 3892, 3527, 3891, 3491, 3330, 170, 2692]. Weighting [2896]. weights [3806, 2190, 3751, 3508, 2279, 295, 1136]. Weisfeiler [3566]. welfare [3678]. well [3524, 2060]. well-formed [2060]. well-quasi-ordering [3524]. where [2406, 57]. whether [3513, 3261]. Which [1730, 718, 625, 759, 37, 492, 744, 922]. while [271, 2930]. White [914]. Whitney [3224]. who [3650]. whose [3943, 444]. Wi [3147]. Wi-Fi [3147]. wide [2507, 3859, 2628]. Width [1254, 3219, 2352, 2999, 962, 3823, 3929, 3477, 3524, 2237, 3521, 2583, 1885, 3737]. WiMax [3240]. Win [3209]. Win-win [3209]. window [3514, 3642, 2706]. windows [2667]. winner [3806, 2548]. winners [3768]. winning [2978]. Wire [2058, 613]. wire-routing [613]. wirelength [3884]. Wireless [3112, 2937, 3811, 3113, 2578, 2576, 3119, 3117, 2817, 3008, 2623, 2604, 3934, 2938, 2315, 3736, 2767, 2769, 2768, 2864, 2929, 2939, 2770, 3025, 2931].

Wise [1637, 3256, 2808]. within [1031, 1590, 2922, 269, 2563, 2353, 2298, 2014, 698]. without [1962, 1458, 2472, 989, 3503, 227, 3676, 2688, 746, 1155, 3585, 2690, 1676, 3187, 2906, 3254, 3023]. Witness [1257, 1824]. witnesses [1210]. witnessing [1929]. Witold [860]. WMN [3119]. WMN-GA [3119]. WoLLIC [2915, 2504, 3301]. Wong [3107]. Won’t [1251]. Word [3673, 3167, 2888, 2777, 3774, 3101]. Words [1722, 1476, 3790, 2685, 2798, 550, 2946, 2440, 3258, 3102, 2787, 2071, 420, 3484]. work [2992]. Workflow [2525, 2518, 2514, 2517, 2060, 2522, 2741]. workflow-based [2518]. workflows [2520, 2521, 2516, 2515, 2519]. Workload [2251, 3006]. Workshop [2525, 3301]. workspace [3041]. world [3698, 1180, 3497, 2597]. worlds [2596]. worm [2402, 2837]. Wormhole [1671, 2403]. wormholes [2938]. Worst [2475, 2286]. Worst-case [2475]. worst-order [2286]. WPP [2179]. wrapping [1217]. Write [1366, 1611, 2105]. Writing [168, 69]. WSN [3051]. WSNs [2940]. WSTS [2149].

Xiang [3187]. XML [2359, 2632, 1882, 2262, 2852, 2463, 2631, 2633, 2260, 2457, 1844, 2594, 2857, 2634, 2274, 2261, 2271, 1838, 2635, 2340, 2475, 2643, 2459, 2259, 2722, 2519]. XML-A [2259]. XML-enabled [2340]. XOR [3191, 1676]. XPath [3332, 2262, 2475, 2643, 2720].

years [3378]. yield [2433]. yields [1974, 1222]. Yugoslavia [85]. Yutsis [3905].

Z [3438, 1337, 1657]. Z-Compressed [1337]. Zero [1522, 675, 1714, 1610, 1055, 2159, 265]. Zero-inequivalence [675]. Zero-Knowledge [1610, 1055]. zeros [396]. Zheng [3187]. Zhou [3187]. Zigzag [3331]. ZPP [2235, 1793].

References

Ginsburg:1967:BCF

[1] Seymour Ginsburg and Michael A. Harrison. Bracketed context-free languages. Journal of Computer and System Sciences, 1(1):1–23, April 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                              URL http://www.sciencedirect.
                                              com/science/article/pii/S0022000067800035.

Bucy:1967:LPM

[2] R. S. Bucy. Linear positive machines. Journal of Computer and System Sciences, 1(1):24–28, April 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/S0022000067800047.

Rosen:1967:CSL

[3] J. B. Rosen. Chebyshev solution of large linear systems. Journal of Computer and System Sciences, 1(1):29–43, April 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/S0022000067800059.

Ritter:1967:PMS

[4] K. Ritter. A parametric method for solving certain nonconcave maximization problems. Journal of Computer and System Sciences, 1(1): 44–54, April 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/S0022000067800060.

Krohn:1967:MAT

[5] Kenneth Krohn, Richard Mateosian, and John Rhodes. Methods of the algebraic theory of machines: I. Decomposition theorem for generalized machines; properties preserved under series and parallel compositions of machines. Journal of Computer and System Sciences, 1(1): 55–85, April 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800072.

Blum:1967:SPF

[6] E. K. Blum. Stationary points of functionals in pre-Hilbert spaces. Journal of Computer and System Sciences, 1(1):86–90, April 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800084.

Balakrishnan:1967:FSS

[7] A. V. Balakrishnan. Foundations of the state-space theory of continuous systems. I. Journal of Computer and System Sciences, 1(1): 91–116, April 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800096.

Anonymous:1967:EB

[8] Anonymous. Editorial Board. Journal of Computer and System Sciences, 1(1):CO2, April 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000067800023.

Krohn:1967:APA

[9] Kenneth Krohn, Rudolph Langer, and John Rhodes. Algebraic principles for the analysis of a biochemical system. Journal of Computer and System Sciences, 1(2):119– 136, August 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000067800102.

Giveon:1967:SPF

[10] Yehoshafat Give’on. On some properties of the free monoids with applications to automata theory. Journal of Computer and System Sciences, 1 (2):137–154, August 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000067800114.

Hartmanis:1967:HIL

[11] Juris Hartmanis and Wayne A. Davis. Homomorphic images of linear sequential machines. Journal of Computer and System Sciences, 1(2):155– 165, August 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000067800126.

Hopcroft:1967:NSA

[12] J. E. Hopcroft and J. D. Ullman. Nonerasing stack automata. Journal of Computer and System Sciences, 1 (2):166–186, August 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800138.

Scott:1967:SDS

[13] Dana Scott. Some definitional suggestions for automata theory. Journal of Computer and System Sciences, 1(2):187–212, August 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200006780014X.

Dantzig:1967:GUB

[14] G. B. Dantzig and R. M. Van Slyke. Generalized upper bounding techniques. Journal of Computer and System Sciences, 1(3):213– 226, October 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800151.

Guinn:1967:SGO

[15] T. Guinn. Solutions of generalized optimization problems. Journal of Computer and System Sciences, 1(3): 227–234, October 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800163.

Guinn:1967:FON

[16] T. Guinn. First-order necessary conditions for generalized optimization problems. Journal of Computer and System Sciences, 1(3):235– 240, October 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800175.

Ritter:1967:DMS

[17] K. Ritter. A decomposition method for structured quadratic programming problems. Journal of Computer and System Sciences, 1(3):241– 260, October 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800187.

Axelband:1967:SOP

[18] E. I. Axelband. A solution to the optimal pursuit problem for distributed parameter systems. Journal of Computer and System Sciences, 1 (3):261–286, October 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800199.

Kleinrock:1967:DAS

[19] Leonard Kleinrock and Edward G. Coffman. Distribution of attained service in time-shared systems. Journal of Computer and System Sciences, 1 (3):287–298, October 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800205.

Rosenberg:1967:MFA

[20] Arnold L. Rosenberg. Multitape finite automata with rewind instructions. Journal of Computer and System Sciences, 1(3):299– 315, October 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800217.

Thatcher:1967:CDT

[21] J. W. Thatcher. Characterizing derivation trees of context-free grammars through a generalization of finite automata theory. Journal of Computer and System Sciences, 1(4): 317–322, December 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800229.

Beltrami:1967:IDC

[22] E. J. Beltrami. On infinite-dimensional convex programs. Journal of Computer and System Sciences, 1(4):323– 329, December 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800230.

Eggert:1967:ADS

[23] D. Eggert and P. Varaiya. Affine dynamical systems. Journal of Computer and System Sciences, 1(4):330– 348, December 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800242.

Bucy:1967:GTR

[24] R. S. Bucy. Global theory of the Riccati equation. Journal of Computer and System Sciences, 1(4):349– 361, December 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000067800254.

Blum:1967:CEE

[25] E. K. Blum. The computation of eigenvalues and eigenvector of a completely continuous self-adjoint operator. Journal of Computer and System Sciences, 1(4):362–370, December 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000067800266.

Paz:1967:FSF

[26] A. Paz. Fuzzy star functions, probabilistic automata, and their approximation by nonprobabilistic automata. Journal of Computer and System Sciences, 1(4):371–390, December 1967. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000067800278.

Balakrishnan:1967:SEI

[27] A. V. Balakrishnan and J. L. Lions. State estimation for infinitedimensional systems. Journal of Computer and System Sciences, 1(4): 391–403, December 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200006780028X.

Anonymous:1967:AI

[28] Anonymous. Author index. Journal of Computer and System Sciences, 1(4):404, December 1967. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000067800291.

Anonymous:1968:EB

[29] Anonymous. Editorial Board. Journal of Computer and System Sciences, 2(1):i, June 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800029.

Hopcroft:1968:DSA

[30] J. E. Hopcroft and J. D. Ullman. Deterministic stack automata and the quotient operator. Journal of Computer and System Sciences, 2 (1):1–12, June 1968. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800030.

Wagner:1968:BAM

[31] Eric G. Wagner. Bounded action machines: Toward an abstract theory of computer structure. Journal of Computer and System Sciences, 2(1):13–75, June 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800042.

Schumitzky:1968:EBM

[32] Alan Schumitzky. On the equivalence between matrix Riccati equations and Fredholm resolvents. Journal of Computer and System Sciences, 2(1):76–87, June 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800054.

Fischer:1968:MOW

[33] Patrick C. Fischer and Arnold L. Rosenberg. Multitape one-way nonwriting automata. Journal of Computer and System Sciences, 2(1): 88–101, June 1968. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800066.

Balakrishnan:1968:NCT

[34] A. V. Balakrishnan. A new computing technique in system identification. Journal of Computer and System Sciences, 2(1):102–116, June 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000068800078.

Hartmanis:1968:TRB

[35] J. Hartmanis. Tape-reversal bounded Turing machine computations. Journal of Computer and System Sciences, 2(2):117–135, August 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800273.

Fischer:1968:RTR

[36] Patrick C. Fischer. The reduction of tape reversals for off-line onetape Turing machines. Journal of Computer and System Sciences, 2 (2):136–147, August 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800285.

Rose:1968:CWP

[37] Gene F. Rose. Closures which preserve finiteness in families of languages. Journal of Computer and System Sciences, 2(2):148–168, August 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000068800297.

Alter:1968:NCP

[38] Ronald Alter. On the nonexistence of close-packed double Hammingerror-correcting codes on q = 7 symbols. Journal of Computer and System Sciences, 2(2):169– 176, August 1968. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800303.

Antosiewicz:1968:NMB

[39] H. A. Antosiewicz. Newton’s method and boundary value problems. Journal of Computer and System Sciences, 2(2):177–202, August 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800315.

Danchick:1968:FRG

[40] Roy Danchick. Further results on generalized predictor-corrector methods. Journal of Computer and System Sciences, 2(2):203–218, August 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000068800327.

Muller:1968:FDA

[41] David E. Muller and Gianfranco R. Putzolu. Frequency of decomposability among machines with a large number of states. Journal of Computer and System Sciences, 2(3):219– 227, October 1968. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200006880008X.

Ginsburg:1968:DBL

[42] Seymour Ginsburg and Edwin H. Spanier. Derivation-bounded languages. Journal of Computer and System Sciences, 2(3):228–250, October 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000068800091.

Kimura:1968:EAC

[43] Izumi Kimura. Extensions of asynchronous circuits and the delay problem I. Good extensions and the delay problem of the first kind. Journal of Computer and System Sciences, 2 (3):251–287, October 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800108.

Blaquiere:1968:GOS

[44] Austin Blaquière and Françoise Gérard. On the geometry of optimal strategies in two-person games of kind. Journal of Computer and System Sciences, 2 (3):288–304, October 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200006880011X.

Walkup:1968:NDR

[45] David W. Walkup and Roger J.- B. Wets. A note on decision rules for stochastic programs. Journal of Computer and System Sciences, 2(3): 305–311, October 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800121.

Putzolu:1968:PAM

[46] Gianfranco R. Putzolu. Probabilistic aspects of machine decomposition theory. Journal of Computer and System Sciences, 2(3):312– 331, October 1968. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800133.

Collatz:1968:MEP

[47] L. Collatz. Multiparametric eigenvalue problems in inner-product spaces. Journal of Computer and System Sciences, 2(4):333–341, December 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000068800339.

Demjanov:1968:ASM

[48] Vladimir F. Demjanov. Algorithms for some minimax problems. Journal of Computer and System Sciences, 2 (4):342–380, December 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800340.

Gallaire:1968:ILS

[49] H. Gallaire, J. N. Gray, M. A. Harrison, and G. T. Herman. Infinite linear sequential machines. Journal of Computer and System Sciences, 2(4): 381–419, December 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800352.

Dempster:1968:MAT

[50] J. R. H. Dempster. On the multiplying ability of two-way automata. Journal of Computer and System Sciences, 2 (4):420–426, December 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800364.

Paull:1968:SEC

[51] Marvin C. Paull and Stephen H. Unger. Structural equivalence of context-free grammars. Journal of Computer and System Sciences, 2(4): 427–463, December 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000068800376.

Allen:1968:CNS

[52] Dennis Allen. On a characterization of the nonregular set of primes. Journal of Computer and System Sciences, 2 (4):464–467, December 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000068800388.

Anonymous:1968:AI

[53] Anonymous. Author index. Journal of Computer and System Sciences, 2(4):468, December 1968. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200006880039X.

Anonymous:1969:EB

[54] Anonymous. Editorial Board. Journal of Computer and System Sciences, 3(1):i, February 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800024.

Pasquali:1969:CNS

[55] A. Pasquali. On the convergence of nonlinear simultaneous displacements. Journal of Computer and System Sciences, 3(1):1– 7, February 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800036.

Breuer:1969:CEC

[56] Melvin A. Breuer. Combinatorial equivalence of (0, 1) circulant matrices. Journal of Computer and System Sciences, 3(1):8– 23, February 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800048.

Yavin:1969:OCC

[57] Y. Yavin. Optimal control for a class of distributed parameter systems where the cost functions are norms. Journal of Computer and System Sciences, 3(1):24–36, February 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200006980005X.

Aho:1969:SDT

[58] A. V. Aho and J. D. Ullman. Syntax directed translations and the pushdown assembler. Journal of Computer and System Sciences, 3(1):37– 56, February 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000069800061.

Neustadt:1969:GTE

[59] Lucien W. Neustadt. A general theory of extremals. Journal of Computer and System Sciences, 3(1):57– 92, February 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000069800073.

Deuel:1969:TVL

[60] D. R. Deuel. Time-varying linear sequential machines. I. Journal of Computer and System Sciences, 3 (1):93–118, February 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000069800085.

Manna:1969:CP

[61] Zohar Manna. The correctness of programs. Journal of Computer and System Sciences, 3(2): 119–127, May 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800097.

Butz:1969:CHS

[62] Arthur R. Butz. Convergence with Hilbert’s space filling curve. Journal of Computer and System Sciences, 3(2):128–146, May 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800103.

Karp:1969:PPS

[63] Richard M. Karp and Raymond E. Miller. Parallel program schemata. Journal of Computer and System Sciences, 3(2):147–195, May 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000069800115.

Greibach:1969:CAO

[64] Sheila Greibach. Checking automata and one-way stack languages. Journal of Computer and System Sciences, 3(2):196–217, May 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800127.

Dennis:1969:CTC

[65] J. E. Dennis and A. A. Goldstein. Cubature and the Tchakaloff cone. Journal of Computer and System Sciences, 3(2):218–220, May 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800139.

Ito:1969:ESS

[66] Ryuichi Ito. Every semilinear set is a finite union of disjoint linear sets. Journal of Computer and System Sciences, 3(2):221–231, May 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800140.

Greibach:1969:SCG

[67] Sheila Greibach and John Hopcroft. Scattered context grammars. Journal of Computer and System Sciences, 3 (3):233–247, August 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800152.

Blum:1969:TTS

[68] E. K. Blum. Towards a theory of semantics and compilers for programming languages. Journal of Computer and System Sciences, 3(3):248– 275, August 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800164.

Mager:1969:WPA

[69] George Mager. Writing pushdown acceptors. Journal of Computer and System Sciences, 3(3):276– 318, August 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800176.

Aho:1969:PSD

[70] A. V. Aho and J. D. Ullman. Properties of syntax directed translations. Journal of Computer and System Sciences, 3(3):319–334, August 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800188.

Grigoriadis:1969:DMS

[71] M. D. Grigoriadis and K. Ritter. A decomposition method for structured linear and nonlinear programs. Journal of Computer and System Sciences, 3 (4):335–360, November 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800267.

Kaplan:1969:REE

[72] Donald M. Kaplan. Regular expressions and the equivalence of programs. Journal of Computer and System Sciences, 3(4):361–386, November 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800279.

Morris:1969:EII

[73] Alfred H. Morris. Elementary indefinite integration theory for the computer. Journal of Computer and System Sciences, 3(4):387–408, November 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800280.

Mizumoto:1969:SCF

[74] Masaharu Mizumoto, Junichi Toyoda, and Kohkichi Tanaka. Some considerations on fuzzy automata. Journal of Computer and System Sciences, 3 (4):409–422, November 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800292.

Baer:1969:RAM

[75] R. M. Baer and E. H. Spanier. Referenced automata and metaregular families. Journal of Computer and System Sciences, 3(4):423– 446, November 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800309.

Cleave:1969:PRA

[76] John Cleave. The primitive recursive analysis of ordinary differential equations and the complexity of their solutions. Journal of Computer and System Sciences, 3(4): 447–455, November 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800310.

Wegbreit:1969:GCS

[77] Ben Wegbreit. A generator of contextsensitive languages. Journal of Computer and System Sciences, 3(4):456– 461, November 1969. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800322.

Anonymous:1969:AI

[78] Anonymous. Author index. Journal of Computer and System Sciences, 3(4):462, November 1969. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000069800334.

Strong:1970:DBC

[79] H. R. Strong. Depth-bounded computation. Journal of Computer and System Sciences, 4(1):1– 14, February 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800083.

Keller:1970:NMU

[80] Herbert B. Keller. Newton’s method under mild differentiability conditions. Journal of Computer and System Sciences, 4(1):15–28, February 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800095.

Goldstein:1970:NOC

[81] A. A. Goldstein and J. S. Meditch. Nonlinear optimal controls. Journal of Computer and System Sciences, 4 (1):29–37, February 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800101.

Conti:1970:CPP

[82] R. Conti. A convex programming problem in Banach spaces and applications to optimum control theory. Journal of Computer and System Sciences, 4 (1):38–49, February 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800113.

Fischer:1970:TRS

[83] Patrick C. Fischer, Albert R. Meyer, and Arnold L. Rosenberg. Timerestricted sequence generation. Journal of Computer and System Sciences, 4(1):50–73, February 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800125.

Orgass:1970:SRC

[84] R. J. Orgass. Some results concerning proofs of statements about programs. Journal of Computer and System Sciences, 4(1):74–88, February 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800137.

Anonymous:1970:IFI

[85] Anonymous. International Federation for Information Processing: IFIP Congress 71 Ljubljana, Yugoslavia, 23– 28 August 1971: Announcement. Journal of Computer and System Sciences, 4(1):89–92, February 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800149.

Anonymous:1970:EB

[86] Anonymous. Editorial Board. Journal of Computer and System Sciences, 4(1):CO2, February 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800071.

Preparata:1970:GNO

[87] Franco P. Preparata and David E. Muller. Generation of near-optimal universal Boolean functions. Journal of Computer and System Sciences, 4(2):93–102, April 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800022.

McNabb:1970:FOI

[88] Alex McNabb and Alan Schumitzky. Factorization of operators II: a nonlinear Volterra method for numerical solution of linear Fredholm equations. Journal of Computer and System Sciences, 4(2): 103–128, April 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800034.

Aho:1970:CPP

[89] A. V. Aho, J. D. Ullman, and J. E. Hopcroft. On the computational power of pushdown automata. Journal of Computer and System Sciences, 4 (2):129–136, April 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800046.

Yamada:1970:CPP

[90] H. Yamada and S. Amoroso. A completeness problem for pattern generation in tessellation automata. Journal of Computer and System Sciences, 4(2):137–176, April 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800058.

Savitch:1970:RBN

[91] Walter J. Savitch. Relationships between nondeterministic and deterministic tape complexities. Journal of Computer and System Sciences, 4 (2):177–192, April 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007080006X.

Rose:1970:AFP

[92] Gene F. Rose. Abstract families of processors. Journal of Computer and System Sciences, 4(3): 193–204, June 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800204.

Milner:1970:EPS

[93] R. Milner. Equivalences on program schemes. Journal of Computer and System Sciences, 4(3):

205–219, June 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800216.

Luckham:1970:FCP

[94] D. C. Luckham, D. M. R. Park, and M. S. Paterson. On formalised computer programs. Journal of Computer and System Sciences, 4(3): 220–249, June 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800228.

Temam:1970:RAS

[95] R. Temam. Remarks on the approximation of some nonlinear elliptic equations. Journal of Computer and System Sciences, 4(3): 250–259, June 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007080023X.

Cohen:1970:GPS

[96] Rina S. Cohen and J. A. Brzozowski. General properties of star height of regular events. Journal of Computer and System Sciences, 4 (3):260–280, June 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800241.

Cohen:1970:SHC

[97] Rina S. Cohen. Star height of certain families of regular events. Journal of Computer and System Sciences, 4(3):281–297, June 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800253.

Liu:1970:CSS

[98] L. Y. Liu and P. Weiner. A characterization of semilinear sets. Journal of Computer and System Sciences, 4 (4):299–307, August 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800150.

Ginsburg:1970:PA

[99] Seymour Ginsburg and Sheila Greibach. Principal AFL. Journal of Computer and System Sciences, 4(4):308– 338, August 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800162.

Thatcher:1970:GSM

[100] James W. Thatcher. Generalized 2 sequential machine maps. Journal of Computer and System Sciences, 4 (4):339–367, August 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800174.

Hartmanis:1970:WMS

[101] J. Hartmanis and J. E. Hopcroft. What makes some language theory problems undecidable. Journal of Computer and System Sciences, 4 (4):368–376, August 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800186.

Moore:1970:SLR

[102] Ramon E. Moore. On the stability of linear recurrence equations with arbitrary time lags. Journal of Computer and System Sciences, 4 (4):377–383, August 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800198.

Smith:1970:EDF

[103] William Bridges Smith. Error detection in formal languages. Journal of Computer and System Sciences, 4 (5):385–405, October 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007080040X. See

comment [248].

Doner:1970:TAS

[104] John Doner. Tree acceptors and some of their applications. Journal of Computer and System Sciences, 4 (5):406–451, October 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800411.

Basu:1970:SSD

[105] Sanat K. Basu. On the structure of subrecursive degrees. Journal of Computer and System Sciences, 4(5): 452–464, October 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800423.

Miller:1970:RFT

[106] Webb Miller. Recursive function theory and numerical analysis. Journal of Computer and System Sciences, 4 (5):465–472, October 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800435.

Angel:1970:IID

[107] Edward Angel. Invariant imbedding, difference equations, and elliptic boundary value problems. Journal of Computer and System Sciences, 4 (5):473–491, October 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800447.

Kasai:1970:HBC

[108] Takumi Kasai. An hierarchy between context-free and context-sensitive languages. Journal of Computer and System Sciences, 4(5):492–508, October 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000070800459.

Perrie:1970:URA

[109] A. L. Perrie. Uniform rational approximation with osculatory interpolation. Journal of Computer and System Sciences, 4(6):509–522, December 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800265.

Aho:1970:CTW

[110] A. V. Aho and J. D. Ullman. A characterization of two-way deterministic classes of languages. Journal of Computer and System Sciences, 4(6): 523–538, December 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800277.

Giveon:1970:ASL

[111] Y. Give’on and Y. Zalcstein. Algebraic structures in linear systems theory. Journal of Computer and System Sciences, 4(6):539– 556, December 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800289.

Sakawa:1970:AET

[112] Yoshiyuki Sakawa. An application of the epsilon technique to the solution of pursuit and evasion problems. Journal of Computer and System Sciences, 4 (6):557–569, December 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000070800290.

Kapur:1970:CPA

[113] Kailash C. Kapur and Richard M. Van Slyke. Cutting plane algorithms and state space constrained linear optimal control problems. Journal of Computer and System Sciences, 4(6):

570–605, December 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800307.

Book:1970:TTB

[114] Ronald V. Book, Sheila A. Greibach, and Ben Wegbreit. Time- and tape-bounded Turing acceptors and AFLs. Journal of Computer and System Sciences, 4(6):606–621, December 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000070800319.

Book:1970:TBT

[115] Ronald V. Book, Sheila A. Greibach, Oscar H. Ibarra, and Ben Wegbreit. Tape-bounded Turing acceptors and principal AFLs. Journal of Computer and System Sciences, 4(6):622– 625, December 1970. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800320.

Anonymous:1970:AI

[116] Anonymous. Author index. Journal of Computer and System Sciences, 4(6):626, December 1970. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000070800332.

Cohen:1971:DDS

[117] Rina S. Cohen and J. A. Brzozowski. Dot-depth of star-free events. Journal of Computer and System Sciences, 5(1):1–16, February 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007180003X.

Aberth:1971:CEM

[118] Oliver Aberth. The concept of effective method applied to computational problems of linear algebra. Journal of Computer and System Sciences, 5 (1):17–25, February 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800041.

Carlyle:1971:RSF

[119] J. W. Carlyle and A. Paz. Realizations by stochastic finite automata. Journal of Computer and System Sciences, 5(1):26–40, February 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800053.

Brzozowski:1971:CNE

[120] J. A. Brzozowski, K. Culík, and A. Gabrielian. Classification of noncounting events. Journal of Computer and System Sciences, 5(1):41– 53, February 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800065.

Hartmanis:1971:ULS

[121] J. Hartmanis and F. D. Lewis. The use of lists in the study of undecidable problems in automata theory. Journal of Computer and System Sciences, 5(1):54–66, February 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800077.

Engeler:1971:AA

[122] Erwin Engeler. Algorithmic approximations. Journal of Computer and System Sciences, 5(1):67– 82, February 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800089.

Anonymous:1971:EB

[123] Anonymous. Editorial Board. Journal of Computer and System Sciences, 5(1):CO2, February 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800028.

Goldstein:1971:IRN

[124] A. A. Goldstein, F. C. Johnson, and Jurgen Spiess. On the isolation of the roots of a non-linear operator. Journal of Computer and System Sciences, 5(2):83–87, April 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800284.

Ibarra:1971:CST

[125] Oscar H. Ibarra. Characterizations of some tape and time complexity classes of Turing machines in terms of multihead and auxiliary stack automata. Journal of Computer and System Sciences, 5 (2):88–117, April 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800296.

Ausiello:1971:ACC

[126] Giorgio Ausiello. Abstract computational complexity and cycling computations. Journal of Computer and System Sciences, 5(2): 118–128, April 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800302.

Kimura:1971:EAC

[127] Izumi Kimura. Extensions of asynchronous circuits and the delay problem. Part II: Spike-free extensions and the delay problem of the second kind. Journal of Computer and System Sciences, 5(2): 129–162, April 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800314.

Balakrishnan:1971:CAM

[128] A. V. Balakrishnan. A computational approach to the maximum principle. Journal of Computer and System Sciences, 5(2): 163–191, April 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800326.

Rosenberg:1971:DGA

[129] Arnold L. Rosenberg. Data graphs and addressing schemes. Journal of Computer and System Sciences, 5 (3):193–238, June 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007180034X.

Manna:1971:MTP

[130] Zohar Manna. Mathematical theory of partial correctness. Journal of Computer and System Sciences, 5 (3):239–253, June 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800351.

Strong:1971:TRE

[131] H. R. Strong. Translating recursion equations into flow charts. Journal of Computer and System Sciences, 5 (3):254–285, June 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800363.

Lewis:1971:EIC

[132] F. D. Lewis. The enumerability and invariance of complexity classes. Journal of Computer and System Sciences, 5(3):286–303, June 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800375.

Ullian:1971:TTC

[133] J. S. Ullian. Three theorems concerning principal AFLs. Journal of Computer and System Sciences, 5 (3):304–314, June 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800387.

Constable:1971:SPL

[134] Robert L. Constable. Subrecursive programming languages II on program size. Journal of Computer and System Sciences, 5(3): 315–334, June 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800399.

Anonymous:1971:A

[135] Anonymous. Announcement. Journal of Computer and System Sciences, 5(3):335, June 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800405.

Karp:1971:SSI

[136] Richard M. Karp. Special SIGACT issue. Journal of Computer and System Sciences, 5(3):i+(1*), June 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000071800338.

Revesz:1971:UCS

[137] György Révész. Unilateral context sensitive grammars and leftto-right parsing. Journal of Computer and System Sciences, 5(4):337– 352, August 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800223.

Gruska:1971:CCF

[138] Jozef Gruska. A characterization of context-free languages. Journal of Computer and System Sciences, 5 (4):353–364, August 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800235.

Ginsburg:1971:ASP

[139] Seymour Ginsburg and Edwin H. Spanier. AFL with the semilinear property. Journal of Computer and System Sciences, 5(4):365– 396, August 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800247.

Book:1971:TBG

[140] Ronald V. Book. Time-bounded grammars and their languages. Journal of Computer and System Sciences, 5 (4):397–429, August 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800259.

Rosen:1971:MEB

[141] J. B. Rosen. Minimum error bounds for multidimensional spline approximation. Journal of Computer and System Sciences, 5(4):430– 452, August 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800260.

Anonymous:1971:E

[142] Anonymous. Erratum. Journal of Computer and System Sciences, 5 (4):453, August 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800272.

Amoroso:1971:TSR

[143] Serafino Amoroso and Gerald Cooper. Tessellation structures for reproduction of arbitrary patterns. Journal of Computer and System Sciences, 5 (5):455–464, October 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800090.

Grzymala-Busse:1971:OPF

[144] Jerzy W. Grzymala-Busse. Operationpreserving functions and autonomous factors of finite automata. Journal of Computer and System Sciences, 5 (5):465–474, October 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800107.

Fleck:1971:TTD

[145] A. C. Fleck. Towards a theory of data structures. Journal of Computer and System Sciences, 5(5):475– 488, October 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800119.

Herman:1971:EFS

[146] G. T. Herman. Every finite sequential machine is linearly realizable. Journal of Computer and System Sciences, 5 (5):489–510, October 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800120.

Commoner:1971:MDG

[147] F. Commoner, A. W. Holt, S. Even, and A. Pnueli. Marked directed graphs. Journal of Computer and System Sciences, 5(5):511–523, October 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000071800132.

Nicholson:1971:ATF

[148] Peter J. Nicholson. Algebraic theory of finite Fourier transforms. Journal of Computer and System Sciences, 5 (5):524–547, October 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800144.

Siromoney:1971:FTC

[149] Rani Siromoney. Finite-turn checking automata. Journal of Computer and System Sciences, 5(6):549– 559, December 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800156.

Pratt:1971:PGG

[150] Terrence W. Pratt. Pair grammars, graph languages and stringto-graph translations. Journal of Computer and System Sciences, 5(6): 560–595, December 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800168.

Culik:1971:AFR

[151] Karel Culík. Abstract families of relations. Journal of Computer and System Sciences, 5(6):596– 606, December 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007180017X.

Fenner:1971:FIM

[152] Trevor I. Fenner and Georghios Loizou. On fully indecomposable matrices. Journal of Computer and System Sciences, 5(6):607– 622, December 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800181.

Ostrand:1971:PRT

[153] Thomas J. Ostrand. Pattern reproduction in tessellation automata of arbitrary dimension. Journal of Computer and System Sciences, 5(6): 623–628, December 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000071800193.

McWhirter:1971:SE

[154] I. P. McWhirter. Substitution expressions. Journal of Computer and System Sciences, 5(6):629–637, December 1971. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007180020X.

Anonymous:1971:AI

[155] Anonymous. Author index. Journal of Computer and System Sciences, 5(6):638, December 1971. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000071800211.

Heindel:1972:CPM

[156] Lee E. Heindel. Computation of powers of multivariate polynomials over the integers. Journal of Computer and System Sciences, 6(1):1– 8, February 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800370.

Goldstine:1972:SBL

[157] Jonathan Goldstine. Substitution and bounded languages. Journal of Computer and System Sciences, 6 (1):9–29, February 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800382.

Greibach:1972:SOF

[158] Sheila A. Greibach. Syntactic operators on full semiAFLs. Journal of Computer and System Sciences, 6 (1):30–76, February 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800394.

Kang:1972:EON

[159] Andy N. C. Kang. Effective operators with no strong gaps *. Journal of Computer and System Sciences, 6 (1):77–87, February 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800400.

Itkin:1972:PSE

[160] V. E. Itkin, V. E. Itkin, and Z. Zwinogrodzki. On program schemata equivalence. Journal of Computer and System Sciences, 6 (1):88–101, February 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800412.

Anonymous:1972:EB

[161] Anonymous. Editorial Board. Journal of Computer and System Sciences, 6(1):CO2, February 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800369.

Carlyle:1972:EAS

[162] Jack W. Carlyle, Sheila Greibach, Juris Hartmanis, and E. K. Blum. Eleventh Annual Switching and Automata Theory Symposium. Journal of Computer and System Sciences, 6(2):i, April 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800151.

Gale:1972:PTS

[163] David Gale and Richard M. Karp. A phenomenon in the theory of sorting. Journal of Computer and System Sciences, 6(2):103–115, April 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000072800163.

Paterson:1972:TBT

[164] Michael S. Paterson. Tape bounds for time-bounded Turing machines. Journal of Computer and System Sciences, 6(2):116–124, April 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800175.

Aho:1972:EPS

[165] A. V. Aho and J. D. Ullman. Equivalence of programs with structured variables. Journal of Computer and System Sciences, 6(2): 125–137, April 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800187.

Kameda:1972:PAC

[166] T. Kameda. Pushdown automata with counters. Journal of Computer and System Sciences, 6(2): 138–150, April 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800199.

Zalcstein:1972:LTL

[167] Yechezkel Zalcstein. Locally testable languages. Journal of Computer and System Sciences, 6(2):151–167, April 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000072800205.

Giuliano:1972:WSA

[168] Joseph Alphonso Giuliano. Writing stack acceptors. Journal of Computer and System Sciences, 6 (2):168–204, April 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800217.

Spira:1972:CLP

[169] P. M. Spira. Complete linear proofs of systems of linear inequalities. Journal of Computer and System Sciences, 6(3):205–216, June 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800023.

Stanat:1972:HTW

[170] Donald F. Stanat. A homomorphism theorem for weighted contextfree grammars. Journal of Computer and System Sciences, 6(3): 217–232, June 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800035.

Smith:1972:RTL

[171] Alvy Ray Smith. Real-time language recognition by one-dimensional cellular automata. Journal of Computer and System Sciences, 6(3): 233–253, June 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800047.

Cleave:1972:CSC

[172] J. P. Cleave. Combinatorial systems I. Cylindrical problems. Journal of Computer and System Sciences, 6(3):254–266, June 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800059.

Eickel:1972:RBD

[173] Jürgen Eickel and Jacques Loeckx. The relation between derivations and syntactical structures in phrasestructure grammars. Journal of Computer and System Sciences, 6 (3):267–282, June 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800060.

Butzbach:1972:CDM

[174] Philippe Butzbach. Les 1-convexes du monoïde commutatif N^{n}. (French) [The 1-convexes of the commutative monoid N^{n}]. Journal of Computer and System Sciences, 6(3): 283–289, June 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800072.

Anonymous:1972:N

[175] Anonymous. Notice. Journal of Computer and System Sciences, 6 (3):290, June 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800084.

Golberg:1972:ESI

[176] Michael A. Golberg. The equivalence of several initial value methods for solving integral equations. Journal of Computer and System Sciences, 6 (4):291–297, August 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800229.

Schultz:1972:DTA

[177] Martin H. Schultz. Discrete Tchebycheff approximation for multivariate splines. Journal of Computer and System Sciences, 6(4):298– 304, August 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800230.

King:1972:IOT

[178] James C. King and Robert W. Floyd. An interpretation-oriented theorem prover over integers. Journal of Computer and System Sciences, 6 (4):305–323, August 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800242.

Rajlich:1972:APG

[179] Vaclav Rajlich. Absolutely parallel grammars and two-way finitestate transducers. Journal of Computer and System Sciences, 6(4):324– 342, August 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800254.

Bucy:1972:REB

[180] R. S. Bucy. The Riccati equation and its bounds. Journal of Computer and System Sciences, 6 (4):343–353, August 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800266.

McCormick:1972:GAO

[181] Steve F. McCormick. A general approach to one-step iterative methods with application to eigenvalue problems. Journal of Computer and System Sciences, 6(4):354– 372, August 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800278.

Richardson:1972:TLT

[182] D. Richardson. Tessellations with local transformations. Journal of Computer and System Sciences, 6(5): 373–388, October 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800096.

Enderton:1972:DCC

[183] H. B. Enderton. Degrees of computational complexity. Journal of Computer and System Sciences, 6(5): 389–396, October 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800102.

Preparata:1972:CVL

[184] F. P. Preparata and R. T. Yeh. Continuously valued logic. Journal of Computer and System Sciences, 6(5): 397–418, October 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800114.

Muller:1972:GTI

[185] D. E. Muller and R. E. Miller. A generalization of the theory of incompletely specified machines. Journal of Computer and System Sciences, 6 (5):419–447, October 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800126.

Amoroso:1972:DPS

[186] S. Amoroso and Y. N. Patt. Decision procedures for surjectivity and injectivity of parallel maps for tessellation structures. Journal of Computer and System Sciences, 6(5):448– 464, October 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800138.

Katzenelson:1972:MAL

[187] Jacob Katzenelson. The Markov algorithm as a language parserlinear bounds. Journal of Computer and System Sciences, 6(5):465– 478, October 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007280014X.

Fischer:1972:SSI

[188] Michael J. Fischer. Special SIGACT issue. Journal of Computer and System Sciences, 6(6):479, December 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S002200007280028X.

Constable:1972:SPS

[189] R. L. Constable and S. S. Muchnick. Subrecursive program schemata I and II: I. Undecidable equivalence problems; II. Decidable equivalence problems. Journal of Computer and System Sciences, 6(6):480– 537, December 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800291.

Wise:1972:GOR

[190] David S. Wise. Generalized overlap resolvable grammars and their parsers. Journal of Computer and System Sciences, 6(6):538–572, December 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800308.

Aho:1972:OLP

[191] A. V. Aho and J. D. Ullman. Optimization of LR(k) parsers. Journal of Computer and System Sciences, 6 (6):573–602, December 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007280031X.

Machtey:1972:ALL

[192] Michael Machtey. Augmented loop languages and classes of computable functions. Journal of Computer and System Sciences, 6(6):603–624, December 1972. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800321.

Munro:1972:EEP

[193] Ian Munro and Allan Borodin. Efficient evaluation of polynomial forms. Journal of Computer and System Sciences, 6(6):625–638, December 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000072800333.

Rabin:1972:PSP

[194] Michael O. Rabin. Proving simultaneous positivity of linear forms. Journal of Computer and System Sciences, 6 (6):639–650, December 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800345.

Anonymous:1972:AI

[195] Anonymous. Author index. Journal of Computer and System Sciences, 6(6):651, December 1972. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000072800357.

Ginsburg:1973:AGF

[196] Seymour Ginsburg and Sheila Greibach. On AFL generators for finitely encoded AFA. Journal of Computer and System Sciences, 7(1):1– 27, February 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800479.

Ibarra:1973:TWM

[197] Oscar H. Ibarra. On two-way multihead automata. Journal of Computer and System Sciences, 7(1):28– 36, February 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800480.

Abe:1973:WGS

[198] Norihiro Abe, Masaharu Mizumoto, Jun ichi Toyoda, and Kohkichi Tanaka. Web grammars and several graphs. Journal of Computer and System Sciences, 7(1):37– 65, February 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800492.

Culik:1973:LRG

[199] Karel Culik and Rina Cohen. LRregular grammars — an extension of LR(k) grammars. Journal of Computer and System Sciences, 7(1):66– 96, February 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800509.

Feliciangeli:1973:APG

[200] Horacio Feliciangeli and Gabor T. Herman. Algorithms for producing grammars from sample derivations: a common problem of formal language theory and developmental biology. Journal of Computer and System Sciences, 7(1):97–118, February 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800510.

Anonymous:1973:EB

[201] Anonymous. Editorial Board. Journal of Computer and System Sciences, 7(1):CO2, February 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800467.

Garland:1973:PSR

[202] Stephen J. Garland and David C. Luckham. Program schemes, recursion schemes, and formal languages. Journal of Computer and System Sciences, 7(2):119–160, April 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800406.

Zeigler:1973:EDI

[203] Bernard P. Zeigler. Every discrete input machine is linearly simulatable. Journal of Computer and System Sciences, 7(2): 161–167, April 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800418.

Hartmanis:1973:GTC

[204] J. Hartmanis and H. Walter. Group theoretic characterization of linear permutation automata. Journal of Computer and System Sciences, 7 (2):168–188, April 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007380042X.

Munro:1973:OAP

[205] Ian Munro and Michael Paterson. Optimal algorithms for parallel polynomial evaluation. Journal of Computer and System Sciences, 7(2): 189–198, April 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800431.

Igarashi:1973:EGT

[206] Yoshihide Igarashi and Namio Honda. On the extension of Gladkij’s Theorem and the hierarchies of languages. Journal of Computer and System Sciences, 7(2):199–217, April 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800443.

Bird:1973:EPD

[207] Malcolm Bird. The equivalence problem for deterministic two-tape automata. Journal of Computer and System Sciences, 7(2):218–236, April 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000073800455.

Harrison:1973:SDG

[208] Michael A. Harrison and Ivan M. Havel. Strict deterministic grammars. Journal of Computer and System Sciences, 7(3):237–277, June 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007380008X.

Arbib:1973:CDM

[209] Michael A. Arbib. Coproducts and decomposable machines. Journal of Computer and System Sciences, 7 (3):278–287, June 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800091.

Kohavi:1973:ECB

[210] Zvi Kohavi and Joseph Winograd. Establishing certain bounds concerning finite automata. Journal of Computer and System Sciences, 7 (3):288–299, June 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800108.

Masunaga:1973:STA

[211] Yoshifumi Masunaga, Shoichi Noguchi, and Juro Oizumi. A structure theory of automata characterized by groups. Journal of Computer and System Sciences, 7(3): 300–305, June 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007380011X.

Cheung:1973:ASE

[212] To yat Cheung. Approximate solutions and error bounds for quasilinear elliptic boundary value problems. Journal of Computer and System Sciences, 7(3):306–322, June 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800121.

Hopcroft:1973:AIT

[213] J. E. Hopcroft and R. E. Tarjan. A V log V algorithm for isomorphism of triconnected planar graphs. Journal of Computer and System Sciences, 7(3):323–331, June 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800133.

Rosenberg:1973:SSI

[214] Arnold L. Rosenberg, Edward K. Blum, and Patrick C. Fischer. Special SIGACT issue: Fourth Annual ACM Symposium on Theory of Computing.

Journal of Computer and System Sciences, 7(4):333, August 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800261.

Kung:1973:BME

[215] H. T. Kung. A bound on the multiplicative efficiency of iteration. Journal of Computer and System Sciences, 7(4):334–342, August 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800273.

Cook:1973:HNT

[216] Stephen A. Cook. A hierarchy for nondeterministic time complexity. Journal of Computer and System Sciences, 7(4):343–353, August 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800285.

Cook:1973:TBR

[217] Stephen A. Cook and Robert A. Reckhow. Time bounded random access machines. Journal of Computer and System Sciences, 7(4):354– 375, August 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800297.

Schnorr:1973:PCE

[218] C. P. Schnorr. Process complexity and effective random tests. Journal of Computer and System Sciences, 7 (4):376–388, August 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800303.

Savitch:1973:MRA

[219] Walter J. Savitch. Maze recognizing automata and nondeterministic tape complexity. Journal of Computer and System Sciences, 7(4):389– 403, August 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800315.

Walker:1973:CFR

[220] S. A. Walker and H. R. Strong. Characterizations of flowchartable recursions. Journal of Computer and System Sciences, 7(4):404–447, August 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000073800327.

Blum:1973:TBS

[221] Manuel Blum, Robert W. Floyd, Vaughan Pratt, Ronald L. Rivest, and Robert E. Tarjan. Time bounds for selection. Journal of Computer and System Sciences, 7(4):448– 461, August 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800339.

Blattner:1973:UEP

[222] Meera Blattner. The unsolvability of the equality problem for sentential forms of context-free grammars. Journal of Computer and System Sciences, 7(5):463–468, October 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800029.

Horowitz:1973:ECP

[223] Ellis Horowitz. The efficient calculation of powers of polynomials. Journal of Computer and System Sciences, 7 (5):469–480, October 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800030.

Yaku:1973:CCC

[224] Takeo Yaku. The constructibility of a configuration in a cellular automaton. Journal of Computer and System Sciences, 7(5):481– 496, October 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800042.

Hughes:1973:MOD

[225] C. E. Hughes. Many-one degrees associated with semi-Thue systems. Journal of Computer and System Sciences, 7(5):497–505, October 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800054.

Konheim:1973:DQL

[226] Alan G. Konheim and Bernd Meister. Distributions of queue lengths and waiting times in a loop with two-way traffic. Journal of Computer and System Sciences, 7(5):506– 521, October 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800066.

Kartashev:1973:SAR

[227] Svetlana P. Kartashev. State assignment for realizing modular inputfree sequential logical networks without invertors. Journal of Computer and System Sciences, 7(5):522– 542, October 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800078.

Rutledge:1973:PSA

[228] J. D. Rutledge. Program schemata as automata. I. Journal of Computer and System Sciences, 7(6):543– 578, December 1973. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800340.

Rosebrugh:1973:CTP

[229] Robert D. Rosebrugh and Derick Wood. A characterization theorem for n-parallel right linear languages. Journal of Computer and System Sciences, 7(6):579–582, December 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000073800352.

Boasson:1973:TIT

[230] L. Boasson. Two iteration theorems for some families of languages. Journal of Computer and System Sciences, 7(6):583–596, December 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800364.

Nudds:1973:CPB

[231] D. Nudds. Conversion of programs between computers: Interpreters, simulators and compilers. Journal of Computer and System Sciences, 7(6): 597–614, December 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800376.

Nozaki:1973:TEG

[232] Akihiro Nozaki. Two entropies of a generalized sorting problem. Journal of Computer and System Sciences, 7 (6):615–621, December 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000073800388.

Anonymous:1973:AI

[233] Anonymous. Author index. Journal of Computer and System Sciences, 7(6):622, December 1973. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007380039X.

Walljasper:1974:LDB

[234] S. J. Walljasper. Left-derivation bounded languages. Journal of Computer and System Sciences, 8(1):1– 7, February 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800188.

Hart:1974:ADP

[235] Johnson Murdoch Hart. Ambiguity and decision problems for local adjunct languages. Journal of Computer and System Sciences, 8 (1):8–21, February 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007480019X.

Bird:1974:TP

[236] Richard Bird. On transformations of programs. Journal of Computer and System Sciences, 8(1):22– 35, February 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800206.

Crestin:1974:SGA

[237] J. P. Crestin. Structure des grammaires d’ambiguïté bornée. (French) [Structure of bounded ambiguous grammars]. Journal of Computer and System Sciences, 8(1):36–40, February 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800218.

Guha:1974:PSM

[238] Ratan K. Guha and Raymond T. Yeh. On periodicity of sequential machines. Journal of Computer and System Sciences, 8(1):41– 70, February 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007480022X.

Lewis:1974:PSF

[239] Harry R. Lewis. Program schemata and the first-order decision problem. Journal of Computer and System Sciences, 8(1):71–83, February 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800231.

Ibaraki:1974:CDO

[240] Toshihide Ibaraki. Classes of discrete optimization problems and their decision problems. Journal of Computer and System Sciences, 8(1):84– 116, February 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800243.

Anonymous:1974:EBa

[241] Anonymous. Editorial Board. Journal of Computer and System Sciences, 8(1):CO2, February 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800176.

Choueka:1974:TAT

[242] Yaacov Choueka. Theories of automata on ω-tapes: a simplified approach. Journal of Computer and System Sciences, 8(2): 117–141, April 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800516.

Khabbaz:1974:GHL

[243] Nabil A. Khabbaz. A geometric hierarchy of languages. Journal of Computer and System Sciences, 8 (2):142–157, April 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800528.

Cremers:1974:VL

[244] A. B. Cremers and O. Mayer. On vector languages. Journal of Computer and System Sciences, 8(2): 158–166, April 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007480053X.

Igarashi:1974:DMA

[245] Yoshihide Igarashi and Namio Honda. Deterministic multitape automata computations. Journal of Computer and System Sciences, 8(2): 167–189, April 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800541.

Fischer:1974:PM

[246] Patrick C. Fischer and John C. Warkentin. Predecessor machines. Journal of Computer and System Sciences, 8(2):190–219, April 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000074800553.

Blum:1974:SEP

[247] E. K. Blum and G. H. Rodrigue. Solution of eigenvalue problems in Hilbert spaces by a gradient method. Journal of Computer and System Sciences, 8(2):220–237, April 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800565.

Revesz:1974:CPE

[248] György Révész. Comment on the paper “Error detection in formal languages”. Journal of Computer and System Sciences, 8(2): 238–242, April 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800577. See

[103].

Ginsburg:1974:ESC

[249] Seymour Ginsburg and Gene F. Rose. The equivalence of stackcounter acceptors and quasi-realtime stack-counter acceptors. Journal of Computer and System Sciences, 8 (2):243–269, April 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800589.

Strong:1974:TAI

[250] H. Raymond Strong, Edward K. Blum, Richard M. Karp, and Arnold L. Rosenberg. Thirteenth annual IEEE Symposium on Switching and Automata Theory. Journal of Computer and System Sciences, 8(3): 271, June 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800255.

Johnson:1974:FAP

[251] David S. Johnson. Fast algorithms for bin packing. Journal of Computer and System Sciences, 8(3): 272–314, June 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800267.

Baker:1974:RBM

[252] Brenda S. Baker and Ronald V. Book. Reversal-bounded multipushdown machines. Journal of Computer and System Sciences, 8(3): 315–332, June 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800279.

Goguen:1974:HCT

[253] J. A. Goguen. On homomorphisms, correctness, termination, unfoldments, equivalence of flow diagram programs. Journal of Computer and System Sciences, 8(3): 333–365, June 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800280.

Borodin:1974:FMT

[254] A. Borodin and R. Moenck. Fast modular transforms. Journal of Computer and System Sciences, 8 (3):366–386, June 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800292.

Kfoury:1974:TSR

[255] A. J. Kfoury. Translatability of schemas over restricted interpretations. Journal of Computer and System Sciences, 8(3):387–408, June 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000074800309.

Maibaum:1974:GAF

[256] T. S. E. Maibaum. A generalized approach to formal languages. Journal of Computer and System Sciences, 8(3):409–439, June 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800310.

Anonymous:1974:AIa

[257] Anonymous. Author index. Journal of Computer and System Sciences, 8(3):440, June 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800322.

Gill:1974:MEF

[258] Arthur Gill and Lawrence T. Kou. Multiple-entry finite automata. Journal of Computer and System Sciences, 9(1):1–19, August 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800346.

Meyer:1974:AMP

[259] Gerad G. L. Meyer. Algorithm model for penalty functions-type iterative procedures. Journal of Computer and System Sciences, 9(1):20– 30, August 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800358.

Maruoka:1974:CPO

[260] Akira Maruoka and Masayuki Kimura. Completeness problem of one-dimensional binary scope-3 tessellation automata. Journal of Computer and System Sciences, 9(1):31–47, August 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200007480036X.

Norris:1974:DIS

[261] Eugene M. Norris. Difunctionally induced state machines. Journal of Computer and System Sciences, 9 (1):48–55, August 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800371.

Hwang:1974:CDF

[262] Kai Hwang. Cyclic decomposition of finite stochastic systems. Journal of Computer and System Sciences, 9(1):56–68, August 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800383.

Robertson:1974:CCP

[263] Edward L. Robertson. Complexity classes of partial recursive functions. Journal of Computer and System Sciences, 9(1):69– 87, August 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800395.

Ginsburg:1974:IAF

[264] Seymour Ginsburg and Edwin H. Spanier. On incomparable abstract family of languages (AFL). Journal of Computer and System Sciences, 9 (1):88–108, August 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800401.

Heindel:1974:CAA

[265] Lee E. Heindel. Congruence arithmetic algorithms for polynomial real zero determination. Journal of Computer and System Sciences, 9(1):109– 128, August 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800413.

Anonymous:1974:EBb

[266] Anonymous. Editorial Board. Journal of Computer and System Sciences, 9(1):CO2, August 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800334.

Rouchaleau:1974:LDS

[267] Yves Rouchaleau and Bostwick F. Wyman. Linear dynamical systems over integral domains. Journal of Computer and System Sciences, 9(2): 129–142, October 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800024.

Lynch:1974:AHP

[268] Nancy Lynch. Approximations to the halting problem. Journal of Computer and System Sciences, 9(2): 143–150, October 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800036.

Daley:1974:EDS

[269] R. P. Daley. The extent and density of sequences within the minimalprogram complexity hierarchies. Journal of Computer and System Sciences, 9(2):151–163, October 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800048.

Kosaraju:1974:WSA

[270] S. Rao Kosaraju. 1-way stack automaton with jumps. Journal of Computer and System Sciences, 9(2): 164–176, October 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007480005X.

Kasai:1974:TFW

[271] Takumi Kasai. Translatability of flowcharts into while programs. Journal of Computer and System Sciences, 9(2):177–195, October 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800061.

Elias:1974:MTM

[272] Peter Elias. Minimum times and memories needed to compute the values of a function. Journal of Computer and System Sciences, 9(2):196– 212, October 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800073.

Book:1974:CCC

[273] Ronald V. Book. Comparing complexity classes. Journal of Computer and System Sciences, 9(2):213– 229, October 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800085.

Aho:1974:SSI

[274] Alfred V. Aho, Edward K. Blum, and P. C. Fischer. Special SIGACT issue: Fifth Annual ACM Symposium on Theory of Computing. Journal of Computer and System Sciences, 9 (3):231, December 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800425.

Kosaraju:1974:ASP

[275] S. Rao Kosaraju. Analysis of structured programs. Journal of Computer and System Sciences, 9(3):232– 255, December 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800437.

Johnson:1974:AAC

[276] David S. Johnson. Approximation algorithms for combinatorial problems. Journal of Computer and System Sciences, 9(3):256–278, December 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800449.

Lewis:1974:AT

[277] P. M. Lewis, D. J. Rosenkrantz, and R. E. Stearns. Attributed translations. Journal of Computer and System Sciences, 9(3):279– 307, December 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000074800450.

Cook:1974:OTS

[278] Stephen A. Cook. An observation on time-storage trade off. Journal of Computer and System Sciences, 9(3): 308–316, December 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800462.

Fischer:1974:FLI

[279] Michael J. Fischer and Larry J. Stockmeyer. Fast on-line integer multiplication. Journal of Computer and System Sciences, 9(3):317– 331, December 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800474.

Vuillemin:1974:COI

[280] Jean Vuillemin. Correct and optimal implementations of recursion in a simple programming language. Journal of Computer and System Sciences, 9 (3):332–354, December 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800486.

Tarjan:1974:TFG

[281] R. Endre Tarjan. Testing flow graph reducibility. Journal of Computer and System Sciences, 9(3):355– 365, December 1974. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800498.

Anonymous:1974:AIb

[282] Anonymous. Author index. Journal of Computer and System Sciences, 9(3):366, December 1974. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000074800504.

Anonymous:1975:EBa

[283] Anonymous. Editorial Board. Journal of Computer and System Sciences, 10(1):i, February 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800110.

Goguen:1975:DTM

[284] J. A. Goguen. Discrete-time machines in closed monoidal categories. I. Journal of Computer and System Sciences, 10(1):1–43, February 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800122.

Harnik:1975:EPP

[285] Victor Harnik. Effective proper procedures and universal classes of program schemata. Journal of Computer and System Sciences, 10(1):44– 61, February 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800134.

Sudborough:1975:TBC

[286] I. H. Sudborough. On tape-bounded complexity classes and multihead finite automata. Journal of Computer and System Sciences, 10(1):62– 76, February 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800146.

Amoroso:1975:SCC

[287] Serafino Amoroso, Gerald Cooper, and Yale Patt. Some clarifications of the concept of a Gardenof-Eden configuration. Journal of Computer and System Sciences, 10 (1):77–82, February 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800158.

Owings:1975:SCS

[288] James C. Owings. Splitting a contextsensitive set. Journal of Computer and System Sciences, 10(1):83– 87, February 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007580016X.

Kiel:1975:TWT

[289] Donald I. Kiel. Two-way atransducers and AFL. Journal of Computer and System Sciences, 10 (1):88–109, February 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800171.

Ashcroft:1975:PAA

[290] E. A. Ashcroft. Proving assertions about parallel programs. Journal of Computer and System Sciences, 10 (1):110–135, February 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800183.

Joshi:1975:TAG

[291] Aravind K. Joshi, Leon S. Levy, and Masako Takahashi. Tree adjunct grammars. Journal of Computer and System Sciences, 10(1): 136–163, February 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800195.

Ginsburg:1975:UEA

[292] Seymour Ginsburg, Jonathan Goldstine, and Sheila Greibach. Uniformly erasable AFL. Journal of Computer and System Sciences, 10 (2):165–182, April 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800389.

Machtey:1975:DHS

[293] Michael Machtey. On the density of honest subrecursive classes. Journal of Computer and System Sciences, 10(2):183–199, April 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800390.

Salomon:1975:DMP

[294] Kenneth B. Salomon. The decidability of a mapping problem for generalized sequential machines with final states. Journal of Computer and System Sciences, 10(2): 200–218, April 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800407.

Mizumoto:1975:VKA

[295] M. Mizumoto, J. Toyoda, and K. Tanaka. Various kinds of automata with weights. Journal of Computer and System Sciences, 10 (2):219–236, April 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800419.

Greibach:1975:OCL

[296] S. A. Greibach. One counter languages and the IRS condition. Journal of Computer and System Sciences, 10(2):237–247, April 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800420.

Hamilton:1975:RTS

[297] Wallace L. Hamilton and John R. Mertens. Reproduction in tessellation structures. Journal of Computer and System Sciences, 10(2): 248–252, April 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800432.

Nakamura:1975:CS

[298] Akira Nakamura. On causal ω^{2}- systems. Journal of Computer and System Sciences, 10(2):253– 265, April 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800444.

Alagic:1975:NST

[299] Suad Alagić. Natural state transformations. Journal of Computer and System Sciences, 10(2):266– 307, April 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800456.

Valiant:1975:GCF

[300] Leslie G. Valiant. General contextfree recognition in less than cubic time. Journal of Computer and System Sciences, 10(2):308– 315, April 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800468.

Beeri:1975:TWN

[301] C. Beeri. Two-way nested stack automata are equivalent to two-way stack automata. Journal of Computer and System Sciences, 10(3): 317–339, June 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800043.

Valiant:1975:DOC

[302] Leslie G. Valiant and Michael S. Paterson. Deterministic one-counter automata. Journal of Computer and System Sciences, 10(3):340– 350, June 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800055.

Goldstine:1975:SIF

[303] Jonathan Goldstine. Some independent families of one-letter languages. Journal of Computer and System Sciences, 10(3):351–369, June 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800067.

Germano:1975:PCC

[304] G. Germano and A. Maggiolo-Schettini. Proving a compiler correct: a simple approach. Journal of Computer and System Sciences, 10(3):370–383, June 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800079.

Ullman:1975:NCS

[305] J. D. Ullman. NP-complete scheduling problems. Journal of Computer and System Sciences, 10(3): 384–393, June 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800080.

McCloskey:1975:AFL

[306] Sister Teresemarie McCloskey. Abstract families of length-preserving processors. Journal of Computer and System Sciences, 10(3):394– 427, June 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800092.

Anonymous:1975:AI

[307] Anonymous. Author index. Journal of Computer and System Sciences, 10(3):428, June 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800109.

Anonymous:1975:EBb

[308] Anonymous. Editorial Board. Journal of Computer and System Sciences, 11(1):i, August 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007580047X.

Rovan:1975:PCB

[309] Branislav Rovan. Proving containment of bounded AFL. Journal of Computer and System Sciences, 11 (1):1–55, August 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800481.

Ibarra:1975:HTM

[310] Oscar H. Ibarra and Sartaj K. Sahni. Hierarchies of Turing machines with restricted tape alphabet size. Journal of Computer and System Sciences, 11(1):56–67, August 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800493.

Jones:1975:SBR

[311] Neil D. Jones. Space-bounded reducibility among combinatorial problems. Journal of Computer and System Sciences, 11(1):68–85, August 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200007580050X.

Cremers:1975:CFG

[312] Armin Cremers and Seymour Ginsburg. Context-free grammar forms. Journal of Computer and System Sciences, 11(1):86–117, August 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000075800511.

Freedman:1975:SBP

[313] A. R. Freedman and R. E. Ladner. Space bounds for processing contentless inputs. Journal of Computer and System Sciences, 11(1): 118–128, August 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800523.

Rozenberg:1975:SPC

[314] G. Rozenberg and K. P. Lee. Some properties of the class of L languages with interactions. Journal of Computer and System Sciences, 11 (1):129–147, August 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800535.

Arbib:1975:EF

[315] Michael A. Arbib and E. K. Blum. Editors’ foreword. Journal of Computer and System Sciences, 11(2):

149, October 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007580064X.

Nishio:1975:FTC

[316] Hidenosuke Nishio and Youichi Kobuchi. Fault tolerant cellular spaces. Journal of Computer and System Sciences, 11 (2):150–170, October 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800651.

Harao:1975:FTC

[317] Masateru Harao and Shoichi Noguchi. Fault tolerant cellular automata. Journal of Computer and System Sciences, 11(2):171–185, October 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800663.

Rajlich:1975:DDS

[318] Václav Rajlich. Dynamics of discrete systems and pattern reproduction. Journal of Computer and System Sciences, 11(2):186– 202, October 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800675.

VanLeeuwen:1975:TCC

[319] Jan Van Leeuwen. The tapecomplexity of context-independent developmental languages. Journal of Computer and System Sciences, 11 (2):203–211, October 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800687.

Ehrig:1975:GGA

[320] Hartmut Ehrig and Karl Wilhelm Tischer. Graph grammars and applications to specialization and evolution in biology. Journal of Computer and System Sciences, 11(2): 212–236, October 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800699.

Hollerer:1975:FCA

[321] W. O. Höllerer and R. Vollmar. On “forgetful” cellular automata. Journal of Computer and System Sciences, 11 (2):237–251, October 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800705.

Walker:1975:FGR

[322] Adrian Walker. Formal grammars and the regeneration capability of biological systems. Journal of Computer and System Sciences, 11(2): 252–261, October 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800717.

Reedy:1975:ADS

[323] Ann Reedy and Walter J. Savitch. Ambiguity in the developmental systems of Lindenmayer. Journal of Computer and System Sciences, 11 (2):262–283, October 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800729.

Baer:1975:CA

[324] Robert M. Baer. Computation by assembly. Journal of Computer and System Sciences, 11(3): 285–294, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800547.

Best:1975:ACD

[325] Michael J. Best and Klaus Ritter. An accelerated conjugate direction method to solve linearly constrained minimization problems. Journal of Computer and System Sciences, 11(3): 295–322, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800559.

deBakker:1975:CIA

[326] J. W. de Bakker and L. G. L. T. Meertens. On the completeness of the inductive assertion method. Journal of Computer and System Sciences, 11 (3):323–357, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800560.

Rosen:1975:PEC

[327] Barry K. Rosen. Program equivalence and context-free grammars. Journal of Computer and System Sciences, 11 (3):358–374, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800572.

Sontag:1975:SQR

[328] Eduardo D. Sontag. On some questions of rationality and decidability. Journal of Computer and System Sciences, 11 (3):375–381, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800584.

DiGregorio:1975:RCA

[329] S. Di Gregorio and G. Trautteur. On reversibility in cellular automata. Journal of Computer and System Sciences, 11(3): 382–391, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800596.

Ehrenfeucht:1975:PD

[330] Andrzej Ehrenfeucht. Practical decidability. Journal of Computer and System Sciences, 11(3):392– 396, December 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000075800602.

Anonymous:1975:E

[331] Anonymous. Erratum. Journal of Computer and System Sciences, 11 (3):397, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800626.

Blum:1975:EN

[332] E. K. Blum. Editor’s note. Journal of Computer and System Sciences, 11(3):397, December 1975. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800614.

Anonymous:1975:AIV

[333] Anonymous. Author index for volume 11. Journal of Computer and System Sciences, 11(3): 398, December 1975. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000075800638.

Anonymous:1976:EB

[334] Anonymous. Editorial Board. Journal of Computer and System Sciences, 12(1):i, February 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800128.

Yaku:1976:SNP

[335] Takeo Yaku. Surjectivity of nondeterministic parallel maps induced bynondeterministic cellular automata. Journal of Computer and System Sciences, 12(1):1–5, February 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007680013X.

Richardson:1976:CSR

[336] D. Richardson. Continuous selfreproduction. Journal of Computer and System Sciences, 12(1):6– 12, February 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800141.

Helton:1976:SAA

[337] F. Joanne Helton. The semigroup of an algebra automaton. Journal of Computer and System Sciences, 12 (1):13–24, February 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800153.

Savage:1976:NSA

[338] Sam Savage, Peter Weiner, and A. Bagchi. Neighborhood search algorithms for guaranteeing optimal traveling salesman tours must be inefficient. Journal of Computer and System Sciences, 12(1):25– 35, February 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800165.

Nasu:1976:CPO

[339] Masakazu Nasu and Namio Honda. A completeness property of onedimensional tessellation automata. Journal of Computer and System Sciences, 12(1):36–48, February 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000076800177.

Ezawa:1976:IL

[340] Yoshinori Ezawa, Masaharu Mizumoto, Junichi Toyoda, and Kokichi Tanaka. Interactive languages. Journal of Computer and System Sciences, 12(1):49–63, February 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800189.

Hart:1976:DLP

[341] Johnson M. Hart. The derivation language of a phrase structure grammar. Journal of Computer and System Sciences, 12(1):64– 79, February 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800190.

Weihrauch:1976:CCP

[342] Klaus Weihrauch. The computational complexity of program schemata. Journal of Computer and System Sciences, 12(1):80–107, February 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800207.

Meyer:1976:SCC

[343] R. R. Meyer. Sufficient conditions for the convergence of monotonic mathematical programming algorithms. Journal of Computer and System Sciences, 12(1):108– 121, February 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800219.

Lewis:1976:CR

[344] F. D. Lewis. On computational reducibility. Journal of Computer and System Sciences, 12(1):122– 131, February 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800220.

Constable:1976:SSI

[345] Robert L. Constable, E. K. Blum, and M. A. Harrison. Special SIGACT issue: Sixth Annual ACM Symposium on Theory of Computing. Journal of Computer and System Sciences, 12(2):133, April 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800335.

Ehrenfeucht:1976:CMR

[346] Andrzej Ehrenfeucht and Paul Zeiger. Complexity measures for regular expressions. Journal of Computer and System Sciences, 12(2):134– 146, April 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800347.

Mehlhorn:1976:PAS

[347] Kurt Mehlhorn. Polynomial and abstract subrecursive classes. Journal of Computer and System Sciences, 12 (2):147–178, April 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800359.

Courcelle:1976:CRE

[348] Bruno Courcelle and Jean Vuillemin. Completeness results for the equivalence of recursive schemas. Journal of Computer and System Sciences, 12(2):179–197, April 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800360.

Pratt:1976:CPV

[349] Vaughan R. Pratt and Larry J. Stockmeyer. A characterization of the power of vector machines. Journal of Computer and System Sciences, 12 (2):198–221, April 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800372.

Hunt:1976:ECC

[350] Harry B. Hunt, Daniel J. Rosenkrantz, and Thomas G. Szymanski. On the equivalence, containment, and covering problems for the regular and context-free languages. Journal of Computer and System Sciences, 12 (2):222–268, April 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800384.

Ehrig:1976:SAR

[351] Hartmut Ehrig and Hans-Jörg Kreowski. Systematic approach to reduction and minimization in automata and system theory. Journal of Computer and System Sciences, 12(3): 269–304, June 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800025.

Bloom:1976:ECF

[352] Stephen L. Bloom and Calvin C. Elgot. The existence and construction of free iterative theories. Journal of Computer and System Sciences, 12(3):305–318, June 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800037.

Bruno:1976:BSJ

[353] J. Bruno, E. G. Coffman, and D. B. Johnson. On batch scheduling of jobs with stochastic service times and cost structures on a single server. Journal of Computer and System Sciences, 12(3):319–335, June 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800049.

Breitbart:1976:SBC

[354] Y. Breitbart. Some bounds on the complexity of predicate recognition by finite automata. Journal of Computer and System Sciences, 12(3): 336–349, June 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800050.

Galil:1976:NME

[355] Zvi Galil and Janos Simon. A note on multiple-entry finite automata. Journal of Computer and System Sciences, 12(3):350–351, June 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800062.

Hashiguchi:1976:PCE

[356] K. Hashiguchi and N. Honda. Properties of code events and homomorphisms over regular events. Journal of Computer and System Sciences, 12(3):352–367, June 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800074.

Alton:1976:NPO

[357] Donald A. Alton. Nonexistence of program optimizers in several abstract settings. Journal of Computer and System Sciences, 12(3): 368–393, June 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800086.

Anderson:1976:APT

[358] Peter Gordon Anderson. Another proof of the theorem on pattern reproduction in tessellation structures. Journal of Computer and System Sciences, 12(3):394–398, June 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800098.

Goldstine:1976:BA

[359] J. Goldstine. Bounded AFLs. Journal of Computer and System Sciences, 12(3):399–419, June 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800104.

Anonymous:1976:AIVa

[360] Anonymous. Author index for volume 12. Journal of Computer and System Sciences, 12(3): 420, June 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800116.

Peterson:1976:CSS

[361] James L. Peterson. Computation sequence sets. Journal of Computer and System Sciences, 13(1): 1–24, August 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800475.

Cook:1976:SRD

[362] Stephen Cook and Ravi Sethi. Storage requirements for deterministic polynomial-time recognizable languages. Journal of Computer and System Sciences, 13(1):25–37, August 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000076800487.

Taniguchi:1976:REP

[363] Kenichi Taniguchi and Tadao Kasami. A result on the equivalence problem for deterministic pushdown automata. Journal of Computer and System Sciences, 13(1):38–50, August 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000076800499.

Fischer:1976:SNF

[364] Patrick C. Fischer, Edward L. Robertson, and Lawrence V. Saxton. On the sequential nature of functions. Journal of Computer and System Sciences, 13(1):51–68, August 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800505.

Dobkin:1976:NLB

[365] David Dobkin. A nonlinear lower bound on linear search tree programs for solving knapsack problems. Journal of Computer and System Sciences, 13(1):69–73, August 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800517.

Masunaga:1976:CAD

[366] Yoshifumi Masunaga, Shoichi Noguchi, and Juro Oizumi. A characterization of automata and a direct product decomposition. Journal of Computer and System Sciences, 13(1): 74–89, August 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800529.

Rozenberg:1976:CFG

[367] G. Rozenberg and A. Salomaa. Context-free grammars with graphcontrolled tables. Journal of Computer and System Sciences, 13(1):

90–99, August 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800530.

Lynch:1976:CCE

[368] Nancy Lynch. Complexity-class-encoding sets. Journal of Computer and System Sciences, 13(1): 100–118, August 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800542.

Baer:1976:HPL

[369] Robert M. Baer and Jan van Leeuwen. The halting problem for linear Turing assemblers. Journal of Computer and System Sciences, 13(2): 119–135, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800232.

Amoroso:1976:IPM

[370] Serafino Amoroso and Irving J. Epstein. Indecomposable parallel maps in tessellation structures. Journal of Computer and System Sciences, 13 (2):136–142, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800244. See

note [468].

Shibata:1976:SAA

[371] Yukio Shibata. On the structure of abelian automata. Journal of Computer and System Sciences, 13(2):

143–152, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800256.

Ibarra:1976:UDS

[372] Oscar H. Ibarra and Chul E. Kim. A useful device for showing the solvability of some decision problems. Journal of Computer and System Sciences, 13 (2):153–160, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800268.

Nozaki:1976:HAA

[373] A. Nozaki. Hazard analysis of asynchronous circuits in Muller– Bartky’s sense. Journal of Computer and System Sciences, 13(2): 161–171, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007680027X.

Laing:1976:AI

[374] Richard Laing. Automaton introspection. Journal of Computer and System Sciences, 13(2):172– 183, October 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800281.

Schonhage:1976:FM

[375] A. Schönhage, M. Paterson, and N. Pippenger. Finding the median. Journal of Computer and System Sciences, 13(2):184–199, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000076800293.

Bloom:1976:VOA

[376] Stephen L. Bloom. Varieties of ordered algebras. Journal of Computer and System Sciences, 13(2): 200–212, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007680030X.

McCormick:1976:SCR

[377] S. F. McCormick. Some convergence results on the method of gradients for Ax = λBx. Journal of Computer and System Sciences, 13 (2):213–222, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800311.

Arnold:1976:TDP

[378] André Arnold and Max Dauchet. Un théorème de duplication pour les forêts algébriques. (French) [A duplication theorem for algebraic forests]. Journal of Computer and System Sciences, 13 (2):223–244, October 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800323.

Rounds:1976:EF

[379] William C. Rounds, E. K. Blum, and M. A. Harrison. Editors’ foreword. Journal of Computer and System Sciences, 13(3):245, December 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800396.

Perrault:1976:ILT

[380] C. Raymond Perrault. Intercalation lemmas for tree transducer languages. Journal of Computer and System Sciences, 13(3):246– 277, December 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800402.

Valiant:1976:GTP

[381] Leslie G. Valiant. Graph-theoretic properties in computational complexity. Journal of Computer and System Sciences, 13(3):278– 285, December 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800414.

Aho:1976:NLR

[382] A. V. Aho and J. D. Ullman. Node listings for reducible flow graphs. Journal of Computer and System Sciences, 13 (3):286–299, December 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000076800426.

Miller:1976:RHT

[383] Gary L. Miller. Riemann’s hypothesis and tests for primality. Journal of Computer and System Sciences, 13 (3):300–317, December 1976. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800438.

Hunt:1976:CMC

[384] Harry B. Hunt and Thomas G. Szymanski. Complexity metatheorems for context-free grammar problems. Journal of Computer and System Sciences, 13(3):318– 334, December 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007680044X.

Booth:1976:TCO

[385] Kellogg S. Booth and George S. Lueker. Testing for the consecutive ones property, interval graphs, and graph planarity using PQ-tree algorithms. Journal of Computer and System Sciences, 13(3):335– 379, December 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800451.

Anonymous:1976:AIVb

[386] Anonymous. Author index for volume 13. Journal of Computer and System Sciences, 13(3): 380, December 1976. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000076800463.

Anonymous:1977:EBa

[387] Anonymous. Editorial Board. Journal of Computer and System Sciences, 14(1):i, February 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800378.

Fleischner:1977:EMT

[388] Herbert Fleischner. On the equivalence of Mealy-type and Moore-type automata and a relation between reducibility and Moore-reducibility. Journal of Computer and System Sciences, 14(1):1–16, February 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S002200007780038X.

Sikdar:1977:TM

[389] Kripasindhu Sikdar. On topological machines. Journal of Computer and System Sciences, 14(1):17– 48, February 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800391.

Lipton:1977:SCC

[390] R. J. Lipton, R. E. Miller, and L. Snyder. Synchronization and computing capabilities of linear asynchronous structures. Journal of Computer and System Sciences, 14(1):49– 72, February 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800408.

Seiferas:1977:TSS

[391] Joel I. Seiferas. Techniques for separating space complexity classes. Journal of Computer and System Sciences, 14(1):73–99, February 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200007780041X.

Seiferas:1977:RRS

[392] Joel I. Seiferas. Relating refined space complexity classes. Journal of Computer and System Sciences, 14 (1):100–129, February 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800421.

Rayward-Smith:1977:HEM

[393] V. J. Rayward-Smith. Hypergrammars: an extension of macrogrammars. Journal of Computer and System Sciences, 14(1):130– 149, February 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800433.

Hartfiel:1977:SCL

[394] D. J. Hartfiel and C. J. Maxson. A semigroup characterization of a linearly realizable automaton over GF(p). Journal of Computer and System Sciences, 14(1): 150–155, February 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800445.

Rosenberg:1977:SCA

[395] Arnold L. Rosenberg. On storing concatenable arrays. Journal of Computer and System Sciences, 14 (2):157–174, April 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007780010X.

Rodrigue:1977:MCA

[396] Garry Rodrigue. The method of closest approach for finding nonisolate zeros of functions. Journal of Computer and System Sciences, 14 (2):175–186, April 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800111.

Wotschke:1977:DLN

[397] Detlef Wotschke. Degree-languages: a new concept of acceptance. Journal of Computer and System Sciences, 14(2):187–209, April 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800123.

Plaisted:1977:SCP

[398] David Alan Plaisted. Sparse complex polynomials and polynomial reducibility. Journal of Computer and System Sciences, 14(2):210– 221, April 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800135.

Fagin:1977:AMR

[399] Ronald Fagin. Asymptotic miss ratios over independent references. Journal of Computer and System Sciences, 14(2):222–250, April 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800147.

Bloom:1977:SVI

[400] Stephen L. Bloom, Susanna Ginali, and Joseph D. Rutledge. Scalar and vector iteration. Journal of Computer and System Sciences, 14 (2):251–256, April 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800159.

Ae:1977:DCP

[401] Tadashi Ae. Direct or cascade product of pushdown automata. Journal of Computer and System Sciences, 14(2):257–263, April 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800160.

Geller:1977:CPFa

[402] Matthew M. Geller and Michael A. Harrison. Characteristic parsing: a framework for producing compact deterministic parsers, I. Journal of Computer and System Sciences, 14 (3):265–317, June 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800172.

Geller:1977:CPFb

[403] Matthew M. Geller and Michael A. Harrison. Characteristic parsing: a framework for producing compact deterministic parsers, II. Journal of Computer and System Sciences, 14 (3):318–343, June 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800184.

Friedman:1977:EPD

[404] Emily P. Friedman. Equivalence problems for deterministic contextfree languages and monadic recursion schemes. Journal of Computer and System Sciences, 14(3): 344–359, June 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800196.

Ching:1977:DRP

[405] W. S. Ching and B. F. Wyman. Duality and the regulator problem for linear systems over commutative rings. Journal of Computer and System Sciences, 14(3):360–368, June 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800202.

Maibaum:1977:E

[406] T. S. E. Maibaum. Erratum. Journal of Computer and System Sciences, 14(3):369, June 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800214.

Anonymous:1977:AIVa

[407] Anonymous. Author index for volume 14. Journal of Computer and System Sciences, 14(3): 370, June 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800226.

Anonymous:1977:EBb

[408] Anonymous. Editorial Board. Journal of Computer and System Sciences, 15(1):i, August 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800238.

Watanabe:1977:AA

[409] Toshimasa Watanabe and Shoichi Noguchi. The amalgamation of automata. Journal of Computer and System Sciences, 15(1):1–16, August 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200007780024X.

Tsuji:1977:RDG

[410] T. Tsuji, J. Toyoda, and K. Tanaka. Relational data graphs and some properties of them. Journal of Computer and System Sciences, 15(1): 17–34, August 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800251.

Greibach:1977:CSC

[411] S. A. Greibach. Control sets on context-free grammar forms. Journal of Computer and System Sciences, 15(1):35–98, August 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800263.

Nijholt:1977:CPG

[412] Anton Nijholt. On the covering of parsable grammars. Journal of Computer and System Sciences, 15 (1):99–110, August 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800275.

Atanasiu:1977:CGT

[413] Adrian Atanasiu. C-grammars and tree-codifications. Journal of Computer and System Sciences, 15(1): 111–120, August 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800287.

Sato:1977:CRB

[414] Tadakazu Sato and Namio Honda. Certain relations between properties of maps of tessellation automata. Journal of Computer and System Sciences, 15(2):121–145, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800020.

Reusch:1977:LRF

[415] Bernd Reusch. Linear realization of finite automata. Journal of Computer and System Sciences, 15(2): 146–168, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800032.

Cohen:1977:TLC

[416] Rina S. Cohen and Arie Y. Gold. Theory of ω-languages: I: Characterizations of ω-context-free languages. Journal of Computer and System Sciences, 15(2):169–184, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800044.

Cohen:1977:TLI

[417] Rina S. Cohen and Arie Y. Gold. Theory of ω-languages. II: a study of various models of ω-type generation and recognition. Journal of Computer and System Sciences, 15 (2):185–208, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800056.

Wand:1977:CWP

[418] Mitchell Wand. A characterization of weakest preconditions. Journal of Computer and System Sciences, 15 (2):209–212, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800068.

Toffoli:1977:CCU

[419] Tommaso Toffoli. Computation and construction universality of reversible cellular automata. Journal of Computer and System Sciences, 15(2):

213–231, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200007780007X.

Revesz:1977:APD

[420] György Révész. Algebraic properties of derivation words. Journal of Computer and System Sciences, 15 (2):232–240, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800081.

Jones:1977:SBR

[421] Neil D. Jones. Space-bounded reducibility among combinatorial problems. Journal of Computer and System Sciences, 15(2):241, October 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000077800093.

Willson:1977:LSC

[422] Stephen J. Willson. Limiting shapes for configurations. Journal of Computer and System Sciences, 15(3): 243–261, December 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800299.

Cremers:1977:SCF

[423] Armin Cremers, Seymour Ginsburg, and Edwin H. Spanier. The structure of context-free grammatical families. Journal of Computer and System Sciences, 15(3): 262–279, December 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800305.

Burkhard:1977:ART

[424] Walter A. Burkhard. Associative retrieval trie hash-coding. Journal of Computer and System Sciences, 15(3): 280–299, December 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800317.

Fong:1977:FDF

[425] Amelia C. Fong and Jeffrey D. Ullman. Finding the depth of a flow graph. Journal of Computer and System Sciences, 15(3): 300–309, December 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800329.

Blattner:1977:SVT

[426] Meera Blattner and Tom Head. Singlevalued a-transducers. Journal of Computer and System Sciences, 15(3): 310–327, December 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000077800330.

Engelfriet:1977:IO

[427] Joost Engelfriet and Erik Meineche Schmidt. IO and OI. I. Journal of Computer and System Sciences, 15(3): 328–353, December 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800342.

Golan:1977:CG

[428] Igal Golan. Conditional grammars. Journal of Computer and System Sciences, 15(3):354–371, December 1977. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000077800354.

Anonymous:1977:AIVb

[429] Anonymous. Author index for volume 15. Journal of Computer and System Sciences, 15(3): 372, December 1977. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000077800366.

Baker:1978:NPF

[430] Theodore P. Baker. “Natural” properties of flowchart step-counting measures. Journal of Computer and System Sciences, 16(1):1–22, February 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900478.

Smyth:1978:PD

[431] M. B. Smyth. Power domains. Journal of Computer and System Sciences, 16(1):23–36, February 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200007890048X.

Brzozowski:1978:DDH

[432] J. A. Brzozowski and R. Knast. The dot-depth hierarchy of star-free languages is infinite. Journal of Computer and System Sciences, 16(1):37– 55, February 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900491.

Klove:1978:CUL

[433] Torleiv Kløve. On complements of unary L languages. Journal of Computer and System Sciences, 16 (1):56–66, February 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900508.

Engelfriet:1978:IOI

[434] Joost Engelfriet and Erik Meineche Schmidt. IO and OI. II. Journal of Computer and System Sciences, 16 (1):67–99, February 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007890051X.

Berger:1978:IRB

[435] J. Berger and C. Pair. Inference for regular bilanguages. Journal of Computer and System Sciences, 16 (1):100–122, February 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900521.

Chandra:1978:EF

[436] Ashok Chandra, E. K. Blum, and M. A. Harrison. Editors’ foreword. Journal of Computer and System Sciences, 16(2):123, April 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900405.

Lipton:1978:EPS

[437] Richard J. Lipton and Larry J. Stockmeyer. Evaluation of polynomials with super-preconditioning. Journal of Computer and System Sciences, 16 (2):124–139, April 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900417.

Galil:1978:PRR

[438] Zvi Galil. Palindrome recognition in real time by a multitape Turing machine. Journal of Computer and System Sciences, 16(2): 140–157, April 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900429.

Paterson:1978:LU

[439] M. S. Paterson and M. N. Wegman. Linear unification. Journal of Computer and System Sciences, 16 (2):158–167, April 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900430.

Manders:1978:NCD

[440] Kenneth L. Manders and Leonard Adleman. NP-complete decision problems for binary quadratics. Journal of Computer and System Sciences, 16(2):168–184, April 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900442.

Schaefer:1978:CST

[441] Thomas J. Schaefer. On the complexity of some two-person perfectinformation games. Journal of Computer and System Sciences, 16(2): 185–225, April 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900454.

Guibas:1978:ADH

[442] Leo J. Guibas and Endre Szemerédi. The analysis of double hashing. Journal of Computer and System Sciences, 16(2):226–274, April 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900466.

Cohen:1978:CDP

[443] Rina S. Cohen and Arie Y. Gold. ωcomputations on deterministic pushdown machines. Journal of Computer and System Sciences, 16(3): 275–300, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900193.

Jonassen:1978:TAW

[444] Arne T. Jonassen and Donald E. Knuth. A trivial algorithm whose analysis isn’t. Journal of Computer and System Sciences, 16(3): 301–322, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200007890020X. See

continuation [3943].

Oppen:1978:UBC

[445] Derek C. Oppen. A 2^{2^{2^{pn}}} upper bound on the complexity of Presburger Arithmetic. Journal of Computer and System Sciences, 16(3): 323–332, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900211.

Chin:1978:AUM

[446] Francis Chin and David Houck. Algorithms for updating minimal spanning trees. Journal of Computer and System Sciences, 16(3): 333–344, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900223.

Maurer:1978:EF

[447] H. A. Maurer, A. Salomaa, and D. Wood. ET0L forms. Journal of Computer and System Sciences, 16(3):345–361, June 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900235.

Elgot:1978:ASR

[448] Calvin C. Elgot, Stephen L. Bloom, and Ralph Tindell. On the algebraic structure of rooted trees. Journal of Computer and System Sciences, 16(3):362–399, June 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900247.

Schuler:1978:IDF

[449] P. F. Schuler. Inductive definability in formal language theory. Journal of Computer and System Sciences, 16(3):400–412, June 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900259.

Dobkin:1978:LBL

[450] David Dobkin and Richard J. Lipton. A lower bound of 12n^{2} on linear search programs for the Knapsack problem. Journal of Computer and System Sciences, 16(3): 413–417, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900260.

Hartmanis:1978:PTI

[451] J. Hartmanis and L. Berman. On polynomial time isomorphisms of some new complete sets. Journal of Computer and System Sciences, 16(3): 418–422, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900272.

Vitanyi:1978:IDP

[452] Paul M. B. Vitányi and Walter J. Savitch. On inverse deterministic pushdown transductions. Journal of Computer and System Sciences, 16 (3):423–444, June 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900284.

Barto:1978:NPR

[453] Andrew G. Barto. A note on pattern reproduction in tessellation structures. Journal of Computer and System Sciences, 16(3): 445–455, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900296.

Wotschke:1978:NBO

[454] Detlef Wotschke. Nondeterminism and boolean operations in pda’s. Journal of Computer and System Sciences, 16(3):456–461, June 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900302.

Anonymous:1978:AIVa

[455] Anonymous. Author index for volume 16. Journal of Computer and System Sciences, 16(3): 462, June 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900314.

Inagaki:1978:AAN

[456] Yasuyoshi Inagaki, Toshiki Sakabe, and Teruo Fukumura. Addressable approximations to nonaddressable data graphs. Journal of Computer and System Sciences, 17(1): 1–34, August 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900326.

Jones:1978:ESP

[457] Anita K. Jones and Richard J. Lipton. The enforcement of security policies for computation. Journal of Computer and System Sciences, 17 (1):35–55, August 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900338.

Breidbart:1978:SRS

[458] Seth Breidbart. On splitting recursive sets. Journal of Computer and System Sciences, 17(1): 56–64, August 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007890034X.

Ito:1978:RSC

[459] Masami Ito. A representation of strongly connected automata and its applications. Journal of Computer and System Sciences, 17(1):

65–80, August 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900351.

Choueka:1978:FAD

[460] Yaacov Choueka. Finite automata, definable sets, and regular expressions over ω^{n}-tapes. Journal of Computer and System Sciences, 17 (1):81–97, August 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900363.

Maxson:1978:ELA

[461] Carlton J. Maxson and Kirby C. Smith. Endomorphisms of linear automata. Journal of Computer and System Sciences, 17(1): 98–107, August 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900375.

McColl:1978:CDS

[462] William F. McColl. The circuit depth of symmetric boolean functions. Journal of Computer and System Sciences, 17(1):108–115, August 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900387.

Hegner:1978:DTD

[463] Stephen J. Hegner. Duality theory for discrete-time linear systems. Journal of Computer and System Sciences, 17 (1):116–143, August 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900399.

Pippenger:1978:RNB

[464] Nicholas Pippenger. On rearrangeable and non-blocking switching networks. Journal of Computer and System Sciences, 17(2): 145–162, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900016.

Culik:1978:DHE

[465] K. Culik and Arto Salomaa. On the decidability of homomorphism equivalence for languages. Journal of Computer and System Sciences, 17 (2):163–175, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900028.

Golze:1978:SND

[466] Ulrich Golze. (a-)synchronous (non- )deterministic cell spaces simulating each other. Journal of Computer and System Sciences, 17(2): 176–193, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007890003X.

Reusch:1978:RFA

[467] Bernd Reusch. Realization of finite automata with linear parts. Journal of Computer and System Sciences, 17 (2):194–206, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900041.

Aggarwal:1978:NIP

[468] Sudhir Aggarwal. A note on “indecomposable parallel maps” by S. Amoroso and I. J. Epstein. Journal of Computer and System Sciences, 17 (2):207–208, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900053. See

[370].

Plotkin:1978:UD

[469] G. Plotkin. T^{ω} as a Universal Domain. Journal of Computer and System Sciences, 17(2): 209–236, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900065.

Priese:1978:NAC

[470] Lutz Priese. A note on asynchronous cellular automata. Journal of Computer and System Sciences, 17(2): 237–252, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900077.

Moore:1978:CTO

[471] Dan Moore and John Case. The complexity of total order structures. Journal of Computer and System Sciences, 17(2):253–269, October 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900089.

Lucchesi:1978:CKR

[472] Cláudio L. Lucchesi and Sylvia L. Osborn. Candidate keys for relations. Journal of Computer and System Sciences, 17(2):270– 279, October 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900090.

King:1978:SLS

[473] K. N. King and C. Wrathall. Stack languages and log n space. Journal of Computer and System Sciences, 17 (3):281–299, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900107.

Kobayashi:1978:FSS

[474] Kojiro Kobayashi. The firing squad synchronization problem for a class of polyautomata networks. Journal of Computer and System Sciences, 17(3): 300–318, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900119.

Maibaum:1978:PLT

[475] T. S. E. Maibaum. Pumping lemmas for term languages. Journal of Computer and System Sciences, 17(3): 319–330, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900120.

Ehrenfeucht:1978:SDD

[476] A. Ehrenfeucht and G. Rozenberg. On the structure of derivations in deterministic ET0L systems. Journal of Computer and System Sciences, 17(3): 331–347, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900132.

Milner:1978:TTP

[477] Robin Milner. A theory of type polymorphism in programming. Journal of Computer and System Sciences, 17 (3):348–375, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900144.

Takaoka:1978:DML

[478] Tadao Takaoka. A definition of measures over language space. Journal of Computer and System Sciences, 17 (3):376–387, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000078900156.

Courcelle:1978:SCI

[479] Bruno Courcelle and Irène Guessarian. On some classes of interpretations. Journal of Computer and System Sciences, 17(3): 388–413, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900168.

Rabinovitz:1978:CPL

[480] Moshe Rabinovitz. Closure properties and languages defined by bilinear automata. Journal of Computer and System Sciences, 17(3): 414–423, December 1978. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007890017X.

Anonymous:1978:AIVb

[481] Anonymous. Author index for volume 17. Journal of Computer and System Sciences, 17(3): 424, December 1978. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000078900181.

Butler:1979:DMG

[482] Jon T. Butler. Decomposable maps in general tessellation structures. Journal of Computer and System Sciences, 18(1):1–7, February 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900473.

Nozaki:1979:EPN

[483] A. Nozaki. Equivalence problem of non-deterministic finite automata. Journal of Computer and System Sciences, 18(1):8–17, February 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900485.

Gurari:1979:SDP

[484] Eitan M. Gurari and Oscar H. Ibarra. Some decision problems concerning sequential transducers and checking automata. Journal of Computer and System Sciences, 18(1):18– 34, February 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900497.

Valette:1979:APN

[485] R. Valette. Analysis of petri nets by stepwise refinements. Journal of Computer and System Sciences, 18 (1):35–46, February 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900503.

Maruoka:1979:ISP

[486] Akira Maruoka and Masayuki Kimura. Injectivity and surjectivity of parallel maps for cellular automata. Journal of Computer and System Sciences, 18 (1):47–64, February 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900515.

Rouchaleau:1979:EMR

[487] Yves Rouchaleau and Eduardo D. Sontag. On the existence of minimal realizations of linear dynamical systems over Noetherian integral domains. Journal of Computer and System Sciences, 18(1):65– 75, February 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900527.

Mitchell:1979:LAR

[488] S. L. Mitchell, E. J. Cockayne, and S. T. Hedetniemi. Linear algorithms on recursive representations of trees. Journal of Computer and System Sciences, 18(1):76–85, February 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900539.

Dobkin:1979:CCU

[489] David P. Dobkin and Richard J. Lipton. On the complexity of computations under varying sets of primitives. Journal of Computer and System Sciences, 18(1):86– 91, February 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900540.

Linna:1979:TDR

[490] Matti Linna. Two decidability results for deterministic pushdown automata. Journal of Computer and System Sciences, 18(1):92–107, February 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900552.

Miller:1979:EF

[491] Raymond E. Miller, E. K. Blum, and R. M. Karp. Editors’ foreword. Journal of Computer and System Sciences, 18(2):109, April 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900412.

Tarjan:1979:CAW

[492] Robert Endre Tarjan. A class of algorithms which require nonlinear time to maintain disjoint sets. Journal of Computer and System Sciences, 18 (2):110–127, April 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900424.

Miller:1979:GIG

[493] Gary L. Miller. Graph isomorphism, general remarks. Journal of Computer and System Sciences, 18 (2):128–142, April 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900436.

Carter:1979:UCH

[494] J. Lawrence Carter and Mark N. Wegman. Universal classes of hash functions. Journal of Computer and System Sciences, 18(2): 143–154, April 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900448.

Angluin:1979:FPA

[495] D. Angluin and L. G. Valiant. Fast probabilistic algorithms for Hamiltonian circuits and matchings. Journal of Computer and System Sciences, 18(2):155–193, April 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007990045X.

Fischer:1979:PDL

[496] Michael J. Fischer and Richard E. Ladner. Propositional dynamic logic of regular programs. Journal of Computer and System Sciences, 18 (2):194–211, April 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900461.

Werschulz:1979:OOA

[497] Arthur G. Werschulz. Optimal order for approximation of derivatives. Journal of Computer and System Sciences, 18(3):213–217, June 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900308.

Poplawski:1979:LRG

[498] David A. Poplawski. On LLregular grammars. Journal of Computer and System Sciences, 18(3): 218–227, June 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007990031X.

Ginali:1979:RTF

[499] Susanna Ginali. Regular trees and the free iterative theory. Journal of Computer and System Sciences, 18 (3):228–242, June 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900321.

Ono:1979:UFO

[500] Hiroakira Ono and Akira Nakamura. Undecidability of the first-order arithmetic A[P (x), 2x, x + 1]. Journal of Computer and System Sciences, 18 (3):243–253, June 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900333.

Friedman:1979:MRS

[501] Emily P. Friedman and Sheila A. Greibach. Monadic recursion schemes: The effect of constants. Journal of Computer and System Sciences, 18 (3):254–266, June 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900345.

Paun:1979:FFI

[502] Gheorghe Páun. On the family of finite index matrix languages. Journal of Computer and System Sciences, 18(3):267–280, June 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900357.

Weyuker:1979:MPS

[503] Elaine J. Weyuker. Modifications of the program scheme model. Journal of Computer and System Sciences, 18(3):281–293, June 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900369.

Gordon:1979:CCP

[504] Dan Gordon. Complexity classes of provable recursive functions. Journal of Computer and System Sciences, 18(3):294–303, June 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900370.

Veloso:1979:SRM

[505] Paulo A. S. Veloso and Arthur Gill. Some remarks on multiple-entry finite automata. Journal of Computer and System Sciences, 18(3): 304–306, June 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900382.

Latteux:1979:CRC

[506] Michel Latteux. Cônes rationnels commutatifs. Journal of Computer and System Sciences, 18(3): 307–333, June 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900394.

Anonymous:1979:AIVa

[507] Anonymous. Author index for volume 18. Journal of Computer and System Sciences, 18(3): 334, June 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900400.

Pair:1979:CPS

[508] C. Pair, M. Amirchahy, and D. Neel. Correctness proofs of syntax-directed processing descriptions by attributes. Journal of Computer and System Sciences, 19(1):1–17, August 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000079900096.

Maxson:1979:ALA

[509] Carlton J. Maxson and Kirby C. Smith. Automorphisms of linear automata. Journal of Computer and System Sciences, 19(1): 18–26, August 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900102.

Wand:1979:FAS

[510] Mitchell Wand. Final algebra semantics and data type extensions. Journal of Computer and System Sciences, 19(1):27–44, August 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900114.

Blattner:1979:DED

[511] Meera Blattner and Tom Head. The decidability of equivalence for deterministic finite transducers. Journal of Computer and System Sciences, 19(1):45–49, August 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900126.

Benson:1979:PPN

[512] David B. Benson. Parameter passing in nondeterministic recursive programs. Journal of Computer and System Sciences, 19(1): 50–62, August 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900138.

Harrison:1979:HDL

[513] Michael A. Harrison and Amiram Yehudai. A hierarchy of deterministic languages. Journal of Computer and System Sciences, 19(1): 63–78, August 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007990014X.

Friedman:1979:SDM

[514] E. P. Friedman and S. A. Greibach. Superdeterministic DPDAs: The method of accepting does affect decision problems. Journal of Computer and System Sciences, 19(1): 79–117, August 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900151.

Harju:1979:SRA

[515] Tero Harju. A simulation result for the auxiliary pushdown automata. Journal of Computer and System Sciences, 19 (2):119–132, October 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900230.

Germano:1979:CSF

[516] Giorgio Germano and Andrea Maggiolo-Schettini. Computable stack functions for semantics of stack programs. Journal of Computer and System Sciences, 19(2):133–144, October 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900242.

Gurari:1979:SCM

[517] Eitan M. Gurari and Oscar H. Ibarra. Simple counter machines and numbertheoretic problems. Journal of Computer and System Sciences, 19(2): 145–162, October 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900254.

Adamek:1979:LFP

[518] Jirí Adámek and Václav Koubek. Least fixed point of a functor. Journal of Computer and System Sciences, 19 (2):163–178, October 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900266.

Rozenberg:1979:RES

[519] G. Rozenberg and D. Vermeir. On recursion in ETOL systems. Journal of Computer and System Sciences, 19 (2):179–196, October 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900278.

Paul:1979:TH

[520] W. J. Paul. On time hierarchies. Journal of Computer and System Sciences, 19(2):197–202, October 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200007990028X.

Head:1979:CLS

[521] Tom Head. Codeterministic Lindenmayer schemes and systems. Journal of Computer and System Sciences, 19 (2):203–210, October 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900291.

Kanetkar:1979:GCT

[522] Sharad V. Kanetkar and Meghanad D. Wagh. Group character tables in discrete transform theory. Journal of Computer and System Sciences, 19(3): 211–221, December 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900011.

Hunt:1979:OCR

[523] H. B. Hunt. Observations on the complexity of regular expression problems. Journal of Computer and System Sciences, 19(3): 222–236, December 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900023.

Ehrenfeucht:1979:ESR

[524] A. Ehrenfeucht, G. Rozenberg, and D. Vermeir. On ET0L systems with rank. Journal of Computer and System Sciences, 19(3):237– 255, December 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000079900035.

Book:1979:RM

[525] Ronald V. Book, Sheila A. Greibach, and Celia Wrathall. Reset machines. Journal of Computer and System Sciences, 19(3):256–276, December 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000079900047.

Naude:1979:UR

[526] G. Naudé. Universal realization. Journal of Computer and System Sciences, 19(3):277–289, December 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000079900059.

Francez:1979:SNC

[527] Nissim Francez, C. A. R. Hoare, Daniel J. Lehmann, and Willem P. De Roever. Semantics of nondeterminism, concurrency, and communication. Journal of Computer and System Sciences, 19(3):290– 308, December 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900060.

Nozaki:1979:SUN

[528] A. Nozaki. Sorting using networks of deques. Journal of Computer and System Sciences, 19(3): 309–315, December 1979. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900072.

Anonymous:1979:AIVb

[529] Anonymous. Author index for volume 19. Journal of Computer and System Sciences, 19(3): 316, December 1979. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000079900084.

Kasai:1980:CTC

[530] Takumi Kasai and Akeo Adachi. A characterization of time complexity by simple loop programs. Journal of Computer and System Sciences, 20 (1):1–17, February 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008090001X.

Masek:1980:FAC

[531] William J. Masek and Michael S. Paterson. A faster algorithm computing string edit distances. Journal of Computer and System Sciences, 20 (1):18–31, February 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900021.

Brzozowski:1980:LRT

[532] J. A. Brzozowski and Faith E. Fich. Languages of R-trivial monoids. Journal of Computer and System Sciences, 20(1):32–49, February 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900033.

Gallant:1980:FML

[533] John Gallant, David Maier, and James Astorer. On finding minimal length superstrings. Journal of Computer and System Sciences, 20 (1):50–58, February 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900045.

Miyano:1980:OWW

[534] Satoru Miyano. One-way weak-stack-counter automata. Journal of Computer and System Sciences, 20(1):59– 76, February 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900057.

Gonzalez:1980:CCB

[535] Teofilo Gonzalez and Joseph Ja’Ja’. On the complexity of computing bilinear forms with 0, 1 constants. Journal of Computer and System Sciences, 20(1):77–95, February 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900069.

Hennessy:1980:PSF

[536] Matthew C. B. Hennessy. A proof system for the first-order relational calculus. Journal of Computer and System Sciences, 20(1): 96–110, February 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900070.

Dietz:1980:RTE

[537] Paul Dietz and S. Rao Kosaraju. Recognition of topological equivalence of patterns by array automata. Journal of Computer and System Sciences, 20 (1):111–116, February 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900082.

Lipton:1980:EF

[538] Richard Lipton, E. K. Blum, and A. V. Aho. Editors’ foreword. Journal of Computer and System Sciences, 20(2):117, April 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900550.

Tompa:1980:TST

[539] Martin Tompa. Time-space tradeoffs for computing functions, using connectivity properties of their circuits. Journal of Computer and System Sciences, 20(2):118–132, April 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900562.

Lapaugh:1980:SHP

[540] Andrea S. Lapaugh and Ronald L. Rivest. The subgraph homeomorphism problem. Journal of Computer and System Sciences, 20(2): 133–149, April 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900574.

Engelfriet:1980:TTS

[541] Joost Engelfriet, Grzegorz Rozenberg, and Giora Slutzki. Tree transducers, L systems, and two-way machines. Journal of Computer and System Sciences, 20(2):150–202, April 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900586.

Miller:1980:FSS

[542] Raymond E. Miller and Chee K. Yap. On formulating simultaneity for studying parallelism and synchronization. Journal of Computer and System Sciences, 20(2): 203–218, April 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900598.

Lewis:1980:NDP

[543] John M. Lewis and Mihalis Yannakakis. The node-deletion problem for hereditary properties is NP-complete. Journal of Computer and System Sciences, 20(2):219–230, April 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900604.

Pratt:1980:NOM

[544] Vaughan R. Pratt. A near-optimal method for reasoning about action. Journal of Computer and System Sciences, 20(2):231–254, April 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000080900616.

Filotti:1980:AIC

[545] I. S. Filotti. An algorithm for imbedding cubic graphs in the torus. Journal of Computer and System Sciences, 20(2):255–276, April 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900628.

Ginzburg:1980:VAS

[546] A. Ginzburg and M. Yoeli. Vector addition systems and regular languages. Journal of Computer and System Sciences, 20(3): 277–284, June 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900094.

Boasson:1980:AL

[547] Luc Boasson and Maurice Nivat. Adherences of languages. Journal of Computer and System Sciences, 20 (3):285–309, June 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900100.

Paredaens:1980:IIC

[548] Jan Paredaens. The interaction of integrity constraints in an information system. Journal of Computer and System Sciences, 20(3): 310–329, June 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900112.

Arbib:1980:FST

[549] Michael A. Arbib and Ernest G. Manes. Foundations of system theory: The Hankel matrix. Journal of Computer and System Sciences, 20(3):330–378, June 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900124.

Culik:1980:TSC

[550] Karel Culik and Arto Salomaa. Test sets and checking words for homomorphism equivalence. Journal of Computer and System Sciences, 20 (3):379–395, June 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900136.

Rivest:1980:CEB

[551] R. L. Rivest, A. R. Meyer, D. J. Kleitman, K. Winklmann, and J. Spencer. Coping with errors in binary search procedures. Journal of Computer and System Sciences, 20(3):396– 404, June 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900148.

Anonymous:1980:AIVa

[552] Anonymous. Author index for volume 20. Journal of Computer and System Sciences, 20(3): 405, June 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008090015X.

Lehmann:1980:AO

[553] Daniel J. Lehmann. On the algebra of order. Journal of Computer and System Sciences, 21(1): 1–23, August 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900380.

Mavrides:1980:FSS

[554] Lazaros P. Mavrides. A finite step-size procedure for the F–W method. Journal of Computer and System Sciences, 21(1):24–29, August 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900392.

Seki:1980:DSS

[555] Shigeko Seki and Youichi Kobuchi. On DIL schemes simulated by D0L schemes. Journal of Computer and System Sciences, 21(1):30–45, August 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000080900409.

Angluin:1980:FPC

[556] Dana Angluin. Finding patterns common to a set of strings. Journal of Computer and System Sciences, 21 (1):46–62, August 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900410.

Fujishige:1980:EPG

[557] Satoru Fujishige. An efficient PQgraph algorithm for solving the graphrealization problem. Journal of Computer and System Sciences, 21(1): 63–86, August 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900422.

Head:1980:TMS

[558] Tom Head. A-transducers and the monotonicity of IL schemes. Journal of Computer and System Sciences, 21(1):87–91, August 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900434.

Jaffe:1980:ERP

[559] Jeffrey M. Jaffe. The equivalence of R.E. program schemes and data flow schemes. Journal of Computer and System Sciences, 21(1): 92–109, August 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900446.

Maurer:1980:CFG

[560] H. A. Maurer, A. Salomaa, and D. Wood. Context-free grammar forms with strict interpretations. Journal of Computer and System Sciences, 21 (1):110–135, August 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900458.

Ausiello:1980:SPR

[561] G. Ausiello, A. D’Atri, and M. Protasi. Structure preserving reductions among convex optimization problems. Journal of Computer and System Sciences, 21(1):136–153, August 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008090046X.

Snyder:1980:EF

[562] Lawrence Snyder, E. K. Blum, and Nicholas Pippenger. Editor’s foreword. Journal of Computer and System Sciences, 21(2): 155, October 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900318.

Chandra:1980:CQR

[563] Ashok K. Chandra and David Harel. Computable queries for relational data bases. Journal of Computer and System Sciences, 21(2): 156–178, October 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008090032X.

Ladner:1980:CPS

[564] Richard E. Ladner. The complexity of problems in systems of communicating sequential processes. Journal of Computer and System Sciences, 21 (2):179–194, October 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900331.

Sipser:1980:LBS

[565] Michael Sipser. Lower bounds on the size of sweeping automata. Journal of Computer and System Sciences, 21 (2):195–202, October 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000080900343.

Galil:1980:AMF

[566] Zvi Galil and Amnon Naamad. An O(EV log^{2} V) algorithm for the maximal flow problem. Journal of Computer and System Sciences, 21(2): 203–217, October 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900355.

Ruzzo:1980:TSB

[567] Walter L. Ruzzo. Tree-size bounded alternation. Journal of Computer and System Sciences, 21(2):218– 235, October 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900367.

Munro:1980:IDS

[568] J. Ian Munro and Hendra Suwanda. Implicit data structures for fast search and update. Journal of Computer and System Sciences, 21(2): 236–250, October 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900379.

Lynch:1980:SLP

[569] Nancy A. Lynch. Straight-line program length as a parameter for complexity analysis. Journal of Computer and System Sciences, 21(3): 251–280, December 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900240.

Naude:1980:DTD

[570] G. Naudé. A duality theory for decomposable systems in a category. Journal of Computer and System Sciences, 21 (3):281–291, December 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900252.

Brandenburg:1980:MES

[571] Franz-Josef Brandenburg. Multiple equality sets and post machines. Journal of Computer and System Sciences, 21(3):292–316, December 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000080900264.

Lewis:1980:CRC

[572] Harry R. Lewis. Complexity results for classes of quantificational formulas. Journal of Computer and System Sciences, 21(3):317– 353, December 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900276.

Gonnet:1980:OAH

[573] Gaston H. Gonnet. Open-addressing hashing with unequal-probability keys. Journal of Computer and System Sciences, 21(3):354–367, December 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000080900288.

Maier:1980:ADR

[574] David Maier, Alberto O. Mendelzon, Fereidoon Sadri, and Jeffrey Dullman. Adequacy of decompositions of relational databases. Journal of Computer and System Sciences, 21(3): 368–379, December 1980. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008090029X.

Anonymous:1980:AIVb

[575] Anonymous. Author index for volume 21. Journal of Computer and System Sciences, 21(3): 380, December 1980. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000080900306.

Moran:1981:SRR

[576] Shlomo Moran. Some results on relativized deterministic and nondeterministic time hierarchies. Journal of Computer and System Sciences, 22(1):1–8, February 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900179.

Nakamura:1981:RFS

[577] Tsuyoshi Nakamura. Regularity and firing sequences of computation graphs. Journal of Computer and System Sciences, 22(1):9– 28, February 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900180.

Willson:1981:GPO

[578] Stephen J. Willson. Growth patterns of ordered cellular automata. Journal of Computer and System Sciences, 22(1):29–41, February 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900192.

Ruohonen:1981:DDD

[579] Keijo Ruohonen. The decidability of the DOL-DTOL equivalence problem. Journal of Computer and System Sciences, 22(1):42– 52, February 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900209.

Chew:1981:NSL

[580] Paul Chew and Michael Machtey. A note on structure and looking back applied to the relative complexity of computable functions. Journal of Computer and System Sciences, 22 (1):53–59, February 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900210.

Ohtsuki:1981:MAG

[581] Tatsuo Ohtsuki, Hajimu Mori, Toshinobu Kashiwabara, and Toshio Fujisawa. On minimal augmentation of a graph to obtain an interval graph. Journal of Computer and System Sciences, 22(1):60– 97, February 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900222.

Melville:1981:ISR

[582] Robert Melville. An improved simulation result for ink-bounded Turing machines. Journal of Computer and System Sciences, 22(1): 98–105, February 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900234.

Riccardi:1981:ICS

[583] Gregory A. Riccardi. The independence of control structures in abstract programming systems. Journal of Computer and System Sciences, 22(2):107–143, April 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900246.

Janssens:1981:DPN

[584] D. Janssens and G. Rozenberg. Decision problems for node label controlled graph grammars. Journal of Computer and System Sciences, 22 (2):144–177, April 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900258.

Workman:1981:SPC

[585] David A. Workman. SR(s, k) parsers: a class of shift-reduce boundedcontext parsers. Journal of Computer and System Sciences, 22(2): 178–197, April 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008190026X.

Georgeff:1981:ITS

[586] M. P. Georgeff. Interdependent translation schemes. Journal of Computer and System Sciences, 22(2): 198–219, April 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900271.

Gurari:1981:CDP

[587] Eitan M. Gurari and Oscar H. Ibarra. The complexity of decision problems for finite-turn multicounter machines. Journal of Computer and System Sciences, 22(2): 220–229, April 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900283.

Savage:1981:ATT

[588] John E. Savage. Area-time tradeoffs for matrix multiplication and related problems in VLSI models. Journal of Computer and System Sciences, 22(2):230–242, April 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900295.

Baker:1981:ELL

[589] Theodore P. Baker. Extending lookahead for LR parsers. Journal of Computer and System Sciences, 22 (2):243–259, April 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900301.

Kosaraju:1981:EF

[590] S. Rao Kosaraju and Albert R. Meyer. Editor’s foreword. Journal of Computer and System Sciences, 22(3):261, June 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900313.

Meyer:1981:DOC

[591] Albert R. Meyer. Dedicated to our colleague and friend Michael Machtey. Journal of Computer and System Sciences, 22(3):262– 264, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900325.

Wegman:1981:NHF

[592] Mark N. Wegman and J. Lawrence Carter. New hash functions and their use in authentication and set equality. Journal of Computer and System Sciences, 22(3): 265–279, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900337.

Brassard:1981:TLT

[593] Gilles Brassard. A time-luck tradeoff in relativized cryptography. Journal of Computer and System Sciences, 22(3):280–311, June 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900349.

Paul:1981:TVS

[594] W. Paul and R. Reischuk. On time versus space II. Journal of Computer and System Sciences, 22 (3):312–327, June 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900350.

Kozen:1981:SPP

[595] Dexter Kozen. Semantics of probabilistic programs. Journal of Computer and System Sciences, 22(3): 328–350, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900362.

Borodin:1981:TST

[596] Allan Borodin, Michael J. Fischer, David G. Kirkpatrick, Nancy A. Lynch, and Martin Tompa. A timespace tradeoff for sorting on nonoblivious machines. Journal of Computer and System Sciences, 22(3): 351–364, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900374.

Ruzzo:1981:UCC

[597] Walter L. Ruzzo. On uniform circuit complexity. Journal of Computer and System Sciences, 22(3): 365–383, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900386.

Immerman:1981:NQB

[598] Neil Immerman. Number of quantifiers is better than number of tape cells. Journal of Computer and System Sciences, 22(3):384– 406, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900398.

Gabber:1981:ECL

[599] Ofer Gabber and Zvi Galil. Explicit constructions of linear-sized superconcentrators. Journal of Computer and System Sciences, 22(3): 407–420, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900404.

Simon:1981:DIU

[600] Janos Simon. Division in idealized unit cost RAMs. Journal of Computer and System Sciences, 22 (3):421–441, June 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900416.

Machtey:1981:RRV

[601] Michael Machtey and Paul Young. Remarks on recursion versus diagonalization and exponentially difficult problems. Journal of Computer and System Sciences, 22(3): 442–453, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900428.

Pippenger:1981:CCA

[602] Nicholas Pippenger. Computational complexity of algebraic functions. Journal of Computer and System Sciences, 22(3):454–470, June 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/002200008190043X.

See correction [953].

Anonymous:1981:AIVa

[603] Anonymous. Author index for volume 22. Journal of Computer and System Sciences, 22(3): 471, June 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900441.

Maurer:1981:CCF

[604] H. A. Maurer, A. Salomaa, and D. Wood. Completeness of contextfree grammar forms. Journal of Computer and System Sciences, 23 (1):1–10, August 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900015.

Huet:1981:CPC

[605] Gérard Huet. A complete proof of correctness of the Knuth–Bendix completion algorithm. Journal of Computer and System Sciences, 23 (1):11–21, August 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900027.

Nakamura:1981:SAT

[606] Katsuhiko Nakamura. Synchronous to asynchronous transformation of polyautomata. Journal of Computer and System Sciences, 23(1): 22–37, August 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900039.

Werschulz:1981:MOL

[607] Arthur G. Werschulz. On maximal order for local and global numerical problems. Journal of Computer and System Sciences, 23(1): 38–48, August 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900040.

Back:1981:CRP

[608] R. J. R. Back. On correct refinement of programs. Journal of Computer and System Sciences, 23 (1):49–68, August 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900052.

Gallier:1981:RCA

[609] Jean H. Gallier. Recursion-closed algebraic theories. Journal of Computer and System Sciences, 23(1): 69–105, August 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900064.

Tompa:1981:C

[610] Martin Tompa. Corrigendum. Journal of Computer and System Sciences, 23(1):106, August 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900076.

Savitch:1981:EF

[611] Walter J. Savitch, B. S. Baker, and R. E. Miller. Editor’s foreword. Journal of Computer and System Sciences, 23(2):107, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900088.

Paul:1981:ITA

[612] Wolfgang J. Paul, Joel I. Seiferas, and Janos Simon. An informationtheoretic approach to time bounds for on-line computation. Journal of Computer and System Sciences, 23 (2):108–126, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008190009X.

Tompa:1981:OSW

[613] Martin Tompa. An optimal solution to a wire-routing problem. Journal of Computer and System Sciences, 23 (2):127–150, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900106.

Pippenger:1981:PAP

[614] Nicholas Pippenger. Pebbling with an auxiliary pushdown. Journal of Computer and System Sciences, 23 (2):151–165, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900118.

Overmars:1981:MCP

[615] Mark H. Overmars and Jan van Leeuwen. Maintenance of configurations in the plane. Journal of Computer and System Sciences, 23 (2):166–204, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008190012X.

Joseph:1981:IRC

[616] Deborah Joseph and Paul Young. Independence results in computer science? Journal of Computer and System Sciences, 23(2):205– 222, October 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900131.

Ehrig:1981:CAI

[617] Hartmut Ehrig and Bernd Mahr. Complexity of algebraic implementations for abstract data types. Journal of Computer and System Sciences, 23 (2):223–253, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900143.

Lynch:1981:UBS

[618] Nancy A. Lynch. Upper bounds for static resource allocation in a distributed system. Journal of Computer and System Sciences, 23(2): 254–278, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900155.

Meyer:1981:DDL

[619] Albert R. Meyer and Rohit Parikh. Definability in dynamic logic. Journal of Computer and System Sciences, 23 (2):279–298, October 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900167.

Valk:1981:PNR

[620] Rüdiger Valk and Guy Vidal-Naquet. Petri nets and regular languages. Journal of Computer and System Sciences, 23(3):299–325, December 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000081900672.

Selman:1981:SON

[621] Alan L. Selman. Some observations on NP real numbers and Pselective sets. Journal of Computer and System Sciences, 23(3): 326–332, December 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900684.

Snyder:1981:TCT

[622] Lawrence Snyder. Theft and conspiracy in the take-grant protection model. Journal of Computer and System Sciences, 23(3): 333–347, December 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900696.

Leiss:1981:CRR

[623] Ernst Leiss. The complexity of restricted regular expressions and the synthesis problem for finite automata. Journal of Computer and System Sciences, 23(3):348– 354, December 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900702.

Fredman:1981:QTV

[624] Michael L. Fredman and Dennis J. Volper. Query time versus redundancy trade-offs for range queries. Journal of Computer and System Sciences, 23 (3):355–365, December 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000081900714.

Oyamaguchi:1981:EPT

[625] Michio Oyamaguchi, Yasuyoshi Inagaki, and Namio Honda. The equivalence problem for two dpda’s, one of which is a finite-turn or onecounter machine. Journal of Computer and System Sciences, 23(3): 366–382, December 1981. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900726.

Aiello:1981:EIL

[626] Luigia Aiello and Gianfranco Print. An efficient interpreter for the lambdacalculus. Journal of Computer and System Sciences, 23(3):383– 424, December 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000081900738.

Anonymous:1981:AIVb

[627] Anonymous. Author index for volume 23. Journal of Computer and System Sciences, 23(3): 425, December 1981. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008190074X.

Laskowski:1982:CLB

[628] Sharon J. Laskowski. Computing lower bounds on tensor rank over finite fields. Journal of Computer and System Sciences, 24(1):1– 14, February 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900526.

Ko:1982:MVP

[629] Ker-I Ko. The maximum value problem and NP real numbers. Journal of Computer and System Sciences, 24 (1):15–35, February 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900538.

Pittl:1982:LP

[630] Jan Pittl. On LLP(k) parsers. Journal of Computer and System Sciences, 24(1):36–68, February 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008290054X.

Ibarra:1982:SDP

[631] Oscar H. Ibarra and Shlomo Moran. On some decision problems for RAM programs. Journal of Computer and System Sciences, 24(1):69– 81, February 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900551.

Pippenger:1982:SD

[632] Nicholas Pippenger. Superconcentrators of depth 2. Journal of Computer and System Sciences, 24(1):82– 90, February 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900563.

Yehudai:1982:HRT

[633] Amiram Yehudai. A hierarchy of real-time deterministic languages and their equivalence. Journal of Computer and System Sciences, 24(1): 91–100, February 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900575.

Lynch:1982:AVD

[634] Nancy A. Lynch. Accessibility of values as a determinant of relative complexity in algebras. Journal of Computer and System Sciences, 24 (1):101–113, February 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900587.

Adamek:1982:TCF

[635] Jiri Adamek, Evelyn Nelson, and Jan Reiterman. Tree constructions of free continuous algebras. Journal of Computer and System Sciences, 24 (1):114–146, February 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900599.

Anonymous:1982:A

[636] Anonymous. Announcement. Journal of Computer and System Sciences, 24(1):147, February 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900605.

Nijholt:1982:EPL

[637] Anton Nijholt. The equivalence problem for LL- and LR-regular grammars. Journal of Computer and System Sciences, 24(2):149– 161, April 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900447.

Rodeh:1982:FMD

[638] Michael Rodeh. Finding the median distributively. Journal of Computer and System Sciences, 24(2): 162–166, April 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900459.

Albert:1982:TCP

[639] J. Albert and K. Culik. Tree correspondence problems. Journal of Computer and System Sciences, 24 (2):167–179, April 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900460.

Goldman:1982:QA

[640] Jerry Goldman and Steven Homer. Quadratic automata. Journal of Computer and System Sciences, 24 (2):180–196, April 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900472.

Frederickson:1982:CSR

[641] Greg N. Frederickson and Donald B. Johnson. The complexity of selection and ranking in X + Y and matrices with sorted columns. Journal of Computer and System Sciences, 24(2):197–208, April 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900484.

JaJa:1982:CCS

[642] Joseph Ja’Ja’. The computational complexity of a set of quadratic functions. Journal of Computer and System Sciences, 24(2):209– 223, April 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900496.

Long:1982:NSO

[643] Timothy J. Long. A note on sparse oracles for NP. Journal of Computer and System Sciences, 24(2): 224–232, April 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900502.

Hashiguchi:1982:LTF

[644] K. Hashiguchi. Limitedness theorem on finite automata with distance functions. Journal of Computer and System Sciences, 24(2): 233–244, April 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900514.

Maggiolo-Schettini:1982:PTS

[645] Andrea Maggiolo-Schettini and Jozef Winkowski. Processes of transforming structures. Journal of Computer and System Sciences, 24(3): 245–282, June 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900290.

Engelfriet:1982:SMV

[646] Joost Engelfriet and Gilberto Filé. Simple multi-visit attribute grammars. Journal of Computer and System Sciences, 24(3):283–314, June 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000082900307.

Ginsburg:1982:PDT

[647] Seymour Ginsburg, Jonathan Goldstine, and Edwin H. Spanier. A prime decomposition theorem for grammatical families. Journal of Computer and System Sciences, 24 (3):315–361, June 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900319.

Machida:1982:SCL

[648] Hajime Machida and Takumi Kasai. Space complexity in on-line computation. Journal of Computer and System Sciences, 24(3): 362–372, June 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900320.

Thuraisingham:1982:ROO

[649] M. B. Thuraisingham. Representation of one-one degrees by decision problems for system functions. Journal of Computer and System Sciences, 24(3):373–377, June 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900332.

Joseph:1982:C

[650] Deborah Joseph and Paul Young. Corrigendum. Journal of Computer and System Sciences, 24(3): 378, June 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900344.

Anonymous:1982:AIVa

[651] Anonymous. Author index for volume 24. Journal of Computer and System Sciences, 24(3): 379, June 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900356.

Selman:1982:EFa

[652] Alan L. Selman. Editor’s foreword. Journal of Computer and System Sciences, 25(1):1, August 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900071.

Yannakakis:1982:AD

[653] Mihalis Yannakakis and Christos H. Papadimitriou. Algebraic dependencies. Journal of Computer and System Sciences, 25(1):2–41, August 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000082900083.

Luks:1982:IGB

[654] Eugene M. Luks. Isomorphism of graphs of bounded valence can be tested in polynomial time. Journal of Computer and System Sciences, 25(1):42–65, August 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900095.

Rabin:1982:PME

[655] Michael O. Rabin. N-process mutual exclusion with bounded waiting by 4·Log_{2}N-valued shared variable. Journal of Computer and System Sciences, 25(1):66–75, August 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900101.

Immerman:1982:ULB

[656] Neil Immerman. Upper and lower bounds for first order expressibility. Journal of Computer and System Sciences, 25(1):76–98, August 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000082900113.

Chandra:1982:SCR

[657] Ashok Chandra and David Harel. Structure and complexity of relational queries. Journal of Computer and System Sciences, 25(1): 99–128, August 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900125.

Selman:1982:EFb

[658] Alan L. Selman. Editor’s foreword. Journal of Computer and System Sciences, 25(2):129, October 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000082900010.

Mahaney:1982:SCS

[659] Stephen R. Mahaney. Sparse complete sets for NP: Solution of a conjecture of Berman and Hartmanis. Journal of Computer and System Sciences, 25 (2):130–143, October 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900022.

Harel:1982:PLE

[660] David Harel, Dexter Kozen, and Rohit Parikh. Process logic: Expressiveness, decidability, completeness. Journal of Computer and System Sciences, 25(2):144–170, October 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900034.

Flajolet:1982:AHB

[661] Philippe Flajolet and Andrew Odlyzko. The average height of binary trees and other simple trees. Journal of Computer and System Sciences, 25 (2):171–213, October 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900046.

Bentley:1982:GCR

[662] Jon Louis Bentley and Donna J. Brown. A general class of resource tradeoffs. Journal of Computer and System Sciences, 25(2): 214–238, October 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900058.

Huet:1982:PIE

[663] Gerard Huet and Jean-Marie Hullot. Proofs by induction in equational theories with constructors. Journal of Computer and System Sciences, 25 (2):239–266, October 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008290006X.

Bergstra:1982:ECH

[664] J. A. Bergstra and J. V. Tucker. Expressiveness and the completeness of Hoare’s logic. Journal of Computer and System Sciences, 25(3): 267–284, December 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900137.

Kasai:1982:HBM

[665] Takumi Kasai and Raymond E. Miller. Homomorphisms between models of parallel computation. Journal of Computer and System Sciences, 25(3): 285–331, December 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900149.

Hoffmann:1982:SGG

[666] Christoph M. Hoffmann. Subcomplete generalizations of graph isomorphism. Journal of Computer and System Sciences, 25(3):332– 359, December 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900150.

Thomas:1982:CRE

[667] Wolfgang Thomas. Classifying regular events in symbolic logic. Journal of Computer and System Sciences, 25 (3):360–376, December 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900162.

Kuich:1982:ACS

[668] W. Kuich. An algebraic characterization of some principal regulated rational cones. Journal of Computer and System Sciences, 25(3): 377–401, December 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900174.

Ben-Ari:1982:DPD

[669] Mordechai Ben-Ari, Joseph Y. Halpern, and Amir Pnueli. Deterministic propositional dynamic logic: Finite models, complexity, and completeness. Journal of Computer and System Sciences, 25 (3):402–417, December 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000082900186.

Engelfriet:1982:CPO

[670] Joost Engelfriet and Sven Skyum. The copying power of one-state tree transducers. Journal of Computer and System Sciences, 25(3): 418–435, December 1982. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900198.

Anonymous:1982:AIVb

[671] Anonymous. Author index for volume 25. Journal of Computer and System Sciences, 25(3): 436, December 1982. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000082900204.

Comer:1983:EUO

[672] Douglas Comer. Effects of updates on optimality in tries. Journal of Computer and System Sciences, 26 (1):1–13, February 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008390017X.

Latteux:1983:LCF

[673] Michel Latteux. Langages à un compteur. (French) [One-counter languages]. Journal of Computer and System Sciences, 26(1):14–33, February 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900181.

Ehrenfeucht:1983:CFN

[674] A. Ehrenfeucht, J. Engelfriet, and G. Rozenberg. Context free normal systems and ETOL systems. Journal of Computer and System Sciences, 26(1):34–46, February 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900193.

Ibarra:1983:ZIP

[675] Oscar H. Ibarra and Brian S. Leininger. On the zero-inequivalence problem far loop programs. Journal of Computer and System Sciences, 26(1):47– 64, February 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008390020X.

Carlson:1983:SST

[676] David A. Carlson and John E. Savage. Size-space tradeoffs for oblivious computations. Journal of Computer and System Sciences, 26(1):65– 81, February 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900211.

Albert:1983:CST

[677] Jürgen Albert and Derick Wood. Checking sets, test sets, rich languages and commutatively closed languages. Journal of Computer and System Sciences, 26(1):82–91, February 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900223.

Riffelmacher:1983:MCA

[678] Dave Riffelmacher. Multiplicative complexity and algebraic structure. Journal of Computer and System Sciences, 26(1):92–106, February 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900235.

Watanabe:1983:TSF

[679] Toshimasa Watanabe and Akira Nakamura. On the transformation semigroups of finite automata. Journal of Computer and System Sciences, 26 (1):107–138, February 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900247.

Grefenstette:1983:NSF

[680] John J. Grefenstette. Network structure and the firing squad synchronization problem. Journal of Computer and System Sciences, 26(1): 139–152, February 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900259.

Ukkonen:1983:LBS

[681] Esko Ukkonen. Lower bounds on the size of deterministic parsers. Journal of Computer and System Sciences, 26(2):153–170, April 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900107.

Ginsburg:1983:EGF

[682] Seymour Ginsburg, Jonathan Goldstine, and Edwin H. Spanier. On the equality of grammatical families. Journal of Computer and System Sciences, 26(2):171–196, April 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900119.

Asano:1983:ECP

[683] Takao Asano and Tomio Hirata. Edgecontraction problems. Journal of Computer and System Sciences, 26 (2):197–208, April 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900120.

Ko:1983:SRW

[684] Ker-I Ko. On self-reducibility and weak P-selectivity. Journal of Computer and System Sciences, 26(2): 209–221, April 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900132.

Harel:1983:PDL

[685] David Harel, Amir Pnueli, and Jonathan Stavi. Propositional dynamic logic of nonregular programs. Journal of Computer and System Sciences, 26(2):222–243, April 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900144.

Sippu:1983:CLK

[686] Seppo Sippu and Eljas Soisalon-Soininen. On the complexity of LL(k) testing. Journal of Computer and System Sciences, 26(2): 244–268, April 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900156.

Tchuente:1983:CBF

[687] Maurice Tchuente. Computation of Boolean functions on networks of binary automata. Journal of Computer and System Sciences, 26(2): 269–277, April 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900168.

Demillo:1983:EF

[688] Richard Demillo, Ronald Rivest, and E. K. Blum. Editors’ foreword. Journal of Computer and System Sciences, 26(3):279, June 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900016.

Galil:1983:TSO

[689] Zvi Galil and Joel Seiferas. Timespace-optimal string matching. Journal of Computer and System Sciences, 26(3):280–294, June 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900028.

Muller:1983:GTE

[690] David E. Muller and Paul E. Schupp. Groups, the theory of ends, and context-free languages. Journal of Computer and System Sciences, 26 (3):295–310, June 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008390003X.

Joseph:1983:PTC

[691] Deborah Joseph. Polynomial time computations in models of ET. Journal of Computer and System Sciences, 26(3):311–338, June 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900041.

Kleitman:1983:AOL

[692] Daniel Kleitman, Frank Thomson Leighton, Margaret Lepley, and Gary L. Miller. An asymptotically optimal layout for the shuffleexchange graph. Journal of Computer and System Sciences, 26(3): 339–361, June 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900053.

Sleator:1983:DSD

[693] Daniel D. Sleator and Robert Endre Tarjan. A data structure for dynamic trees. Journal of Computer and System Sciences, 26(3):

362–391, June 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900065.

Fich:1983:LBC

[694] Faith E. Fich. Lower bounds for the cycle detection problem. Journal of Computer and System Sciences, 26(3):392–409, June 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900077.

Venkataraman:1983:VCT

[695] K. N. Venkataraman, Ann Yasuhara, and Frank M. Hawrusik. A view of computability on term algebras. Journal of Computer and System Sciences, 26(3):410–471, June 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900089.

Anonymous:1983:AIVa

[696] Anonymous. Author index for volume 26. Journal of Computer and System Sciences, 26(3): 472, June 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900090.

Heilbrunner:1983:TLL

[697] Stephan Heilbrunner. Tests for the LR- , LL-, and LC-regular conditions. Journal of Computer and System Sciences, 27(1):1–13, August 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900260.

Schoning:1983:LHH

[698] Uwe Schöning. A low and a high hierarchy within NP. Journal of Computer and System Sciences, 27 (1):14–28, August 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900272.

Steenstrup:1983:PAA

[699] Martha Steenstrup, Michael A. Arbib, and Ernest G. Manes. Port automata and the algebra of concurrent processes. Journal of Computer and System Sciences, 27(1): 29–50, August 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900284.

Suzuki:1983:MSR

[700] Ichiro Suzuki and Tadao Murata. A method for stepwise refinement and abstraction of Petri nets. Journal of Computer and System Sciences, 27(1):51–76, August 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900296.

Storer:1983:CC

[701] James A. Storer. On the complexity of chess. Journal of Computer and System Sciences, 27(1): 77–100, August 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900302.

Hashiguchi:1983:RTR

[702] K. Hashiguchi. Representation theorems on regular languages. Journal of Computer and System Sciences, 27 (1):101–115, August 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900314.

Miyano:1983:RMP

[703] Satoru Miyano. Remarks on multihead pushdown automata and multihead stack automata. Journal of Computer and System Sciences, 27 (1):116–124, August 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900326.

Ito:1983:LCA

[704] Masanobu Itô, Nobuyasu Ôsato, and Masakazu Nasu. Linear cellular automata over Z_{m}. Journal of Computer and System Sciences, 27(1): 125–140, August 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900338.

Shepherdson:1983:CCE

[705] J. C. Shepherdson. Calvin C. Elgot 1922–1980. Journal of Computer and System Sciences, 27(2): 141–144, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008390034X.

Miller:1983:R

[706] R. E. Miller. Reminiscences 1. Journal of Computer and System Sciences, 27(2):145, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900351.

Blum:1983:R

[707] E. K. Blum. Reminiscences 2. Journal of Computer and System Sciences, 27 (2):146–147, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900363.

Bloom:1983:RIC

[708] Stephen L. Bloom, James W. Thatcher, Eric G. Wagner, and Jesse B. Wright. Recursion and iteration in continuous theories: The “M-construction”. Journal of Computer and System Sciences, 27(2):148–164, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900375.

Schmeck:1983:ACR

[709] Hartmut Schmeck. Algebraic characterization of reducible flowcharts. Journal of Computer and System Sciences, 27(2):165–199, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900387.

Parisi-Presicce:1983:EIC

[710] Francesco Parisi-Presicce. Extensions of iterative congruences on free iterative algebras. Journal of Computer and System Sciences, 27(2): 200–220, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900399.

Troeger:1983:ADS

[711] Douglas R. Troeger. An axiomatization of D-scheme strong equivalence. Journal of Computer and System Sciences, 27(2):221–224, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900405.

Bloom:1983:ASS

[712] Stephen L. Bloom. All solutions of a system of recursion equations in infinite trees and other contraction theories. Journal of Computer and System Sciences, 27(2): 225–255, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900417.

Rabin:1983:TPB

[713] Michael O. Rabin. Transaction protection by beacons. Journal of Computer and System Sciences, 27(2): 256–267, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900429.

Goguen:1983:CRP

[714] J. A. Goguen and J. Meseguer. Correctness of recursive parallel nondeterministic flow programs. Journal of Computer and System Sciences, 27 (2):268–290, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900430.

Esik:1983:AIT

[715] Z. Ésik. Algebras of iteration theories. Journal of Computer and System Sciences, 27(2):291– 303, October 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900442.

Blum:1983:IDT

[716] Edward K. Blum and Francesco Parisi-Presicce. Implementation of data types by algebraic methods. Journal of Computer and System Sciences, 27 (2):304–330, October 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900454.

Hull:1983:AJD

[717] Richard Hull. Acyclic join dependency and data base projections. Journal of Computer and System Sciences, 27 (3):331–349, December 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900466.

Lynch:1983:TDA

[718] Nancy A. Lynch and Michael J. Fischer. A technique for decomposing algorithms which use a single shared variable. Journal of Computer and System Sciences, 27(3): 350–377, December 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900478.

Hodgson:1983:NFA

[719] Bernard R. Hodgson and Clement F. Kent. A normal form for arithmetical representation of NP-sets. Journal of Computer and System Sciences, 27 (3):378–388, December 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008390048X.

Astesiano:1983:ITN

[720] Egidio Astesiano and Gerardo Costa. The insensitivity theorem for nonreducing reflexive types. Journal of Computer and System Sciences, 27(3): 389–399, December 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000083900491.

Lorentz:1983:DNF

[721] Richard J. Lorentz and David B. Benson. Deterministic and nondeterministic flowchart interpretations. Journal of Computer and System Sciences, 27 (3):400–433, December 1983. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900508.

Staiger:1983:FSL

[722] Ludwig Staiger. Finite-state ωlanguages. Journal of Computer and System Sciences, 27(3):434– 448, December 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008390051X.

Anonymous:1983:AIVb

[723] Anonymous. Author index for volume 27. Journal of Computer and System Sciences, 27(3): 449, December 1983. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000083900521.

Aho:1984:EF

[724] Alfred V. Aho and E. K. Blum. Editors’ foreword. Journal of Computer and System Sciences, 28(1): 1, February 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900722.

Klug:1984:O

[725] Anthony C. Klug and Alfred V. Aho. Obituary. Journal of Computer and System Sciences, 28(1): 2, February 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900734.

Vardi:1984:IFI

[726] Moshe Y. Vardi. The implication and finite implication problems for typed template dependencies. Journal of Computer and System Sciences, 28(1):3–28, February 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900746.

Casanova:1984:IDT

[727] Marco A. Casanova, Ronald Fagin, and Christos H. Papadimitriou. Inclusion dependencies and their interaction with functional dependencies. Journal of Computer and System Sciences, 28(1):29–59, February 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900758.

Goodman:1984:TPT

[728] Nathan Goodman and Oded Shmueli. The tree projection theorem and relational query processing. Journal of Computer and System Sciences, 28 (1):60–79, February 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008490076X.

Imielinski:1984:RMD

[729] Tomasz Imieliński and Witold Lipski. The relational model of data and cylindric algebras. Journal of Computer and System Sciences, 28 (1):80–102, February 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900771.

Kanellakis:1984:DLH

[730] Paris C. Kanellakis and Christos H. Papadimitriou. Is distributed locking harder? Journal of Computer and System Sciences, 28(1): 103–120, February 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900783.

Graham:1984:IDS

[731] Marc H. Graham and Mihalis Yannakakis. Independent database schemas. Journal of Computer and System Sciences, 28(1):121– 141, February 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900795.

Zaniolo:1984:DRN

[732] Carlo Zaniolo. Database relations with null values. Journal of Computer and System Sciences, 28(1): 142–166, February 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900801.

Johnson:1984:TCC

[733] D. S. Johnson and A. Klug. Testing containment of conjunctive queries under functional and inclusion dependencies. Journal of Computer and System Sciences, 28(1):167– 189, February 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900813.

Lewis:1984:EF

[734] Harry R. A. Lewis. Editor’s foreword. Journal of Computer and System Sciences, 28(2):191, April 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900643.

Feldman:1984:PDL

[735] Yishai A. Feldman and David Harel. A probabilistic dynamic logic. Journal of Computer and System Sciences, 28(2):193–215, April 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900655.

Ruzzo:1984:SBH

[736] Walter L. Ruzzo, Janos Simon, and Martin Tompa. Space-bounded hierarchies and probabilistic computations. Journal of Computer and System Sciences, 28(2):216– 230, April 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900667.

Furer:1984:DSD

[737] Martin Fürer. Data structures for distributed counting. Journal of Computer and System Sciences, 28 (2):231–243, April 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900679.

Papadimitriou:1984:CFS

[738] C. H. Papadimitriou and M. Yannakakis. The complexity of facets (and some facets of complexity). Journal of Computer and System Sciences, 28(2):244–259, April 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900680.

Papadimitriou:1984:CC

[739] Christos H. Papadimitriou and Michael Sipser. Communication complexity. Journal of Computer and System Sciences, 28(2):260–269, April 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000084900692.

Goldwasser:1984:PE

[740] Shafi Goldwasser and Silvio Micali. Probabilistic encryption. Journal of Computer and System Sciences, 28 (2):270–299, April 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900709. See

also preliminary version in 14th STOC, 1982.

Bhatt:1984:FSV

[741] Sandeep N. Bhatt and Frank Thomson Leighton. A framework for solving VLSI graph layout problems. Journal of Computer and System Sciences, 28(2):300–343, April 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900710.

Warmuth:1984:CIS

[742] Manfred K. Warmuth and David Haussler. On the complexity of iterated shuffle. Journal of Computer and System Sciences, 28(3): 345–358, June 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900187.

Low:1984:MAP

[743] Michael C. Low. Minimizing access pointers into trees and arrays. Journal of Computer and System Sciences, 28(3):359–378, June 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900199.

Willard:1984:NTD

[744] Dan E. Willard. New trie data structures which support very fast search operations. Journal of Computer and System Sciences, 28(3): 379–394, June 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900205.

Hunt:1984:CMR

[745] H. B. Hunt and D. J. Rosenkrantz. The complexity of monadic recursion schemes: exponential time bounds. Journal of Computer and System Sciences, 28(3):395– 419, June 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900217.

Lloyd:1984:OLR

[746] Errol L. Lloyd and S. S. Ravi. One-layer routing without component constraints. Journal of Computer and System Sciences, 28(3): 420–438, June 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900229.

Milner:1984:CIS

[747] Robin Milner. A complete inference system for a class of regular behaviours. Journal of Computer and System Sciences, 28(3): 439–466, June 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900230.

Anonymous:1984:AIVa

[748] Anonymous. Author index for volume 28. Journal of Computer and System Sciences, 28(3): 467, June 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900242.

Devroye:1984:EBR

[749] Luc Devroye. Exponential bounds for the running time of a selection algorithm. Journal of Computer and System Sciences, 29(1):

1–7, August 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900096.

Plaisted:1984:CPF

[750] David A. Plaisted. Complete problems in the first-order predicate calculus. Journal of Computer and System Sciences, 29(1): 8–35, August 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900102.

Sugihara:1984:ADC

[751] Kokichi Sugihara. An n log n algorithm for determining the congruity of polyhedra. Journal of Computer and System Sciences, 29 (1):36–47, August 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900114.

Pachl:1984:FPN

[752] Jan K. Pachl. Finding pseudoperipheral nodes in graphs. Journal of Computer and System Sciences, 29 (1):48–53, August 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900126.

Latteux:1984:COC

[753] M. Latteux and G. Rozenberg. Commutative one-counter languages are regular. Journal of Computer and System Sciences, 29(1):

54–57, August 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900138.

Chen:1984:FNR

[754] Tsong Yueh Chen. On the fixpoints of nondeterministic recursive definitions. Journal of Computer and System Sciences, 29(1): 58–79, August 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008490014X.

Kroger:1984:GNO

[755] F. Kröger. A generalized nexttime operator in temporal logic. Journal of Computer and System Sciences, 29(1):80–98, August 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900151.

Troeger:1984:WEC

[756] Douglas R. Troeger. Weak equivalence in a class of structured program schemes. Journal of Computer and System Sciences, 29(1): 99–117, August 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000084900163.

Savitch:1984:CNS

[757] Walter J. Savitch and Patrick W. Dymond. Consistency in nondeterministic storage. Journal of Computer and System Sciences, 29(1):

118–132, August 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900175.

Hsu:1984:NAL

[758] W. J. Hsu and M. W. Du. New algorithms for the LCS problem. Journal of Computer and System Sciences, 29 (2):133–152, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900254.

Restivo:1984:CPL

[759] Antonio Restivo and Christophe Reutenauer. On cancellation properties of languages which are supports of rational power series. Journal of Computer and System Sciences, 29 (2):153–159, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900266.

Meyer:1984:EAL

[760] Albert R. Meyer and Jerzy Tiuryn. Equivalences among logics of programs. Journal of Computer and System Sciences, 29(2):160–170, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000084900278.

Miyajima:1984:IMT

[761] Hiromi Miyajima, Masateru Harao, and Shoichi Noguchi. Indecomposable maps in tessellation structures of arbitrary dimension. Journal of Computer and System Sciences, 29 (2):171–182, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008490028X.

Yesha:1984:TST

[762] Yaacov Yesha. Time-space tradeoffs for matrix multiplication and the Discrete Fourier Transform on any general sequential random-access computer. Journal of Computer and System Sciences, 29(2):183– 197, October 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900291.

Porat:1984:THN

[763] Sara Porat and Michael Yoeli. Towards a hierarchy of nets. Journal of Computer and System Sciences, 29 (2):198–206, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900308.

Muro:1984:MVC

[764] Shojiro Muro, Tiko Kameda, and Toshimi Minoura. Multi-version concurrency control scheme for a database system. Journal of Computer and System Sciences, 29(2): 207–224, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008490031X.

Ntafos:1984:CCP

[765] Simeon Ntafos and Teofilo Gonzalez. On the computational complexity of path cover problems. Journal of Computer and System Sciences, 29 (2):225–242, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900321.

Savage:1984:PMV

[766] John E. Savage. The performance of multilective VLSI algorithms. Journal of Computer and System Sciences, 29 (2):243–273, October 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900333.

Reif:1984:CTP

[767] John H. Reif. The complexity of two-player games of incomplete information. Journal of Computer and System Sciences, 29(2):274– 301, October 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900345.

Vitanyi:1984:TTR

[768] Paul M. B. Vitányi. On two-tape realtime computation and queues. Journal of Computer and System Sciences, 29 (3):303–311, December 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900011.

Motoki:1984:ACR

[769] Tatsuya Motoki. An application of Cohen’s result on star height to the theory of control structures. Journal of Computer and System Sciences, 29 (3):312–329, December 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900023.

Atallah:1984:FET

[770] Mikhail Atallah and Uzi Vishkin. Finding Euler tours in parallel. Journal of Computer and System Sciences, 29 (3):330–337, December 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900035.

Goodman:1984:GRC

[771] Nathan Goodman, Oded Shmueli, and Y. C. Tay. GYO reductions, canonical connections, tree and cyclic schemas, and tree projections. Journal of Computer and System Sciences, 29(3): 338–358, December 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900047.

Harel:1984:UP

[772] D. Harel and M. S. Paterson. Undecidability of PDL with L = {a^{2i}|; i ≥ 0}. Journal of Computer and System Sciences, 29(3): 359–365, December 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900059.

Engelfriet:1984:EMG

[773] Joost Engelfriet and Giora Slutzki. Extended macro grammars and stack controlled machines. Journal of Computer and System Sciences, 29(3): 366–408, December 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900060.

Yu:1984:ODT

[774] Clement T. Yu, Z. Meral Ozsoyoglu, and K. Lam. Optimization of distributed tree queries. Journal of Computer and System Sciences, 29(3): 409–445, December 1984. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900072.

Anonymous:1984:AIVb

[775] Anonymous. Author index for volume 29. Journal of Computer and System Sciences, 29(3): 446, December 1984. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000084900084.

Emerson:1985:DPE

[776] E. Allen Emerson and Joseph Y. Halpern. Decision procedures and expressiveness in the temporal logic of branching time. Journal of Computer and System Sciences, 30(1):1– 24, February 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900017.

Schimpf:1985:TPA

[777] Karl M. Schimpf and Jean H. Gallier. Tree pushdown automata. Journal of Computer and System Sciences, 30(1):25–40, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900029.

Reif:1985:MNL

[778] John Reif and A. P. Sistla. A multiprocess network logic with temporal and spatial modalities. Journal of Computer and System Sciences, 30 (1):41–53, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900030.

Chiba:1985:LAE

[779] Norishige Chiba, Takao Nishizeki, Shigenobu Abe, and Takao Ozawa. A linear algorithm for embedding planar graphs using PQ-trees. Journal of Computer and System Sciences, 30 (1):54–76, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900042.

Chrobak:1985:VTD

[780] Marek Chrobak. Variations on the technique of Duris and Galil. Journal of Computer and System Sciences, 30(1):77–85, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900054.

Engelfriet:1985:HHA

[781] Joost Engelfriet. Hierarchies of hyper-AFLs. Journal of Computer and System Sciences, 30(1): 86–115, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900066.

Ladner:1985:SA

[782] Richard E. Ladner and Jeffrey K. Norman. Solitaire automata. Journal of Computer and System Sciences, 30 (1):116–129, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900078.

Borodin:1985:RMS

[783] A. Borodin and J. E. Hopcroft. Routing, merging, and sorting on parallel models of computation. Journal of Computer and System Sciences, 30 (1):130–145, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008590008X.

Anonymous:1985:A

[784] Anonymous. Announcement. Journal of Computer and System Sciences, 30(1):146, February 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900091.

Ruzzo:1985:EF

[785] Walter L. Ruzzo and Edward K. Blum. Editors’ foreword. Journal of Computer and System Sciences, 30(2):147, April 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900108.

Dymond:1985:SDM

[786] Patrick W. Dymond and Martin Tompa. Speedups of deterministic machines by synchronous parallel machines. Journal of Computer and System Sciences, 30(2): 149–161, April 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008590011X.

Kozen:1985:PP

[787] Dexter Kozen. A probabilistic PDL. Journal of Computer and System Sciences, 30(2):162–178, April 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000085900121.

Landau:1985:SRP

[788] Susan Landau and Gary Lee Miller. Solvability by radicals is in polynomial time. Journal of Computer and System Sciences, 30(2): 179–208, April 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900133.

Gabow:1985:LTA

[789] Harold N. Gabow and Robert Endre Tarjan. A linear-time algorithm for a special case of disjoint set union. Journal of Computer and System Sciences, 30(2):209–221, April 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900145.

Chandra:1985:UFC

[790] Ashok K. Chandra, Steven Fortune, and Richard Lipton. Unbounded fan-in circuits and associative functions. Journal of Computer and System Sciences, 30(2): 222–234, April 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900157.

Lenstra:1985:FMP

[791] Arjen K. Lenstra. Factoring multivariate polynomials over finite fields. Journal of Computer and System Sciences, 30(2):235–248, April 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900169.

Gonczarowski:1985:PSG

[792] J. Gonczarowski and E. Shamir. Pattern selector grammars and several parsing algorithms in the contextfree style. Journal of Computer and System Sciences, 30(3):249– 273, June 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900467.

Bahamondf:1985:CFO

[793] Antonio Bahamondf. Cost functors and optimal realization. Journal of Computer and System Sciences, 30 (3):274–290, June 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900479.

Aso:1985:DCL

[794] Hirotomo Aso and Namio Honda. Dynamical characteristics of linear cellular automata. Journal of Computer and System Sciences, 30(3): 291–317, June 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900480.

Main:1985:FSR

[795] Michael G. Main and David B. Benson. Free semiring-representations and nondeterminism. Journal of Computer and System Sciences, 30 (3):318–328, June 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900492.

Lin:1985:APT

[796] Ferng-Ching Lin and I-Chen Wu. Area-period tradeoffs for multiplication of rectangular matrices. Journal of Computer and System Sciences, 30(3):329–342, June 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900509.

Narendran:1985:CCD

[797] Paliath Narendran, Colm Ó’Dúnlaing, and Heinrich Rolletschek. Complexity of certain decision problems about congruential languages. Journal of Computer and System Sciences, 30 (3):343–358, June 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900510.

Kantor:1985:STP

[798] William M. Kantor. Sylow’s theorem in polynomial time. Journal of Computer and System Sciences, 30 (3):359–394, June 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900522.

Book:1985:QRC

[799] Ronald V. Book, Timothy J. Long, and Alan L. Selman. Qualitative relativizations of complexity classes. Journal of Computer and System Sciences, 30(3):395–413, June 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900534.

Anonymous:1985:AIVa

[800] Anonymous. Author index for volume 30. Journal of Computer and System Sciences, 30(3): 414, June 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900546.

Melhem:1985:FAS

[801] Rami Melhem. Formal analysis of a systolic system for finite element stiffness matrices. Journal of Computer and System Sciences, 31(1): 1–27, August 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900625.

Hromkovic:1985:PAA

[802] Juraj Hromkovic. On the power of alternation in automata theory. Journal of Computer and System Sciences, 31(1):28–39, August 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900637.

Kaliski:1985:BCN

[803] Martin E. Kaliski and Quentin L. Klein. Behavior of a class of nonlinear discrete-time systems. Journal of Computer and System Sciences, 31(1):40–62, August 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900649.

Atallah:1985:MPP

[804] Mikhail J. Atallah. A matching problem in the plane. Journal of Computer and System Sciences, 31 (1):63–70, August 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900650.

Engelfriet:1985:MTT

[805] Joost Engelfriet and Heiko Vogler. Macro tree transducers. Journal of Computer and System Sciences, 31 (1):71–146, August 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900662.

Snyder:1985:EF

[806] Lawrence Snyder. Editor’s foreword. Journal of Computer and System Sciences, 31(2):147, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000085900388.

Garbow:1985:SAN

[807] Harold N. Garbow. Scaling algorithms for network problems. Journal of Computer and System Sciences, 31 (2):148–168, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008590039X.

Wilson:1985:RCC

[808] Christopher B. Wilson. Relativized circuit complexity. Journal of Computer and System Sciences, 31(2): 169–181, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900406.

Flajolet:1985:PCA

[809] Philippe Flajolet and G. Nigel Martin. Probabilistic counting algorithms for data base applications. Journal of Computer and System Sciences, 31 (2):182–209, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900418.

Incerpi:1985:IUB

[810] Janet Incerpi and Robert Sedgewick. Improved upper bounds on shellsort. Journal of Computer and System Sciences, 31(2):210–224, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/002200008590042X.

vonzurGathen:1985:IMP

[811] Joachim von zur Gathen. Irreducibility of multivariate polynomials. Journal of Computer and System Sciences, 31 (2):225–264, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900431.

vonzurGathen:1985:FSM

[812] Joachim von zur Gathen and Erich Kaltofen. Factoring sparse multivariate polynomials. Journal of Computer and System Sciences, 31(2): 265–287, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900443.

Papadimitriou:1985:GAN

[813] Christos H. Papadimitriou. Games against nature. Journal of Computer and System Sciences, 31(2):

288–301, October 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900455.

Senizergues:1985:EIP

[814] G. Senizergues. The equivalence and inclusion problems for NTS languages. Journal of Computer and System Sciences, 31(3):303– 331, December 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900558.

Boasson:1985:NLD

[815] L. Boasson and G. Senizergues. NTS languages are deterministic and congruential. Journal of Computer and System Sciences, 31(3):332– 342, December 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008590056X.

Fischer:1985:IBD

[816] Patrick C. Fischer, Lawrence V. Saxton, Stan J. Thomas, and Dirk Van Gucht. Interactions between dependencies and nested relational structures. Journal of Computer and System Sciences, 31(3):343– 354, December 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000085900571.

Bernstein:1985:STR

[817] Philip A. Bernstein and Nathan Goodman. Serializability theory for replicated databases. Journal of Computer and System Sciences, 31(3): 355–374, December 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900583.

Bloom:1985:AST

[818] Stephen L. Bloom and Zoltán Esik. Axiomatizing schemes and their behaviors. Journal of Computer and System Sciences, 31(3):375– 393, December 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900595.

Apostolico:1985:SPS

[819] A. Apostolico and F. P. Preparata. Structural properties of the string statistics problem. Journal of Computer and System Sciences, 31(3): 394–411, December 1985. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900601.

Anonymous:1985:AIVb

[820] Anonymous. Author index for volume 31. Journal of Computer and System Sciences, 31(3): 412, December 1985. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000085900613.

Li:1986:BTU

[821] Shuo-Yen Robert Li. Binary trees and uniform distribution of traffic cutback. Journal of Computer and System Sciences, 32(1):1– 14, February 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900012.

Gacs:1986:RCC

[822] Peter Gács. Reliable computation with cellular automata. Journal of Computer and System Sciences, 32 (1):15–78, February 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900024.

deChampeaux:1986:APW

[823] Dennis de Champeaux. About the Paterson–Wegman linear unification algorithm. Journal of Computer and System Sciences, 32(1):79– 90, February 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900036.

Wegener:1986:TST

[824] Ingo Wegener. Time-space trade-offs for branching programs. Journal of Computer and System Sciences, 32 (1):91–96, February 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900048.

Urzyczyn:1986:DEA

[825] Pawel Urzyczyn. “During’cannot be expressed by “after”. Journal of Computer and System Sciences, 32 (1):97–104, February 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008690005X.

Rosier:1986:MAB

[826] Louis E. Rosier and Hsu-Chun Yen. A multiparameter analysis of the boundedness problem for vector addition systems. Journal of Computer and System Sciences, 32(1): 105–135, February 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900061.

Kantabutra:1986:NAM

[827] V. Kantabutra and S. R. Kosaraju. New algorithms for multilink robot arms. Journal of Computer and System Sciences, 32(1):136–153, February 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000086900073.

DeMillo:1986:EF

[828] Richard A. DeMillo and Edward K. Blum. Editors’ foreword. Journal of Computer and System Sciences, 32(2):155, April 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900243.

Chazelle:1986:RCS

[829] Bernard Chazelle. Reporting and counting segment intersections. Journal of Computer and System Sciences, 32(2):156–182, April 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900255.

Vardi:1986:ATT

[830] Moshe Y. Vardi and Pierre Wolper. Automata-theoretic techniques for modal logics of programs. Journal of Computer and System Sciences, 32 (2):183–221, April 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900267.

Boppana:1986:TFB

[831] Ravi B. Boppana. Threshold functions and bounded depth monotone circuits. Journal of Computer and System Sciences, 32(2): 222–229, April 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900279.

Dolev:1986:PIA

[832] Danny Dolev, Joseph Y. Halpern, and H. Raymond Strong. On the possibility and impossibility of achieving clock synchronization. Journal of Computer and System Sciences, 32 (2):230–250, April 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900280.

Ben-Or:1986:CEA

[833] Michael Ben-Or, Dexter Kozen, and John Reif. The complexity of elementary algebra and geometry. Journal of Computer and System Sciences, 32(2):251–264, April 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900292.

Miller:1986:FSS

[834] Gary L. Miller. Finding small simple cycle separators for 2-connected planar graphs. Journal of Computer and System Sciences, 32(3): 265–279, June 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900309.

Reif:1986:ESA

[835] John H. Reif and Harry R. Lewis. Efficient symbolic analysis of programs. Journal of Computer and System Sciences, 32(3):280–314, June 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000086900310.

Karhumaki:1986:SUP

[836] J. Karhumäki and Y. Maon. A simple undecidable problem: Existential agreement of inverses of two morphisms on a regular language. Journal of Computer and System Sciences, 32(3):315–322, June 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900322.

Bergstra:1986:CRR

[837] J. A. Bergstra and J. W. Klop. Conditional rewrite rules: Confluence and termination. Journal of Computer and System Sciences, 32 (3):323–362, June 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900334.

Hodgson:1986:UNF

[838] Bernard R. Hodgson and Clement F. Kent. Uniform normal form for general time-bounded complexity classes. Journal of Computer and System Sciences, 32(3):363– 369, June 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900346.

Owings:1986:ECF

[839] James C. Owings. Effective choice functions and index sets. Journal of Computer and System Sciences, 32(3):370–373, June 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900358.

Chlebus:1986:DTG

[840] Bogdan S. Chlebus. Dominotiling games. Journal of Computer and System Sciences, 32(3): 374–392, June 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008690036X.

Perrin:1986:FOL

[841] Dominique Perrin and Jean-Eric Pin. First-order logic and starfree sets. Journal of Computer and System Sciences, 32(3):393– 406, June 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900371.

Lakshmipathy:1986:GGP

[842] N. Lakshmipathy and Karl Winklmann. “Global” graph problems tend to be intractable. Journal of Computer and System Sciences, 32 (3):407–428, June 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900383.

Anonymous:1986:AIVa

[843] Anonymous. Author index for volume 32. Journal of Computer and System Sciences, 32(3): 429, June 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900395.

Stockmeyer:1986:EF

[844] Larry Stockmeyer. Editor’s foreword. Journal of Computer and System Sciences, 33(1):1, August 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900401.

Mulmuley:1986:FAS

[845] Ketan Mulmuley. Fully abstract submodels of typed lambda calculi. Journal of Computer and System Sciences, 33(1):2–46, August 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900413.

Fiat:1986:PAN

[846] Amos Fiat and Adi Shamir. Polymorphic arrays: a novel VLSI layout for systolic computers. Journal of Computer and System Sciences, 33(1):47–65, August 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900425.

Munro:1986:IDS

[847] J. Ian Munro. An implicit data structure supporting insertion, deletion, and search in O(log^{2} n) time. Journal of Computer and System Sciences, 33(1):66–74, August 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900437.

Santha:1986:GQR

[848] Miklos Santha and Umesh V. Vazirani. Generating quasi-random sequences from semi-random sources. Journal of Computer and System Sciences, 33(1):75–87, August 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900449.

Greenberg:1986:LBP

[849] Albert G. Greenberg and Alan Weiss. A lower bound for probabilistic algorithms for finite state machines. Journal of Computer and System Sciences, 33(1):88–105, August 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900450.

Paturi:1986:PCC

[850] Ramamohan Paturi and Janos Simon. Probabilistic communication complexity. Journal of Computer and System Sciences, 33(1): 106–123, August 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900462.

Ullman:1986:EF

[851] Jeff Ullman. Editor’s foreword. Journal of Computer and System Sciences, 33(2):125, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900140.

Mannila:1986:DEA

[852] Heikki Mannila and Karijouko Raihä. Design by example: an application of Armstrong relations. Journal of Computer and System Sciences, 33 (2):126–141, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900152.

Vardi:1986:QLD

[853] Moshe Y. Vardi. Querying logical databases. Journal of Computer and System Sciences, 33(2): 142–160, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900164.

Wolfson:1986:DFS

[854] Ouri Wolfson and Mihalis Yannakakis. Deadlock-freedom (and safety) of transactions in a distributed database. Journal of Computer and System Sciences, 33(2): 161–178, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900176.

Ausiello:1986:CPG

[855] Giorgio Ausiello, Alessandro D’Atri, and Marina Moscarini. Chordality properties on graphs and minimal conceptual connections in semantic data models. Journal of Computer and System Sciences, 33(2): 179–202, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900188.

Cosmadakis:1986:PSR

[856] Stavros S. Cosmadakis, Paris C. Kanellakis, and Nicolas Spyratos. Partition semantics for relations. Journal of Computer and System Sciences, 33 (2):203–233, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008690019X.

Connors:1986:EVQ

[857] Tim Connors. Equivalence of views by query capacity. Journal of Computer and System Sciences, 33(2): 234–274, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900206.

Sagiv:1986:COB

[858] Yehoshua Sagiv. Concurrent operations on B^{∗}-trees with overtaking. Journal of Computer and System Sciences, 33(2):275– 296, October 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900218.

Hadzilacos:1986:AAM

[859] Thanasis Hadzilacos and Christos H. Papadimitriou. Algorithmic aspects of multiversion concurrency control. Journal of Computer and System Sciences, 33(2): 297–310, October 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008690022X.

Imielinski:1986:WLJ

[860] T. Imielinski. Witold Lipski, Jr. (1949–1985). Journal of Computer and System Sciences, 33(2): 311, October 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900231.

Grumberg:1986:CRE

[861] Orna Grumberg, Nissim Francez, and Shmuel Katz. A complete rule for equifair termination. Journal of Computer and System Sciences, 33(3): 313–332, December 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900565.

Tarlecki:1986:QVA

[862] Andrzej Tarlecki. Quasi-varieties in abstract algebraic institutions. Journal of Computer and System Sciences, 33 (3):333–360, December 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900577.

Abiteboul:1986:NFN

[863] Serge Abiteboul and Nicole Bidoit. Non First Normal Form relations: an algebra allowing data restructuring. Journal of Computer and System Sciences, 33(3):361–393, December 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000086900589.

Kaliski:1986:TOB

[864] Martin E. Kaliski, S. Yunkap Kwankam, Pamela Halpern, and David Shulman. A theory of orbital behavior in a class of nonlinear systems: Chaos and a signature-based approach. Journal of Computer and System Sciences, 33(3):

394–414, December 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900590.

Janssens:1986:BDP

[865] Dirk Janssens, Grzegorz Rozenberg, and Emo Welzl. The bounded degree problem for NLC grammars is decidable. Journal of Computer and System Sciences, 33(3): 415–422, December 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900607.

Hotz:1986:RTI

[866] Günter Hotz. A representation theorem of infinite dimensional algebras and applications to language theory. Journal of Computer and System Sciences, 33 (3):423–455, December 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900619.

Dahlhaus:1986:MGC

[867] Elias Dahlhaus and Manfred K. Warmuth. Membership for growing context-sensitive grammars is polynomial. Journal of Computer and System Sciences, 33(3):456– 472, December 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000086900620.

Card:1986:ABL

[868] H. C. Card, A. Thanailakis, W. Pries, and R. D. McLeod. Analysis of bounded linear cellular automata based on a method of image charges. Journal of Computer and System Sciences, 33 (3):473–480, December 1986. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900632.

Anonymous:1986:AIVb

[869] Anonymous. Author index for volume 33. Journal of Computer and System Sciences, 33(3): 481, December 1986. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000086900644.

Cantone:1987:DPE

[870] D. Cantone, A. Ferro, and J. T. Schwartz. Decision procedures for elementary sublanguages of set theory. V. Multilevel syllogistic extended by the general union operator. Journal of Computer and System Sciences, 34(1):1–18, February 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900018.

Frederickson:1987:UBT

[871] Greg N. Frederickson. Upper bounds for time-space trade-offs in sorting and selection. Journal of Computer and System Sciences, 34(1):19– 26, February 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008790002X.

Amir:1987:ECF

[872] Amihood Amir. Expressive completeness failure in branching time structures. Journal of Computer and System Sciences, 34(1):27–42, February 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000087900031.

Mirzaian:1987:RRV

[873] Andranik Mirzaian. River routing in VLSI. Journal of Computer and System Sciences, 34(1):43– 54, February 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000087900043.

Howell:1987:ANP

[874] Rodney R. Howell and Louis E. Rosier. An analysis of the nonemptiness problem for classes of reversal-bounded multicounter machines. Journal of Computer and System Sciences, 34 (1):55–74, February 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000087900055.

Ko:1987:NTV

[875] Ker-I Ko and Chin-Ming Hua. A note on the two-variable patternfinding problem. Journal of Computer and System Sciences, 34(1):75– 86, February 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900067.

Vogler:1987:BTT

[876] Heiko Vogler. Basic tree transducers. Journal of Computer and System Sciences, 34(1):87–128, February 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900079.

Abrahamson:1987:SRR

[877] Karl R. Abrahamson. Succinct representation of regular sets using gotos and boolean variables. Journal of Computer and System Sciences, 34 (1):129–148, February 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900080.

Ehrig:1987:EF

[878] Hartmut Ehrig. Editor’s foreword. Journal of Computer and System Sciences, 34(2–3):149, April/June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000087900225.

Sannella:1987:OEA

[879] Donald Sannella and Andrzej Tarlecki. On observational equivalence and algebraic specification. Journal of Computer and System Sciences, 34(2–3): 150–178, April/June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900237.

Padawitz:1987:PPD

[880] Peter Padawitz. Parameter-preserving data type specifications. Journal of Computer and System Sciences, 34(2– 3):179–209, April/June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900249.

Castellani:1987:BAH

[881] Ilaria Castellani. Bisimulations and abstraction homomorphisms. Journal of Computer and System Sciences, 34(2– 3):210–235, April/June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900250.

Broy:1987:STD

[882] Manfred Broy. Specification and top-down design of distributed systems. Journal of Computer and System Sciences, 34(2–3):236–265, April/June 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900262.

Makowsky:1987:WHF

[883] J. A. Makowsky. Why horn formulas matter in computer science: Initial structures and generic examples. Journal of Computer and System Sciences, 34(2–3):266–292, April/June 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900274.

Blum:1987:ASM

[884] E. K. Blum, H. Ehrig, and F. Parisi-Presicce. Algebraic specification of modules and their basic interconnections. Journal of Computer and System Sciences, 34(2–3):293– 339, April/June 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900286.

Poigne:1987:IAD

[885] Axel Poigné and Josef Voss. On the implementation of abstract data types by programming language constructs. Journal of Computer and System Sciences, 34(2–3):340–376, April/June 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900298.

Boehm:1987:AGT

[886] Paul Boehm, Harald-Reto Fonio, and Annegret Habel. Amalgamation of graph transformations: A synchronization mechanism. Journal of Computer and System Sciences, 34(2–3): 377–408, April/June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900304.

Burstall:1987:IDF

[887] R. M. Burstall. Inductively defined functions in functional programming languages. Journal of Computer and System Sciences, 34(2–3): 409–421, April/June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000087900316.

Degano:1987:CHB

[888] Pierpaolo Degano and Ugo Montanari. Concurrent histories: a basis for observing distributed systems. Journal of Computer and System Sciences, 34(2– 3):422–461, April/June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000087900328.

Anonymous:1987:AIVa

[889] Anonymous. Author index for volume 34. Journal of Computer and System Sciences, 34(2–3):462, April/June 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008790033X.

Biskup:1987:DRD

[890] Joachim Biskup and Renate Meyer. Design of relational database schemes by deleting attributes in the canonical decomposition. Journal of Computer and System Sciences, 35(1): 1–22, August 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000087900341.

Peleg:1987:CCD

[891] David Peleg. Communication in concurrent dynamic logic. Journal of Computer and System Sciences, 35 (1):23–58, August 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900353.

Kurshan:1987:CDB

[892] R. P. Kurshan. Complementing deterministic Büchi automata in polynomial time. Journal of Computer and System Sciences, 35(1): 59–71, August 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900365.

Parisi-Presicce:1987:UAM

[893] Francesco Parisi-Presicce. Union and actualization of module specifications: Some compatibility results. Journal of Computer and System Sciences, 35(1):72–95, August 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900377.

Watanabe:1987:ECA

[894] Toshimasa Watanabe and Akira Nakamura. Edge-connectivity augmentation problems. Journal of Computer and System Sciences, 35(1): 96–144, August 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900389.

Toda:1987:SCU

[895] Seinosuke Toda. Σ_{2} SPACE(n) is closed under complement. Journal of Computer and System Sciences, 35 (2):145–152, October 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900092.

Babai:1987:LBR

[896] László Babai, Péter Hajnal, Endre Szemerédi, and György Turán. A lower bound for read-once-only branching programs. Journal of Computer and System Sciences, 35(2): 153–162, October 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900109.

Steffanescu:1987:FTP

[897] Gheorghe Şteffanescu. On flowchart theories. Part I. The deterministic case. Journal of Computer and System Sciences, 35(2):163– 191, October 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900110.

Main:1987:HNG

[898] Michael G. Main and Grzegorz Rozenberg. Handle NLC grammars and R. E. languages. Journal of Computer and System Sciences, 35(2): 192–205, October 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900122.

Winkowski:1987:AP

[899] Józef Winkowski and Andrea Maggiolo-Schettini. An algebra of processes. Journal of Computer and System Sciences, 35(2):206–228, October 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000087900134.

Trnkova:1987:DAT

[900] Vera Trnková and Jan Reiterman. Dynamic algebras with test. Journal of Computer and System Sciences, 35 (2):229–242, October 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900146.

Autebert:1987:GNL

[901] Jean-Michel Autebert, Luc Boasson, and Géraud Sénizergues. Groups and NTS languages. Journal of Computer and System Sciences, 35(2): 243–267, October 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900158.

Fredman:1987:RCA

[902] Michael L. Fredman and Thomas H. Spencer. Refined complexity analysis for heap operations. Journal of Computer and System Sciences, 35(3): 269–284, December 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008790016X.

Otto:1987:DCF

[903] Friedrich Otto. On deciding the confluence of a finite string-rewriting system on a given congruence class. Journal of Computer and System Sciences, 35 (3):285–310, December 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900171.

Beauquier:1987:CCF

[904] J. Beauquier, M. Blattner, and M. Latteux. On commutative contextfree languages. Journal of Computer and System Sciences, 35(3): 311–320, December 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900183.

Bidoit:1987:VAH

[905] N. Bidoit. The verso algebra or how to answer queries with fewer joins. Journal of Computer and System Sciences, 35(3): 321–364, December 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900195.

Benson:1987:ASR

[906] David B. Benson and Irène Guessarian. Algebraic solutions to recursion schemes. Journal of Computer and System Sciences, 35(3): 365–400, December 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900201.

Anonymous:1987:AIVb

[907] Anonymous. Author index for volume 35. Journal of Computer and System Sciences, 35(3): 401, December 1987. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000087900213.

Baldi:1988:GAL

[908] Pierre Baldi. Group actions and learning for a family of automata. Journal of Computer and System Sciences, 36(1):1–15, February 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900177.

Lindsay:1988:AA

[909] Peter A. Lindsay. On alternating ω-automata. Journal of Computer and System Sciences, 36(1):16– 24, February 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900189.

Katsura:1988:AGF

[910] Masashi Katsura. Automorphism groups and factor automata of strongly connected automata. Journal of Computer and System Sciences, 36 (1):25–65, February 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900190.

Esik:1988:IEA

[911] Z. Ésik. Independence of the equational axioms for iteration theories. Journal of Computer and System Sciences, 36(1):66–76, February 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900207.

VanGucht:1988:MNR

[912] Dirk Van Gucht and Patrick C. Fischer. Multilevel nested relational structures. Journal of Computer and System Sciences, 36(1): 77–105, February 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900219.

Odlyzko:1988:EF

[913] Andrew Odlyzko. Editor’s foreword. Journal of Computer and System Sciences, 36(2):107, April 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900220.

Wilber:1988:WPH

[914] Robert Wilber. White pebbles help. Journal of Computer and System Sciences, 36(2):108–124, April 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000088900232.

Gacs:1988:STD

[915] Peter Gács and John Reif. A simple three-dimensional real-time reliable cellular array. Journal of Computer and System Sciences, 36(2): 125–147, April 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900244.

Chung:1988:SOS

[916] F. R. K. Chung, D. J. Hajela, and P. D. Seymour. Self-organizing sequential search and Hilbert’s inequalities. Journal of Computer and System Sciences, 36(2):148–157, April 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900256.

deBakker:1988:TSM

[917] J. W. de Bakker, J.-J. Ch. Meyer, E.- R. Olderog, and J. I. Zucker. Transition systems, metric spaces and ready sets in the semantics of uniform concurrency. Journal of Computer and System Sciences, 36(2): 158–224, April 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900268.

Karp:1988:CPS

[918] Richard M. Karp, Eli Upfal, and Avi Wigderson. The complexity of parallel search. Journal of Computer and System Sciences, 36(2): 225–253, April 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008890027X.

Babai:1988:AMG

[919] László Babai and Shlomo Moran. Arthur–Merlin games: a randomized proof system, and a hierarchy of complexity classes. Journal of Computer and System Sciences, 36 (2):254–276, April 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900281.

Mahaney:1988:EF

[920] Stephen R. Mahaney. Editor’s foreword. Journal of Computer and System Sciences, 36(3):277, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000088900293.

Parberry:1988:PCT

[921] Ian Parberry and Georg Schnitger. Parallel computation with threshold functions. Journal of Computer and System Sciences, 36(3): 278–302, June 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008890030X.

Wilson:1988:MRS

[922] Christopher B. Wilson. A measure of relativized space which is faithful with respect to depth. Journal of Computer and System Sciences, 36(3):303–312, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900311.

Dietzfelbinger:1988:LBA

[923] Martin Dietzfelbinger and Wolfgang Maass. Lower bound arguments with “inaccessible” numbers. Journal of Computer and System Sciences, 36 (3):313–335, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900323.

Allender:1988:IR

[924] Eric W. Allender. Isomorphisms and 1 − L reductions. Journal of Computer and System Sciences, 36 (3):336–350, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900335.

Buss:1988:RAS

[925] Jonathan F. Buss. Relativized alternation and space-bounded computation. Journal of Computer and System Sciences, 36(3):351– 378, June 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900347.

Sipser:1988:ERT

[926] Michael Sipser. Expanders, randomness, or time versus space. Journal of Computer and System Sciences, 36(3):379–383, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900359.

Regan:1988:TPC

[927] Kenneth W. Regan. The topology of provability in complexity theory. Journal of Computer and System Sciences, 36(3):384–432, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900360.

Zachos:1988:PQG

[928] Stathis Zachos. Probabilistic quantifiers and games. Journal of Computer and System Sciences, 36(3): 433–451, June 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900372.

Condon:1988:PGA

[929] Anne Condon and Richard E. Ladner. Probabilistic game automata. Journal of Computer and System Sciences, 36(3):452–489, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900384.

Krentel:1988:COP

[930] Mark W. Krentel. The complexity of optimization problems. Journal of Computer and System Sciences, 36(3):490–509, June 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900396.

Anonymous:1988:AIVa

[931] Anonymous. Author index for volume 36. Journal of Computer and System Sciences, 36(3): 510, June 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900402.

Lagarias:1988:EF

[932] Jeff Lagarias. Editor’s foreword. Journal of Computer and System Sciences, 37(1):1, August 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900414.

Papadimitriou:1988:CFR

[933] Christos H. Papadimitriou and David Wolfe. The complexity of facets resolved. Journal of Computer and System Sciences, 37(1):2–13, August 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000088900426.

Kirousis:1988:CRP

[934] Lefteris M. Kirousis and Christos H. Papadimitriou. The complexity of recognizing polyhedral scenes. Journal of Computer and System Sciences, 37(1):14–38, August 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900438.

Luks:1988:PAS

[935] Eugene M. Luks and Pierre McKenzie. Parallel algorithms for solvable permutation groups. Journal of Computer and System Sciences, 37 (1):39–62, August 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008890044X.

Landau:1988:FSM

[936] Gad M. Landau and Uzi Vishkin. Fast string matching with k differences. Journal of Computer and System Sciences, 37(1):63–78, August 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000088900451.

Johnson:1988:HEL

[937] David S. Johnson, Christos H. Papadimitriou, and Mihalis Yannakakis. How easy is local search? Journal of Computer and System Sciences, 37 (1):79–100, August 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900463.

Li:1988:STP

[938] Ming Li. Simulating two pushdown stores by one tape in O(n^{1.5}\sqrt{log n}) time. Journal of Computer and System Sciences, 37(1):101–116, August 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000088900475.

Kannan:1988:EF

[939] Ravi Kannan. Editors’s foreword. Journal of Computer and System Sciences, 37(2):117, October 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000088900013.

Alon:1988:MTA

[940] Noga Alon and Wolfgang Maasst. Meanders and their applications in lower bounds arguments. Journal of Computer and System Sciences, 37 (2):118–129, October 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900025.

Raghavan:1988:PCD

[941] Prabhakar Raghavan. Probabilistic construction of deterministic algorithms: Approximating packing integer programs. Journal of Computer and System Sciences, 37(2): 130–143, October 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900037.

Chrobak:1988:KBT

[942] Marek Chrobak and Ming Li. k + 1heads are better than k for PDAs. Journal of Computer and System Sciences, 37(2):144–155, October 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000088900049.

Brassard:1988:MDP

[943] Gilles Brassard, David Chaum, and Claude Crépeau. Minimum disclosure proofs of knowledge. Journal of Computer and System Sciences, 37 (2):156–189, October 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900050.

Klein:1988:EPA

[944] Philip N. Klein and John H. Reif. An efficient parallel algorithm for planarity. Journal of Computer and System Sciences, 37(2):190– 246, October 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900062.

Kurtz:1988:CD

[945] Stuart A. Kurtz, Stephen R. Mahaney, and James S. Royer. Collapsing degrees. Journal of Computer and System Sciences, 37(2): 247–268, October 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900074.

Chan:1988:PAR

[946] Tat-Hung Chan. Pushdown automata with reversal-bounded counters. Journal of Computer and System Sciences, 37(3):269–291, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000088900086.

Lloyd:1988:LMF

[947] Errol L. Lloyd, Mary Lou Soffa, and Ching-Chy Wang. On locating minimum feedback vertex sets. Journal of Computer and System Sciences, 37 (3):292–311, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900098.

Schoning:1988:GIL

[948] Uwe Schöning. Graph isomorphism is in the low hierarchy. Journal of Computer and System Sciences, 37(3): 312–323, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900104.

OConnor:1988:UAF

[949] Mary G. O’Connor. An unpredictability approach to finite-state randomness. Journal of Computer and System Sciences, 37(3): 324–336, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900116.

Frederickson:1988:DAS

[950] Greg N. Frederickson. Distributed algorithms for selection in sets. Journal of Computer and System Sciences, 37 (3):337–348, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000088900128.

Howell:1988:CRC

[951] Rodney R. Howell and Louis E. Rosier. Completeness results for conflict-free vector replacement systems. Journal of Computer and System Sciences, 37 (3):349–366, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008890013X.

Salomaa:1988:DTP

[952] Kai Salomaa. Deterministic tree pushdown automata and monadic tree rewriting systems. Journal of Computer and System Sciences, 37(3): 367–394, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900141.

Pippenger:1988:CCC

[953] Nicholas Pippenger. Correction to “Computational complexity of algebraic functions”. Journal of Computer and System Sciences, 37(3): 395–399, December 1988. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900153. See

[602].

Anonymous:1988:AIVb

[954] Anonymous. Author index for volume 37. Journal of Computer and System Sciences, 37(3): 400, December 1988. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000088900165.

Hartmanis:1989:EF

[955] Juris Hartmanis. Editor’s foreword. Journal of Computer and System Sciences, 38(1):1, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900305.

Coffman:1989:PEA

[956] E. G. Coffman and F. T. Leighton. A provably efficient algorithm for dynamic storage allocation. Journal of Computer and System Sciences, 38 (1):2–35, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900317.

Yannakakis:1989:EPG

[957] Mihalis Yannakakis. Embedding planar graphs in four pages. Journal of Computer and System Sciences, 38 (1):36–67, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900329.

Cai:1989:POR

[958] Jin-Yi Cai. With probability one, a random oracle separates PSPACE from the polynomial-time hierarchy. Journal of Computer and System Sciences, 38(1):68–85, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900330.

Driscoll:1989:MDS

[959] James R. Driscoll, Neil Sarnak, Daniel D. Sleator, and Robert E. Tarjan. Making data structures persistent. Journal of Computer and System Sciences, 38(1):86–124, February 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900342.

Ajtai:1989:OPS

[960] Miklós Ajtai, János Komlós, W. L. Steiger, and Endre Szemerédi. Optimal parallel selection has complexity O(LogLogN). Journal of Computer and System Sciences, 38(1): 125–133, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900354.

Galil:1989:NSP

[961] Zvi Galil, Ravi Kannan, and Endre Szemerédi. On nontrivial separators for k-page graphs and simulations by nondeterministic one-tape Turing machines. Journal of Computer and System Sciences, 38(1): 134–149, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900366.

Barrington:1989:BWP

[962] David A. Barrington. Boundedwidth polynomial-size branching programs recognize exactly those languages in NC 1. Journal of Computer and System Sciences, 38(1): 150–164, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900378.

Edelsbrunner:1989:TSA

[963] Herbert Edelsbrunner and Leonidas J. Guibas. Topologically sweeping an arrangement. Journal of Computer and System Sciences, 38(1): 165–194, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008990038X.

Halpern:1989:CRA

[964] Joseph Y. Halpern and Moshe Y. Vardi. The complexity of reasoning about knowledge and time. I. Lower bounds. Journal of Computer and System Sciences, 38(1): 195–237, February 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900391.

Silberschatz:1989:EF

[965] Avi Silberschatz. Editor’s foreword. Journal of Computer and System Sciences, 38(2):239, April 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900019.

Mannila:1989:AGT

[966] Heikki Mannila and Kari-Jouko Räihä. Automatic generation of test data for relational queries. Journal of Computer and System Sciences, 38 (2):240–258, April 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900020.

Naughton:1989:DIR

[967] Jeffrey F. Naughton. Data independent recursion in deductive databases. Journal of Computer and System Sciences, 38(2):259–289, April 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900032.

Bidoit:1989:MJN

[968] Nicole Bidoit and Richard Hull. Minimalism, justification and nonmonotonicity in deductive databases. Journal of Computer and System Sciences, 38(2):290–325, April 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000089900044.

Bancilhon:1989:CCO

[969] François Bancilhon and Setrag Khoshafian. A calculus for complex objects. Journal of Computer and System Sciences, 38(2):326–340, April 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900056.

Afrati:1989:CQE

[970] Foto Afrati, Christos H. Papadimitriou, George Papageorgiou, Athena Roussou, Yehoshua Sagiv, and Jeffrey D. Ullman. On the convergence of query evaluation. Journal of Computer and System Sciences, 38 (2):341–359, April 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900068.

Hadzilacos:1989:DCT

[971] Thanasis Hadzilacos and Nihalis Yannakakis. Deleting completed transactions. Journal of Computer and System Sciences, 38(2): 360–379, April 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008990007X.

Sagiv:1989:CFF

[972] Yehoshua Sagiv and Oded Shmueli. A characterization of finite fdacyclicity. Journal of Computer and System Sciences, 38(2):380– 404, April 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900081.

Olken:1989:RDM

[973] Frank Olken and Doron Rotem. Rearranging data to maximize the efficiency of compression. Journal of Computer and System Sciences, 38 (2):405–430, April 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900093.

Hromkovic:1989:LBL

[974] Juraj Hromkovic, Katsushi Inoue, and Itsuo Takanami. Lower bounds for language recognition on two-dimensional alternating multihead machines. Journal of Computer and System Sciences, 38(3):431–451, June 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200008990010X.

Ibarra:1989:ICT

[975] Oscar H. Ibarra, Tao Jiang, and Jik H. Chang. On iterative and cellular tree arrays. Journal of Computer and System Sciences, 38(3): 452–473, June 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000089900111.

Fle:1989:SCP

[976] Marie-Paule Flé. Serialization of concurrent programs. Journal of Computer and System Sciences, 38 (3):474–493, June 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000089900123.

Pan:1989:FES

[977] Victor Pan and John Reif. Fast and efficient solution of path algebra problems. Journal of Computer and System Sciences, 38(3): 494–510, June 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000089900135.

Rich:1989:PRP

[978] Craig A. Rich. Positive relativizations of the P =? NP problem. Journal of Computer and System Sciences, 38(3):511–523, June 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900147.

Anonymous:1989:AIVa

[979] Anonymous. Author index for volume 38. Journal of Computer and System Sciences, 38(3): 524, June 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900159.

Mahaney:1989:EFa

[980] Steve Mahaney and Alan L. Selman. Editors’ foreword. Journal of Computer and System Sciences, 39(1):1, August 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900160.

Chung:1989:SNT

[981] Moon Jung Chung and B. Ravikumar. Strong nondeterministic Turing reduction — a technique for proving intractability. Journal of Computer and System Sciences, 39(1): 2–20, August 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900172.

Abadi:1989:HIO

[982] Martín Abadi, Joan Feigenbaum, and Joe Kilian. On hiding information from an oracle. Journal of Computer and System Sciences, 39 (1):21–50, August 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900184.

Leivant:1989:DCC

[983] Daniel Leivant. Descriptive characterizations of computational complexity. Journal of Computer and System Sciences, 39(1):51–83, August 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000089900196.

Schoning:1989:PCC

[984] Uwe Schöning. Probabilistic complexity classes and lowness. Journal of Computer and System Sciences, 39 (1):84–100, August 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900202.

Allender:1989:SCE

[985] Eric W. Allender. Some consequences of the existence of pseudorandom generators. Journal of Computer and System Sciences, 39 (1):101–124, August 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900214.

Yap:1989:GEF

[986] Chef Yap. Guest Editor’s foreword. Journal of Computer and System Sciences, 39(2):125, October 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000089900408.

Guibas:1989:OSP

[987] Leonidas J. Guibas and John Hershberger. Optimal shortest path queries in a simple polygon. Journal of Computer and System Sciences, 39 (2):126–152, October 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008990041X.

Dadoun:1989:PCS

[988] N. Dadoun and D. G. Kirkpatrick. Parallel construction of subdivision hierarchies. Journal of Computer and System Sciences, 39(2):153– 165, October 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900421.

Culberson:1989:OCC

[989] Joseph Culberson and Robert A. Reckhow. Orthogonally convex coverings of orthogonal polygons without holes. Journal of Computer and System Sciences, 39(2):166– 204, October 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900433.

Chew:1989:TPG

[990] L. Paul Chew. There are planar graphs almost as good as the complete graph. Journal of Computer and System Sciences, 39(2): 205–219, October 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000089900445.

Suri:1989:CGF

[991] Subhash Suri. Computing geodesic furthest neighbors in simple polygons. Journal of Computer and System Sciences, 39(2):220–235, October 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000089900457.

Sugihara:1989:FPR

[992] Kokichi Sugihara. On finite-precision representations of geometric objects. Journal of Computer and System Sciences, 39(2):236–247, October 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000089900469.

Mahaney:1989:EFb

[993] Steve Mahaney and Alan L. Selman. Editors’ foreword. Journal of Computer and System Sciences, 39 (3):249, December 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000089900226.

Ambos-Spies:1989:HPT

[994] Klaus Ambos-Spies. Honest polynomial time reducibilities and the P=?NP problem. Journal of Computer and System Sciences, 39(3): 250–281, December 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000089900238.

Kadin:1989:PST

[995] Jim Kadin. P^{NP [O(log n)]} and sparse Turing-complete sets for NP. Journal of Computer and System Sciences, 39 (3):282–298, December 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200008990024X.

Hemachandra:1989:SEH

[996] Lane A. Hemachandra. The strong exponential hierarchy collapses. Journal of Computer and System Sciences, 39 (3):299–322, December 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900251.

Greenlaw:1989:OVR

[997] Raymond Greenlaw. Ordered vertex removal and subgraph problems. Journal of Computer and System Sciences, 39 (3):323–342, December 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900263.

America:1989:SRD

[998] Pierre America and Jan Rutten. Solving reflexive domain equations in a category of complete metric spaces. Journal of Computer and System Sciences, 39(3):343– 375, December 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900275.

Robinson:1989:SCG

[999] A. G. Robinson and A. J. Goldman. The set coincidence game: Complexity, attainability, and symmetric strategies. Journal of Computer and System Sciences, 39(3): 376–387, December 1989. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900287.

Anonymous:1989:AIVb

[1000] Anonymous. Author index for volume 39. Journal of Computer and System Sciences, 39(3): 388, December 1989. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000089900299.

VanWyk:1990:EF

[1001] Christopher J. Van Wyk. Editor’s foreword. Journal of Computer and System Sciences, 40(1): 1, February 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090015D.

Yap:1990:GCT

[1002] Chee-Keng Yap. A geometric consistency theorem for a symbolic perturbation scheme. Journal of Computer and System Sciences, 40(1):2– 18, February 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090016E.

Motwani:1990:COP

[1003] Rajeev Motwani, Arvind Raghunathan, and Huzur Saran. Covering orthogonal polygons with star polygons: The perfect graph approach. Journal of Computer and System Sciences, 40(1):19–48, February 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090017F.

Bern:1990:HSR

[1004] Marshall Bern. Hidden surface removal for rectangles. Journal of Computer and System Sciences, 40 (1):49–69, February 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090018G.

Dobkin:1990:ACG

[1005] David Dobkin and Deborah Silver. Applied computational geometry: Towards robust solutions of basic problems. Journal of Computer and System Sciences, 40(1):70– 87, February 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090019H.

Mitchell:1990:MFP

[1006] Joseph S. B. Mitchell. On maximum flows in polyhedral domains. Journal of Computer and System Sciences, 40 (1):88–123, February 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090020L.

Flannery:1990:HSP

[1007] Kevin E. Flannery and Johannes J. Martin. The Hoare and Smyth power domain constructors commute under composition. Journal of Computer and System Sciences, 40(2): 125–135, April 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000090900089.

Grandjean:1990:FOS

[1008] Etienne Grandjean. First-order spectra with one variable. Journal of Computer and System Sciences, 40 (2):136–153, April 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009090009A.

Dassow:1990:SA

[1009] J. Dassow and H. Jürgensen. Soliton automata. Journal of Computer and System Sciences, 40(2): 154–181, April 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009090010I.

Gruska:1990:GSM

[1010] Damas P. Gruska. Genomorphisms of semi-modular lattices. Journal of Computer and System Sciences, 40 (2):182–187, April 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000090900119.

Main:1990:ELC

[1011] Michael G. Main and Grzegorz Rozenberg. Edge-label controlled graph grammars. Journal of Computer and System Sciences, 40(2): 188–228, April 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090012A.

Poigne:1990:POS

[1012] Axel Poigné. Parametrization for order-sorted algebraic specification. Journal of Computer and System Sciences, 40(2):229–268, April 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/002200009090013B.

Lausen:1990:PSL

[1013] Georg Lausen, Eljas Soisalon-Soininen, and Peter Widmayer. On the power of safe locking. Journal of Computer and System Sciences, 40(2): 269–288, April 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090014C.

Kim:1990:PIP

[1014] Changwook Kim. Picture iteration and picture ambiguity. Journal of Computer and System Sciences, 40 (3):289–306, June 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000090900012.

Engelfriet:1990:BGG

[1015] Joost Engelfriet, George Leih, and Emo Welzl. Boundary graph grammars with dynamic edge relabeling. Journal of Computer and System Sciences, 40(3):307–345, June 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000090900023.

Ginsburg:1990:IDO

[1016] Seymour Ginsburg and Dan Tian. Input-dependent-only object histories. Journal of Computer and System Sciences, 40(3):346–375, June 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0022000090900034.

Aalbersberg:1990:CRD

[1017] Ijsbrand Jan Aalbersberg, Joost Engelfriet, and Grzegorz Rozenberg. The complexity of regular DNLC graph languages. Journal of Computer and System Sciences, 40(3): 376–404, June 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000090900045.

Ozsoyoglu:1990:ICS

[1018] Gultekin Ozsoyoglu and Tzong-An Su. On inference control in semantic data models for statistical databases. Journal of Computer and System Sciences, 40(3):405–443, June 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000090900056.

Fulk:1990:NCF

[1019] Mark A. Fulk. A note on A.E. hcomplex functions. Journal of Computer and System Sciences, 40(3): 444–449, June 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000090900067.

Anonymous:1990:AIVa

[1020] Anonymous. Author index for volume 40. Journal of Computer and System Sciences, 40(3): 450, June 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000090900078.

Vardi:1990:EF

[1021] Moshe Y. Vardi. Editor’s foreword.

Journal of Computer and System Sciences, 41(1):1, August 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090030O.

Dechter:1990:DRT

[1022] Rina Dechter. Decomposing a relation into a tree of binary relations. Journal of Computer and System Sciences, 41(1):2–24, August 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090031F.

Atzeni:1990:NIN

[1023] Paolo Atzeni and Maria Cristina De Bernardis. A new interpretation for null values in the weak instance model. Journal of Computer and System Sciences, 41(1): 25–43, August 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090032G.

Kuper:1990:LPS

[1024] Gabriel M. Kuper. Logic programming with sets. Journal of Computer and System Sciences, 41(1): 44–64, August 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090033H.

Fekete:1990:CBL

[1025] Alan Fekete, Nancy Lynch, Michael Merritt, and William Weihl. Commutativitybased locking for nested transactions. Journal of Computer and System Sciences, 41(1):65–156, August 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/002200009090034I.

Beeri:1990:BPS

[1026] Catriel Beeri, Paris Kanellakis, François Bancilhon, and Raghu Ramakrishnan. Bounds on the propagation of selection into logic programs. Journal of Computer and System Sciences, 41 (2):157–180, October 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090035J.

Abiteboul:1990:PLD

[1027] Serge Abiteboul and Victor Vianu.

Procedural languages for database queries and updates. Journal of Computer and System Sciences, 41 (2):181–229, October 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090036K.

Bermudez:1990:PAL

[1028] Manuel E. Bermudez and Karl M.

Schimpf. Practical arbitrary lookahead LR parsing. Journal of Computer and System Sciences, 41(2): 230–250, October 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090037L.

Hemachandra:1990:CR

[1029] Lane A. Hemachandra and Steven Rudich. On the complexity of ranking. Journal of Computer and System Sciences, 41(2):251– 271, October 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090038M.

Ladner:1990:EF

[1030] Richard E. Ladner. Editor’s foreword. Journal of Computer and System Sciences, 41(3):273, December 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/002200009090021C.

Barrington:1990:UWN

[1031] David A. Mix Barrington, Neil Immerman, and Howard Straubing. On uniformity within NC 1. Journal of Computer and System Sciences, 41(3): 274–306, December 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009090022D.

Lutz:1990:PSB

[1032] Jack H. Lutz. Pseudorandom sources for BPP. Journal of Computer and System Sciences, 41(3):307– 320, December 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009090023E.

Shinoda:1990:TPD

[1033] Juichi Shinoda and Theodore A. Slaman. On the theory of the PTIME degrees of the recursive sets. Journal of Computer and System Sciences, 41 (3):321–366, December 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009090024F.

Balcazar:1990:SR

[1034] JoséL. Balcázar. Self-reducibility. Journal of Computer and System Sciences, 41(3):367–388, December 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/002200009090025G.

Longpre:1990:CRF

[1035] Luc Longpré and Paul Young. Cook reducibility is faster than Karp reducibility in NP. Journal of Computer and System Sciences, 41(3): 389–401, December 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090026H.

Halstenberg:1990:RBC

[1036] Bernd Halstenberg and Rüdiger Reischuk. Relations between communication complexity classes. Journal of Computer and System Sciences, 41(3): 402–429, December 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090027I.

Pitt:1990:PPR

[1037] Leonard Pitt and Manfred K. Warmuth. Prediction-preserving reducibility. Journal of Computer and System Sciences, 41(3):430– 467, December 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090028J.

Anonymous:1990:AIVb

[1038] Anonymous. Author index for volume 41. Journal of Computer and System Sciences, 41(3): 468, December 1990. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009090029K.

Broy:1991:DVP

[1039] Manfred Broy and Christian Lengauer. On denotational versus predicative semantics. Journal of Computer and System Sciences, 42(1):1– 29, February 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900376.

Bach:1991:RAS

[1040] Eric Bach. Realistic analysis of some randomized algorithms. Journal of Computer and System Sciences, 42 (1):30–53, February 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900387.

You:1991:UME

[1041] Jia-Hual You. Unification modulo an equality theory for equational logic programming. Journal of Computer and System Sciences, 42 (1):54–75, February 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900398.

Beigel:1991:RCC

[1042] Richard Beigel. Relativized counting classes: Relations among thresholds, parity, and mods. Journal of Computer and System Sciences, 42 (1):76–96, February 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190040C.

Bodlaender:1991:SLB

[1043] H. L. Bodlaender. Some lower bound results for decentralized extremafinding in rings of processors. Journal of Computer and System Sciences, 42 (1):97–118, February 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000091900413.

Li:1991:RBP

[1044] Ming Li and Yaacov Yesha. Resource bounds for parallel computation of threshold and symmetric functions. Journal of Computer and System Sciences, 42(1): 119–137, February 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000091900424.

Giovannetti:1991:KLL

[1045] Elio Giovannetti, Giorgio Levi, Corrado Moiso, and Catuscia Palamidessi. Kernel-LEAF: a logic plus functional language. Journal of Computer and System Sciences, 42(2): 139–185, April 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190009T.

Otto:1991:OFP

[1046] Friedrich Otto and Celia Wrathall.

Overlaps in free partially commutative monoids. Journal of Computer and System Sciences, 42(2): 186–198, April 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900103.

Naughton:1991:OSR

[1047] Jeffrey F. Naughton. One-sided recursions. Journal of Computer and System Sciences, 42(2):199– 236, April 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190011S.

Garzon:1991:ITC

[1048] Max Garzon and Yechezkel Zalcstein. On isomorphism testing of a class of 2-nilpotent groups. Journal of Computer and System Sciences, 42 (2):237–248, April 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190012T.

Anonymous:1991:C

[1049] Anonymous. Corrigendum. Journal of Computer and System Sciences, 42(2):249–251, April 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190013U.

Kosaraju:1991:EF

[1050] S. Rao Kosaraju. Editor’s foreword. Journal of Computer and System Sciences, 42(3):253, June 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190001L.

Duris:1991:TLB

[1051] Pavol Duris and Zvi Galil. Two lower bounds in asynchronous distributed computation. Journal of Computer and System Sciences, 42 (3):254–266, June 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190002M.

Yao:1991:LBR

[1052] Andrew Chi-Chih Yao. Lower bounds to randomized algorithms for graph properties. Journal of Computer and System Sciences, 42(3): 267–287, June 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190003N.

Kanevsky:1991:IAG

[1053] Arkady Kanevsky and Vijaya Ramachandran. Improved algorithms for graph four- connectivity. Journal of Computer and System Sciences, 42(3):288–306, June 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190004O.

Ranade:1991:HES

[1054] Abhiram G. Ranade. How to emulate shared memory. Journal of Computer and System Sciences, 42 (3):307–326, June 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190005P.

Aiello:1991:SZK

[1055] William Aiello and Johan Hastad. Statistical zero-knowledge languages can be recognized in two rounds. Journal of Computer and System Sciences, 42(3):327–345, June 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190006Q.

Gurevich:1991:ACC

[1056] Yuri Gurevich. Average case completeness. Journal of Computer and System Sciences, 42(3):346– 398, June 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190007R.

Anonymous:1991:AIVa

[1057] Anonymous. Author index for volume 42. Journal of Computer and System Sciences, 42(3): 399, June 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190008S.

Yannakakis:1991:EF

[1058] Mihalis Yannakakis. Editor’s foreword. Journal of Computer and System Sciences, 43(1):1, August 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900295.

Hadzilacos:1991:TSO

[1059] Thanasis Hadzilacos and Vassos Hadzilacos. Transaction synchronisation in object bases. Journal of Computer and System Sciences, 43(1): 2–24, August 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000091900309.

Herlihy:1991:HCC

[1060] Maurice P. Herlihy and William E.

Weihl. Hybrid concurrency control for abstract data types. Journal of Computer and System Sciences, 43 (1):25–61, August 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190031Y.

Abiteboul:1991:DED

[1061] Serge Abiteboul and Victor Vianu.

Datalog extensions for database queries and updates. Journal of Computer and System Sciences, 43(1): 62–124, August 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190032Z.

Kolaitis:1991:WNF

[1062] Phokion G. Kolaitis and Christos H.

Papadimitriou. Why not negation by fixpoint? Journal of Computer and System Sciences, 43(1): 125–144, August 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000091900332.

Haddad:1991:CAC

[1063] Ramsey W. Haddad and Jeffrey F. Naughtont. A counting algorithm for a cyclic binary query. Journal of Computer and System Sciences, 43 (1):145–169, August 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900343.

Buneman:1991:SCO

[1064] O. P. Buneman, S. B. Davidson, and A. Watters. A semantics for complex objects and approximate answers. Journal of Computer and System Sciences, 43(1): 170–218, August 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900354.

Hull:1991:EPD

[1065] Richard Hull and Jianwen Su. On the expressive power of database queries with intermediate types. Journal of Computer and System Sciences, 43 (1):219–267, August 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900365.

Abrahamson:1991:TST

[1066] Karl Abrahamson. Time-space tradeoffs for algebraic problems on general sequential machines. Journal of Computer and System Sciences, 43 (2):269–289, October 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190014V.

Alon:1991:LBR

[1067] Noga Alon, Amotz Bar-Noy, Nathan Linial, and David Peleg. A lower bound for radio broadcast. Journal of Computer and System Sciences, 43 (2):290–298, October 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190015W.

Habarra:1991:LRL

[1068] Oscar Habarra and Tao Jiang. Learning regular languages from counterexamples. Journal of Computer and System Sciences, 43(2): 299–316, October 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190016X.

Krizanc:1991:ORL

[1069] Danny Krizanc. Oblivious routing with limited buffer capacity. Journal of Computer and System Sciences, 43 (2):317–327, October 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190017Y.

Engelfriet:1991:SGP

[1070] Joost Engelfriet and Linda Heyker.

The string generating power of contextfree hypergraph grammars. Journal of Computer and System Sciences, 43 (2):328–360, October 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009190018Z.

Stifter:1991:AAV

[1071] Sabine Stifter. An axiomatic approach to Voronoi-diagrams in 3D. Journal of Computer and System Sciences, 43 (2):361–379, October 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900192.

Venkateswaran:1991:PCL

[1072] H. Venkateswaran. Properties that characterize LOGCFL. Journal of Computer and System Sciences, 43 (2):380–404, October 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900206.

Ullman:1991:EF

[1073] Jeffrey D. Ullman. Editor’s foreword. Journal of Computer and System Sciences, 43(3):405, December 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/002200009190021V.

Fiat:1991:IDS

[1074] Amos Fiat, J. Ian Munro, Moni Naor, Alejandro A. Schäffer, Jeanette P. Schmidt, and Alan Siegel. An implicit data structure for searching a multikey table in logarithmic time. Journal of Computer and System Sciences, 43 (3):406–424, December 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190022W.

Papadimitriou:1991:OAC

[1075] Christos H. Papadimitriou and Mihalis Yannakakis. Optimization, approximation, and complexity classes. Journal of Computer and System Sciences, 43 (3):425–440, December 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190023X.

Yannakakis:1991:ECO

[1076] Mihalis Yannakakis. Expressing combinatorial optimization problems by linear programs. Journal of Computer and System Sciences, 43(3): 441–466, December 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190024Y.

Fenton:1991:PSS

[1077] N. E. Fenton and R. W. Whitty.

Program structures: Some new characterisations. Journal of Computer and System Sciences, 43(3): 467–483, December 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009190025Z.

Ito:1991:OIC

[1078] M. Ito, H. Jürgensen, H. J. Shyr, and G. Thierrin. Outfix and infix codes and related classes of languages. Journal of Computer and System Sciences, 43 (3):484–508, December 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900262.

Li:1991:CDR

[1079] Leong Kwan Li. On computing decision regions with neural nets. Journal of Computer and System Sciences, 43 (3):509–512, December 1991. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900273.

Anonymous:1991:AIVb

[1080] Anonymous. Author index for volume 43. Journal of Computer and System Sciences, 43(3): 513, December 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000091900284.

Chan:1992:CTF

[1081] Edward P. F. Chan and Paolo Atzeni. Connection-trap-free database schemes. Journal of Computer and System Sciences, 44(1):1–22, February 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290001Y.

Diaconescu:1992:CAU

[1082] Razvan Diaconescu. Contraction algebras and unification of (infinite) terms. Journal of Computer and System Sciences, 44(1):23– 42, February 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290002Z.

Muller:1992:PRE

[1083] D. E. Muller and F. P. Preparata.

Parallel restructuring and evaluation of expressions. Journal of Computer and System Sciences, 44(1):43– 62, February 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000092900032.

Lengauer:1992:CBC

[1084] T. Lengauer and K. W. Wagner. The correlation between the complexities of the nonhierarchical and hierarchical versions of graph problems. Journal of Computer and System Sciences, 44(1):63–93, February 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000092900043.

Dahlhaus:1992:PMR

[1085] Elias Dahlhaus and Marek Karpinski. Perfect matching for regular graphs is AC^{0}-hard for the general matching problem. Journal of Computer and System Sciences, 44(1): 94–102, February 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000092900054.

Ragde:1992:PTT

[1086] Prabhakar Ragde. Processor-time tradeoffs in PRAM simulations. Journal of Computer and System Sciences, 44(1):103–113, February 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000092900065.

Takahashi:1992:SSL

[1087] Satoshi Takahashi. Self-similarity of linear cellular automata. Journal of Computer and System Sciences, 44 (1):114–140, February 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000092900076.

Sandhu:1992:USS

[1088] Ravinderpal Singh Sandhu. Undecidability of safety for the schematic protection model with cyclic creates. Journal of Computer and System Sciences, 44(1):141–159, February 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0022000092900087.

Herbst:1992:SRT

[1089] Thomas Herbst. Some remarks on a theorem of Sakarovitch. Journal of Computer and System Sciences, 44 (1):160–165, February 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000092900098.

Young:1992:EF

[1090] Paul Young. Editor’s foreword. Journal of Computer and System Sciences, 44(2):167, April 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290017D.

Maass:1992:CTC

[1091] Wolfgang Maass and Theodore A.

Slaman. The complexity types of computable sets. Journal of Computer and System Sciences, 44(2): 168–192, April 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290018E.

Ben-David:1992:TAC

[1092] Shai Ben-David, Benny Chor, Oded Goldreich, and Michel Luby. On the theory of average case complexity. Journal of Computer and System Sciences, 44(2):193– 219, April 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290019F.

Lutz:1992:AEH

[1093] Jack H. Lutz. Almost everywhere high nonuniform complexity. Journal of Computer and System Sciences, 44(2):220–258, April 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290020J.

Feige:1992:MOI

[1094] Uriel Feige and Adi Shamir. Multioracle interactive protocols with constant space verifiers. Journal of Computer and System Sciences, 44 (2):259–271, April 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290021A.

Kobler:1992:TMF

[1095] Johannes Kobler, Uwe Schïng, Seinosuke Toda, and Jacobo Torán. Turing machines with few accepting computations and low sets for PP. Journal of Computer and System Sciences, 44(2):272–286, April 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290022B.

Homer:1992:OSP

[1096] Steven Homer and Alan L. Selman. Oracles for structural properties: The isomorphism problem and public-key cryptography. Journal of Computer and System Sciences, 44 (2):287–301, April 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290023C.

Mayr:1992:CCV

[1097] Ernst W. Mayr and Ashok Subramanian. The complexity of circuit value and network stability. Journal of Computer and System Sciences, 44(2):302–323, April 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290024D.

Lam:1992:RCC

[1098] Tak Wah Lam and Walter L. Ruzzo. Results on communication complexity classes. Journal of Computer and System Sciences, 44(2):324– 342, April 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290025E.

Li:1992:IRK

[1099] Ming Li and Paul M. B. Vitányi.

Inductive reasoning and Kolmogorov complexity. Journal of Computer and System Sciences, 44(2):343– 384, April 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290026F.

Lindell:1992:IPB

[1100] Steven Lindell. The invariant problem for binary string structures and the parallel complexity theory of queries. Journal of Computer and System Sciences, 44(3): 385–410, June 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290010G.

Klapper:1992:DEA

[1101] A. Klapper. Distributed event algebras. Journal of Computer and System Sciences, 44(3):411– 424, June 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000092900117.

Kaminski:1992:FAD

[1102] Michael Kaminski and Shlomit S. Pinter. Finite automata on directed graphs. Journal of Computer and System Sciences, 44(3): 425–446, June 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000092900128.

Silva:1992:BFS

[1103] Manuel Silva and Tadao Murata. B-fairness and structural B-fairness in Petri net models of concurrent systems. Journal of Computer and System Sciences, 44(3): 447–477, June 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0022000092900139.

Barrington:1992:RLN

[1104] David A. Mix Barrington, Kevin Compton, Howard Straubing, and Denis Thérien. Regular languages in NC 1. Journal of Computer and System Sciences, 44(3):478– 499, June 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290014A.

Watanabe:1992:PTO

[1105] Osamu Watanabe. On polynomial time one-truth-table reducibility to a sparse set. Journal of Computer and System Sciences, 44(3): 500–516, June 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290015B.

Anonymous:1992:AIVa

[1106] Anonymous. Author index for volume 44. Journal of Computer and System Sciences, 44(3): 517, June 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290016C.

Wang:1992:DTV

[1107] Xin Wang and Edward K. Blum.

Discrete-time versus continuous-time models of neural networks. Journal of Computer and System Sciences, 45(1):1–19, August 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290038K.

Sontag:1992:FNI

[1108] Eduardo D. Sontag. Feedforward nets for interpolation and classification. Journal of Computer and System Sciences, 45(1): 20–48, August 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290039L.

Honsell:1992:ATT

[1109] Furio Honsell and Simonetta Ronchi Della Rocca. An approximation theorem for topological lambda models and the topological incompleteness of lambda calculus. Journal of Computer and System Sciences, 45 (1):49–75, August 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290040P.

Gyssens:1992:PAN

[1110] Marl Gyssens and Dirk Van Gucht. The powerset algebra as a natural tool to handle nested database relations. Journal of Computer and System Sciences, 45(1):76–103, August 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290041G.

Bar-Yehuda:1992:TCB

[1111] Reuven Bar-Yehuda, Oded Goldreich, and Alon Itai. On the time-complexity of broadcast in multi-hop radio networks: an exponential gap between determinism and randomization. Journal of Computer and System Sciences, 45(1):104–126, August 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290042H.

Stewart:1992:UHP

[1112] Iain A. Stewart. Using the Hamiltonian path operator to capture NP. Journal of Computer and System Sciences, 45(1):127–151, August 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290043I.

Papadimitriou:1992:EF

[1113] Christos H. Papadimitriou. Editor’s foreword. Journal of Computer and System Sciences, 45(2): 153, October 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290044J.

Birget:1992:PST

[1114] Jean-Camille Birget. Positional simulation of two-way automata: Proof of a conjecture of R. Kannan and generalizations. Journal of Computer and System Sciences, 45(2): 154–179, October 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290045K.

Borodin:1992:LBL

[1115] Allan Borodin, Walter L. Ruzzo, and Martin Tompat. Lower bounds on the length of universal traversal sequences. Journal of Computer and System Sciences, 45(2): 180–203, October 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290046L.

Babai:1992:MPP

[1116] László Babai, Noam Nisant, and Márió Szegedy. Multiparty protocols, pseudorandom generators for logspace, and time-space trade-offs. Journal of Computer and System Sciences, 45 (2):204–232, October 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290047M.

Feder:1992:NFP

[1117] Tomás Feder. A new fixed point approach for stable networks and stable marriages. Journal of Computer and System Sciences, 45(2): 233–284, October 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290048N.

Regan:1992:MCP

[1118] Kenneth W. Regan. Minimumcomplexity pairing functions. Journal of Computer and System Sciences, 45 (3):285–295, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290027G.

Lam:1992:TOB

[1119] Tak Lam, Prasoon Tiwari, and Martin Tompa. Trade-offs between communication and space. Journal of Computer and System Sciences, 45(3): 296–315, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290028H.

Weil:1992:CVL

[1120] Pascal Weil. Closure of varieties of languages under products with counter. Journal of Computer and System Sciences, 45(3): 316–339, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009290029I.

Bettayeb:1992:EGH

[1121] Said Bettayeb, Zevi Miller, and I. Hal Sudborough. Embedding grids into hypercubes. Journal of Computer and System Sciences, 45(3): 340–366, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290030M.

Carlstrom:1992:FA

[1122] Ian F. Carlstrom. Foreground automata. Journal of Computer and System Sciences, 45(3):367– 384, December 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290031D.

Downey:1992:NTP

[1123] Rod Downey. Nondiamond theorems for polynomial time reducibility. Journal of Computer and System Sciences, 45(3):385–395, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/002200009290032E.

Krob:1992:MKR

[1124] Daniel Krob. Models of a K-rational identity system. Journal of Computer and System Sciences, 45(3): 396–434, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290033F.

Yuan:1992:DDR

[1125] Li-Yan Yuan and Z. Meral Ozsoyoglu. Design of desirable relational database schemes. Journal of Computer and System Sciences, 45(3): 435–470, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290034G.

Obradovic:1992:CDM

[1126] Zoran Obradovic and Ian Parberry. Computing with discrete multivalued neurons. Journal of Computer and System Sciences, 45(3): 471–492, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290035H.

Schmid:1992:SIF

[1127] U. Schmid and J. Blieberger. Some investigations on FCFS scheduling in hard real time applications. Journal of Computer and System Sciences, 45 (3):493–512, December 1992. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290036I.

Anonymous:1992:AIVb

[1128] Anonymous. Author index for volume 45. Journal of Computer and System Sciences, 45(3): 513, December 1992. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009290037J.

Howell:1993:NSP

[1129] Rodney R. Howell, Louis E. Rosier, and Hsu-Chun Yen. Normal and sinkless petri nets. Journal of Computer and System Sciences, 46(1):1– 26, February 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390046Y.

Jansen:1993:IP

[1130] Klaus Jansen. The interconnection problem. Journal of Computer and System Sciences, 46(1):27– 38, February 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390047Z.

Ziarko:1993:VPR

[1131] Wojciech Ziarko. Variable precision rough set model. Journal of Computer and System Sciences, 46 (1):39–59, February 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000093900482.

Janssens:1993:CGA

[1132] D. Janssens, M. Lens, and G. Rozenberg. Computation graphs for actor grammars. Journal of Computer and System Sciences, 46(1):60– 90, February 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000093900493.

Watanabe:1993:MCA

[1133] Toshimasa Watanabe and Akira Nakamura. A minimum 3-connectivity augmentation of a graph. Journal of Computer and System Sciences, 46 (1):91–128, February 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000093900507.

Hajnal:1993:TCB

[1134] András Hajnal, Wolfgang Maass, Pavel Pudlák, Márió Szegedy, and György Turán. Threshold circuits of bounded depth. Journal of Computer and System Sciences, 46(2): 129–154, April 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390001D.

Dassow:1993:DSA

[1135] J. Dassow and H. Jürgensen. Deterministic soliton automata with at most one cycle. Journal of Computer and System Sciences, 46(2): 155–197, April 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390002E.

Venkatesh:1993:DDN

[1136] Santosh S. Venkatesh. Directed drift: a new linear threshold algorithm for learning binary weights on-line. Journal of Computer and System Sciences, 46(2):198–217, April 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390003F.

Courcelle:1993:HRH

[1137] Bruno Courcelle, Joost Engelfriet, and Grzegorz Rozenberg. Handle-rewriting hypergraph grammars. Journal of Computer and System Sciences, 46 (2):218–270, April 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390004G.

Cohen:1993:EPT

[1138] Joëlle Cohen, Dominique Perrin, and Jean-Eric Pin. On the expressive power of temporal logic. Journal of Computer and System Sciences, 46 (3):271–294, June 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390005H.

Ogiwara:1993:CTF

[1139] Mitsunori Ogiwara and Lane A.

Hemachandra. A complexity theory for feasible closure properties. Journal of Computer and System Sciences, 46(3):295–325, June 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390006I.

Ravi:1993:GTA

[1140] S. S. Ravi and Errol L. Lloyd.

Graph theoretic analysis of PLA folding heuristics. Journal of Computer and System Sciences, 46(3): 326–348, June 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390007J.

Goldsmith:1993:NBI

[1141] Judy Goldsmith, Deborah Joseph, and Paul Young. A note on bi-immunity and p-closeness of pcheatable sets in P/poly. Journal of Computer and System Sciences, 46(3):349–362, June 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390008K.

Hemachandra:1993:CDS

[1142] Lane A. Hemachandra and Albrecht Hoeno. Collapsing degrees via strong computation. Journal of Computer and System Sciences, 46(3): 363–380, June 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390009L.

Bloom:1993:MMIa

[1143] Stephen L. Bloom and Zoltán Ésik. Matrix and matricial iteration theories, Part I. Journal of Computer and System Sciences, 46(3): 381–408, June 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390010T.

Bloom:1993:MMIb

[1144] Stephen L. Bloom and Zoltán Ésik. Matrix and matricial iteration theories, Part II. Journal of Computer and System Sciences, 46(3):

409–439, June 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390011K.

Anonymous:1993:AIVa

[1145] Anonymous. Author index for volume 46. Journal of Computer and System Sciences, 46(3): 440, June 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390012L.

Chandra:1993:EF

[1146] Ashok K. Chandra. Editor’s foreword.

Journal of Computer and System Sciences, 47(1):1, August 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390018R.

Cohen:1993:WSP

[1147] Simona R. Cohen and Ouri Wolfson.

Why a single parallelization strategy is not enough in knowledge bases. Journal of Computer and System Sciences, 47(1):2–44, August 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390019S.

Johnson:1993:BTI

[1148] Theodore Johnson and Dennis Shasha.

B-trees with inserts and deletes: Why free-at-empty is better than merge-at-half. Journal of Computer and System Sciences, 47(1):

45–76, August 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390020W.

Kifer:1993:LPC

[1149] Michael Kifer and James Wu. A logic for programming with complex objects. Journal of Computer and System Sciences, 47(1): 77–120, August 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390021N.

Hull:1993:ACQ

[1150] Richard Hull and Jianwen Su. Algebraic and calculus query languages for recursively typed complex objects. Journal of Computer and System Sciences, 47(1): 121–156, August 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390022O.

Weihl:1993:IRC

[1151] William E. Weihl. The impact of recovery on concurrency control. Journal of Computer and System Sciences, 47(1):157–184, August 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390023P.

VanGelder:1993:AFL

[1152] Allen Van Gelder. The alternating fixpoint of logic programs with negation. Journal of Computer and System Sciences, 47(1): 185–221, August 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390024Q.

Ramakrishnan:1993:LQO

[1153] Raghu Ramakrishnan, Yehoshua Sagiv, Jeffrey D. Ullman, and Moshe Y. Vardi. Logical query optimization by proof-tree transformation. Journal of Computer and System Sciences, 47 (1):222–248, August 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390025R.

Kozen:1993:EF

[1154] Dexter Kozen. Editor’s foreword. Journal of Computer and System Sciences, 47(2):249, October 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390032R.

Luby:1993:RRP

[1155] Michael Luby. Removing randomness in parallel computation without a processor penalty. Journal of Computer and System Sciences, 47 (2):250–286, October 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390033S.

McAllester:1993:NFS

[1156] David McAllester, Prakash Panangaden, and Vasant Shanbhogue.

Nonexpressibility of fairness and signaling. Journal of Computer and System Sciences, 47(2):287– 321, October 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390034T.

Lovasz:1993:CCC

[1157] László Lovasz and Michael Saks. Communication complexity and combinatorial lattice theory. Journal of Computer and System Sciences, 47 (2):322–349, October 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390035U.

Komlos:1993:ECA

[1158] János Komlós and Ramamohan Paturi. Effect of connectivity in an associative memory model. Journal of Computer and System Sciences, 47 (2):350–373, October 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390036V.

Gabow:1993:EF

[1159] Harold N. Gabow. Editor’s foreword. Journal of Computer and System Sciences, 47(3):375, December 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/002200009390037W.

Haastad:1993:DLM

[1160] J. Håstad, A. W. Schrift, and A. Shamir. The discrete logarithm modulo a composite hides 0(n) bits. Journal of Computer and System Sciences, 47(3):376– 404, December 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390038X.

Szegedy:1993:FBS

[1161] Mario Szegedy. Functions with bounded symmetric communication complexity, programs over commutative monoids, and ACC. Journal of Computer and System Sciences, 47(3): 405–423, December 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390039Y.

Fredman:1993:SIT

[1162] Michael L. Fredman and Dan E.

Willard. Surpassing the information theoretic bound with fusion trees. Journal of Computer and System Sciences, 47(3):424– 436, December 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0022000093900404.

Mulmuley:1993:OSD

[1163] Ketan Mulmuley. Output sensitive and dynamic constructions of higher order Voronoi diagrams and levels in arrangements. Journal of Computer and System Sciences, 47(3): 437–458, December 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390041T.

Kao:1993:TOT

[1164] Ming-Yang Kao and Philip N. Klein. Towards overcoming the transitiveclosure bottleneck: Efficient parallel algorithms for planar digraphs. Journal of Computer and System Sciences, 47 (3):459–500, December 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390042U.

Cypher:1993:DSN

[1165] Robert Cypher and C. Greg Plaxton. Deterministic sorting in nearly logarithmic time on the hypercube and related computers. Journal of Computer and System Sciences, 47(3): 501–548, December 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/002200009390043V.

Blum:1993:STC

[1166] Manuel Blum, Michael Luby, and Ronitt Rubinfeld. Self-testing/correcting with applications to numerical problems. Journal of Computer and System Sciences, 47(3):549–595, December 1993. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/002200009390044W.

Anonymous:1993:AIVb

[1167] Anonymous. Author index for volume 47. Journal of Computer and System Sciences, 47(3):

596, December 1993. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/002200009390045X.

Anonymous:1994:EBa

[1168] Anonymous. Editorial Board. Journal of Computer and System Sciences, 48(1):i, February 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800194.

Gereb-Graus:1994:TOW

[1169] Mihály Geréb-Graus and Ming Li.

Three one-way heads cannot do string matching. Journal of Computer and System Sciences, 48(1): 1–8, February 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800200.

Hegner:1994:UCD

[1170] Stephen J. Hegner. Unique complements and decompositions of database schemata. Journal of Computer and System Sciences, 48(1):9– 57, February 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800212.

Engelfriet:1994:HLB

[1171] Joost Engelfriet and Linda Heyker.

Hypergraph languages of bounded degree. Journal of Computer and System Sciences, 48(1):58–89, February 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800224.

Blelloch:1994:PSG

[1172] Guy E. Blelloch and James J. Little. Parallel solutions to geometric problems in the scan model of computation. Journal of Computer and System Sciences, 48(1): 90–115, February 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800236.

Fenner:1994:GDC

[1173] Stephen A. Fenner, Lance J. Fortnow, and Stuart A. Kurtz. Gapdefinable counting classes. Journal of Computer and System Sciences, 48 (1):116–148, February 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800248.

Kari:1994:RSP

[1174] Jarkko Kari. Reversibility and surjectivity problems of cellular automata. Journal of Computer and System Sciences, 48(1):149– 182, February 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000580025X.

Cai:1994:PPT

[1175] Jin yi Cai, Anne Condon, and Richard J. Lipton. PSPACE is provable by two provers in one round. Journal of Computer and System Sciences, 48 (1):183–193, February 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800261.

Alon:1994:SDO

[1176] Noga Alon and Pavel Pudlak. Superconcentrators of depths 2 and 3; odd levels help (rarely). Journal of Computer and System Sciences, 48 (1):194–202, February 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800273.

Li:1994:LFP

[1177] Leong Kwan Li. Learning fixed point patterns by recurrent networks. Journal of Computer and System Sciences, 48(2):203–213, April 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800017.

Berkman:1994:FLA

[1178] Omer Berkman and Uzi Vishkin. Finding level-ancestors in trees. Journal of Computer and System Sciences, 48(2):214–230, April 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800029.

Dietzfelbinger:1994:ELT

[1179] Martin Dietzfelbinger, Miroslaw Kutylowski, and Rüdiger Reischuk. Exact lower time bounds for computing Boolean functions on CREW PRAMs. Journal of Computer and System Sciences, 48(2):231– 254, April 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800030.

Cadoli:1994:CPC

[1180] Marco Cadoli and Maurizio Lenzerini. The complexity of propositional closed world reasoning and circumscription. Journal of Computer and System Sciences, 48(2): 255–310, April 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800042.

Benedek:1994:NL

[1181] Gyora M. Benedek and Alon Itai. Nonuniform learnability. Journal of Computer and System Sciences, 48 (2):311–323, April 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800054.

Homer:1994:RNS

[1182] Steven Homer and Luc Longpré. On reductions of NP sets to sparse sets. Journal of Computer and System Sciences, 48(2):324–336, April 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000005800066.

Stinson:1994:CTU

[1183] D. R. Stinson. Combinatorial techniques for universal hashing. Journal of Computer and System Sciences, 48 (2):337–346, April 1, 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800078.

Siu:1994:PLA

[1184] Kai-Yeung Siu, Amir Dembo, and Thomas Kailath. On the perceptron learning algorithm on data with high precision. Journal of Computer and System Sciences, 48(2): 347–356, April 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000580008X.

Selman:1994:TCC

[1185] Alan L. Selman. A taxonomy of complexity classes of functions. Journal of Computer and System Sciences, 48(2):357–381, April 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800091.

Yannakakis:1994:EF

[1186] Mihalis Yannakakis. Editor’s foreword.

Journal of Computer and System Sciences, 48(3):383, June 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800583.

Bern:1994:PGM

[1187] Marshall Bern, David Eppstein, and John Gilbert. Provably good mesh generation. Journal of Computer and System Sciences, 48(3): 384–409, June 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800595.

Fiat:1994:CSA

[1188] Amos Fiat, Yuval Rabani, and Yiftach Ravid. Competitive kserver algorithms. Journal of Computer and System Sciences, 48(3): 410–428, June 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800601.

Neff:1994:SPP

[1189] C. Andrew Neff. Specified precision polynomial root isolation is in NC. Journal of Computer and System Sciences, 48(3):429– 463, June 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800613.

Kearns:1994:EDF

[1190] Michael J. Kearns and Robert E. Schapire. Efficient distribution-free learning of probabilistic concepts. Journal of Computer and System Sciences, 48(3):464–497, June 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000005800625.

Papadimitriou:1994:CPA

[1191] Christos H. Papadimitriou. On the complexity of the parity argument and other inefficient proofs of existence. Journal of Computer and System Sciences, 48(3): 498–532, June 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800637.

Fredman:1994:TDA

[1192] Michael L. Fredman and Dan E.

Willard. Trans-dichotomous algorithms for minimum spanning trees and shortest paths. Journal of Computer and System Sciences, 48(3): 533–551, June 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800649.

Anonymous:1994:AIVa

[1193] Anonymous. Author index for volume 48. Journal of Computer and System Sciences, 48(3): 552, June 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800650.

Anonymous:1994:EBb

[1194] Anonymous. Editorial Board. Journal of Computer and System Sciences, 49(1):i, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800819.

Toda:1994:SCP

[1195] Seinosuke Toda. Simple characterizations of P(#P) and complete problems. Journal of Computer and System Sciences, 49(1): 1–17, August 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800820.

Sato:1994:GSL

[1196] Tadakazu Sato. Group structured linear cellular automata over Zm. Journal of Computer and System Sciences, 49(1):18–23, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800832.

Chang:1994:ROH

[1197] Richard Chang, Benny Chor, Oded Goldreich, Juris Hartmanis, Johan Håstad, Desh Ranjan, and Pankaj Rohatgi. The random oracle hypothesis is false. Journal of Computer and System Sciences, 49(1): 24–39, August 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800844.

Hattori:1994:TET

[1198] Takashi Hattori. Time-extraction for temporal logic-logic programming and local process time. Journal of Computer and System Sciences, 49 (1):40–56, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800856.

Wanke:1994:CCP

[1199] Egon Wanke. The complexity of connectivity problems on contextfree graph languages. Journal of Computer and System Sciences, 49 (1):57–82, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800868.

Pippenger:1994:SSC

[1200] Nicholas Pippenger. Symmetry in selfcorrecting cellular automata. Journal of Computer and System Sciences, 49(1):83–95, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000580087X.

Jain:1994:CLI

[1201] Sanjay Jain and Arun Sharma. Characterizing language identification by standardizing operations. Journal of Computer and System Sciences, 49 (1):96–107, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800881.

Raghavan:1994:BDG

[1202] Vijay Raghavan. Bounded degree graph inference from walks. Journal of Computer and System Sciences, 49 (1):108–132, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800893.

Fachini:1994:KLC

[1203] E. Fachini and A. Monti. A kleenelike characterization of languages accepted by systolic tree automata. Journal of Computer and System Sciences, 49(1):133–147, August 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000580090X.

Nisan:1994:HVR

[1204] Noam Nisan and Avi Wigderson. Hardness vs randomness. Journal of Computer and System Sciences, 49(2): 149–167, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800431.

Nishida:1994:ISS

[1205] Taishin Nishida and Youichi Kobuchi. Invariant sets for substitution. Journal of Computer and System Sciences, 49 (2):168–174, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800443.

Krishnakumar:1994:CSD

[1206] Narayanan Krishnakumar. On computing serial dependency relations. Journal of Computer and System Sciences, 49(2):175–188, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800455.

Case:1994:VLN

[1207] John Case, Sanjay Jain, and Arun Sharma. Vacillatory learning of nearly minimal size grammars. Journal of Computer and System Sciences, 49 (2):189–207, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800467.

Amir:1994:DDM

[1208] Amihood Amir, Martin Farach, Zvi Galil, Raffaele Giancarlo, and Kunsoo Park. Dynamic dictionary matching. Journal of Computer and System Sciences, 49(2):208– 222, October 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800479.

Kaminski:1994:BTL

[1209] Michael Kaminski. A branching time logic with past operators. Journal of Computer and System Sciences, 49 (2):223–246, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800480.

Karchmer:1994:NDC

[1210] Mauricio Karchmer, Ilan Newman, Mike Saks, and Avi Wigderson. Nondeterministic communication complexity with few witnesses. Journal of Computer and System Sciences, 49 (2):247–257, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800492.

Engelfriet:1994:TPT

[1211] Joost Engelfriet and Heiko Vogler. The translation power of top-down tree-to-graph transducers. Journal of Computer and System Sciences, 49 (2):258–305, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800509.

Kozen:1994:EIP

[1212] Dexter Kozen, Jens Palsberg, and Michael I. Schwartzbach. Efficient inference of partial types. Journal of Computer and System Sciences, 49 (2):306–324, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800510.

Kantabutra:1994:LTN

[1213] Vitit Kantabutra. A linear-time nearoptimum-length triangulation algorithm for convex polygons. Journal of Computer and System Sciences, 49 (2):325–333, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800522.

You:1994:TVS

[1214] Jia-Huai You and Li Yan Yuan. A three-valued semantics for deductive databases and logic programs. Journal of Computer and System Sciences, 49(2):334–361, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800534.

Salomaa:1994:MNP

[1215] Kai Salomaa and Sheng Yu. Measures of nondeterminism for pushdown automata. Journal of Computer and System Sciences, 49(2): 362–374, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800546.

Obradovic:1994:LDM

[1216] Zoran Obradović and Ian Parberry. Learning with discrete multivalued neurons. Journal of Computer and System Sciences, 49(2): 375–390, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800558.

Sugihara:1994:RGW

[1217] Kokichi Sugihara. Robust gift wrapping for the three-dimensional convex hull. Journal of Computer and System Sciences, 49(2):391– 407, October 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000580056X.

Goldberg:1994:LGS

[1218] Leslie Ann Goldberg. Listing graphs that satisfy first-order sentences. Journal of Computer and System Sciences, 49(2):408–424, October 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800571.

Galil:1994:EF

[1219] Zvi Galil. Editor’s foreword. Journal of Computer and System Sciences, 49(3):425, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800662.

Mulmuley:1994:EAH

[1220] Ketan Mulmuley. An efficient algorithm for hidden surface removal, II. Journal of Computer and System Sciences, 49(3):427– 453, December 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800674.

Berger:1994:ENA

[1221] Bonnie Berger, John Rompel, and Peter W. Shor. Efficient NC algorithms for set cover with applications to learning and geometry. Journal of Computer and System Sciences, 49(3): 454–477, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800686.

Motwani:1994:PMY

[1222] Rajeev Motwani, Joseph (Seffi) Naor, and Moni Naor. The probabilistic method yields deterministic parallel algorithms. Journal of Computer and System Sciences, 49(3): 478–516, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800698.

Ramachandran:1994:PTP

[1223] Vijaya Ramachandran and John Reif.

Planarity testing in parallel. Journal of Computer and System Sciences, 49 (3):517–561, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800704.

Ajtai:1994:DVF

[1224] Miklos Ajtai and Yuri Gurevich. Datalog vs first-order logic. Journal of Computer and System Sciences, 49(3): 562–588, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800716.

Fulk:1994:OPS

[1225] Mark Fulk, Sanjay Jain, and Daniel N.

Osherson. Open problems in “systems that learn”. Journal of Computer and System Sciences, 49(3): 589–604, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800728.

Cai:1994:HTD

[1226] Jin yi Cai and Juris Hartmanis.

On Hausdorff and topological dimensions of the Kolmogorov complexity of the real line. Journal of Computer and System Sciences, 49(3): 605–619, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000580073X.

Qian:1994:ALP

[1227] Zhenyu Qian. Another look at parameterization for order-sorted algebraic specifications. Journal of Computer and System Sciences, 49(3): 620–666, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800741.

Boyar:1994:ERC

[1228] Joan Boyar and Kim S. Larsen. Efficient rebalancing of chromatic search trees. Journal of Computer and System Sciences, 49(3): 667–682, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005800753.

Levene:1994:NUR

[1229] Mark Levene and George Loizou. The nested universal relation data model. Journal of Computer and System Sciences, 49(3):683–717, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000005800765.

Durand:1994:SPC

[1230] Bruno Durand. The surjectivity problem for 2d cellular automata. Journal of Computer and System Sciences, 49 (3):718–725, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800777.

Ito:1994:IPF

[1231] Minoru Ito and Grant E. Weddell.

Implication problems for functional constraints on databases supporting complex objects. Journal of Computer and System Sciences, 49(3): 726–768, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800789.

Fellows:1994:SDE

[1232] Michael R. Fellows and Michael A.

Langston. On search, decision, and the efficiency of polynomialtime algorithms. Journal of Computer and System Sciences, 49(3): 769–779, December 1994. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800790.

Anonymous:1994:AIVb

[1233] Anonymous. Author index for volume 49. Journal of Computer and System Sciences, 49(3):780– 781, December 1994. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005800807.

Kalpakis:1995:PLA

[1234] K. Kalpakis and Y. Yesha. On the power of the linear array architecture for performing tree-structured computations. Journal of Computer and System Sciences, 50(1):1– 10, February 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200008571001X.

Karhumaki:1995:PST

[1235] J. Karhumaki, W. Plandowski, and W. Rytter. Polynomial size test sets for context-free languages. Journal of Computer and System Sciences, 50 (1):11–19, February 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710021.

Goldman:1995:CT

[1236] S. A. Goldman and M. J. Kearns. On the complexity of teaching. Journal of Computer and System Sciences, 50 (1):20–31, February 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710033.

Blum:1995:LPF

[1237] A. Blum, L. Hellerstein, and N. Littlestone. Learning in the presence of finitely or infinitely many irrelevant attributes. Journal of Computer and System Sciences, 50(1):32– 40, February 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710045.

Kinber:1995:LMC

[1238] E. Kinber, C. H. Smith, M. Velauthapillai, and R. Wiehagen. On learning multiple concepts in parallel. Journal of Computer and System Sciences, 50(1):41–52, February 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710057.

Jiang:1995:DPP

[1239] T. Jiang, A. Salomaa, K. Salomaa, and S. Yu. Decision problems for patterns. Journal of Computer and System Sciences, 50(1):53– 63, February 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710069.

Parberry:1995:LSP

[1240] I. Parberry. Load sharing with parallel priority queues. Journal of Computer and System Sciences, 50 (1):64–73, February 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710070.

Halpern:1995:EF

[1241] J. Halpern, B. Awerbuch, M. Benor, A. Chandra, J. Feigenbaum, J. Vonzurgathen, L. Guibas, L. Pitt, M. Saks, D. Shmoys, D. Sleator, E. Upfal, U. Vazirani, and A. Yao. Editors’ foreword. Journal of Computer and System Sciences, 50(2): 189, April 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710161.

Beigel:1995:PCU

[1242] R. Beigel, N. Reingold, and D. Spielman. PP is closed under intersection. Journal of Computer and System Sciences, 50(2):191– 202, April 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710173.

Matousek:1995:AOG

[1243] J. Matousek. Approximations and optimal geometric divide-and-conquer. Journal of Computer and System Sciences, 50(2):203–208, April 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000085710185.

Yannakakis:1995:TFS

[1244] M. Yannakakis and D. Lee. Testing finite state machines: Fault detection. Journal of Computer and System Sciences, 50(2):209– 227, April 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710197.

Leighton:1995:FAA

[1245] T. Leighton, F. Makedon, S. Plotkin, C. Stein, E. Tardos, and S. Tragoudas. Fast approximation algorithms for multicommodity flow problems. Journal of Computer and System Sciences, 50(2):228–243, April 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710203.

Borodin:1995:CPL

[1246] A. Borodin, S. Irani, P. Raghavan, and B. Schieber. Competitive paging with locality of reference. Journal of Computer and System Sciences, 50(2):244–258, April 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710215.

Gabow:1995:MAF

[1247] H. N. Gabow. A matroid approach to finding edge connectivity and packing arborescences. Journal of Computer and System Sciences, 50(2): 259–273, April 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710227.

Kaltofen:1995:ENI

[1248] E. Kaltofen. Effective Noether irreducibility forms and applications. Journal of Computer and System Sciences, 50(2):274–295, April 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710239.

Babai:1995:FMC

[1249] L. Babai, G. Cooperman, L. Finkelstein, E. Luks, and A. Seress. Fast Monte Carlo algorithms for permutation groups. Journal of Computer and System Sciences, 50(2):

296–308, April 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710240.

Abiteboul:1995:CFO

[1250] S. Abiteboul and V. Vianu. Computing with first-order logic. Journal of Computer and System Sciences, 50(2):309–335, April 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710252.

Angluin:1995:WWM

[1251] Dana Angluin and Michael Kharitonov. When won’t membership queries help? Journal of Computer and System Sciences, 50(2):336– 355, April 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710264. 23rd

Symposium on the Theory of Computing (New Orleans, LA, 1991).

Immerman:1995:EF

[1252] N. Immerman. Editor’s foreword. Journal of Computer and System Sciences, 50(3):357, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710276.

Chang:1995:UST

[1253] R. Chang, J. Kadin, and P. Rohatgi. On unique satisfiability and the threshold behavior of randomized reductions. Journal of Computer and System Sciences, 50(3): 359–373, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710288.

Barrington:1995:SLB

[1254] D. A. M. Barrington and H. Straubing.

Superlinear lower bounds for boundedwidth branching programs. Journal of Computer and System Sciences, 50(3):374–381, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200008571029X.

Bovet:1995:CCS

[1255] D. P. Bovet, P. Crescenzi, and R. Silvestri. Complexity classes and sparse oracles. Journal of Computer and System Sciences, 50(3): 382–390, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710306.

Kolaitis:1995:APN

[1256] P. G. Kolaitis and M. N. Thakur. Approximation properties of NP minimization classes. Journal of Computer and System Sciences, 50(3): 391–411, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710318.

Gupta:1995:CPW

[1257] S. Gupta. Closure properties and witness reduction. Journal of Computer and System Sciences, 50(3): 412–432, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200008571032X.

Grigni:1995:MSL

[1258] M. Grigni and M. Sipser. Monotone separation of logarithmic space from logarithmic depth. Journal of Computer and System Sciences, 50 (3):433–437, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710331.

Long:1995:EF

[1259] T. Long. Editor’s foreword. Journal of Computer and System Sciences, 50(3):439, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710343.

Beaudry:1995:CMN

[1260] M. Beaudry and P. Mckenzie. Circuits, matrices, and nonassociative computation. Journal of Computer and System Sciences, 50(3): 441–455, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710355.

Green:1995:PMB

[1261] F. Green, J. Kobler, K. W. Regan, T. Schwentick, and J. Toran. The power of the middle bit of a #P function. Journal of Computer and System Sciences, 50(3): 456–467, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710367.

Gavalda:1995:BCA

[1262] R. Gavalda. Bounding the complexity of advice functions. Journal of Computer and System Sciences, 50 (3):468–475, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710379.

Rohatgi:1995:SQR

[1263] P. Rohatgi. Saving queries with randomness. Journal of Computer and System Sciences, 50(3): 476–492, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710380.

Saluja:1995:DCP

[1264] S. Saluja, K. V. Subrahmanyam, and M. N. Thakur. Descriptive complexity of #P functions. Journal of Computer and System Sciences, 50 (3):493–505, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710392.

Condon:1995:IPS

[1265] A. Condon and R. Ladner. Interactive proof systems with polynomially bounded strategies. Journal of Computer and System Sciences, 50 (3):506–518, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710409.

Li:1995:EF

[1266] M. Li. Editor’s foreword. Journal of Computer and System Sciences, 50(3):519, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710410.

Bshouty:1995:LBR

[1267] N. H. Bshouty, T. R. Hancock, and L. Hellerstein. Learning Boolean read-once formulas over generalized bases. Journal of Computer and System Sciences, 50(3): 521–542, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710422.

Mansour:1995:LAD

[1268] Y. Mansour. An O(n^{log log n}) learning algorithm for DNF under the uniform distribution. Journal of Computer and System Sciences, 50 (3):543–550, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710434.

Helmbold:1995:WL

[1269] D. P. Helmbold and M. K. Warmuth. On weak learning. Journal of Computer and System Sciences, 50(3):551–573, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710446.

Daley:1995:BPB

[1270] R. Daley, B. Kalyansundaram, and M. Velauthapillai. Breaking the probability 12 barrier in FIN-type learning. Journal of Computer and System Sciences, 50(3):574– 599, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710458.

Kharitonov:1995:CLB

[1271] M. Kharitonov. Cryptographic lower bounds for learnability of Boolean functions on the uniform distribution. Journal of Computer and System Sciences, 50(3):600–610, June 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200008571046X.

Anonymous:1995:AIVa

[1272] Anonymous. Author index for volume 50. Journal of Computer and System Sciences, 50(3): 611, June 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710471.

Sagiv:1995:EF

[1273] Y. Sagiv. Editor’s foreword. Journal of Computer and System Sciences, 51(1):1, August 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710483.

Kabanza:1995:HIT

[1274] F. Kabanza, J. M. Stevenne, and P. Wolper. Handling infinite temporal data. Journal of Computer and System Sciences, 51(1): 3–17, August 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710495.

Lipton:1995:QSE

[1275] R. J. Lipton and J. F. Naughton. Query size estimation by adaptive sampling. Journal of Computer and System Sciences, 51(1): 18–25, August 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710501.

Kanellakis:1995:CQL

[1276] P. C. Kanellakis, G. M. Kuper, and P. Z. Revesz. Constraint query languages. Journal of Computer and System Sciences, 51(1): 26–52, August 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710513.

Agrawal:1995:CSL

[1277] D. Agrawal and A. Elabbadi. Constrained shared locks for increasing concurrency in databases. Journal of Computer and System Sciences, 51(1):53–63, August 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710525.

Schlipf:1995:EPL

[1278] J. S. Schlipf. The expressive powers of the logic programming semantics. Journal of Computer and System Sciences, 51(1):64–86, August 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000085710537.

Saraiya:1995:ETD

[1279] Y. P. Saraiya. On the efficiency of transforming database logic programs. Journal of Computer and System Sciences, 51(1): 87–109, August 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710549.

Kolaitis:1995:EPD

[1280] P. G. Kolaitis and M. Y. Vardi.

On the expressive power of Datalog: Tools and a case study. Journal of Computer and System Sciences, 51 (1):110–134, August 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710550.

Mendelzon:1995:EF

[1281] A. O. Mendelzon. Editor’s foreword. Journal of Computer and System Sciences, 51(2):135, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000085710562.

Seshadri:1995:ESR

[1282] S. Seshadri and J. F. Naughton. On the expected size of recursive Datalog queries. Journal of Computer and System Sciences, 51(2): 137–148, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710574.

Grumbach:1995:TQL

[1283] S. Grumbach and V. Vianu. Tractable query languages for complex object databases. Journal of Computer and System Sciences, 51(2): 149–167, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710586.

Johnson:1995:TUB

[1284] D. B. Johnson and L. Raab. A tight upper bound on the benefits of replica control protocols. Journal of Computer and System Sciences, 51 (2):168–176, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710598.

Afrati:1995:DVP

[1285] F. Afrati, S. S. Cosmadakis, and M. Yannakakis. On Datalog vs polynomial time. Journal of Computer and System Sciences, 51(2):

177–196, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710604.

Helm:1995:SQO

[1286] R. Helm, K. Marriott, and M. Odersky. Spatial query optimization: From Boolean constraints to range queries. Journal of Computer and System Sciences, 51(2):197– 210, October 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710616.

Garciamolina:1995:NDQ

[1287] H. Garciamolina and K. Salem. Nondeterministic queue operations. Journal of Computer and System Sciences, 51(2):211–222, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710628.

Beeri:1995:SO

[1288] C. Beeri and T. Milo. Subtyping in OODBs. Journal of Computer and System Sciences, 51(2): 223–243, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200008571063X.

Ganguly:1995:EPD

[1289] S. Ganguly, S. Greco, and C. Zaniolo. Extrema predicates in deductive databases. Journal of Computer and System Sciences, 51(2):

244–259, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710641.

Feder:1995:CPG

[1290] T. Feder and R. Motwani. Clique partitions, graph compression and speeding-up algorithms. Journal of Computer and System Sciences, 51 (2):261–272, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710653.

Baliga:1995:LLS

[1291] G. Baliga, J. Case, and S. Jain. Language learning with some negative information. Journal of Computer and System Sciences, 51(2): 273–285, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710665.

Cole:1995:EAA

[1292] R. Cole and O. Zajicek. The expected advantage of asynchrony. Journal of Computer and System Sciences, 51(2):286–300, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710677.

Ali:1995:OAS

[1293] H. H. Ali and H. Elrewini. An optimal algorithm for scheduling interval ordered tasks with communication on N processors. Journal of Computer and System Sciences, 51 (2):301–306, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710689.

Grolmusz:1995:SCC

[1294] V. Grolmusz. Separating the communication complexities of MOD m and MOD p circuits. Journal of Computer and System Sciences, 51 (2):307–313, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710690.

Desel:1995:SPR

[1295] J. Desel and J. Esparza. Shortest paths in reachability graphs. Journal of Computer and System Sciences, 51 (2):314–323, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710707.

Blanchetsadri:1995:SLC

[1296] F. Blanchetsadri. Some logical characterizations of the dot-depth hierarchy and applications. Journal of Computer and System Sciences, 51 (2):324–337, October 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710719.

Maass:1995:EF

[1297] W. Maass. Editor’s foreword. Journal of Computer and System Sciences, 51(3):339, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710720.

Bartal:1995:CAD

[1298] Y. Bartal, A. Fiat, and Y. Rabani. Competitive algorithms for distributed data management. Journal of Computer and System Sciences, 51(3): 341–358, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710732.

Bartal:1995:NAA

[1299] Y. Bartal, A. Fiat, H. Karloff, and R. Vohra. New algorithms for an ancient scheduling problem. Journal of Computer and System Sciences, 51(3): 359–366, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710744.

Blum:1995:FLT

[1300] A. Blum and S. Rudich. Fast learning of k-term DNF formulas with queries. Journal of Computer and System Sciences, 51(3): 367–373, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710756.

Goodrich:1995:PSP

[1301] M. T. Goodrich. Planar separators and parallel polygon triangulation. Journal of Computer and System Sciences, 51 (3):374–389, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710768.

Kahn:1995:ES

[1302] J. Kahn and J. H. Kim. Entropy and sorting. Journal of Computer and System Sciences, 51(3): 390–399, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200008571077X.

Seidel:1995:APS

[1303] R. Seidel. On the all-pairs-shortest-path problem in unweighted undirected graphs. Journal of Computer and System Sciences, 51(3): 400–403, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710781.

Imielinski:1995:CTD

[1304] T. Imielinski, R. Vandermeyden, and K. V. Vadaparty. Complexity tailored design: a new design methodology for databases with incomplete information. Journal of Computer and System Sciences, 51(3): 405–432, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710793.

Abiteboul:1995:MS

[1305] S. Abiteboul, P. Kanellakis, S. Ramaswamy, and E. Waller. Method schemas. Journal of Computer and System Sciences, 51(3):433– 455, December 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200008571080X.

Seo:1995:CAA

[1306] C. Seo, J. Lim, and H. C. Eun. On comparison and analysis of algorithms for multiplication in GF(2^{m}). Journal of Computer and System Sciences, 51 (3):456–459, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710811.

Adamek:1995:CAR

[1307] J. Adamek, E. Nelson, and J. Reiterman. Continuous algebras revisited. Journal of Computer and System Sciences, 51(3):460– 471, December 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710823.

Skodinis:1995:EPE

[1308] K. Skodinis and E. Wanke. Emptiness problems of eNCE graph languages. Journal of Computer and System Sciences, 51(3):472– 485, December 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710835.

Vela:1995:BNR

[1309] C. R. Vela and A. Bahamonde. The best nondeterministic representations of finite orderings. Journal of Computer and System Sciences, 51(3): 486–494, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710847.

Benasher:1995:DTB

[1310] Y. Benasher and I. Newman. Decision trees with Boolean threshold queries. Journal of Computer and System Sciences, 51(3): 495–502, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710859.

Asano:1995:TAC

[1311] T. Asano. An O(n log log n) time algorithm for constructing a graph of maximum connectivity with prescribed degrees. Journal of Computer and System Sciences, 51(3): 503–510, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710860.

Creignou:1995:DTM

[1312] N. Creignou. A dichotomy theorem for maximum generalized satisfiability problems. Journal of Computer and System Sciences, 51(3): 511–522, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000085710872.

Chomicki:1995:FCT

[1313] J. Chomicki and D. Niwinski. On the feasibility of checking temporal integrity constraints. Journal of Computer and System Sciences, 51(3): 523–535, December 1995. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710884.

Anonymous:1995:AIVb

[1314] Anonymous. Author index for volume 51. Journal of Computer and System Sciences, 51(3): 536, December 1995. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000085710896.

Goodrich:1996:EF

[1315] Michael T. Goodrich and Chee Yap. Editors’ foreword. Journal of Computer and System Sciences, 52(1): 1, February 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009690001X.

Eppstein:1996:SBS

[1316] David Eppstein, Zvi Galil, Giuseppe F. Italiano, and Thomas H. Spencer. Separator based sparsification: I. Planarity testing and minimum spanning trees. Journal of Computer and System Sciences, 52(1):3– 27, February 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900021.

Baker:1996:PPM

[1317] Brenda S. Baker. Parameterized pattern matching: Algorithms and applications. Journal of Computer and System Sciences, 52(1):28– 42, February 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900033.

Nisan:1996:RLS

[1318] Noam Nisan and David Zuckerman.

Randomness is linear in space. Journal of Computer and System Sciences, 52(1):43–52, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900045.

Pippenger:1996:SRS

[1319] Nicholas Pippenger. Self-routing superconcentrators. Journal of Computer and System Sciences, 52(1):53– 60, February 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900057.

Cherubini:1996:PTP

[1320] Alessandra Cherubini and Pierluigi San Pietro. A polynomial-time parsing algorithm for K-depth languages. Journal of Computer and System Sciences, 52(1):61–79, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900069.

Chen:1996:SPS

[1321] Tsong Yueh Chen. On the structural properties of the set of fixpoints for nondeterministic recursive definitions. Journal of Computer and System Sciences, 52(1):80– 86, February 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900070.

LaPoutre:1996:LBU

[1322] Han La Poutré. Lower bounds for the union-find and the split-find problem on pointer machines. Journal of Computer and System Sciences, 52 (1):87–99, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900082.

Krishnamurthy:1996:FTS

[1323] Ravi Krishnamurthy, Raghu Ramakrishnan, and Oded Shmueli. A framework for testing safety and effective computability. Journal of Computer and System Sciences, 52(1): 100–124, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900094.

Libkin:1996:SRQ

[1324] Leonid Libkin and Limsoon Wong. Semantic representations and query languages for or-sets. Journal of Computer and System Sciences, 52(1): 125–142, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900100.

Astesiano:1996:FCD

[1325] Egidio Astesiano and Elena Zucca. A free construction of dynamic terms. Journal of Computer and System Sciences, 52(1):143–156, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000096900112.

Willard:1996:ARQ

[1326] Dan E. Willard. Applications of range query theory to relational data base join and selection operations. Journal of Computer and System Sciences, 52 (1):157–169, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900124.

Ehrenfeucht:1996:FLR

[1327] Andrzej Ehrenfeucht, Joost Engelfriet, and Grzegorz Rozenberg. Finite languages for the representation of finite graphs. Journal of Computer and System Sciences, 52(1): 170–184, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900136.

Honkala:1996:PSL

[1328] Juha Honkala. On Parikh slender languages and power series. Journal of Computer and System Sciences, 52 (1):185–190, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900148.

Fortnow:1996:GS

[1329] Lance Fortnow and Tomoyuki Yamakami. Generic separations. Journal of Computer and System Sciences, 52 (1):191–197, February 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009690015X.

Maass:1996:EF

[1330] Wolfgang Maass. Editor’s foreword. Journal of Computer and System Sciences, 52(2):199, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900161.

Schapire:1996:LSM

[1331] Robert E. Schapire and Linda M. Sellie. Learning sparse multivariate polynomials over a field with queries and counterexamples. Journal of Computer and System Sciences, 52 (2):201–213, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900173.

Kummer:1996:SDI

[1332] Martin Kummer and Frank Stephan. On the structure of degrees of inferability. Journal of Computer and System Sciences, 52(2):214– 238, April 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900185.

Simon:1996:GBN

[1333] Hans Ulrich Simon. General bounds on the number of examples needed for learning probabilistic concepts. Journal of Computer and System Sciences, 52(2):239–254, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900197.

Goldman:1996:TSL

[1334] Sally A. Goldman and H. David Mathias. Teaching a smarter learner. Journal of Computer and System Sciences, 52(2):255–267, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900203.

Bshouty:1996:AQM

[1335] Nader H. Bshouty, Sally A. Goldman, Thomas R. Hancock, and Sleiman Matar. Asking questions to minimize errors. Journal of Computer and System Sciences, 52(2): 268–286, April 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900215.

Kshemkalyani:1996:TII

[1336] Ajay D. Kshemkalyani. Temporal interactions of intervals in distributed systems. Journal of Computer and System Sciences, 52(2):

287–298, April 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900227.

Amir:1996:LSF

[1337] Amihood Amir, Gary Benson, and Martin Farach. Let sleeping files lie: Pattern matching in Z-compressed files. Journal of Computer and System Sciences, 52(2):299–307, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000096900239.

Schuler:1996:SAC

[1338] Rainer Schuler and Tomoyuki Yamakami. Structural average case complexity. Journal of Computer and System Sciences, 52(2):308– 327, April 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900240.

Vaishnavi:1996:DBB

[1339] Vijay K. Vaishnavi. On k-dimensional balanced binary trees. Journal of Computer and System Sciences, 52 (2):328–348, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900252.

Gibson:1996:ECT

[1340] Gavin J. Gibson. Exact classification with two-layer neural nets. Journal of Computer and System Sciences, 52(2):349–356, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900264.

Andries:1996:ICD

[1341] Marc Andries and Jan Paredaens.

On instance-completeness for database query languages involving object creation. Journal of Computer and System Sciences, 52(2):357– 373, April 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900276.

Fang:1996:LBP

[1342] Shao C. Fang and Santosh S.

Venkatesh. Learning binary perceptrons perfectly efficiently. Journal of Computer and System Sciences, 52 (2):374–389, April 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900288.

Gasarch:1996:EF

[1343] William Gasarch and Ming Li. Editors’ foreword. Journal of Computer and System Sciences, 52(3): 391, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009690029X.

Jain:1996:ICL

[1344] Sanjay Jain and Arun Sharma.

The intrinsic complexity of language identification. Journal of Computer and System Sciences, 52(3):

393–402, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900306.

Kummer:1996:IPP

[1345] Martin Kummer and Frank Stephan. Inclusion problems in parallel learning and games. Journal of Computer and System Sciences, 52(3): 403–420, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900318.

Bshouty:1996:OQS

[1346] Nader H. Bshouty, Richard Cleve, Ricard Gavaldà, Sampath Kannan, and Christino Tamon. Oracles and queries that are sufficient for exact learning. Journal of Computer and System Sciences, 52(3): 421–433, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009690032X.

Bartlett:1996:FSL

[1347] Peter L. Bartlett, Philip M. Long, and Robert C. Williamson. Fatshattering and the learnability of realvalued functions. Journal of Computer and System Sciences, 52(3): 434–452, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900331.

Dobkin:1996:CMB

[1348] David P. Dobkin, Dimitrios Gunopulos, and Wolfgang Maass. Computing the maximum bichromatic discrepancy, with applications to computer graphics and machine learning. Journal of Computer and System Sciences, 52(3):453–470, June 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900343.

Frazier:1996:LCI

[1349] Michael Frazier, Sally Goldman, Nina Mishra, and Leonard Pitt. Learning from a consistently ignorant teacher. Journal of Computer and System Sciences, 52(3): 471–492, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900355.

Beeri:1996:EF

[1350] Catriel Beeri. Editor’s foreword. Journal of Computer and System Sciences, 52(3):493, June 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900367.

Wong:1996:NFC

[1351] Limsoon Wong. Normal forms and conservative extension properties for query languages over collection types. Journal of Computer and System Sciences, 52(3):495–505, June 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900379.

Morishita:1996:EVG

[1352] Shinichi Morishita. An extension of Van Gelder’s alternating fixpoint to magic programs. Journal of Computer and System Sciences, 52(3): 506–521, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900380.

Hirst:1996:CRR

[1353] Tirza Hirst and David Harel. Completeness results for recursive data bases. Journal of Computer and System Sciences, 52(3):522–536, June 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000096900392.

VandenBussche:1996:RPR

[1354] Jan Van den Bussche, Dirk van Gucht, and Gottfried Vossen. Reflective programming in the relational algebra. Journal of Computer and System Sciences, 52(3): 537–549, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900409.

Haas:1996:SCE

[1355] Peter J. Haas, Jeffrey F. Naughton, S. Seshadri, and Arun N. Swami. Selectivity and cost estimation for joins based on random sampling. Journal of Computer and System Sciences, 52(3):550–569, June 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900410.

Grumbach:1996:TTA

[1356] Stéphane Grumbach and Tova Milo.

Towards tractable algebras for bags. Journal of Computer and System Sciences, 52(3):570–588, June 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000096900422.

Kanellakis:1996:IDM

[1357] Paris Kanellakis, Sridhar Ramaswamy, Darren E. Vengroff, and Jeffrey Scott Vitter. Indexing for data models with constraints and classes. Journal of Computer and System Sciences, 52(3):589–612, June 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900434.

Anonymous:1996:AIVa

[1358] Anonymous. Author index for volume 52. Journal of Computer and System Sciences, 52(3): 613, June 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900446.

Vikas:1996:AAG

[1359] Narayan Vikas. An O(n) algorithm for Abelian p-group isomorphism and an O(n log n) algorithm for Abelian group isomorphism. Journal of Computer and System Sciences, 53(1):1–9, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900458.

Shallit:1996:APM

[1360] Jeffrey Shallit and Yuri Breitbart. Automaticity i: Properties of a measure of descriptional complexity. Journal of Computer and System Sciences, 53(1):10–25, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009690046X.

Baliga:1996:LAC

[1361] Ganesh Baliga and John Case. Learnability: Admissible, co-finite, and hypersimple languages. Journal of Computer and System Sciences, 53 (1):26–32, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900471.

Drewes:1996:LTA

[1362] Frank Drewes. Language theoretic and algorithmic properties of d-dimensional collages and patterns in a grid. Journal of Computer and System Sciences, 53(1):33–60, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900483.

Moran:1996:CC

[1363] Shlomo Moran, Gadi Taubenfeld, and Irit Yadin. Concurrent counting. Journal of Computer and System Sciences, 53(1):61–78, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900495.

Sieling:1996:NLB

[1364] Detlef Sieling. New lower bounds and hierarchy results for restricted branching programs. Journal of Computer and System Sciences, 53 (1):79–87, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900501.

Lange:1996:ILP

[1365] Steffen Lange and Thomas Zeugmann. Incremental learning from positive data. Journal of Computer and System Sciences, 53(1): 88–103, August 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900513.

Fich:1996:LPP

[1366] Faith E. Fich, Russell Impagliazzo, Bruce Kapron, Valerie King, and Miroslaw Kutylowski. Limits on the power of parallel random access machines with weak forms of write conflict resolution. Journal of Computer and System Sciences, 53(1): 104–111, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900525.

Gao:1996:TSP

[1367] Feng Gao. Towards structured parallel computing on architectureindependent parallel algorithm design for distributed-memory architectures. Journal of Computer and System Sciences, 53(1):112– 128, August 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900537.

Gabow:1996:ETP

[1368] Harold N. Gabow and Ying Xu. Efficient theoretic and practical algorithms for linear matroid intersection problems. Journal of Computer and System Sciences, 53(1): 129–147, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900549.

Dietz:1996:SEV

[1369] Paul F. Dietz. A space efficient variant of path copying for partially persistent sorted sets. Journal of Computer and System Sciences, 53 (1):148–152, August 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900550.

Homer:1996:EF

[1370] Steven Homer. Editor’s foreword. Journal of Computer and System Sciences, 53(2):153, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900562.

Agrawal:1996:PTI

[1371] Manindra Agrawal and Somenath Biswas. Polynomial-time isomorphism of 1 − L-complete sets. Journal of Computer and System Sciences, 53 (2):155–160, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900574.

Papadimitriou:1996:LNC

[1372] Christos H. Papadimitriou and Mihalis Yannakakis. On limited nondeterminism and the complexity of the V–C dimension. Journal of Computer and System Sciences, 53(2): 161–170, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900586.

Ogihara:1996:CPC

[1373] Mitsunori Ogihara, Thomas Thierauf, Seinosuke Toda, and Osamu Watanabe. On closure properties of #P in the context of P F · #P. Journal of Computer and System Sciences, 53 (2):171–179, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900598.

Hirst:1996:TIL

[1374] Tirza Hirst and David Harel. Taking it to the limit: On infinite variants of NP-complete problems. Journal of Computer and System Sciences, 53 (2):180–193, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900604.

Hemaspaandra:1996:PSS

[1375] Edith Hemaspaandra, Ashish V. Naik, Mitsunori Ogihara, and Alan L. Selman. P-selective sets and reducing search to decision vs selfreducibility. Journal of Computer and System Sciences, 53(2):194– 209, October 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900616.

Buhrman:1996:PSS

[1376] Harry Buhrman and Leen Torenvliet. P-selective self-reducible sets: a new characterization of P. Journal of Computer and System Sciences, 53 (2):210–217, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900628.

Dymond:1996:PVA

[1377] Patrick W. Dymond, Faith E. Fich, Naomi Nishimura, Prabhakar Ragde, and Walter L. Ruzzo. Pointers versus arithmetic in PRAMs. Journal of Computer and System Sciences, 53 (2):218–232, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009690063X.

Schoning:1996:EFS

[1378] Uwe Schöning. Editor’s foreword:

Special issue on the 9th Annual Structure in Complexity Theory Conference in Amsterdam, June 28– July 1, 1994. Journal of Computer and System Sciences, 53(2): 233, October 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900641.

Kautz:1996:RRO

[1379] Steven M. Kautz and Peter Bro Miltersen. Relative to a random oracle, NP is not small. Journal of Computer and System Sciences, 53 (2):235–250, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900653.

Tardos:1996:MPE

[1380] Gábor Tardos. Multi-prover encoding schemes and three-prover proof systems. Journal of Computer and System Sciences, 53(2): 251–260, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900665.

Buhrman:1996:RSM

[1381] Harry Buhrman and Pekka Orponen. Random strings make hard instances. Journal of Computer and System Sciences, 53(2):261– 266, October 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900677.

Agrawal:1996:ICW

[1382] Manindra Agrawal. On the isomorphism conjecture for weak reducibilities. Journal of Computer and System Sciences, 53(2):267– 282, October 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900689.

Compton:1996:LDC

[1383] Kevin J. Compton and Erich Grädel. Logical definability of counting functions. Journal of Computer and System Sciences, 53(2):283–297, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000096900690.

Chang:1996:QCC

[1384] Richard Chang. On the query complexity of clique size and maximum satisfiability. Journal of Computer and System Sciences, 53(2): 298–313, October 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900707.

Anonymous:1996:EF

[1385] Anonymous. Editors’ foreword. Journal of Computer and System Sciences, 53(3):315, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900719.

Mitzenmacher:1996:BGR

[1386] Michael Mitzenmacher. Bounds on the greedy routing algorithm for array networks. Journal of Computer and System Sciences, 53(3): 317–327, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900720.

Schwabe:1996:IPD

[1387] Eric J. Schwabe and Ian M. Sutherland. Improved parity-declustered layouts for disk arrays. Journal of Computer and System Sciences, 53(3): 328–343, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900732.

Reid-Miller:1996:LRL

[1388] Margaret Reid-Miller. List ranking and list scan on the Cray C90. Journal of Computer and System Sciences, 53 (3):344–356, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900744.

Ghosh:1996:DLB

[1389] Bhaskar Ghosh and S. Muthukrishnan. Dynamic load balancing by random matchings. Journal of Computer and System Sciences, 53(3): 357–370, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900756.

Tamaki:1996:CMT

[1390] Hisao Tamaki. Construction of the mesh and the torus tolerating a large number of faults. Journal of Computer and System Sciences, 53(3): 371–379, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900768.

Leoncini:1996:PCG

[1391] M. Leoncini. On the parallel complexity of Gaussian elimination with pivoting. Journal of Computer and System Sciences, 53(3): 380–394, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009690077X.

Halperin:1996:ORL

[1392] Shay Halperin and Uri Zwick. An optimal randomised logarithmic time connectivity algorithm for the EREW PRAM. Journal of Computer and System Sciences, 53(3):395– 416, December 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900781.

Gibbons:1996:ELC

[1393] Phillip B. Gibbons, Yossi Matias, and Vijaya Ramachandran. Efficient lowcontention parallel algorithms. Journal of Computer and System Sciences, 53 (3):417–442, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900793.

Meghini:1996:OPL

[1394] Carlo Meghini and Costantino Thanos.

An optimal predicate locking scheduler. Journal of Computer and System Sciences, 53(3):443–468, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200009690080X.

Goerdt:1996:TU

[1395] Andreas Goerdt. A threshold for unsatisfiability. Journal of Computer and System Sciences, 53(3): 469–486, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900811.

Kari:1996:MMS

[1396] Lila Kari and Gabriel Thierrin. Maximal and minimal solutions to language equations. Journal of Computer and System Sciences, 53(3): 487–496, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900823.

Eiter:1996:CNC

[1397] Thomas Eiter and Georg Gottlob.

The complexity of nested counterfactuals and iterated knowledge base revisions. Journal of Computer and System Sciences, 53(3):

497–512, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900835.

Jiang:1996:OWH

[1398] Tao Jiang and Ming Li. k one-way heads cannot do stringmatching. Journal of Computer and System Sciences, 53(3):513– 524, December 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900847.

Blair:1996:RRG

[1399] Jean R. S. Blair and Errol L. Lloyd. River routing with a generalized model. Journal of Computer and System Sciences, 53(3): 525–544, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900859.

Chari:1996:CUR

[1400] Suresh Chari and Pankaj Rohatgi. On completeness under random reductions. Journal of Computer and System Sciences, 53(3):545– 555, December 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000096900860.

Engelfriet:1996:RDC

[1401] Joost Engelfriet and Vincent van Oostrom. Regular description of contextfree graph languages. Journal of Computer and System Sciences, 53(3): 556–574, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900872.

Grzymala-Busse:1996:PTT

[1402] Jerzy W. Grzymala-Busse and Soe Than. Partition triples: a tool for reduction of data sets. Journal of Computer and System Sciences, 53(3): 575–582, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900884.

Jain:1996:PSP

[1403] Sanjay Jain. Program synthesis in the presence of infinite number of inaccuracies. Journal of Computer and System Sciences, 53(3): 583–591, December 1996. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900896.

Anonymous:1996:AIVb

[1404] Anonymous. Author index for volume 53. Journal of Computer and System Sciences, 53(3): 592, December 1996. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000096900902.

Fagin:1997:EF

[1405] Ronald Fagin. Editor’s foreword. Journal of Computer and System Sciences, 54(1):1, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914483.

Beeri:1997:CFP

[1406] Catriel Beeri and Tova Milo. Comparison of functional and predicative query paradigms. Journal of Computer and System Sciences, 54 (1):3–33, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914495.

VandenBussche:1997:SAO

[1407] Jan Van den Bussche and Dirk Van Gucht. A semideterministic approach to object creation and nondeterminism in database queries. Journal of Computer and System Sciences, 54 (1):34–47, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914501.

Papadimitriou:1997:TBS

[1408] Christos H. Papadimitriou and Mihalis Yannakakis. Tie-breaking semantics and structural totality. Journal of Computer and System Sciences, 54 (1):48–60, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914513.

Chaudhuri:1997:ERN

[1409] Surajit Chaudhuri and Moshe Y. Vardi. On the equivalence of recursive and nonrecursive Datalog programs. Journal of Computer and System Sciences, 54(1):61–78, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914525.

Ross:1997:MAD

[1410] Kenneth A. Ross and Yehoshua Sagiv. Monotonic aggregation in deductive databases. Journal of Computer and System Sciences, 54(1):79– 97, February 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914537.

Grahne:1997:KT

[1411] Gösta Grahne, Alberto O. Mendelzon, and Peter Z. Revesz. Knowledgebase transformations. Journal of Computer and System Sciences, 54 (1):98–112, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914549.

vanderMeyden:1997:CQI

[1412] Ron van der Meyden. The complexity of querying indefinite data about linearly ordered domains. Journal of Computer and System Sciences, 54 (1):113–135, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914550.

Bonner:1997:RMR

[1413] Anthony J. Bonner and Tomasz Imielinski. Reusing and modifying rulebases by predicate substitution. Journal of Computer and System Sciences, 54(1):136– 166, February 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914562.

Maass:1997:EF

[1414] Wolfgang Maass. Editor’s foreword. Journal of Computer and System Sciences, 54(1):167, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000097914768.

Karpinski:1997:PBV

[1415] Marek Karpinski and Angus Macintyre. Polynomial bounds for VC dimension of sigmoidal and general Pfaffian neural networks. Journal of Computer and System Sciences, 54 (1):169–176, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009791477X.

Koiran:1997:WVB

[1416] Pascal Koiran. A weak version of the Blum, Shub, and Smale model. Journal of Computer and System Sciences, 54 (1):177–189, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914781.

Koiran:1997:NNQ

[1417] Pascal Koiran and Eduardo D. Sontag. Neural networks with quadratic VC dimension. Journal of Computer and System Sciences, 54(1): 190–198, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914793.

Turan:1997:CBF

[1418] György Turán and Farrokh Vatan.

On the computation of Boolean functions by analog circuits of bounded fan-in. Journal of Computer and System Sciences, 54(1): 199–212, February 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009791480X.

Sipser:1997:EF

[1419] Michael Sipser. Editor’s foreword.

Journal of Computer and System Sciences, 54(2):213, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914811.

Lapidot:1997:FPM

[1420] Dror Lapidot and Adi Shamir. Fully parallelized multi-prover protocols for NEXP-time. Journal of Computer and System Sciences, 54(2): 215–220, April 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097912381.

Radhakrishnan:1997:BLB

[1421] Jaikumar Radhakrishnan. Better lower bounds for monotone threshold formulas. Journal of Computer and System Sciences, 54(2): 221–226, April 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097912873.

Johnson:1997:CCP

[1422] Donald B. Johnson and Panagiotis Metaxas. Connected components in O(log^{3/2} n) parallel time for the CREW PRAM. Journal of Computer and System Sciences, 54(2): 227–242, April 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097912915.

Galil:1997:APS

[1423] Zvi Galil and Oded Margalit. All pairs shortest paths for graphs with small integer length edges. Journal of Computer and System Sciences, 54 (2):243–254, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097913854.

Alon:1997:EAP

[1424] Noga Alon, Zvi Galil, and Oded Margalit. On the exponent of the all pairs shortest path problem. Journal of Computer and System Sciences, 54(2):255–262, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009791388X.

Epstein:1997:EF

[1425] David Epstein. Editor’s foreword.

Journal of Computer and System Sciences, 54(2):263, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914690.

Leighton:1997:BBS

[1426] Tom Leighton, Yuan Ma, and C. Greg Plaxton. Breaking the Θ(n log^{2} n) barrier for sorting with faults. Journal of Computer and System Sciences, 54(2):265–304, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914707.

Feige:1997:STS

[1427] Uriel Feige. A spectrum of timespace trade-offs for undirected s– t connectivity. Journal of Computer and System Sciences, 54(2): 305–316, April 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914719.

Arora:1997:HAO

[1428] Sanjeev Arora, László Babai, Jacques Stern, and Z. Sweedyk. The hardness of approximate optima in lattices, codes, and systems of linear equations. Journal of Computer and System Sciences, 54(2):317–331, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914720.

Leiserson:1997:ECA

[1429] Charles E. Leiserson, Satish Rao, and Sivan Toledo. Efficient out-of-core algorithms for linear relaxation using blocking covers. Journal of Computer and System Sciences, 54 (2):332–344, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914732.

Ben-Amram:1997:WCW

[1430] Amir M. Ben-Amram and Zvi Galil. When can we sort in o(n log n) time? Journal of Computer and System Sciences, 54(2):345– 370, April 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914744.

Blum:1997:LIC

[1431] Avrim L. Blum and Ravindran Kannan. Learning an intersection of a constant number of halfspaces over a uniform distribution. Journal of Computer and System Sciences, 54 (2):371–380, April 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914756.

Allender:1997:EF

[1432] Eric Allender. Editor’s foreword. Journal of Computer and System Sciences, 54(3):381, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914823.

Agrawal:1997:DND

[1433] Manindra Agrawal. DSPACE(n) = ? NSPACE(n): a degree theoretic characterization. Journal of Computer and System Sciences, 54(3): 383–392, June 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914835.

Buhrman:1997:EKR

[1434] Harry Buhrman and Elvira Mayordomo. An excursion to the Kolmogorov random strings. Journal of Computer and System Sciences, 54 (3):393–399, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914847.

Etessami:1997:CQS

[1435] Kousha Etessami. Counting quantifiers, successor relations, and logarithmic space. Journal of Computer and System Sciences, 54(3): 400–411, June 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914859.

Rogers:1997:ICH

[1436] John Rogers. The isomorphism conjecture holds and one-way functions exist relative to an oracle. Journal of Computer and System Sciences, 54(3):412–423, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914860.

Royer:1997:SVS

[1437] James S. Royer. Semantics vs syntax vs computations: Machine models for type-2 polynomial-time bounded functionals. Journal of Computer and System Sciences, 54 (3):424–436, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914872.

Tran:1997:PIE

[1438] Nicholas Tran. On P-immunity of exponential time complete sets. Journal of Computer and System Sciences, 54(3):437–440, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914884.

Sacca:1997:EPS

[1439] Domenico Saccà. The expressive powers of stable models for bound and unbound DATALOG queries. Journal of Computer and System Sciences, 54(3):441–464, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009791446X.

Cai:1997:FPT

[1440] Liming Cai and Jianer Chen. On fixed-parameter tractability and approximability of NP optimization problems. Journal of Computer and System Sciences, 54(3):465– 474, June 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914902.

Ben-Asher:1997:GAO

[1441] Yosi Ben-Asher and Ilan Newman. Geometric approach for optimal routing on a mesh with buses. Journal of Computer and System Sciences, 54(3):475–486, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914926.

Mathias:1997:MIT

[1442] H. David Mathias. A model of interactive teaching. Journal of Computer and System Sciences, 54(3): 487–501, June 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914914.

Meinke:1997:CTE

[1443] Karl Meinke. A completeness theorem for the expressive power of higherorder algebraic specifications. Journal of Computer and System Sciences, 54(3):502–519, June 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914896.

Anonymous:1997:AIVa

[1444] Anonymous. Author index for volume 54. Journal of Computer and System Sciences, 54(3): 520, June 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915099.

Beigel:1997:GEF

[1445] Richard Beigel and Roberto Tamassia. Guest Editors’ foreword. Journal of Computer and System Sciences, 55(1):1, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915294.

Henzinger:1997:FSP

[1446] Monika R. Henzinger, Philip Klein, Satish Rao, and Sairam Subramanian. Faster shortest-path algorithms for planar graphs. Journal of Computer and System Sciences, 55(1):3–23, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097914938.

Razborov:1997:NP

[1447] Alexander A. Razborov and Steven Rudich. Natural proofs. Journal of Computer and System Sciences, 55 (1):24–35, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009791494X.

Yao:1997:DTC

[1448] Andrew Chi-Chih Yao. Decision tree complexity and Betti numbers. Journal of Computer and System Sciences, 55(1):36–43, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914951.

Hariharan:1997:OPS

[1449] Ramesh Hariharan. Optimal parallel suffix tree construction. Journal of Computer and System Sciences, 55(1):44–69, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914963.

Li:1997:EF

[1450] Ming Li and Paul Vitányi. Editors’ foreword. Journal of Computer and System Sciences, 55(1): 71, August 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914999.

Solomonoff:1997:DAP

[1451] Ray J. Solomonoff. The discovery of algorithmic probability. Journal of Computer and System Sciences, 55(1):73–88, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915002.

Rissanen:1997:SCL

[1452] J. Rissanen. Stochastic complexity in learning. Journal of Computer and System Sciences, 55(1): 89–95, August 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915014.

Vovk:1997:LAP

[1453] V. G. Vovk. Learning about the parameter of the Bernoulli model. Journal of Computer and System Sciences, 55(1):96–104, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915026.

Yamanishi:1997:LML

[1454] Kenji Yamanishi. On-line maximum likelihood prediction with respect to general loss functions. Journal of Computer and System Sciences, 55 (1):105–118, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915038.

Freund:1997:DTG

[1455] Yoav Freund and Robert E. Schapire. A decision-theoretic generalization of on-line learning and an application to boosting. Journal of Computer and System Sciences, 55(1): 119–139, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009791504X.

Schuurmans:1997:CRV

[1456] Dale Schuurmans. Characterizing rational versus exponential learning curves. Journal of Computer and System Sciences, 55(1):140– 160, August 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915051.

Gurvits:1997:ALC

[1457] Leonid Gurvits and Pascal Koiran.

Approximation and learning of convex superpositions. Journal of Computer and System Sciences, 55(1): 161–170, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915063.

Ben-David:1997:LDT

[1458] Shai Ben-David and Michael Lindenbaum. Learning distributions by their density levels: A paradigm for learning without a teacher. Journal of Computer and System Sciences, 55 (1):171–182, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915075.

Case:1997:LRF

[1459] John Case, Susanne Kaufmann, Efim Kinber, and Martin Kummer. Learning recursive functions from approximations. Journal of Computer and System Sciences, 55(1): 183–196, August 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915087.

Patnaik:1997:DFP

[1460] Sushant Patnaik and Neil Immerman. Dyn-FO: a parallel, dynamic complexity class. Journal of Computer and System Sciences, 55(2): 199–209, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915208.

Afrati:1997:BAF

[1461] Foto N. Afrati. Bounded arity fDatalog(/=) queries on graphs. Journal of Computer and System Sciences, 55 (2):210–228, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009791521X.

Faloutsos:1997:RUI

[1462] Christos Faloutsos and Ibrahim Kamel. Relaxing the uniformity and independence assumptions using the concept of fractal dimension. Journal of Computer and System Sciences, 55 (2):229–240, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915221.

Libkin:1997:QLB

[1463] Leonid Libkin and Limsoon Wong. Query languages for bags and aggregate functions. Journal of Computer and System Sciences, 55(2): 241–272, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915233.

Grumbach:1997:FRD

[1464] Stéphane Grumbach and Jianwen Su.

Finitely representable databases. Journal of Computer and System Sciences, 55(2):273–298, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915245.

Suciu:1997:QLN

[1465] Dan Suciu and Val Tannen. A query language for NC. Journal of Computer and System Sciences, 55(2): 299–321, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915257.

Suciu:1997:CEC

[1466] Dan Suciu and Jan Paredaens. The complexity of the evaluation of complex algebra expressions. Journal of Computer and System Sciences, 55 (2):322–343, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915269.

Krishnaswamy:1997:RSA

[1467] V. Krishnaswamy, D. Agrawal, J. L.

Bruno, and A. El Abbadi. Relative serializability: an approach for relaxing the atomicity of transactions. Journal of Computer and System Sciences, 55 (2):344–354, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915270.

Chaudhuri:1997:CDA

[1468] Surajit Chaudhuri and Phokion G. Kolaitis. Can Datalog be approximated? Journal of Computer and System Sciences, 55(2): 355–369, October 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915282.

Goldwasser:1997:EF

[1469] Shafi Goldwasser and S. Rao Kosaraju. Editors’ foreword. Journal of Computer and System Sciences, 55(3): 371, December 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915397.

Sinha:1997:EOB

[1470] Rakesh Kumar Sinha and Jayram S. Thathachar. Efficient oblivious branching programs for threshold and mod functions. Journal of Computer and System Sciences, 55(3): 373–384, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915300.

Kleitman:1997:DRB

[1471] Dan Kleitman, Tom Leighton, and Yuan Ma. On the design of reliable Boolean circuits that contain partially unreliable gates. Journal of Computer and System Sciences, 55(3): 385–401, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915312.

Karger:1997:RCS

[1472] David R. Karger and Daphne Koller.

(de)randomized construction of small sample spaces in NC. Journal of Computer and System Sciences, 55(3): 402–413, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915324.

Jackson:1997:EMQ

[1473] Jeffrey C. Jackson. An efficient membership-query algorithm for learning DNF with respect to the uniform distribution. Journal of Computer and System Sciences, 55(3): 414–440, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915336.

Cohen:1997:SEF

[1474] Edith Cohen. Size-estimation framework with applications to transitive closure and reachability. Journal of Computer and System Sciences, 55(3): 441–453, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915348.

Weihe:1997:MFP

[1475] Karsten Weihe. Maximum (s, t)-flows in planar networks in O(|V | log |V |) time. Journal of Computer and System Sciences, 55(3):454–475, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000097915385.

Berstel:1997:SMW

[1476] Jean Berstel and L. Boasson. The set of minimal words of a contextfree language is context-free. Journal of Computer and System Sciences, 55 (3):477–488, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914975.

Engelfriet:1997:LDC

[1477] Joost Engelfriet and Vincent van Oostrom. Logical description of contextfree graph languages. Journal of Computer and System Sciences, 55(3): 489–503, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915105.

Boyar:1997:ARC

[1478] Joan Boyar, Rolf Fagerberg, and Kim S. Larsen. Amortization results for chromatic search trees, with an application to priority queues. Journal of Computer and System Sciences, 55 (3):504–521, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915117.

Fromentin:1997:SGS

[1479] Eddy Fromentin and Michel Raynal. Shared global states in distributed computations. Journal of Computer and System Sciences, 55(3): 522–528, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915129.

Yan:1997:STP

[1480] G. Y. Yan, A. Albrecht, G. H. F.

Young, and C. K. Wong. The Steiner tree problem in orientation metrics. Journal of Computer and System Sciences, 55(3): 529–546, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915130.

Hemaspaandra:1997:USC

[1481] Lane A. Hemaspaandra and Mitsunori Ogihara. Universally serializable computation. Journal of Computer and System Sciences, 55(3): 547–560, December 1997. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915166.

Anonymous:1997:AIVb

[1482] Anonymous. Author index for volume 55. Journal of Computer and System Sciences, 55(3): 561, December 1997. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915403.

Ginsburg:1998:RSO

[1483] Seymour Ginsburg and X. Sean Wang.

Regular sequence operations and their use in database queries. Journal of Computer and System Sciences, 56 (1):1–26, February 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915142.

Satta:1998:TIS

[1484] Giorgio Satta. Trading independent for synchronized parallelism in finite copying parallel rewriting systems. Journal of Computer and System Sciences, 56(1):27–45, February 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915154.

Weber:1998:TSV

[1485] Andreas Weber. Transforming a single-valued transducer into a mealy machine. Journal of Computer and System Sciences, 56(1):46– 59, February 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915178.

Manzini:1998:ILC

[1486] Giovanni Manzini and Luciano Margara. Invertible linear cellular automata over Z_{m}:: Algorithmic and dynamical aspects. Journal of Computer and System Sciences, 56(1):60– 67, February 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009791535X.

Rastogi:1998:CNE

[1487] Rajeev Rastogi, Sharad Mehrotra, Yuri Breitbart, Henry F. Korth, and Avi Silberschatz. On correctness of nonserializable executions. Journal of Computer and System Sciences, 56 (1):68–82, February 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915361.

Rajasekaran:1998:TRT

[1488] Sanguthevar Rajasekaran and Shibu Yooseph. TAL recognition in O(M(n^{2})) time. Journal of Computer and System Sciences, 56(1):83– 89, February 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915373.

Duris:1998:LBM

[1489] Pavol Duris and José D. P. Rolim.

Lower bounds on the multiparty communication complexity. Journal of Computer and System Sciences, 56 (1):90–95, February 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915476.

Ladner:1998:CPR

[1490] Richard E. Ladner, Anthony LaMarca, and Ewan Tempero. Counting protocols for reliable end-to-end transmission. Journal of Computer and System Sciences, 56(1): 96–111, February 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009791549X.

Bshouty:1998:IAR

[1491] Daoud Bshouty and Nader H. Bshouty. On interpolating arithmetic read-once formulas with exponentiation. Journal of Computer and System Sciences, 56 (1):112–124, February 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915506.

Kortelainen:1998:RAC

[1492] Juha Kortelainen. Remarks about commutative context-free languages. Journal of Computer and System Sciences, 56(1):125–129, February 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000097915531.

Shapire:1998:GEF

[1493] Robert E. Shapire. Guest Editor’s foreword. Journal of Computer and System Sciences, 56(2): 131, April 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915543.

Ron:1998:LUA

[1494] Dana Ron, Yoram Singer, and Naftali Tishby. On the learnability and usage of acyclic probabilistic finite automata. Journal of Computer and System Sciences, 56(2): 133–152, April 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915555.

Vovk:1998:GPE

[1495] V. Vovk. A game of prediction with expert advice. Journal of Computer and System Sciences, 56(2): 153–173, April 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915567.

Bartlett:1998:PLU

[1496] Peter L. Bartlett and Philip M.

Long. Prediction, learning, uniform convergence, and scale-sensitive dimensions. Journal of Computer and System Sciences, 56(2):174– 190, April 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915579.

Aslam:1998:SSS

[1497] Javed A. Aslam and Scott E. Decatur. Specification and simulation of statistical query algorithms for efficiency and noise tolerance. Journal of Computer and System Sciences, 56(2):191–208, April 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915580.

Blum:1998:LUB

[1498] Avrim Blum, Prasad Chalasani, Sally A. Goldman, and Donna K. Slonim. Learning with unreliable boundary queries. Journal of Computer and System Sciences, 56(2): 209–222, April 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915592.

Kim:1998:CCR

[1499] Jeong Han Kim and James R. Roche. Covering cubes by random half cubes, with applications to binary neural networks. Journal of Computer and System Sciences, 56 (2):223–252, April 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915609.

Ilie:1998:TRP

[1500] Lucian Ilie and Arto Salomaa. 2testability and relabelings produce everything. Journal of Computer and System Sciences, 56(3):253– 262, June 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915488.

Berg:1998:LBP

[1501] Christer Berg and Staffan Ulfberg. A lower bound for perceptrons and an oracle separation of the PP^{P H} hierarchy. Journal of Computer and System Sciences, 56 (3):263–271, June 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009791552X.

Sharma:1998:NBI

[1502] Arun Sharma. A note on batch and incremental learnability. Journal of Computer and System Sciences, 56 (3):272–276, June 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915689.

Ben-David:1998:LRF

[1503] Shai Ben-David and Eli Dichterman. Learning with restricted focus of attention. Journal of Computer and System Sciences, 56(3): 277–298, June 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915690.

Malvestuto:1998:FAQ

[1504] Francesco Mario Malvestuto and Marina Moscarini. A fast algorithm for query optimization in universalrelation databases. Journal of Computer and System Sciences, 56(3): 299–309, June 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915707.

Bshouty:1998:AEL

[1505] Nader Bshouty and Lisa Hellerstein.

Attribute-efficient learning in query and mistake-bound models. Journal of Computer and System Sciences, 56(3):310–319, June 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915719.

Chen:1998:OCA

[1506] Lin Chen. Optimal circular arc representations: Properties, recognition, and construction. Journal of Computer and System Sciences, 56 (3):320–331, June 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915720.

Engelfriet:1998:EBT

[1507] Joost Engelfriet and Heiko Vogler. The equivalence of bottom-up and topdown tree-to-graph transducers. Journal of Computer and System Sciences, 56(3):332–356, June 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915732.

Anonymous:1998:AIVa

[1508] Anonymous. Author index for volume 56. Journal of Computer and System Sciences, 56(3): 357, June 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009891594X.

Gabow:1998:GEF

[1509] Harold N. Gabow. Guest Editor’s foreword. Journal of Computer and System Sciences, 57(1): 1, August 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915744.

Beame:1998:RCN

[1510] Paul Beame, Stephen Cook, Jeff Edmonds, Russell Impagliazzo, and Toniann Pitassi. The relative complexity of NP search problems. Journal of Computer and System Sciences, 57(1):3–19, August 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915756.

Diaconis:1998:WDW

[1511] P. Diaconis and L. Saloff-Coste. What do we know about the Metropolis Algorithm? Journal of Computer and System Sciences, 57(1): 20–36, August 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915768.

Miltersen:1998:DSA

[1512] Peter Bro Miltersen, Noam Nisan, Shmuel Safra, and Avi Wigderson. On data structures and asymmetric communication complexity. Journal of Computer and System Sciences, 57(1):37–49, August 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009891577X.

Kavraki:1998:RQP

[1513] Lydia E. Kavraki, Jean-Claude Latombe, Rajeev Motwani, and Prabhakar Raghavan. Randomized query processing in robot path planning. Journal of Computer and System Sciences, 57(1):50–60, August 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915781.

Kleinberg:1998:ADP

[1514] Jon Kleinberg and Éva Tardos. Approximations for the disjoint paths problem in high-diameter planar networks. Journal of Computer and System Sciences, 57(1):61–73, August 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000098915793.

Andersson:1998:SLT

[1515] Arne Andersson, Torben Hagerup, Stefan Nilsson, and Rajeev Raman. Sorting in linear time? Journal of Computer and System Sciences, 57 (1):74–93, August 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009891580X.

Henzinger:1998:WDA

[1516] Thomas A. Henzinger, Peter W. Kopke, Anuj Puri, and Pravin Varaiya. What’s decidable about hybrid automata? Journal of Computer and System Sciences, 57(1): 94–124, August 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915811.

Cai:1998:EF

[1517] Jin-Yi Cai and Alan Selman. Editors’ foreword. Journal of Computer and System Sciences, 57(2):

125, October 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915823.

Agrawal:1998:RCC

[1518] Manindra Agrawal, Eric Allender, and Steven Rudich. Reductions in circuit complexity: an isomorphism theorem and a gap theorem. Journal of Computer and System Sciences, 57 (2):127–143, October 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915835.

Barland:1998:IPF

[1519] Ian Barland, Phokion G. Kolaitis, and Madhukar N. Thakur. Integer programming as a framework for optimization and approximability. Journal of Computer and System Sciences, 57 (2):144–161, October 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915847.

Buss:1998:GDB

[1520] Samuel R. Buss and Toniann Pitassi.

Good degree bounds on Nullstellensatz refutations of the induction principle. Journal of Computer and System Sciences, 57(2): 162–171, October 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915859.

Bach:1998:DMA

[1521] Eric Bach, Anne Condon, Elton Glaser, and Celena Tanguay. DNA models and algorithms for NP-complete problems. Journal of Computer and System Sciences, 57(2):172– 186, October 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915860.

Feige:1998:ZKC

[1522] Uriel Feige and Joe Kilian. Zero knowledge and the chromatic number. Journal of Computer and System Sciences, 57(2):187–199, October 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915872.

Caussinus:1998:NNC

[1523] Hervé Caussinus, Pierre McKenzie, Denis Thérien, and Heribert Vollmer. Nondeterministic NC^{1} computation. Journal of Computer and System Sciences, 57(2): 200–212, October 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915884.

vanMelkebeek:1998:DRB

[1524] Dieter van Melkebeek. Deterministic and randomized bounded truthtable reductions of P, NL, and L to sparse sets. Journal of Computer and System Sciences, 57(2): 213–232, October 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915896.

Abiteboul:1998:GEF

[1525] Serge Abiteboul. Guest Editor’s foreword. Journal of Computer and System Sciences, 57(3):233, December 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915616.

Bonner:1998:SDT

[1526] Anthony Bonner and Giansalvatore Mecca. Sequences, Datalog, and transducers. Journal of Computer and System Sciences, 57(3): 234–259, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915628.

Chaudhri:1998:SLP

[1527] Vinay K. Chaudhri and Vassos Hadzilacos. Safe locking policies for dynamic databases. Journal of Computer and System Sciences, 57(3): 260–271, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009891563X.

Consens:1998:AQT

[1528] Mariano P. Consens and Tova Milo.

Algebras for querying text regions: Expressive power and optimization. Journal of Computer and System Sciences, 57(3):272–288, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000098915641.

Dong:1998:ABF

[1529] Guozhu Dong and Jianwen Su. Arity bounds in first-order incremental evaluation and definition of polynomial time database queries. Journal of Computer and System Sciences, 57 (3):289–308, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915653.

Keidar:1998:IRD

[1530] Idit Keidar and Danny Dolev. Increasing the resilience of distributed and replicated database systems. Journal of Computer and System Sciences, 57 (3):309–324, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915665.

Picouet:1998:SEI

[1531] Philippe Picouet and Victor Vianu. Semantics and expressiveness issues in active databases. Journal of Computer and System Sciences, 57(3): 325–355, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915677.

Jain:1998:LR

[1532] Sanjay Jain. Learning with refutation. Journal of Computer and System Sciences, 57(3):356–365, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000098915914.

Hagerup:1998:CMF

[1533] Torben Hagerup, Jyrki Katajainen, Naomi Nishimura, and Prabhakar Ragde. Characterizing multiterminal flow networks and computing flows in networks of small treewidth. Journal of Computer and System Sciences, 57 (3):366–375, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915926.

Auer:1998:AHR

[1534] Peter Auer, Philip M. Long, and Aravind Srinivasan. Approximating hyper-rectangles: Learning and pseudorandom sets. Journal of Computer and System Sciences, 57(3): 376–388, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915938.

Asarin:1998:ATC

[1535] Eugene Asarin and Oded Maler.

Achilles and the tortoise climbing up the arithmetical hierarchy. Journal of Computer and System Sciences, 57 (3):389–398, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098916014.

Schulman:1998:TPC

[1536] Leonard J. Schulman. A three-party communication problem. Journal of Computer and System Sciences, 57(3): 399–401, December 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916038.

Anonymous:1998:AIVb

[1537] Anonymous. Author index for volume 57. Journal of Computer and System Sciences, 57(3): 402, December 1998. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916075.

Hull:1999:EF

[1538] Richard Hull. Editor’s foreword. Journal of Computer and System Sciences, 58(1):iv, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916203.

Fukuda:1999:MOA

[1539] Takeshi Fukuda, Yasuhiko Morimoto, Shinichi Morishita, and Takeshi Tokuyama. Mining optimized association rules for numeric attributes. Journal of Computer and System Sciences, 58(1):1–12, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915951.

Jagadish:1999:MCB

[1540] H. V. Jagadish, Alberto O. Mendelzon, and Inderpal Singh Mumick. Managing conflicts between rules. Journal of Computer and System Sciences, 58(1):13–28, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915963.

Papadimitriou:1999:TQS

[1541] CH. Papadimitriou, D. Suciu, and V. Vianu. Topological queries in spatial databases. Journal of Computer and System Sciences, 58(1):29– 53, February 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915975.

Abiteboul:1999:TCV

[1542] Serge Abiteboul, Laurent Herr, and Jan Van den Bussche. Temporal connectives versus explicit timestamps to query temporal databases. Journal of Computer and System Sciences, 58 (1):54–68, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098915987.

Levy:1999:AQU

[1543] Alon Y. Levy, Anand Rajaraman, and Jeffrey D. Ullman. Answering queries using limited external query processors. Journal of Computer and System Sciences, 58(1):69– 82, February 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915999.

Fagin:1999:CFI

[1544] Ronald Fagin. Combining fuzzy information from multiple systems. Journal of Computer and System Sciences, 58(1):83–99, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916002.

Henzinger:1999:EF

[1545] Monika R. Henzinger and Ronitt Rubinfeld. Editors’ foreword. Journal of Computer and System Sciences, 58(1):100, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915415.

Blum:1999:CFA

[1546] Avrim Blum, R. Ravi, and Santosh Vempala. A constant-factor approximation algorithm for the k-MST problem. Journal of Computer and System Sciences, 58(1): 101–108, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000097915427.

Kearns:1999:BAT

[1547] Michael Kearns and Yishay Mansour. On the boosting ability of top-down decision tree learning algorithms. Journal of Computer and System Sciences, 58 (1):109–128, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915439.

Kushilevitz:1999:CLS

[1548] Eyal Kushilevitz, Rafail Ostrovsky, and Adi Rosén. Characterizing linear size circuits in terms of privacy. Journal of Computer and System Sciences, 58 (1):129–136, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915440.

Alon:1999:SCA

[1549] Noga Alon, Yossi Matias, and Mario Szegedy. The space complexity of approximating the frequency moments. Journal of Computer and System Sciences, 58(1): 137–147, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915452.

Nisan:1999:ERS

[1550] Noam Nisan and Amnon Ta-Shma. Extracting randomness: a survey and new constructions. Journal of Computer and System Sciences, 58(1): 148–173, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915464.

Ben-Dor:1999:NIB

[1551] Amir Ben-Dor, Anna Karlin, Nathan Linial, and Yuri Rabinovich. A note on the influence of an ϵ-biased random source. Journal of Computer and System Sciences, 58(1): 174–176, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097915518.

Schaefer:1999:DVC

[1552] Marcus Schaefer. Deciding the Vapnik–Cervonenkis dimension is PΣ_{3}^{p}- complete. Journal of Computer and System Sciences, 58(1):177– 182, February 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098916026.

Ratsaby:1999:LRF

[1553] Joel Ratsaby and Vitaly Maiorov. On the learnability of rich function classes. Journal of Computer and System Sciences, 58(1): 183–192, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009891604X.

Arora:1999:PTA

[1554] Sanjeev Arora, David Karger, and Marek Karpinski. Polynomial time approximation schemes for dense instances of NP-hard problems. Journal of Computer and System Sciences, 58 (1):193–210, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098916051.

Case:1999:MML

[1555] John Case and Mark A. Fulk. Maximal machine learnable classes. Journal of Computer and System Sciences, 58 (1):211–214, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009791518X.

Antonoiu:1999:SSD

[1556] Gheorghe Antonoiu and Pradip K. Srimani. A self-stabilizing distributed algorithm to find the median of a tree graph. Journal of Computer and System Sciences, 58(1): 215–221, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098916063.

Lund:1999:PAD

[1557] Carsten Lund, Steven Phillips, and Nick Reingold. Paging against a distribution and IP networking. Journal of Computer and System Sciences, 58 (1):222–231, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000097914987.

Goldberg:1999:APB

[1558] Leslie Ann Goldberg and Philip D.

MacKenzie. Analysis of practical backoff protocols for contention resolution with multiple servers. Journal of Computer and System Sciences, 58 (1):232–258, February 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098915902.

Agarwal:1999:GEF

[1559] Pankaj K. Agarwal and Stephen Cook. Guest Editors’ foreword. Journal of Computer and System Sciences, 58 (2):259, April 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916318.

Beals:1999:AMG

[1560] Robert Beals. Algorithms for matrix groups and the Tits alternative. Journal of Computer and System Sciences, 58(2):260–279, April 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916142.

Cai:1999:SHS

[1561] Jin-Yi Cai and D. Sivakumar. Sparse hard sets for P: Resolution of a conjecture of Hartmanis. Journal of Computer and System Sciences, 58 (2):280–296, April 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916154.

Dey:1999:TCS

[1562] Tamal K. Dey and Sumanta Guha. Transforming curves on surfaces. Journal of Computer and System Sciences, 58(2):297–325, April 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916191.

Haken:1999:ELB

[1563] Armin Haken and Stephen A. Cook.

An exponential lower bound for the size of monotone real circuits. Journal of Computer and System Sciences, 58(2):326–335, April 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098916178.

Naor:1999:STA

[1564] Moni Naor and Omer Reingold. Synthesizers and their application to the parallel construction of pseudorandom functions. Journal of Computer and System Sciences, 58(2): 336–375, April 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009891618X.

Saks:1999:BD

[1565] Michael Saks and Shiyu Zhou.

BP_{H}SPACE(S) ⊆ DSPACE (S^{3/2}). Journal of Computer and System Sciences, 58(2):376–403, April 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000098916166.

Anonymous:1999:EF

[1566] Anonymous. Editor’s foreword. Journal of Computer and System Sciences, 58(3):405–406, June 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916409.

Papadimitriou:1999:CDQ

[1567] Christos H. Papadimitriou and Mihalis Yannakakis. On the complexity of database queries. Journal of Computer and System Sciences, 58 (3):407–427, June 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916264.

Abiteboul:1999:RPQ

[1568] Serge Abiteboul and Victor Vianu. Regular path queries with constraints. Journal of Computer and System Sciences, 58(3):428– 452, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916276.

Mecca:1999:CP

[1569] Giansalvatore Mecca and Paolo Atzeni. Cut and paste. Journal of Computer and System Sciences, 58(3): 453–482, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916239.

Gyssens:1999:CGQ

[1570] Marc Gyssens, Jan Van den Bussche, and Dirk Van Gucht. Complete geometric query languages. Journal of Computer and System Sciences, 58(3):483–511, June 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916306.

Albert:1999:EKR

[1571] Joseph Albert, Yannis Ioannidis, and Raghu Ramakrishnan. Equivalence of keyed relational schemas by conjunctive queries. Journal of Computer and System Sciences, 58(3): 512–534, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916288.

Dumortier:1999:DSS

[1572] Freddy Dumortier, Marc Gyssens, Luc Vandeurzen, and Dirk Van Gucht. On the decidability of semilinearity for semialgebraic sets and its implications for spatial databases. Journal of Computer and System Sciences, 58(3):535–571, June 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916252. See

correction [1605].

Buss:1999:CCS

[1573] Jonathan F. Buss, Gudmund S. Frandsen, and Jeffrey O. Shallit. The computational complexity of some problems of linear algebra. Journal of Computer and System Sciences, 58 (3):572–596, June 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098916087.

Manzini:1999:ALC

[1574] Giovanni Manzini and Luciano Margara. Attractors of linear cellular automata. Journal of Computer and System Sciences, 58(3):

597–610, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916099.

Dassow:1999:SCM

[1575] Jürgen Dassow and Victor Mitrana. Stack cooperation in multistack pushdown automata. Journal of Computer and System Sciences, 58(3): 611–621, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916105.

Chen:1999:TBJ

[1576] Jianer Chen, Donald K. Friesen, and Hao Zheng. Tight bound on Johnson’s algorithm for maximum satisfiability. Journal of Computer and System Sciences, 58(3): 622–640, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916129.

Krawczyk:1999:CFS

[1577] Adam Krawczyk. The complexity of finding a second Hamiltonian cycle in cubic graphs. Journal of Computer and System Sciences, 58 (3):641–647, June 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000098916117.

Hemaspaandra:1999:CST

[1578] Lane A. Hemaspaandra and Jörg Rothe. Creating strong, total, commutative, associative one-way functions from any one-way function in complexity theory. Journal of Computer and System Sciences, 58(3): 648–659, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000098916130.

Anonymous:1999:AIVa

[1579] Anonymous. Author index for volume 58. Journal of Computer and System Sciences, 58(3): 660, June 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916410.

Ibarra:1999:TPD

[1580] Oscar H. Ibarra and Jianwen Su.

A technique for proving decidability of containment and equivalence of linear constraint queries. Journal of Computer and System Sciences, 59(1):1–28, August 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916240.

Breitbart:1999:RCD

[1581] Yuri Breitbart and Henry F. Korth. Replication and consistency in a distributed environment. Journal of Computer and System Sciences, 59(1):29–69, August 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916392.

Chen:1999:LUT

[1582] Zhixiang Chen and Foued Ameur. The learnability of unions of two rectangles in the two-dimensional discretized space. Journal of Computer and System Sciences, 59(1): 70–83, August 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916215.

Chaudhuri:1999:SSA

[1583] Pranay Chaudhuri. A self-stabilizing algorithm for detecting fundamental cycles in a graph. Journal of Computer and System Sciences, 59 (1):84–93, August 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916227.

Baudinet:1999:CGD

[1584] Marianne Baudinet, Jan Chomicki, and Pierre Wolper. Constraintgenerating dependencies. Journal of Computer and System Sciences, 59 (1):94–115, August 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009991632X.

Grahne:1999:RAS

[1585] Gösta Grahne, Matti Nykänen, and Esko Ukkonen. Reasoning about strings in databases. Journal of Computer and System Sciences, 59 (1):116–162, August 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916331.

Feigenbaum:1999:GEF

[1586] Joan Feigenbaum and Alan Selman.

Guest Editors’ foreword. Journal of Computer and System Sciences, 59(2):163, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916458.

Allender:1999:IMC

[1587] Eric Allender, Klaus Reinhardt, and Shiyu Zhou. Isolation, matching, and counting uniform and nonuniform upper bounds. Journal of Computer and System Sciences, 59(2): 164–181, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009991646X.

Buhrman:1999:TQ

[1588] Harry Buhrman and Lance Fortnow. Two queries. Journal of Computer and System Sciences, 59(2): 182–194, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916471.

Beimel:1999:ABP

[1589] Amos Beimel and Anna Gál. On arithmetic branching programs. Journal of Computer and System Sciences, 59 (2):195–220, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916483.

Cai:1999:ASW

[1590] Jin-Yi Cai and Ajay Nerurkar. Approximating the SVP to within a factor (1 + 1/ dim ϵ) is NP-hard under randomized reductions. Journal of Computer and System Sciences, 59 (2):221–239, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916495.

Fortnow:1999:CLQ

[1591] Lance Fortnow and John Rogers. Complexity limitations on quantum computation. Journal of Computer and System Sciences, 59(2): 240–252, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916513.

Goldreich:1999:CIS

[1592] Oded Goldreich and Madhu Sudan. Computational indistinguishability: a sample hierarchy. Journal of Computer and System Sciences, 59(2): 253–269, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916525.

Sivakumar:1999:MCS

[1593] D. Sivakumar. On membership comparable sets. Journal of Computer and System Sciences, 59(2): 270–280, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916537.

Watrous:1999:SBQ

[1594] John Watrous. Space-bounded quantum complexity. Journal of Computer and System Sciences, 59(2): 281–326, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916550.

Buhrman:1999:HSH

[1595] Harry Buhrman and Dieter van Melkebeek. Hard sets are hard to find. Journal of Computer and System Sciences, 59(2):327–345, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916501.

Raz:1999:AMG

[1596] Ran Raz, Gábor Tardos, Oleg Verbitsky, and Nikolai Vereshchagin. Arthur–Merlin games in Boolean decision trees. Journal of Computer and System Sciences, 59(2): 346–372, October 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916549.

Gaubert:1999:PNL

[1597] Stéphane Gaubert and Alessandro Giua. Petri net languages and infinite subsets of N^{m}. Journal of Computer and System Sciences, 59(3): 373–391, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916343.

Cesa-Bianchi:1999:ATG

[1598] Nicolò Cesa-Bianchi. Analysis of two gradient-based algorithms for online regression. Journal of Computer and System Sciences, 59(3): 392–411, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916355.

Corradini:1999:MNR

[1599] Flavio Corradini, Rocco De Nicola, and Anna Labella. Models of nondeterministic regular expressions. Journal of Computer and System Sciences, 59 (3):412–449, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916367.

Chen:1999:GII

[1600] Lin Chen. Graph isomorphism and identification matrices: Sequential algorithms. Journal of Computer and System Sciences, 59(3): 450–475, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916422.

Jancar:1999:PNR

[1601] Petr Jancar, Javier Esparza, and Faron Moller. Petri nets and regular processes. Journal of Computer and System Sciences, 59(3): 476–503, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916434.

Weems:1999:BVM

[1602] B. P. Weems. Bistable versions of the marriages and roommates problems. Journal of Computer and System Sciences, 59(3):504– 520, December 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916574.

Liberti:1999:SIC

[1603] Leo Liberti. Structure of the invertible CA transformations group. Journal of Computer and System Sciences, 59 (3):521–536, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916598.

Barbuti:1999:SMC

[1604] Roberto Barbuti, Nicoletta De Francesco, Antonella Santone, and Gigliola Vaglini. Selective Mu-calculus and formula-based equivalence of transition systems. Journal of Computer and System Sciences, 59(3): 537–556, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916604.

Anonymous:1999:CVN

[1605] Anonymous. Corrigendum: Volume 58, Number 3 (1999), in Article No.

“On the Decidability of Semilinearity for Semialgebraic Sets and Its Implications for Spatial Databases,” by Freddy Dumortier, Marc Gyssens, Luc Vandeurzen, and Dirk Van Gucht, pages 535–571. Journal of Computer and System Sciences, 59(3): 557–562, December 1999. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916963. See

[1572].

Anonymous:1999:AIVb

[1606] Anonymous. Author index for volume 59. Journal of Computer and System Sciences, 59(3): 563, December 1999. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916975.

Ma:2000:IDP

[1607] Bin Ma and Lusheng Wang. On the inapproximability of disjoint paths and minimum Steiner forest with bandwidth constraints. Journal of Computer and System Sciences, 60 (1):1–12, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916616.

Fokkink:2000:RBB

[1608] Wan Fokkink. Rooted branching bisimulation as a congruence. Journal of Computer and System Sciences, 60 (1):13–37, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009991663X.

Cao:2000:GES

[1609] Jinde Cao. Global exponential stability and periodic solutions of delayed cellular neural networks. Journal of Computer and System Sciences, 60(1):38–46, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916586.

Okamoto:2000:RBS

[1610] Tatsuaki Okamoto. On relationships between statistical zero-knowledge proofs. Journal of Computer and System Sciences, 60(1):47–108, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000099916641.

Lange:2000:DIR

[1611] Klaus-Jörn Lange and Rolf Niedermeier. Data independence of read, write, and control structures in PRAM computations. Journal of Computer and System Sciences, 60(1): 109–144, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916653.

Bekesi:2000:LTP

[1612] József Békési, Gábor Galambos, and Hans Kellerer. A 5/4 linear time bin packing algorithm. Journal of Computer and System Sciences, 60 (1):145–160, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916677.

Servedio:2000:CSC

[1613] Rocco A. Servedio. Computational sample complexity and attributeefficient learning. Journal of Computer and System Sciences, 60(1): 161–178, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916665.

Cao:2000:EDA

[1614] Jinde Cao and Qing Tao. Estimation of the domain of attraction and the convergence rate of a Hopfield associative memory and an application. Journal of Computer and System Sciences, 60 (1):179–186, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916793.

Baier:2000:DBS

[1615] Christel Baier, Bettina Engelen, and Mila Majster-Cederbaum. Deciding bisimilarity and similarity for probabilistic processes. Journal of Computer and System Sciences, 60(1): 187–231, February 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916835.

Raghavan:2000:EF

[1616] Vijay Raghavan. Editor’s foreword. Journal of Computer and System Sciences, 60(2):233, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916689.

Case:2000:RLA

[1617] John Case, Sanjay Jain, Matthias Ott, Arun Sharma, and Frank Stephan. Robust learning aided by context. Journal of Computer and System Sciences, 60(2):234–257, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916379.

Birkendorf:2000:SRA

[1618] Andreas Birkendorf, Norbert Klasner, Christian Kuhlmann, and Hans U. Simon. Structural results about exact learning with unspecified attribute values. Journal of Computer and System Sciences, 60(2): 258–277, April 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916380.

Rosin:2000:SCM

[1619] Christopher D. Rosin. Sample complexity of model-based search. Journal of Computer and System Sciences, 60(2):278–301, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916446.

Reischuk:2000:ACO

[1620] Rüdiger Reischuk and Thomas Zeugmann. An average-case optimal one-variable pattern language learner. Journal of Computer and System Sciences, 60(2):302– 335, April 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916690. Eleventh

Annual Conference on Computational Learning Theory (Madison, WI, 1998).

Balcazar:2000:EF

[1621] José L. Balcázar. Editor’s foreword. Journal of Computer and System Sciences, 60(2):336, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916707.

Fortnow:2000:TST

[1622] Lance Fortnow. Time-space tradeoffs for satisfiability. Journal of Computer and System Sciences, 60 (2):337–353, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916719.

Lange:2000:RSE

[1623] Klaus-Jörn Lange, Pierre McKenzie, and Alain Tapp. Reversible space equals deterministic space. Journal of Computer and System Sciences, 60 (2):354–367, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916720.

Moore:2000:CEN

[1624] Cristopher Moore, Denis Thérien, François Lemieux, Joshua Berman, and Arthur Drisko. Circuits and expressions with nonassociative gates. Journal of Computer and System Sciences, 60(2):368–394, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916732.

Agrawal:2000:TAA

[1625] Manindra Agrawal, Eric Allender, and Samir Datta. On TC^{0}, AC^{0}, and arithmetic circuits. Journal of Computer and System Sciences, 60 (2):395–421, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916756.

Beigel:2000:CPP

[1626] Richard Beigel and Bin Fu. Circuits over PP and PL. Journal of Computer and System Sciences, 60 (2):422–441, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916768.

Hammer:2000:ISE

[1627] Daniel Hammer, Andrei Romashchenko, Alexander Shen, and Nikolai Vereshchagin. Inequalities for Shannon entropy and Kolmogorov complexity. Journal of Computer and System Sciences, 60(2):442–464, April 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009991677X.

Downey:2000:URL

[1628] Rod Downey and André Nies. Undecidability results for low complexity time classes. Journal of Computer and System Sciences, 60(2): 465–479, April 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916781.

Anonymous:2000:PAFa

[1629] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 60(2): 480, April 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917148.

Graham:2000:GEF

[1630] Fan Chung Graham. Guest Editor’s foreword. Journal of Computer and System Sciences, 60(3): 481, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200009991680X.

Aiello:2000:APR

[1631] William Aiello, Eyal Kushilevitz, Rafail Ostrovsky, and Adi Rosén. Adaptive packet routing for bursty adversarial traffic. Journal of Computer and System Sciences, 60(3): 482–509, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916811.

Feige:2000:ABV

[1632] Uriel Feige. Approximating the bandwidth via volume respecting embeddings. Journal of Computer and System Sciences, 60(3): 510–539, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916823.

Goldreich:2000:LNL

[1633] Oded Goldreich and Shafi Goldwasser.

On the limits of nonapproximability of lattice problems. Journal of Computer and System Sciences, 60 (3):540–563, June 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916860.

Calinescu:2000:IAA

[1634] Gruia Calinescu, Howard Karloff, and Yuval Rabani. An improved approximation algorithm for MUL-TIWAY CUT. Journal of Computer and System Sciences, 60(3): 564–574, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916872.

Khanna:2000:IDB

[1635] Sanjeev Khanna and Shiyu Zhou. On indexed data broadcast. Journal of Computer and System Sciences, 60 (3):575–591, June 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916884.

Gertner:2000:PDP

[1636] Yael Gertner, Yuval Ishai, Eyal Kushilevitz, and Tal Malkin. Protecting data privacy in private information retrieval schemes. Journal of Computer and System Sciences, 60(3):592–629, June 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916896.

Broder:2000:MWI

[1637] Andrei Z. Broder, Moses Charikar, Alan M. Frieze, and Michael Mitzenmacher. Min-wise independent permutations. Journal of Computer and System Sciences, 60(3):630– 659, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916902.

Ajtai:2000:CMN

[1638] Miklos Ajtai, Ronald Fagin, and Larry Stockmeyer. The closure of monadic NP. Journal of Computer and System Sciences, 60(3): 660–716, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916914.

Ergun:2000:SC

[1639] Funda Ergün, Sampath Kannan, S. Ravi Kumar, Ronitt Rubinfeld, and Mahesh Viswanathan. Spotcheckers. Journal of Computer and System Sciences, 60(3):717– 751, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916926.

Anonymous:2000:PAFb

[1640] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 60(3): 752, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000091715X.

Anonymous:2000:AIVa

[1641] Anonymous. Author index for volume 60. Journal of Computer and System Sciences, 60(3): 753, June 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917161.

Bloem:2000:CTT

[1642] Roderick Bloem and Joost Engelfriet.

A comparison of tree transductions defined by monadic second order logic and by attribute grammars. Journal of Computer and System Sciences, 61(1):1–50, August 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916847.

Catalano:2000:NES

[1643] Dario Catalano and Rosario Gennaro. New efficient and secure protocols for verifiable signature sharing and other applications. Journal of Computer and System Sciences, 61(1):51–80, August 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916859.

Chari:2000:IAA

[1644] Suresh Chari, Pankaj Rohatgi, and Aravind Srinivasan. Improved algorithms via approximations of probability distributions. Journal of Computer and System Sciences, 61(1): 81–107, August 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916951.

Paun:2000:CM

[1645] Gheorghe Paun. Computing with membranes. Journal of Computer and System Sciences, 61(1):108– 143, August 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916938.

Anonymous:2000:PAFc

[1646] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 61(1): 144, August 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917173.

Paredaens:2000:GEF

[1647] Jan Paredaens. Guest Editor’s foreword. Journal of Computer and System Sciences, 61(2):145, October 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000000917069.

Buneman:2000:PCS

[1648] Peter Buneman, Wenfei Fan, and Scott Weinstein. Path constraints in semistructured databases. Journal of Computer and System Sciences, 61 (2):146–193, October 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917100.

Agarwal:2000:ESL

[1649] Pankaj K. Agarwal, Lars Arge, Jeff Erickson, Paulo G. Franciosa, and Jeffrey Scott Vitter. Efficient searching with linear constraints. Journal of Computer and System Sciences, 61 (2):194–216, October 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917094.

Papadimitriou:2000:LSI

[1650] Christos H. Papadimitriou, Prabhakar Raghavan, Hisao Tamaki, and Santosh Vempala. Latent semantic indexing: a probabilistic analysis. Journal of Computer and System Sciences, 61 (2):217–235, October 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917112.

Abiteboul:2000:RTE

[1651] Serge Abiteboul, Victor Vianu, Brad Fordham, and Yelena Yesha. Relational transducers for electronic commerce. Journal of Computer and System Sciences, 61(2):236– 269, October 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917082.

Segoufin:2000:QSD

[1652] Luc Segoufin and Victor Vianu. Querying spatial databases via topological invariants. Journal of Computer and System Sciences, 61(2): 270–301, October 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917124.

Kolaitis:2000:CQC

[1653] Phokion G. Kolaitis and Moshe Y. Vardi. Conjunctive-query containment and constraint satisfaction. Journal of Computer and System Sciences, 61 (2):302–332, October 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917136.

Anonymous:2000:PAFd

[1654] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 61(2): 333, October 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917355.

Ambos-Spies:2000:SNC

[1655] Klaus Ambos-Spies and Levke Bentzien. Separating NP-completeness notions under strong hypotheses. Journal of Computer and System Sciences, 61(3): 335–361, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916744.

Bellare:2000:SCB

[1656] Mihir Bellare, Joe Kilian, and Phillip Rogaway. The security of the cipher block chaining message authentication code. Journal of Computer and System Sciences, 61(3): 362–399, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200009991694X.

Quaglia:2000:NZC

[1657] Francesco Quaglia, Roberto Baldoni, and Bruno Ciciani. On the no-Z-cycle property in distributed executions. Journal of Computer and System Sciences, 61(3):400– 427, December 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917203.

Kearns:2000:TPS

[1658] Michael Kearns and Dana Ron. Testing problems with sublearning sample complexity. Journal of Computer and System Sciences, 61(3): 428–456, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000099916562.

Cheng:2000:FTM

[1659] Sheng-Tzong Cheng, Chia-Mei Chen, and Satish K. Tripathi. A faulttolerance model for multiprocessor real-time systems. Journal of Computer and System Sciences, 61(3): 457–477, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917045.

Eiter:2000:DHT

[1660] Thomas Eiter, Toshihide Ibaraki, and Kazuhisa Makino. On the difference of Horn theories. Journal of Computer and System Sciences, 61(3): 478–507, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917197.

Larsen:2000:ATR

[1661] Kim S. Larsen. AVL trees with relaxed balance. Journal of Computer and System Sciences, 61(3): 508–522, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917057.

Takeuchi:2000:LPP

[1662] Jun ichi Takeuchi, Naoki Abe, and Shun ichi Amari. The lobpass problem. Journal of Computer and System Sciences, 61(3): 523–557, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917185.

Bollig:2000:AOB

[1663] Beate Bollig and Ingo Wegener.

Asymptotically optimal bounds for OBDDs and the solution of some basic OBDD problems. Journal of Computer and System Sciences, 61(3): 558–579, December 2000. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917331.

Anonymous:2000:PAFe

[1664] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 61(3): 580, December 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917380.

Anonymous:2000:AIVb

[1665] Anonymous. Author index for volume 61. Journal of Computer and System Sciences, 61(3): 581, December 2000. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917392.

Ghilezan:2001:FIT

[1666] Silvia Ghilezan. Full intersection types and topologies in Lambda calculus. Journal of Computer and System Sciences, 62(1):1–14, February 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917033.

Giannotti:2001:SEP

[1667] Fosca Giannotti, Dino Pedreschi, and Carlo Zaniolo. Semantics and expressive power of nondeterministic constructs in deductive databases. Journal of Computer and System Sciences, 62(1):15–42, February 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916999.

Liberatore:2001:BRU

[1668] Paolo Liberatore and Marco Schaerf. Belief revision and update: Complexity of model checking. Journal of Computer and System Sciences, 62 (1):43–72, February 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000916982.

Philippou:2001:PCA

[1669] Anna Philippou and David Walker. A process-calculus analysis of concurrent operations on B-trees. Journal of Computer and System Sciences, 62 (1):73–122, February 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917215.

Goldman:2001:ALG

[1670] Sally A. Goldman, Stephen S. Kwek, and Stephen D. Scott. Agnostic learning of geometric patterns. Journal of Computer and System Sciences, 62 (1):123–151, February 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917239.

Cole:2001:BSV

[1671] Richard J. Cole, Bruce M. Maggs, and Ramesh K. Sitaraman. On the benefit of supporting virtual channels in wormhole routers. Journal of Computer and System Sciences, 62 (1):152–177, February 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000091701X.

Jain:2001:RLR

[1672] Sanjay Jain, Carl Smith, and Rolf Wiehagen. Robust learning is rich. Journal of Computer and System Sciences, 62(1):178–212, February 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000000917008.

Anonymous:2001:PAFa

[1673] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 62(1): 213, February 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917453.

Fortnow:2001:GEF

[1674] Lance Fortnow. Guest Editor’s foreword. Journal of Computer and System Sciences, 62(2):215, March 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000000917240.

Rojas:2001:CAG

[1675] J. Maurice Rojas. Computational arithmetic geometry: I. Sentences nearly in the polynomial hierarchy. Journal of Computer and System Sciences, 62(2):216– 235, March 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917288.

Sudan:2001:PGX

[1676] Madhu Sudan, Luca Trevisan, and Salil Vadhan. Pseudorandom generators without the XOR lemma. Journal of Computer and System Sciences, 62(2):236–266, March 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917306.

Buss:2001:LGB

[1677] Sam Buss, Dima Grigoriev, Russell Impagliazzo, and Toniann Pitassi. Linear gaps between degrees for the polynomial calculus modulo distinct primes. Journal of Computer and System Sciences, 62(2): 267–289, March 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917264.

Schaefer:2001:GRT

[1678] Marcus Schaefer. Graph Ramsey theory and the polynomial hierarchy. Journal of Computer and System Sciences, 62(2):290– 322, March 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000091729X.

Ponzio:2001:CCP

[1679] Stephen J. Ponzio, Jaikumar Radhakrishnan, and S. Venkatesh. The communication complexity of pointer chasing. Journal of Computer and System Sciences, 62(2):323– 355, March 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917318.

Allender:2001:LBP

[1680] Eric Allender, Michael Saks, and Igor Shparlinski. A lower bound for primality. Journal of Computer and System Sciences, 62(2): 356–366, March 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917252.

Impagliazzo:2001:CS

[1681] Russell Impagliazzo and Ramamohan Paturi. On the complexity of k-SAT. Journal of Computer and System Sciences, 62(2): 367–375, March 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917276.

Watrous:2001:QSC

[1682] John Watrous. Quantum simulations of classical random walks and undirected graph connectivity. Journal of Computer and System Sciences, 62 (2):376–391, March 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000091732X.

Anonymous:2001:PAFb

[1683] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 62(2): 392, March 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917465.

Awerbuch:2001:CRV

[1684] Baruch Awerbuch, Yossi Azar, Serge Plotkin, and Orli Waarts. Competitive routing of virtual circuits with unknown duration. Journal of Computer and System Sciences, 62 (3):385–397, May 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000099916628.

Reif:2001:POS

[1685] John H. Reif. Parallel output-sensitive algorithms for combinatorial and linear algebra problems. Journal of Computer and System Sciences, 62 (3):398–412, May 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917409.

Case:2001:SLT

[1686] John Case and Sanjay Jain. Synthesizing learners tolerating computable noisy data. Journal of Computer and System Sciences, 62(3): 413–441, May 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917367.

Blondel:2001:SSL

[1687] Vincent D. Blondel, Olivier Bournez, Pascal Koiran, and John N. Tsitsiklis. The stability of saturated linear dynamical systems is undecidable. Journal of Computer and System Sciences, 62(3):442–462, May 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917379.

Buccafurri:2001:AFH

[1688] Francesco Buccafurri, Thomas Eiter, Georg Gottlob, and Nicola Leone. On ACTL formulas having linear counterexamples. Journal of Computer and System Sciences, 62(3): 463–515, May 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917343.

Li:2001:IBS

[1689] Yi Li, Philip M. Long, and Aravind Srinivasan. Improved bounds on the sample complexity of learning. Journal of Computer and System Sciences, 62(3):516–527, May 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917410.

Cao:2001:EDA

[1690] Jinde Cao and Qing Tao. Estimation on domain of attraction and convergence rate of Hopfield continuous feedback neural networks. Journal of Computer and System Sciences, 62(3):528–534, May 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917227.

Anonymous:2001:PAFc

[1691] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 62 (3):535, May 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917581.

Ishihara:2001:RCR

[1692] Yasunori Ishihara, Shougo Shimizu, Hiroyuki Seki, and Minoru Ito. Refinements of complexity results on type consistency for object-oriented databases. Journal of Computer and System Sciences, 62(4):537– 564, June 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917507.

Hromkovic:2001:TRE

[1693] Juraj Hromkovic, Sebastian Seibert, and Thomas Wilke. Translating regular expressions into small ϵ-free nondeterministic finite automata. Journal of Computer and System Sciences, 62(4):565–588, June 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917489.

Aumann:2001:LCT

[1694] Yonatan Aumann, Johan Håstad, Michael O. Rabin, and Madhu Sudan. Linear-consistency testing. Journal of Computer and System Sciences, 62(4):589–607, June 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917477.

Kim:2001:SRS

[1695] Changwook Kim. Separation results for separated apex NLC and NCE graph languages. Journal of Computer and System Sciences, 62 (4):608–628, June 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917490.

Lautemann:2001:DCA

[1696] Clemens Lautemann, Pierre McKenzie, Thomas Schwentick, and Heribert Vollmer. The descriptive complexity approach to LOGCFL. Journal of Computer and System Sciences, 62(4):629–652, June 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917422.

Hutter:2001:NEB

[1697] Marcus Hutter. New error bounds for Solomonoff prediction. Journal of Computer and System Sciences, 62(4):653–667, June 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917434.

Kesten:2001:VAA

[1698] Yonit Kesten, Amir Pnueli, and Moshe Y. Vardi. Verification by augmented abstraction: The automatatheoretic view. Journal of Computer and System Sciences, 62(4): 668–690, June 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000000917446.

Anonymous:2001:PAFd

[1699] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 62(4): 691, June 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917702.

Anonymous:2001:AIVa

[1700] Anonymous. Author index for volume 62. Journal of Computer and System Sciences, 62(4): 692, June 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917714.

Graham:2001:EF

[1701] Fan Chung Graham. Editor’s foreword. Journal of Computer and System Sciences, 63(1):1, August 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917519.

Kleinberg:2001:FRL

[1702] Jon Kleinberg, Yuval Rabani, and Éva Tardos. Fairness in routing and load balancing. Journal of Computer and System Sciences, 63(1): 2–20, August 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917520.

Feigenbaum:2001:SCM

[1703] Joan Feigenbaum, Christos H. Papadimitriou, and Scott Shenker. Sharing the cost of multicast transmissions. Journal of Computer and System Sciences, 63(1):21–41, August 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000001917544.

Kumar:2001:RSP

[1704] Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, and Andrew Tomkins. Recommendation systems: a probabilistic analysis. Journal of Computer and System Sciences, 63 (1):42–61, August 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000191757X.

Goel:2001:CFT

[1705] Ashish Goel, Adam Meyerson, and Serge Plotkin. Combining fairness with throughput: Online routing with multiple objectives. Journal of Computer and System Sciences, 63 (1):62–79, August 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917556.

Hanlon:2001:CCM

[1706] Philip J. Hanlon, Dean Chung, Siddhartha Chatterjee, Daniela Genius, Alvin R. Lebeck, and Erin Parker. The combinatorics of cache misses during matrix multiplication. Journal of Computer and System Sciences, 63 (1):80–126, August 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917568.

Adler:2001:CUE

[1707] Micah Adler and Tom Leighton. Compression using efficient multicasting. Journal of Computer and System Sciences, 63(1):127– 145, August 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917532.

Anonymous:2001:PAFe

[1708] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 63(1): 146, August 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000191788X.

Pudlak:2001:GEF

[1709] Pavel Pudlák. Guest Editor’s foreword. Journal of Computer and System Sciences, 63(2):147, September 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000191760X.

Ambainis:2001:CCE

[1710] Andris Ambainis, Harry Buhrman, William Gasarch, Bala Kalyanasundaram, and Leen Torenvliet. The communication complexity of enumeration, elimination, and selection. Journal of Computer and System Sciences, 63(2): 148–185, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917611.

Barrington:2001:CSP

[1711] David Mix Barrington, Peter Kadau, Klaus-Jörn Lange, and Pierre McKenzie. On the complexity of some problems on groups input as multiplication tables. Journal of Computer and System Sciences, 63(2): 186–200, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917647.

Berthiaume:2001:QKC

[1712] André Berthiaume, Wim van Dam, and Sophie Laplante. Quantum Kolmogorov complexity. Journal of Computer and System Sciences, 63(2): 201–221, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917659.

Giel:2001:BPS

[1713] Oliver Giel. Branching program size is almost linear in formula size. Journal of Computer and System Sciences, 63 (2):222–235, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917623.

Kabanets:2001:EAH

[1714] Valentine Kabanets. Easiness assumptions and hardness tests: Trading time for zero error. Journal of Computer and System Sciences, 63(2): 236–252, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917635.

Mulmuley:2001:LBS

[1715] Ketan Mulmuley and Pradyut Shah. A lower bound for the shortest path problem. Journal of Computer and System Sciences, 63(2): 253–267, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917660.

Tourlakis:2001:TST

[1716] Iannis Tourlakis. Time-space tradeoffs for SAT on nonuniform machines. Journal of Computer and System Sciences, 63(2):268–287, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000001917672.

Yang:2001:ICE

[1717] Ke Yang. Integer circuit evaluation is PSPACE-complete. Journal of Computer and System Sciences, 63(2): 288–303, September 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917684.

Anonymous:2001:PAFf

[1718] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 63(2): 304, September 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917891.

Jain:2001:LLT

[1719] Sanjay Jain, Efim Kinber, and Rolf Wiehagen. Language learning from texts: Degrees of intrinsic complexity and their characterizations. Journal of Computer and System Sciences, 63 (3):305–354, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917593.

Gupta:2001:QNN

[1720] Sanjay Gupta and R. K. P. Zia. Quantum neural networks. Journal of Computer and System Sciences, 63(3): 355–383, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917696.

Leung:2001:TLB

[1721] Hing Leung. Tight lower bounds on the size of sweeping automata. Journal of Computer and System Sciences, 63 (3):384–393, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917830.

Bedon:2001:LWD

[1722] Nicolas Bedon. Logic over words on denumerable ordinals. Journal of Computer and System Sciences, 63(3): 394–431, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917829.

Chung:2001:DRP

[1723] Fan Chung, Mark Garrett, Ronald Graham, and David Shallcross. Distance realization problems with applications to Internet tomography. Journal of Computer and System Sciences, 63(3):432–448, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000001917854.

Lokam:2001:SMM

[1724] Satyanarayana V. Lokam. Spectral methods for matrix rigidity with applications to size-depth trade-offs and communication complexity. Journal of Computer and System Sciences, 63 (3):449–473, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917866.

Peleg:2001:LCV

[1725] David Peleg and Eilon Reshef. Low complexity variants of the arrow distributed directory. Journal of Computer and System Sciences, 63(3): 474–485, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917878.

Sethuraman:2001:PTA

[1726] Jay Sethuraman and Chung-Piaw Teo.

A polynomial-time algorithm for the bistable roommates problem. Journal of Computer and System Sciences, 63 (3):486–497, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000191791X.

Haastad:2001:SSL

[1727] Johan Håstad. A slight sharpening of LMN. Journal of Computer and System Sciences, 63(3): 498–508, November 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918033.

Anonymous:2001:PAFg

[1728] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 63(3): 509, November 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918325.

Motwani:2001:GEF

[1729] Rajeev Motwani. Guest Editor’s foreword. Journal of Computer and System Sciences, 63(4): 511, December 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917775.

Impagliazzo:2001:WPS

[1730] Russell Impagliazzo, Ramamohan Paturi, and Francis Zane. Which problems have strongly exponential complexity? Journal of Computer and System Sciences, 63(4): 512–530, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000191774X.

delaVega:2001:RAS

[1731] W. Fernandez de la Vega and Claire Kenyon. A randomized approximation scheme for metric MAX-CUT. Journal of Computer and System Sciences, 63 (4):531–541, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917726.

Beame:2001:TST

[1732] Paul Beame, T. S. Jayram, and Michael Saks. Time-space tradeoffs for branching programs. Journal of Computer and System Sciences, 63(4): 542–572, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917787.

Adler:2001:PAC

[1733] Micah Adler and Bruce M. Maggs.

Protocols for asymmetric communication channels. Journal of Computer and System Sciences, 63(4): 573–596, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917799.

Umans:2001:MED

[1734] Christopher Umans. The minimum equivalent DNF problem and shortest implicants. Journal of Computer and System Sciences, 63(4): 597–611, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917751.

Russell:2001:PIL

[1735] Alexander Russell and David Zuckerman. Perfect information leader election in log^{∗} n + O(1) rounds. Journal of Computer and System Sciences, 63 (4):612–626, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917763.

Indyk:2001:ANN

[1736] Piotr Indyk. On approximate nearest neighbors under l_{∞} norm. Journal of Computer and System Sciences, 63 (4):627–638, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917817.

Feige:2001:HSG

[1737] Uriel Feige and Joe Kilian. Heuristics for semirandom graph problems. Journal of Computer and System Sciences, 63(4):639–671, December 2001. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000001917738.

Impagliazzo:2001:RVT

[1738] Russell Impagliazzo and Avi Wigderson. Randomness vs time: Derandomization under a uniform assumption. Journal of Computer and System Sciences, 63(4):672– 688, December 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917805.

Anonymous:2001:PAFh

[1739] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 63(4): 689, December 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918379.

Anonymous:2001:AIVb

[1740] Anonymous. Author index for volume 63. Journal of Computer and System Sciences, 63(4): 690, December 2001. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918380.

Goldman:2002:GEF

[1741] Sally Goldman. Guest Editor’s foreword. Journal of Computer and System Sciences, 64(1):1, February 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000001917921.

Balcazar:2002:NAC

[1742] José L. Balcázar, Jorge Castro, and David Guijarro. A new abstract combinatorial dimension for exact learning via queries. Journal of Computer and System Sciences, 64(1):2– 21, February 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917945.

Ben-David:2002:CCD

[1743] Shai Ben-David, Nadav Eiron, and Hans Ulrich Simon. The computational complexity of densest region detection. Journal of Computer and System Sciences, 64(1):22– 47, February 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917970.

Auer:2002:ASC

[1744] Peter Auer, Nicolò Cesa-Bianchi, and Claudio Gentile. Adaptive and self-confident on-line learning algorithms. Journal of Computer and System Sciences, 64(1):48–75, February 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917957.

Forster:2002:REI

[1745] Jürgen Forster and Manfred K. Warmuth. Relative expected instantaneous loss bounds. Journal of Computer and System Sciences, 64(1): 76–102, February 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001917982.

Mansour:2002:BUB

[1746] Yishay Mansour and David McAllester. Boosting using branching programs. Journal of Computer and System Sciences, 64(1):103–112, February 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000001917969.

Freund:2002:DGB

[1747] Yoav Freund and Manfred Opper. Drifting games and Brownian motion. Journal of Computer and System Sciences, 64(1): 113–132, February 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918021.

Bartlett:2002:EAB

[1748] Peter L. Bartlett and Jonathan Baxter. Estimation and approximation bounds for gradient-based reinforcement learning. Journal of Computer and System Sciences, 64(1): 133–150, February 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917933.

Anonymous:2002:PAFa

[1749] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 64(1): 151, February 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918398.

Fulk:2002:IIA

[1750] Mark Fulk. Inductive inference with additional information. Journal of Computer and System Sciences, 64 (2):153–159, March 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000000917021.

Bazgan:2002:EAA

[1751] Cristina Bazgan, Miklos Santha, and Zsolt Tuza. Efficient approximation algorithms for the subset-sums equality problem. Journal of Computer and System Sciences, 64(2): 160–170, March 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917842.

Grosse-Rhode:2002:FSS

[1752] Martin Große-Rhode, Francesco Parisi Presicce, and Marta Simeoni. Formal software specification with refinements and modules of typed graph transformation systems. Journal of Computer and System Sciences, 64 (2):171–218, March 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918008.

Garofalakis:2002:CLS

[1753] Minos Garofalakis, Yannis Ioannidis, Banu Özden, and Avi Silberschatz. Competitive on-line scheduling of continuous-media streams. Journal of Computer and System Sciences, 64(2):219–248, March 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917994.

Drewes:2002:HGT

[1754] Frank Drewes, Berthold Hoffmann, and Detlef Plump. Hierarchical graph transformation. Journal of Computer and System Sciences, 64(2): 249–283, March 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001917908.

Westerdale:2002:MAC

[1755] T. H. Westerdale. Minimality of an automaton cascade decomposition for learning system environments. Journal of Computer and System Sciences, 64(2):284–307, March 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000191801X.

Bridson:2002:CFL

[1756] Martin R. Bridson and Robert H.

Gilman. Context-free languages of sub-exponential growth. Journal of Computer and System Sciences, 64 (2):308–310, March 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918045.

Hemaspaandra:2002:RNS

[1757] Lane A. Hemaspaandra, Mitsunori Ogihara, and Gerd Wechsung. Reducing the number of solutions of NP functions. Journal of Computer and System Sciences, 64(2): 311–328, March 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000191815X.

Helary:2002:ICA

[1758] J. M. Hélary, A. Mostefaoui, and M. Raynal. Interval consistency of asynchronous distributed computations. Journal of Computer and System Sciences, 64(2):329– 349, March 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918197.

Engelfriet:2002:OSL

[1759] Joost Engelfriet and Sebastian Maneth.

Output string languages of compositions of deterministic macro tree transducers. Journal of Computer and System Sciences, 64(2):350– 395, March 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918161.

Diekert:2002:LEC

[1760] Volker Diekert and Paul Gastin. LTL is expressively complete for Mazurkiewicz traces. Journal of Computer and System Sciences, 64 (2):396–418, March 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918173.

Sieling:2002:LBL

[1761] Detlef Sieling. Lower bounds for linearly transformed OBDDs and FBDDs. Journal of Computer and System Sciences, 64(2):419– 438, March 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918185.

Anonymous:2002:PAFb

[1762] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 64(2): 439, March 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918532.

Ioannidis:2002:GEF

[1763] Yannis Ioannidis and Christos Papadimitriou. Guest Editors’ foreword. Journal of Computer and System Sciences, 64(3):441–442, May 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000002918337.

Calvanese:2002:RRE

[1764] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. Rewriting of regular expressions and regular path queries. Journal of Computer and System Sciences, 64(3):443– 465, May 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918057.

Cosmadakis:2002:ICR

[1765] Stavros Cosmadakis. Inherent complexity of recursive queries. Journal of Computer and System Sciences, 64(3):466–495, May 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918069.

Dawson:2002:MSP

[1766] Steven Dawson, Sabrina De Capitani di Vimercati, Patrick Lincoln, and Pierangela Samarati. Maximizing sharing of protected information. Journal of Computer and System Sciences, 64(3):496–541, May 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918070.

Ganti:2002:FMD

[1767] Venkatesh Ganti, Johannes Gehrke, Raghu Ramakrishnan, and Wei-Yin Loh. A framework for measuring differences in data characteristics. Journal of Computer and System Sciences, 64(3):542–578, May 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918082.

Gottlob:2002:HDT

[1768] Georg Gottlob, Nicola Leone, and Francesco Scarcello. Hypertree decompositions and tractable queries. Journal of Computer and System Sciences, 64(3):579–627, May 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918094.

Benedikt:2002:AOC

[1769] Michael Benedikt and Leonid Libkin. Aggregate operators in constraint query languages. Journal of Computer and System Sciences, 64(3): 628–654, May 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918100.

Kanza:2002:QII

[1770] Yaron Kanza, Werner Nutt, and Yehoshua Sagiv. Querying incomplete information in semistructured data. Journal of Computer and System Sciences, 64(3): 655–693, May 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000001918112.

VandenBussche:2002:PTI

[1771] Jan Van den Bussche and Emmanuel Waller. Polymorphic type inference for the relational algebra. Journal of Computer and System Sciences, 64(3):694–718, May 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918124.

Alon:2002:TJS

[1772] Noga Alon, Phillip B. Gibbons, Yossi Matias, and Mario Szegedy. Tracking join and self-join sizes in limited storage. Journal of Computer and System Sciences, 64(3): 719–747, May 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000001918136.

Anonymous:2002:PAFc

[1773] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 64 (3):748, May 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918568.

Khanna:2002:GEF

[1774] Sanjeev Khanna. Guest Editor’s foreword. Journal of Computer and System Sciences, 64(4): 749, June 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918258.

Ambainis:2002:QLB

[1775] Andris Ambainis. Quantum lower bounds by quantum arguments. Journal of Computer and System Sciences, 64(4):750–767, June 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000291826X.

Boneh:2002:FSI

[1776] Dan Boneh. Finding smooth integers in short intervals using CRT decoding. Journal of Computer and System Sciences, 64(4):768– 784, June 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918271.

Charikar:2002:QSP

[1777] Moses Charikar, Ronald Fagin, Venkatesan Guruswami, Jon Kleinberg, Prabhakar Raghavan, and Amit Sahai. Query strategies for priced information. Journal of Computer and System Sciences, 64(4): 785–819, June 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918283.

Kempe:2002:CIP

[1778] David Kempe, Jon Kleinberg, and Amit Kumar. Connectivity and inference problems for temporal networks. Journal of Computer and System Sciences, 64(4):820– 842, June 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918295.

Maciel:2002:NPW

[1779] Alexis Maciel, Toniann Pitassi, and Alan R. Woods. A new proof of the weak pigeonhole principle. Journal of Computer and System Sciences, 64(4):843–872, June 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918301.

Barkol:2002:TLB

[1780] Omer Barkol and Yuval Rabani.

Tighter lower bounds for nearest neighbor search and related problems in the cell probe model. Journal of Computer and System Sciences, 64 (4):873–896, June 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918313.

Anonymous:2002:PAFd

[1781] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 64(4): 897, June 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000291857X.

Anonymous:2002:AIV

[1782] Anonymous. Author index for volume 64. Journal of Computer and System Sciences, 64(4): 898, June 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918581.

Cowen:2002:GEF

[1783] Lenore Cowen, Ronald Fagin, Joe Kilian, and Jon Kleinberg. Guest Editor’s foreword. Journal of Computer and System Sciences, 65(1): 1, August 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918349.

Ajtai:2002:DVN

[1784] Miklós Ajtai. Determinism versus nondeterminism for linear time RAMs with memory restrictions. Journal of Computer and System Sciences, 65(1):2–37, August 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918210.

Beame:2002:OBP

[1785] Paul Beame and Faith E. Fich. Optimal bounds for the predecessor problem and related problems. Journal of Computer and System Sciences, 65(1):38–72, August 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918222.

Li:2002:FSR

[1786] Ming Li, Bin Ma, and Lusheng Wang. Finding similar regions in many sequences. Journal of Computer and System Sciences, 65(1):

73–96, August 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918234.

Raz:2002:EAR

[1787] Ran Raz, Omer Reingold, and Salil Vadhan. Extracting all the randomness and reducing the error in Trevisan’s extractors. Journal of Computer and System Sciences, 65 (1):97–128, August 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918246.

Charikar:2002:CFA

[1788] Moses Charikar, Sudipto Guha, Éva Tardos, and David B. Shmoys. A constant-factor approximation algorithm for the k-median problem. Journal of Computer and System Sciences, 65(1):129–149, August 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918829.

King:2002:FDA

[1789] Valerie King and Garry Sagert. A fully dynamic algorithm for maintaining the transitive closure. Journal of Computer and System Sciences, 65 (1):150–167, August 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918830.

Anonymous:2002:PAFe

[1790] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 65(1): 168, August 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918817.

Stewart:2002:RBC

[1791] Charles Stewart. Reducibility between classes of port graph grammar. Journal of Computer and System Sciences, 65 (2):169–223, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918143.

Nakamura:2002:OLB

[1792] Atsuyoshi Nakamura and Naoki Abe. Online learning of binary and n-ary relations over clustered domains. Journal of Computer and System Sciences, 65 (2):224–256, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918209.

Arvind:2002:NLR

[1793] V. Arvind and Johannes Köbler. New lowness results for ZPP NP and other complexity classes. Journal of Computer and System Sciences, 65(2): 257–277, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918350.

Harju:2002:SDP

[1794] Tero Harju, Oscar Ibarra, Juhani Karhumäki, and Arto Salomaa. Some decision problems concerning semilinearity and commutation. Journal of Computer and System Sciences, 65(2): 278–294, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918362.

Willard:2002:AHM

[1795] Dan E. Willard. An algorithm for handling many relational calculus queries efficiently. Journal of Computer and System Sciences, 65(2): 295–331, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918489.

Berman:2002:CPM

[1796] Piotr Berman, Marek Karpinski, Lawrence L. Larmore, Wojciech Plandowski, and Wojciech Rytter. On the complexity of pattern matching for highly compressed twodimensional texts. Journal of Computer and System Sciences, 65(2): 332–350, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918520.

Skodinis:2002:NRE

[1797] K. Skodinis and E. Wanke. Node replacements in embedding normal form. Journal of Computer and System Sciences, 65(2):351–376, September 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918519.

Honkala:2002:EPD

[1798] Juha Honkala. The equivalence problem for DF0L languages and power series. Journal of Computer and System Sciences, 65(2): 377–392, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918507.

Pighizzini:2002:UCF

[1799] Giovanni Pighizzini, Jeffrey Shallit, and Ming wei Wang. Unary contextfree grammars and pushdown automata, descriptional complexity and auxiliary space lower bounds. Journal of Computer and System Sciences, 65 (2):393–414, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918556.

Mason:2002:GEC

[1800] Llew Mason, Peter L. Bartlett, and Mostefa Golea. Generalization error of combined classifiers. Journal of Computer and System Sciences, 65(2): 415–438, September 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002918544.

Anonymous:2002:PAFf

[1801] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 65(2): 439, September 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002918842.

Anonymous:2002:AS

[1802] Anonymous. Advertisement #sciencejobs N. Journal of Computer and System Sciences, 65(3): ii, November 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000405.

Li:2002:F

[1803] Ming Li, Pavel Pevzner, and Ron Shamir. Foreword. Journal of Computer and System Sciences, 65 (3):441, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000028.

El-Mabrouk:2002:RAG

[1804] Nadia El-Mabrouk. Reconstructing an ancestral genome using minimum segments duplications and reversals. Journal of Computer and System Sciences, 65(3):442– 464, November 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000200003X.

Lin:2002:LCS

[1805] Guohui Lin, Zhi-Zhong Chen, Tao Jiang, and Jianjun Wen. The longest common subsequence problem for sequences with nested arc annotations. Journal of Computer and System Sciences, 65(3): 465–480, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000041.

Lagergren:2002:CPR

[1806] J. Lagergren. Combining polynomial running time and fast convergence for the disk-covering method. Journal of Computer and System Sciences, 65(3): 481–493, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000053.

Jaitly:2002:MRH

[1807] Deep Jaitly, Paul Kearney, Guohui Lin, and Bin Ma. Methods for reconstructing the history of tandem repeats and their application to the human genome. Journal of Computer and System Sciences, 65(3): 494–507, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000065.

Moret:2002:STA

[1808] Bernard M. E. Moret, Jijun Tang, Li-San Wang, and Tandy Warnow. Steps toward accurate reconstructions of phylogenies from gene-order data. Journal of Computer and System Sciences, 65 (3):508–525, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000077.

Pandurangan:2002:RMP

[1809] Gopal Pandurangan and H. Ramesh.

The restriction mapping problem revisited. Journal of Computer and System Sciences, 65(3):526– 544, November 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000089.

Lyngso:2002:CSP

[1810] Rune B. Lyngsø and Christian N. S.

Pedersen. The consensus string problem and the complexity of comparing hidden Markov models. Journal of Computer and System Sciences, 65 (3):545–569, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000090.

Lin:2002:EAL

[1811] Yaw-Ling Lin, Tao Jiang, and Kun-Mao Chao. Efficient algorithms for locating the lengthconstrained heaviest segments with applications to biomolecular sequence analysis. Journal of Computer and System Sciences, 65(3):570– 586, November 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000107.

Tesler:2002:EAM

[1812] Glenn Tesler. Efficient algorithms for multichromosomal genome rearrangements. Journal of Computer and System Sciences, 65(3): 587–609, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002000119.

Anonymous:2002:PAa

[1813] Anonymous. Papers to appear. Journal of Computer and System Sciences, 65(3):610, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002000417.

Anonymous:2002:CAa

[1814] Anonymous. C3 — advertisement #09111. Journal of Computer and System Sciences, 65(3): ibc, November 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002000430.

Anonymous:2002:CTa

[1815] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 65(3):obc, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002000442.

Anonymous:2002:EBa

[1816] Anonymous. Editorial board. Journal of Computer and System Sciences, 65(3):ifc, November 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000387.

Anonymous:2002:A

[1817] Anonymous. Advertisement. Journal of Computer and System Sciences, 65(4):ii, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000200048X.

Anonymous:2002:VT

[1818] Anonymous. Volume TOC. Journal of Computer and System Sciences, 65 (4):III–V, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000533.

Sudan:2002:F

[1819] Madhu Sudan. Foreword. Journal of Computer and System Sciences, 65 (4):611, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000181.

Forster:2002:LLB

[1820] Jürgen Forster. A linear lower bound on the unbounded error probabilistic communication complexity. Journal of Computer and System Sciences, 65 (4):612–625, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000193.

Atserias:2002:MSN

[1821] Albert Atserias, Nicola Galesi, and Pavel Pudlák. Monotone simulations of non-monotone proofs. Journal of Computer and System Sciences, 65(4): 626–638, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000200020X.

Shpilka:2002:APS

[1822] Amir Shpilka. Affine projections of symmetric polynomials. Journal of Computer and System Sciences, 65(4): 639–659, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000211.

Mossel:2002:CAV

[1823] Elchanan Mossel and Christopher Umans. On the complexity of approximating the VC dimension. Journal of Computer and System Sciences, 65 (4):660–671, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000223.

Impagliazzo:2002:SEW

[1824] Russell Impagliazzo, Valentine Kabanets, and Avi Wigderson. In search of an easy witness: exponential time vs. probabilistic polynomial time. Journal of Computer and System Sciences, 65 (4):672–694, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000247.

Hesse:2002:UCD

[1825] William Hesse, Eric Allender, and David A. Mix Barrington. Uniform constant-depth threshold circuits for division and iterated multiplication. Journal of Computer and System Sciences, 65(4):695– 716, December 2002. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000259. See

corrigendum [2926].

Koucky:2002:UTS

[1826] Michal Koucký. Universal traversal sequences with backtracking. Journal of Computer and System Sciences, 65 (4):717–726, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000235.

Anonymous:2002:PAb

[1827] Anonymous. Papers to appear. Journal of Computer and System Sciences, 65(4):727, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000491.

Anonymous:2002:VAI

[1828] Anonymous. Volume author index. Journal of Computer and System Sciences, 65(4):728, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000002000508.

Anonymous:2002:CAb

[1829] Anonymous. C3 — advertisement.

Journal of Computer and System Sciences, 65(4):ibc, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000002000557.

Anonymous:2002:CTb

[1830] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 65(4):obc, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002000569.

Anonymous:2002:EBb

[1831] Anonymous. Editorial board. Journal of Computer and System Sciences, 65(4):ifc, December 2002. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000002000466.

Anonymous:2003:APP

[1832] Anonymous. Advertisement #praise 5x8 els.pdf. Journal of Computer and System Sciences, 66(1):ii, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000187.

Anonymous:2003:PN

[1833] Anonymous. Publisher’s note. Journal of Computer and System Sciences, 66(1):iii–iv, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000229.

Gottlob:2003:GEF

[1834] Georg Gottlob. Guest Editor’s foreword. Journal of Computer and System Sciences, 66(1):1, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000002000260.

Lynch:2003:AAA

[1835] James F. Lynch. Analysis and application of adaptive sampling. Journal of Computer and System Sciences, 66(1):2–19, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000272.

Millstein:2003:QCD

[1836] Todd Millstein, Alon Halevy, and Marc Friedman. Query containment for data integration systems. Journal of Computer and System Sciences, 66 (1):20–39, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000284.

Spielmann:2003:VRT

[1837] Marc Spielmann. Verification of relational transducers for electronic commerce. Journal of Computer and System Sciences, 66(1):40–65, February 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000296.

Milo:2003:TXT

[1838] Tova Milo, Dan Suciu, and Victor Vianu. Typechecking for XML transformers. Journal of Computer and System Sciences, 66(1):66– 97, February 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000302.

Chen:2003:GSS

[1839] Zhiyuan Chen, Flip Korn, Nick Koudas, and S. Muthukrishnan. Generalized substring selectivity estimation. Journal of Computer and System Sciences, 66(1):98–132, February 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000314.

Grumbach:2003:CMA

[1840] Stéphane Grumbach and Leonardo Tininini. On the content of materialized aggregate views. Journal of Computer and System Sciences, 66 (1):133–168, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000338.

Benedikt:2003:RCQ

[1841] Michael Benedikt, Martin Grohe, Leonid Libkin, and Luc Segoufin. Reachability and connectivity queries in constraint databases. Journal of Computer and System Sciences, 66 (1):169–206, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000200034X.

Agarwal:2003:IMP

[1842] Pankaj K. Agarwal, Lars Arge, and Jeff Erickson. Indexing moving points. Journal of Computer and System Sciences, 66(1):207– 243, February 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000351.

Kleinberg:2003:ABA

[1843] Jon Kleinberg, Christos Papadimitriou, and Prabhakar Raghavan. Auditing Boolean attributes. Journal of Computer and System Sciences, 66 (1):244–253, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000363.

Fan:2003:ICX

[1844] Wenfei Fan and Jérôme Siméon. Integrity constraints for XML. Journal of Computer and System Sciences, 66 (1):254–291, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000002000326.

Anonymous:2003:PAa

[1845] Anonymous. Papers to appear. Journal of Computer and System Sciences, 66(1):292, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000199.

Anonymous:2003:AH

[1846] Anonymous. Advertisement #HA08111M. Journal of Computer and System Sciences, 66(1):ibc, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000003000205.

Anonymous:2003:EBa

[1847] Anonymous. Ed. Board. Journal of Computer and System Sciences, 66(1):ifc, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000163.

Anonymous:2003:Ta

[1848] Anonymous.  TOC.  Journal of

Computer and System Sciences, 66 (1):obc, February 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000217.

Anonymous:2003:AHFa

[1849] Anonymous. Advertisement #HA05062O

                                        (FTP file HA05062ABCNOP.pdf.sea).

Journal of Computer and System Sciences, 66(2):ii, March 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000540.

Carchiolo:2003:TML

[1850] Vincenza Carchiolo, Michele Malgeri, and Giuseppe Mangioni. TTL: a modular language for hardware/software systems design. Journal of Computer and System Sciences, 66(2): 293–315, March 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000023.

Tini:2003:ASS

[1851] Simone Tini. An axiomatic semantics for the synchronous language Gentzen. Journal of Computer and System Sciences, 66(2): 316–348, March 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000035.

Kellerer:2003:EFP

[1852] Hans Kellerer, Renata Mansini, Ulrich Pferschy, and Maria Grazia Speranza. An efficient fully polynomial approximation scheme for the subset-sum problem. Journal of Computer and System Sciences, 66 (2):349–370, March 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000060.

Kiwi:2003:ATE

[1853] Marcos Kiwi, Frédéric Magniez, and Miklos Santha. Approximate testing with error relative to input size. Journal of Computer and System Sciences, 66(2):371–392, March 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000047.

Reynolds:2003:CTL

[1854] M. Reynolds. The complexity of the temporal logic with “until” over general linear time. Journal of Computer and System Sciences, 66(2): 393–426, March 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000059.

Anonymous:2003:PAb

[1855] Anonymous. Papers to appear. Journal of Computer and System Sciences, 66(2):427, March 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000552.

Anonymous:2003:CAB

[1856] Anonymous. C3 — advertisement #better 5x8 els (FTP file #better 5x8 els.pdf) in generic folder. Journal of Computer and System Sciences, 66(2):ibc, March 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000576.

Anonymous:2003:CCE

[1857] Anonymous. C1–C2 Ed. Board. Journal of Computer and System Sciences, 66(2):ifc, March 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000527.

Anonymous:2003:CTa

[1858] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 66(2):obc, March 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000588.

Anonymous:2003:B

[1859] Anonymous. Blank. Journal of Computer and System Sciences, 66 (3):III, May 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000886.

Anonymous:2003:AHFb

[1860] Anonymous. Advertisement #HA03022M (FTP file HA03022LMS.sea). Journal of Computer and System Sciences, 66(3):IV, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000928.

Anonymous:2003:AHFc

[1861] Anonymous. Advertisement #HA05062O (FTP file HA05062ABCNOP.pdf.sea). Journal of Computer and System Sciences, 66(3):V, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300093X.

Anonymous:2003:AAFa

[1862] Anonymous. Advertisement #AM01053M (FTP file HA01053LMS.zip). Journal of Computer and System Sciences, 66(3):VI, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000916.

Kobayashi:2003:QMP

[1863] Hirotada Kobayashi and Keiji Matsumoto. Quantum multi-prover interactive proof systems with limited prior entanglement. Journal of Computer and System Sciences, 66 (3):429–450, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000357.

Geffert:2003:TBR

[1864] Viliam Geffert. Translation of binary regular expressions into nondeterministic ϵ-free automata with O(n log n) transitions. Journal of Computer and System Sciences, 66 (3):451–472, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000369.

Sauerhoff:2003:GVV

[1865] Martin Sauerhoff. Guess-and-verify versus unrestricted nondeterminism for OBDDs and one-way Turing machines. Journal of Computer and System Sciences, 66(3):473– 495, May 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000370.

Ben-David:2003:DAM

[1866] Shai Ben-David, Nadav Eiron, and Philip M. Long. On the difficulty of approximately maximizing agreements. Journal of Computer and System Sciences, 66(3): 496–514, May 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000382.

Gaujal:2003:BTF

[1867] Bruno Gaujal, Stefan Haar, and Jean Mairesse. Blocking a transition in a free choice net and what it tells about its throughput. Journal of Computer and System Sciences, 66 (3):515–548, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000394.

Jenner:2003:CRG

[1868] Birgit Jenner, Johannes Köbler, Pierre McKenzie, and Jacobo Torán. Completeness results for graph isomorphism. Journal of Computer and System Sciences, 66(3):549– 566, May 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000424. See

corrigendum [2185].

Henzinger:2003:SMU

[1869] Monika R. Henzinger and Stefano Leonardi. Scheduling multicasts on unit-capacity trees and meshes. Journal of Computer and System Sciences, 66(3):567–611, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000436.

Anonymous:2003:PAc

[1870] Anonymous. Papers to appear. Journal of Computer and System Sciences, 66(3):612, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000862.

Anonymous:2003:CAHa

[1871] Anonymous. C3 — advertisement #HA06043M (FTP file HA06043LMSU.zip). Journal of Computer and System Sciences, 66(3):ibc, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000898.

Anonymous:2003:CEBa

[1872] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 66(3):ifc, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000837.

Anonymous:2003:CTb

[1873] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 66(3):obc, May 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000904.

Anonymous:2003:VTa

[1874] Anonymous. Volume TOC. Journal of Computer and System Sciences, 66(4):III–V, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001028.

Anonymous:2003:AHFd

[1875] Anonymous. Advertisement #HA03022M (FTP file HA03022LMS.sea). Journal of Computer and System Sciences, 66(4):IX, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001089.

Anonymous:2003:AHFe

[1876] Anonymous. Advertisement #HA05062N

(FTP file HA05062ABCNOP.pdf.sea).

Journal of Computer and System Sciences, 66(4):X, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001090.

Anonymous:2003:AAFb

[1877] Anonymous. Advertisement #AM01053M (FTP file AM01053LMS.zip). Journal of Computer and System Sciences, 66(4):XI, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001041.

Buneman:2003:GEF

[1878] Peter Buneman. Guest Editor’s foreword. Journal of Computer and System Sciences, 66(4):613, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000003000242.

Fagin:2003:OAA

[1879] Ronald Fagin, Amnon Lotem, and Moni Naor. Optimal aggregation algorithms for middleware. Journal of Computer and System Sciences, 66 (4):614–656, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000266.

Larsen:2003:RMW

[1880] Kim S. Larsen. Relaxed multi-way trees with group updates. Journal of Computer and System Sciences, 66(4):657–670, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000278.

Achlioptas:2003:DFR

[1881] Dimitris Achlioptas. Databasefriendly random projections: Johnson– Lindenstrauss with binary coins. Journal of Computer and System Sciences, 66(4):671–687, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000254.

Alon:2003:XDV

[1882] Noga Alon, Tova Milo, Frank Neven, Dan Suciu, and Victor Vianu. XML with data values: typechecking revisited. Journal of Computer and System Sciences, 66(4): 688–727, June 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000321.

Dalvi:2003:PMQ

[1883] Nilesh N. Dalvi, Sumit K. Sanghai, Prasan Roy, and S. Sudarshan. Pipelining in multi-query optimization. Journal of Computer and System Sciences, 66(4):728–762, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300031X.

Ferragina:2003:TDS

[1884] Paolo Ferragina, Nick Koudas, S. Muthukrishnan, and Divesh Srivastava. Twodimensional substring indexing. Journal of Computer and System Sciences, 66(4):763–774, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300028X.

Gottlob:2003:RMG

[1885] Georg Gottlob, Nicola Leone, and Francesco Scarcello. Robbers, marshals, and guards: game theoretic and logical characterizations of hypertree width. Journal of Computer and System Sciences, 66(4): 775–808, June 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000308.

Rajaraman:2003:QWU

[1886] Anand Rajaraman and Jeffrey D. Ullman. Querying websites using compact skeletons. Journal of Computer and System Sciences, 66(4): 809–851, June 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000291.

Anonymous:2003:PAd

[1887] Anonymous. Papers to appear. Journal of Computer and System Sciences, 66(4):852, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000989.

Anonymous:2003:VAIa

[1888] Anonymous. Volume author index. Journal of Computer and System Sciences, 66(4):853, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000990.

Anonymous:2003:CAHb

[1889] Anonymous. C3 — advertisement #HA06043M (FTP file HA06043LMSU.zip). Journal of Computer and System Sciences, 66(4):ibc, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001053.

Anonymous:2003:CEBb

[1890] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 66(4):ifc, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000953.

Anonymous:2003:CTc

[1891] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 66(4):obc, June 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001065.

Anonymous:2003:AHFf

[1892] Anonymous. Ad #HA06043M (FTP file HA06043LMSU.zip). Journal of Computer and System Sciences, 67 (1):ii, August 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001144.

Anonymous:2003:AEF

[1893] Anonymous. Ad #econdirectM (FTP file econdirectLMS.zip). Journal of Computer and System Sciences, 67 (1):III, August 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300117X.

Rim:2003:SRN

[1894] Hwakyung Rim, Ju wook Jang, and Sungchun Kim. A simple reduction of non-uniformity in dynamic load balancing of quantized loads on hypercube multiprocessors and hiding balancing overheads. Journal of Computer and System Sciences, 67(1):1–25, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000333.

Dolev:2003:SLL

[1895] Shlomi Dolev and Sergio Rajsbaum. Stability of long-lived consensus. Journal of Computer and System Sciences, 67(1):26–45, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000631.

Damaschke:2003:PAE

[1896] Peter Damaschke. On parallel attribute-efficient learning. Journal of Computer and System Sciences, 67(1):46–62, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000473.

Gasarch:2003:CTP

[1897] William Gasarch, Evan Golub, and Clyde Kruskal. Constant time parallel sorting: an empirical view. Journal of Computer and System Sciences, 67(1):63–91, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000400.

Cohen:2003:CCM

[1898] Edith Cohen, Haim Kaplan, and Uri Zwick. Connection caching: model and algorithms. Journal of Computer and System Sciences, 67(1): 92–126, August 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000412.

Meinke:2003:PTH

[1899] K. Meinke. Proof theory of higherorder equations: conservativity, normal forms and term rewriting. Journal of Computer and System Sciences, 67(1):127–173, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000485.

Allouche:2003:RPC

[1900] J.-P. Allouche and G. Skordev. Remarks on permutive cellular automata. Journal of Computer and System Sciences, 67(1):174– 182, August 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000497.

Chrobak:2003:PSO

[1901] Marek Chrobak, Leah Epstein, John Noga, Jiří Sgall, Robvan Stee, Tomáš Tichýd, and Nodari Vakhania. Preemptive scheduling in overloaded systems. Journal of Computer and System Sciences, 67(1): 183–197, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000709.

Ranade:2003:SLC

[1902] Abhiram G. Ranade. Scheduling loosely connected task graphs. Journal of Computer and System Sciences, 67(1):198–208, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000710.

Anonymous:2003:PAe

[1903] Anonymous. Papers to appear. Journal of Computer and System Sciences, 67(1):209, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001156.

Anonymous:2003:CAA

[1904] Anonymous. C3 — ad #AM01053M (FTP file AM01053LMS.zip). Journal of Computer and System Sciences, 67(1):ibc, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001181.

Anonymous:2003:CEBc

[1905] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 67(1):ifc, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001120.

Anonymous:2003:CTd

[1906] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 67(1):obc, August 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001193.

Anonymous:2003:AAFc

[1907] Anonymous. Advertisement: AS007M (FTP file AS007M.pdf). Journal of Computer and System Sciences, 67 (2):ii, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001247.

Anonymous:2003:AAFd

[1908] Anonymous. Advertisement: AS004M (FTP file AS004M.pdf). Journal of Computer and System Sciences, 67 (2):III, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001272.

Reif:2003:GEF

[1909] John H. Reif. Guest Editor’s foreword. Journal of Computer and System Sciences, 67(2):211, September 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000072.

Alon:2003:RSA

[1910] Noga Alon, W. Fernandez de la Vega, Ravi Kannan, and Marek Karpinski. Random sampling and approximation of MAX-CSPs. Journal of Computer and System Sciences, 67(2): 212–243, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000084.

Basu:2003:CBN

[1911] Saugata Basu. Computing the Betti numbers of arrangements via spectral sequences. Journal of Computer and System Sciences, 67(2): 244–262, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000096.

Chien:2003:CAA

[1912] Steve Chien, Lars Rasmussen, and Alistair Sinclair. Clifford algebras and approximating the permanent. Journal of Computer and System Sciences, 67 (2):263–290, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000102.

Cryan:2003:PTA

[1913] Mary Cryan and Martin Dyer. A polynomial-time algorithm to approximately count contingency tables when the number of rows is constant. Journal of Computer and System Sciences, 67 (2):291–310, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300014X.

Deng:2003:CPE

[1914] Xiaotie Deng, Christos Papadimitriou, and Shmuel Safra. On the complexity of price equilibria. Journal of Computer and System Sciences, 67(2): 311–324, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000114.

Arora:2003:FAC

[1915] Sanjeev Arora and Subhash Khot. Fitting algebraic curves to noisy data. Journal of Computer and System Sciences, 67(2):325–340, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000003000126.

Roughgarden:2003:PAI

[1916] Tim Roughgarden. The price of anarchy is independent of the network topology. Journal of Computer and System Sciences, 67(2): 341–364, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000448.

Schaefer:2003:RSG

[1917] Marcus Schaefer, Eric Sedgwick, and Daniel Stefankovic. Recognizing string graphs in NP. Journal of Computer and System Sciences, 67(2): 365–380, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300045X.

Brodal:2003:OFS

[1918] Gerth Stølting Brodal, George Lagogiannis, Christos Makris, Athanasios Tsakalidis, and Kostas Tsichlas. Optimal finger search trees in the pointer machine. Journal of Computer and System Sciences, 67(2): 381–418, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000138.

Umans:2003:PRG

[1919] Christopher Umans. Pseudo-random generators for all hardnesses. Journal of Computer and System Sciences, 67 (2):419–440, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000461.

Achlioptas:2003:AAG

[1920] Dimitris Achlioptas and Cristopher Moore. Almost all graphs with average degree 4 are 3-colorable. Journal of Computer and System Sciences, 67 (2):441–471, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000300120X.

Anonymous:2003:PAf

[1921] Anonymous. Papers to appear. Journal of Computer and System Sciences, 67(2):472, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001259.

Anonymous:2003:CAF

[1922] Anonymous. C3 — AS008M (FTP file AS008M.pdf). Journal of Computer and System Sciences, 67(2): ibc, September 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001284.

Anonymous:2003:CTe

[1923] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 67 (2):obc, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001296.

Anonymous:2003:EBb

[1924] Anonymous. Ed. Board. Journal of Computer and System Sciences, 67(2):ifc, September 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001223.

Anonymous:2003:SFF

[1925] Anonymous. sdmathM (FTP file sdmathLMS.zip). Journal of Computer and System Sciences, 67(3): ii, November 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001582.

Guruswami:2003:NOH

[1926] Venkatesan Guruswami, Sanjeev Khanna, Rajmohan Rajaraman, Bruce Shepherd, and Mihalis Yannakakis. Nearoptimal hardness results and approximation algorithms for edge-disjoint paths and related problems. Journal of Computer and System Sciences, 67 (3):473–496, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000667.

Ly:2003:AGD

[1927] Olivier Ly. Automatic graphs and d0lsequences of finite graphs. Journal of Computer and System Sciences, 67 (3):497–545, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000655.

Jain:2003:ICL

[1928] Sanjay Jain and Efim Kinber. Intrinsic complexity of learning geometrical concepts from positive data. Journal of Computer and System Sciences, 67 (3):546–607, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000679.

Homan:2003:OWP

[1929] Christopher M. Homan and Mayur Thakur. One-way permutations and self-witnessing languages. Journal of Computer and System Sciences, 67(3): 608–622, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000680.

Li:2003:CVW

[1930] Chunguang Li, Xiaofeng Liao, and Juebang Yu. Complex-valued wavelet network. Journal of Computer and System Sciences, 67(3):623– 632, November 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000692.

Srinivasan:2003:ACR

[1931] Aravind Srinivasan. On the approximability of clique and related maximization problems. Journal of Computer and System Sciences, 67(3): 633–651, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001107.

Anonymous:2003:PAg

[1932] Anonymous. Papers to appear. Journal of Computer and System Sciences, 67(3):652, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001594.

Anonymous:2003:EBc

[1933] Anonymous. Ed. Board. Journal of Computer and System Sciences, 67(3):ifc, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001569.

Anonymous:2003:HFF

[1934] Anonymous. HA11093M (FTP file HA11093LMS.zip). Journal of Computer and System Sciences, 67(3): ibc, November 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001612.

Anonymous:2003:Tb

[1935] Anonymous.  TOC.  Journal of

Computer and System Sciences, 67 (3):obc, November 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001624.

Anonymous:2003:A

[1936] Anonymous. Advertisement. Journal of Computer and System Sciences, 67(4):ii, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001661.

Anonymous:2003:VTb

[1937] Anonymous. Volume TOC. Journal of Computer and System Sciences, 67 (4):III–V, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001715.

Chen:2003:FGE

[1938] Jianer Chen and Michael R. Fellows. Foreword from the guest editors. Journal of Computer and System Sciences, 67(4):653, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001545.

Cesati:2003:TWP

[1939] Marco Cesati. The Turing way to parameterized complexity. Journal of Computer and System Sciences, 67(4): 654–685, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000734.

Cotta:2003:FSP

[1940] Carlos Cotta and Pablo Moscato.

The k-feature set problem is W [2]- complete. Journal of Computer and System Sciences, 67(4):686– 690, December 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000813.

Cheetham:2003:SLF

[1941] James Cheetham, Frank Dehne, Andrew Rau-Chaplin, Ulrike Stege, and Peter J. Taillon. Solving large FPT problems on coarse-grained parallel machines. Journal of Computer and System Sciences, 67(4): 691–706, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000758.

Anand:2003:CCK

[1942] R. Sai Anand, Thomas Erlebach, Alexander Hall, and Stamatis Stefanakos. Call control with k rejections. Journal of Computer and System Sciences, 67(4):707– 722, December 2003. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000300076X.

Gramm:2003:FPA

[1943] Jens Gramm and Rolf Niedermeier. A fixed-parameter algorithm for minimum quartet inconsistency. Journal of Computer and System Sciences, 67 (4):723–741, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000771.

Makowsky:2003:PCK

[1944] J. A. Makowsky and J. P. Mariño. The parametrized complexity of knot polynomials. Journal of Computer and System Sciences, 67(4): 742–756, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000801.

Pietrzak:2003:PCF

[1945] Krzysztof Pietrzak. On the parameterized complexity of the fixed alphabet shortest common supersequence and longest common subsequence problems. Journal of Computer and System Sciences, 67(4): 757–771, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003000783.

Heusch:2003:IFP

[1946] Peter Heusch, Stefan Porschen, and Ewald Speckenmeyer. Improving a fixed parameter tractability time bound for the shadow problem. Journal of Computer and System Sciences, 67 (4):772–788, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000795.

Cai:2003:ESP

[1947] Liming Cai and David Juedes. On the existence of subexponential parameterized algorithms. Journal of Computer and System Sciences, 67(4): 789–807, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000746.

Alber:2003:GSP

[1948] Jochen Alber, Henning Fernau, and Rolf Niedermeier. Graph separators: a parameterized view. Journal of Computer and System Sciences, 67(4): 808–832, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000722.

Chen:2003:CMV

[1949] Jianer Chen and Iyad A. Kanj. Constrained minimum vertex cover in bipartite graphs: complexity and parameterized algorithms. Journal of Computer and System Sciences, 67(4): 833–847, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001533.

Anonymous:2003:PAh

[1950] Anonymous. Papers to appear. Journal of Computer and System Sciences, 67(4):848, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001673.

Anonymous:2003:VAIb

[1951] Anonymous. Volume author index. Journal of Computer and System Sciences, 67(4):849, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000003001685.

Anonymous:2003:CA

[1952] Anonymous. C3 — advertisement. Journal of Computer and System Sciences, 67(4):ibc, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000003001739.

Anonymous:2003:CTf

[1953] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 67(4):obc, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001740.

Anonymous:2003:EBd

[1954] Anonymous. Ed. Board. Journal of Computer and System Sciences, 67(4):ifc, December 2003. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001648.

Anonymous:2004:Aa

[1955] Anonymous. Advertisement. Journal of Computer and System Sciences, 68(1):ii, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001909.

Anonymous:2004:Ab

[1956] Anonymous. Advertisement. Journal of Computer and System Sciences, 68(1):III, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001934.

Anonymous:2004:Ac

[1957] Anonymous. Advertisement. Journal of Computer and System Sciences, 68(1):IV, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300196X.

Anonymous:2004:Ad

[1958] Anonymous. Advertisement. Journal of Computer and System Sciences, 68(1):V, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001971.

Mateescu:2004:SHP

[1959] Alexandru Mateescu, Arto Salomaa, and Sheng Yu. Subword histories and Parikh matrices. Journal of Computer and System Sciences, 68 (1):1–21, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001478.

Choi:2004:SAE

[1960] Kwok Pui Choi and Louxin Zhang. Sensitivity analysis and efficient method for identifying optimal spaced seeds. Journal of Computer and System Sciences, 68(1):22–40, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000003001430.

Le:2004:AVD

[1961] Ngȯc-Minh Lê. Abstract Voronoi diagram in 3-space. Journal of Computer and System Sciences, 68(1):41– 79, February 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001302.

Becchetti:2004:ASM

[1962] Luca Becchetti, Stefano Leonardi, and S. Muthukrishnan. Average stretch without migration. Journal of Computer and System Sciences, 68(1):80– 95, February 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300134X.

Downey:2004:RR

[1963] Rod G. Downey, Denis R. Hirschfeldt, and Geoff LaForte. Randomness and reducibility. Journal of Computer and System Sciences, 68(1): 96–114, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001314.

Li:2004:SMD

[1964] Deng-Feng Li. Some measures of dissimilarity in intuitionistic fuzzy structures. Journal of Computer and System Sciences, 68(1):115– 122, February 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001338.

Anceaume:2004:NSC

[1965] E. Anceaume, A. Fernández, A. Mostefaoui, G. Neiger, and M. Raynal. A necessary and sufficient condition for transforming limited accuracy failure detectors. Journal of Computer and System Sciences, 68(1): 123–133, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300148X.

Merkle:2004:TL

[1966] Wolfgang Merkle and Frank Stephan.

Trees and learning. Journal of Computer and System Sciences, 68(1): 134–156, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001491.

Anderson:2004:MPE

[1967] James H. Anderson and Anand Srinivasan. Mixed Pfair/ERfair scheduling of asynchronous periodic tasks. Journal of Computer and System Sciences, 68 (1):157–204, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001508.

Bshouty:2004:MEP

[1968] Nader H. Bshouty, Jeffrey C. Jackson, and Christino Tamon. More efficient PAC-learning of DNF with membership queries under the uniform distribution. Journal of Computer and System Sciences, 68(1): 205–234, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001764.

Anonymous:2004:PAa

[1969] Anonymous. Papers to appear. Journal of Computer and System Sciences, 68(1):235, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001910.

Anonymous:2004:CA

[1970] Anonymous. C3 — advertisement. Journal of Computer and System Sciences, 68(1):ibc, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000003001946.

Anonymous:2004:CEBa

[1971] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 68(1):ifc, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001880.

Anonymous:2004:CTa

[1972] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 68(1):obc, February 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001958.

Kannan:2004:GEF

[1973] Sampath Kannan and Mihalis Yannakakis. Guest editors’ foreword. Journal of Computer and System Sciences, 68(2):237, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001351.

Achlioptas:2004:STP

[1974] Dimitris Achlioptas, Paul Beame, and Michael Molloy. A sharp threshold in proof complexity yields lower bounds for satisfiability search. Journal of Computer and System Sciences, 68 (2):238–268, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001466.

Chazelle:2004:LBI

[1975] Bernard Chazelle and Ding Liu.

Lower bounds for intersection searching and fractional cascading in higher dimension. Journal of Computer and System Sciences, 68(2):269– 284, March 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001405.

Grohe:2004:CCN

[1976] Martin Grohe. Computing crossing numbers in quadratic time. Journal of Computer and System Sciences, 68 (2):285–302, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001387.

Klivans:2004:LDT

[1977] Adam R. Klivans and Rocco A. Servedio. Learning DNF in time 2^{O(n^{1/3})}. Journal of Computer and System Sciences, 68(2):303– 318, March 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001363.

Schaefer:2004:DSG

[1978] Marcus Schaefer and Daniel Štefankovič. Decidability of string graphs. Journal of Computer and System Sciences, 68(2):319–334, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001375.

Dunagan:2004:OOR

[1979] John Dunagan and Santosh Vempala. Optimal outlier removal in high-dimensional spaces. Journal of Computer and System Sciences, 68 (2):335–373, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001442.

deAlfaro:2004:QSO

[1980] Luca de Alfaro and Rupak Majumdar. Quantitative solution of omegaregular games. Journal of Computer and System Sciences, 68(2): 374–397, March 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001429.

Ambainis:2004:NPL

[1981] Andris Ambainis. A new protocol and lower bounds for quantum coin flipping. Journal of Computer and System Sciences, 68(2): 398–416, March 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001417.

Charikar:2004:CMS

[1982] Moses Charikar and Rina Panigrahy.

Clustering to minimize the sum of cluster diameters. Journal of Computer and System Sciences, 68(2): 417–441, March 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001399.

Goemans:2004:AAM

[1983] Michel X. Goemans and David P. Williamson. Approximation algorithms for Max-3-Cut and other problems via complex semidefinite programming. Journal of Computer and System Sciences, 68(2): 442–470, March 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000003001454.

Anonymous:2004:PAb

[1984] Anonymous. Papers to appear. Journal of Computer and System Sciences, 68(2):471, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000121.

Anonymous:2004:CEBb

[1985] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 68(2):ifc, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000091.

Anonymous:2004:CTb

[1986] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 68(2):obc, March 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000169.

Vikas:2004:CCC

[1987] Narayan Vikas. Computational complexity of compaction to irreflexive cycles. Journal of Computer and System Sciences, 68(3): 473–496, May 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000345.

Palacz:2004:AHG

[1988] Wojciech Palacz. Algebraic hierarchical graph transformation. Journal of Computer and System Sciences, 68(3):497–520, May 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003000643.

Kuusela:2004:LSI

[1989] P. Kuusela and D. Ocone. Learning with side information: PAC learning bounds. Journal of Computer and System Sciences, 68(3): 521–545, May 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001326.

Grandjean:2004:GPC

[1990] Etienne Grandjean and Frédéric Olive.

Graph properties checkable in linear time in the number of vertices. Journal of Computer and System Sciences, 68(3):546–597, May 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001521.

Sun:2004:DTS

[1991] Hung-Min Sun, Bing-Chang Chen, and Her-Tyan Yeh. On the design of time-stamped signatures. Journal of Computer and System Sciences, 68(3):598–610, May 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001776.

Drewes:2004:BSG

[1992] Frank Drewes and Joost Engelfriet. Branching synchronization grammars with nested tables. Journal of Computer and System Sciences, 68(3): 611–656, May 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001752.

Homan:2004:TLB

[1993] Christopher M. Homan. Tight lower bounds on the ambiguity of strong, total, associative, oneway functions. Journal of Computer and System Sciences, 68(3): 657–674, May 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001788.

Duris:2004:PNV

[1994] Pavol Duris, Juraj Hromkovic, and Katsushi Inoue. On the power of nondeterminism and Las Vegas randomization for two-dimensional finite automata. Journal of Computer and System Sciences, 68(3): 675–699, May 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001867.

Anonymous:2004:PAc

[1995] Anonymous. Papers to appear. Journal of Computer and System Sciences, 68(3):700, May 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000261.

Anonymous:2004:CEBc

[1996] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 68(3):ifc, May 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000236.

Anonymous:2004:CTc

[1997] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 68(3):obc, May 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000303.

Anonymous:2004:VTa

[1998] Anonymous. Volume TOC. Journal of Computer and System Sciences, 68(4):III–V, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000400056X.

Cohen:2004:GEF

[1999] Edith Cohen and Venkatesan Guruswami. Guest editors’ foreword. Journal of Computer and System Sciences, 68(4):701, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300179X.

Bar-Yossef:2004:ISA

[2000] Ziv Bar-Yossef, T. S. Jayram, Ravi Kumar, and D. Sivakumar. An information statistics approach to data stream and communication complexity. Journal of Computer and System Sciences, 68(4):702–732, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001855.

Feldman:2004:DTL

[2001] Jon Feldman and David R. Karger. Decoding turbo-like codes via linear programming. Journal of Computer and System Sciences, 68(4): 733–752, June 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001843.

Fischer:2004:TJ

[2002] Eldar Fischer, Guy Kindler, Dana Ron, Shmuel Safra, and Alex Samorodnitsky. Testing juntas. Journal of Computer and System Sciences, 68 (4):753–787, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001831.

Franceschini:2004:IBT

[2003] Gianni Franceschini, Roberto Grossi, J. Ian Munro, and Linda Pagli. Implicit B-trees: a new data structure for the dictionary problem. Journal of Computer and System Sciences, 68(4):788–807, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000300182X.

Klivans:2004:LIT

[2004] Adam R. Klivans, Ryan O’Donnell, and Rocco A. Servedio. Learning intersections and thresholds of halfspaces. Journal of Computer and System Sciences, 68(4):808– 840, June 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001818.

Vempala:2004:SAL

[2005] Santosh Vempala and Grant Wang.

A spectral algorithm for learning mixture models. Journal of Computer and System Sciences, 68(4): 841–860, June 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000003001806.

Anonymous:2004:PAd

[2006] Anonymous. Papers to appear. Journal of Computer and System Sciences, 68(4):861, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000522.

Anonymous:2004:VAIa

[2007] Anonymous. Volume author index.

Journal of Computer and System Sciences, 68(4):862, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000534.

Anonymous:2004:CEBd

[2008] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 68(4):ifc, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000492.

Anonymous:2004:CTd

[2009] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 68(4):obc, June 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000601.

Condon:2004:GEF

[2010] Anne Condon. Guest Editor’s foreword. Journal of Computer and System Sciences, 69(1):1–2, August 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000004000029.

Razborov:2004:RLB

[2011] Alexander A. Razborov. Resolution lower bounds for perfect matching principles. Journal of Computer and System Sciences, 69(1): 3–27, August 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000030.

Green:2004:CBP

[2012] Frederic Green. The correlation between parity and quadratic polynomials mod 3. Journal of Computer and System Sciences, 69(1): 28–44, August 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000042.

Feige:2004:ILC

[2013] Uriel Feige and Daniele Micciancio. The inapproximability of lattice and coding problems with preprocessing. Journal of Computer and System Sciences, 69(1): 45–67, August 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000054.

ODonnell:2004:HAW

[2014] Ryan O’Donnell. Hardness amplification within NP. Journal of Computer and System Sciences, 69(1): 68–94, August 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000066.

Anonymous:2004:PAe

[2015] Anonymous. Papers to appear. Journal of Computer and System Sciences, 69(1):95, August 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000716.

Anonymous:2004:CEBe

[2016] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 69(1):ifc, August 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000686.

Anonymous:2004:CTe

[2017] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 69(1):obc, August 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000753.

Hitchcock:2004:SDN

[2018] John M. Hitchcock, Jack H. Lutz, and Elvira Mayordomo. Scaled dimension and nonuniform complexity. Journal of Computer and System Sciences, 69 (2):97–122, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000300151X.

Case:2004:RLR

[2019] John Case, Sanjay Jain, Frank Stephan, and Rolf Wiehagen. Robust learning-rich and poor. Journal of Computer and System Sciences, 69 (2):123–165, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000212.

Demaine:2004:AAC

[2020] Erik D. Demaine, Mohammad Taghi Hajiaghayi, Naomi Nishimura, Prabhakar Ragde, and Dimitrios M. Thilikos. Approximation algorithms for classes of graphs excluding singlecrossing graphs as minors. Journal of Computer and System Sciences, 69(2): 166–195, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000194.

Chawla:2004:AWS

[2021] Deepak Chawla, Lin Li, and Stephen Scott. On approximating weighted sums with exponentially many terms. Journal of Computer and System Sciences, 69(2):196–234, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000004000078.

Alexeev:2004:MDT

[2022] Boris Alexeev. Minimal DFA for testing divisibility. Journal of Computer and System Sciences, 69(2): 235–243, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000200.

Barnum:2004:LBQ

[2023] Howard Barnum and Michael Saks.

A lower bound on the quantum query complexity of read-once functions. Journal of Computer and System Sciences, 69(2):244–258, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200000400039X.

Damm:2004:RBC

[2024] Carsten Damm, Matthias Krause, Christoph Meinel, and Stephan Waack. On relations between counting communication complexity classes. Journal of Computer and System Sciences, 69 (2):259–280, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000340.

Lomonosov:2004:NQG

[2025] Andrew Lomonosov, Meera Sitharam, and Kihong Park. Network QoS games: stability vs optimality tradeoff. Journal of Computer and System Sciences, 69 (2):281–302, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000339.

Anonymous:2004:PAf

[2026] Anonymous. Papers to appear. Journal of Computer and System Sciences, 69(2):303, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000844.

Anonymous:2004:C

[2027] Anonymous. Contents. Journal of Computer and System Sciences, 69 (2):obc, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000400087X.

Anonymous:2004:CEBf

[2028] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 69(2):ifc, September 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000820.

Khanna:2004:SIA

[2029] Sanjeev Khanna and Aravind Srinivasan. Special issue: 35th Annual ACM Symposium on Theory of Computing. Journal of Computer and System Sciences, 69(3): 305, November 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000881.

Beier:2004:RKE

[2030] Rene Beier and Berthold Vöcking.

Random knapsack in expected polynomial time. Journal of Computer and System Sciences, 69(3): 306–329, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000431.

Thorup:2004:IPQ

[2031] Mikkel Thorup. Integer priority queues with decrease key in constant time and the single source shortest paths problem. Journal of Computer and System Sciences, 69(3): 330–353, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000400042X.

Alon:2004:TSD

[2032] Noga Alon and Asaf Shapira. Testing subgraphs in directed graphs. Journal of Computer and System Sciences, 69 (3):354–382, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000479.

Azar:2004:OOR

[2033] Yossi Azar, Edith Cohen, Amos Fiat, Haim Kaplan, and Harald Räcke. Optimal oblivious routing in polynomial time. Journal of Computer and System Sciences, 69(3): 383–394, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000625.

Kerenidis:2004:ELB

[2034] Iordanis Kerenidis and Ronald de Wolf. Exponential lower bound for 2-query locally decodable codes via a quantum argument. Journal of Computer and System Sciences, 69(3): 395–420, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000467.

Mossel:2004:LFK

[2035] Elchanan Mossel, Ryan O’Donnell, and Rocco A. Servedio. Learning functions of k relevant variables. Journal of Computer and System Sciences, 69 (3):421–434, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000418.

Jayram:2004:CPL

[2036] T. S. Jayram, Subhash Khot, Ravi Kumar, and Yuval Rabani. Cellprobe lower bounds for the partial match problem. Journal of Computer and System Sciences, 69(3): 435–447, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000455.

Gurvits:2004:CCQ

[2037] Leonid Gurvits. Classical complexity and quantum entanglement. Journal of Computer and System Sciences, 69 (3):448–484, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000893.

Fakcharoenphol:2004:TBA

[2038] Jittat Fakcharoenphol, Satish Rao, and Kunal Talwar. A tight bound on approximating arbitrary metrics by tree metrics. Journal of Computer and System Sciences, 69(3): 485–497, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000637.

Anonymous:2004:PAg

[2039] Anonymous. Papers to appear. Journal of Computer and System Sciences, 69(3):498, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000400100X.

Anonymous:2004:CEBg

[2040] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 69(3):ifc, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000984.

Anonymous:2004:CTf

[2041] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 69(3):obc, November 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001035.

Anonymous:2004:VTb

[2042] Anonymous. Volume TOC. Journal of Computer and System Sciences, 69 (4):III–V, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001138.

Ogihara:2004:RSI

[2043] Mitsunori Ogihara and Till Tantau. On the reducibility of sets inside NP to sets with low information content. Journal of Computer and System Sciences, 69 (4):499–524, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000352.

Gusfield:2004:LTA

[2044] Dan Gusfield and Jens Stoye. Linear time algorithms for finding and representing all the tandem repeats in a string. Journal of Computer and System Sciences, 69(4): 525–546, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000364.

Bar-Yehuda:2004:ADS

[2045] Reuven Bar-Yehuda and Zehavit Kehat. Approximating the dense setcover problem. Journal of Computer and System Sciences, 69(4): 547–561, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000388.

Bach:2004:ODQ

[2046] Eric Bach, Susan Coppersmith, Marcel Paz Goldschen, Robert Joynt, and John Watrous. One-dimensional quantum walks with absorbing boundaries. Journal of Computer and System Sciences, 69(4):562–592, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000004000376.

Maass:2004:CPC

[2047] Wolfgang Maass and Henry Markram.

On the computational power of circuits of spiking neurons. Journal of Computer and System Sciences, 69(4): 593–616, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000406.

Vandeurzen:2004:ELL

[2048] Luc Vandeurzen, Marc Gyssens, and Dirk Van Gucht. An expressive language for linear spatial database queries. Journal of Computer and System Sciences, 69(4): 617–655, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000443.

Szeider:2004:MUF

[2049] Stefan Szeider. Minimal unsatisfiable formulas with bounded clausevariable difference are fixed-parameter tractable. Journal of Computer and System Sciences, 69(4):656– 674, December 2004. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000613.

Chen:2004:CSM

[2050] Ning Chen, Xiaotie Deng, and Xiaoming Sun. On complexity of singleminded auction. Journal of Computer and System Sciences, 69(4): 675–687, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000649.

Huang:2004:GRS

[2051] He Huang, Jinde Cao, and Yuzhong Qu. Global robust stability of delayed neural networks with a class of general activation functions. Journal of Computer and System Sciences, 69 (4):688–700, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000662.

Anonymous:2004:PAh

[2052] Anonymous. Papers to appear. Journal of Computer and System Sciences, 69(4):701, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001096.

Anonymous:2004:VAIb

[2053] Anonymous. Volume author index. Journal of Computer and System Sciences, 69(4):702, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000004001102.

Anonymous:2004:CEBh

[2054] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 69(4):ifc, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001072.

Anonymous:2004:CTg

[2055] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 69(4):obc, December 2004. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001163.

Flesca:2005:POR

[2056] Sergio Flesca and Sergio Greco. Partially ordered regular languages for graph queries. Journal of Computer and System Sciences, 70(1):1– 25, February 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000935.

Huang:2005:EFI

[2057] Chongfu Huang and Claudio Moraga. Extracting fuzzy if-then rules by using the information matrix technique. Journal of Computer and System Sciences, 70(1):26– 52, February 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000650.

Legenstein:2005:WLC

[2058] Robert A. Legenstein and Wolfgang Maass. Wire length as a circuit complexity measure. Journal of Computer and System Sciences, 70(1):53– 72, February 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000765.

Li:2005:MDM

[2059] Deng-Feng Li. Multiattribute decision making models and methods using intuitionistic fuzzy sets. Journal of Computer and System Sciences, 70 (1):73–85, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000777.

Son:2005:EWC

[2060] Jin Hyun Son, Jung Sun Kim, and Myoung Ho Kim. Extracting the workflow critical path from the extended well-formed workflow schema. Journal of Computer and System Sciences, 70 (1):86–106, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000790.

Zhu:2005:PBB

[2061] Wenxing Zhu. A provable better branch and bound method for a nonconvex integer quadratic programming problem. Journal of Computer and System Sciences, 70(1): 107–117, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000400090X.

Even:2005:PEA

[2062] Guy Even, Peter-M. Seidel, and Warren E. Ferguson. A parametric error analysis of Goldschmidt’s division algorithm. Journal of Computer and System Sciences, 70(1): 118–139, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000960.

Hsieh:2005:EPP

[2063] Sun-Yuan Hsieh. Efficiently parallelizable problems on a class of decomposable graphs. Journal of Computer and System Sciences, 70(1): 140–156, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004000959.

Kari:2005:LEM

[2064] Lila Kari and Stavros Konstantinidis. Language equations, maximality and error-detection. Journal of Computer and System Sciences, 70 (1):157–178, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001047.

Anonymous:2005:PAa

[2065] Anonymous. Papers to appear. Journal of Computer and System Sciences, 70(1):179, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001436.

Anonymous:2005:CEBa

[2066] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 70(1):ifc, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001187.

Anonymous:2005:CTa

[2067] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 70(1):obc, February 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001229.

Barrington:2005:FOE

[2068] David A. Mix Barrington, Neil Immerman, Clemens Lautemann, Nicole Schweikardt, and Denis Thérien. First-order expressibility of languages with neutral letters or: The Crane Beach conjecture. Journal of Computer and System Sciences, 70(2): 101–127, March 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000807.

Goldwasser:2005:LTA

[2069] Michael H. Goldwasser, Ming-Yang Kao, and Hsueh-I Lu. Linear-time algorithms for computing maximumdensity sequence segments with bioinformatics applications. Journal of Computer and System Sciences, 70 (2):128–144, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000923.

Albers:2005:PLR

[2070] Susanne Albers, Lene M. Favrholdt, and Oliver Giel. On paging with locality of reference. Journal of Computer and System Sciences, 70 (2):145–175, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000947.

Qiu:2005:PMC

[2071] Daowen Qiu and Huaiqing Wang. A probabilistic model of computing with words. Journal of Computer and System Sciences, 70(2): 176–200, March 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001059.

Lee:2005:DIC

[2072] Jia Lee, Susumu Adachi, Ferdinand Peper, and Shinro Mashiko. Delay-insensitive computation in asynchronous cellular automata. Journal of Computer and System Sciences, 70 (2):201–220, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001357.

Neven:2005:AGU

[2073] Frank Neven. Attribute grammars for unranked trees as a query language for structured documents. Journal of Computer and System Sciences, 70 (2):221–257, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001345.

Santone:2005:LAT

[2074] Antonella Santone and Gigliola Vaglini. A local approach for temporal model checking of Java bytecode. Journal of Computer and System Sciences, 70 (2):258–281, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001448.

Anonymous:2005:PAb

[2075] Anonymous. Papers to appear. Journal of Computer and System Sciences, 70(2):282, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000061.

Anonymous:2005:CEBb

[2076] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 70(2):CO2, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000036.

Anonymous:2005:CTb

[2077] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 70(2):CO4, March 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000097.

Karp:2005:GEF

[2078] Richard Karp, Ming Li, Pavel Pevzner, and Ron Shamir. Guest editors’ foreword. Journal of Computer and System Sciences, 70(3): 283, May 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001473.

Wang:2005:NAS

[2079] Lusheng Wang, Daming Zhu, Xiaowen Liu, and Shaohan Ma. An o(n2) algorithm for signed translocation. Journal of Computer and System Sciences, 70(3):284–299, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001485.

Hartman:2005:AAS

[2080] Tzvika Hartman and Roded Sharan. A 1.5-approximation algorithm for sorting by transpositions and transreversals. Journal of Computer and System Sciences, 70(3): 300–320, May 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001497.

Kaplan:2005:SSP

[2081] Haim Kaplan and Elad Verbin. Sorting signed permutations by reversals, revisited. Journal of Computer and System Sciences, 70(3): 321–341, May 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001503.

Buhler:2005:DSS

[2082] Jeremy Buhler, Uri Keich, and Yanni Sun. Designing seeds for similarity search in genomic DNA. Journal of Computer and System Sciences, 70(3):342–363, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001515.

Brejova:2005:VSE

[2083] Brona Brejová, Daniel G. Brown, and Tomás Vinar. Vector seeds: an extension to spaced seeds. Journal of Computer and System Sciences, 70 (3):364–380, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001527.

Gusfield:2005:OER

[2084] Dan Gusfield. Optimal, efficient reconstruction of root-unknown phylogenetic networks with constrained and structured recombination. Journal of Computer and System Sciences, 70(3):381–398, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001539.

Aumann:2005:DOM

[2085] Yonatan Aumann, Efrat Manisterski, and Zohar Yakhini. Designing optimally multiplexed SNP genotyping assays. Journal of Computer and System Sciences, 70(3): 399–417, May 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001540.

Ma:2005:EAP

[2086] Bin Ma, Kaizhong Zhang, and Chengzhi Liang. An effective algorithm for peptide de novo sequencing from MS/MS spectra. Journal of Computer and System Sciences, 70(3):418–430, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001552.

Anonymous:2005:PAc

[2087] Anonymous. Papers to appear. Journal of Computer and System Sciences, 70(3):431, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000218.

Anonymous:2005:CEBc

[2088] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 70(3):CO2, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000188.

Anonymous:2005:CTc

[2089] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 70(3):CO4, May 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000243.

Anonymous:2005:VT

[2090] Anonymous. Volume TOC. Journal of Computer and System Sciences, 70(4):III–V, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000395.

Sloan:2005:SIC

[2091] Robert Sloan. Special issue on computational learning theory. Journal of Computer and System Sciences, 70(4):433–434, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001266.

Hellerstein:2005:ELD

[2092] Lisa Hellerstein and Vijay Raghavan. Exact learning of DNF formulas using DNF hypotheses. Journal of Computer and System Sciences, 70(4):435–470, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001278.

Bshouty:2005:ELB

[2093] Nader H. Bshouty, Jeffrey C. Jackson, and Christino Tamon. Exploring learnability between exact and PAC. Journal of Computer and System Sciences, 70(4):471–484, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000400128X.

Yang:2005:NLB

[2094] Ke Yang. New lower bounds for statistical query learning. Journal of Computer and System Sciences, 70 (4):485–509, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001291.

Zilles:2005:IPU

[2095] Sandra Zilles. Increasing the power of uniform inductive learners. Journal of Computer and System Sciences, 70(4):510–538, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001308.

Vyugin:2005:PCI

[2096] Michael V. Vyugin and Vladimir V. V’yugin. Predictive complexity and information. Journal of Computer and System Sciences, 70(4): 539–554, June 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000400131X.

Dasgupta:2005:PGH

[2097] Sanjoy Dasgupta and Philip M. Long. Performance guarantees for hierarchical clustering. Journal of Computer and System Sciences, 70(4): 555–569, June 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001321.

Fortnow:2005:PD

[2098] Lance Fortnow and Jack H. Lutz. Prediction and dimension. Journal of Computer and System Sciences, 70 (4):570–589, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001333.

Anonymous:2005:PAd

[2099] Anonymous. Papers to appear. Journal of Computer and System Sciences, 70(4):590, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000358.

Anonymous:2005:VAI

[2100] Anonymous. Volume author index.

Journal of Computer and System Sciences, 70(4):591, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000500036X.

Anonymous:2005:CEBd

[2101] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 70(4):CO2, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000322.

Anonymous:2005:CTd

[2102] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 70(4):CO4, June 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000413.

Koch:2005:GBS

[2103] Manuel Koch, L. V. Mancini, and Francesco Parisi-Presicce. Graph-based specification of access control policies. Journal of Computer and System Sciences, 71(1):1–33, July 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000400145X.

Lanotte:2005:MHS

[2104] Ruggero Lanotte and Andrea Maggiolo-Schettini. Monotonic hybrid systems. Journal of Computer and System Sciences, 71(1):34–69, July 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001461.

Huang:2005:SSA

[2105] Tetz C. Huang. A self-stabilizing algorithm for the shortest path problem assuming read/write atomicity. Journal of Computer and System Sciences, 71(1):70–85, July 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000139.

Reif:2005:EPF

[2106] John H. Reif. Efficient parallel factorization and solution of structured and unstructured linear systems. Journal of Computer and System Sciences, 71(1):86–143, July 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000127.

Anonymous:2005:PAe

[2107] Anonymous. Papers to appear. Journal of Computer and System Sciences, 71(1):144, July 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000516.

Anonymous:2005:CEBe

[2108] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 71(1):CO2, July 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000450.

Anonymous:2005:CTe

[2109] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 71(1):CO4, July 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000504.

Berman:2005:TAR

[2110] Piotr Berman, Bhaskar DasGupta, and Ming-Yang Kao. Tight approximability results for test set problems in bioinformatics. Journal of Computer and System Sciences, 71 (2):145–162, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000140.

Xu:2005:IAM

[2111] Baowen Xu, Yuming Zhou, and Hongmin Lu. An improved accuracy measure for rough sets. Journal of Computer and System Sciences, 71 (2):163–173, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000152.

Gutin:2005:KPD

[2112] Gregory Gutin, Ton Kloks, Chuan Min Lee, and Anders Yeo. Kernels in planar digraphs. Journal of Computer and System Sciences, 71(2): 174–184, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000164.

Kiltz:2005:TCL

[2113] Eike Kiltz and Hans Ulrich Simon. Threshold circuit lower bounds on cryptographic functions. Journal of Computer and System Sciences, 71 (2):185–212, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000292.

Beimel:2005:GCI

[2114] Amos Beimel, Yuval Ishai, and Eyal Kushilevitz. General constructions for information-theoretic private information retrieval. Journal of Computer and System Sciences, 71(2): 213–247, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000309.

Anonymous:2005:PAf

[2115] Anonymous. Papers to appear. Journal of Computer and System Sciences, 71(2):248, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000590.

Anonymous:2005:CEBf

[2116] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 71(2):CO2, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000565.

Anonymous:2005:CTf

[2117] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 71(2):CO4, August 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000620.

Goldman:2005:SIL

[2118] Sally Goldman. Special issue: Learning theory 2003. Journal of Computer and System Sciences, 71(3): 249, October 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001369.

Bshouty:2005:LDR

[2119] Nader H. Bshouty, Elchanan Mossel, Ryan O’Donnell, and Rocco A. Servedio. Learning DNF from random walks. Journal of Computer and System Sciences, 71(3): 250–265, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001370.

Kalai:2005:BPN

[2120] Adam Tauman Kalai and Rocco A. Servedio. Boosting in the presence of noise. Journal of Computer and System Sciences, 71(3): 266–290, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001382.

Kalai:2005:EAO

[2121] Adam Kalai and Santosh Vempala. Efficient algorithms for online decision problems. Journal of Computer and System Sciences, 71(3): 291–307, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001394.

Case:2005:LCR

[2122] John Case, Sanjay Jain, Franco Montagna, Giulia Simi, and Andrea Sorbi. On learning to coordinate: random bits help, insightful normal forms, and competency isomorphisms. Journal of Computer and System Sciences, 71 (3):308–332, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000004001400.

Hein:2005:MMC

[2123] Matthias Hein, Olivier Bousquet, and Bernhard Schölkopf. Maximal margin classification for metric spaces. Journal of Computer and System Sciences, 71 (3):333–359, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001412.

Charikar:2005:CQI

[2124] Moses Charikar, Venkatesan Guruswami, and Anthony Wirth. Clustering with qualitative information. Journal of Computer and System Sciences, 71(3):360–383, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004001424.

Anonymous:2005:PAg

[2125] Anonymous. Papers to appear. Journal of Computer and System Sciences, 71(3):384, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000814.

Anonymous:2005:CEBg

[2126] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 71(3):CO2, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000784.

Anonymous:2005:CTg

[2127] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 71(3):CO4, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000838.

Anonymous:2005:LT

[2128] Anonymous. Learning theory 2003. Journal of Computer and System Sciences, 71(3):??, October 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Alber:2005:RST

[2129] Jochen Alber, Hongbing Fan, Michael R. Fellows, Henning Fernau, Rolf Niedermeier, Fran Rosamond, and Ulrike Stege. A refined search tree technique for dominating set on planar graphs. Journal of Computer and System Sciences, 71(4): 385–405, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000789.

Vikas:2005:CEC

[2130] Narayan Vikas. A complete and equal computational complexity classification of compaction and retraction to all graphs with at most four vertices and some general results. Journal of Computer and System Sciences, 71(4): 406–439, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000004000911.

Goldstine:2005:MNP

[2131] Jonathan Goldstine, Hing Leung, and Detlef Wotschke. Measuring nondeterminism in pushdown automata. Journal of Computer and System Sciences, 71(4):440– 466, November 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000437.

Dubhashi:2005:FDA

[2132] Devdatt Dubhashi, Alessandro Mei, Alessandro Panconesi, Jaikumar Radhakrishnan, and Aravind Srinivasan. Fast distributed algorithms for (weakly) connected dominating sets and linearsize skeletons. Journal of Computer and System Sciences, 71(4): 467–479, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000500053X.

Karaata:2005:OSS

[2133] Mehmet Hakan Karaata. An optimal self-stabilizing strarvation-free alternator. Journal of Computer and System Sciences, 71(4):480– 494, November 2005. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000541.

Kolliopoulos:2005:AAC

[2134] Stavros G. Kolliopoulos and Neal E.

Young. Approximation algorithms for covering/packing integer programs. Journal of Computer and System Sciences, 71(4):495–505, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000005000656.

Honkala:2005:BUE

[2135] Juha Honkala. An n^{2}-bound for the ultimate equivalence problem of certain D0L systems over an nletter alphabet. Journal of Computer and System Sciences, 71(4): 506–519, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000668.

Woelfel:2005:BOS

[2136] Philipp Woelfel. Bounds on the OBDDsize of integer multiplication via universal hashing. Journal of Computer and System Sciences, 71(4): 520–534, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000500067X.

Gori:2005:VFF

[2137] Roberta Gori and Giorgio Levi. On the verification of finite failure. Journal of Computer and System Sciences, 71 (4):535–575, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000681.

Anonymous:2005:PAh

[2138] Anonymous. Papers to appear. Journal of Computer and System Sciences, 71(4):576, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005001030.

Anonymous:2005:CEBh

[2139] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 71(4):CO2, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001005.

Anonymous:2005:CTh

[2140] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 71 (4):CO4, November 2005. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001066.

Dooly:2006:RVM

[2141] Daniel R. Dooly, Sally A. Goldman, and Stephen S. Kwek. Realvalued multiple-instance learning with queries. Journal of Computer and System Sciences, 72(1):1–15, February 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000693.

Gandhi:2006:IAA

[2142] Rajiv Gandhi, Eran Halperin, Samir Khuller, Guy Kortsarz, and Aravind Srinivasan. An improved approximation algorithm for vertex cover with hard capacities. Journal of Computer and System Sciences, 72 (1):16–33, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000747.

Flum:2006:BFP

[2143] Jörg Flum, Martin Grohe, and Mark Weyer. Bounded fixed-parameter tractability and log2n nondeterministic bits. Journal of Computer and System Sciences, 72(1):34– 71, February 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000500070X.

Khardon:2006:SLQ

[2144] Roni Khardon and Marta Arias. The subsumption lattice and query learning. Journal of Computer and System Sciences, 72(1):72–94, February 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000759.

Hutter:2006:SPB

[2145] Marcus Hutter. Sequential predictions based on algorithmic complexity. Journal of Computer and System Sciences, 72(1):95–117, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000723.

Andrews:2006:MMF

[2146] Matthew Andrews and Lisa Zhang. Minimizing maximum fiber requirement in optical networks. Journal of Computer and System Sciences, 72 (1):118–131, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000735.

Epstein:2006:OPS

[2147] Leah Epstein and Tamir Tassa. Optimal preemptive scheduling for general target functions. Journal of Computer and System Sciences, 72 (1):132–162, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000760.

McKenzie:2006:MFT

[2148] Pierre McKenzie, Thomas Schwentick, Denis Thérien, and Heribert Vollmer. The many faces of a translation. Journal of Computer and System Sciences, 72(1):163–179, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000005000887.

Geeraerts:2006:EEC

[2149] G. Geeraerts, J.-F. Raskin, and L. Van Begin. Expand, enlarge and check: New algorithms for the coverability problem of WSTS. Journal of Computer and System Sciences, 72(1): 180–203, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000875.

Anonymous:2006:PAa

[2150] Anonymous. Papers to appear. Journal of Computer and System Sciences, 72(1):204, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001200.

Anonymous:2006:CEBa

[2151] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 72(1):CO2, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005001170.

Anonymous:2006:CTa

[2152] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 72(1):CO4, February 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005001236.

Chekuri:2006:SIF

[2153] Chandra Chekuri and Daniele Micciancio. Special issue: FOCS 2003. Journal of Computer and System Sciences, 72(2):205, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000905.

Khot:2006:HAS

[2154] Subhash Khot. Hardness of approximating the shortest vector problem in high ℓ_{p} norms. Journal of Computer and System Sciences, 72 (2):206–219, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000930.

Ambainis:2006:PDV

[2155] Andris Ambainis. Polynomial degree vs. quantum query complexity. Journal of Computer and System Sciences, 72(2):220–238, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000899.

Mihail:2006:CCP

[2156] Milena Mihail, Christos Papadimitriou, and Amin Saberi. On certain connectivity properties of the Internet topology. Journal of Computer and System Sciences, 72(2): 239–251, March 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000942.

Bhatnagar:2006:SPS

[2157] Nayantara Bhatnagar, Parikshit Gopalan, and Richard J. Lipton. Symmetric polynomials over Z_{m} and simultaneous communication protocols. Journal of Computer and System Sciences, 72(2):252–285, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000917.

Impagliazzo:2006:LRA

[2158] Russell Impagliazzo and Bruce M.

Kapron. Logics for reasoning about cryptographic constructions. Journal of Computer and System Sciences, 72 (2):286–320, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000929.

Barak:2006:LBN

[2159] Boaz Barak, Yehuda Lindell, and Salil Vadhan. Lower bounds for nonblack-box zero knowledge. Journal of Computer and System Sciences, 72 (2):321–391, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000954.

Lovasz:2006:SAC

[2160] László Lovász and Santosh Vempala. Simulated annealing in convex bodies and an o*(n4) volume algorithm. Journal of Computer and System Sciences, 72(2):392–417, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005000966.

Anonymous:2006:PAb

[2161] Anonymous. Papers to appear. Journal of Computer and System Sciences, 72(2):418, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000600016X.

Anonymous:2006:CEBb

[2162] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 72(2):CO2, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000134.

Anonymous:2006:CTb

[2163] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 72(2):CO4, March 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000195.

Graham:2006:F

[2164] Fan Chung Graham. Foreword. Journal of Computer and System Sciences, 72(3):419, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001078.

Graham:2006:BON

[2165] Fan Chung Graham. A brief overview of network algorithms. Journal of Computer and System Sciences, 72 (3):420–424, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001145.

Plaxton:2006:AAH

[2166] C. Greg Plaxton. Approximation algorithms for hierarchical location problems. Journal of Computer and System Sciences, 72(3): 425–443, May 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000500111X.

Cole:2006:HMC

[2167] Richard Cole, Yevgeniy Dodis, and Tim Roughgarden. How much can taxes help selfish routing? Journal of Computer and System Sciences, 72(3):444–467, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005001121.

Chen:2006:MMA

[2168] Jiangzhuo Chen, Rajmohan Rajaraman, and Ravi Sundaram. Meet and merge: Approximation algorithms for confluent flows. Journal of Computer and System Sciences, 72(3):468–489, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005001133.

Elsasser:2006:RCR

[2169] R. Elsässer and L. Gasieniec. Radio communication in random graphs. Journal of Computer and System Sciences, 72(3):490–506, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000005001108.

Anonymous:2006:PAc

[2170] Anonymous. Papers to appear. Journal of Computer and System Sciences, 72(3):507, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000390.

Anonymous:2006:CEBc

[2171] Anonymous. C2 — Ed. Board. Journal of Computer and System Sciences, 72(3):CO2, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000365.

Anonymous:2006:CTc

[2172] Anonymous. C4 — TOC. Journal of Computer and System Sciences, 72(3):CO4, May 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000419.

Engebretsen:2006:TBM

[2173] Lars Engebretsen and Marek Karpinski. TSP with bounded metrics. Journal of Computer and System Sciences, 72(4):509–546, June 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001285.

Demri:2006:PAS

[2174] S. Demri, F. Laroussinie, and Ph.

Schnoebelen. A parametric analysis of the state-explosion problem in model checking. Journal of Computer and System Sciences, 72(4): 547–575, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001297.

Gross-Amblard:2006:UGS

[2175] David Gross-Amblard and Michel de Rougemont. Uniform generation in spatial constraint databases and applications. Journal of Computer and System Sciences, 72(4): 576–591, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005000978.

Hassin:2006:APC

[2176] Yehuda Hassin and David Peleg. Average probe complexity in quorum systems. Journal of Computer and System Sciences, 72(4): 592–616, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000500125X.

Genest:2006:ISH

[2177] Blaise Genest, Anca Muscholl, Helmut Seidl, and Marc Zeitoun. Infinitestate high-level MSCs: Model-checking and realizability. Journal of Computer and System Sciences, 72(4): 617–647, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000500098X.

Elkin:2006:SAT

[2178] Michael Elkin and Guy Kortsarz. Sublogarithmic approximation for telephone multicast. Journal of Computer and System Sciences, 72(4): 648–659, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000005001303.

Spakowski:2006:LWU

[2179] Holger Spakowski and Rahul Tripathi. LWPP and WPP are not uniformly gap-definable. Journal of Computer and System Sciences, 72 (4):660–689, June 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000031.

Shao:2006:SDT

[2180] Zuhua Shao. Security of the design of time-stamped signatures. Journal of Computer and System Sciences, 72(4):690–705, June 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000600002X.

Chunlin:2006:DMD

[2181] Li Chunlin and Li Layuan. A distributed multiple dimensional QoS constrained resource scheduling optimization policy in computational grid. Journal of Computer and System Sciences, 72(4):706–726, June 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000006000110.

Liao:2006:PAS

[2182] I-En Liao, Cheng-Chi Lee, and Min-Shiang Hwang. A password authentication scheme over insecure networks. Journal of Computer and System Sciences, 72(4): 727–740, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001157.

Xia:2006:FLM

[2183] Hui-Cheng Xia, Deng-Feng Li, Ji-Yan Zhou, and Jian-Ming Wang. Fuzzy LINMAP method for multiattribute decision making under fuzzy environments. Journal of Computer and System Sciences, 72(4): 741–759, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001261.

Hitchcock:2006:DER

[2184] John M. Hitchcock and N. V. Vinodchandran. Dimension, entropy rates, and compression. Journal of Computer and System Sciences, 72 (4):760–782, June 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001248.

Jenner:2006:CCR

[2185] Birgit Jenner, Johannes Köbler, Pierre McKenzie, and Jacobo Torán. Corrigendum to “Completeness results for graph isomorphism” [J. Comput. System Sci. 66 (2003) 549–566]. Journal of Computer and System Sciences, 72(4):783, June 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000005001273. See

[1868].

Anonymous:2006:PAFa

[2186] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 72(4): 784, June 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000596.

Anonymous:2006:EBa

[2187] Anonymous. Editorial Board. Journal of Computer and System Sciences, 72(4):CO2, June 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000560.

Charikar:2006:GEF

[2188] Moses Charikar. Guest Editor’s foreword. Journal of Computer and System Sciences, 72(5):785, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000006000043.

Ta-Shma:2006:ERM

[2189] Amnon Ta-Shma, David Zuckerman, and Shmuel Safra. Extractors from Reed–Muller codes. Journal of Computer and System Sciences, 72(5): 786–812, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000055.

Demetrescu:2006:FDA

[2190] Camil Demetrescu and Giuseppe F.

Italiano. Fully dynamic all pairs shortest paths with real edge weights. Journal of Computer and System Sciences, 72(5):813–837, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000067.

Fleischer:2006:IRA

[2191] Lisa Fleischer, Kamal Jain, and David P. Williamson. Iterative rounding 2-approximation algorithms for minimum-cost vertex connectivity problems. Journal of Computer and System Sciences, 72(5): 838–867, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000079.

Fakcharoenphol:2006:PGN

[2192] Jittat Fakcharoenphol and Satish Rao. Planar graphs, negative weight edges, shortest paths, and near linear time. Journal of Computer and System Sciences, 72(5): 868–889, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000080.

Bartal:2006:RTT

[2193] Yair Bartal, Béla Bollobás, and Manor Mendel. Ramsey-type theorems for metric spaces with applications to online problems. Journal of Computer and System Sciences, 72(5): 890–921, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000092.

Roughgarden:2006:SBP

[2194] Tim Roughgarden. On the severity of Braess’s paradox: Designing networks for selfish users is hard. Journal of Computer and System Sciences, 72(5):922–953, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000109.

Anonymous:2006:PAFb

[2195] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 72(5): 954, August 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000729.

Anonymous:2006:EBb

[2196] Anonymous. Editorial Board. Journal of Computer and System Sciences, 72(5):CO2, August 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000699.

Sabharwal:2006:NNS

[2197] Yogish Sabharwal, Nishant Sharma, and Sandeep Sen. Nearest neighbors search using point location in balls with applications to approximate Voronoi decompositions. Journal of Computer and System Sciences, 72(6): 955–977, September 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000600078X.

DiRaimondo:2006:PST

[2198] Mario Di Raimondo and Rosario Gennaro. Provably secure threshold password-authenticated key exchange. Journal of Computer and System Sciences, 72(6):978–1001, September 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000600033X.

Sun:2006:PBA

[2199] Hung-Min Sun and Her-Tyan Yeh. Password-based authentication and key distribution protocols with perfect forward secrecy. Journal of Computer and System Sciences, 72(6): 1002–1011, September 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000481.

Parnas:2006:TPT

[2200] Michal Parnas, Dana Ron, and Ronitt Rubinfeld. Tolerant property testing and distance approximation. Journal of Computer and System Sciences, 72(6): 1012–1042, September 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000444.

Bohler:2006:EBP

[2201] Elmar Böhler, Christian Glaßer, and Daniel Meister. Error-bounded probabilistic computations between MA and AM. Journal of Computer and System Sciences, 72(6):1043– 1076, September 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000778.

Aspnes:2006:ISV

[2202] James Aspnes, Kevin Chang, and Aleksandr Yampolskiy. Inoculation strategies for victims of viruses and the sum-of-squares partition problem. Journal of Computer and System Sciences, 72(6):1077–1093, September 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000328.

Srinivasan:2006:ORB

[2203] Anand Srinivasan and James H. Anderson. Optimal rate-based scheduling on multiprocessors. Journal of Computer and System Sciences, 72(6): 1094–1117, September 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000432.

Anonymous:2006:Ca

[2204] Anonymous. Contents. Journal of Computer and System Sciences, 72 (6):CO4, September 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000869.

Anonymous:2006:EBc

[2205] Anonymous. Editorial Board. Journal of Computer and System Sciences, 72(6):CO2, September 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000821.

Awan:2006:PME

[2206] Irfan Awan. Performance modelling and evaluation of computer systems. Journal of Computer and System Sciences, 72(7):1119– 1120, November 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000298.

Ferro:2006:AML

[2207] Armando Ferro, Igor Delgado, Alex Muñoz, and Fidel Liberal. An analytical model for loss estimation in network traffic analysis systems. Journal of Computer and System Sciences, 72(7): 1121–1133, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000201.

Botta:2006:SPM

[2208] Alessio Botta, Donato Emma, Antonio Pescapé, and Giorgio Ventre. Systematic performance modeling and characterization of heterogeneous IP networks. Journal of Computer and System Sciences, 72(7): 1134–1143, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000213.

Er:2006:COC

[2209] Inn Inn Er and Winston K. G. Seah. Clustering overhead and convergence time analysis of the mobility-based multi-hop clustering algorithm for mobile ad hoc networks. Journal of Computer and System Sciences, 72(7): 1144–1155, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000225.

Konrad:2006:DMA

[2210] Almudena Konrad, Ben Y. Zhao, and Anthony D. Joseph. Determining model accuracy of network traces. Journal of Computer and System Sciences, 72(7):1156– 1171, November 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000237.

Pereira:2006:PES

[2211] E. Grishikashvili Pereira, R. Pereira, and A. Taleb-Bendiab. Performance evaluation for self-healing distributed services and fault detection mechanisms. Journal of Computer and System Sciences, 72(7): 1172–1182, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000249.

Ikeda:2006:PEI

[2212] Makoto Ikeda, Leonard Barolli, Akio Koyama, Arjan Durresi, Giuseppe De Marco, and Jiro Iwashige. Performance evaluation of an intelligent CAC and routing framework for multimedia applications in broadband networks. Journal of Computer and System Sciences, 72(7): 1183–1200, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000250.

Tan:2006:PSL

[2213] Guang Tan and Stephen A. Jarvis. Prediction of short-lived TCP transfer latency on bandwidth asymmetric links. Journal of Computer and System Sciences, 72(7):1201– 1210, November 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000262.

Itaya:2006:HSP

[2214] Satoshi Itaya, Naohiro Hayashibara, Tomoya Enokido, and Makoto Takizawa. HAMS: Scalable peer-to-peer multimedia streaming model in heterogeneous networks. Journal of Computer and System Sciences, 72(7): 1211–1225, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000420.

Younas:2006:TCP

[2215] Muhammad Younas and Kuo-Ming Chao. A tentative commit protocol for composite web services. Journal of Computer and System Sciences, 72(7): 1226–1237, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000274.

Guan:2006:CQD

[2216] L. Guan, M. E. Woodward, and I. U. Awan. Control of queueing delay in a buffer with timevarying arrival rate. Journal of Computer and System Sciences, 72(7):

1238–1248, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000286.

Anonymous:2006:PAFc

[2217] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 72(7): 1249, November 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000973.

Anonymous:2006:Cb

[2218] Anonymous. Contents. Journal of Computer and System Sciences, 72 (7):CO4, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000997.

Anonymous:2006:EBd

[2219] Anonymous. Editorial Board. Journal of Computer and System Sciences, 72(7):CO2, November 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000948.

Anonymous:2006:CV

[2220] Anonymous. Contents of volume 72. Journal of Computer and System Sciences, 72(8):III–VII, December 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001243.

Grov:2006:TNR

[2221] Jon Grov and Ragnar Normann. Technical note: Replication graphs revisited. Journal of Computer and System Sciences, 72(8):1251– 1261, December 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000341.

Chlebus:2006:RGA

[2222] Bogdan S. Chlebus and Dariusz R. Kowalski. Robust gossiping with an application to consensus. Journal of Computer and System Sciences, 72(8): 1262–1281, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000894.

Elkin:2006:FDP

[2223] Michael Elkin. A faster distributed protocol for constructing a minimum spanning tree. Journal of Computer and System Sciences, 72(8): 1282–1308, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000808.

Chung:2006:MDL

[2224] Fan Chung, Ronald Graham, Ranjita Bhagwan, Stefan Savage, and Geoffrey M. Voelker. Maximizing data locality in distributed systems. Journal of Computer and System Sciences, 72(8):1309–1316, December 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000791.

Barrett:2006:CRP

[2225] Christopher L. Barrett, Harry B. Hunt, Madhav V. Marathe, S. S. Ravi, Daniel J. Rosenkrantz, and Richard E. Stearns. Complexity of reachability problems for finite discrete dynamical systems. Journal of Computer and System Sciences, 72(8): 1317–1345, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000600047X.

Chen:2006:SCL

[2226] Jianer Chen, Xiuzhen Huang, Iyad A.

Kanj, and Ge Xia. Strong computational lower bounds via parameterized complexity. Journal of Computer and System Sciences, 72(8): 1346–1367, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000675.

Takada:2006:CUP

[2227] Yousuke Takada, Teijiro Isokawa, Ferdinand Peper, and Nobuyuki Matsui. Construction universality in purely asynchronous cellular automata. Journal of Computer and System Sciences, 72(8): 1368–1385, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000663.

Guo:2006:CBF

[2228] Jiong Guo, Jens Gramm, Falk Hüffner, Rolf Niedermeier, and Sebastian Wernicke. Compressionbased fixed-parameter algorithms for feedback vertex set and edge bipartization. Journal of Computer and System Sciences, 72(8):1386– 1396, December 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000316.

Gradara:2006:DED

[2229] Sara Gradara, Antonella Santone, and Maria Luisa Villani. DELFIN +: an efficient deadlock detection tool for CCS processes. Journal of Computer and System Sciences, 72(8): 1397–1412, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000456.

Anonymous:2006:PAFd

[2230] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 72(8): 1413, December 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001206.

Anonymous:2006:AIV

[2231] Anonymous. Author index for volume 72. Journal of Computer and System Sciences, 72(8):1414– 1415, December 2006. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001218.

Anonymous:2006:Cc

[2232] Anonymous. Contents. Journal of Computer and System Sciences, 72 (8):CO4, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001267.

Anonymous:2006:EBe

[2233] Anonymous. Editorial Board. Journal of Computer and System Sciences, 72(8):CO2, December 2006. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001176.

Bruyere:2007:ALO

[2234] Véronique Bruyère and Olivier Carton.

Automata on linear orderings. Journal of Computer and System Sciences, 73(1):1–24, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001322.

Cai:2007:Z

[2235] Jin-Yi Cai. S_{2}^{p} ⊆ ZPP^{NP}. Journal of Computer and System Sciences, 73 (1):25–35, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000600105X.

Glasser:2007:LPT

[2236] Christian Glaßer. Languages polylogtime reducible to dot-depth 1/2. Journal of Computer and System Sciences, 73(1):36–56, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001024.

Downey:2007:OPP

[2237] Rodney G. Downey and Catherine McCartin. Online promise problems with online width metrics. Journal of Computer and System Sciences, 73 (1):57–72, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000900.

Hemaspaandra:2007:DVS

[2238] Edith Hemaspaandra and Lane A. Hemaspaandra. Dichotomy for voting systems. Journal of Computer and System Sciences, 73(1):73– 83, February 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000924.

Lin:2007:MFD

[2239] Lin Lin, Xue-Hai Yuan, and Zun-Quan Xia. Multicriteria fuzzy decisionmaking methods based on intuitionistic fuzzy sets. Journal of Computer and System Sciences, 73(1):84– 88, February 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000468.

Jain:2007:LLU

[2240] Sanjay Jain, Yen Kaow Ng, and Tiong Seng Tay. Learning languages in a union. Journal of Computer and System Sciences, 73(1): 89–108, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000304.

Harizanov:2007:LVS

[2241] Valentina S. Harizanov and Frank Stephan. On the learnability of vector spaces. Journal of Computer and System Sciences, 73(1): 109–122, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000912.

Cheng:2007:QCS

[2242] Sheng-Tzong Cheng and Ming-Hung Tao. Quantum cooperative search algorithm for 3-SAT. Journal of Computer and System Sciences, 73(1): 123–136, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001012.

Sosik:2007:MCC

[2243] Petr Sosík and Alfonso Rodríguez-Patón. Membrane computing and complexity theory: A characterization of PSPACE. Journal of Computer and System Sciences, 73(1): 137–152, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001048.

Anonymous:2007:PAFa

[2244] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 73(1): 153, February 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001541.

Anonymous:2007:EBa

[2245] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(1):CO2, February 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001516.

Anonymous:2007:VNF

[2246] Anonymous. Volume 73, number 1, February 2007. Journal of Computer and System Sciences, 73(1): CO4, February 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001577.

Gill:2007:GEF

[2247] Christopher Gill and Oleg Sokolsky. Guest editors’ foreword. Journal of Computer and System Sciences, 73(2):155, March 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000547.

Alur:2007:DSE

[2248] Rajeev Alur and Arun Chandrashekharapuram. Dispatch sequences for embedded control models. Journal of Computer and System Sciences, 73 (2):156–170, March 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000535.

Balasubramanian:2007:PIC

[2249] Krishnakumar Balasubramanian, Jaiganesh Balasubramanian, Jeff Parsons, Aniruddha Gokhale, and Douglas C. Schmidt. A platform-independent component modeling language for distributed realtime and embedded systems. Journal of Computer and System Sciences, 73(2):171–185, March 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000600050X.

Pellizzoni:2007:HAA

[2250] Rodolfo Pellizzoni and Giuseppe Lipari. Holistic analysis of asynchronous real-time transactions with earliest deadline scheduling. Journal of Computer and System Sciences, 73 (2):186–206, March 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006000523.

Wandeler:2007:WCM

[2251] Ernesto Wandeler and Lothar Thiele.

Workload correlations in multi-processor hard real-time systems. Journal of Computer and System Sciences, 73 (2):207–224, March 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000511.

Andersson:2007:EAC

[2252] Björn Andersson and Cecilia Ekelin. Exact admission-control for integrated aperiodic and periodic tasks. Journal of Computer and System Sciences, 73(2):225–241, March 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006000493.

Anonymous:2007:PAFb

[2253] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 73(2): 242, March 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001632.

Anonymous:2007:EBb

[2254] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(2):CO2, March 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001607.

Anonymous:2007:VNM

[2255] Anonymous. Volume 73, number 2, March 2007. Journal of Computer and System Sciences, 73(2): CO4, March 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000600167X.

Benedikt:2007:DT

[2256] Michael Benedikt. Database theory 2004. Journal of Computer and System Sciences, 73(3):243– 244, May 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001061.

Bawa:2007:PVD

[2257] Mayank Bawa, Aristides Gionis, Hector Garcia-Molina, and Rajeev Motwani. The price of validity in dynamic networks. Journal of Computer and System Sciences, 73(3): 245–264, May 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001115.

Cohen:2007:SDA

[2258] Edith Cohen and Haim Kaplan.

Spatially-decaying aggregation over a network. Journal of Computer and System Sciences, 73(3): 265–288, May 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001103.

Schwentick:2007:AXS

[2259] Thomas Schwentick. Automata for XML-a survey. Journal of Computer and System Sciences, 73(3): 289–315, May 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001085.

Davidson:2007:PXC

[2260] Susan Davidson, Wenfei Fan, and Carmem Hara. Propagating XML constraints to relations. Journal of Computer and System Sciences, 73 (3):316–361, May 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001152. See

erratum [2359].

Martens:2007:FTT

[2261] Wim Martens and Frank Neven. Frontiers of tractability for typechecking simple XML transformations. Journal of Computer and System Sciences, 73(3):362–390, May 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001139.

Bar-Yossef:2007:MRX

[2262] Ziv Bar-Yossef, Marcus Fontoura, and Vanja Josifovski. On the memory requirements of XPath evaluation over XML streams. Journal of Computer and System Sciences, 73 (3):391–441, May 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001073.

Deutsch:2007:SVD

[2263] Alin Deutsch, Liying Sui, and Victor Vianu. Specification and verification of data-driven Web applications. Journal of Computer and System Sciences, 73(3):442–474, May 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001140.

Scarcello:2007:WHD

[2264] Francesco Scarcello, Gianluigi Greco, and Nicola Leone. Weighted hypertree decompositions and optimal query plans. Journal of Computer and System Sciences, 73(3): 475–506, May 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001097.

Miklau:2007:FAI

[2265] Gerome Miklau and Dan Suciu. A formal analysis of information disclosure in data exchange. Journal of Computer and System Sciences, 73(3):507–534, May 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001127.

Anonymous:2007:PAFc

[2266] Anonymous. Papers to appear in forthcoming issues. Journal of Computer and System Sciences, 73 (3):535, May 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001784.

Anonymous:2007:Ca

[2267] Anonymous. Contents. Journal of Computer and System Sciences, 73 (3):CO4, May 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001802.

Anonymous:2007:EBc

[2268] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(3):CO2, May 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001759.

Suciu:2007:SID

[2269] Dan Suciu. Special issue on database theory. Journal of Computer and System Sciences, 73(4): 537, June 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001383.

Leinders:2007:CDS

[2270] Dirk Leinders and Jan Van den Bussche. On the complexity of division and set joins in the relational algebra. Journal of Computer and System Sciences, 73(4):538–549, June 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001395.

Martens:2007:MXS

[2271] Wim Martens and Joachim Niehren. On the minimization of XML schemas and tree automata for unranked trees. Journal of Computer and System Sciences, 73(4):550–583, June 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000006001401.

Faber:2007:MST

[2272] Wolfgang Faber, Gianluigi Greco, and Nicola Leone. Magic sets and their application to data integration. Journal of Computer and System Sciences, 73(4):584–609, June 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001413.

Fuxman:2007:FOQ

[2273] Ariel Fuxman and Renée J. Miller.

First-order query rewriting for inconsistent databases. Journal of Computer and System Sciences, 73 (4):610–635, June 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001425.

Kolahi:2007:DPN

[2274] Solmaz Kolahi. Dependency-preserving normalization of relational and XML data. Journal of Computer and System Sciences, 73(4):636–647, June 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000006001437.

Cohen:2007:IAC

[2275] Sara Cohen and Yehoshua Sagiv. An incremental algorithm for computing ranked full disjunctions. Journal of Computer and System Sciences, 73 (4):648–668, June 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001449.

Foster:2007:ESD

[2276] J. Nathan Foster, Michael B. Greenwald, Christian Kirkegaard, Benjamin C. Pierce, and Alan Schmitt. Exploiting schemas in data synchronization. Journal of Computer and System Sciences, 73(4): 669–689, June 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000006001450.

Anonymous:2007:Cb

[2277] Anonymous. Contents. Journal of Computer and System Sciences, 73 (4):obc, June 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000372.

Anonymous:2007:EBd

[2278] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(4):ifc, June 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000335.

Jonsson:2007:MHC

[2279] Peter Jonsson and Andrei Krokhin. Maximum H-colourable subdigraphs and constraint optimization with arbitrary weights. Journal of Computer and System Sciences, 73(5): 691–702, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000050.

Afrati:2007:UVG

[2280] Foto N. Afrati, Chen Li, and Jeffrey D.

Ullman. Using views to generate efficient evaluation plans for queries. Journal of Computer and System Sciences, 73(5):703–724, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001486.

Anthony:2007:GEF

[2281] Martin Anthony. On the generalization error of fixed combinations of classifiers. Journal of Computer and System Sciences, 73(5): 725–734, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001462.

Glasser:2007:AMI

[2282] Christian Glaßer, Mitsunori Ogihara, A. Pavan, Alan L. Selman, and Liyu Zhang. Autoreducibility, mitoticity, and immunity. Journal of Computer and System Sciences, 73 (5):735–754, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001498.

Hajiaghayi:2007:SIL

[2283] MohammadTaghi Hajiaghayi and Naomi Nishimura. Subgraph isomorphism, log-bounded fragmentation, and graphs of (locally) bounded treewidth. Journal of Computer and System Sciences, 73(5):755– 768, August 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000049.

Downey:2007:USS

[2284] Rod Downey, Denis R. Hirschfeldt, and Geoff LaForte. Undecidability of the structure of the Solovay degrees of c.e. reals. Journal of Computer and System Sciences, 73 (5):769–787, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000037.

Ishihara:2007:ELB

[2285] Yasunori Ishihara, Toshiyuki Morita, Hiroyuki Seki, and Minoru Ito. An equational logic based approach to the security problem against inference attacks on object-oriented databases. Journal of Computer and System Sciences, 73(5):788– 817, August 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000025.

Boyar:2007:RWO

[2286] Joan Boyar, Lene M. Favrholdt, and Kim S. Larsen. The relative worstorder ratio applied to paging. Journal of Computer and System Sciences, 73(5):818–843, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000177.

Anonymous:2007:Cc

[2287] Anonymous.  Contents.  Journal

of Computer and System Sciences, 73(5):obc, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700044X.

Anonymous:2007:EBe

[2288] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(5):ifc, August 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000402.

Kuper:2007:RIL

[2289] Gabriel M. Kuper and Jianwen Su. A representation independent language for planar spatial databases with Euclidean distance. Journal of Computer and System Sciences, 73(6): 845–874, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001310.

Leonardi:2007:ATF

[2290] Stefano Leonardi and Danny Raz. Approximating total flow time on parallel machines. Journal of Computer and System Sciences, 73(6):

875–891, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001474.

Chen:2007:GCC

[2291] Jianer Chen, Iyad A. Kanj, Ljubomir Perković, Eric Sedgwick, and Ge Xia. Genus characterizes the complexity of certain graph problems: Some tight results. Journal of Computer and System Sciences, 73(6): 892–907, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001723.

Gramlich:2007:MNR

[2292] Gregor Gramlich and Georg Schnitger. Minimizing nfa’s and regular expressions. Journal of Computer and System Sciences, 73(6): 908–923, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000006001735.

Balcazar:2007:GDQ

[2293] José L. Balcázar, Jorge Castro, David Guijarro, Johannes Köbler, and Wolfgang Lindner. A general dimension for query learning. Journal of Computer and System Sciences, 73(6): 924–940, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000190.

Godfrey:2007:VDR

[2294] Parke Godfrey and Jarek Gryz. View disassembly: a rewrite that extracts portions of views. Journal of Computer and System Sciences, 73(6): 941–961, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000189.

Charles:2007:CLV

[2295] Denis Xavier Charles. Counting lattice vectors. Journal of Computer and System Sciences, 73(6): 962–972, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700030X.

Buhrman:2007:ICC

[2296] Harry Buhrman, Hartmut Klauck, Nikolai Vereshchagin, and Paul Vitányi. Individual communication complexity. Journal of Computer and System Sciences, 73(6): 973–985, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000311.

Kavitha:2007:LTA

[2297] T. Kavitha. Linear time algorithms for Abelian group isomorphism and related problems. Journal of Computer and System Sciences, 73(6): 986–996, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000293.

Ma:2007:NOM

[2298] Bin Ma, Lusheng Wang, and Ming Li. Near optimal multiple alignment within a band in polynomial time. Journal of Computer and System Sciences, 73(6):997–1011, September 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000281.

Belohlavek:2007:FFS

[2299] Radim Belohlavek, Jirí Dvorák, and Jan Outrata. Fast factorization by similarity in formal concept analysis of data with fuzzy attributes. Journal of Computer and System Sciences, 73(6): 1012–1022, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000384.

Anonymous:2007:Cd

[2300] Anonymous. Contents. Journal of Computer and System Sciences, 73 (6):obc, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000967.

Anonymous:2007:EBf

[2301] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(6):ifc, September 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000700092X.

Karp:2007:SIC

[2302] Richard Karp, Ming Li, Pavel Pevzner, and Ron Shamir. Special issue on computational molecular biology. Journal of Computer and System Sciences, 73(7):1023, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700027X.

Ma:2007:CSS

[2303] Bin Ma and Ming Li. On the complexity of the spaced seeds. Journal of Computer and System Sciences, 73(7): 1024–1034, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000268.

Farach-Colton:2007:OSS

[2304] Martin Farach-Colton, Gad M. Landau, S. Cenk Sahinalp, and Dekel Tsur. Optimal spaced seeds for faster approximate string matching. Journal of Computer and System Sciences, 73(7): 1035–1044, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000256.

Cui:2007:AAU

[2305] Yun Cui, Lusheng Wang, and Daming Zhu. A 1.75-approximation algorithm for unsigned translocation distance. Journal of Computer and System Sciences, 73(7):1045– 1059, November 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000244.

Brejova:2007:MPA

[2306] Brona Brejová, Daniel G. Brown, and Tomás Vinar. The most probable annotation problem in HMMs and its application to bioinformatics. Journal of Computer and System Sciences, 73(7): 1060–1077, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000232.

Moran:2007:EAC

[2307] Shlomo Moran and Sagi Snir. Efficient approximation of convex recolorings. Journal of Computer and System Sciences, 73(7):1078– 1089, November 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000220.

Hwang:2007:IPC

[2308] F. K. Hwang and F. H. Chang. The identification of positive clones in a general inhibitor model. Journal of Computer and System Sciences, 73(7): 1090–1094, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000219.

Landau:2007:TAL

[2309] Gad M. Landau, Eugene Myers, and Michal Ziv-Ukelson. Two algorithms for LCS consecutive suffix alignment. Journal of Computer and System Sciences, 73(7):

1095–1117, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000207.

Anonymous:2007:Ce

[2310] Anonymous. Contents. Journal of Computer and System Sciences, 73 (7):obc, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001158.

Anonymous:2007:EBg

[2311] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(7):ifc, November 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001110.

Sarbazi-Azad:2007:NBC

[2312] H. Sarbazi-Azad and L. M. Mackenzie.

Network-based computing. Journal of Computer and System Sciences, 73(8): 1119–1120, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000062.

Alzeidi:2007:NAM

[2313] N. Alzeidi, A. Khonsari, M. Ould-Khaoua, and L. Mackenzie. A new approach to model virtual channels in interconnection networks. Journal of Computer and System Sciences, 73(8): 1121–1130, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000074.

Safaei:2007:CDA

[2314] F. Safaei, A. Khonsari, M. Fathy, and M. Ould-Khaoua. Communication delay analysis of fault-tolerant pipelined circuit switching in torus. Journal of Computer and System Sciences, 73(8): 1131–1144, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000086.

Jiao:2007:EEW

[2315] Yu Jiao and Ali R. Hurson. Energyefficient wireless information retrieval. Journal of Computer and System Sciences, 73(8):1145–1163, December 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000098.

Moreno-Vozmediano:2007:PIR

[2316] Rafael Moreno-Vozmediano, Krishna Nadiminti, Srikumar Venugopal, Ana B. Alonso-Conde, Hussein Gibbins, and Rajkumar Buyya. Portfolio and investment risk analysis on global grids. Journal of Computer and System Sciences, 73(8):1164–1175, December 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000104.

Subrata:2007:ALT

[2317] Riky Subrata, Albert Y. Zomaya, and Bjorn Landfeldt. Artificial life techniques for load balancing in computational grids. Journal of Computer and System Sciences, 73(8): 1176–1190, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000116.

Lu:2007:PDL

[2318] Kai Lu, Riky Subrata, and Albert Y.

Zomaya. On the performance-driven load distribution for heterogeneous computational grids. Journal of Computer and System Sciences, 73(8): 1191–1206, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000128.

Jin:2007:PAP

[2319] Xiaolong Jin and Geyong Min. Performance analysis of priority scheduling mechanisms under heterogeneous network traffic. Journal of Computer and System Sciences, 73(8): 1207–1220, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700013X.

Goudarzi:2007:UCN

[2320] Maziar Goudarzi, Naser MohammadZadeh, and Shaahin Hessabi. Using on-chip networks to implement polymorphism in the co-design of object-oriented embedded systems. Journal of Computer and System Sciences, 73(8):1221–1231, December 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000141.

Xiao:2007:SPC

[2321] Wenjun Xiao and Behrooz Parhami. Structural properties of Cayley digraphs with applications to mesh and pruned torus interconnection networks. Journal of Computer and System Sciences, 73(8):1232– 1239, December 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000153.

Boukerche:2007:THA

[2322] Azzedine Boukerche, Raed A. Al-Shaikh, and Mirela Sechi Moretti Annoni Notare. Towards highly available and scalable high performance clusters. Journal of Computer and System Sciences, 73(8):1240– 1251, December 2007. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000165.

Anonymous:2007:Cf

[2323] Anonymous. Contents. Journal of Computer and System Sciences, 73 (8):obc, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007001353.

Anonymous:2007:EBh

[2324] Anonymous. Editorial Board. Journal of Computer and System Sciences, 73(8):ifc, December 2007. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001316.

Dasgupta:2008:SIL

[2325] Sanjoy Dasgupta. Special issue on learning theory. Journal of Computer and System Sciences, 74(1): 1, February 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000645.

Bisht:2008:LEA

[2326] Laurence Bisht, Nader H. Bshouty, and Lawrance Khoury. Learning with errors in answers to membership queries. Journal of Computer and System Sciences, 74(1):2– 15, February 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000566.

Alekhnovich:2008:CPL

[2327] Misha Alekhnovich, Mark Braverman, Vitaly Feldman, Adam R. Klivans, and Toniann Pitassi. The complexity of properly learning simple concept classes. Journal of Computer and System Sciences, 74(1):16– 34, February 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000578.

Klivans:2008:LIH

[2328] Adam R. Klivans and Rocco A. Servedio. Learning intersections of halfspaces with a margin. Journal of Computer and System Sciences, 74 (1):35–48, February 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000700058X.

Kleinberg:2008:UMM

[2329] Jon Kleinberg and Mark Sandler. Using mixture models for collaborative filtering. Journal of Computer and System Sciences, 74(1):49– 69, February 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000591.

Kempe:2008:DAS

[2330] David Kempe and Frank McSherry. A decentralized algorithm for spectral analysis. Journal of Computer and System Sciences, 74(1):70– 83, February 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000608.

McAllester:2008:CFD

[2331] David McAllester, Michael Collins, and Fernando Pereira. Case-factor diagrams for structured probabilistic modeling. Journal of Computer and System Sciences, 74(1):84– 96, February 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700061X.

Awerbuch:2008:OLO

[2332] Baruch Awerbuch and Robert Kleinberg. Online linear optimization and adaptive routing. Journal of Computer and System Sciences, 74(1): 97–114, February 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000621.

Kakade:2008:DCN

[2333] Sham M. Kakade and Dean P. Foster. Deterministic calibration and Nash equilibrium. Journal of Computer and System Sciences, 74(1): 115–130, February 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000633.

Anonymous:2008:Ca

[2334] Anonymous. Contents. Journal of Computer and System Sciences, 74 (1):obc, February 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001523.

Anonymous:2008:EBa

[2335] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(1):ifc, February 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001481.

Younas:2008:SIW

[2336] Muhammad Younas, Elhadi Shakshuki, and Kuo-Ming Chao. Special issue: Web and mobile information systems. Journal of Computer and System Sciences, 74(2): 131–132, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000554.

Rupert:2008:WCA

[2337] Maya Rupert, Amjad Rattrout, and Salima Hassas. The web from a complex adaptive systems perspective. Journal of Computer and System Sciences, 74(2):133– 145, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000451.

Chua:2008:EDI

[2338] H. N. Chua, S. D. Scott, and Y. W. Choi. Enabling device independent co-browsing through adaptive view points. Journal of Computer and System Sciences, 74(2): 146–157, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000463.

Eidsvik:2008:IRT

[2339] Arne Ketil Eidsvik, Randi Karlsen, Gordon Blair, and Paul Grace. Interfacing remote transaction services using UPnP. Journal of Computer and System Sciences, 74(2):

158–169, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000475.

Pardede:2008:XDU

[2340] Eric Pardede, J. Wenny Rahayu, and David Taniar. XML data update management in XML-enabled database. Journal of Computer and System Sciences, 74(2):170– 195, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000487.

Zang:2008:OEA

[2341] Binyu Zang, Yinsheng Li, Wei Xie, Zhuangjian Chen, Chen-Fang Tsai, and Christopher Laing. An ontological engineering approach for automating inspection and quarantine at airports. Journal of Computer and System Sciences, 74(2): 196–210, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000499.

Lin:2008:CCQ

[2342] Wei-Li Lin, Chi-Chun Lo, Kuo-Ming Chao, and Muhammad Younas. Consumer-centric QoS-aware selection of web services. Journal of Computer and System Sciences, 74 (2):211–231, March 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000505.

Awan:2008:PAM

[2343] Irfan Awan, Shakeel Ahmad, and Bashir Ahmad. Performance analysis of multimedia based web traffic with QoS constraints. Journal of Computer and System Sciences, 74 (2):232–242, March 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000517.

Anane:2008:SDC

[2344] Rachid Anane, Sukhvir Dhillon, and Behzad Bordbar. Stateless data concealment for distributed systems. Journal of Computer and System Sciences, 74(2):243–254, March 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000529.

Zhang:2008:TAA

[2345] Hao Lan Zhang, Clement H. C. Leung, and Gitesh K. Raikundalia. Topological analysis of AOCD-based agent networks and experimental results. Journal of Computer and System Sciences, 74(2):255–278, March 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000530.

Shakshuki:2008:DMA

[2346] Elhadi Shakshuki, Hsiang-Hwa Koo, Darcy Benoit, and Daniel Silver. A distributed multi-agent meeting scheduler. Journal of Computer and System Sciences, 74(2):279– 296, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000542.

Anonymous:2008:Cb

[2347] Anonymous. Contents. Journal of Computer and System Sciences, 74 (2):obc, March 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700164X.

Anonymous:2008:EBb

[2348] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(2):ifc, March 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001596.

Buhrman:2008:F

[2349] Harry Buhrman. Foreword. Journal of Computer and System Sciences, 74(3):297, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000906.

ODonnell:2008:EPP

[2350] Ryan O’Donnell and Rocco A. Servedio. Extremal properties of polynomial threshold functions. Journal of Computer and System Sciences, 74(3):298–312, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700089X.

Aaronson:2008:QCC

[2351] Scott Aaronson. Quantum certificate complexity. Journal of Computer and System Sciences, 74(3): 313–322, May 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000888.

Atserias:2008:CCR

[2352] Albert Atserias and Víctor Dalmau. A combinatorial characterization of resolution width. Journal of Computer and System Sciences, 74(3): 323–334, May 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000876.

Khot:2008:VCM

[2353] Subhash Khot and Oded Regev. Vertex cover might be hard to approximate to within 2 − ϵ. Journal of Computer and System Sciences, 74 (3):335–349, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000864.

Merkle:2008:CSS

[2354] Wolfgang Merkle. The complexity of stochastic sequences. Journal of Computer and System Sciences, 74 (3):350–357, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000852.

Fortnow:2008:PSD

[2355] Lance Fortnow, A. Pavan, and Samik Sengupta. Proving SAT does not have small circuits with an application to the two queries problem. Journal of Computer and System Sciences, 74(3):358–363, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000840.

Sen:2008:LBP

[2356] Pranab Sen and S. Venkatesh. Lower bounds for predecessor searching in the cell probe model. Journal of Computer and System Sciences, 74 (3):364–385, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000839.

Calabro:2008:CUS

[2357] Chris Calabro, Russell Impagliazzo, Valentine Kabanets, and Ramamohan Paturi. The complexity of unique k-SAT: an isolation lemma for k-CNFs. Journal of Computer and System Sciences, 74(3): 386–393, May 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000827.

Sieling:2008:MDT

[2358] Detlef Sieling. Minimization of decision trees is hard to approximate. Journal of Computer and System Sciences, 74(3):394–403, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000815.

Davidson:2008:EPX

[2359] Susan Davidson, Wenfei Fan, and Carmem Hara. Erratum to “Propagating XML constraints to relations” [J. Comput. System Sci. 73 (2007) 316–361]. Journal of Computer and System Sciences, 74(3): 404–405, May 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007001419. See

[2260].

Anonymous:2008:Cc

[2360] Anonymous. Contents. Journal of Computer and System Sciences, 74 (3):obc, May 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000081.

Anonymous:2008:EBc

[2361] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(3):ifc, May 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000044.

Anonymous:2008:OCS

[2362] Anonymous. Obituary for Carl Smith. Journal of Computer and System Sciences, 74(4):407, June 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000670.

Gasarch:2008:MIC

[2363] William Gasarch. Memorial issue for Carl Smith. Journal of Computer and System Sciences, 74 (4):408, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000803.

Carlucci:2008:NUS

[2364] Lorenzo Carlucci, John Case, Sanjay Jain, and Frank Stephan. Nonu-shaped vacillatory and team learning. Journal of Computer and System Sciences, 74(4):409–430, June 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000007000797.

Jain:2008:LLP

[2365] Sanjay Jain and Efim Kinber. Learning languages from positive data and negative counterexamples. Journal of Computer and System Sciences, 74(4):431–456, June 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000785.

Ambainis:2008:PTP

[2366] Andris Ambainis. Probabilistic and team PFIN-type learning: General properties. Journal of Computer and System Sciences, 74(4): 457–489, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000773.

Gasarch:2008:IAQ

[2367] William I. Gasarch and Andrew C. Y. Lee. Inferring answers to queries. Journal of Computer and System Sciences, 74(4):490–512, June 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700075X.

Kalyanasundaram:2008:TTM

[2368] Bala Kalyanasundaram and Mahendran Velauthapillai. Taming teams with mind changes. Journal of Computer and System Sciences, 74(4): 513–526, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000748.

Barzdins:2008:LBL

[2369] Janis Barzdins, Rusins Freivalds, and Carl H. Smith. Learning with belief levels. Journal of Computer and System Sciences, 74(4): 527–545, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000736.

Angluin:2008:LHG

[2370] Dana Angluin and Jiang Chen. Learning a hidden graph using O(log n) queries per edge. Journal of Computer and System Sciences, 74(4): 546–556, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000724.

Leskes:2008:VAN

[2371] Boaz Leskes and Leen Torenvliet.

The value of agreement a new boosting algorithm. Journal of Computer and System Sciences, 74(4): 557–586, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000712.

Kruskal:2008:OVP

[2372] Clyde P. Kruskal. The orchard visibility problem and some variants. Journal of Computer and System Sciences, 74(4):587–597, June 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000700.

Kruskal:2008:CNP

[2373] Clyde P. Kruskal. The chromatic number of the plane: The bounded case. Journal of Computer and System Sciences, 74(4): 598–627, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000694.

Gasarch:2008:FLF

[2374] William Gasarch, James Glenn, and Clyde P. Kruskal. Finding large 3free sets I: The small n case. Journal of Computer and System Sciences, 74(4):628–655, June 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000682.

Anonymous:2008:Cd

[2375] Anonymous. Contents. Journal of Computer and System Sciences, 74 (4):obc, June 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000299.

Anonymous:2008:EBd

[2376] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(4):ifc, June 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000800024X.

Xiang:2008:CPA

[2377] Tao Xiang, Kwok wo Wong, and Xiaofeng Liao. Cryptanalysis of a password authentication scheme over insecure networks. Journal of Computer and System Sciences, 74(5): 657–661, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007000657.

Kortsarz:2008:TAA

[2378] Guy Kortsarz and Zeev Nutov. Tight approximation algorithm for connectivity augmentation problems. Journal of Computer and System Sciences, 74(5):662–670, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000669.

Bonizzoni:2008:ACC

[2379] Paola Bonizzoni, Gianluca Della Vedova, Riccardo Dondi, and Tao Jiang. On the approximation of correlation clustering and consensus clustering. Journal of Computer and System Sciences, 74(5): 671–696, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000980.

Milo:2008:CSD

[2380] Tova Milo, Assaf Sagi, and Elad Verbin. Compact samples for data dissemination. Journal of Computer and System Sciences, 74(5): 697–720, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001213.

Cohen:2008:UTS

[2381] David Cohen, Peter Jeavons, and Marc Gyssens. A unified theory of structural tractability for constraint satisfaction problems. Journal of Computer and System Sciences, 74 (5):721–743, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001225.

Bender:2008:IBS

[2382] Michael A. Bender, Dongdong Ge, Simai He, Haodong Hu, Ron Y. Pinter, Steven Skiena, and Firas Swidan.

Improved bounds on sorting by lengthweighted reversals. Journal of Computer and System Sciences, 74(5): 744–774, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001298.

Arias:2008:CRR

[2383] Marta Arias, Lenore J. Cowen, and Kofi A. Laing. Compact roundtrip routing with topology-independent node names. Journal of Computer and System Sciences, 74(5): 775–795, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001407.

Crochemore:2008:MRS

[2384] Maxime Crochemore and Lucian Ilie. Maximal repetitions in strings. Journal of Computer and System Sciences, 74(5):796–807, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001420.

Gasieniec:2008:FPG

[2385] Leszek Gasieniec, Ralf Klasing, Russell Martin, Alfredo Navarra, and Xiaohui Zhang. Fast periodic graph exploration with constant memory. Journal of Computer and System Sciences, 74(5):808–822, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001432.

Guerraoui:2008:GCI

[2386] Rachid Guerraoui and Petr Kuznetsov.

The gap in circumventing the impossibility of consensus. Journal of Computer and System Sciences, 74 (5):823–830, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001456.

Nicolas:2008:HOS

[2387] François Nicolas and Eric Rivals.

Hardness of optimal spaced seed design. Journal of Computer and System Sciences, 74(5):831–849, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000007001444.

Moran:2008:CRS

[2388] Shlomo Moran and Sagi Snir. Convex recolorings of strings and trees: Definitions, hardness results and algorithms. Journal of Computer and System Sciences, 74(5):850– 869, August 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001468.

Li:2008:FAT

[2389] Qiang Li and Huimin Xie. Finite automata for testing composition-based reconstructibility of sequences. Journal of Computer and System Sciences, 74(5):870–874, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001572.

Bazgan:2008:ASB

[2390] Cristina Bazgan, Zsolt Tuza, and Daniel Vanderpooten. Approximation of satisfactory bisection problems. Journal of Computer and System Sciences, 74(5):875– 883, August 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000020.

Kutz:2008:SMH

[2391] Martin Kutz, Khaled Elbassioni, Irit Katriel, and Meena Mahajan. Simultaneous matchings: Hardness and approximation. Journal of Computer and System Sciences, 74(5): 884–897, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000159.

Jonsson:2008:CCA

[2392] Peter Jonsson and Andrei Krokhin. Computational complexity of auditing finite attributes in statistical databases. Journal of Computer and System Sciences, 74(5): 898–909, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000160.

McCann:2008:FTC

[2393] Mark McCann and Nicholas Pippenger. Fault tolerance in cellular automata at high fault rates. Journal of Computer and System Sciences, 74 (5):910–918, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000172.

Diaz:2008:EAC

[2394] Josep Díaz, Maria Serna, and Dimitrios M. Thilikos. Efficient algorithms for counting parameterized list H-colorings. Journal of Computer and System Sciences, 74(5): 919–937, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000184.

Anonymous:2008:Ce

[2395] Anonymous.  Contents.  Journal

of Computer and System Sciences, 74(5):obc, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000457.

Anonymous:2008:EBe

[2396] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(5):ifc, August 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000408.

Ould-Khaoua:2008:PAE

[2397] M. Ould-Khaoua and G. Min. Performance analysis and evaluation of parallel, cluster, and grid computing systems. Journal of Computer and System Sciences, 74(6): 939–941, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001092.

Filippopoulos:2008:TCP

[2398] D. Filippopoulos and H. Karatza. A two-class parallel system with general service times of the parallel class. Journal of Computer and System Sciences, 74(6):942–964, September 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007000992.

Saini:2008:PES

[2399] Subhash Saini, Robert Ciotti, Brian T. N. Gunney, Thomas E. Spelce, Alice Koniges, Don Dossa, Panagiotis Adamidis, Rolf Rabenseifner, Sunil R. Tiyyagura, and Matthias Mueller. Performance evaluation of supercomputers using HPCC and IMB benchmarks. Journal of Computer and System Sciences, 74(6): 965–982, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001006.

Schaeli:2008:SAP

[2400] Basile Schaeli, Sebastian Gerlach, and Roger D. Hersch. A simulator for adaptive parallel applications. Journal of Computer and System Sciences, 74(6): 983–999, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001018.

Kiasari:2008:APC

[2401] A. E. Kiasari and H. Sarbazi-Azad.

Analytic performance comparison of hypercubes and star graphs with implementation constraints. Journal of Computer and System Sciences, 74(6): 1000–1012, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000700102X.

Bradley:2008:ADI

[2402] Jeremy T. Bradley, Stephen T.

Gilmore, and Jane Hillston. Analysing distributed Internet worm attacks using continuous state-space approximation of process algebra models. Journal of Computer and System Sciences, 74(6):1013–1032, September 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001031.

Alzeidi:2008:NGM

[2403] N. Alzeidi, M. Ould-Khaoua, and A. Khonsari. A new general method to compute virtual channels occupancy probabilities in wormhole networks. Journal of Computer and System Sciences, 74(6):1033–1042, September 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001043.

Boukerche:2008:ADG

[2404] Azzedine Boukerche, Yunfeng Gu, and Regina B. Araujo. An adaptive dynamic grid-based approach to DDM for large-scale distributed simulation systems. Journal of Computer and System Sciences, 74(6): 1043–1054, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001055.

Jin:2008:HFD

[2405] Xiaolong Jin and Geyong Min. A heuristic flow-decomposition approach for generalized processor sharing under self-similar traffic. Journal of Computer and System Sciences, 74(6): 1055–1066, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001067.

Thomas:2008:CJA

[2406] Nigel Thomas. Comparing job allocation schemes where service demand is unknown. Journal of Computer and System Sciences, 74(6): 1067–1081, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001079.

Steffenel:2008:FAC

[2407] Luiz Angelo Steffenel and Grégory Mounié. A framework for adaptive collective communications for heterogeneous hierarchical computing systems. Journal of Computer and System Sciences, 74(6):1082– 1093, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001080.

Anonymous:2008:EBf

[2408] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(6):ifc, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000548.

Anonymous:2008:EBg

[2409] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(6):obc, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000597.

Nakamura:2008:LRR

[2410] Katsuhiko Nakamura. Languages not recognizable in real time by onedimensional cellular automata. Journal of Computer and System Sciences, 74(7):1095–1102, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000007000979.

See erratum [2469].

Creignou:2008:SIB

[2411] Nadia Creignou, Phokion Kolaitis, and Bruno Zanuttini. Structure identification of Boolean relations and plain bases for co-clones. Journal of Computer and System Sciences, 74(7): 1103–1115, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000196.

Domaratzki:2008:LBT

[2412] Michael Domaratzki and Kai Salomaa. Lower bounds for the transition complexity of NFAs. Journal of Computer and System Sciences, 74(7): 1116–1130, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000214.

Grigorian:2008:EPM

[2413] Haik Grigorian and Samvel Shoukourian. The equivalence problem of multidimensional multitape automata. Journal of Computer and System Sciences, 74(7):1131– 1138, November 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000202.

Gangal:2008:PCS

[2414] Devdatta Gangal and Abhiram Ranade. Precedence constrained scheduling in (2 − 7/(3p + 1))· optimal. Journal of Computer and System Sciences, 74(7): 1139–1146, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000226.

Cohen:2008:GAM

[2415] Sara Cohen, Benny Kimelfeld, and Yehoshua Sagiv. Generating all maximal induced subgraphs for hereditary and connected-hereditary graph properties. Journal of Computer and System Sciences, 74(7):1147– 1159, November 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000512.

Yang:2008:TFM

[2416] Guomin Yang, Duncan S. Wong, Huaxiong Wang, and Xiaotie Deng. Twofactor mutual authentication based on smart cards and passwords. Journal of Computer and System Sciences, 74(7): 1160–1172, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000482.

Glasser:2008:CUD

[2417] Christian Glaßer, Alan L. Selman, Stephen Travers, and Klaus W. Wagner. The complexity of unions of disjoint sets. Journal of Computer and System Sciences, 74(7): 1173–1187, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000494.

Chen:2008:IAF

[2418] Jianer Chen, Fedor V. Fomin, Yang Liu, Songjian Lu, and Yngve Villanger. Improved algorithms for feedback vertex set problems. Journal of Computer and System Sciences, 74(7): 1188–1198, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000500.

Gairing:2008:NED

[2419] Martin Gairing, Thomas Lücking, Marios Mavronicolas, Burkhard Monien, and Manuel Rode. Nash equilibria in discrete routing games with convex latency functions. Journal of Computer and System Sciences, 74(7): 1199–1225, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000524.

Anonymous:2008:Cf

[2420] Anonymous. Contents. Journal of Computer and System Sciences, 74 (7):obc, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000822.

Anonymous:2008:EBh

[2421] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(7):ifc, November 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000780.

Auer:2008:GEI

[2422] Peter Auer and Philip M. Long. Guest Editors’ introduction: Special issue on learning theory. Journal of Computer and System Sciences, 74 (8):1227, December 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007001286.

Kalnishkan:2008:WAA

[2423] Yuri Kalnishkan and Michael V. Vyugin. The weak aggregating algorithm and weak mixability. Journal of Computer and System Sciences, 74(8): 1228–1244, December 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001237.

Balcan:2008:RMD

[2424] Maria-Florina Balcan, Avrim Blum, Jason D. Hartline, and Yishay Mansour. Reducing mechanism design to algorithm design via machine learning. Journal of Computer and System Sciences, 74(8): 1245–1270, December 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001249.

Awerbuch:2008:CCL

[2425] Baruch Awerbuch and Robert Kleinberg. Competitive collaborative learning. Journal of Computer and System Sciences, 74(8):1271– 1288, December 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000007001250.

Belkin:2008:TTF

[2426] Mikhail Belkin and Partha Niyogi.

Towards a theoretical foundation for Laplacian-based manifold methods. Journal of Computer and System Sciences, 74(8):1289–1308, December 2008. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000007001274.

Strehl:2008:AMB

[2427] Alexander L. Strehl and Michael L. Littman. An analysis of model-based interval estimation for Markov decision processes. Journal of Computer and System Sciences, 74(8): 1309–1331, December 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000767.

Anonymous:2008:Cg

[2428] Anonymous. Contents. Journal of Computer and System Sciences, 74 (8):obc, December 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000937.

Anonymous:2008:EBi

[2429] Anonymous. Editorial Board. Journal of Computer and System Sciences, 74(8):ifc, December 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000883.

Hellerstein:2009:SIL

[2430] Lisa Hellerstein and Hans U. Simon. Special issue: Learning theory 2006. Journal of Computer and System Sciences, 75(1):1, January 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000640.

Klivans:2009:CHL

[2431] Adam R. Klivans and Alexander A. Sherstov. Cryptographic hardness for learning intersections of halfspaces. Journal of Computer and System Sciences, 75(1): 2–12, January 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000706.

Feldman:2009:HAT

[2432] Vitaly Feldman. Hardness of approximate two-level logic minimization and PAC learning with membership queries. Journal of Computer and System Sciences, 75(1): 13–26, January 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000800069X.

Fortnow:2009:ELA

[2433] Lance Fortnow and Adam R. Klivans. Efficient learning algorithms yield circuit lower bounds. Journal of Computer and System Sciences, 75 (1):27–36, January 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000688.

Rubinstein:2009:SOI

[2434] Benjamin I. P. Rubinstein, Peter L. Bartlett, and J. Hyam Rubinstein. Shifting: One-inclusion mistake bounds and sample compression. Journal of Computer and System Sciences, 75(1):37–59, January 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000676. See

corrigendum [2501].

Angluin:2009:LCI

[2435] Dana Angluin, James Aspnes, Jiang Chen, and Yinghua Wu. Learning a circuit by injecting values. Journal of Computer and System Sciences, 75(1):60–77, January 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000664.

Balcan:2009:AAL

[2436] Maria-Florina Balcan, Alina Beygelzimer, and John Langford. Agnostic active learning. Journal of Computer and System Sciences, 75(1): 78–89, January 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000652.

Anonymous:2009:Ca

[2437] Anonymous.  Contents.  Journal

of Computer and System Sciences, 75(1):obc, January 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008001074.

Anonymous:2009:EBa

[2438] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(1):ifc, January 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008001025.

Katz:2009:ECR

[2439] Jonathan Katz and Chiu-Yuen Koo.

On expected constant-round protocols for Byzantine agreement. Journal of Computer and System Sciences, 75 (2):91–112, February 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000718.

Kari:2009:PBW

[2440] Lila Kari and Shinnosuke Seki.

On pseudoknot-bordered words and their properties. Journal of Computer and System Sciences, 75(2): 113–121, February 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000800072X.

Epstein:2009:BBM

[2441] Leah Epstein and Asaf Levin. Better bounds for minimizing SONET ADMs. Journal of Computer and System Sciences, 75(2):122– 136, February 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000731.

Mahajan:2009:PAB

[2442] Meena Mahajan, Venkatesh Raman, and Somnath Sikdar. Parameterizing above or below guaranteed values. Journal of Computer and System Sciences, 75(2):

137–153, February 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000743.

Anonymous:2009:Cb

[2443] Anonymous. Contents. Journal of Computer and System Sciences, 75 (2):obc, February 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008001232.

Anonymous:2009:EBb

[2444] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(2):ifc, February 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008001189.

Linhart:2009:FPM

[2445] Chaim Linhart and Ron Shamir. Faster pattern matching with character classes using prime number encoding. Journal of Computer and System Sciences, 75(3): 155–162, May 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000008000755.

Brandt:2009:SCP

[2446] Felix Brandt, Felix Fischer, and Markus Holzer. Symmetries and the complexity of pure Nash equilibrium. Journal of Computer and System Sciences, 75(3):163– 177, May 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000846.

Martin:2009:IOF

[2447] Cristian Martín, Mikel Larrea, and Ernesto Jiménez. Implementing the Omega failure detector in the crashrecovery failure model. Journal of Computer and System Sciences, 75 (3):178–189, May 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000800086X.

Auletta:2009:PVO

[2448] Vincenzo Auletta, Roberto De Prisco, Paolo Penna, and Giuseppe Persiano. The power of verification for oneparameter agents. Journal of Computer and System Sciences, 75(3): 190–211, May 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008000858.

Anonymous:2009:Cc

[2449] Anonymous. Contents. Journal of Computer and System Sciences, 75 (3):obc, May 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000087.

Anonymous:2009:EBc

[2450] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(3):ifc, May 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000038.

Clementi:2009:BDR

[2451] Andrea E. F. Clementi, Angelo Monti, Francesco Pasquale, and Riccardo Silvestri. Broadcasting in dynamic radio networks. Journal of Computer and System Sciences, 75(4): 213–230, June 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008001153.

Bodlaender:2009:DAC

[2452] Hans L. Bodlaender, Michael R. Fellows, and Dimitrios M. Thilikos. Derivation of algorithms for cutwidth and related graph layout parameters. Journal of Computer and System Sciences, 75(4): 231–244, June 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008001001.

Allender:2009:CSP

[2453] Eric Allender, Michael Bauland, Neil Immerman, Henning Schnoor, and Heribert Vollmer. The complexity of satisfiability problems: Refining Schaefer’s theorem. Journal of Computer and System Sciences, 75 (4):245–254, June 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008001141.

Nishimura:2009:AQF

[2454] Harumichi Nishimura and Tomoyuki Yamakami. An application of quantum finite automata to interactive proof systems. Journal of Computer and System Sciences, 75(4): 255–269, June 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000008001165.

Anonymous:2009:Cd

[2455] Anonymous. Contents. Journal of Computer and System Sciences, 75 (4):obc, June 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000221.

Anonymous:2009:EBd

[2456] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(4):ifc, June 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000178.

Engelfriet:2009:DET

[2457] Joost Engelfriet, Sebastian Maneth, and Helmut Seidl. Deciding equivalence of top-down XML transformations in polynomial time. Journal of Computer and System Sciences, 75 (5):271–286, August 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000117.

Ashley:2009:AFC

[2458] Mary Ashley, Tanya Berger-Wolf, Piotr Berman, Wanpracha Chaovalitwongse, Bhaskar DasGupta, and Ming-Yang Kao. On approximating four covering and packing problems. Journal of Computer and System Sciences, 75 (5):287–302, August 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000129.

Sakr:2009:XCT

[2459] Sherif Sakr. XML compression techniques: a survey and comparison. Journal of Computer and System Sciences, 75(5):303–322, August 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000142.

Anonymous:2009:Ce

[2460] Anonymous.  Contents.  Journal

of Computer and System Sciences, 75(5):obc, August 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000373.

Anonymous:2009:EBe

[2461] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(5):ifc, August 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000324.

Bshouty:2009:UDD

[2462] Nader H. Bshouty, Yi Li, and Philip M.

Long. Using the doubling dimension to analyze the generalization of learning algorithms. Journal of Computer and System Sciences, 75(6): 323–335, September 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000130.

Cautis:2009:RAX

[2463] Bogdan Cautis, Serge Abiteboul, and Tova Milo. Reasoning about XML update constraints. Journal of Computer and System Sciences, 75(6): 336–358, September 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000154.

Amir:2009:PMA

[2464] Amihood Amir, Yonatan Aumann, Gary Benson, Avivit Levy, Ohad Lipsky, Ely Porat, Steven Skiena, and Uzi Vishne. Pattern matching with address errors: Rearrangement distances. Journal of Computer and System Sciences, 75(6): 359–370, September 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000257.

Anonymous:2009:Cf

[2465] Anonymous. Contents. Journal of Computer and System Sciences, 75 (6):obc, September 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000543.

Anonymous:2009:EBf

[2466] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(6):ifc, September 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000900049X.

Chen:2009:IVF

[2467] Ting-Yu Chen and Jih-Chang Wang. Interval-valued fuzzy permutation method and experimental analysis on cardinal and ordinal evaluations. Journal of Computer and System Sciences, 75(7):371– 387, November 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000269.

Devi:2009:ICB

[2468] UmaMaheswari C. Devi and James H. Anderson. Improved conditions for bounded tardiness under EPDF Pfair multiprocessor scheduling. Journal of Computer and System Sciences, 75(7): 388–420, November 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000270.

Nakamura:2009:ELR

[2469] Katsuhiko Nakamura. Erratum to “Languages not recognizable in real time by one-dimensional cellular automata” [J. Comput. System Sci. 74 (7) (2008) 1095–1102]. Journal of Computer and System Sciences, 75 (7):421, November 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000385. See

[2410].

Anonymous:2009:Cg

[2470] Anonymous. Contents. Journal of Computer and System Sciences, 75 (7):obc, November 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000671.

Anonymous:2009:EBg

[2471] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(7):ifc, November 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000634.

Bodlaender:2009:PPK

[2472] Hans L. Bodlaender, Rodney G.

Downey, Michael R. Fellows, and Danny Hermelin. On problems without polynomial kernels. Journal of Computer and System Sciences, 75(8): 423–434, December 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000282.

Razgon:2009:ASF

[2473] Igor Razgon and Barry O’Sullivan.

Almost 2-SAT is fixed-parameter tractable. Journal of Computer and System Sciences, 75(8):435– 450, December 2009. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000294.

Lin:2009:EAI

[2474] Tzu-Chin Lin, Chung-Chin Kuo, Yong-Hsiang Hsieh, and Biing-Feng Wang. Efficient algorithms for the inverse sorting problem with bound constraints under the l_{∞}-norm and the Hamming distance. Journal of Computer and System Sciences, 75(8): 451–464, December 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000397.

Ramanan:2009:WCO

[2475] Prakash Ramanan. Worst-case optimal algorithm for XPath evaluation over XML streams. Journal of Computer and System Sciences, 75(8): 465–485, December 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000555.

Anonymous:2009:Ch

[2476] Anonymous. Contents. Journal of Computer and System Sciences, 75 (8):obc, December 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000853.

Anonymous:2009:EBh

[2477] Anonymous. Editorial Board. Journal of Computer and System Sciences, 75(8):ifc, December 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000804.

Belohlavek:2010:EF

[2478] Radim Belohlavek and Rudolf Kruse.

Editor’s foreword. Journal of Computer and System Sciences, 76(1): 1–2, February 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000403.

Belohlavek:2010:DOF

[2479] Radim Belohlavek and Vilem Vychodil. Discovery of optimal factors in binary data via a novel method of matrix decomposition. Journal of Computer and System Sciences, 76 (1):3–20, February 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000415.

Borgelt:2010:CIA

[2480] Christian Borgelt. A conditional independence algorithm for learning undirected graphical models. Journal of Computer and System Sciences, 76 (1):21–33, February 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000427.

Hajek:2010:GMM

[2481] Petr Hájek, Martin Holena, and Jan Rauch. The GUHA method and its meaning for data mining. Journal of Computer and System Sciences, 76 (1):34–48, February 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000439.

Hullermeier:2010:PAR

[2482] Eyke Hüllermeier and Johannes Fürnkranz. On predictive accuracy and risk minimization in pairwise label ranking. Journal of Computer and System Sciences, 76(1):49– 62, February 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000440.

Liu:2010:FCF

[2483] Guoqing Liu and Vladik Kreinovich. Fast convolution and Fast Fourier Transform under interval and fuzzy uncertainty. Journal of Computer and System Sciences, 76(1):63– 76, February 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000452.

Steinbrecher:2010:VFF

[2484] Matthias Steinbrecher and Rudolf Kruse. Visualizing and fuzzy filtering for discovering temporal trajectories of association rules. Journal of Computer and System Sciences, 76 (1):77–87, February 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000464.

Gouda:2010:PEA

[2485] Karam Gouda, Mosab Hassaan, and Mohammed J. Zaki. Prism: an effective approach for frequent sequence mining via prime-block encoding. Journal of Computer and System Sciences, 76(1):88–102, February 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000476.

Anonymous:2010:Ca

[2486] Anonymous. Contents. Journal of Computer and System Sciences, 76 (1):obc, February 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000900110X.

Anonymous:2010:EBa

[2487] Anonymous. Editorial Board. Journal of Computer and System Sciences, 76(1):ifc, February 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001068.

Samer:2010:CSB

[2488] Marko Samer and Stefan Szeider.

Constraint satisfaction with bounded treewidth revisited. Journal of Computer and System Sciences, 76(2): 103–114, March 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000300.

Clifford:2010:PMD

[2489] Raphaël Clifford, Klim Efremenko, Ely Porat, and Amir Rothschild. Pattern matching with don’t cares and few errors. Journal of Computer and System Sciences, 76(2): 115–124, March 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000567.

Verbeek:2010:RRR

[2490] H. M. W. Verbeek, M. T. Wynn, W. M. P. van der Aalst, and A. H. M. ter Hofstede. Reduction rules for reset/inhibitor nets. Journal of Computer and System Sciences, 76(2): 125–143, March 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000579.

Daligault:2010:FAK

[2491] Jean Daligault, Gregory Gutin, Eun Jung Kim, and Anders Yeo. FPT algorithms and kernels for the directed k-leaf problem. Journal of Computer and System Sciences, 76 (2):144–152, March 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000592.

Anonymous:2010:Cb

[2492] Anonymous. Contents. Journal of Computer and System Sciences, 76 (2):obc, March 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001160.

Anonymous:2010:EBb

[2493] Anonymous. Editorial Board. Journal of Computer and System Sciences, 76(2):ifc, March 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001123.

Anonymous:2010:OOE

[2494] Anonymous. Obituary for our editor Arthur W. Burks. Journal of Computer and System Sciences, 76 (3–4):153, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001019.

Bjorner:2010:CDC

[2495] Nikolaj Bjørner, Andreas Blass, and Yuri Gurevich. Contentdependent chunking for differential compression, the local maximum approach. Journal of Computer and System Sciences, 76(3–4):154– 203, May/June 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000580.

Dom:2010:AFP

[2496] Michael Dom, Jiong Guo, and Rolf Niedermeier. Approximation and fixedparameter algorithms for consecutive ones submatrix problems. Journal of Computer and System Sciences, 76(3– 4):204–221, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000609.

Fraigniaud:2010:CAA

[2497] Pierre Fraigniaud, David Ilcinkas, and Andrzej Pelc. Communication algorithms with advice. Journal of Computer and System Sciences, 76(3–4): 222–232, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000610.

Jain:2010:NOL

[2498] Sanjay Jain and Frank Stephan. Numberings optimal for learning. Journal of Computer and System Sciences, 76(3– 4):233–250, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000749.

Okhotin:2010:DPL

[2499] Alexander Okhotin. Decision problems for language equations. Journal of Computer and System Sciences, 76(3– 4):251–266, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000750.

Dyer:2010:ATB

[2500] Martin Dyer, Leslie Ann Goldberg, and Mark Jerrum. An approximation trichotomy for Boolean #CSP. Journal of Computer and System Sciences, 76(3– 4):267–277, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000762.

Rubinstein:2010:CSO

[2501] Benjamin I. P. Rubinstein, Peter L.

Bartlett, and J. Hyam Rubinstein. Corrigendum to “Shifting: One-inclusion mistake bounds and sample compression” [J. Comput. System Sci. 75 (1) (2009) 37–59]. Journal of Computer and System Sciences, 76(3–4): 278–280, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000280. See

[2434].

Anonymous:2010:Cc

[2502] Anonymous. Contents. Journal of Computer and System Sciences, 76 (3–4):obc, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000383.

Anonymous:2010:EBc

[2503] Anonymous. Editorial Board. Journal of Computer and System Sciences, 76(3–4):ifc, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000346.

Hodges:2010:EFJ

[2504] Wilfrid Hodges and Ruy de Queiroz.

Editors’ foreword for JCSS WoL-LIC 2008. Journal of Computer and System Sciences, 76(5):281– 282, August 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000993.

Calves:2010:MAE

[2505] Christophe Calvès and Maribel Fernández. Matching and alpha-equivalence check for nominal terms. Journal of Computer and System Sciences, 76(5): 283–301, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000920.

Danvy:2010:IDS

[2506] Olivier Danvy and Jacob Johannsen. Inter-deriving semantic artifacts for object-oriented programming. Journal of Computer and System Sciences, 76 (5):302–323, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000932.

Dawar:2010:HPQ

[2507] Anuj Dawar. Homomorphism preservation on quasi-wide classes. Journal of Computer and System Sciences, 76 (5):324–332, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000944. See

corrigendum [3859].

Figueira:2010:FSI

[2508] Santiago Figueira, Daniel Gorín, and Rafael Grimson. On the formal semantics of IF-like logics. Journal of Computer and System Sciences, 76 (5):333–346, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000956.

Koiran:2010:ALB

[2509] Pascal Koiran, Jürgen Landes, Natacha Portier, and Penghui Yao. Adversary lower bounds for nonadaptive quantum algorithms. Journal of Computer and System Sciences, 76 (5):347–355, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009000968.

Mezhirov:2010:ARS

[2510] Ilya Mezhirov and Nikolay Vereshchagin. On abstract resource semantics and computability logic. Journal of Computer and System Sciences, 76 (5):356–372, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200000900097X.

Robaldo:2010:IIM

[2511] Livio Robaldo. Interpretation and inference with maximal referential terms. Journal of Computer and System Sciences, 76(5):373– 388, August 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000981.

Anonymous:2010:Cd

[2512] Anonymous.  Contents.  Journal

of Computer and System Sciences, 76(5):obc, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000644.

Anonymous:2010:EBd

[2513] Anonymous. Editorial Board. Journal of Computer and System Sciences, 76(5):ifc, August 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000607.

Chen:2010:SIS

[2514] Jinjun Chen and Lizhe Wang. Special issue: Scientific workflow 2009. Journal of Computer and System Sciences, 76(6):389, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009001172.

Yang:2010:IFA

[2515] Ping Yang, Shiyong Lu, Mikhail I. Gofman, and Zijiang Yang. Information flow analysis of scientific workflows. Journal of Computer and System Sciences, 76(6): 390–402, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009001184.

Tolosana-Calasanz:2010:UAC

[2516] Rafael Tolosana-Calasanz, José Ángel Bañares, Pedro Álvarez, Joaquín Ezpeleta, and Omer Rana. An uncoordinated asynchronous checkpointing model for hierarchical scientific workflows. Journal of Computer and System Sciences, 76(6): 403–415, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001196.

Dou:2010:CSA

[2517] Wanchun Dou, J. Leon Zhao, and Shaokun Fan. A collaborative scheduling approach for service-driven scientific workflow execution. Journal of Computer and System Sciences, 76(6): 416–427, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001202.

Callaghan:2010:SWB

[2518] Scott Callaghan, Ewa Deelman, Dan Gunter, Gideon Juve, Philip Maechling, Christopher Brooks, Karan Vahi, Kevin Milner, Robert Graves, Edward Field, David Okaya, and Thomas Jordan. Scaling up workflowbased applications. Journal of Computer and System Sciences, 76(6): 428–446, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001214.

Zinn:2010:PXD

[2519] Daniel Zinn, Shawn Bowers, Sven Köhler, and Bertram Ludäscher. Parallelizing XML data-streaming workflows via MapReduce. Journal of Computer and System Sciences, 76(6): 447–463, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009001226.

Chen:2010:LTC

[2520] Jinjun Chen and Yun Yang. Localising temporal constraints in scientific workflows. Journal of Computer and System Sciences, 76(6): 464–474, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009001238.

Ma:2010:ESS

[2521] Jiangang Ma, Jinli Cao, and Yanchun Zhang. Efficiently supporting secure and reliable collaboration in scientific workflows. Journal of Computer and System Sciences, 76(6): 475–489, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200000900124X.

Sroka:2010:FST

[2522] Jacek Sroka, Jan Hidders, Paolo Missier, and Carole Goble. A formal semantics for the Taverna 2 workflow model. Journal of Computer and System Sciences, 76(6): 490–508, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001251.

Anonymous:2010:Ce

[2523] Anonymous. Contents. Journal of Computer and System Sciences, 76 (6):obc, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000723.

Anonymous:2010:EBe

[2524] Anonymous. Editorial Board. Journal of Computer and System Sciences, 76(6):ifc, September 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000681.

Anonymous:2010:IWW

[2525] Anonymous. The 2nd International Workshop on Workflow Management and Application in Grid Environments & The 3rd International Workshop on Workflow Management and Applications in Grid Environments. Journal of Computer and System Sciences, 76 (6):??, September 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Arvind:2010:ICT

[2526] V. Arvind, Bireswar Das, and Partha Mukhopadhyay. Isomorphism and canonization of tournaments and hypertournaments. Journal of Computer and System Sciences, 76(7): 509–523, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000774.

Abu-Khzam:2010:KAH

[2527] Faisal N. Abu-Khzam. A kernelization algorithm for d-hitting set. Journal of Computer and System Sciences, 76 (7):524–531, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000786.

Wehler:2010:SPB

[2528] Joachim Wehler. Simplified proof of the blocking theorem for freechoice Petri nets. Journal of Computer and System Sciences, 76(7): 532–537, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000907.

Kuijpers:2010:TDD

[2529] Bart Kuijpers and Walied Othman. Trajectory databases: Data models, uncertainty and complete query languages. Journal of Computer and System Sciences, 76(7): 538–560, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000009000919.

Hemaspaandra:2010:GMS

[2530] Edith Hemaspaandra, Henning Schnoor, and Ilka Schnoor. Generalized modal satisfiability. Journal of Computer and System Sciences, 76(7): 561–578, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001007.

Chen:2010:CGS

[2531] Hubie Chen and Omer Giménez.

Causal graphs and structurally restricted planning. Journal of Computer and System Sciences, 76(7): 579–592, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001020.

Fernau:2010:CTC

[2532] Henning Fernau, Michael Kaufmann, and Mathias Poths. Comparing trees via crossing minimization. Journal of Computer and System Sciences, 76(7): 593–608, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001032.

Ciric:2010:FRE

[2533] Miroslav Ćirić, Aleksandar Stamenković, Jelena Ignjatović, and Tatjana Petković. Fuzzy relation equations and reduction of fuzzy automata. Journal of Computer and System Sciences, 76(7):609– 633, November 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001044.

Hermann:2010:CCP

[2534] Miki Hermann and Reinhard Pichler. Counting complexity of propositional abduction. Journal of Computer and System Sciences, 76(7): 634–649, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000009001263.

Cohen:2010:AFV

[2535] Nathann Cohen, Fedor V. Fomin, Gregory Gutin, Eun Jung Kim, Saket Saurabh, and Anders Yeo. Algorithm for finding k-vertex out-trees and its application to k-internal outbranching problem. Journal of Computer and System Sciences, 76(7): 650–662, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000024.

Fisteus:2010:HCN

[2536] Jesus Arias Fisteus, Norberto Fernández García, Luis Sánchez Fernández, and Carlos Delgado Kloos. Hashing and canonicalizing notation 3 graphs. Journal of Computer and System Sciences, 76(7):663– 685, November 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000048.

Muthukrishnan:2010:TRG

[2537] S. Muthukrishnan and Gopal Pandurangan. Thresholding random geometric graph properties motivated by ad hoc sensor networks. Journal of Computer and System Sciences, 76(7): 686–696, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000036.

Anonymous:2010:Cf

[2538] Anonymous. Contents. Journal of Computer and System Sciences, 76 (7):obc, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000802.

Anonymous:2010:EBf

[2539] Anonymous. Editorial Board. Journal of Computer and System Sciences, 76(7):ifc, November 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000760.

Yu:2010:IAF

[2540] Chih-Chiang Yu, Chien-Hsin Lin, and Biing-Feng Wang. Improved algorithms for finding length-bounded two vertexdisjoint paths in a planar graph and minmax k vertex-disjoint paths in a directed acyclic graph. Journal of Computer and System Sciences, 76(8): 697–708, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000140.

Eisenbrand:2010:CFL

[2541] Friedrich Eisenbrand, Fabrizio Grandoni, Thomas Rothvoß, and Guido Schäfer. Connected facility location via random facility sampling and core detouring. Journal of Computer and System Sciences, 76(8):709– 726, December 2010. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000152.

Fellows:2010:PCS

[2542] Michael R. Fellows, Jiong Guo, and Iyad Kanj. The parameterized complexity of some minimum label problems. Journal of Computer and System Sciences, 76(8): 727–740, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000267.

Kim:2010:DCP

[2543] Saehwa Kim. Dual ceiling protocol for real-time synchronization under preemption threshold scheduling. Journal of Computer and System Sciences, 76 (8):741–750, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000279.

Tsur:2010:SHF

[2544] Dekel Tsur. Sequencing by hybridization in few rounds. Journal of Computer and System Sciences, 76(8): 751–758, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000292.

Dekel:2010:ICR

[2545] Ofer Dekel, Felix Fischer, and Ariel D. Procaccia. Incentive compatible regression learning. Journal of Computer and System Sciences, 76(8): 759–777, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000309.

Tan:2010:PAD

[2546] Tony Tan. On pebble automata for data languages with decidable emptiness problem. Journal of Computer and System Sciences, 76(8): 778–791, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000310.

Glasser:2010:SEI

[2547] Christian Glaßer. Space-efficient informational redundancy. Journal of Computer and System Sciences, 76(8): 792–811, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000322.

Betzler:2010:TDP

[2548] Nadja Betzler and Britta Dorn. Towards a dichotomy for the Possible Winner problem in elections based on scoring rules. Journal of Computer and System Sciences, 76(8): 812–836, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000449.

Masopust:2010:SRC

[2549] Tomás Masopust. Simple restriction in context-free rewriting. Journal of Computer and System Sciences, 76(8): 837–846, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000437.

Chen:2010:CSS

[2550] Hubie Chen and Martin Grohe. Constraint satisfaction with succinctly specified relations. Journal of Computer and System Sciences, 76(8): 847–860, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000450.

Salomaa:2010:SBP

[2551] Arto Salomaa. Subword balance, position indices and power sums. Journal of Computer and System Sciences, 76 (8):861–871, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000541.

Gutin:2010:BPA

[2552] Gregory Gutin, Eun Jung Kim, Matthias Mnich, and Anders Yeo. Betweenness parameterized above tight lower bound. Journal of Computer and System Sciences, 76(8): 872–878, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000553.

Gamarnik:2010:DAA

[2553] David Gamarnik and Dmitriy Katz.

A deterministic approximation algorithm for computing the permanent of a 0,1 matrix. Journal of Computer and System Sciences, 76(8): 879–883, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000565.

Dolev:2010:WCM

[2554] Shlomi Dolev, Ronen I. Kat, and Elad M. Schiller. When consensus meets self-stabilization. Journal of Computer and System Sciences, 76(8): 884–900, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000838.

Anonymous:2010:Cg

[2555] Anonymous. Contents. Journal of Computer and System Sciences, 76 (8):obc, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001182.

Anonymous:2010:EBg

[2556] Anonymous. Editorial Board. Journal of Computer and System Sciences, 76(8):ifc, December 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001145.

Cai:2011:F

[2557] Jin-Yi Cai and Alan L. Selman. Foreword. Journal of Computer and System Sciences, 77(1):1–2, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000010000991.

Agrawal:2011:ICC

[2558] Manindra Agrawal. The isomorphism conjecture for constant depth reductions. Journal of Computer and System Sciences, 77(1):3–13, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000010000875.

Allender:2011:PRR

[2559] Eric Allender, Michal Koucký, Detlef Ronneburger, and Sambuddha Roy. The pervasive reach of resourcebounded Kolmogorov complexity in computational complexity theory. Journal of Computer and System Sciences, 77(1):14–40, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000010000887.

Cai:2011:HAA

[2560] Jin-Yi Cai and Pinyan Lu. Holographic algorithms: From art to science. Journal of Computer and System Sciences, 77(1):41–61, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000899.

Dubey:2011:HRA

[2561] Chandan Dubey, Uriel Feige, and Walter Unger. Hardness results for approximating the bandwidth. Journal of Computer and System Sciences, 77(1):62–90, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000905.

Fortnow:2011:IIC

[2562] Lance Fortnow and Rahul Santhanam.

Infeasibility of instance compression and succinct PCPs for NP. Journal of Computer and System Sciences, 77 (1):91–106, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000917.

Gopalan:2011:HAW

[2563] Parikshit Gopalan and Venkatesan Guruswami. Hardness amplification within NP against deterministic algorithms. Journal of Computer and System Sciences, 77(1): 107–121, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000929.

Karp:2011:HAC

[2564] Richard M. Karp. Heuristic algorithms in computational molecular biology. Journal of Computer and System Sciences, 77(1): 122–128, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000930.

Khot:2011:HLI

[2565] Subhash Khot and Rishi Saket. On the hardness of learning intersections of two halfspaces. Journal of Computer and System Sciences, 77(1): 129–141, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000942.

Buchfuhrer:2011:CBF

[2566] David Buchfuhrer and Christopher Umans. The complexity of Boolean formula minimization. Journal of Computer and System Sciences, 77 (1):142–153, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000954.

Valiant:2011:BMM

[2567] Leslie G. Valiant. A bridging model for multi-core computing. Journal of Computer and System Sciences, 77 (1):154–166, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000966.

Raz:2011:MFM

[2568] Ran Raz and Amir Yehudayoff. Multilinear formulas, maximal-partition discrepancy and mixed-sources extractors. Journal of Computer and System Sciences, 77(1):167– 190, January 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000978.

Kamp:2011:DES

[2569] Jesse Kamp, Anup Rao, Salil Vadhan, and David Zuckerman. Deterministic extractors for smallspace sources. Journal of Computer and System Sciences, 77(1): 191–220, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001000098X.

Anonymous:2011:Ca

[2570] Anonymous.  Contents.  Journal

of Computer and System Sciences, 77(1):obc, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001000139X.

Anonymous:2011:EBa

[2571] Anonymous. Editorial Board. Journal of Computer and System Sciences, 77(1):ifc, January 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001352.

Anane:2011:ESI

[2572] Rachid Anane and Muhammad Younas.

Editorial for the special issue on “adaptivity in heterogeneous environments”. Journal of Computer and System Sciences, 77(2):221– 222, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000139.

Lin:2011:MGQ

[2573] Wei-Li Lin, Chi-Chun Lo, Kuo-Ming Chao, and Nick Godwin. Multi-group QoS consensus for web services. Journal of Computer and System Sciences, 77(2):223–243, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001000005X.

Glasner:2011:ATP

[2574] Christian Glasner and Jens Volkert. Adaps — a three-phase adaptive prediction system for the run-time of jobs based on user behaviour. Journal of Computer and System Sciences, 77 (2):244–261, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000061.

Caminero:2011:NAH

[2575] Agustín Caminero, Omer Rana, Blanca Caminero, and Carmen Carrión. Network-aware heuristics for interdomain meta-scheduling in grids. Journal of Computer and System Sciences, 77(2):262–281, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000073.

Bajaber:2011:ADR

[2576] Fuad Bajaber and Irfan Awan. Adaptive decentralized re-clustering protocol for wireless sensor networks. Journal of Computer and System Sciences, 77(2):282–292, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000085.

Randles:2011:DRR

[2577] Martin Randles, David Lamb, E. Odat, and A. Taleb-Bendiab. Distributed redundancy and robustness in complex systems. Journal of Computer and System Sciences, 77(2): 293–304, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000097.

Anthony:2011:AMS

[2578] Richard Anthony and Mona Ghassemian. Automatic mobility status estimation in wireless self-organised networks. Journal of Computer and System Sciences, 77(2):305– 321, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000103.

Bessler:2011:UTS

[2579] Sandford Bessler, Alexander Fischer, Eva Kühn, Richard Mordinyi, and Slobodanka Tomic. Using tuple-spaces to manage the storage and dissemination of spatial-temporal content. Journal of Computer and System Sciences, 77(2):322–331, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000115.

Ameedeen:2011:MIM

[2580] Mohamed A. Ameedeen, Behzad Bordbar, and Rachid Anane. Model interoperability via model driven development. Journal of Computer and System Sciences, 77(2):332– 347, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000127.

Shi:2011:FEV

[2581] Q. Shi, N. Zhang, and M. Merabti. Fair exchange of valuable information: a generalised framework. Journal of Computer and System Sciences, 77 (2):348–371, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001315.

Arrighi:2011:UPC

[2582] Pablo Arrighi, Vincent Nesme, and Reinhard Werner. Unitarity plus causality implies localizability. Journal of Computer and System Sciences, 77(2):372–378, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000589.

Fu:2011:TAW

[2583] Bin Fu. Theory and application of width bounded geometric separators. Journal of Computer and System Sciences, 77(2): 379–392, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000577.

Manna:2011:CRG

[2584] M. Manna, F. Scarcello, and N. Leone.

On the complexity of regular-grammars with integer attributes. Journal of Computer and System Sciences, 77 (2):393–421, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001000084X.

Gutin:2011:PAP

[2585] Gregory Gutin, Eun Jung Kim, Stefan Szeider, and Anders Yeo. A probabilistic approach to problems parameterized above or below tight bounds. Journal of Computer and System Sciences, 77(2):422–429, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000851.

Myrvold:2011:EGE

[2586] Wendy Myrvold and William Kocay. Errors in graph embedding algorithms. Journal of Computer and System Sciences, 77(2):430– 438, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000863.

Epstein:2011:GCR

[2587] Leah Epstein, Asaf Levin, and Gerhard J. Woeginger. Graph coloring with rejection. Journal of Computer and System Sciences, 77(2):

439–447, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001005.

Anonymous:2011:Cb

[2588] Anonymous. Contents. Journal of Computer and System Sciences, 77 (2):obc, March 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001492.

Anonymous:2011:EBb

[2589] Anonymous. Editorial Board. Journal of Computer and System Sciences, 77(2):ifc, March 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001455.

Arenas:2011:GES

[2590] Marcelo Arenas. Guest editorial: Special issue on database theory. Journal of Computer and System Sciences, 77(3):449, May 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000462.

Bjorklund:2011:CQC

[2591] Henrik Björklund, Wim Martens, and Thomas Schwentick. Conjunctive query containment over trees. Journal of Computer and System Sciences, 77(3):450–472, May 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000474.

Dalvi:2011:QMV

[2592] Nilesh Dalvi, Christopher Re, and Dan Suciu. Queries and materialized views on probabilistic databases. Journal of Computer and System Sciences, 77(3):473–490, May 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000486.

Geerts:2011:RCQ

[2593] Floris Geerts and Jan Van den Bussche.

Relational completeness of query languages for annotated databases. Journal of Computer and System Sciences, 77(3):491–504, May 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000498.

Gelade:2011:SPB

[2594] Wouter Gelade and Frank Neven. Succinctness of pattern-based schema languages for XML. Journal of Computer and System Sciences, 77(3): 505–519, May 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000504.

Gutierrez:2011:FSW

[2595] Claudio Gutierrez, Carlos A. Hurtado, Alberto O. Mendelzon, and Jorge Pérez. Foundations of Semantic Web databases. Journal of Computer and System Sciences, 77 (3):520–541, May 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000516.

Libkin:2011:DES

[2596] Leonid Libkin and Cristina Sirangelo. Data exchange and schema mappings in open and closed worlds. Journal of Computer and System Sciences, 77(3):542–571, May 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010000528.

Rosati:2011:FCC

[2597] Riccardo Rosati. On the finite controllability of conjunctive query answering in databases under openworld assumption. Journal of Computer and System Sciences, 77(3): 572–594, May 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001000053X.

Anonymous:2011:Cc

[2598] Anonymous. Contents. Journal of Computer and System Sciences, 77 (3):obc, May 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000122.

Anonymous:2011:EBc

[2599] Anonymous. Editorial Board. Journal of Computer and System Sciences, 77(3):ifc, May 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000080.

Xhafa:2011:SIF

[2600] Fatos Xhafa and Leonard Barolli.

Special issue foreword: Advances in pervasive computing systems and applications. Journal of Computer and System Sciences, 77(4): 595–596, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000164.

Siewe:2011:CCA

[2601] François Siewe, Hussein Zedan, and Antonio Cau. The calculus of contextaware ambients. Journal of Computer and System Sciences, 77(4): 597–620, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000176.

Papamarkos:2011:PME

[2602] George Papamarkos, Alexandra Poulovassilis, and Peter T. Wood. Performance modelling of event-condition-action rules in P2P networks. Journal of Computer and System Sciences, 77(4):621–636, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000188.

Xuan:2011:VBR

[2603] Kefeng Xuan, Geng Zhao, David Taniar, Wenny Rahayu, Maytham Safar, and Bala Srinivasan. Voronoibased range and continuous range query processing in mobile databases. Journal of Computer and System Sciences, 77(4):637–651, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200001000019X.

Ernst:2011:DEF

[2604] Jason B. Ernst and Mieso K. Denko. The design and evaluation of fair scheduling in wireless mesh networks. Journal of Computer and System Sciences, 77(4): 652–664, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000206.

Wang:2011:QEP

[2605] Jie Wang, Lin Guan, Lee Booi Lim, X. G. Wang, A. Grigg, I. Awan, I. Philips, and X. Chi. QoS enhancements and performance analysis for delay sensitive applications. Journal of Computer and System Sciences, 77(4):665–676, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000218.

Durresi:2011:NTM

[2606] Arjan Durresi, Mimoza Durresi, and Leonard Barolli. Network trust management in emergency situations. Journal of Computer and System Sciences, 77(4):677–686, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001000022X.

Dong:2011:FDC

[2607] Hai Dong, Farookh Khadeer Hussain, and Elizabeth Chang. A framework for discovering and classifying ubiquitous services in digital health ecosystems. Journal of Computer and System Sciences, 77 (4):687–704, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000231.

Sheltami:2011:VSA

[2608] Tarek R. Sheltami, Elhadi M. Shakshuki, and Hussein T. Mouftah. Video streaming application over WEAC protocol in MANET. Journal of Computer and System Sciences, 77 (4):705–719, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000243.

Hsu:2011:MRB

[2609] Hui-Huang Hsu and Hsuan-Ting Liao.

A mobile RFID-based tour system with instant microblogging. Journal of Computer and System Sciences, 77(4):720–727, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010000255.

Chow:2011:ANP

[2610] Timothy Y. Chow. Almost-natural proofs. Journal of Computer and System Sciences, 77(4):728–737, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000010001017.

Terwijn:2011:NNI

[2611] Sebastiaan A. Terwijn, Leen Torenvliet, and Paul M. B. Vitányi. Nonapproximability of the normalized information distance. Journal of Computer and System Sciences, 77(4): 738–742, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001029.

Charlier:2011:RRN

[2612] Émilie Charlier, Marion Le Gonidec, and Michel Rigo. Representing real numbers in a generalized numeration system. Journal of Computer and System Sciences, 77(4): 743–759, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001042.

Jain:2011:ISU

[2613] Sanjay Jain, Frank Stephan, and Jason Teutsch. Index sets and universal numberings. Journal of Computer and System Sciences, 77(4): 760–773, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001030.

Betzler:2011:APP

[2614] Nadja Betzler, Jiong Guo, Christian Komusiewicz, and Rolf Niedermeier. Average parameterization and partial kernelization for computing medians. Journal of Computer and System Sciences, 77(4): 774–789, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001078.

Wang:2011:PSE

[2615] Ren-Chiun Wang, Wen-Shenq Juang, and Chin-Laung Lei. Provably secure and efficient identification and key agreement protocol with user anonymity. Journal of Computer and System Sciences, 77(4): 790–798, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001066.

Fellows:2011:ULB

[2616] Michael R. Fellows, Guillaume Fertin, Danny Hermelin, and Stéphane Vialette. Upper and lower bounds for finding connected motifs in vertexcolored graphs. Journal of Computer and System Sciences, 77(4): 799–811, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001054.

Calude:2011:RLC

[2617] Cristian S. Calude, Nicholas J. Hay, and Frank Stephan. Representation of left-computable ϵ-random reals. Journal of Computer and System Sciences, 77(4):812–819, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001000108X.

Blondel:2011:SRM

[2618] Vincent Blondel, Stéphane Gaubert, and Natacha Portier. The set of realizations of a max-plus linear sequence is semi-polyhedral. Journal of Computer and System Sciences, 77(4):820–833, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001000125X.

Anonymous:2011:Cd

[2619] Anonymous. Contents. Journal of Computer and System Sciences, 77 (4):obc, July 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000237.

Anonymous:2011:EBd

[2620] Anonymous. Editorial Board. Journal of Computer and System Sciences, 77(4):ifc, July 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000195.

Guan:2011:PME

[2621] Lin Guan, Xingang Wang, Irfan Awan, and Waltenegus Dargie. Performance modeling and evaluation of heterogeneous computer networks. Journal of Computer and System Sciences, 77(5): 835–836, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001248.

Mapp:2011:EGL

[2622] Glenford Mapp, Dhawal Thakker, and Orhan Gemikonakli. Exploring gate-limited analytical models for high-performance network storage servers. Journal of Computer and System Sciences, 77(5): 837–851, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001091.

Dargie:2011:SPA

[2623] Waltenegus Dargie and Alexander Schill. Stability and performance analysis of randomly deployed wireless networks. Journal of Computer and System Sciences, 77(5): 852–860, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001108.

Clegg:2011:CMP

[2624] Richard G. Clegg, Raul Landa, and Miguel Rio. Criticisms of modelling packet traffic using long-range dependence (extended version). Journal of Computer and System Sciences, 77(5): 861–868, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001000111X.

Hong:2011:EFU

[2625] Jinpyo Hong and Hwangnam Kim. An empirical framework for user mobility models: Refining and modeling user registration patterns. Journal of Computer and System Sciences, 77(5): 869–883, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001121.

Subramanian:2011:SPS

[2626] Sureshkumar V. Subramanian and Rudra Dutta. A study of performance and scalability metrics of a SIP proxy server — a practical approach. Journal of Computer and System Sciences, 77 (5):884–897, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001194.

Lim:2011:CMQ

[2627] L. B. Lim, L. Guan, A. Grigg, I. W. Phillips, X. G. Wang, and I. U. Awan. Controlling mean queuing delay under multi-class bursty and correlated traffic. Journal of Computer and System Sciences, 77(5): 898–916, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001224.

Noh:2011:BEA

[2628] Dong Kun Noh and Kyungtae Kang. Balanced energy allocation scheme for a solar-powered sensor system and its effects on network-wide performance. Journal of Computer and System Sciences, 77(5):917–932, September 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001236.

Anonymous:2011:Ce

[2629] Anonymous. Contents. Journal of Computer and System Sciences, 77 (5):obc, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000651.

Anonymous:2011:EBe

[2630] Anonymous. Editorial Board. Journal of Computer and System Sciences, 77(5):ifc, September 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000614.

Cuzzocrea:2011:DXP

[2631] Alfredo Cuzzocrea and David Taniar.

Distributed XML processing and management: Theory and practice. Journal of Computer and System Sciences, 77 (6):933–935, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000365.

Abiteboul:2011:DXD

[2632] S. Abiteboul, G. Gottlob, and M. Manna. Distributed XML design. Journal of Computer and System Sciences, 77(6):936–964, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000011000328.

Cuzzocrea:2011:PPO

[2633] Alfredo Cuzzocrea and Elisa Bertino. Privacy preserving OLAP over distributed XML data: A theoreticallysound secure-multiparty-computation approach. Journal of Computer and System Sciences, 77(6):965– 987, November 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001100033X.

Greco:2011:CCX

[2634] Sergio Greco, Francesco Gullo, Giovanni Ponti, and Andrea Tagarelli. Collaborative clustering of XML documents. Journal of Computer and System Sciences, 77(6):988– 1008, November 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000341.

Moreira:2011:DCC

[2635] Leonardo O. Moreira, Flávio R. C. Sousa, and Javam C. Machado. A distributed concurrency control mechanism for XML data. Journal of Computer and System Sciences, 77(6): 1009–1022, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000353.

Jukna:2011:MRC

[2636] Stasys Jukna and Georg Schnitger. Min-rank conjecture for logdepth circuits. Journal of Computer and System Sciences, 77(6): 1023–1038, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000778.

Choi:2011:ATU

[2637] Sung-Soon Choi, Kyomin Jung, and Jeong Han Kim. Almost tight upper bound for finding Fourier coefficients of bounded pseudo-Boolean functions. Journal of Computer and System Sciences, 77(6):1039– 1053, November 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001261.

Wu:2011:NAM

[2638] Bang Ye Wu. A note on approximating the min-max vertex disjoint paths on directed acyclic graphs. Journal of Computer and System Sciences, 77(6): 1054–1057, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001273.

Kanj:2011:IMP

[2639] Iyad Kanj, Michael J. Pelsmajer, Marcus Schaefer, and Ge Xia. On the induced matching problem. Journal of Computer and System Sciences, 77(6): 1058–1070, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001285.

Bessy:2011:KFA

[2640] Stéphane Bessy, Fedor V. Fomin, Serge Gaspers, Christophe Paul, Anthony Perez, Saket Saurabh, and Stéphan Thomassé. Kernels for feedback arc set in tournaments. Journal of Computer and System Sciences, 77(6): 1071–1078, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001297.

Afrati:2011:SUV

[2641] Foto Afrati and Rada Chirkova. Selecting and using views to compute aggregate queries. Journal of Computer and System Sciences, 77(6): 1079–1107, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001327.

Dragan:2011:SSG

[2642] Feodor F. Dragan, Fedor V. Fomin, and Petr A. Golovach. Spanners in sparse graphs. Journal of Computer and System Sciences, 77(6): 1108–1119, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000010001303.

Ramanan:2011:MLB

[2643] Prakash Ramanan. Memory lower bounds for XPath evaluation over XML streams. Journal of Computer and System Sciences, 77(6): 1120–1140, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001339.

Fellows:2011:GNT

[2644] Michael R. Fellows, Jiong Guo, Hannes Moser, and Rolf Niedermeier. A generalization of Nemhauser and Trotter’s local optimization theorem. Journal of Computer and System Sciences, 77(6): 1141–1158, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001000142X.

Amini:2011:IBP

[2645] Omid Amini, Fedor V. Fomin, and Saket Saurabh. Implicit branching and parameterized partial cover problems. Journal of Computer and System Sciences, 77(6):1159– 1171, November 2011. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000010001431.

Alvarez:2011:EPG

[2646] Carme Àlvarez, Joaquim Gabarro, and Maria Serna. Equilibria problems on games: Complexity versus succinctness. Journal of Computer and System Sciences, 77(6): 1172–1197, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001100002X.

Anonymous:2011:Cf

[2647] Anonymous. Contents. Journal of Computer and System Sciences, 77 (6):ibc, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000924.

Anonymous:2011:Cg

[2648] Anonymous. Contents. Journal of Computer and System Sciences, 77 (6):obc, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000936.

Anonymous:2011:EBf

[2649] Anonymous. Editorial Board. Journal of Computer and System Sciences, 77(6):ifc, November 2011. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000894.

Palopoli:2012:EF

[2650] Luigi Palopoli. Editor’s foreword. Journal of Computer and System Sciences, 78(1):1, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000377.

Conitzer:2012:COO

[2651] Vincent Conitzer and Tuomas Sandholm. Computing optimal outcomes under an expressive representation of settings with externalities. Journal of Computer and System Sciences, 78(1):2–14, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000389.

Dobzinski:2012:TRM

[2652] Shahar Dobzinski, Noam Nisan, and Michael Schapira. Truthful randomized mechanisms for combinatorial auctions. Journal of Computer and System Sciences, 78(1): 15–25, January 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000390.

Calvanese:2012:VBQ

[2653] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Riccardo Rosati. View-based query answering in description logics: Semantics and complexity. Journal of Computer and System Sciences, 78 (1):26–46, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000407.

Eiter:2012:CQA

[2654] Thomas Eiter, Magdalena Ortiz, and Mantas Simkus. Conjunctive query answering in the description logic SH using knots. Journal of Computer and System Sciences, 78(1): 47–85, January 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000419.

Pearce:2012:STK

[2655] D. Pearce and A. Valverde. Synonymous theories and knowledge representations in answer set programming. Journal of Computer and System Sciences, 78(1):86–104, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000011000420.

Albagli:2012:MNB

[2656] Sivan Albagli, Rachel Ben-Eliyahu-Zohary, and Solomon E. Shimony. Markov network based ontology matching. Journal of Computer and System Sciences, 78(1):105– 118, January 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000432.

Armando:2012:ABA

[2657] Alessandro Armando, Enrico Giunchiglia, Marco Maratea, and Serena Elisa Ponta. An action-based approach to the formal specification and automatic analysis of business processes under authorization constraints. Journal of Computer and System Sciences, 78 (1):119–141, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000444.

Holbl:2012:ITP

[2658] Marko Hölbl, Tatjana Welzer, and Bostjan Brumen. An improved two-party identity-based authenticated key agreement protocol using pairings. Journal of Computer and System Sciences, 78(1): 142–150, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000031.

Gutin:2012:ETP

[2659] Gregory Gutin, Leo van Iersel, Matthias Mnich, and Anders Yeo. Every ternary permutation constraint satisfaction problem parameterized above average has a kernel with a quadratic number of variables. Journal of Computer and System Sciences, 78(1): 151–163, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000055.

Chen:2012:TSA

[2660] Zhi-Zhong Chen, Bin Ma, and Lusheng Wang. A three-string approach to the closest string problem. Journal of Computer and System Sciences, 78 (1):164–178, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000043.

Mathieson:2012:EGS

[2661] Luke Mathieson and Stefan Szeider. Editing graphs to satisfy degree constraints: A parameterized approach. Journal of Computer and System Sciences, 78(1): 179–191, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000067.

Preparata:2012:APA

[2662] Franco P. Preparata. Accurate and precise aggregation counting. Journal of Computer and System Sciences, 78 (1):192–197, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000171.

Bjorklund:2012:TFN

[2663] Henrik Björklund and Wim Martens. The tractability frontier for NFA minimization. Journal of Computer and System Sciences, 78(1): 198–210, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000456.

Chen:2012:KKC

[2664] Jianer Chen and Jie Meng. A 2 k kernel for the cluster editing problem. Journal of Computer and System Sciences, 78 (1):211–220, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000468.

Ipate:2012:LFC

[2665] Florentin Ipate. Learning finite cover automata from queries. Journal of Computer and System Sciences, 78 (1):221–244, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001100047X.

Jansen:2012:SMT

[2666] Klaus Jansen and Hu Zhang. Scheduling malleable tasks with precedence constraints. Journal of Computer and System Sciences, 78(1):245– 259, January 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000481.

Braverman:2012:OSS

[2667] Vladimir Braverman, Rafail Ostrovsky, and Carlo Zaniolo. Optimal sampling from sliding windows. Journal of Computer and System Sciences, 78 (1):260–272, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000493.

Peng:2012:NEI

[2668] Yung-Hsing Peng, Chang-Biau Yang, Chiou-Ting Tseng, and Chiou-Yi Hor. A new efficient indexing algorithm for one-dimensional real scaled patterns. Journal of Computer and System Sciences, 78(1): 273–278, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001100050X.

Feldman:2012:IAA

[2669] Moran Feldman, Guy Kortsarz, and Zeev Nutov. Improved approximation algorithms for Directed Steiner Forest. Journal of Computer and System Sciences, 78(1): 279–292, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000584.

Dassow:2012:RCF

[2670] Jürgen Dassow and Tomás Masopust. On restricted context-free grammars. Journal of Computer and System Sciences, 78(1):293– 304, January 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000572.

Gaspers:2012:UFA

[2671] Serge Gaspers and Gregory B. Sorkin. A universally fastest algorithm for Max 2-Sat, Max 2-CSP, and everything in between. Journal of Computer and System Sciences, 78(1): 305–335, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000596.

Bomhoff:2012:BBD

[2672] Matthijs Bomhoff, Walter Kern, and Georg Still. On bounded block decomposition problems for under-specified systems of equations. Journal of Computer and System Sciences, 78 (1):336–347, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000705.

Kohlas:2012:GLC

[2673] Jürg Kohlas, Marc Pouly, and Cesar Schneuwly. Generic local computation. Journal of Computer and System Sciences, 78(1):348– 369, January 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000717.

Talaei-Khoei:2012:MAU

[2674] Amir Talaei-Khoei, Terje Solvoll, Pradeep Ray, and Nandan Parameshwaran. Maintaining awareness using policies; enabling agents to identify relevance of information. Journal of Computer and System Sciences, 78 (1):370–391, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000729.

Anonymous:2012:Ca

[2675] Anonymous.  Contents.  Journal

of Computer and System Sciences, 78(1):obc, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001267.

Anonymous:2012:EBa

[2676] Anonymous. Editorial Board. Journal of Computer and System Sciences, 78(1):ifc, January 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001100122X.

Veith:2012:SIG

[2677] Helmut Veith. Special issue: Games in verification. Journal of Computer and System Sciences, 78(2):

393, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000560.

Chatterjee:2012:SSR

[2678] Krishnendu Chatterjee and Thomas A. Henzinger. A survey of stochastic ω-regular games. Journal of Computer and System Sciences, 78 (2):394–413, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000511.

Shoham:2012:MVM

[2679] Sharon Shoham and Orna Grumberg. Multi-valued model checking games. Journal of Computer and System Sciences, 78(2):414– 429, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000523.

Heljanko:2012:SPG

[2680] Keijo Heljanko, Misa Keinänen, Martin Lange, and Ilkka Niemelä. Solving parity games by a reduction to SAT. Journal of Computer and System Sciences, 78(2):430–440, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000535.

Jobstmann:2012:FFF

[2681] Barbara Jobstmann, Stefan Staber, Andreas Griesmayer, and Roderick Bloem. Finding and fixing faults. Journal of Computer and System Sciences, 78(2):441–460, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000547.

Lohrey:2012:MCH

[2682] Markus Lohrey. Model-checking hierarchical structures. Journal of Computer and System Sciences, 78 (2):461–490, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000559.

DiGiacomo:2012:DTM

[2683] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, and Henk Meijer. Drawing a tree as a minimum spanning tree approximation. Journal of Computer and System Sciences, 78 (2):491–503, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000730.

Tian:2012:RCA

[2684] Jing Tian and Xianzhong Zhao. Representations of commutative asynchronous automata. Journal of Computer and System Sciences, 78(2): 504–516, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000742.

Bedon:2012:SET

[2685] Nicolas Bedon and Chloé Rispal. Schützenberger and Eilenberg theorems for words on linear orderings. Journal of Computer and System Sciences, 78(2):517– 536, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000754.

Kapoutsis:2012:SCR

[2686] Christos Kapoutsis, Richard Královic, and Tobias Mömke. Size complexity of rotating and sweeping automata. Journal of Computer and System Sciences, 78(2):537–558, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000766.

Weller:2012:MDG

[2687] Mathias Weller, Christian Komusiewicz, Rolf Niedermeier, and Johannes Uhlmann. On making directed graphs transitive. Journal of Computer and System Sciences, 78(2): 559–574, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001100078X.

Lara:2012:PME

[2688] Pedro Lara, Fábio Borges, Renato Portugal, and Nadia Nedjah. Parallel modular exponentiation using load balancing without precomputation. Journal of Computer and System Sciences, 78(2):575–582, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000791.

Deutch:2012:STQ

[2689] Daniel Deutch and Tova Milo. A structural/temporal query language for business processes. Journal of Computer and System Sciences, 78 (2):583–609, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001036.

Shao:2012:STA

[2690] Zehui Shao, Fei Deng, Meilian Liang, and Xiaodong Xu. On sets without k-term arithmetic progression. Journal of Computer and System Sciences, 78(2):610–618, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001024.

Jansson:2012:USR

[2691] Jesper Jansson, Kunihiko Sadakane, and Wing-Kin Sung. Ultra-succinct representation of ordered trees with applications. Journal of Computer and System Sciences, 78(2): 619–631, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001012.

Yuster:2012:ASP

[2692] Raphael Yuster. Approximate shortest paths in weighted graphs. Journal of Computer and System Sciences, 78(2):632–637, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001000.

Bulatov:2012:EH

[2693] Andrei A. Bulatov, Víctor Dalmau, Martin Grohe, and Dániel Marx. Enumerating homomorphisms. Journal of Computer and System Sciences, 78 (2):638–650, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001100105X.

Katz:2012:TSP

[2694] Jonathan Katz, Philip MacKenzie, Gelareh Taban, and Virgil Gligor. Two-server password-only authenticated key exchange. Journal of Computer and System Sciences, 78 (2):651–669, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001048.

Kawarabayashi:2012:LTA

[2695] Ken ichi Kawarabayashi and Yusuke Kobayashi. A linear time algorithm for the induced disjoint paths problem in planar graphs. Journal of Computer and System Sciences, 78 (2):670–680, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001103.

Bulatov:2012:CWU

[2696] Andrei Bulatov, Martin Dyer, Leslie Ann Goldberg, Markus Jalsenius, Mark Jerrum, and David Richerby. The complexity of weighted and unweighted #CSP. Journal of Computer and System Sciences, 78(2): 681–688, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001474.

Anonymous:2012:Cb

[2697] Anonymous. Contents. Journal of Computer and System Sciences, 78 (2):ibc, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000074.

Anonymous:2012:Cc

[2698] Anonymous. Contents. Journal of Computer and System Sciences, 78 (2):obc, March 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000086.

Anonymous:2012:EBb

[2699] Anonymous. Editorial Board. Journal of Computer and System Sciences, 78(2):ifc, March 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000037.

Jiang:2012:ACL

[2700] Minghui Jiang. Approximability of constrained LCS. Journal of Computer and System Sciences, 78(3): 689–697, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001085.

Fomin:2012:FAF

[2701] Fedor V. Fomin, Daniel Lokshtanov, Venkatesh Raman, Saket Saurabh, and B. V. Raghavendra Rao. Faster algorithms for finding and counting subgraphs. Journal of Computer and System Sciences, 78(3): 698–706, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001073.

Fellows:2012:LSB

[2702] Michael R. Fellows, Fedor V. Fomin, Daniel Lokshtanov, Frances Rosamond, Saket Saurabh, and Yngve Villanger. Local search: Is bruteforce avoidable? Journal of Computer and System Sciences, 78(3): 707–719, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001097.

Lin:2012:IAA

[2703] Guohui Lin, Randy Goebel, Zhong Li, and Lusheng Wang. An improved approximation algorithm for the complementary maximal strip recovery problem. Journal of Computer and System Sciences, 78(3): 720–730, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001206.

Durand:2012:FPT

[2704] Bruno Durand, Andrei Romashchenko, and Alexander Shen. Fixed-point tile sets and their applications. Journal of Computer and System Sciences, 78(3):731–764, May 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001358.

Datta:2012:SCP

[2705] Samir Datta, Raghav Kulkarni, Raghunath Tewari, and N. V. Vinodchandran. Space complexity of perfect matching in bounded genus bipartite graphs. Journal of Computer and System Sciences, 78(3): 765–779, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001100136X.

Nagy:2012:CSS

[2706] Benedek Nagy and Friedrich Otto.

On CD-systems of stateless deterministic R-automata with window size one. Journal of Computer and System Sciences, 78(3): 780–806, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001541.

Lin:2012:AAE

[2707] Tianrong Lin. Another approach to the equivalence of measure-many oneway quantum finite automata and its application. Journal of Computer and System Sciences, 78(3): 807–821, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001200030X. See

corrigendum [3109].

Bouajjani:2012:EF

[2708] Ahmed Bouajjani, David Harel, and Lenore Zuck. Editor’s foreword. Journal of Computer and System Sciences, 78(3):822, May 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000808.

Apt:2012:VOO

[2709] Krzysztof R. Apt, Frank S. de Boer, Ernst-Rüdiger Olderog, and Stijn de Gouw. Verification of objectoriented programs: A transformational approach. Journal of Computer and System Sciences, 78(3): 823–852, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001100081X.

Balaban:2012:VML

[2710] Ittai Balaban, Amir Pnueli, Yaniv Sa’ar, and Lenore D. Zuck. Verification of multi-linked heaps. Journal of Computer and System Sciences, 78(3):853–876, May 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000821.

Benveniste:2012:NSS

[2711] Albert Benveniste, Timothy Bourke, Benoît Caillaud, and Marc Pouzet. Non-standard semantics of hybrid systems modelers. Journal of Computer and System Sciences, 78(3): 877–910, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001061.

Bloem:2012:SRD

[2712] Roderick Bloem, Barbara Jobstmann, Nir Piterman, Amir Pnueli, and Yaniv Sa’ar. Synthesis of reactive(1) designs. Journal of Computer and System Sciences, 78(3): 911–938, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000869.

Borger:2012:AAS

[2713] Egon Börger, Antonio Cisternino, and Vincenzo Gervasi. Ambient abstract state machines with applications. Journal of Computer and System Sciences, 78(3):939–959, May 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011000833.

Francez:2012:WDT

[2714] Nissim Francez. When are different type-logical semantic definitions defining equivalent meanings? Journal of Computer and System Sciences, 78(3):960–969, May 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000845.

Harel:2012:SSB

[2715] David Harel and Itai Segall. Synthesis from scenario-based specifications. Journal of Computer and System Sciences, 78(3):970– 980, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000870.

Kupferman:2012:OA

[2716] Orna Kupferman, Amir Pnueli, and Moshe Y. Vardi. Once and for all. Journal of Computer and System Sciences, 78(3):981–996, May 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011000857.

Anonymous:2012:Cd

[2717] Anonymous. Contents. Journal of Computer and System Sciences, 78 (3):obc, May 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000396.

Anonymous:2012:EBc

[2718] Anonymous. Editorial Board. Journal of Computer and System Sciences, 78(3):ifc, May 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000359.

Wei:2012:IOO

[2719] Chunfu Wei, Zheng Pei, and Huamin Li. An induced OWA operator in coal mine safety evaluation. Journal of Computer and System Sciences, 78(4):997–1005, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001449.

Ramanan:2012:RXQ

[2720] Prakash Ramanan. Rewriting XPath queries using materialized XPath views. Journal of Computer and System Sciences, 78(4):1006–1025, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000011001462.

Link:2012:CMD

[2721] Sebastian Link. Characterisations of multivalued dependency implication over undetermined universes. Journal of Computer and System Sciences, 78(4):1026–1044, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001577.

Vincent:2012:IPC

[2722] M. W. Vincent, J. Liu, and M. Mohania. The implication problem for ‘closest node’ functional dependencies in complete XML documents. Journal of Computer and System Sciences, 78(4):1045–1098, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000293.

Lou:2012:NAA

[2723] Xiaowen Lou and Daming Zhu. A new approximation algorithm for cutand-paste sorting of unsigned circular permutations. Journal of Computer and System Sciences, 78(4): 1099–1114, July 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000311.

Gajardo:2012:TSC

[2724] Anahí Gajardo, Jarkko Kari, and Andrés Moreira. On time-symmetry in cellular automata. Journal of Computer and System Sciences, 78 (4):1115–1126, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000323.

Wen:2012:VSF

[2725] Yicheng Wen and Asok Ray. Vector space formulation of probabilistic finite state automata. Journal of Computer and System Sciences, 78 (4):1127–1141, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000335.

Beimel:2012:CED

[2726] Amos Beimel, Yeow Meng Chee, Huaxiong Wang, and Liang Feng Zhang. Communication-efficient distributed oblivious transfer. Journal of Computer and System Sciences, 78 (4):1142–1157, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000487.

Iqbal:2012:IRD

[2727] Rahat Iqbal and Nazaraf Shah. Information retrieval, decision making process and user needs. Journal of Computer and System Sciences, 78 (4):1158–1159, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001115.

Gemmell:2012:RRS

[2728] Jonathan Gemmell, Thomas Schimoler, Bamshad Mobasher, and Robin Burke. Resource recommendation in social annotation systems: A linearweighted hybrid approach. Journal of Computer and System Sciences, 78 (4):1160–1174, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001127.

Eirinaki:2012:FBO

[2729] Magdalini Eirinaki, Shamita Pisal, and Japinder Singh. Feature-based opinion mining and ranking. Journal of Computer and System Sciences, 78 (4):1175–1184, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001139.

Shen:2012:PMU

[2730] Haifeng Shen and Mark Reilly. Personalized multi-user view and content synchronization and retrieval in real-time mobile social software applications. Journal of Computer and System Sciences, 78(4):1185– 1203, July 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001140.

Grzywaczewski:2012:TSI

[2731] Adam Grzywaczewski and Rahat Iqbal. Task-specific information retrieval systems for software engineers. Journal of Computer and System Sciences, 78 (4):1204–1218, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001152.

Rahman:2012:SAD

[2732] Fazilatur Rahman and Jawed Siddiqi. Semantic annotation of digital music. Journal of Computer and System Sciences, 78(4):1219– 1231, July 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001164.

Fenza:2012:HAC

[2733] G. Fenza, D. Furno, and V. Loia. Hybrid approach for context-aware service discovery in healthcare domain. Journal of Computer and System Sciences, 78(4):1232–1247, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001176.

Qureshi:2012:HMC

[2734] Pir Abdul Rasool Qureshi and Nasrullah Memon. Hybrid model of content extraction. Journal of Computer and System Sciences, 78(4): 1248–1257, July 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001188.

Iqbal:2012:CBI

[2735] Kashif Iqbal, Michael O. Odetayo, and Anne James. Content-based image retrieval approach for biometric security using colour, texture and shape features controlled by fuzzy heuristics. Journal of Computer and System Sciences, 78(4): 1258–1277, July 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001100119X.

Anonymous:2012:Ce

[2736] Anonymous. Contents. Journal of Computer and System Sciences, 78 (4):ibc, July 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000761.

Anonymous:2012:Cf

[2737] Anonymous. Contents. Journal of Computer and System Sciences, 78 (4):obc, July 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001200061X.

Anonymous:2012:EBd

[2738] Anonymous. Editorial Board. Journal of Computer and System Sciences, 78(4):ifc, July 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000578.

Chen:2012:SIC

[2739] Jinjun Chen and Lizhe Wang. Special issue: Cloud computing 2011. Journal of Computer and System Sciences, 78(5):1279, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001589.

Wu:2012:SBA

[2740] Linlin Wu, Saurabh Kumar Garg, and Rajkumar Buyya. SLA-based admission control for a Software-as-a-Service provider in Cloud computing environments. Journal of Computer and System Sciences, 78(5): 1280–1299, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001590.

Tolosana-Calasanz:2012:EQS

[2741] Rafael Tolosana-Calasanz, José Ángel Bañares, Congduc Pham, and Omer F. Rana. Enforcing QoS in scientific workflow systems enacted over cloud infrastructures. Journal of Computer and System Sciences, 78(5): 1300–1315, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001607.

Qi:2012:QAC

[2742] Lianyong Qi, Wanchun Dou, Xuyun Zhang, and Jinjun Chen. A QoSaware composition method supporting cross-platform service invocation in cloud environment. Journal of Computer and System Sciences, 78(5): 1316–1329, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001619.

Villegas:2012:CFL

[2743] David Villegas, Norman Bobroff, Ivan Rodero, Javier Delgado, Yanbin Liu, Aditya Devarakonda, Liana Fong, S. Masoud Sadjadi, and Manish Parashar. Cloud federation in a layered service model. Journal of Computer and System Sciences, 78(5): 1330–1344, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001620.

He:2012:DDP

[2744] Jing He, Yanchun Zhang, Guangyan Huang, Yong Shi, and Jie Cao. Distributed data possession checking for securing multiple replicas in geographically-dispersed clouds. Journal of Computer and System Sciences, 78(5):1345–1358, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000011001632.

Chadwick:2012:PPA

[2745] David W. Chadwick and Kaniz Fatema. A privacy preserving authorisation system for the cloud. Journal of Computer and System Sciences, 78(5): 1359–1373, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001644.

Zhang:2012:HPB

[2746] Gaofeng Zhang, Yun Yang, and Jinjun Chen. A historical probability based noise generation strategy for privacy protection in cloud computing. Journal of Computer and System Sciences, 78(5):1374– 1381, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001656.

Marozzo:2012:PMP

[2747] Fabrizio Marozzo, Domenico Talia, and Paolo Trunfio. P2P-MapReduce: Parallel data processing in dynamic Cloud environments. Journal of Computer and System Sciences, 78(5): 1382–1402, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001668.

Rakhlin:2012:F

[2748] Alexander Rakhlin. Foreword. Journal of Computer and System Sciences, 78(5):1403, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000207.

Cesa-Bianchi:2012:CB

[2749] Nicolò Cesa-Bianchi and Gábor Lugosi. Combinatorial bandits. Journal of Computer and System Sciences, 78(5): 1404–1422, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000219.

Dalalyan:2012:SRL

[2750] A. S. Dalalyan and A. B. Tsybakov. Sparse regression learning by aggregation and Langevin Monte-Carlo. Journal of Computer and System Sciences, 78(5):1423– 1443, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000220.

Feldman:2012:CCS

[2751] Vitaly Feldman. A complete characterization of statistical query learning with applications to evolvability. Journal of Computer and System Sciences, 78(5):1444–1459, September 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000232.

Hsu:2012:SAL

[2752] Daniel Hsu, Sham M. Kakade, and Tong Zhang. A spectral algorithm for learning hidden Markov models. Journal of Computer and System Sciences, 78(5):1460–1480, September 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000244.

Kalai:2012:RAL

[2753] Adam Tauman Kalai, Varun Kanade, and Yishay Mansour. Reliable agnostic learning. Journal of Computer and System Sciences, 78(5): 1481–1495, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000256.

Kpotufe:2012:TBR

[2754] Samory Kpotufe and Sanjoy Dasgupta. A tree-based regressor that adapts to intrinsic dimension. Journal of Computer and System Sciences, 78(5): 1496–1515, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000268.

Lazaric:2012:LSI

[2755] Alessandro Lazaric and Rémi Munos. Learning with stochastic inputs and adversarial outputs. Journal of Computer and System Sciences, 78(5): 1516–1537, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001200027X.

Yue:2012:ADB

[2756] Yisong Yue, Josef Broder, Robert Kleinberg, and Thorsten Joachims.

The K-armed dueling bandits problem. Journal of Computer and System Sciences, 78(5):1538–1556, September 2012. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000281.

Jordan:2012:TUC

[2757] Charles Jordan and Thomas Zeugmann. Testable and untestable classes of first-order formulae. Journal of Computer and System Sciences, 78(5): 1557–1578, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001200089X.

Chebolu:2012:CAC

[2758] Prasad Chebolu, Leslie Ann Goldberg, and Russell Martin. The complexity of approximately counting stable roommate assignments. Journal of Computer and System Sciences, 78(5): 1579–1605, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000499.

Dorn:2012:CSD

[2759] Frederic Dorn, Fedor V. Fomin, and Dimitrios M. Thilikos. Catalan structures and dynamic programming in Hminor-free graphs. Journal of Computer and System Sciences, 78(5): 1606–1622, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000505.

Carlucci:2012:LOB

[2760] Lorenzo Carlucci, Sanjay Jain, and Frank Stephan. Learning with ordinal-bounded memory from positive data. Journal of Computer and System Sciences, 78(5):1623– 1636, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000876.

Chen:2012:MTV

[2761] Jianer Chen, Jia-Hao Fan, Iyad Kanj, Yang Liu, and Fenghui Zhang. Multicut in trees viewed through the eyes of vertex cover. Journal of Computer and System Sciences, 78(5): 1637–1650, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000864.

Lohrey:2012:PRA

[2762] Markus Lohrey, Sebastian Maneth, and Manfred Schmidt-Schauß. Parameter reduction and automata evaluation for grammar-compressed trees. Journal of Computer and System Sciences, 78(5): 1651–1669, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000888.

Anonymous:2012:Cg

[2763] Anonymous. Contents. Journal of Computer and System Sciences, 78 (5):ibc, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001109.

Anonymous:2012:Ch

[2764] Anonymous. Contents. Journal of Computer and System Sciences, 78 (5):obc, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001110.

Anonymous:2012:EBe

[2765] Anonymous. Editorial Board. Journal of Computer and System Sciences, 78(5):ifc, September 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001079.

Jin:2012:SIM

[2766] Qun Jin and Yufeng Wang. Special issue on multidisciplinary emerging networks and systems. Journal of Computer and System Sciences, 78(6): 1671–1672, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001371.

Kajioka:2012:AAR

[2767] Shinsuke Kajioka, Naoki Wakamiya, and Masayuki Murata. Autonomous and adaptive resource allocation among multiple nodes and multiple applications in heterogeneous wireless networks. Journal of Computer and System Sciences, 78(6): 1673–1685, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001383.

Lin:2012:EEA

[2768] Chi Lin, Guowei Wu, Feng Xia, Mingchu Li, Lin Yao, and Zhongyi Pei. Energy efficient ant colony algorithms for data aggregation in wireless sensor networks. Journal of Computer and System Sciences, 78(6): 1686–1702, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001395.

Kohno:2012:IAI

[2769] Eitaro Kohno, Tomoya Okazaki, Mario Takeuchi, Tomoyuki Ohta, Yoshiaki Kakuda, and Masaki Aida. Improvement of assurance including security for wireless sensor networks using dispersed data transmission. Journal of Computer and System Sciences, 78(6): 1703–1715, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001401.

Waluyo:2012:TBE

[2770] Agustinus Borgy Waluyo, David Taniar, Wenny Rahayu, Ailixier Aikebaier, Makoto Takizawa, and Bala Srinivasan. Trustworthy-based efficient data broadcast model for P2P interaction in resource-constrained wireless environments. Journal of Computer and System Sciences, 78(6): 1716–1736, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001413.

Fan:2012:BBR

[2771] Xinxin Fan, Mingchu Li, Jianhua Ma, Yizhi Ren, Hui Zhao, and Zhiyuan Su. Behavior-based reputation management in P2P filesharing networks. Journal of Computer and System Sciences, 78(6): 1737–1750, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001437.

Dohi:2012:AMC

[2772] Tadashi Dohi and Toshikazu Uemura. An adaptive mode control algorithm of a scalable intrusion tolerant architecture. Journal of Computer and System Sciences, 78(6): 1751–1774, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001201.

Chen:2012:SUP

[2773] Hong Chen, Xiaokang Zhou, and Qun Jin. Socialized ubiquitous personal study: Toward an individualized information portal. Journal of Computer and System Sciences, 78(6): 1775–1792, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001425.

Fernau:2012:GE

[2774] Henning Fernau and Carlos Martín-Vide. From the Guest Editors. Journal of Computer and System Sciences, 78(6):1793–1794, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000011001450.

Hovland:2012:IPR

[2775] Dag Hovland. The inclusion problem for regular expressions. Journal of Computer and System Sciences, 78(6): 1795–1813, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001486.

Kutrib:2012:RPA

[2776] Martin Kutrib and Andreas Malcher. Reversible pushdown automata. Journal of Computer and System Sciences, 78(6):1814–1827, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000011001498.

Crochemore:2012:MNC

[2777] M. Crochemore, C. S. Iliopoulos, M. Kubica, J. Radoszewski, W. Rytter, and T. Waleń. The maximal number of cubic runs in a word. Journal of Computer and System Sciences, 78(6): 1828–1836, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001504.

Reghizzi:2012:OPV

[2778] Stefano Crespi Reghizzi and Dino Mandrioli. Operator precedence and the visibly pushdown property. Journal of Computer and System Sciences, 78(6): 1837–1867, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001516.

Policriti:2012:RNA

[2779] Alberto Policriti, Alexandru I. Tomescu, and Francesco Vezzi. A randomized Numerical Aligner (rNA). Journal of Computer and System Sciences, 78(6): 1868–1882, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001528.

Arrighi:2012:IUD

[2780] Pablo Arrighi and Jonathan Grattage.

Intrinsically universal n-dimensional quantum cellular automata. Journal of Computer and System Sciences, 78(6): 1883–1898, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001100153X.

Buszkowski:2012:PGL

[2781] Wojciech Buszkowski, Zhe Lin, and Katarzyna Moroz. Pregroup grammars with letter promotions: Complexity and context-freeness. Journal of Computer and System Sciences, 78(6): 1899–1909, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000011001553.

Jain:2012:LAC

[2782] Sanjay Jain, Qinglong Luo, and Frank Stephan. Learnability of automatic classes. Journal of Computer and System Sciences, 78(6): 1910–1927, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000011001565.

Anonymous:2012:Ci

[2783] Anonymous. Contents. Journal of Computer and System Sciences, 78 (6):ibc, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001304.

Anonymous:2012:Cj

[2784] Anonymous. Contents. Journal of Computer and System Sciences, 78 (6):obc, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001316.

Anonymous:2012:EBf

[2785] Anonymous. Editorial Board. Journal of Computer and System Sciences, 78(6):ifc, November 2012. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001274.

Fomin:2013:LVK

[2786] Fedor V. Fomin, Serge Gaspers, Saket Saurabh, and Stéphan Thomassé. A linear vertex kernel for maximum internal spanning tree. Journal of Computer and System Sciences, 79 (1):1–6, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000906.

Manea:2013:HCF

[2787] Florin Manea and Catalin Tiseanu.

The hardness of counting full words compatible with partial words. Journal of Computer and System Sciences, 79(1):7–22, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000918.

Demri:2013:CBP

[2788] Stéphane Demri, Marcin Jurdziński, Oded Lachish, and Ranko Lazić. The covering and boundedness problems for branching vector addition systems. Journal of Computer and System Sciences, 79(1):23– 38, February 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001200092X.

Jansen:2013:PFN

[2789] Klaus Jansen, Stefan Kratsch, Dániel Marx, and Ildikó Schlotter. Bin packing with fixed number of bins revisited. Journal of Computer and System Sciences, 79(1):39–49, February 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000943.

DasGupta:2013:CNC

[2790] Bhaskar DasGupta and Devendra Desai. On the complexity of Newman’s community finding approach for biological and social networks. Journal of Computer and System Sciences, 79 (1):50–67, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000931.

Goldberg:2013:ATP

[2791] Leslie Ann Goldberg and Mark Jerrum. Approximating the Tutte polynomial of a binary matroid and other related combinatorial polynomials. Journal of Computer and System Sciences, 79(1):68–78, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000955.

Bodirsky:2013:DCS

[2792] Manuel Bodirsky and Víctor Dalmau. Datalog and constraint satisfaction with infinite templates. Journal of Computer and System Sciences, 79 (1):79–100, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001213.

Honkala:2013:SEP

[2793] Juha Honkala. The sequence equivalence problem for primitive D0L systems. Journal of Computer and System Sciences, 79(1):101– 110, February 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001225.

Kesselman:2013:SST

[2794] Alexander Kesselman, Kirill Kogan, Sergey Nemzer, and Michael Segal. Space and speed tradeoffs in TCAM hierarchical packet classification. Journal of Computer and System Sciences, 79 (1):111–121, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001237.

Gong:2013:NOT

[2795] Longyan Gong, Jingxin Pan, Beibei Liu, and Shengmei Zhao. A novel one-time password mutual authentication scheme on sharing renewed finite random sub-passwords. Journal of Computer and System Sciences, 79 (1):122–130, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001249.

Liu:2013:SGD

[2796] Peide Liu. Some generalized dependent aggregation operators with intuitionistic linguistic numbers and their application to group decision making. Journal of Computer and System Sciences, 79(1): 131–143, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001250.

Marx:2013:CIM

[2797] Dániel Marx. Completely inapproximable monotone and antimonotone parameterized problems. Journal of Computer and System Sciences, 79 (1):144–151, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001365.

Cao:2013:PAC

[2798] Yongzhi Cao, Lirong Xia, and Mingsheng Ying. Probabilistic automata for computing with words. Journal of Computer and System Sciences, 79 (1):152–172, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001377.

Anonymous:2013:Ca

[2799] Anonymous. Contents. Journal of Computer and System Sciences, 79 (1):obc, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001572.

Anonymous:2013:EBa

[2800] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(1):ifc, February 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001535.

Jin:2013:SIF

[2801] Xiaolong Jin, Ahmed Y. Al-Dubai, and Laurence T. Yang. Special issue: Frontiers and advance topics of computer and information technology. Journal of Computer and System Sciences, 79 (2):173–174, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000967.

Ye:2013:RCO

[2802] John M. Ye, Man Cao, Zening Qu, and Tianzhou Chen. Regional cache organization for NoC based manycore processors. Journal of Computer and System Sciences, 79(2): 175–186, March 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000979.

Li:2013:IDP

[2803] Dingding Li, Hai Jin, Xiaofei Liao, Yu Zhang, and Bingbing Zhou. Improving disk I/O performance in a virtualized system. Journal of Computer and System Sciences, 79(2): 187–200, March 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012000980.

Chen:2013:THC

[2804] Gang Chen, Yongwei Wu, Jie Wu, and Weimin Zheng. TopCluster: a hybrid cluster model to support dynamic deployment in Grid. Journal of Computer and System Sciences, 79 (2):201–215, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001055.

Li:2013:SMA

[2805] You Li, Kaiyong Zhao, Xiaowen Chu, and Jiming Liu. Speeding up kmeans algorithm by GPUs. Journal of Computer and System Sciences, 79 (2):216–229, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012000992.

Baraglia:2013:MCJ

[2806] Ranieri Baraglia, Gabriele Capannini, Patrizio Dazzi, and Giancarlo Pagano. A multi-criteria job scheduling framework for large computing farms. Journal of Computer and System Sciences, 79(2):230–244, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001006.

Zheng:2013:AAE

[2807] Lihong Zheng, Xiangjian He, Bijan Samali, and Laurence T. Yang. An algorithm for accuracy enhancement of license plate recognition. Journal of Computer and System Sciences, 79 (2):245–255, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001018.

Ma:2013:SFM

[2808] Jun Ma, Long Zheng, Mianxiong Dong, Xiangjian He, Minyi Guo, Yuichi Yaguchi, and Ryuichi Oka.

A segmentation-free method for image classification based on pixelwise matching. Journal of Computer and System Sciences, 79(2): 256–268, March 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001043.

Xu:2013:AMP

[2809] Junfeng Xu, Keqiu Li, and Geyong Min. Asymmetric multi-path division communications in underwater acoustic networks with fading channels. Journal of Computer and System Sciences, 79(2): 269–278, March 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001200102X.

Wang:2013:SSD

[2810] Guojun Wang, Fengshun Yue, and Qin Liu. A secure self-destructing scheme for electronic data. Journal of Computer and System Sciences, 79 (2):279–290, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001031.

Baraglia:2013:PPR

[2811] Ranieri Baraglia, Patrizio Dazzi, Matteo Mordacchini, and Laura Ricci. A peer-to-peer recommender system for self-emerging user communities based on gossip overlays. Journal of Computer and System Sciences, 79 (2):291–308, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001195.

Anonymous:2013:Cb

[2812] Anonymous. Contents. Journal of Computer and System Sciences, 79 (2):obc, March 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001651.

Anonymous:2013:EBb

[2813] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(2):ifc, March 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001614.

Cuzzocrea:2013:TPA

[2814] Alfredo Cuzzocrea. Theoretical and practical aspects of warehousing, querying and mining sensor and streaming data. Journal of Computer and System Sciences, 79(3): 309–311, May 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001444.

Nehme:2013:MRQ

[2815] Rimma V. Nehme, Karen Works, Chuan Lei, Elke A. Rundensteiner, and Elisa Bertino. Multi-route query processing and optimization. Journal of Computer and System Sciences, 79(3):312–329, May 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001468.

Works:2013:OAM

[2816] Karen Works, Elke A. Rundensteiner, and Emmanuel Agu. Optimizing adaptive multi-route query processing via time-partitioned indices. Journal of Computer and System Sciences, 79 (3):330–348, May 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001200147X.

Chatzimilioudis:2013:NDF

[2817] Georgios Chatzimilioudis, Alfredo Cuzzocrea, Dimitrios Gunopulos, and Nikos Mamoulis. A novel distributed framework for optimizing query routing trees in wireless sensor networks via optimal operator placement. Journal of Computer and System Sciences, 79(3):349–368, May 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001493.

Gaber:2013:ISA

[2818] Mohamed Medhat Gaber, Shonali Krishnaswamy, Brett Gillick, Hasnain Al-Taiar, Nicholas Nicoloudis, Jonathan Liono, and Arkady Zaslavsky. Interactive self-adaptive clutter-aware visualisation for mobile data mining. Journal of Computer and System Sciences, 79(3):369–382, May 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001456.

Tang:2013:TAS

[2819] Lu-An Tang, Xiao Yu, Sangkyum Kim, Quanquan Gu, Jiawei Han, Alice Leung, and Thomas La Porta. Trustworthiness analysis of sensor data in cyber-physical systems. Journal of Computer and System Sciences, 79 (3):383–401, May 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001481.

Anonymous:2013:Cc

[2820] Anonymous. Contents. Journal of Computer and System Sciences, 79 (3):obc, May 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001300007X.

Anonymous:2013:EBc

[2821] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(3):ifc, May 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000032.

Sarbazi-Azad:2013:MCS

[2822] H. Sarbazi-Azad and N. Bagherzadeh. Multicore computing systems: Architecture, programming tools, and applications. Journal of Computer and System Sciences, 79(4): 403–405, June 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001511.

Mohaqeqi:2013:UAO

[2823] Morteza Mohaqeqi and Mehdi Kargahi. Utility accrual object distribution in MPSoC real-time embedded systems. Journal of Computer and System Sciences, 79(4): 406–420, June 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001390.

Wang:2013:SLB

[2824] Chifeng Wang, Wen-Hsiang Hu, and Nader Bagherzadeh. Scalable load balancing congestion-aware networkon-chip router architecture. Journal of Computer and System Sciences, 79(4):421–439, June 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001432.

Rahmani:2013:DPE

[2825] Amir-Mohammad Rahmani, Pasi Liljeberg, Juha Plosila, and Hannu Tenhunen. Developing a power-efficient and low-cost 3D NoC using smart GALS-based vertical channels. Journal of Computer and System Sciences, 79(4):440–456, June 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001407.

Mahabadi:2013:REA

[2826] A. Mahabadi, S. M. Zahedi, and A. Khonsari. Reliable energy-aware application mapping and voltagefrequency island partitioning for GALS-based NoC. Journal of Computer and System Sciences, 79(4): 457–474, June 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001420.

Ebrahimi:2013:CBT

[2827] Masoumeh Ebrahimi, Masoud Daneshtalab, Pasi Liljeberg, Juha Plosila, and Hannu Tenhunen. Clusterbased topologies for 3D Networkson-Chip using advanced inter-layer bus architecture. Journal of Computer and System Sciences, 79(4): 475–491, June 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001419.

Arjomand:2013:EGB

[2828] Mohammad Arjomand, S. Hamid Amiri, and Hamid Sarbazi-Azad. Efficient genetic based topological mapping using analytical models for onchip networks. Journal of Computer and System Sciences, 79(4): 492–513, June 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001200150X.

Anonymous:2013:Cd

[2829] Anonymous. Contents. Journal of Computer and System Sciences, 79 (4):obc, June 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300038X.

Anonymous:2013:EBd

[2830] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(4):ifc, June 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000342.

Wang:2013:SIJ

[2831] Guojun Wang, Jianhua Ma, Xiaolin (Andy) Li, and Athanasios V. Vasilakos. Special issue of JCSS on UbiSafe computing and communications. Journal of Computer and System Sciences, 79(5): 515–517, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001699.

Qiu:2013:SAO

[2832] Meikang Qiu, Lei Zhang, Zhong Ming, Zhi Chen, Xiao Qin, and Laurence T. Yang. Security-aware optimization for ubiquitous computing systems with SEAT graph approach. Journal of Computer and System Sciences, 79 (5):518–529, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001705.

Belwal:2013:FIT

[2833] Chaitanya Belwal, Albert M. K.

Cheng, and Bo Liu. Feasibility interval for the transactional event handlers of P-FRP. Journal of Computer and System Sciences, 79(5): 530–541, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001742.

Zhang:2013:EQI

[2834] Xuyun Zhang, Chang Liu, Surya Nepal, and Jinjun Chen. An efficient quasi-identifier index based approach for privacy preservation over incremental data sets on cloud. Journal of Computer and System Sciences, 79(5):542–555, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001766.

Yan:2013:PRS

[2835] Zheng Yan, Yu Chen, and Yue Shen. A practical reputation system for pervasive social chatting. Journal of Computer and System Sciences, 79 (5):556–572, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001717.

Zhang:2013:UTC

[2836] Jun Zhang, Yang Xiang, Wanlei Zhou, and Yu Wang. Unsupervised traffic classification using flow statistical properties and IP packet payload. Journal of Computer and System Sciences, 79(5): 573–585, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001729.

Peng:2013:MDW

[2837] Sancheng Peng, Guojun Wang, and Shui Yu. Modeling the dynamics of worm propagation using twodimensional cellular automata in smartphones. Journal of Computer and System Sciences, 79(5): 586–595, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001754.

Su:2013:RTM

[2838] Xing Su, Minjie Zhang, Yi Mu, and Quan Bai. A robust trust model for service-oriented systems. Journal of Computer and System Sciences, 79 (5):596–608, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001730.

Wang:2013:PGV

[2839] Jianxin Wang, Yongjie Yang, Jiong Guo, and Jianer Chen. Planar graph vertex partition for linear problem kernels. Journal of Computer and System Sciences, 79(5): 609–621, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000012001389.

Freydenberger:2013:IDG

[2840] Dominik D. Freydenberger and Daniel Reidenbach. Inferring descriptive generalisations of formal languages. Journal of Computer and System Sciences, 79(5):622–639, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001687.

Chatterjee:2013:SIC

[2841] Krishnendu Chatterjee, Luca de Alfaro, and Thomas A. Henzinger. Strategy improvement for concurrent reachability and turn-based stochastic safety games. Journal of Computer and System Sciences, 79(5): 640–657, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000012001778.

Alazemi:2013:SEP

[2842] Hamed M. K. Alazemi and Anton Cerný. Several extensions of the Parikh matrix L-morphism. Journal of Computer and System Sciences, 79 (5):658–668, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000287.

Chang:2013:EHL

[2843] Nai-Wen Chang and Sun-Yuan Hsieh. {2, 3}-extraconnectivities of hypercubelike networks. Journal of Computer and System Sciences, 79(5): 669–688, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000238.

Demri:2013:SUV

[2844] Stéphane Demri. On selective unboundedness of VASS. Journal of Computer and System Sciences, 79 (5):689–713, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300024X.

Bournez:2013:CPD

[2845] Olivier Bournez, Daniel S. Graça, and Emmanuel Hainry. Computation with perturbed dynamical systems. Journal of Computer and System Sciences, 79(5):714–724, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000469.

Khandekar:2013:SND

[2846] Rohit Khandekar, Guy Kortsarz, and Zeev Nutov. On some network design problems with degree constraints. Journal of Computer and System Sciences, 79(5):725– 736, August 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000299.

Anonymous:2013:Ce

[2847] Anonymous.  Contents.  Journal

of Computer and System Sciences, 79(5):obc, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000548.

Anonymous:2013:EBe

[2848] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(5):ifc, August 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000500.

Barcelo:2013:EF

[2849] Pablo Barceló. Editor’s foreword. Journal of Computer and System Sciences, 79(6):737–738, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000013000135.

Abiteboul:2013:EDS

[2850] Serge Abiteboul, Balder ten Cate, and Yannis Katsis. On the equivalence of distributed systems with queries and communication. Journal of Computer and System Sciences, 79(6): 739–762, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000111.

Arenas:2013:LPT

[2851] Marcelo Arenas, Jorge Pérez, Juan Reutter, and Cristian Riveros. The language of plain SO-tgds: Composition, inversion and structural properties. Journal of Computer and System Sciences, 79(6):763–784, September 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000123.

Bojanczyk:2013:SXD

[2852] Mikolaj Bojańczyk, Leszek A. Kolodziejczyk, and Filip Murlak. Solutions in XML data exchange. Journal of Computer and System Sciences, 79(6): 785–815, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000147.

Calvanese:2013:SSM

[2853] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. On simplification of schema mappings. Journal of Computer and System Sciences, 79(6): 816–834, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000159.

Kharlamov:2013:CMB

[2854] Evgeny Kharlamov, Dmitriy Zheleznyakov, and Diego Calvanese. Capturing model-based ontology evolution at the instance level: The case of DL-Lite. Journal of Computer and System Sciences, 79(6):835–872, September 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000160.

Deng:2013:APS

[2855] Ting Deng, Wenfei Fan, Leonid Libkin, and Yinghui Wu. On the aggregation problem for synthesized Web services. Journal of Computer and System Sciences, 79(6): 873–891, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000172.

Freydenberger:2013:ESA

[2856] Dominik D. Freydenberger and Nicole Schweikardt. Expressiveness and static analysis of extended conjunctive regular path queries. Journal of Computer and System Sciences, 79(6): 892–909, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000184.

Gelade:2013:SXS

[2857] Wouter Gelade, Tomasz Idziaszek, Wim Martens, Frank Neven, and Jan Paredaens. Simplifying XML schema: Single-type approximations of regular tree languages. Journal of Computer and System Sciences, 79(6): 910–936, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000196.

Meier:2013:SQO

[2858] Michael Meier, Michael Schmidt, Fang Wei, and Georg Lausen. Semantic query optimization in the presence of types. Journal of Computer and System Sciences, 79(6): 937–957, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000202.

Maslowski:2013:DCC

[2859] Dany Maslowski and Jef Wijsen. A dichotomy in the complexity of counting database repairs. Journal of Computer and System Sciences, 79(6): 958–983, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000214.

Pichler:2013:TCA

[2860] Reinhard Pichler and Sebastian Skritek. Tractable counting of the answers to conjunctive queries. Journal of Computer and System Sciences, 79(6): 984–1001, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000226.

Anonymous:2013:Cf

[2861] Anonymous. Contents. Journal of Computer and System Sciences, 79 (6):obc, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000858.

Anonymous:2013:EBf

[2862] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(6):ifc, September 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000810.

Enokido:2013:SII

[2863] Tomoya Enokido, Leonard Barolli, and Makoto Takizawa. Special issue on intelligent and innovative information systems. Journal of Computer and System Sciences, 79(7):

1003–1004, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000251.

Ng:2013:ESS

[2864] Joseph Kee-Yin Ng, Kam-Yiu Lam, Quan Jia Cheng, and Kevin Chin Yiu Shum. An effective signal strengthbased wireless location estimation system for tracking indoor mobile users. Journal of Computer and System Sciences, 79(7):1005– 1016, November 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000263.

Taniar:2013:TNN

[2865] David Taniar and Wenny Rahayu. A taxonomy for nearest neighbour queries in spatial databases. Journal of Computer and System Sciences, 79(7): 1017–1039, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000275.

Fazal:2013:IAA

[2866] Anhar Fazal, Farookh Khadeer Hussain, and Tharam S. Dillon. An innovative approach for automatically grading spelling in essays using rubricbased scoring. Journal of Computer and System Sciences, 79(7): 1040–1056, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000317.

Casado:2013:MDC

[2867] Rubén Casado, Muhammad Younas, and Javier Tuya. Multi-dimensional criteria for testing web services transactions. Journal of Computer and System Sciences, 79(7):1057– 1076, November 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000305.

Uehara:2013:AOR

[2868] Minoru Uehara. Advanced orthogonal RAID based on virtual disks. Journal of Computer and System Sciences, 79(7):1077–1085, November 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000470.

Xhafa:2013:ESS

[2869] Fatos Xhafa, Xavier Herrero, Admir Barolli, Leonard Barolli, and Makoto Takizawa. Evaluation of struggle strategy in genetic algorithms for ground stations scheduling problem. Journal of Computer and System Sciences, 79(7): 1086–1100, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000445.

Koyama:2013:ROO

[2870] Akio Koyama and Hiroyuki Suzuki.

Real object-oriented communication method for ad hoc networks. Journal of Computer and System Sciences, 79(7): 1101–1112, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000457.

Mullner:2013:CDR

[2871] Nils Müllner, Oliver Theel, and Martin Fränzle. Combining decomposition and reduction for state space analysis of a self-stabilizing system. Journal of Computer and System Sciences, 79(7): 1113–1125, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000329.

McCann:2013:FTC

[2872] Mark McCann and Nicholas Pippenger. Fault tolerance in cellular automata at low fault rates. Journal of Computer and System Sciences, 79(7): 1126–1143, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000482.

Egidi:2013:BSS

[2873] Lavinia Egidi and Giovanni Manzini. Better spaced seeds using quadratic residues. Journal of Computer and System Sciences, 79(7):1144– 1155, November 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000664.

Souza:2013:RCG

[2874] Uéverton dos Santos Souza, Fábio Protti, and Maise Dantas da Silva. Revisiting the complexity of and/or graph solution. Journal of Computer and System Sciences, 79(7): 1156–1163, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000779.

Kowalski:2013:LEA

[2875] Dariusz R. Kowalski and Andrzej Pelc.

Leader election in ad hoc radio networks: a keen ear helps. Journal of Computer and System Sciences, 79(7): 1164–1180, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000792.

Feng:2013:MCQ

[2876] Yuan Feng, Nengkun Yu, and Mingsheng Ying. Model checking quantum Markov chains. Journal of Computer and System Sciences, 79(7): 1181–1198, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000780.

Anonymous:2013:Cg

[2877] Anonymous. Contents. Journal of Computer and System Sciences, 79 (7):ibc, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001013.

Anonymous:2013:Ch

[2878] Anonymous. Contents. Journal of Computer and System Sciences, 79 (7):obc, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001025.

Anonymous:2013:EBg

[2879] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(7):ifc, November 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000986.

Shen:2013:ATC

[2880] Yanming Shen, Keqiu Li, and Weisong Shi. Advanced topics on cloud computing. Journal of Computer and System Sciences, 79(8): 1199, December 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000603.

Wu:2013:HAN

[2881] Wei Wu, Jianying Zhou, Yang Xiang, and Li Xu. How to achieve nonrepudiation of origin with privacy protection in cloud computing. Journal of Computer and System Sciences, 79(8): 1200–1213, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000640.

Shi:2013:DOA

[2882] Xuanhua Shi, Hongbo Jiang, Ligang He, Hai Jin, Chonggang Wang, Bo Yu, and Xueguang Chen. Developing an optimized application hosting framework in Clouds. Journal of Computer and System Sciences, 79(8): 1214–1229, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000615.

Gao:2013:MOA

[2883] Yongqiang Gao, Haibing Guan, Zhengwei Qi, Yang Hou, and Liang Liu. A multi-objective ant colony system algorithm for virtual machine placement in cloud computing. Journal of Computer and System Sciences, 79(8): 1230–1242, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000627.

Ahmed:2013:SRD

[2884] Waseem Ahmed and Yong Wei Wu. A survey on reliability in distributed systems. Journal of Computer and System Sciences, 79(8): 1243–1255, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000652.

Wu:2013:PMH

[2885] Xingfu Wu and Valerie Taylor. Performance modeling of hybrid MPI/ OpenMP scientific applications on large-scale multicore supercomputers. Journal of Computer and System Sciences, 79(8):1256–1268, December 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000639.

Stewart:2013:MNT

[2886] Iain A. Stewart. Multiswapped networks and their topological and algorithmic properties. Journal of Computer and System Sciences, 79(8): 1269–1286, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001116.

Mann:2013:ACC

[2887] Zoltán Ádám Mann and Anikó Szajkó. Average-case complexity of backtrack search for coloring sparse random graphs. Journal of Computer and System Sciences, 79(8):1287– 1301, December 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001128.

Benedikt:2013:BRW

[2888] Michael Benedikt, Gabriele Puppis, and Cristian Riveros. Bounded repairability of word languages. Journal of Computer and System Sciences, 79(8):1302–1321, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000013001104.

Kenthapadi:2013:DLI

[2889] Krishnaram Kenthapadi, Nina Mishra, and Kobbi Nissim. Denials leak information: Simulatable auditing. Journal of Computer and System Sciences, 79(8):1322–1340, December 2013. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300113X.

Caballer:2013:EEC

[2890] Miguel Caballer, Carlos de Alfonso, Fernando Alvarruiz, and Germán Moltó. EC3: Elastic Cloud Computing Cluster. Journal of Computer and System Sciences, 79(8): 1341–1351, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001141.

Anonymous:2013:Ci

[2891] Anonymous. Contents. Journal of Computer and System Sciences, 79 (8):obc, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001335.

Anonymous:2013:EBh

[2892] Anonymous. Editorial Board. Journal of Computer and System Sciences, 79(8):ifc, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001293.

Ventura:2014:FID

[2893] S. Ventura, C. Romero, and A. Abraham. Foreword: Intelligent data analysis. Journal of Computer and System Sciences, 80(1):1–2, February 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000676.

Baena-Garcia:2014:SAS

[2894] Manuel Baena-García, José M. Carmona-Cejudo, and Rafael Morales-Bueno. String analysis by sliding positioning strategy. Journal of Computer and System Sciences, 80(1):3– 15, February 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000688.

Shelokar:2014:TOS

[2895] Prakash Shelokar, Arnaud Quirin, and Óscar Cordón. Three-objective subgraph mining using multiobjective evolutionary programming. Journal of Computer and System Sciences, 80 (1):16–26, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300069X.

Smiti:2014:WDA

[2896] Abir Smiti and Zied Elouedi. WCOID-DG: an approach for case base maintenance based on weighting, clustering, outliers, internal detection and Dbsan-Gmeans. Journal of Computer and System Sciences, 80(1):27– 38, February 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000706.

Cerri:2014:HML

[2897] Ricardo Cerri, Rodrigo C. Barros, and André C. P. L. F. de Carvalho. Hierarchical multi-label classification using local neural networks. Journal of Computer and System Sciences, 80 (1):39–56, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000718.

Ortigosa:2014:PUP

[2898] Alvaro Ortigosa, Rosa M. Carro, and José Ignacio Quiroga. Predicting user personality by mining social interactions in Facebook. Journal of Computer and System Sciences, 80 (1):57–71, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300072X.

DeMeo:2014:MLG

[2899] Pasquale De Meo, Emilio Ferrara, Giacomo Fiumara, and Alessandro Provetti. Mixing local and global information for community detection in large networks. Journal of Computer and System Sciences, 80(1):72– 87, February 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000767.

Otero:2014:BAM

[2900] José Otero, Luciano Sánchez, Inés Couso, and Ana Palacios. Bootstrap analysis of multiple repetitions of experiments using an interval-valued multiple comparison procedure. Journal of Computer and System Sciences, 80 (1):88–100, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000731.

Garcia:2014:FCA

[2901] David García, Antonio González, and Raúl Pérez. A feature construction approach for genetic iterative rule learning algorithm. Journal of Computer and System Sciences, 80 (1):101–117, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000755.

Martinez-Ballesteros:2014:DGA

[2902] M. Martínez-Ballesteros, I. A. Nepomuceno-Chamorro, and J. C. Riquelme. Discovering gene association networks by multi-objective evolutionary quantitative association rules. Journal of Computer and System Sciences, 80 (1):118–136, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013000743.

Brazdil:2014:BTM

[2903] Tomás Brázdil, Václav Brozek, Vojtech Forejt, and Antonín Kucera. Branching-time model-checking of probabilistic pushdown automata. Journal of Computer and System Sciences, 80(1):139–156, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000013001396.

Fomin:2014:PCC

[2904] Fedor V. Fomin and Petr A. Golovach.

Parameterized complexity of connected even/odd subgraph problems. Journal of Computer and System Sciences, 80 (1):157–179, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001402.

Ganian:2014:LBC

[2905] Robert Ganian, Petr Hliněný, Alexander Langer, Jan Obdržálek, Peter Rossmanith, and Somnath Sikdar. Lower bounds on the complexity of MSO1 model-checking. Journal of Computer and System Sciences, 80(1): 180–194, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001438.

Xue:2014:LDP

[2906] Kaiping Xue, Peilin Hong, and Changsha Ma. A lightweight dynamic pseudonym identity based authentication and key agreement protocol without verification tables for multiserver architecture. Journal of Computer and System Sciences, 80(1): 195–206, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001426.

Giannopoulou:2014:ECI

[2907] Archontia C. Giannopoulou, Iosif Salem, and Dimitris Zoros. Effective computation of immersion obstructions for unions of graph classes. Journal of Computer and System Sciences, 80 (1):207–216, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001463.

Cai:2014:CCW

[2908] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. The complexity of complex weighted Boolean #CSP. Journal of Computer and System Sciences, 80 (1):217–236, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001414.

Wan:2014:PDM

[2909] Shuping Wan and Jiuying Dong. A possibility degree method for intervalvalued intuitionistic fuzzy multiattribute group decision making. Journal of Computer and System Sciences, 80(1):237–256, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000013001451.

Romano:2014:SRT

[2910] Paolo Romano, Roberto Palmieri, Francesco Quaglia, Nuno Carvalho, and Luis Rodrigues. On speculative replication of transactional systems. Journal of Computer and System Sciences, 80(1):257– 276, February 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300144X.

Durand:2014:CWC

[2911] Arnaud Durand and Stefan Mengel.

The complexity of weighted counting for acyclic conjunctive queries. Journal of Computer and System Sciences, 80 (1):277–296, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001475.

Goos:2014:SMT

[2912] Mika Göös, Tuomo Lempiäinen, Eugen Czeizler, and Pekka Orponen. Search methods for tile sets in patterned DNA self-assembly. Journal of Computer and System Sciences, 80 (1):297–319, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001499.

Anonymous:2014:Ca

[2913] Anonymous. Contents. Journal of Computer and System Sciences, 80 (1):obc, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001566.

Anonymous:2014:EBa

[2914] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(1):ifc, February 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001529.

Dawar:2014:EFW

[2915] Anuj Dawar and Ruy de Queiroz. Editor’s foreword: WoLLIC 2010. Journal of Computer and System Sciences, 80(2):321–322, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000901.

Abbasi:2014:ALR

[2916] Naeem Abbasi, Osman Hasan, and Sofiène Tahar. An approach for lifetime reliability analysis using theorem proving. Journal of Computer and System Sciences, 80(2): 323–345, March 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000913.

Barthwal:2014:MSC

[2917] Aditi Barthwal and Michael Norrish. A mechanisation of some context-free language theory in HOL4. Journal of Computer and System Sciences, 80 (2):346–362, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000925.

Kotek:2014:RTH

[2918] T. Kotek and J. A. Makowsky. A representation theorem for (q- )holonomic sequences. Journal of Computer and System Sciences, 80 (2):363–374, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000937.

Latte:2014:BTL

[2919] Markus Latte and Martin Lange.

Branching-time logics with path relativisation. Journal of Computer and System Sciences, 80(2):375– 389, March 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000949.

Morrill:2014:DLA

[2920] Glyn Morrill and Oriol Valentín. Displacement logic for anaphora. Journal of Computer and System Sciences, 80(2):390–409, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000950.

Reiser:2014:IVF

[2921] R. H. S. Reiser, B. C. Bedregal, and G. A. A. dos Reis. Intervalvalued fuzzy coimplications and related dual interval-valued conjugate functions. Journal of Computer and System Sciences, 80(2): 410–425, March 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013000962.

Creignou:2014:BRW

[2922] Nadia Creignou, Odile Papini, Reinhard Pichler, and Stefan Woltran. Belief revision within fragments of propositional logic. Journal of Computer and System Sciences, 80(2): 427–449, March 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001487.

Kontorovich:2014:DUD

[2923] Aryeh Kontorovich and Ari Trachtenberg. Deciding unique decodability of bigram counts via finite automata. Journal of Computer and System Sciences, 80(2): 450–456, March 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001670.

Ionescu:2014:RCI

[2924] Adrian Ionescu and Ernst L. Leiss. On the role of complementation in implicit language equations and relations. Journal of Computer and System Sciences, 80(2):457–467, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001505.

Fomin:2014:PSM

[2925] Fedor V. Fomin, Bart M. P. Jansen, and Michal Pilipczuk. Preprocessing subgraph and minor problems: When does a small vertex cover help? Journal of Computer and System Sciences, 80(2):468–495, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001682.

Hesse:2014:CUC

[2926] William Hesse, Eric Allender, and David A. Mix Barrington. Corrigendum to “Uniform constant-depth threshold circuits for division and iterated multiplication” [J. Comput. System Sci. 65 (4) (2002) 695–716]. Journal of Computer and System Sciences, 80(2):496–497, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001657. See

[1825].

Anonymous:2014:Cb

[2927] Anonymous. Contents. Journal of Computer and System Sciences, 80 (2):obc, March 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001761.

Anonymous:2014:EBb

[2928] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(2):ifc, March 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001724.

Pathan:2014:DAW

[2929] Al-Sakib Khan Pathan, Zubair Muhammad Fadlullah, Mostafa M. Fouda, and Hae Young Lee. Defending against wireless network intrusion. Journal of Computer and System Sciences, 80(3):499–501, May 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001153.

Wahid:2014:ENH

[2930] Alif Wahid, Christopher Leckie, and Chenfeng Zhou. Estimating the number of hosts corresponding to an intrusion alert while preserving privacy. Journal of Computer and System Sciences, 80(3): 502–519, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001165.

Zhao:2014:TAH

[2931] Yifan Zhao, Swee-Won Lo, Robert H. Deng, and Xuhua Ding. Technique for authenticating H.264/SVC and its performance evaluation over wireless mobile networks. Journal of Computer and System Sciences, 80 (3):520–532, May 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001177.

Yang:2014:BEB

[2932] Yang Yang. Broadcast encryption based non-interactive key distribution in MANETs. Journal of Computer and System Sciences, 80(3): 533–545, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001189.

Wang:2014:NDH

[2933] Zhiwei Wang, Guozi Sun, and Danwei Chen. A new definition of homomorphic signature for identity management in mobile cloud computing. Journal of Computer and System Sciences, 80(3):546–553, May 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001190.

Mafra:2014:ADI

[2934] P. M. Mafra, J. S. Fraga, and A. O.

Santin. Algorithms for a distributed IDS in MANETs. Journal of Computer and System Sciences, 80(3): 554–570, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001207.

Perez:2014:BRB

[2935] Manuel Gil Pérez, Félix Gómez Mármol, Gregorio Martínez Pérez, and Antonio F. Skarmeta Gómez. Building a reputation-based bootstrapping mechanism for newcomers in collaborative alert systems. Journal of Computer and System Sciences, 80 (3):571–590, May 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001219.

Rehman:2014:AIA

[2936] Saeed Ur Rehman, Kevin W. Sowerby, and Colin Coghill. Analysis of impersonation attacks on systems using RF fingerprinting and lowend receivers. Journal of Computer and System Sciences, 80(3): 591–601, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001220.

Han:2014:MAT

[2937] Guangjie Han, Jinfang Jiang, Lei Shu, Jianwei Niu, and Han-Chieh Chao. Management and applications of trust in wireless sensor networks: a survey. Journal of Computer and System Sciences, 80(3): 602–617, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001232.

Giannetsos:2014:LLD

[2938] Thanassis Giannetsos and Tassos Dimitriou. LDAC: a localized and decentralized algorithm for efficiently countering wormholes in mobile wireless networks. Journal of Computer and System Sciences, 80(3): 618–643, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001244.

Shafiei:2014:DMS

[2939] H. Shafiei, A. Khonsari, H. Derakhshi, and P. Mousavi. Detection and mitigation of sinkhole attacks in wireless sensor networks. Journal of Computer and System Sciences, 80(3):644–653, May 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000013001256.

Conti:2014:CWD

[2940] M. Conti, R. Di Pietro, and A. Spognardi. Clone wars: Distributed detection of clone attacks in mobile WSNs. Journal of Computer and System Sciences, 80(3): 654–669, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001268.

Maccari:2014:BEO

[2941] Leonardo Maccari and Renato Lo Cigno. Betweenness estimation in OLSR-based multi-hop networks for distributed filtering. Journal of Computer and System Sciences, 80(3): 670–685, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300127X.

Anonymous:2014:Cc

[2942] Anonymous. Contents. Journal of Computer and System Sciences, 80 (3):obc, May 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001992.

Anonymous:2014:EBc

[2943] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(3):ifc, May 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001955.

Crowston:2014:SMT

[2944] R. Crowston, M. Fellows, G. Gutin, M. Jones, E. J. Kim, F. Rosamond, I. Z. Ruzsa, S. Thomassé, and A. Yeo.

Satisfying more than half of a system of linear equations over GF(2): a multivariate approach. Journal of Computer and System Sciences, 80 (4):687–696, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001700.

Hemaspaandra:2014:COM

[2945] Edith Hemaspaandra, Lane A. Hemaspaandra, and Jörg Rothe. The complexity of online manipulation of sequential elections. Journal of Computer and System Sciences, 80(4): 697–710, June 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001694.

Fici:2014:SBS

[2946] Gabriele Fici. On the structure of bispecial Sturmian words. Journal of Computer and System Sciences, 80(4):711–719, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001888.

Bohm:2014:BER

[2947] Stanislav Böhm, Stefan Göller, and Petr Jancar. Bisimulation equivalence and regularity for real-time one-counter automata. Journal of Computer and System Sciences, 80 (4):720–743, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001906.

Kralovic:2014:IVF

[2948] Richard Královic. Infinite vs. finite size-bounded randomized computations. Journal of Computer and System Sciences, 80(4):744– 765, June 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001918.

Buss:2014:USN

[2949] Sam Buss and Michael Soltys. Unshuffling a square is NP-hard. Journal of Computer and System Sciences, 80(4):766–776, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001300189X.

Jain:2014:RLA

[2950] Sanjay Jain, Eric Martin, and Frank Stephan. Robust learning of automatic classes of languages. Journal of Computer and System Sciences, 80(4):777–795, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013001931.

Gomez-Rodriguez:2014:FSB

[2951] Carlos Gómez-Rodríguez. Finding the smallest binarization of a CFG is NP-hard. Journal of Computer and System Sciences, 80(4): 796–805, June 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000013002055.

Case:2014:ALF

[2952] John Case, Sanjay Jain, Yuh Shin Ong, Pavel Semukhin, and Frank Stephan. Automatic learners with feedback queries. Journal of Computer and System Sciences, 80(4): 806–820, June 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001300192X.

Tosi:2014:DCB

[2953] Stefania Tosi, Sara Casolari, and Michele Colajanni. Detecting correlation between server resources for system management. Journal of Computer and System Sciences, 80 (4):821–836, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000038.

Duan:2014:CPM

[2954] Qi Duan and Jinhui Xu. On the connectivity preserving minimum cut problem. Journal of Computer and System Sciences, 80(4): 837–848, June 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001400004X.

Civril:2014:CSS

[2955] A. Çivril. Column subset selection problem is UG-hard. Journal of Computer and System Sciences, 80 (4):849–859, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000051.

Dolev:2014:RMS

[2956] Danny Dolev, Matthias Függer, Markus Posch, Ulrich Schmid, Andreas Steininger, and Christoph Lenzen. Rigorously modeling self-stabilizing faulttolerant circuits: an ultra-robust clocking scheme for systems-on-chip. Journal of Computer and System Sciences, 80(4):860–900, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000026.

Anonymous:2014:Cd

[2957] Anonymous. Contents. Journal of Computer and System Sciences, 80 (4):obc, June 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000117.

Anonymous:2014:EBd

[2958] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(4):ifc, June 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000075.

Capone:2014:PN

[2959] Rebecca Capone. Publisher’s note.

Journal of Computer and System Sciences, 80(5):901, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000014000415.

Chen:2014:SID

[2960] Jinjun Chen and Jianxun Liu. Special issue: Dependable and secure computing. Journal of Computer and System Sciences, 80(5): 902, August 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001400021X.

Xie:2014:MMB

[2961] Yongming Xie and Guojun Wang. Message matching-based greedy behavior detection in delay tolerant networks. Journal of Computer and System Sciences, 80(5): 903–915, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000130.

Nagarajan:2014:TED

[2962] Aarthi Nagarajan, Vijay Varadharajan, and Nathan Tarr. Trust enhanced distributed authorisation for web services. Journal of Computer and System Sciences, 80(5): 916–934, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000142.

Bailey:2014:SAF

[2963] Christopher Bailey, David W. Chadwick, and Rogério de Lemos. Selfadaptive federated authorization infrastructures. Journal of Computer and System Sciences, 80(5): 935–952, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000154.

Shalan:2014:SCF

[2964] Atef Shalan and Mohammad Zulkernine. Software control flow error detection and correlation with system performance deviation. Journal of Computer and System Sciences, 80 (5):953–972, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000166.

Jang-Jaccard:2014:SET

[2965] Julian Jang-Jaccard and Surya Nepal.

A survey of emerging threats in cybersecurity. Journal of Computer and System Sciences, 80(5): 973–993, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000178.

Zhao:2014:SFG

[2966] Jiaqi Zhao, Lizhe Wang, Jie Tao, Jinjun Chen, Weiye Sun, Rajiv Ranjan, Joanna Kołodziej, Achim Streit, and Dimitrios Georgakopoulos. A security framework in G-Hadoop for big data computing across distributed Cloud data centres. Journal of Computer and System Sciences, 80(5): 994–1007, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400018X.

Zhang:2014:HAS

[2967] Xuyun Zhang, Chang Liu, Surya Nepal, Chi Yang, Wanchun Dou, and Jinjun Chen. A hybrid approach for scalable sub-tree anonymization over big data using MapReduce on cloud. Journal of Computer and System Sciences, 80(5): 1008–1020, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000191.

Wang:2014:ITC

[2968] Yu Wang, Yang Xiang, Jun Zhang, Wanlei Zhou, and Bailin Xie. Internet traffic clustering with side information. Journal of Computer and System Sciences, 80(5):1021– 1036, August 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000208.

Anonymous:2014:Ce

[2969] Anonymous.  Contents.  Journal

of Computer and System Sciences, 80(5):obc, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000348.

Anonymous:2014:EBe

[2970] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(5):ifc, August 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000300.

Anonymous:2014:IIC

[2971] Anonymous. The 9th IEEE International Conference on Dependable, Autonomic and Secure Computing. Journal of Computer and System Sciences, 80(5):??, August 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Beklemishev:2014:EF

[2972] Lev Beklemishev, Ruy De Queiroz, and Andre Scedrov. Editors’ foreword. Journal of Computer and System Sciences, 80(6):1037, September 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000439.

Achilleos:2014:CQJ

[2973] Antonis Achilleos. A complexity question in justification logic. Journal of Computer and System Sciences, 80(6): 1038–1045, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000440.

Bucheli:2014:RPA

[2974] Samuel Bucheli, Roman Kuznets, and Thomas Studer. Realizing public announcements by justifications. Journal of Computer and System Sciences, 80(6):1046–1066, September 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000452.

Clouston:2014:NLT

[2975] Ranald Clouston. Nominal Lawvere theories: a category theoretic account of equational theories with names. Journal of Computer and System Sciences, 80(6): 1067–1086, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000464.

Ferreira:2014:PHG

[2976] F. M. Ferreira, C. M. Freire, M. R. F. Benevides, L. Menasché Schechter, and A. T. Martins. Polynomial hierarchy graph properties in hybrid logic. Journal of Computer and System Sciences, 80(6):1087– 1101, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000476.

Figueira:2014:IFL

[2977] Santiago Figueira, Daniel Gorín, and Rafael Grimson. Independence friendly logic with classical negation via flattening is a second-order logic with weak dependencies. Journal of Computer and System Sciences, 80(6): 1102–1118, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000488.

Gutierrez:2014:DCG

[2978] Julian Gutierrez and Glynn Winskel. On the determinacy of concurrent games on event structures with infinite winning sets. Journal of Computer and System Sciences, 80(6): 1119–1137, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400049X.

Kanovich:2014:MRF

[2979] Max Kanovich. Multiset rewriting over Fibonacci and Tribonacci numbers. Journal of Computer and System Sciences, 80(6):1138– 1151, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000506.

Kontinen:2014:CDS

[2980] Juha Kontinen and Jakub Szymanik. A characterization of definability of second-order generalized quantifiers with applications to non-definability. Journal of Computer and System Sciences, 80(6): 1152–1162, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000518.

Statman:2014:PTU

[2981] Rick Statman. On polymorphic types of untyped terms. Journal of Computer and System Sciences, 80(6): 1163–1173, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400052X.

Anonymous:2014:Cf

[2982] Anonymous. Contents. Journal of Computer and System Sciences, 80 (6):obc, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000774.

Anonymous:2014:EBf

[2983] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(6):ifc, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000737.

Aracena:2014:MNF

[2984] J. Aracena, A. Richard, and L. Salinas. Maximum number of fixed points in AND-OR-NOT networks. Journal of Computer and System Sciences, 80(7): 1175–1190, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000695.

Ameloot:2014:PDP

[2985] Tom J. Ameloot and Jan Van den Bussche. Positive Dedalus programs tolerate non-causality. Journal of Computer and System Sciences, 80(7): 1191–1213, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000129.

Cohen:2014:AES

[2986] Edith Cohen, Nick Duffield, Haim Kaplan, Carstent Lund, and Mikkel Thorup. Algorithms and estimators for summarization of unaggregated data streams. Journal of Computer and System Sciences, 80(7): 1214–1244, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000531.

Amir:2014:RL

[2987] Amihood Amir, Alberto Apostolico, Gad M. Landau, Avivit Levy, Moshe Lewenstein, and Ely Porat. Range LCP. Journal of Computer and System Sciences, 80(7):1245–1253, November 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000221.

Lopez:2014:PSD

[2988] Jesús García López, Florin Manea, and Victor Mitrana. Prefix-suffix duplication. Journal of Computer and System Sciences, 80(7):1254– 1265, November 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000233.

Laoutaris:2014:BBC

[2989] Nikolaos Laoutaris, Laura Poplawski, Rajmohan Rajaraman, Ravi Sundaram, and Shang-Hua Teng. Bounded Budget Connection (BBC) games or how to make friends and influence people, on a budget. Journal of Computer and System Sciences, 80(7): 1266–1284, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000579.

Bazgan:2014:PCF

[2990] Cristina Bazgan, Morgan Chopin, Marek Cygan, Michael R. Fellows, Fedor V. Fomin, and Erik Jan van Leeuwen. Parameterized complexity of firefighting. Journal of Computer and System Sciences, 80(7): 1285–1297, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000245.

Broy:2014:VIA

[2991] Manfred Broy. Verifying of interface assertions for infinite state Mealy machines. Journal of Computer and System Sciences, 80(7): 1298–1322, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000257.

Talaei-Khoei:2014:MDA

[2992] Amir Talaei-Khoei, Suchada Vichitvanichphong, Terje Solvoll, Pradeep Ray, and Amir Hossein Ghapanchi. A methodology to develop awareness in computer supported collaborative work using policies. Journal of Computer and System Sciences, 80(7): 1323–1338, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000269.

Li:2014:MBI

[2993] Chunguang Li, Shouzhen Zeng, Tiejun Pan, and Leina Zheng. A method based on induced aggregation operators and distance measures to multiple attribute decision making under 2-tuple linguistic environment. Journal of Computer and System Sciences, 80(7): 1339–1349, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000270.

Jiang:2014:LKC

[2994] Haitao Jiang and Binhai Zhu. A linear kernel for the complementary maximal strip recovery problem. Journal of Computer and System Sciences, 80(7): 1350–1358, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000282.

Glasser:2014:PCB

[2995] Christian Glaßer, Stephen Travers, and Klaus W. Wagner. Perfect correspondences between dot-depth and polynomial-time hierarchies. Journal of Computer and System Sciences, 80(7): 1359–1373, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000403.

Fomin:2014:SBF

[2996] Fedor V. Fomin and Yngve Villanger.

Searching for better fill-in. Journal of Computer and System Sciences, 80(7): 1374–1383, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000543.

Mnich:2014:BMC

[2997] Matthias Mnich, Geevarghese Philip, Saket Saurabh, and Ondrej Suchý. Beyond Max-Cut: λ-extendible properties parameterized above the Poljak– Turzík bound. Journal of Computer and System Sciences, 80(7): 1384–1403, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000555.

Kianpour:2014:NDB

[2998] Moein Kianpour, Reza Sabbaghi-Nadooshan, and Keivan Navi. A novel design of 8-bit adder/subtractor by quantum-dot cellular automata. Journal of Computer and System Sciences, 80(7):1404–1414, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000014000567.

Atserias:2014:BWQ

[2999] Albert Atserias and Sergi Oliva. Bounded-width QBF is PSPACEcomplete. Journal of Computer and System Sciences, 80(7):1415– 1429, November 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000580.

Fomin:2014:TBP

[3000] Fedor V. Fomin, Stefan Kratsch, Marcin Pilipczuk, Michal Pilipczuk, and Yngve Villanger. Tight bounds for parameterized complexity of cluster editing with a small number of clusters. Journal of Computer and System Sciences, 80(7):1430– 1447, November 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000592.

Flesca:2014:CCQ

[3001] Sergio Flesca, Filippo Furfaro, and Francesco Parisi. Consistency checking and querying in probabilistic databases under integrity constraints. Journal of Computer and System Sciences, 80(7): 1448–1489, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000701.

Anonymous:2014:Cg

[3002] Anonymous. Contents. Journal of Computer and System Sciences, 80 (7):obc, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000841.

Anonymous:2014:EBg

[3003] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(7):ifc, November 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000804.

Xiang:2014:ESI

[3004] Yang Xiang, Ivan Stojmenovic, and Peter Mueller. Editorial: Special issue on theory and applications in parallel and distributed computing systems. Journal of Computer and System Sciences, 80(8): 1491–1492, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000609.

Wang:2014:BAC

[3005] Hua Wang, Lili Sun, and Elisa Bertino. Building access control policy model for privacy preserving and testing policy conflicting problems. Journal of Computer and System Sciences, 80(8): 1493–1503, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000610.

Sallam:2014:PWM

[3006] Ahmed Sallam, Kenli Li, Aijia Ouyang, and Zhiyong Li. Proactive workload management in dynamic virtualized environments. Journal of Computer and System Sciences, 80(8): 1504–1517, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000622.

Zhou:2014:SAC

[3007] Lan Zhou, Vijay Varadharajan, and Michael Hitchens. Secure administration of cryptographic role-based access control for large-scale cloud storage systems. Journal of Computer and System Sciences, 80(8): 1518–1533, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000634.

Chen:2014:DSW

[3008] Zhide Chen, Cheng Qiao, Yihui Qiu, Li Xu, and Wei Wu. Dynamics stability in wireless sensor networks active defense model. Journal of Computer and System Sciences, 80(8): 1534–1548, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000646.

Wei:2014:EEF

[3009] Guiyi Wei, Rongxing Lu, and Jun Shao. EFADS: Efficient, flexible and anonymous data sharing protocol for cloud computing with proxy re-encryption. Journal of Computer and System Sciences, 80(8): 1549–1562, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000658.

Yang:2014:SCB

[3010] Chi Yang, Xuyun Zhang, Changmin Zhong, Chang Liu, Jian Pei, Kotagiri Ramamohanarao, and Jinjun Chen. A spatiotemporal compression based approach for efficient big data processing on Cloud. Journal of Computer and System Sciences, 80(8): 1563–1583, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001400066X.

Jung:2014:OTP

[3011] Younchan Jung and Enrique Festijo. One-time packet key exchange scheme for secure real-time multimedia applications. Journal of Computer and System Sciences, 80(8): 1584–1596, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000671.

Gambs:2014:AAG

[3012] Sébastien Gambs, Marc-Olivier Killijian, and Miguel Núñez del Prado Cortez. De-anonymization attack on geolocated data. Journal of Computer and System Sciences, 80(8):1597– 1614, December 2014. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000683.

Anonymous:2014:EBh

[3013] Anonymous. Editorial Board. Journal of Computer and System Sciences, 80(8):ifc, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001400110X.

Hu:2015:ISI

[3014] Jia Hu, Seetharami Seelam, and Laurent Lefevre. Introduction to special issue on high performance computing architectures and systems. Journal of Computer and System Sciences, 81(1):1–2, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001184.

Liao:2015:NMA

[3015] Xiaofei Liao, Hai Jin, Shizhan Yu, and Yu Zhang. A novel memory allocation scheme for memory energy reduction in virtualization environment. Journal of Computer and System Sciences, 81(1):3–15, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001056.

Paulino:2015:HPS

[3016] Hervé Paulino and Eduardo Marques.

Heterogeneous programming with Single Operation Multiple Data. Journal of Computer and System Sciences, 81(1):16–37, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001081.

Soner:2015:IPB

[3017] Seren Soner and Can Özturan. Integer programming based heterogeneous CPU-GPU cluster schedulers for SLURM resource manager. Journal of Computer and System Sciences, 81 (1):38–56, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000981.

Ma:2015:MSU

[3018] Jianliang Ma, Licheng Yu, John M. Ye, and Tianzhou Chen. MCMG simulator: a unified simulation framework for CPU and graphic GPU. Journal of Computer and System Sciences, 81 (1):57–71, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001044.

Chakroun:2015:THA

[3019] I. Chakroun and N. Melab. Towards a heterogeneous and adaptive parallel branch-and-bound algorithm. Journal of Computer and System Sciences, 81(1):72–84, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000993.

Al-Aqrabi:2015:CBF

[3020] Hussain Al-Aqrabi, Lu Liu, Richard Hill, and Nick Antonopoulos. Cloud BI: Future of business intelligence in the Cloud. Journal of Computer and System Sciences, 81(1):85– 96, February 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001007.

Khalaf:2015:NEV

[3021] Mustafa Bani Khalaf, Ahmed Y. Al-Dubai, and Geyong Min. New efficient velocity-aware probabilistic route discovery schemes for high mobility ad hoc networks. Journal of Computer and System Sciences, 81(1):

97–109, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001019.

Wang:2015:TDR

[3022] Jianjiang Wang, Xiaomin Zhu, Laurence T. Yang, Jianghan Zhu, and Manhao Ma. Towards dynamic realtime scheduling for multiple earth observation satellites. Journal of Computer and System Sciences, 81(1): 110–124, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001032.

Zheng:2015:SPS

[3023] Minghui Zheng, Yang Xiang, and Huihua Zhou. A strong provably secure IBE scheme without bilinear map. Journal of Computer and System Sciences, 81(1): 125–131, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400107X. See

comment [3187].

Li:2015:TTD

[3024] Wei Li, Demetres D. Kouvatsos, and Rod J. Fretwell. Towards time domain invariant QoS measures for queues with correlated traffic. Journal of Computer and System Sciences, 81 (1):132–144, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001068.

Xu:2015:ORV

[3025] Yuwei Xu, Jeremiah D. Deng, Mariusz Nowostawski, and Martin K. Purvis. Optimized routing for video streaming in multi-hop wireless networks using analytical capacity estimation. Journal of Computer and System Sciences, 81 (1):145–157, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001020.

Berenbrink:2015:RDI

[3026] Petra Berenbrink, Colin Cooper, Tom Friedetzky, Tobias Friedrich, and Thomas Sauerwald. Randomized diffusion for indivisible loads. Journal of Computer and System Sciences, 81 (1):159–185, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000713.

Kim:2015:SEF

[3027] Eun Jung Kim, Christophe Paul, and Geevarghese Philip. A singleexponential FPT algorithm for the K4-minor cover problem. Journal of Computer and System Sciences, 81 (1):186–207, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000786.

Zhai:2015:DIC

[3028] Yanhui Zhai, Deyu Li, and Kaishe Qu. Decision implication canonical basis: a logical perspective. Journal of Computer and System Sciences, 81 (1):208–218, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000877.

Clementi:2015:PFG

[3029] Andrea Clementi and Riccardo Silvestri. Parsimonious flooding in geometric random-walks. Journal of Computer and System Sciences, 81 (1):219–233, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000889.

Bampas:2015:NVR

[3030] Evangelos Bampas, Davide Bilò, Guido Drovandi, Luciano Gualà, Ralf Klasing, and Guido Proietti. Network verification via routing table queries. Journal of Computer and System Sciences, 81 (1):234–248, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000890.

Kortelainen:2015:SGR

[3031] Juha Kortelainen. On simple generators of recursively enumerable languages. Journal of Computer and System Sciences, 81(1):249– 257, February 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000907.

Baier:2015:LPK

[3032] Christel Baier, Marcus Daum, Benjamin Engel, Hermann Härtig, Joachim Klein, Sascha Klüppelholz, Steffen Märcker, Hendrik Tews, and Marcus Völp. Locks: Picking key methods for a scalable quantitative analysis. Journal of Computer and System Sciences, 81 (1):258–287, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000919.

Brazdil:2015:RAP

[3033] Tomás Brázdil, Stefan Kiefer, Antonín Kucera, and Ivana Hutarová Vareková. Runtime analysis of probabilistic programs with unbounded recursion. Journal of Computer and System Sciences, 81(1):288– 310, February 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000920.

Chen:2015:CAC

[3034] Xi Chen, Martin Dyer, Leslie Ann Goldberg, Mark Jerrum, Pinyan Lu, Colin McQuillan, and David Richerby. The complexity of approximating conservative counting CSPs. Journal of Computer and System Sciences, 81 (1):311–329, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000932.

Goldberg:2015:APF

[3035] Leslie Ann Goldberg, Mark Jerrum, and Colin McQuillan. Approximating the partition function of planar two-state spin systems. Journal of Computer and System Sciences, 81 (1):330–358, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000944.

Anonymous:2015:Ca

[3036] Anonymous. Contents. Journal of Computer and System Sciences, 81 (1):obc, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001354.

Anonymous:2015:EBa

[3037] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(1):ifc, February 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001202.

Qiu:2015:EMC

[3038] Daowen Qiu, Lvzhou Li, Paulo Mateus, and Amilcar Sernadas. Exponentially more concise quantum recognition of non-RMM regular languages. Journal of Computer and System Sciences, 81(2):359–375, March 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014000956.

Guan:2015:TCC

[3039] Xuechong Guan. Two Cartesian closed categories of information algebras. Journal of Computer and System Sciences, 81(2): 376–382, March 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014000968.

Lafuente:2015:COE

[3040] Alberto Lafuente, Mikel Larrea, Iratxe Soraluze, and Roberto Cortiñas. Communication-optimal eventually perfect failure detection in partially synchronous systems. Journal of Computer and System Sciences, 81 (2):383–397, March 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001400097X.

De:2015:PSL

[3041] Minati De, Subhas C. Nandy, and Sasanka Roy. Prune-and-search with limited workspace. Journal of Computer and System Sciences, 81(2): 398–414, March 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001172.

Torma:2015:UEC

[3042] Ilkka Törmä. A uniquely ergodic cellular automaton. Journal of Computer and System Sciences, 81(2): 415–442, March 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001366.

Bova:2015:CEE

[3043] Simone Bova and Hubie Chen. The complexity of equivalence, entailment, and minimization in existential positive logic. Journal of Computer and System Sciences, 81(2): 443–457, March 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001378.

Possignolo:2015:QAQ

[3044] Rafael Trapani Possignolo, Cintia Borges Margi, and Paulo S. L. M. Barreto. Quantum-assisted QD-CFS signatures. Journal of Computer and System Sciences, 81(2): 458–467, March 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400138X.

Honkala:2015:PMR

[3045] Juha Honkala. Products of matrices and recursively enumerable sets. Journal of Computer and System Sciences, 81(2):468–472, March 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001391.

Sosik:2015:LCD

[3046] Petr Sosík and Ludek Cienciala. A limitation of cell division in tissue P systems by PSPACE. Journal of Computer and System Sciences, 81 (2):473–484, March 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400141X.

Anonymous:2015:EBb

[3047] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(2):ifc, March 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400155X.

Sheng:2015:ACA

[3048] Quan Z. Sheng and Elhadi M. Shakshuki. Ambient and contextaware services. Journal of Computer and System Sciences, 81(3): 485–486, May 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001457.

Sahli:2015:ABF

[3049] Nabil Sahli, Nafaa Jabeura, and Mohamad Badra. Agent-based framework for sensor-to-sensor personalization. Journal of Computer and System Sciences, 81(3):487– 495, May 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001469.

Abid:2015:SDR

[3050] Mohamed Amine Abid, Abdelfettah Belghith, and Khalil Drira. SARP: a dynamically readjustable period size proactive routing protocol for MANETs. Journal of Computer and System Sciences, 81(3): 496–515, May 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001470.

Alkhdour:2015:CLO

[3051] Tayseer Alkhdour, Elhadi Shakshuki, Uthman Baroudi, and Shokri Selim. A cross layer optimization modeling for a periodic WSN application. Journal of Computer and System Sciences, 81(3):516–532, May 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001482.

Yu:2015:MDD

[3052] Jian Yu, Quan Z. Sheng, Joshua K. Y.

Swee, Jun Han, Chengfei Liu, and Talal H. Noor. Model-driven development of adaptive web service processes with aspects and rules. Journal of Computer and System Sciences, 81(3):533–552, May 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001494.

Gunn:2015:DHT

[3053] Tyler Gunn and John Anderson.

Dynamic heterogeneous team formation for robotic urban search and rescue. Journal of Computer and System Sciences, 81(3): 553–567, May 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001500.

Knapen:2015:SIO

[3054] Luk Knapen, Irith Ben-Arroyo Hartman, Daniel Keren, Ansar-Ul-Haque Yasar, Sungjin Cho, Tom Bellemans, Davy Janssens, and Geert Wets. Scalability issues in optimal assignment for carpooling. Journal of Computer and System Sciences, 81(3): 568–584, May 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001512.

Lebrun:2015:MMI

[3055] Yoann Lebrun, Emmanuel Adam, René Mandiau, and Christophe Kolski. A model for managing interactions between tangible and virtual agents on an RFID interactive tabletop: Case study in traffic simulation. Journal of Computer and System Sciences, 81(3):585–598, May 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001524.

Yang:2015:PPM

[3056] Liang Huai Yang, Yanzhu Zhao, Yulei Fan, Yihua Zhu, and Jian Yu. Peak power modeling for join algorithms in DBMS. Journal of Computer and System Sciences, 81 (3):599–614, May 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001536.

Anonymous:2015:Cb

[3057] Anonymous. Contents. Journal of Computer and System Sciences, 81 (3):obc, May 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001974.

Anonymous:2015:EBc

[3058] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(3):ifc, May 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001937.

Shi:2015:FSE

[3059] Q. Shi, N. Zhang, and M. Merabti.

Fair signature exchange via delegation on ubiquitous networks. Journal of Computer and System Sciences, 81(4):615–631, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001421.

Erdelyi:2015:CCBa

[3060] Gábor Erdélyi, Michael R. Fellows, Jörg Rothe, and Lena Schend. Control complexity in Bucklin and fallback voting: A theoretical analysis. Journal of Computer and System Sciences, 81(4):632–660, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001433.

Erdelyi:2015:CCBb

[3061] Gábor Erdélyi, Michael R. Fellows, Jörg Rothe, and Lena Schend. Control complexity in Bucklin and fallback voting: An experimental analysis. Journal of Computer and System Sciences, 81(4):661–670, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001445.

Schwabe:2015:OLD

[3062] Eric J. Schwabe and Ian M. Sutherland. Optimizing large data transfers in parity-declustered data layouts. Journal of Computer and System Sciences, 81(4):671–691, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001603.

Alon:2015:PSS

[3063] Noga Alon, Hagit Attiya, Shlomi Dolev, Swan Dubois, Maria Potop-Butucaru, and Sébastien Tixeuil. Practically stabilizing SWMR atomic memory in message-passing systems. Journal of Computer and System Sciences, 81(4):692–701, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001615.

Jerrum:2015:PCC

[3064] Mark Jerrum and Kitty Meeks. The parameterised complexity of counting connected subgraphs and graph motifs. Journal of Computer and System Sciences, 81(4): 702–716, June 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001627.

Zhou:2015:SBI

[3065] Xun Zhou, Jing He, Guangyan Huang, and Yanchun Zhang. SVD-based incremental approaches for recommender systems. Journal of Computer and System Sciences, 81(4):

717–733, June 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001706.

Wang:2015:FAC

[3066] Lijuan Wang, Jun Shen, and Junzhou Luo. Facilitating an ant colony algorithm for multi-objective dataintensive service provision. Journal of Computer and System Sciences, 81(4):734–746, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001718.

Fernau:2015:MPA

[3067] Henning Fernau, Pinar Heggernes, and Yngve Villanger. A multi-parameter analysis of hard problems on deterministic finite automata. Journal of Computer and System Sciences, 81 (4):747–765, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001913.

Bredereck:2015:EIV

[3068] Robert Bredereck, Jiehua Chen, Sepp Hartung, Christian Komusiewicz, Rolf Niedermeier, and Ondřej Suchý. On explaining integer vectors by few homogeneous segments. Journal of Computer and System Sciences, 81 (4):766–782, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001986.

Zhong:2015:NLM

[3069] Jianghua Zhong and Dongdai Lin. A new linearization method for nonlinear feedback shift registers. Journal of Computer and System Sciences, 81(4):783–796, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014002001.

Anonymous:2015:EBd

[3070] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(4):ifc, June 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000033.

Pattinson:2015:E

[3071] Dirk Pattinson and Lutz Schröder. From the Editors. Journal of Computer and System Sciences, 81(5): 797–798, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000136.

Abramsky:2015:LBK

[3072] Samson Abramsky and Jonathan Zvesper. From Lawvere to Brandenburger– Keisler: Interactive forms of diagonalization and self-reference. Journal of Computer and System Sciences, 81 (5):799–812, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001639.

Adamek:2015:FFT

[3073] Jirí Adámek, Stefan Milius, Lawrence S.

Moss, and Henning Urbat. On finitary functors and their presentations. Journal of Computer and System Sciences, 81(5):813–833, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001640.

Bacci:2015:SOS

[3074] Giorgio Bacci and Marino Miculan. Structural operational semantics for continuous state stochastic transition systems. Journal of Computer and System Sciences, 81(5): 834–858, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001652.

Jacobs:2015:TSD

[3075] Bart Jacobs, Alexandra Silva, and Ana Sokolova. Trace semantics via determinization. Journal of Computer and System Sciences, 81(5): 859–879, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001676.

Marti:2015:LEC

[3076] Johannes Marti and Yde Venema.

Lax extensions of coalgebra functors and their logic. Journal of Computer and System Sciences, 81 (5):880–900, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001688.

Sobocinski:2015:RPC

[3077] Pawel Sobociński. Relational presheaves, change of base and weak simulation. Journal of Computer and System Sciences, 81(5):901– 910, August 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001400169X.

Winter:2015:CFC

[3078] Joost Winter, Marcello M. Bonsangue, and Jan J. M. M. Rutten. Contextfree coalgebras. Journal of Computer and System Sciences, 81(5): 911–939, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001664.

Anonymous:2015:EBe

[3079] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(5):ifc, August 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001500015X.

James:2015:OSP

[3080] Anne James. Optimisation, security, privacy and trust in ebusiness systems. Journal of Computer and System Sciences, 81(6): 941–942, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400172X.

Abraham:2015:PGM

[3081] Goodhead Tomvie Abraham, Anne James, and Norlaily Yaacob. Prioritygrouping method for parallel multischeduling in grid. Journal of Computer and System Sciences, 81(6): 943–957, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001731.

Al-Daraiseh:2015:MAS

[3082] Ahmad Al-Daraiseh, Eyas El-Qawasmeh, and Nazaraf Shah. Multi-agent system for energy consumption optimisation in higher education institutions. Journal of Computer and System Sciences, 81 (6):958–965, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001743.

Tsai:2015:DGP

[3083] Chen-Fang Tsai. Dynamic grey platform for efficient forecasting management. Journal of Computer and System Sciences, 81(6):966–980, September 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001755.

Bulajoul:2015:INI

[3084] Waleed Bul’ajoul, Anne James, and Mandeep Pannu. Improving network intrusion detection system performance through quality of service configuration and parallel technology. Journal of Computer and System Sciences, 81 (6):981–999, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001767.

Casino:2015:AAP

[3085] Fran Casino, Josep Domingo-Ferrer, Constantinos Patsakis, Domènec Puig, and Agusti Solanas. A k-anonymous approach to privacy preserving collaborative filtering. Journal of Computer and System Sciences, 81(6): 1000–1011, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001779.

Wang:2015:MBD

[3086] Ping Wang and Yu-Shih Wang. Malware behavioural detection and vaccine development by using a support vector model classifier. Journal of Computer and System Sciences, 81(6): 1012–1026, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001780.

Chang:2015:VAG

[3087] Shan Chang, Li Lu, Xiaoqiang Liu, Hui Song, and Qingsong Yao. Vulnerability aware graphs for RFID protocol security benchmarking. Journal of Computer and System Sciences, 81(6): 1027–1041, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001792.

Kumar:2015:IAB

[3088] Neeraj Kumar, Rahat Iqbal, Sudip Misra, and Joel J. P. C. Rodrigues. An intelligent approach for building a secure decentralized public key infrastructure in VANET. Journal of Computer and System Sciences, 81(6): 1042–1058, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001809.

Anonymous:2015:Cc

[3089] Anonymous. Contents. Journal of Computer and System Sciences, 81 (6):obc, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000392.

Anonymous:2015:EBf

[3090] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(6):ifc, September 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000318.

Almarza:2015:NNI

[3091] Javier Ignacio Almarza and Santiago Figueira. Normality in noninteger bases and polynomial time randomness. Journal of Computer and System Sciences, 81(7):1059– 1087, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000434.

Augustine:2015:DAD

[3092] John Augustine, Gopal Pandurangan, Peter Robinson, and Eli Upfal. Distributed agreement in dynamic peerto-peer networks. Journal of Computer and System Sciences, 81(7): 1088–1109, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001408.

Shambour:2015:ERS

[3093] Qusai Shambour and Jie Lu. An effective recommender system by unifying user and item trust information for B2B applications. Journal of Computer and System Sciences, 81(7): 1110–1126, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001998.

Cohen:2015:VVE

[3094] David A. Cohen, Martin C. Cooper, Guillaume Escamocher, and Stanislav Zivný. Variable and value elimination in binary constraint satisfaction via forbidden patterns. Journal of Computer and System Sciences, 81(7): 1127–1143, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000082.

Li:2015:HMQ

[3095] Lvzhou Li and Yuan Feng. On hybrid models of quantum finite automata. Journal of Computer and System Sciences, 81(7):1144– 1158, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000094.

Ye:2015:MRP

[3096] Jhih-Hong Ye and Biing-Feng Wang.

On the minmax regret path median problem on trees. Journal of Computer and System Sciences, 81(7): 1159–1170, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000100.

Fearnley:2015:SSS

[3097] John Fearnley, Doron Peled, and Sven Schewe. Synthesis of succinct systems. Journal of Computer and System Sciences, 81(7): 1171–1193, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000203.

Albers:2015:MPS

[3098] Susanne Albers, Antonios Antoniadis, and Gero Greiner. On multi-processor speed scaling with migration. Journal of Computer and System Sciences, 81(7):1194– 1209, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000239.

Fernandes:2015:GSM

[3099] Cristina G. Fernandes and Maya Stein. Geodesic stability for memoryless binary long-lived consensus. Journal of Computer and System Sciences, 81(7): 1210–1220, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000240.

Refsdal:2015:SRS

[3100] Atle Refsdal, Ragnhild Kobro Runde, and Ketil Stølen. Stepwise refinement of sequence diagrams with soft realtime constraints. Journal of Computer and System Sciences, 81(7): 1221–1251, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000252.

Salvati:2015:MMW

[3101] Sylvain Salvati. MIX is a 2-MCFL and the word problem in Z^{2} is captured by the IO and the OI hierarchies. Journal of Computer and System Sciences, 81(7): 1252–1277, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000264.

Libkin:2015:RED

[3102] Leonid Libkin, Tony Tan, and Domagoj Vrgoc. Regular expressions for data words. Journal of Computer and System Sciences, 81(7):

1278–1297, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000276.

Manea:2015:CPP

[3103] Florin Manea, Mike Müller, and Dirk Nowotka. Cubic patterns with permutations. Journal of Computer and System Sciences, 81(7):1298– 1310, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000288.

Backstrom:2015:CPC

[3104] Christer Bäckström, Peter Jonsson, Sebastian Ordyniak, and Stefan Szeider. A complete parameterized complexity analysis of bounded planning. Journal of Computer and System Sciences, 81(7):1311– 1332, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001500029X.

Crescenzi:2015:SCF

[3105] Pierluigi Crescenzi, Daniel Gildea, Andrea Marino, Gianluca Rossi, and Giorgio Satta. Synchronous contextfree grammars and optimal linear parsing strategies. Journal of Computer and System Sciences, 81(7): 1333–1356, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000409.

Vychodil:2015:MFD

[3106] Vilem Vychodil. Monoidal functional dependencies. Journal of Computer and System Sciences, 81(7): 1357–1372, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000410.

Ivanyos:2015:GWS

[3107] Gábor Ivanyos, Marek Karpinski, Youming Qiao, and Miklos Santha. Generalized Wong sequences and their applications to Edmonds’ problems. Journal of Computer and System Sciences, 81(7):1373– 1386, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000446.

Barmpalias:2015:IVB

[3108] George Barmpalias, Rod G. Downey, and Michael McInerney. Integer valued betting strategies and Turing degrees. Journal of Computer and System Sciences, 81(7):1387– 1412, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000458.

Lin:2015:CAA

[3109] Tianrong Lin. Corrigendum to “Another approach to the equivalence of measure-many one-way quantum infinite automata and its application” [J. Comput. Syst. Sci. 78 (3) (2012) 807–821]. Journal of Computer and System Sciences, 81(7):

1413, November 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000112. See

[2707].

Anonymous:2015:Cd

[3110] Anonymous. Contents. Journal of Computer and System Sciences, 81 (7):obc, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000537.

Anonymous:2015:EBg

[3111] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(7):ifc, November 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000495.

Enokido:2015:JCS

[3112] Tomoya Enokido, Leonard Barolli, and Makoto Takizawa. Journal of Computer and System Sciences special issue on reliability and optimization for wireless networking and cloud computing. Journal of Computer and System Sciences, 81(8): 1415–1416, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001810.

Xhafa:2015:SMR

[3113] Fatos Xhafa, Christian Sánchez, Admir Barolli, and Makoto Takizawa. Solving mesh router nodes placement problem in wireless mesh networks by tabu search algorithm. Journal of Computer and System Sciences, 81(8): 1417–1428, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001822.

Brocardo:2015:AVM

[3114] Marcelo Luiz Brocardo, Issa Traore, and Isaac Woungang. Authorship verification of e-mail and tweet messages applied for continuous authentication. Journal of Computer and System Sciences, 81(8):1429– 1440, December 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001834.

Nakamura:2015:RAR

[3115] Shigenari Nakamura, Dilawaer Duolikun, and Makoto Takizawa. Readabortion (RA) based synchronization protocols to prevent illegal information flow. Journal of Computer and System Sciences, 81(8): 1441–1451, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001846.

Hernandez-Ramos:2015:SSA

[3116] José L. Hernández-Ramos, M. Victoria Moreno, Jorge Bernal Bernabé, Dan García Carrillo, and Antonio F. Skarmeta. SAFIR: Secure access framework for IoT-enabled services on smart buildings. Journal of Computer and System Sciences, 81(8):

1452–1463, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001858.

Castiglione:2015:MEE

[3117] Arcangelo Castiglione, Francesco Palmieri, Ugo Fiore, Aniello Castiglione, and Alfredo De Santis. Modeling energy-efficient secure communications in multi-mode wireless mobile devices. Journal of Computer and System Sciences, 81(8):1464– 1478, December 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001400186X.

Khan:2015:IBR

[3118] Amin M. Khan, Ümit C. Büyüksahin, and Felix Freitag. Incentive-based resource assignment and regulation for collaborative cloud services in community networks. Journal of Computer and System Sciences, 81(8): 1479–1495, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000014001871.

Barolli:2015:NPW

[3119] Admir Barolli, Tetsuya Oda, Makoto Ikeda, Leonard Barolli, Fatos Xhafa, and Vincenzo Loia. Node placement for wireless mesh networks: Analysis of WMN-GA system simulation results for different parameters and distributions. Journal of Computer and System Sciences, 81(8): 1496–1507, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001883.

Taniar:2015:TRQ

[3120] David Taniar and Wenny Rahayu. A taxonomy for region queries in spatial databases. Journal of Computer and System Sciences, 81(8): 1508–1531, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001895.

Li:2015:SDS

[3121] Jin Li, Xiaofeng Chen, Fatos Xhafa, and Leonard Barolli. Secure deduplication storage systems supporting keyword search. Journal of Computer and System Sciences, 81(8): 1532–1541, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000014001901.

Garibay-Martinez:2015:TPP

[3122] Ricardo Garibay-Martínez, Geoffrey Nelissen, Luis Lino Ferreira, and Luís Miguel Pinho. Task partitioning and priority assignment for distributed hard real-time systems. Journal of Computer and System Sciences, 81(8): 1542–1555, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000665.

Bulteau:2015:PFH

[3123] Laurent Bulteau, Guillaume Fertin, and Irena Rusu. Pancake flipping is hard. Journal of Computer and System Sciences, 81(8): 1556–1574, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000124.

Bienvenu:2015:SFT

[3124] Laurent Bienvenu, Rod Downey, André Nies, and Wolfgang Merkle. Solovay functions and their applications in algorithmic randomness. Journal of Computer and System Sciences, 81(8): 1575–1591, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000422.

Becher:2015:NA

[3125] Verónica Becher, Olivier Carton, and Pablo Ariel Heiber. Normality and automata. Journal of Computer and System Sciences, 81(8): 1592–1613, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001500046X.

Horvath:2015:LPF

[3126] Gábor Horváth and Chrystopher L.

Nehaniv. Length of polynomials over finite groups. Journal of Computer and System Sciences, 81(8): 1614–1622, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000471.

Boyer:2015:LSC

[3127] L. Boyer, M. Delacourt, V. Poupet, M. Sablik, and G. Theyssier. μ-Limit sets of cellular automata from a computational complexity perspective. Journal of Computer and System Sciences, 81(8):1623–1647, December 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000616.

Jeandel:2015:HCE

[3128] Emmanuel Jeandel and Pascal Vanier. Hardness of conjugacy, embedding and factorization of multidimensional subshifts. Journal of Computer and System Sciences, 81(8):1648– 1664, December 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000604.

Karpinski:2015:NIB

[3129] Marek Karpinski, Michael Lampis, and Richard Schmied. New inapproximability bounds for TSP. Journal of Computer and System Sciences, 81(8): 1665–1677, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000641.

Belohlavek:2015:BAB

[3130] Radim Belohlavek and Martin Trnecka. From-below approximations in Boolean matrix factorization: Geometry and new algorithm. Journal of Computer and System Sciences, 81(8): 1678–1697, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001500063X.

Fortune:2015:EGL

[3131] Steven Fortune. Equivalence and generalization in a layered network model. Journal of Computer and System Sciences, 81(8):1698– 1714, December 2015. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000653.

Dereniowski:2015:CML

[3132] Dariusz Dereniowski, Wieslaw Kubiak, and Yori Zwols. The complexity of minimum-length path decompositions. Journal of Computer and System Sciences, 81(8): 1715–1747, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000732.

Anonymous:2015:Ce

[3133] Anonymous. Contents. Journal of Computer and System Sciences, 81 (8):obc, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000860.

Anonymous:2015:EBh

[3134] Anonymous. Editorial Board. Journal of Computer and System Sciences, 81(8):ifc, December 2015. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000823.

Mandreoli:2016:JCS

[3135] Federica Mandreoli, Riccardo Martoglia, and Wilma Penzo. Journal of computer and system sciences special issue on query answering on graph-structured data. Journal of Computer and System Sciences, 82 (1A):1–2, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000938.

Vasilyeva:2016:AWE

[3136] Elena Vasilyeva, Maik Thiele, Christof Bornhövd, and Wolfgang Lehner. Answering “why empty?” and “why so many?” queries in graph databases. Journal of Computer and System Sciences, 82(1A):3– 22, February 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000690.

Wei-Kleiner:2016:TDB

[3137] Fang Wei-Kleiner. Tree decompositionbased indexing for efficient shortest path and nearest neighbors query answering on graphs. Journal of Computer and System Sciences, 82 (1A):23–44, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000707.

Holzschuher:2016:QGD

[3138] Florian Holzschuher and René Peinl.

Querying a graph database — language selection and performance considerations. Journal of Computer and System Sciences, 82(1A):45– 68, February 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000689.

Grandi:2016:DCH

[3139] Fabio Grandi. Dynamic class hierarchy management for multi-version ontology-based personalization. Journal of Computer and System Sciences, 82(1A):69–90, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000628.

Anonymous:2016:Ca

[3140] Anonymous. Contents. Journal of Computer and System Sciences, 82 (1A):obc, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000999.

Anonymous:2016:EBa

[3141] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(1A):ifc, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000951.

Wang:2016:SIM

[3142] Yufeng Wang, Qun Jin, and Athanasios V. Vasilakos. Special issue on mobile social networking and computing in proximity (MSNP). Journal of Computer and System Sciences, 82 (1B):91–92, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001038.

Xia:2016:UPB

[3143] Feng Xia, Hannan Bin Liaqat, Ahmedin Mohammed Ahmed, Li Liu, Jianhua Ma, Runhe Huang, and Amr Tolba. User popularity-based packet scheduling for congestion control in ad-hoc social networks. Journal of Computer and System Sciences, 82 (1B):93–112, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000756.

Fu:2016:DCI

[3144] Songling Fu, Ligang He, Xiangke Liao, and Chenlin Huang. Developing the cloud-integrated data replication framework in decentralized online social networks. Journal of Computer and System Sciences, 82(1B): 113–129, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000720.

Wang:2016:CRT

[3145] Shangguang Wang, Lin Huang, Ching-Hsien Hsu, and Fangchun Yang. Collaboration reputation for trustworthy Web service selection in social networks. Journal of Computer and System Sciences, 82(1B): 130–143, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000719.

Jiang:2016:ERT

[3146] Zhibin Jiang, Ching-Hsien Hsu, Daqiang Zhang, and Xiaolei Zou. Evaluating rail transit timetable using big passengers’ data. Journal of Computer and System Sciences, 82(1B): 144–155, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000926. See

corrigendum [3199].

Du:2016:ELM

[3147] Shaoyong Du, Jingyu Hua, Yue Gao, and Sheng Zhong. EV-Linker: Mapping eavesdropped Wi-Fi packets to individuals via electronic and visual signal matching. Journal of Computer and System Sciences, 82(1B): 156–172, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000677.

Anonymous:2016:Cb

[3148] Anonymous. Contents. Journal of Computer and System Sciences, 82 (1B):obc, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001099.

Anonymous:2016:EBb

[3149] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(1B):ifc, February 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001051.

Palesi:2016:SIE

[3150] Maurizio Palesi, Mario Collotta, Masoud Daneshtalab, and Pradip Bose. Special issue on energy efficient methods and systems in the emerging cloud era. Journal of Computer and System Sciences, 82(2): 173, March 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001178.

Li:2016:PPM

[3151] Keqin Li. Power and performance management for parallel computations in clouds and data centers. Journal of Computer and System Sciences, 82 (2):174–190, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000744.

Kumar:2016:HTA

[3152] Mohan Raj Velayudhan Kumar and Shriram Raghunathan. Heterogeneity and thermal aware adaptive heuristics for energy efficient consolidation of virtual machines in infrastructure clouds. Journal of Computer and System Sciences, 82(2): 191–212, March 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001500080X.

Dabrowski:2016:EEG

[3153] Konrad K. Dabrowski, Petr A.

Golovach, Pim van ’t Hof, and Daniël Paulusma. Editing to Eulerian graphs. Journal of Computer and System Sciences, 82(2): 213–228, March 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001117.

Fletcher:2016:SCN

[3154] George H. L. Fletcher, Marc Gyssens, Jan Paredaens, Dirk Van Gucht, and Yuqing Wu. Structural characterizations of the navigational expressiveness of relation algebras on a tree. Journal of Computer and System Sciences, 82(2):229–259, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001105.

Liu:2016:CBT

[3155] Zhao Liu, Jianxi Fan, and Xiaohua Jia. Complete binary trees embeddings in Möbius cubes. Journal of Computer and System Sciences, 82 (2):260–281, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001026.

Karakoyunlu:2016:EUM

[3156] Cengiz Karakoyunlu and John A. Chandy. Exploiting user metadata for energy-aware node allocation in a cloud storage system. Journal of Computer and System Sciences, 82 (2):282–309, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001014.

Dolev:2016:SCC

[3157] Danny Dolev, Keijo Heljanko, Matti Järvisalo, Janne H. Korhonen, Christoph Lenzen, Joel Rybicki, Jukka Suomela, and Siert Wieringa. Synchronous counting and computational algorithm design. Journal of Computer and System Sciences, 82(2): 310–332, March 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001002.

Ganian:2016:MKS

[3158] Robert Ganian, Friedrich Slivovsky, and Stefan Szeider. Meta-kernelization with structural parameters. Journal of Computer and System Sciences, 82 (2):333–346, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015000914.

Bulatov:2016:CCS

[3159] Andrei A. Bulatov. Conservative constraint satisfaction re-revisited. Journal of Computer and System Sciences, 82(2):347–356, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000793.

Belohlavek:2016:BCS

[3160] Radim Belohlavek and Jan Konecny.

Bases of closure systems over residuated lattices. Journal of Computer and System Sciences, 82(2): 357–365, March 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000781.

Paul:2016:LKR

[3161] Christophe Paul, Anthony Perez, and Stéphan Thomassé. Linear kernel for Rooted Triplet Inconsistency and other problems based on conflict packing technique. Journal of Computer and System Sciences, 82 (2):366–379, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001500077X.

Jerabek:2016:IFM

[3162] Emil Jerábek. Integer factoring and modular square roots. Journal of Computer and System Sciences, 82 (2):380–394, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015000768.

Anonymous:2016:Cc

[3163] Anonymous. Contents. Journal of Computer and System Sciences, 82 (2):obc, March 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001233.

Anonymous:2016:EBc

[3164] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(2):ifc, March 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001191.

Chirkova:2016:CSE

[3165] Rada Chirkova. Combined-semantics equivalence of conjunctive queries: Decidability and tractability results. Journal of Computer and System Sciences, 82(3):395–465, May 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000015001129.

Xu:2016:SAP

[3166] Lingling Xu, Fangguo Zhang, Willy Susilo, and Yamin Wen. Solutions to the anti-piracy problem in oblivious transfer. Journal of Computer and System Sciences, 82(3): 466–476, May 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001142.

Badkobeh:2016:CME

[3167] Golnaz Badkobeh and Maxime Crochemore.

Computing maximal-exponent factors in an overlap-free word. Journal of Computer and System Sciences, 82 (3):477–487, May 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001282.

Shachnai:2016:RFU

[3168] Hadas Shachnai and Meirav Zehavi.

Representative families: a unified tradeoff-based approach. Journal of Computer and System Sciences, 82 (3):488–502, May 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001294.

Abu-Khzam:2016:DRC

[3169] Faisal N. Abu-Khzam, Cristina Bazgan, Morgan Chopin, and Henning Fernau. Data reductions and combinatorial bounds for improved approximation algorithms. Journal of Computer and System Sciences, 82 (3):503–520, May 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001336.

Froese:2016:EHS

[3170] Vincent Froese, René van Bevern, Rolf Niedermeier, and Manuel Sorge. Exploiting hidden structure in selecting dimensions that distinguish vectors. Journal of Computer and System Sciences, 82(3):

521–535, May 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001361.

Edwards:2016:FPS

[3171] Keith J. Edwards. A faster polynomialspace algorithm for Max 2-CSP. Journal of Computer and System Sciences, 82(3):536–550, May 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001385.

Codish:2016:SNI

[3172] Michael Codish, Luís Cruz-Filipe, Michael Frank, and Peter Schneider-Kamp. Sorting nine inputs requires twenty-five comparisons. Journal of Computer and System Sciences, 82 (3):551–563, May 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001397.

Anonymous:2016:Cd

[3173] Anonymous. Contents. Journal of Computer and System Sciences, 82 (3):obc, May 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001452.

Anonymous:2016:EBd

[3174] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(3):ifc, May 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001415.

Chi:2016:SIT

[3175] Chi-Hung Chi, Can Wang, and Yu Zheng. Special issue on trajectorybased behaviour analytics. Journal of Computer and System Sciences, 82(4):565, June 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000052.

Thomason:2016:ILG

[3176] Alasdair Thomason, Nathan Griffiths, and Victor Sanchez. Identifying locations from geospatial trajectories. Journal of Computer and System Sciences, 82(4):566– 581, June 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001312.

Huang:2016:DSR

[3177] Guangyan Huang, Jing He, Wanlei Zhou, Guang-Li Huang, Limin Guo, Xiangmin Zhou, and Feiyi Tang. Discovery of stop regions for understanding repeat travel behaviors of moving objects. Journal of Computer and System Sciences, 82(4): 582–593, June 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001348.

Chu:2016:SRC

[3178] Victor W. Chu, Raymond K. Wong, Fang Chen, Simon Fong, and Patrick C. K. Hung. Self-regularized causal structure discovery for trajectorybased networks. Journal of Computer and System Sciences, 82(4): 594–609, June 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001300.

La:2016:CFT

[3179] Hyun Jung La. A conceptual framework for trajectory-based medical analytics with IoT contexts. Journal of Computer and System Sciences, 82(4):610–626, June 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001500135X.

Anonymous:2016:Ce

[3180] Anonymous. Contents. Journal of Computer and System Sciences, 82 (4):obc, June 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000118.

Anonymous:2016:EBe

[3181] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(4):ifc, June 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000076.

Albers:2016:LUL

[3182] Susanne Albers and Sonja Lauer.

On list update with locality of reference. Journal of Computer and System Sciences, 82(5):627– 653, August 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001166.

Baldoni:2016:ISO

[3183] Roberto Baldoni, Silvia Bonomi, and Michel Raynal. Implementing set objects in dynamic distributed systems. Journal of Computer and System Sciences, 82(5): 654–689, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001130.

Cai:2016:BHS

[3184] Jin-Yi Cai, Andreas Galanis, Leslie Ann Goldberg, Heng Guo, Mark Jerrum, Daniel Štefankovič, and Eric Vigoda. #BIS-hardness for 2-spin systems on bipartite bounded degree graphs in the tree non-uniqueness region. Journal of Computer and System Sciences, 82 (5):690–711, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000015001324.

Chawla:2016:CBA

[3185] Priyanka Chawla, Inderveer Chana, and Ajay Rana. Cloud-based automatic test data generation framework. Journal of Computer and System Sciences, 82(5):712–738, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000016000027.

Teppan:2016:TFP

[3186] Erich Christian Teppan, Gerhard Friedrich, and Georg Gottlob. Tractability frontiers of the partner units configuration problem. Journal of Computer and System Sciences, 82 (5):739–755, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000167.

Vergnaud:2016:CSP

[3187] Damien Vergnaud. Comment on “A strong provably secure IBE scheme without bilinear map” by M. Zheng, Y. Xiang and H. Zhou [J. Comput. Syst. Sci. 81 (2015) 125–131]. Journal of Computer and System Sciences, 82(5):756–757, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000040. See

[3023].

Kratsch:2016:PKS

[3188] Stefan Kratsch. On polynomial kernels for sparse integer linear programs. Journal of Computer and System Sciences, 82(5):758– 766, August 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000039.

Cheng:2016:EFT

[3189] Chia-Wen Cheng and Sun-Yuan Hsieh.

Edge-fault-tolerant pancyclicity and bipancyclicity of Cartesian product graphs with faulty edges. Journal of Computer and System Sciences, 82 (5):767–781, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000179.

Wang:2016:EAR

[3190] Biing-Feng Wang, Jhih-Hong Ye, and Pei-Jung Chen. Efficient algorithms for the round-trip 1-center and 1median problems. Journal of Computer and System Sciences, 82(5): 782–792, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000155.

Find:2016:SSX

[3191] Magnus Find, Mika Göös, Matti Järvisalo, Petteri Kaski, Mikko Koivisto, and Janne H. Korhonen. Separating OR, SUM, and XOR circuits. Journal of Computer and System Sciences, 82(5): 793–801, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000143.

Dereniowski:2016:BCT

[3192] Dariusz Dereniowski, Adrian Kosowski, Dominik Pajak, and Przemyslaw Uznański. Bounds on the cover time of parallel rotor walks. Journal of Computer and System Sciences, 82 (5):802–816, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000180.

Papoutsakis:2016:ATS

[3193] Ioannis Papoutsakis. On approximating tree spanners that are breadth first search trees. Journal of Computer and System Sciences, 82(5): 817–825, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000234.

Carapelle:2016:SEC

[3194] Claudia Carapelle, Alexander Kartzow, and Markus Lohrey. Satisfiability of ECTL^{∗} with constraints. Journal of Computer and System Sciences, 82(5):826–855, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001600012X.

Hannula:2016:FGI

[3195] Miika Hannula, Juha Kontinen, and Sebastian Link. On the finite and general implication problems of independence atoms and keys. Journal of Computer and System Sciences, 82 (5):856–877, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000222.

Chatterjee:2016:WDA

[3196] Krishnendu Chatterjee, Martin Chmelík, and Mathieu Tracol. What is decidable about partially observable Markov decision processes with ωregular objectives. Journal of Computer and System Sciences, 82(5): 878–911, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000246.

Jonsson:2016:CSS

[3197] Peter Jonsson and Johan Thapper.

Constraint satisfaction and semilinear expansions of addition over the rationals and the reals. Journal of Computer and System Sciences, 82 (5):912–928, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000271.

Dolev:2016:HSH

[3198] Danny Dolev, Matthias Függer, Christoph Lenzen, Martin Perner, and Ulrich Schmid. HEX: Scaling honeycombs is easier than scaling clock trees. Journal of Computer and System Sciences, 82(5): 929–956, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001600026X.

Jiang:2016:CER

[3199] Zhibin Jiang, Ching-Hsien Hsu, Daqiang Zhang, and Xiaolei Zou. Corrigendum to “Evaluating rail transit timetable using big passengers’ data” [J. Comput. Syst. Sci. 82 (1, Part B) (2016) 144–155]. Journal of Computer and System Sciences, 82(5):957, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000015001373. See

[3146].

Anonymous:2016:Cf

[3200] Anonymous.  Contents.  Journal

of Computer and System Sciences, 82(5):obc, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300083.

Anonymous:2016:EBf

[3201] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(5):ifc, August 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300046.

Grandjean:2016:LAL

[3202] Etienne Grandjean and Frédéric Olive. A logical approach to locality in pictures languages. Journal of Computer and System Sciences, 82(6): 959–1006, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000258.

Jez:2016:ESI

[3203] Artur Jez and Alexander Okhotin. Equations over sets of integers with addition only. Journal of Computer and System Sciences, 82(6): 1007–1019, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000131.

Bringmann:2016:PCD

[3204] Karl Bringmann, Danny Hermelin, Matthias Mnich, and Erik Jan van Leeuwen. Parameterized complexity dichotomy for Steiner Multicut. Journal of Computer and System Sciences, 82(6):1020–1043, September 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000283.

Hasan:2016:AAW

[3205] Mohammad Khairul Hasan and Kyung-Yong Chwa. Approximation algorithms for the weighted t-uniform sparsest cut and some other graph partitioning problems. Journal of Computer and System Sciences, 82(6): 1044–1063, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000295.

Lee:2016:HVN

[3206] Troy Lee, Nikos Leonardos, Michael Saks, and Fengming Wang. Hellinger volume and number-on-the-forehead communication complexity. Journal of Computer and System Sciences, 82(6): 1064–1074, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016000301.

vanIersel:2016:KHN

[3207] Leo van Iersel, Steven Kelk, and Celine Scornavacca. Kernelizations for the hybridization number problem on multiple nonbinary trees. Journal of Computer and System Sciences, 82(6): 1075–1089, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300010.

Song:2016:ETF

[3208] Bosheng Song, Mario J. Pérez-Jiménez, and Linqiang Pan. An efficient time-free solution to SAT problem by P systems with proteins on membranes. Journal of Computer and System Sciences, 82(6): 1090–1099, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300113.

Froese:2016:WWK

[3209] Vincent Froese, André Nichterlein, and Rolf Niedermeier. Win-win kernelization for degree sequence completion problems. Journal of Computer and System Sciences, 82(6): 1100–1111, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300125.

Pal:2016:FPG

[3210] Dipali Pal, Praveen Rao, Vasil Slavov, and Anas Katib. Fast processing of graph queries on a large database of small and medium-sized data graphs. Journal of Computer and System Sciences, 82(6):1112– 1143, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300149.

Goldberg:2016:ACL

[3211] Leslie Ann Goldberg, Rob Gysel, and John Lapinskas. Approximately counting locally-optimal structures. Journal of Computer and System Sciences, 82(6):1144– 1160, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300137.

Anonymous:2016:Cg

[3212] Anonymous. Contents. Journal of Computer and System Sciences, 82 (6):obc, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300265.

Anonymous:2016:EBg

[3213] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(6):ifc, September 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300228.

Ferrucci:2016:MRQ

[3214] L. Ferrucci, L. Ricci, M. Albano, R. Baraglia, and M. Mordacchini. Multidimensional range queries on hierarchical Voronoi overlays. Journal of Computer and System Sciences, 82(7): 1161–1179, November 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300204.

Reingold:2016:NTT

[3215] Omer Reingold and Shai Vardi. New techniques and tighter bounds for local computation algorithms. Journal of Computer and System Sciences, 82(7): 1180–1200, November 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300332.

Fulop:2016:LTT

[3216] Zoltán Fülöp and Andreas Maletti. Linking theorems for tree transducers. Journal of Computer and System Sciences, 82(7):1201–1222, November 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300290.

Anonymous:2016:Ch

[3217] Anonymous. Contents. Journal of Computer and System Sciences, 82 (7):obc, November 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300393.

Anonymous:2016:EBh

[3218] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(7):ifc, November 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300356.

Anthony:2016:MCC

[3219] Martin Anthony and Joel Ratsaby.

Multi-category classifiers and sample width. Journal of Computer and System Sciences, 82(8):1223– 1231, December 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300150.

Cabessa:2016:EPF

[3220] Jérémie Cabessa and Alessandro E. P.

Villa. Expressive power of first-order recurrent neural networks determined by their attractor dynamics. Journal of Computer and System Sciences, 82(8): 1232–1250, December 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300186.

Czygrinow:2016:DCS

[3221] Andrzej Czygrinow, Michal Hanćkowiak, Edyta Szymańska, and Wojciech Wawrzyniak. On the distributed complexity of the semi-matching problem. Journal of Computer and System Sciences, 82(8):1251–1267, December 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300277.

Miasnikov:2016:GCC

[3222] Alexei Miasnikov and Alexander Ushakov. Generic case completeness. Journal of Computer and System Sciences, 82(8):1268–1282, December 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300289.

Barmpalias:2016:OAB

[3223] George Barmpalias, Nan Fang, and Andrew Lewis-Pye. Optimal asymptotic bounds on the oracle use in computations from Chaitin’s Omega. Journal of Computer and System Sciences, 82(8): 1283–1299, December 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300307.

Hu:2016:EJD

[3224] Xiaocheng Hu, Miao Qiao, and Yufei Tao. I/O-efficient join dependency testing, Loomis–Whitney join, and triangle enumeration. Journal of Computer and System Sciences, 82(8): 1300–1315, December 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300319.

Liu:2016:SSE

[3225] Xuejiao Liu, Yingjie Xia, Wenzhi Chen, Yang Xiang, Mohammad Mehedi Hassan, and Abdulhameed Alelaiwi. SEMD: Secure and efficient message dissemination with policy enforcement in VANET. Journal of Computer and System Sciences, 82(8):1316– 1328, December 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300320.

Aizikowitz:2016:LCG

[3226] Tamar Aizikowitz and Michael Kaminski. LR (0) conjunctive grammars and deterministic synchronized alternating pushdown automata. Journal of Computer and System Sciences, 82(8): 1329–1359, December 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001630040X.

Rosenke:2016:ECP

[3227] Christian Rosenke. The exact complexity of projective image matching. Journal of Computer and System Sciences, 82(8):1360–1387, December 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300423.

Anonymous:2016:Ci

[3228] Anonymous. Contents. Journal of Computer and System Sciences, 82 (8):obc, December 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300575.

Anonymous:2016:EBi

[3229] Anonymous. Editorial Board. Journal of Computer and System Sciences, 82(8):ifc, December 2016. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300538.

Choo:2017:SIC

[3230] Kim-Kwang Raymond Choo, Jemal Abawajy, and Rafiqul Islam. Special issue on cyber security in the critical infrastructure: Advances and future directions. Journal of Computer and System Sciences, 83(1): 1–2, February 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300708.

Gonzalez-Granadillo:2017:PBA

[3231] G. Gonzalez-Granadillo, J. Garcia-Alfaro, and H. Debar. A polytopebased approach to measure the impact of events against critical infrastructures. Journal of Computer and System Sciences, 83(1):3– 21, February 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000192.

Puthal:2017:DPN

[3232] Deepak Puthal, Surya Nepal, Rajiv Ranjan, and Jinjun Chen. A dynamic prime number based efficient security mechanism for big sensing data streams. Journal of Computer and System Sciences, 83(1):22– 42, February 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016000209.

Grill:2017:RFP

[3233] Martin Grill, Tomás Pevný, and Martin Rehak. Reducing false positives of network anomaly detection by local adaptive multivariate smoothing. Journal of Computer and System Sciences, 83(1):43–57, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300022.

Anwar:2017:MPE

[3234] Adnan Anwar, Abdun Naser Mahmood, and Mark Pickering. Modeling and performance evaluation of stealthy false data injection attacks on smart grid in the presence of corrupted measurements. Journal of Computer and System Sciences, 83 (1):58–72, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300174.

Lin:2017:NUD

[3235] Chao Lin, Wei Wu, Xinyi Huang, and Li Xu. A new universal designated verifier transitive signature scheme for big graph data. Journal of Computer and System Sciences, 83 (1):73–83, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300435.

Ford:2017:SEP

[3236] Vitaly Ford, Ambareen Siraj, and Mohammad Ashiqur Rahman. Secure and efficient protection of consumer privacy in advanced metering infrastructure supporting finegrained data analysis. Journal of Computer and System Sciences, 83 (1):84–100, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300472.

Haar:2017:OCA

[3237] Stefan Haar, Serge Haddad, Tarek Melliti, and Stefan Schwoon. Optimal constructions for active diagnosis. Journal of Computer and System Sciences, 83 (1):101–120, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300198.

Gutin:2017:RPP

[3238] Gregory Gutin, Magnus Wahlström, and Anders Yeo. Rural postman parameterized by the number of components of required edges. Journal of Computer and System Sciences, 83 (1):121–131, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300411.

Diaz:2017:CMD

[3239] Josep Diaz, Olli Pottonen, Maria Serna, and Erik Jan van Leeuwen. Complexity of metric dimension on planar graphs. Journal of Computer and System Sciences, 83(1):

132–158, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300484.

Son:2017:SCM

[3240] Le Hoang Son and Pham Huy Thong.

Soft computing methods for WiMax network planning on 3D geographical information systems. Journal of Computer and System Sciences, 83 (1):159–179, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300514.

Gogacz:2017:CCT

[3241] Tomasz Gogacz and Jerzy Marcinkowski.

Converging to the chase — a tool for finite controllability. Journal of Computer and System Sciences, 83 (1):180–206, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300617.

Onn:2017:HTM

[3242] Shmuel Onn. Huge tables and multicommodity flows are fixed-parameter tractable via unimodular integer Carathéodory. Journal of Computer and System Sciences, 83(1): 207–214, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300629.

Anonymous:2017:Ca

[3243] Anonymous. Contents. Journal of Computer and System Sciences, 83 (1):obc, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300770.

Anonymous:2017:EBa

[3244] Anonymous. Editorial Board. Journal of Computer and System Sciences, 83(1):ifc, February 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300733.

Etscheid:2017:PKW

[3245] Michael Etscheid, Stefan Kratsch, Matthias Mnich, and Heiko Röglin. Polynomial kernels for weighted problems. Journal of Computer and System Sciences, 84(??):1–10, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000016300460.

Day:2017:CPP

[3246] Joel D. Day, Daniel Reidenbach, and Markus L. Schmid. Closure properties of pattern languages. Journal of Computer and System Sciences, 84(??):11–31, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300605.

Civril:2017:SAP

[3247] A. Çivril. Sparse approximation is provably hard under coherent dictionaries. Journal of Computer and System Sciences, 84(??): 32–43, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300587.

Chang:2017:LBM

[3248] Ching-Lueh Chang. A lower bound for metric 1-median selection. Journal of Computer and System Sciences, 84(??):44–51, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300691.

Jonsson:2017:SPC

[3249] Peter Jonsson, Victor Lagerkvist, Gustav Nordh, and Bruno Zanuttini. Strong partial clones and the time complexity of SAT problems. Journal of Computer and System Sciences, 84(??):52–78, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001630068X.

Epstein:2017:AVS

[3250] Leah Epstein and Asaf Levin. An AF-PTAS for variable sized bin packing with general activation costs. Journal of Computer and System Sciences, 84(??):79–96, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300678.

Xiao:2017:GNT

[3251] Mingyu Xiao. On a generalization of Nemhauser and Trotter’s local optimization theorem. Journal of Computer and System Sciences, 84 (??):97–106, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300666.

Gutin:2017:PCA

[3252] Gregory Gutin, Mark Jones, and Bin Sheng. Parameterized complexity of the k-arc Chinese Postman Problem. Journal of Computer and System Sciences, 84(??): 107–119, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300654.

Panolan:2017:PCC

[3253] Fahad Panolan, Geevarghese Philip, and Saket Saurabh. On the parameterized complexity of b-chromatic number. Journal of Computer and System Sciences, 84(??):120– 131, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300927.

Zakablukov:2017:AGC

[3254] Dmitry V. Zakablukov. On asymptotic gate complexity and depth of reversible circuits without additional memory. Journal of Computer and System Sciences, 84(??): 132–143, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300903.

Brazdil:2017:TPS

[3255] Tomás Brázdil, Krishnendu Chatterjee, Vojtech Forejt, and Antonín Kucera. Trading performance for stability in Markov decision processes. Journal of Computer and System Sciences, 84 (??):144–170, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300897.

Feigenblat:2017:MWI

[3256] Guy Feigenblat, Ely Porat, and Ariel Shiftan. d-k-min-wise independent family of hash functions. Journal of Computer and System Sciences, 84 (??):171–184, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300848.

Bundala:2017:ODS

[3257] Daniel Bundala, Michael Codish, Luís Cruz-Filipe, Peter Schneider-Kamp, and Jakub Závodný. Optimal-depth sorting networks. Journal of Computer and System Sciences, 84(??): 185–204, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300836.

Kociumaka:2017:FAA

[3258] T. Kociumaka, J. Radoszewski, and W. Rytter. Fast algorithms for Abelian periods in words and greatest common divisor queries. Journal of Computer and System Sciences, 84 (??):205–218, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300824.

Gajarsky:2017:KUS

[3259] Jakub Gajarský, Petr Hliněný, Jan Obdržálek, Sebastian Ordyniak, Felix Reidl, Peter Rossmanith, Fernando Sánchez Villaamil, and Somnath Sikdar. Kernelization using structural parameters on sparse graph classes. Journal of Computer and System Sciences, 84(??): 219–242, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300812.

Schewe:2017:SPG

[3260] Sven Schewe. Solving parity games in big steps. Journal of Computer and System Sciences, 84(??): 243–262, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300952.

Kari:2017:DWR

[3261] Lila Kari and Steffen Kopecki. Deciding whether a regular language is generated by a splicing system. Journal of Computer and System Sciences, 84 (??):263–287, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300940.

Cassaigne:2017:SMA

[3262] Julien Cassaigne, Nicolas Ollinger, and Rodrigo Torres-Avilés. A small minimal aperiodic reversible Turing machine. Journal of Computer and System Sciences, 84(??): 288–301, March 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301003.

Anonymous:2017:Cb

[3263] Anonymous.  Contents.  Journal

of Computer and System Sciences, 84(??):obc, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301118.

Anonymous:2017:EBb

[3264] Anonymous. Editorial Board. Journal of Computer and System Sciences, 84(??):ifc, March 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301076.

Khorandi:2017:SOC

[3265] Sina Mahmoodi Khorandi and Mohsen Sharifi. Scheduling of online computeintensive synchronized jobs on high performance virtual clusters. Journal of Computer and System Sciences, 85(??):1–17, May 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301052.

Jansen:2017:TKF

[3266] Bart M. P. Jansen. Turing kernelization for finding long paths and cycles in restricted graph classes. Journal of Computer and System Sciences, 85(??):18–37, May 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301040.

Gaspers:2017:BHC

[3267] Serge Gaspers, Neeldhara Misra, Sebastian Ordyniak, Stefan Szeider, and Stanislav Zivný. Backdoors into heterogeneous classes of SAT and CSP. Journal of Computer and System Sciences, 85(??):38–56, May 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301039.

Benedikt:2017:DRF

[3268] M. Benedikt, J. Engelfriet, and S. Maneth. Determinacy and rewriting of functional top-down and MSO tree transformations. Journal of Computer and System Sciences, 85(??): 57–73, May 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001630112X.

Saberifar:2017:CFR

[3269] Fatemeh Zahra Saberifar, Ali Mohades, Mohammadreza Razzazi, and Jason M. O’Kane. Combinatorial filter reduction: Special cases, approximation, and fixed-parameter tractability. Journal of Computer and System Sciences, 85(??):74–92, May 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301192.

Bottoni:2017:NPM

[3270] Paolo Bottoni, Anna Labella, and Victor Mitrana. Networks of polarized multiset processors. Journal of Computer and System Sciences, 85(??):93–103, May 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301180.

Koehler:2017:IDT

[3271] Henning Koehler and Sebastian Link.

Inclusion dependencies and their interaction with functional dependencies in SQL. Journal of Computer and System Sciences, 85(??): 104–131, May 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301167.

Manea:2017:EEL

[3272] Florin Manea, Mike Müller, Dirk Nowotka, and Shinnosuke Seki. The extended equation of Lyndon and Schützenberger. Journal of Computer and System Sciences, 85(??): 132–167, May 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301155.

Dabrowski:2017:EPG

[3273] Konrad K. Dabrowski, Petr A. Golovach, Pim van ’t Hof, Daniël Paulusma, and Dimitrios M. Thilikos. Editing to a planar graph of given degrees. Journal of Computer and System Sciences, 85(??):168–182, May 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016301210.

Anonymous:2017:Cc

[3274] Anonymous. Contents. Journal of Computer and System Sciences, 85 (??):obc, May 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016301271.

Anonymous:2017:EBc

[3275] Anonymous. Editorial Board. Journal of Computer and System Sciences, 85(??):ifc, May 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016301234.

Ogiela:2017:OSI

[3276] Marek R. Ogiela, Fatos Xhafa, and Leonard Barolli. Optimization and security issues for next generation networks — editorial. Journal of Computer and System Sciences, 86 (??):1–2, June 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300107.

Sharma:2017:DBD

[3277] Vishal Sharma, Hsing-Chung Chen, and Rajesh Kumar. Driver behaviour detection and vehicle rating using multi-UAV coordinated vehicular networks. Journal of Computer and System Sciences, 86(??): 3–32, June 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001630099X.

Waluyo:2017:TDD

[3278] Agustinus Borgy Waluyo, David Taniar, Wenny Rahayu, and Bala Srinivasan. Trustworthy data delivery in mobile P2P network. Journal of Computer and System Sciences, 86(??):33–48, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300873.

Miao:2017:PVD

[3279] Meixia Miao, Jianfeng Wang, Jianfeng Ma, and Willy Susilo. Publicly verifiable databases with efficient insertion/deletion operations. Journal of Computer and System Sciences, 86(??):49–58, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300642.

Lee:2017:TTA

[3280] Joonghwan Lee, Jae Woo Seo, Hoon Ko, and Hyoungshick Kim. TARD: Temporary Access Rights Delegation for guest network devices. Journal of Computer and System Sciences, 86(??):59–69, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300599.

Czyzowicz:2017:CFN

[3281] Jurek Czyzowicz, Dariusz Dereniowski, Leszek Gasieniec, Ralf Klasing, Adrian Kosowski, and Dominik Pajak. Collision-free network exploration. Journal of Computer and System Sciences, 86(??): 70–81, June 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301209.

Koutsos:2017:PCV

[3282] Adrien Koutsos and Victor Vianu. Process-centric views of data-driven business artifacts. Journal of Computer and System Sciences, 86(??): 82–107, June 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301362.

Turner:2017:TMM

[3283] Jacob Turner. Tensors masquerading as matchgates: Relaxing planarity restrictions on Pfaffian circuits. Journal of Computer and System Sciences, 86(??):108–116, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301295.

Kuijpers:2017:RDC

[3284] Bart Kuijpers and Bart Moelans. On the realisability of double-cross matrices by polylines in the plane. Journal of Computer and System Sciences, 86(??):117–135, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301283.

Ganardi:2017:CST

[3285] Moses Ganardi, Danny Hucke, Artur Jez, Markus Lohrey, and Eric Noeth. Constructing small tree grammars and small circuits for formulas. Journal of Computer and System Sciences, 86 (??):136–158, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301374.

Bockenhauer:2017:ACS

[3286] Hans-Joachim Böckenhauer, Dennis Komm, Rastislav Královic, and Richard Královic. On the advice complexity of the k-server problem. Journal of Computer and System Sciences, 86(??):159–170, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300016.

Bille:2017:FCS

[3287] Philip Bille, Inge Li Gørtz, Patrick Hagge Cording, Benjamin Sach, Hjalte Wedel Vildhøj, and Søren Vind. Fingerprints in compressed strings. Journal of Computer and System Sciences, 86(??):171–180, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300028.

Woryna:2017:GGB

[3288] Adam Woryna. On groups generated by bi-reversible automata: The two-state case over a changing alphabet. Journal of Computer and System Sciences, 86(??): 181–190, June 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300041.

Kim:2017:PAM

[3289] Eun Jung Kim, Christophe Paul, Ignasi Sau, and Dimitrios M. Thilikos. Parameterized algorithms for min-max multiway cut and list digraph homomorphism. Journal of Computer and System Sciences, 86 (??):191–206, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001730003X.

Okhotin:2017:SCO

[3290] Alexander Okhotin and Kai Salomaa. State complexity of operations on input-driven pushdown automata. Journal of Computer and System Sciences, 86(??):207– 228, June 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300077.

Anonymous:2017:Cd

[3291] Anonymous. Contents. Journal of Computer and System Sciences, 86 (??):obc, June 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300168.

Anonymous:2017:EBd

[3292] Anonymous. Editorial Board. Journal of Computer and System Sciences, 86(??):ifc, June 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300120.

Feree:2017:GSA

[3293] Hugo Férée. Game semantics approach to higher-order complexity. Journal of Computer and System Sciences, 87(??):1–15, August 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300090.

deHaan:2017:PCC

[3294] Ronald de Haan and Stefan Szeider. Parameterized complexity classes beyond para-NP. Journal of Computer and System Sciences, 87(??): 16–57, August 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300089.

Murawski:2017:RPR

[3295] A. S. Murawski, S. J. Ramsay, and N. Tzevelekos. Reachability in pushdown register automata. Journal of Computer and System Sciences, 87 (??):58–83, August 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300272.

Holzer:2017:DMC

[3296] Stephan Holzer, Thomas Locher, Yvonne Anne Pignolet, and Roger Wattenhofer. Deterministic multichannel information exchange. Journal of Computer and System Sciences, 87(??):84–103, August 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300259.

Fulla:2017:PVC

[3297] Peter Fulla and Stanislav Zivný. On planar valued CSPs. Journal of Computer and System Sciences, 87 (??):104–118, August 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300338.

Bjorklund:2017:NSP

[3298] Andreas Björklund, Thore Husfeldt, Petteri Kaski, and Mikko Koivisto. Narrow sieves for parameterized paths and packings. Journal of Computer and System Sciences, 87(??): 119–139, August 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300314.

Anonymous:2017:Ce

[3299] Anonymous. Contents. Journal of Computer and System Sciences, 87 (??):obc, August 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300491.

Anonymous:2017:EBe

[3300] Anonymous. Editorial Board. Journal of Computer and System Sciences, 87(??):ifc, August 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300454.

Libkin:2017:WLL

[3301] Leonid Libkin, Ulrich Kohlenbach, and Ruy J. G. B. de Queiroz. 20th workshop on logic, language, information and computation — WoL-LIC 2013. Journal of Computer and System Sciences, 88(??):1–2, September 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300673.

Zeume:2017:DCQ

[3302] Thomas Zeume and Thomas Schwentick.

Dynamic conjunctive queries. Journal of Computer and System Sciences, 88 (??):3–26, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300430.

Hamadou:2017:QLP

[3303] Sardaouna Hamadou, Catuscia Palamidessi, and Vladimiro Sassone. Quantifying leakage in the presence of unreliable sources of information. Journal of Computer and System Sciences, 88 (??):27–52, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300429.

Caicedo:2017:DOB

[3304] Xavier Caicedo, George Metcalfe, Ricardo Rodríguez, and Jonas Rogger. Decidability of order-based modal logics. Journal of Computer and System Sciences, 88(??): 53–74, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300417.

Czerwinski:2017:DDD

[3305] Wojciech Czerwiński, Claire David, Katja Losemann, and Wim Martens. Deciding definability by deterministic regular expressions. Journal of Computer and System Sciences, 88 (??):75–89, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300405.

Engstrom:2017:DLG

[3306] Fredrik Engström, Juha Kontinen, and Jouko Väänänen. Dependence logic with generalized quantifiers: Axiomatizations. Journal of Computer and System Sciences, 88(??):

90–102, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300387.

Ebbing:2017:BDL

[3307] Johannes Ebbing, Lauri Hella, Peter Lohmann, and Jonni Virtema. Boolean dependence logic and partially-ordered connectives. Journal of Computer and System Sciences, 88(??):103– 125, September 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300375.

Alechina:2017:MCR

[3308] Natasha Alechina, Brian Logan, Hoang Nga Nguyen, and Franco Raimondi. Model-checking for resourcebounded ATL with production and consumption of resources. Journal of Computer and System Sciences, 88(??): 126–144, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300363.

Aracena:2017:FPC

[3309] Julio Aracena, Adrien Richard, and Lilian Salinas. Fixed points in conjunctive networks and maximal independent sets in graph contractions. Journal of Computer and System Sciences, 88(??):145– 163, September 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300533.

Barcelo:2017:DCC

[3310] Pablo Barceló and Gaëlle Fontaine. On the data complexity of consistent query answering over graph databases. Journal of Computer and System Sciences, 88(??):164– 194, September 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300521. See

corrections [3933].

Chitnis:2017:FEA

[3311] Rajesh Chitnis, Fedor V. Fomin, Daniel Lokshtanov, Pranabendu Misra, M. S. Ramanujan, and Saket Saurabh. Faster exact algorithms for some terminal set problems. Journal of Computer and System Sciences, 88(??): 195–207, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300569.

Zhou:2017:ECH

[3312] Jin-Xin Zhou. On g-extra connectivity of hypercube-like networks. Journal of Computer and System Sciences, 88(??): 208–219, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300557.

Kogan:2017:IPO

[3313] Kirill Kogan, Alejandro López-Ortiz, Sergey I. Nikolenko, and Alexander V. Sirotkin. The impact of processing order on performance: A taxonomy of semi-FIFO policies. Journal of Computer and System Sciences, 88(??):

220–235, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300545.

Chatterjee:2017:HST

[3314] Krishnendu Chatterjee and Yaron Velner. Hyperplane separation technique for multidimensional meanpayoff games. Journal of Computer and System Sciences, 88(??): 236–259, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300582.

Xiao:2017:LKS

[3315] Mingyu Xiao. Linear kernels for separating a graph into components of bounded size. Journal of Computer and System Sciences, 88(??): 260–270, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300570.

Siddique:2017:FVS

[3316] Umair Siddique and Sofiène Tahar.

Formal verification of stability and chaos in periodic optical systems. Journal of Computer and System Sciences, 88(??):271–289, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000017300624.

Anonymous:2017:Cf

[3317] Anonymous. Contents. Journal of Computer and System Sciences, 88 (??):obc, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300739.

Anonymous:2017:EBf

[3318] Anonymous. Editorial Board. Journal of Computer and System Sciences, 88(??):ifc, September 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300697.

Anonymous:2017:JCS

[3319] Anonymous. Journal of Computer and System Sciences special issue on theoretical and algorithmic foundation for big data. Journal of Computer and System Sciences, 89 (??):1–3, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300946.

Sun:2017:BFT

[3320] Dawei Sun, Guangyan Zhang, Chengwen Wu, Keqin Li, and Weimin Zheng. Building a fault tolerant framework with deadline guarantee in big data stream computing environments. Journal of Computer and System Sciences, 89(??):4–23, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016301143.

Khoshkbarforoushha:2017:EMS

[3321] Alireza Khoshkbarforoushha, Alireza Khosravian, and Rajiv Ranjan. Elasticity management of streaming data analytics flows on clouds. Journal of Computer and System Sciences, 89 (??):24–40, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301131.

Wang:2017:AIA

[3322] Zhiwei Wang, Cheng Cao, Nianhua Yang, and Victor Chang. ABE with improved auxiliary input for big data security. Journal of Computer and System Sciences, 89(??): 41–50, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301349.

Huang:2017:SFC

[3323] Cheng Huang, Rongxing Lu, and Kim-Kwang Raymond Choo. Secure and flexible cloud-assisted association rule mining over horizontally partitioned databases. Journal of Computer and System Sciences, 89 (??):51–63, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301337.

Liang:2017:DIH

[3324] Wei Liang, Chunhua Hu, Min Wu, and Qun Jin. A data intensive heuristic approach to the two-stage streaming scheduling problem. Journal of Computer and System Sciences, 89 (??):64–79, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300053.

Li:2017:CCD

[3325] Tong Li, Zheli Liu, Jin Li, Chunfu Jia, and Kuan-Ching Li. CDPS: a cryptographic data publishing system. Journal of Computer and System Sciences, 89(??):80–91, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016301313.

Xiao:2017:EMB

[3326] Guoqing Xiao, Kenli Li, Xu Zhou, and Keqin Li. Efficient monochromatic and bichromatic probabilistic reverse top- k query processing for uncertain big data. Journal of Computer and System Sciences, 89(??): 92–113, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300459.

Ahmed:2017:CST

[3327] Aly Ahmed and Alex Thomo. Computing source-to-target shortest paths for complex networks in RDBMS. Journal of Computer and System Sciences, 89 (??):114–129, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300915.

Song:2017:GSB

[3328] Weijing Song, Lizhe Wang, Yang Xiang, and Albert Y. Zomaya. Geographic spatiotemporal big data correlation analysis via the Hilbert-Huang transformation. Journal of Computer and System Sciences, 89(??): 130–141, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300806.

Sun:2017:PKE

[3329] Shi-Feng Sun, Dawu Gu, Udaya Parampalli, Yu Yu, and Baodong Qin. Public key encryption resilient to leakage and tampering attacks. Journal of Computer and System Sciences, 89 (??):142–156, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300326.

Shachnai:2017:MFW

[3330] Hadas Shachnai and Meirav Zehavi. A multivariate framework for weighted FPT algorithms. Journal of Computer and System Sciences, 89(??): 157–189, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300612.

Gong:2017:ZDC

[3331] Xueqing Gong and Chi Wan Sung.

Zigzag decodable codes: Linear-time erasure codes with applications to data storage. Journal of Computer and System Sciences, 89(??): 190–208, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300636.

Abriola:2017:ADX

[3332] Sergio Abriola, María Emilia Descotte, Raul Fervari, and Santiago Figueira. Axiomatizations for downward XPath on data trees. Journal of Computer and System Sciences, 89(??): 209–245, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300661.

Park:2017:DPC

[3333] Jung-Heum Park, Hee-Chul Kim, and Hyeong-Seok Lim. Disjoint path covers with path length constraints in restricted hypercube-like graphs. Journal of Computer and System Sciences, 89 (??):246–269, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001730065X.

Brzozowski:2017:CSF

[3334] Janusz A. Brzozowski and Marek Szykula. Complexity of suffix-free regular languages. Journal of Computer and System Sciences, 89(??): 270–287, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300818.

Salo:2017:DUQ

[3335] Ville Salo. Decidability and universality of quasiminimal subshifts. Journal of Computer and System Sciences, 89 (??):288–314, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300892.

Cohen:2017:PTA

[3336] Nathann Cohen, Daniel Gonçalves, Eun Jung Kim, Christophe Paul, Ignasi Sau, Dimitrios M. Thilikos, and Mathias Weller. A polynomialtime algorithm for outerplanar diameter improvement. Journal of Computer and System Sciences, 89(??): 315–327, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300880.

Stewart:2017:CDD

[3337] Iain A. Stewart. On the combinatorial design of data centre network topologies. Journal of Computer and System Sciences, 89(??): 328–348, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300879.

Barmpalias:2017:DHP

[3338] George Barmpalias and Andrew Lewis-Pye. Differences of halting probabilities. Journal of Computer and System Sciences, 89(??):349– 360, November 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300867.

Feng:2017:PDC

[3339] Yuan Feng and Lijun Zhang. Precisely deciding CSL formulas through approximate model checking for CTMCs. Journal of Computer and System Sciences, 89(??):361– 371, November 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300855.

Groz:2017:ETM

[3340] B. Groz and S. Maneth. Efficient testing and matching of deterministic regular expressions. Journal of Computer and System Sciences, 89(??): 372–399, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300843.

Huang:2017:SSS

[3341] Kaibin Huang, Raylin Tso, and Yu-Chi Chen. Somewhat semantic secure public key encryption with filtered-equality-test in the standard model and its extension to searchable encryption. Journal of Computer and System Sciences, 89(??): 400–409, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300831.

Dabrowski:2017:CDF

[3342] Konrad K. Dabrowski, François Dross, and Daniël Paulusma. Colouring diamond-free graphs. Journal of Computer and System Sciences, 89(??): 410–431, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300922.

Yang:2017:CCA

[3343] Yongjie Yang and Jiong Guo. The control complexity of r-approval: From the single-peaked case to the general case. Journal of Computer and System Sciences, 89(??):432– 449, November 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300910.

Guermeur:2017:NSS

[3344] Yann Guermeur. L_{p}-norm Sauer– Shelah lemma for margin multicategory classifiers. Journal of Computer and System Sciences, 89(??): 450–473, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300909.

Anonymous:2017:EBg

[3345] Anonymous. Editorial Board. Journal of Computer and System Sciences, 89(??):ifc, November 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001730096X.

Ding:2017:SOM

[3346] Yong Ding, Zheng Xu, Jun Ye, and Kim-Kwang Raymond Choo. Secure outsourcing of modular exponentiations under single untrusted programme model. Journal of Computer and System Sciences, 90(??):1– 13, December 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301179.

Yin:2017:QPE

[3347] Hui Yin, Zheng Qin, Lu Ou, and Keqin Li. A query privacy-enhanced and secure search scheme over encrypted data in cloud computing. Journal of Computer and System Sciences, 90 (??):14–27, December 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301301.

Alabdulatif:2017:PPA

[3348] Abdulatif Alabdulatif, Heshan Kumarage, Ibrahim Khalil, and Xun Yi. Privacy-preserving anomaly detection in cloud with lightweight homomorphic encryption. Journal of Computer and System Sciences, 90 (??):28–45, December 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300284.

Au:2017:GFS

[3349] Man Ho Au, Tsz Hon Yuen, Joseph K. Liu, Willy Susilo, Xinyi Huang, Yang Xiang, and Zoe L. Jiang. A general framework for secure sharing of personal health records in cloud system. Journal of Computer and System Sciences, 90(??): 46–62, December 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300296.

Chakraborty:2017:BNO

[3350] Sankardeep Chakraborty, Venkatesh Raman, and Srinivasa Rao Satti. Biconnectivity, st-numbering and other applications of DFS using O(n) bits. Journal of Computer and System Sciences, 90(??):63– 79, December 2017. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300934.

Goldberg:2017:QCA

[3351] Paul W. Goldberg and Stefano Turchetta. Query complexity of approximate equilibria in anonymous games. Journal of Computer and System Sciences, 90(??): 80–98, December 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001730106X.

Bednarova:2017:BLO

[3352] Zuzana Bednárová, Viliam Geffert, Carlo Mereghetti, and Beatrice Palano. Boolean language operations on nondeterministic automata with a pushdown of constant height. Journal of Computer and System Sciences, 90 (??):99–114, December 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301071.

Leporati:2017:CCT

[3353] Alberto Leporati, Luca Manzoni, Giancarlo Mauri, Antonio E. Porreca, and Claudio Zandron. Characterising the complexity of tissue P systems with fission rules. Journal of Computer and System Sciences, 90(??): 115–128, December 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301083.

Anonymous:2017:EBh

[3354] Anonymous. Editorial Board. Journal of Computer and System Sciences, 90(??):ifc, December 2017. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001730123X.

Goles:2018:CTD

[3355] Eric Goles, Pedro Montealegre, Kévin Perrot, and Guillaume Theyssier. On the complexity of two-dimensional signed majority cellular automata. Journal of Computer and System Sciences, 91(??):1–32, February 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000017301198.

Komarath:2018:PMC

[3356] Balagopal Komarath, Jayalal Sarma, and Saurabh Sawlani. Pebbling meets coloring: Reversible pebble game on trees. Journal of Computer and System Sciences, 91(??): 33–41, February 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301186.

Choudhury:2018:RJM

[3357] Anamitra Roy Choudhury, Syamantak Das, Naveen Garg, and Amit Kumar. Rejecting jobs to minimize load and maximum flow-time. Journal of Computer and System Sciences, 91 (??):42–68, February 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301150.

Stout:2018:WIR

[3358] Quentin F. Stout. Weighted L_{∞} isotonic regression. Journal of Computer and System Sciences, 91(??): 69–81, February 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301216.

Kucera:2018:GFC

[3359] Antonín Kucera and Richard Mayr.

A generic framework for checking semantic equivalences between pushdown automata and finite-state automata. Journal of Computer and System Sciences, 91(??):82– 103, February 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301368.

Berkholz:2018:SCP

[3360] Christoph Berkholz and Oleg Verbitsky. On the speed of constraint propagation and the time complexity of arc consistency testing. Journal of Computer and System Sciences, 91 (??):104–114, February 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301356.

Anonymous:2018:EBa

[3361] Anonymous. Editorial Board. Journal of Computer and System Sciences, 91(??):ifc, February 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301411.

Barmpalias:2018:ORC

[3362] George Barmpalias and Andrew Lewis-Pye. Optimal redundancy in computations from random oracles. Journal of Computer and System Sciences, 92(??):1–8, March 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301095.

Agrawal:2018:KDC

[3363] Akanksha Agrawal, Saket Saurabh, Roohani Sharma, and Meirav Zehavi. Kernels for deletion to classes of acyclic digraphs. Journal of Computer and System Sciences, 92(??): 9–21, March 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301174.

Kanj:2018:PAR

[3364] Iyad Kanj, Christian Komusiewicz, Manuel Sorge, and Erik Jan van Leeuwen. Parameterized algorithms for recognizing monopolar and 2subcolorable graphs. Journal of Computer and System Sciences, 92(??): 22–47, March 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301149.

Ivanyos:2018:CTE

[3365] Gábor Ivanyos, Raghav Kulkarni, Youming Qiao, Miklos Santha, and Aarthi Sundaram. On the complexity of trial and error for constraint satisfaction problems. Journal of Computer and System Sciences, 92 (??):48–64, March 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301137.

Kostylev:2018:CQG

[3366] Egor V. Kostylev, Juan L. Reutter, and Domagoj Vrgoc. Containment of queries for graphs with data. Journal of Computer and System Sciences, 92(??):65–91, March 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001730137X.

Chen:2018:AIS

[3367] Li-Hsuan Chen, Dun-Wei Cheng, Sun-Yuan Hsieh, Ling-Ju Hung, Ralf Klasing, Chia-Wei Lee, and Bang Ye Wu. Approximability and inapproximability of the star p-hub center problem with parameterized triangle inequality. Journal of Computer and System Sciences, 92(??): 92–112, March 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301605.

Gebler:2018:SSU

[3368] Daniel Gebler and Simone Tini. SOS specifications for uniformly continuous operators. Journal of Computer and System Sciences, 92(??): 113–151, March 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301599.

Yang:2018:PCM

[3369] Wangdong Yang, Kenli Li, and Keqin Li. A parallel computing method using blocked format with optimal partitioning for SpMV on GPU. Journal of Computer and System Sciences, 92 (??):152–170, March 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301587.

Bonnet:2018:TAT

[3370] Édouard Bonnet, Michael Lampis, and Vangelis Th. Paschos. Timeapproximation trade-offs for inapproximable problems. Journal of Computer and System Sciences, 92(??): 171–180, March 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301575.

Anonymous:2018:EBb

[3371] Anonymous. Editorial Board. Journal of Computer and System Sciences, 92(??):ifc, March 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302155.

Anonymous:2018:EBc

[3372] Anonymous. Editorial Board. Journal of Computer and System Sciences, 93(??):ii, May 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302349.

Wrochna:2018:RBB

[3373] Marcin Wrochna. Reconfiguration in bounded bandwidth and tree-depth. Journal of Computer and System Sciences, 93(??):1–10, May 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302246.

Chierichetti:2018:DPC

[3374] Flavio Chierichetti, Jon Kleinberg, and Sigal Oren. On discrete preferences and coordination. Journal of Computer and System Sciences, 93(??):11–29, May 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302234.

Fellows:2018:PAF

[3375] Michael R. Fellows, Ariel Kulik, Frances Rosamond, and Hadas Shachnai. Parameterized approximation via fidelity preserving transformations. Journal of Computer and System Sciences, 93(??):30–40, May 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000017302222.

Abu-Affash:2018:DPA

[3376] A. Karim Abu-Affash, Paz Carmi, and Anat Parush Tzur. Dual power assignment via second Hamiltonian cycle. Journal of Computer and System Sciences, 93(??): 41–53, May 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017302106.

Anonymous:2018:EBd

[3377] Anonymous. Editorial Board. Journal of Computer and System Sciences, 94 ((supplement)):ii, June 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001830093X.

Segal:2018:JCS

[3378] Michael Segal. Journal of Computer and System Science: 50 years of celebration. In memory of Professor Edward Blum. Journal of Computer and System Sciences, 94 ((supplement)):1, June 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018300898.

Fellows:2018:BHE

[3379] Michael Fellows and Frances Rosamond. A brief history of Edward K. Blum and the Journal of Computer and System Sciences. Journal of Computer and System Sciences, 94((supplement)):

2–10, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018300886.

Gottlob:2018:TPC

[3380] Georg Gottlob, Gianluigi Greco, and Francesco Scarcello. Tree projections and constraint optimization problems: Fixed-parameter tractability and parallel algorithms. Journal of Computer and System Sciences, 94((supplement)):11–40, June 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200001730226X.

Lam:2018:SOH

[3381] William Lam, Kalev Kask, Javier Larrosa, and Rina Dechter. Subproblem ordering heuristics for AND/OR bestfirst search. Journal of Computer and System Sciences, 94((supplement)): 41–62, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302118.

Canetti:2018:TSP

[3382] Ran Canetti, Ling Cheung, Dilsun Kaynar, Moses Liskov, Nancy Lynch, Olivier Pereira, and Roberto Segala. Task-structured probabilistic I/O automata. Journal of Computer and System Sciences, 94((supplement)): 63–97, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301393.

Afrati:2018:CMU

[3383] Foto N. Afrati, Shantanu Sharma, Jonathan R. Ullman, and Jeffrey D. Ullman. Computing marginals using MapReduce. Journal of Computer and System Sciences, 94((supplement)): 98–117, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300260.

Agarwal:2018:RMQ

[3384] Pankaj K. Agarwal, Nirman Kumar, Stavros Sintos, and Subhash Suri. Range-max queries on uncertain data. Journal of Computer and System Sciences, 94((supplement)): 118–134, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017301381.

Buhmann:2018:ROP

[3385] J. M. Buhmann, A. Y. Gronskiy, M. Mihalák, T. Pröger, R. Šrámek, and P. Widmayer. Robust optimization in the presence of uncertainty: A generic approach. Journal of Computer and System Sciences, 94((supplement)): 135–166, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001730212X.

Goldberg:2018:TUC

[3386] Paul W. Goldberg and Christos H. Papadimitriou. Towards a unified complexity theory of total functions. Journal of Computer and System Sciences, 94((supplement)): 167–192, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302878.

Pelc:2018:UIM

[3387] Andrzej Pelc. Use of information, memory and randomization in asynchronous gathering. Journal of Computer and System Sciences, 94((supplement)): 193–205, June 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302131.

Anonymous:2018:EBe

[3388] Anonymous. Editorial Board. Journal of Computer and System Sciences, 95(??):ii, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018302101.

Lin:2018:RBE

[3389] Limei Lin, Sun-Yuan Hsieh, Li Xu, Shuming Zhou, and Riqing Chen. The relationship between extra connectivity and conditional diagnosability of regular graphs under the PMC model. Journal of Computer and System Sciences, 95(??):1–18, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302258.

Gallego:2018:GID

[3390] Antonio-Javier Gallego, Damián López, and Jorge Calera-Rubio. Grammatical inference of directed acyclic graph languages with polynomial time complexity. Journal of Computer and System Sciences, 95(??): 19–34, August 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302866.

Trejo:2018:AAD

[3391] Kristal K. Trejo, Julio B. Clempner, and Alexander S. Poznyak. Adapting attackers and defenders patrolling strategies: a reinforcement learning approach for Stackelberg security games. Journal of Computer and System Sciences, 95(??): 35–54, August 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018300035.

Guan:2018:DQM

[3392] Ji Guan, Yuan Feng, and Mingsheng Ying. Decomposition of quantum Markov chains and its applications. Journal of Computer and System Sciences, 95(??):55– 68, August 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301223.

Gutin:2018:DDP

[3393] Gregory Gutin, Felix Reidl, Magnus Wahlström, and Meirav Zehavi. Designing deterministic polynomialspace algorithms by color-coding multivariate polynomials. Journal of Computer and System Sciences, 95 (??):69–85, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301211.

Gutin:2018:DBD

[3394] Gregory Gutin, Felix Reidl, and Magnus Wahlström. k-distinct in- and out-branchings in digraphs. Journal of Computer and System Sciences, 95 (??):86–97, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001830120X.

Bienvenu:2018:AIP

[3395] Laurent Bienvenu, Santiago Figueira, Benoit Monin, and Alexander Shen. Algorithmic identification of probabilities is hard. Journal of Computer and System Sciences, 95(??): 98–108, August 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301193.

Ke:2018:UIV

[3396] Yuping Ke, Yixin Cao, Xiating Ouyang, Wenjun Li, and Jianxin Wang. Unit interval vertex deletion: Fewer vertices are relevant. Journal of Computer and System Sciences, 95 (??):109–121, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301181.

Lokshtanov:2018:RSG

[3397] Daniel Lokshtanov, Amer E. Mouawad, Fahad Panolan, M. S. Ramanujan, and Saket Saurabh. Reconfiguration on sparse graphs. Journal of Computer and System Sciences, 95 (??):122–131, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018301442.

Brimkov:2018:DGT

[3398] Valentin E. Brimkov and Reneta P. Barneva. Discrete geometry and topology and their applications to imaging sciences. Journal of Computer and System Sciences, 95(??): 132–133, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018302046.

Gonzalez-Diaz:2018:TTC

[3399] Rocio Gonzalez-Diaz, Maria-Jose Jimenez, and Belen Medrano. Topological tracking of connected components in image sequences. Journal of Computer and System Sciences, 95 (??):134–142, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018300047.

Slapal:2018:CAI

[3400] Josef Slapal. Categorical aspects of inducing closure operators on graphs by sets of walks. Journal of Computer and System Sciences, 95(??): 143–150, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300247.

Bhunre:2018:TAV

[3401] Piyush Kanti Bhunre and Partha Bhowmick. Topological analysis of voxelized objects by discrete geodesic Reeb graph. Journal of Computer and System Sciences, 95(??):151– 164, August 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301344.

Han:2018:HPM

[3402] Sang-Eon Han. Homotopic properties of an MA-digitization of 2D Euclidean spaces. Journal of Computer and System Sciences, 95(??): 165–176, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301101.

Nasser:2018:DPD

[3403] Hayat Nasser, Phuc Ngo, and Isabelle Debled-Rennesson. Dominant point detection based on discrete curve structure and applications. Journal of Computer and System Sciences, 95 (??):177–192, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301162.

Mazo:2018:ODT

[3404] Loïc Mazo and Étienne Baudrier.

Object digitization up to a translation. Journal of Computer and System Sciences, 95(??):193–203, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000017301113.

Sarkar:2018:FLR

[3405] Apurba Sarkar, Arindam Biswas, Mousumi Dutt, and Arnab Bhattacharya. Finding a largest rectangle inside a digital object and rectangularization. Journal of Computer and System Sciences, 95(??): 204–217, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300648.

Zrour:2018:OCS

[3406] Rita Zrour, Gaelle Largeteau-Skapin, and Eric Andres. Optimal consensus set for digital Flake hyperspheres in n D. Journal of Computer and System Sciences, 95(??): 218–231, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302556.

Fernau:2018:SPP

[3407] Henning Fernau, Meenakshi Paramasivan, Markus L. Schmid, and Gnanaraj Thomas D. Simple picture processing based on finite automata and regular grammars. Journal of Computer and System Sciences, 95 (??):232–258, August 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301204.

Anonymous:2018:EBf

[3408] Anonymous. Editorial Board. Journal of Computer and System Sciences, 96(??):ii, September 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001830535X.

Chen:2018:OEA

[3409] Lin Chen, Klaus Jansen, and Guochuan Zhang. On the optimality of exact and approximation algorithms for scheduling problems. Journal of Computer and System Sciences, 96(??): 1–32, September 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018302320.

Dosa:2018:TAA

[3410] György Dósa and Leah Epstein. The tight asymptotic approximation ratio of first fit for bin packing with cardinality constraints. Journal of Computer and System Sciences, 96 (??):33–49, September 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018302319.

Dolev:2018:PSS

[3411] Shlomi Dolev, Chryssis Georgiou, Ioannis Marcoullis, and Elad M. Schiller. Practically-self-stabilizing virtual synchrony. Journal of Computer and System Sciences, 96(??): 50–73, September 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001830432X.

Lucas:2018:DPF

[3412] Salvador Lucas, José Meseguer, and Raúl Gutiérrez. The 2D dependency pair framework for conditional rewrite systems. Part I: Definition and basic processors. Journal of Computer and System Sciences, 96(??): 74–106, September 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018304318.

Golovnev:2018:LGE

[3413] Alexander Golovnev, Edward A. Hirsch, Alexander Knop, and Alexander S. Kulikov. On the limits of gate elimination. Journal of Computer and System Sciences, 96(??): 107–119, September 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018305166.

Anonymous:2018:EBg

[3414] Anonymous. Editorial Board. Journal of Computer and System Sciences, 97(??):ii, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018306512.

Goyal:2018:FES

[3415] Prachi Goyal, Pranabendu Misra, Fahad Panolan, Geevarghese Philip, and Saket Saurabh. Finding even subgraphs even faster. Journal of Computer and System Sciences, 97(??):1– 13, November 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018302289.

Dalmau:2018:TCC

[3416] Víctor Dalmau, Andrei Krokhin, and Rajsekar Manokaran. Towards a characterization of constant-factor approximable finite-valued CSPs. Journal of Computer and System Sciences, 97 (??):14–27, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018302307.

Shi:2018:PAM

[3417] Feng Shi, Jianer Chen, Qilong Feng, and Jianxin Wang. A parameterized algorithm for the maximum agreement forest problem on multiple rooted multifurcating trees. Journal of Computer and System Sciences, 97(??): 28–44, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018302290.

Cheraghchi:2018:AML

[3418] Mahdi Cheraghchi, Elena Grigorescu, Brendan Juba, Karl Wimmer, and Ning Xie. AC^{0} · MOD_{2} lower bounds for the Boolean inner product. Journal of Computer and System Sciences, 97 (??):45–59, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001830518X.

Mertzios:2018:SBE

[3419] George B. Mertzios and Paul G. Spirakis. Strong bounds for evolution in networks. Journal of Computer and System Sciences, 97(??): 60–82, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018304781.

Cai:2018:OSE

[3420] Guangya Cai and Daowen Qiu. Optimal separation in exact query complexities for Simon’s problem. Journal of Computer and System Sciences, 97 (??):83–93, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018305178.

Ye:2018:IAM

[3421] Jhih-Hong Ye, Chih-Yu Li, and Biing-Feng Wang. An improved algorithm for the minmax regret path centdian problem on trees. Journal of Computer and System Sciences, 97(??): 94–105, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018306019.

Balliu:2018:WCV

[3422] Alkida Balliu, Gianlorenzo D’Angelo, Pierre Fraigniaud, and Dennis Olivetti. What can be verified locally? Journal of Computer and System Sciences, 97 (??):106–120, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018306032.

Eiben:2018:SEF

[3423] Eduard Eiben, Robert Ganian, and O joung Kwon. A single-exponential fixed-parameter algorithm for distancehereditary vertex deletion. Journal of Computer and System Sciences, 97(??): 121–146, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018306044.

Filiot:2018:VPT

[3424] Emmanuel Filiot, Jean-François Raskin, Pierre-Alain Reynier, Frédéric Servais, and Jean-Marc Talbot. Visibly pushdown transducers. Journal of Computer and System Sciences, 97(??): 147–181, November 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018306007.

Becker:2018:AGC

[3425] Florent Becker, Mathieu Chapelle, Jérôme Durand-Lose, Vincent Levorato, and Maxime Senot. Abstract geometrical computation 8: Small machines, accumulations and rationality. Journal of Computer and System Sciences, 97(??):182– 198, November 2018. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018306020.

Anonymous:2018:EBh

[3426] Anonymous. Editorial Board. Journal of Computer and System Sciences, 98(??):ii, December 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018307736.

Liu:2018:TMS

[3427] Xiao Liu, Mianxiong Dong, Kaoru Ota, Laurence T. Yang, and Anfeng Liu. Trace malicious source to guarantee cyber security for mass monitor critical infrastructure. Journal of Computer and System Sciences, 98 (??):1–26, December 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016300885.

Tosh:2018:EEG

[3428] Deepak Tosh, Shamik Sengupta, Charles A. Kamhoua, and Kevin A. Kwiat. Establishing evolutionary game models for CYBer security information EXchange (CYBEX). Journal of Computer and System Sciences, 98 (??):27–52, December 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001630085X.

Cohen:2018:AST

[3429] Nachshon Cohen and Zeev Nutov. Approximating Steiner trees and forests with minimum number of Steiner points. Journal of Computer and System Sciences, 98(??): 53–64, December 2018. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018307001.

Anonymous:2019:EBa

[3430] Anonymous. Editorial Board. Journal of Computer and System Sciences, 99(??):ii, February 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308596.

Wang:2019:ERD

[3431] Zheng Wang. An elastic and resiliency defense against DDoS attacks on the critical DNS authoritative infrastructure. Journal of Computer and System Sciences, 99 (??):1–26, February 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001730082X.

Anantharamu:2019:PLD

[3432] Lakshmi Anantharamu, Bogdan S.

Chlebus, Dariusz R. Kowalski, and Mariusz A. Rokicki. Packet latency of deterministic broadcasting in adversarial multiple access channels. Journal of Computer and System Sciences, 99 (??):27–52, February 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018306998.

Gajarsky:2019:PSC

[3433] Jakub Gajarský, Petr Hlinený, Martin Koutecký, and Shmuel Onn. Parameterized shifted combinatorial optimization. Journal of Computer and System Sciences, 99(??): 53–71, February 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018306986.

Didimo:2019:HPA

[3434] Walter Didimo, Giuseppe Liotta, and Maurizio Patrignani. HV-planarity: Algorithms and complexity. Journal of Computer and System Sciences, 99 (??):72–90, February 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018307839.

Epstein:2019:PGF

[3435] Leah Epstein and Asaf Levin. On the performance guarantee of First Fit for sum coloring. Journal of Computer and System Sciences, 99 (??):91–105, February 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018307827.

Anonymous:2019:EBb

[3436] Anonymous. Editorial Board. Journal of Computer and System Sciences, 100(??):ii, March 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018308948.

Mazowiecki:2019:CCR

[3437] Filip Mazowiecki and Cristian Riveros. Copyless cost-register automata: Structure, expressiveness, and closure properties. Journal of Computer and System Sciences, 100(??):1– 29, March 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018307815.

Potapov:2019:VSR

[3438] Igor Potapov and Pavel Semukhin.

Vector and scalar reachability problems in SL(2,Z). Journal of Computer and System Sciences, 100(??): 30–43, March 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308110.

Burdick:2019:EPE

[3439] Douglas Burdick, Ronald Fagin, Phokion G. Kolaitis, Lucian Popa, and Wang-Chiew Tan. Expressive power of entity-linking frameworks. Journal of Computer and System Sciences, 100(??):44–69, March 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308092.

Ben:2019:DEC

[3440] Yue Ben, Rohit Chadha, A. Prasad Sistla, and Mahesh Viswanathan. Decidable and expressive classes of probabilistic automata. Journal of Computer and System Sciences, 100(??): 70–95, March 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308109.

Doyen:2019:CSM

[3441] Laurent Doyen, Thierry Massart, and Mahsa Shirmohammadi. The complexity of synchronizing Markov decision processes. Journal of Computer and System Sciences, 100(??): 96–129, March 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018308468.

Anonymous:2019:EBc

[3442] Anonymous. Editorial Board. Journal of Computer and System Sciences, 101(??):ii, May 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018309061.

Gutin:2019:PCE

[3443] Gregory Gutin, M. S. Ramanujan, Felix Reidl, and Magnus Wahlström. Path-contractions, edge deletions and connectivity preservation. Journal of Computer and System Sciences, 101(??):1–20, May 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018300758.

Bergougnoux:2019:CMT

[3444] Benjamin Bergougnoux, Florent Capelli, and Mamadou Moustapha Kanté. Counting minimal transversals of βacyclic hypergraphs. Journal of Computer and System Sciences, 101(??): 21–30, May 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308900.

Hemaspaandra:2019:RTR

[3445] Lane A. Hemaspaandra and Daniel Rubery. Recursion-theoretic ranking and compression. Journal of Computer and System Sciences, 101 (??):31–41, May 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018300163.

Gawrychowski:2019:HSR

[3446] Pawel Gawrychowski, Florin Manea, Robert Mercas, and Dirk Nowotka. Hide and seek with repetitions. Journal of Computer and System Sciences, 101(??):42–67, May 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308973.

Jain:2019:CVL

[3447] Sanjay Jain, Alexei Miasnikov, and Frank Stephan. The complexity of verbal languages over groups. Journal of Computer and System Sciences, 101(??):68–85, May 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308985.

Cabessa:2019:CCA

[3448] Jérémie Cabessa and Olivier Finkel.

Computational capabilities of analog and evolving neural networks over infinite input streams. Journal of Computer and System Sciences, 101 (??):86–99, May 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301703.

Takahashi:2019:BRD

[3449] Norikazu Takahashi, Daiki Hirata, Shuji Jimbo, and Hiroaki Yamamoto. Band-restricted diagonally dominant matrices: Computational complexity and application. Journal of Computer and System Sciences, 101(??): 100–111, May 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018309024.

Anonymous:2019:EBd

[3450] Anonymous. Editorial Board. Journal of Computer and System Sciences, 102(??):ii, June 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300133.

Balogh:2019:OAR

[3451] János Balogh, József Békési, György Dósa, Jirí Sgall, and Rob van Stee. The optimal absolute ratio for online bin packing. Journal of Computer and System Sciences, 102(??): 1–17, June 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300029.

Anonymous:2019:PJ

[3452] Anonymous. Pages 1–86 (June 2019). Journal of Computer and System Sciences, 102(??):1–86, June 2019. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Michail:2019:TCF

[3453] Othon Michail, George Skretas, and Paul G. Spirakis. On the transformation capability of feasible mechanisms for programmable matter. Journal of Computer and System Sciences, 102(??):18–39, June 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018300436.

Choo:2019:SIC

[3454] Kim-Kwang Raymond Choo, Jemal Abawajy, and Rafiqul Islam. Special issue on cybersecurity in the critical infrastructure: Advances and future directions. Journal of Computer and System Sciences, 102(??): 40–41, June 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300630.

Luo:2019:GGC

[3455] Cheng Luo, Bo Zhang, Yang Xiang, and Man Qi. Gaussian-gamma collaborative filtering: a hierarchical Bayesian model for recommender systems. Journal of Computer and System Sciences, 102(??):42–56, June 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300351.

Dereniowski:2019:CDS

[3456] Dariusz Dereniowski, Andrzej Lingas, Dorota Osula, Mia Persson, and Paweł Zyliński. Clearing directed subgraphs by mobile agents: Variations on covering with paths. Journal of Computer and System Sciences, 102(??):57–68, June 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018309000.

Fluschnik:2019:CRC

[3457] Till Fluschnik, Marco Morik, and Manuel Sorge. The complexity of routing with collision avoidance. Journal of Computer and System Sciences, 102(??):69–86, June 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300030.

Anonymous:2019:EBe

[3458] Anonymous. Editorial Board. Journal of Computer and System Sciences, 103(??):ii, August 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300212.

Alvarez:2019:FSI

[3459] Nicolás Álvarez, Verónica Becher, and Olivier Carton. Finite-state independence and normal sequences. Journal of Computer and System Sciences, 103(??):1–17, August 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300169.

Anonymous:2019:PA

[3460] Anonymous. Pages 1–78 (August 2019). Journal of Computer and System Sciences, 103(??):1–78, August 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Drucker:2019:CRU

[3461] Nir Drucker, Hsi-Ming Ho, Joël Ouaknine, Michal Penn, and Ofer Strichman. Cyclic-routing of unmanned aerial vehicles. Journal of Computer and System Sciences, 103(??): 18–45, August 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302830.

Akrida:2019:TFT

[3462] Eleni C. Akrida, Jurek Czyzowicz, Leszek Gasieniec, Lukasz Kuszner, and Paul G. Spirakis. Temporal flows in temporal networks. Journal of Computer and System Sciences, 103 (??):46–60, August 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301090.

Bentert:2019:PAT

[3463] Matthias Bentert, Till Fluschnik, André Nichterlein, and Rolf Niedermeier. Parameterized aspects of triangle enumeration. Journal of Computer and System Sciences, 103(??): 61–77, August 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300170.

Anonymous:2019:EBf

[3464] Anonymous. Editorial Board. Journal of Computer and System Sciences, 104(??):ii, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300376.

Dediu:2019:SIS

[3465] Adrian-Horia Dediu, Carlos Martín-Vide, and Bianca Truthe. Special issue: Selected papers of the 9th International Conference on Language and Automata Theory and Applications, LATA 2015. Journal of Computer and System Sciences, 104 (??):1–4, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300339.

Allender:2019:CRF

[3466] Eric Allender and Ian Mertz. Complexity of regular functions. Journal of Computer and System Sciences, 104 (??):5–16, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016301015.

Autili:2019:ASA

[3467] Marco Autili, Paola Inverardi, Romina Spalazzese, Massimo Tivoli, and Filippo Mignosi. Automated synthesis of application-layer connectors from automata-based specifications. Journal of Computer and System Sciences, 104 (??):17–40, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300248.

Babari:2019:NTW

[3468] Parvaneh Babari and Manfred Droste.

A Nivat theorem for weighted picture automata and weighted MSO logics. Journal of Computer and System Sciences, 104(??):41–57, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000017300302.

Bailly:2019:RSG

[3469] Raphaël Bailly, Guillaume Rabusseau, and François Denis. Recognizable series on graphs and hypergraphs. Journal of Computer and System Sciences, 104 (??):58–81, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301563.

Beyersdorff:2019:GCT

[3470] Olaf Beyersdorff, Leroy Chew, and Karteek Sreenivasaiah. A game characterisation of tree-like Q-resolution size. Journal of Computer and System Sciences, 104(??):82–101, September 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301350.

Bilotta:2019:RRS

[3471] S. Bilotta, E. Pergola, R. Pinzani, and S. Rinaldi. Recurrence relations, succession rules and the positivity problem. Journal of Computer and System Sciences, 104(??): 102–118, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300861.

Bjorklund:2019:EEW

[3472] Johanna Björklund, Frank Drewes, and Niklas Zechner. Efficient enumeration of weighted tree languages over the tropical semiring. Journal of Computer and System Sciences, 104(??): 119–130, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001730034X.

Boichut:2019:TMP

[3473] Yohan Boichut, Jacques Chabin, and Pierre Réty. Towards more precise rewriting approximations. Journal of Computer and System Sciences, 104(??):131–148, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000017300065.

Boiret:2019:LUT

[3474] Adrien Boiret, Vincent Hugot, Joachim Niehren, and Ralf Treinen. Logics for unordered trees with data constraints. Journal of Computer and System Sciences, 104(??):149– 164, September 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300017.

Cazaux:2019:LID

[3475] Bastien Cazaux, Thierry Lecroq, and Eric Rivals. Linking indexing data structures to de Bruijn graphs: Construction and update. Journal of Computer and System Sciences, 104(??): 165–183, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300502.

Codish:2019:SNE

[3476] Michael Codish, Luís Cruz-Filipe, Thorsten Ehlers, Mike Müller, and Peter Schneider-Kamp. Sorting networks: To the end and back again. Journal of Computer and System Sciences, 104(??):184–201, September 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300162.

Dabrowski:2019:BCW

[3477] Konrad K. Dabrowski, Shenwei Huang, and Daniël Paulusma. Bounding clique-width via perfect graphs. Journal of Computer and System Sciences, 104(??):202–215, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000016300496.

Dai:2019:TPA

[3478] H. K. Dai and Z. Wang. Two parallel algorithms for finding all minimal maximum subsequences. Journal of Computer and System Sciences, 104(??): 216–243, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018308997.

Eremondi:2019:IOD

[3479] Joey Eremondi, Oscar H. Ibarra, and Ian McQuillan. Insertion operations on deterministic reversal-bounded counter machines. Journal of Computer and System Sciences, 104(??): 244–257, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001830117X.

Gamard:2019:CMS

[3480] Guilhem Gamard and Gwenaël Richomme. Coverability and multiscale coverability on infinite pictures. Journal of Computer and System Sciences, 104(??):258–277, September 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000017300594.

Jacquemard:2019:OVC

[3481] Florent Jacquemard and Michael Rusinowitch. One-variable context-free hedge automata. Journal of Computer and System Sciences, 104(??): 278–296, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000016301027.

Janicki:2019:CIS

[3482] Ryszard Janicki, Jetty Kleijn, Maciej Koutny, and Lukasz Mikulski.

Classifying invariant structures of step traces. Journal of Computer and System Sciences, 104(??):297– 322, September 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017300600.

Sulzmann:2019:DPD

[3483] Martin Sulzmann and Peter Thiemann. Derivatives and partial derivatives for regular shuffle expressions. Journal of Computer and System Sciences, 104(??):323– 341, September 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016301325.

Vorel:2019:CRC

[3484] Vojtech Vorel and Adam Roman. Complexity of road coloring with prescribed reset words. Journal of Computer and System Sciences, 104(??): 342–358, September 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000016300447.

Yoshinaka:2019:DLC

[3485] Ryo Yoshinaka. Distributional learning of conjunctive grammars and contextual binary feature grammars. Journal of Computer and System Sciences, 104(??):359–374, September 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017301125.

Anonymous:2019:EBg

[3486] Anonymous. Editorial Board. Journal of Computer and System Sciences, 105(??):ii, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300479.

Freydenberger:2019:DRE

[3487] Dominik D. Freydenberger and Markus L. Schmid. Deterministic regular expressions with back-references. Journal of Computer and System Sciences, 105 (??):1–39, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018301818.

Anonymous:2019:PN

[3488] Anonymous. Pages 1–242 (November 2019). Journal of Computer and System Sciences, 105(??):1–242, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Ramaswamy:2019:SCR

[3489] Vidhya Ramaswamy, Jayalal Sarma, and K. S. Sunil. Space complexity of reachability testing in labelled graphs. Journal of Computer and System Sciences, 105(??): 40–53, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018301120.

Gyssens:2019:CSQ

[3490] Marc Gyssens, Jelle Hellings, Jan Paredaens, Dirk Van Gucht, Jef Wijsen, and Yuqing Wu. Calculi for symmetric queries. Journal of Computer and System Sciences, 105(??): 54–86, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018305713.

Sakai:2019:BDC

[3491] Takayuki Sakai, Kazuhisa Seto, Suguru Tamaki, and Junichi Teruyama. Bounded depth circuits with weighted symmetric gates: Satisfiability, lower bounds and compression. Journal of Computer and System Sciences, 105 (??):87–103, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302659.

Park:2019:FTE

[3492] Jung-Heum Park, Hyeong-Seok Lim, and Hee-Chul Kim. Fault-tolerant embedding of starlike trees into restricted hypercube-like graphs. Journal of Computer and System Sciences, 105 (??):104–115, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001930025X.

Bartal:2019:NDA

[3493] Yair Bartal, Arnold Filtser, and Ofer Neiman. On notions of distortion and an almost minimum spanning tree with constant average distortion. Journal of Computer and System Sciences, 105 (??):116–129, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018300709.

Bachmeier:2019:MDH

[3494] Georg Bachmeier, Felix Brandt, Christian Geist, Paul Harrenstein, Keyvan Kardel, Dominik Peters, and Hans Georg Seedig. k-majority digraphs and the hardness of voting with a constant number of voters. Journal of Computer and System Sciences, 105 (??):130–157, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300261.

Eiben:2019:APD

[3495] Eduard Eiben, Danny Hermelin, and M. S. Ramanujan. On approximate preprocessing for domination and hitting subgraphs with connected deletion sets. Journal of Computer and System Sciences, 105(??): 158–170, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018307098.

Atserias:2019:GSP

[3496] Albert Atserias, Phokion G. Kolaitis, and Simone Severini. Generalized satisfiability problems via operator assignments. Journal of Computer and System Sciences, 105(??): 171–198, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300406.

Demaine:2019:SSC

[3497] Erik D. Demaine, Felix Reidl, Peter Rossmanith, Fernando Sánchez Villaamil, Somnath Sikdar, and Blair D. Sullivan. Structural sparsity of complex networks: Bounded expansion in random models and realworld graphs. Journal of Computer and System Sciences, 105(??): 199–241, November 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300418.

Anonymous:2019:EBh

[3498] Anonymous. Editorial Board. Journal of Computer and System Sciences, 106(??):ii, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300601.

Filiot:2019:DWE

[3499] Emmanuel Filiot, Nicolas Mazzocchi, and Jean-François Raskin. Decidable weighted expressions with Presburger combinators. Journal of Computer and System Sciences, 106(??): 1–22, December 2019. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001930042X.

Anonymous:2019:PD

[3500] Anonymous. Pages 1–144 (December 2019). Journal of Computer and System Sciences, 106(??):1–144, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Fluschnik:2019:PCM

[3501] Till Fluschnik, Stefan Kratsch, Rolf Niedermeier, and Manuel Sorge. The parameterized complexity of the minimum shared edges problem. Journal of Computer and System Sciences, 106 (??):23–48, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300431.

Farrugia:2019:DRD

[3502] Ashley Farrugia, Leszek Gasieniec, Lukasz Kuszner, and Eduardo Pacheco. Deterministic rendezvous with different maps. Journal of Computer and System Sciences, 106(??): 49–59, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018301594.

Gaspers:2019:CSF

[3503] Serge Gaspers, Shenwei Huang, and Daniël Paulusma. Colouring square-free graphs without long induced paths. Journal of Computer and System Sciences, 106(??): 60–79, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018308146.

Kosowski:2019:DAM

[3504] Adrian Kosowski and Dominik Pajak. Does adding more agents make a difference? A case study of cover time for the rotor-router. Journal of Computer and System Sciences, 106 (??):80–93, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018300126.

Koyano:2019:OSC

[3505] Hitoshi Koyano, Morihiro Hayashida, and Tatsuya Akutsu. Optimal string clustering based on a Laplace-like mixture and EM algorithm on a set of strings. Journal of Computer and System Sciences, 106(??): 94–128, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300509.

Liskiewicz:2019:PLT

[3506] Maciej Liśkiewicz, Matthias Lutter, and Rüdiger Reischuk. Proper learning of k-term DNF formulas from satisfying assignments. Journal of Computer and System Sciences, 106(??): 129–144, December 2019. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300510.

Anonymous:2020:EBa

[3507] Anonymous. Editorial Board. Journal of Computer and System Sciences, 107(??):ii, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300716.

Guo:2020:CPB

[3508] Heng Guo and Tyson Williams. The complexity of planar Boolean #CSP with complex weights. Journal of Computer and System Sciences, 107 (??):1–27, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300534.

Anonymous:2020:PF

[3509] Anonymous. Pages 1–166 (February 2020). Journal of Computer and System Sciences, 107(??):1–166, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Frati:2020:LDO

[3510] Fabrizio Frati, Maurizio Patrignani, and Vincenzo Roselli. LR-drawings of ordered rooted binary trees and near-linear area drawings of outerplanar graphs. Journal of Computer and System Sciences, 107(??): 28–53, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000017302404.

Jimenez:2020:CUO

[3511] Ernesto Jiménez, José Luis López-Presa, and Javier Martín-Rueda. Consensus using omega in asynchronous systems with unknown membership and degenerative Byzantine failures. Journal of Computer and System Sciences, 107(??):54– 71, February 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300522.

Zschoche:2020:CFS

[3512] Philipp Zschoche, Till Fluschnik, Hendrik Molter, and Rolf Niedermeier. The complexity of finding small separators in temporal graphs. Journal of Computer and System Sciences, 107 (??):72–92, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300546.

Gajser:2020:VWO

[3513] David Gajser. Verifying whether one-tape Turing machines run in linear time. Journal of Computer and System Sciences, 107(??):93– 107, February 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300650.

Akrida:2020:TVC

[3514] Eleni C. Akrida, George B. Mertzios, Paul G. Spirakis, and Viktor Zamaraev. Temporal vertex cover with a sliding time window. Journal of Computer and System Sciences, 107(??): 108–123, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018308687.

Niskanen:2020:DCL

[3515] R. Niskanen, I. Potapov, and J. Reichert. On decidability and complexity of low-dimensional robot games. Journal of Computer and System Sciences, 107(??):124–141, February 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300649.

Hromkovic:2020:WOK

[3516] Juraj Hromkovic and Peter Rossmanith. What one has to know when attacking P vs. NP. Journal of Computer and System Sciences, 107(??): 142–155, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300558.

Anshu:2020:CMP

[3517] Anurag Anshu, Peter Høyer, Mehdi Mhalla, and Simon Perdrix. Contextuality in multipartite pseudo-telepathy graph games. Journal of Computer and System Sciences, 107(??): 156–165, February 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200001930056X.

Anonymous:2020:EBb

[3518] Anonymous. Editorial Board. Journal of Computer and System Sciences, 108(??):ii, March 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000019300789.

Cheng:2020:VSE

[3519] Chia-Wen Cheng, Sun-Yuan Hsieh, and Ralf Klasing. Vulnerability of super extra edge-connected graphs. Journal of Computer and System Sciences, 108(??):1–9, March 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301661.

Anonymous:2020:PMa

[3520] Anonymous. Pages 1–136 (March 2020). Journal of Computer and System Sciences, 108(??):1–136, March 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Erickson:2020:RBW

[3521] Alejandro Erickson, Javier Navaridas, and Iain A. Stewart. Relating the bisection width of dual-port, servercentric datacenter networks and the solution of edge isoperimetric problems in graphs. Journal of Computer and System Sciences, 108(??): 10–28, March 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300662.

Bagan:2020:TRS

[3522] Guillaume Bagan, Angela Bonifati, and Benoit Groz. A trichotomy for regular simple path queries on graphs. Journal of Computer and System Sciences, 108(??):29–48, March 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018305312.

Bienvenu:2020:LSO

[3523] Laurent Bienvenu and Rod Downey.

On low for speed oracles. Journal of Computer and System Sciences, 108 (??):49–63, March 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018305828.

Dabrowski:2020:CWW

[3524] Konrad K. Dabrowski, Vadim V. Lozin, and Daniël Paulusma. Clique-width and well-quasi-ordering of trianglefree graph classes. Journal of Computer and System Sciences, 108(??): 64–91, March 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018305087.

Aisenberg:2020:DTP

[3525] James Aisenberg, Maria Luisa Bonet, and Sam Buss. 2-D Tucker is PPA complete. Journal of Computer and System Sciences, 108(??): 92–103, March 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300674.

Bordihn:2020:DNR

[3526] Henning Bordihn and Victor Mitrana. On the degrees of non-regularity and non-context-freeness. Journal of Computer and System Sciences, 108 (??):104–117, March 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018304537.

Michaliszyn:2020:NDW

[3527] Jakub Michaliszyn and Jan Otop. Nondeterministic weighted automata evaluated over Markov chains. Journal of Computer and System Sciences, 108 (??):118–136, March 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300741.

Anonymous:2020:PMb

[3528] Anonymous. Pages 1–144 (May 2020).

Journal of Computer and System Sciences, 109(??):1–144, May 2020. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2020:EBc

[3529] Anonymous. Editorial Board. Journal of Computer and System Sciences, 109(??):ii, May 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300064.

Bridoux:2020:CSA

[3530] Florian Bridoux, Alonso Castillo-Ramirez, and Maximilien Gadouleau. Complete simulation of automata networks. Journal of Computer and System Sciences, 109(??):1– 21, May 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300820.

Jancar:2020:DSF

[3531] Petr Jancar. Deciding semantic finiteness of pushdown processes and first-order grammars w.r.t. bisimulation equivalence. Journal of Computer and System Sciences, 109(??): 22–44, May 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300844.

Kolay:2020:FGB

[3532] Sudeshna Kolay, Pranabendu Misra, M. S. Ramanujan, and Saket Saurabh. Faster graph bipartization. Journal of Computer and System Sciences, 109(??):45–55, May 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018307050.

Baste:2020:HMB

[3533] Julien Baste, Ignasi Sau, and Dimitrios M. Thilikos. Hitting minors on bounded treewidth graphs. III. Lower bounds. Journal of Computer and System Sciences, 109(??): 56–77, May 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300832.

Ding:2020:LTV

[3534] Hu Ding and Jinhui Xu. Learning the truth vector in high dimensions. Journal of Computer and System Sciences, 109(??):78–94, May 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018305270.

Backens:2020:BAC

[3535] Miriam Backens, Andrei Bulatov, Leslie Ann Goldberg, Colin McQuillan, and Stanislav Zivný. Boolean approximate counting CSPs with weak conservativity, and implications for ferromagnetic two-spin. Journal of Computer and System Sciences, 109 (??):95–125, May 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018304203.

Okrasa:2020:SAV

[3536] Karolina Okrasa and Paweł Rzazewski.

Subexponential algorithms for variants of the homomorphism problem in string graphs. Journal of Computer and System Sciences, 109(??): 126–144, May 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000019300856.

Anonymous:2020:PJ

[3537] Anonymous. Pages 1–54 (June 2020).

Journal of Computer and System Sciences, 110(??):1–54, June 2020. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2020:EBd

[3538] Anonymous. Editorial Board. Journal of Computer and System Sciences, 110(??):ii, June 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300179.

Eiben:2020:UDP

[3539] Eduard Eiben, Robert Ganian, and Sebastian Ordyniak. Using decompositionparameters for QBF: Mind the prefix! Journal of Computer and System Sciences, 110(??):1– 21, June 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300015.

Kupavskii:2020:WEN

[3540] Andrey Kupavskii and Nikita Zhivotovskiy. When are epsilon-nets small? Journal of Computer and System Sciences, 110(??):22– 36, June 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300027.

Rosenke:2020:CSE

[3541] Christian Rosenke and Ronny Tredup. The complexity of synthesizing elementary net systems relative to natural parameters. Journal of Computer and System Sciences, 110(??): 37–54, June 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200002030009X.

Anonymous:2020:PA

[3542] Anonymous. Pages 1–84 (August 2020). Journal of Computer and System Sciences, 111(??):1–84, August 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2020:EBe

[3543] Anonymous. Editorial Board. Journal of Computer and System Sciences, 111(??):ii, August 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300271.

Loff:2020:CPP

[3544] Bruno Loff, Nelma Moreira, and Rogério Reis. The computational power of parsing expression grammars. Journal of Computer and System Sciences, 111(??):1–21, August 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S002200002030012X.

Melo:2020:MAS

[3545] Alexsander A. Melo, Celina M. H.

Figueiredo, and Uéverton S. Souza. A multivariate analysis of the strict terminal connection problem. Journal of Computer and System Sciences, 111(??):22–41, August 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018307645.

DeMarco:2020:SNA

[3546] Gianluca De Marco, Tomasz Jurdziński, Dariusz R. Kowalski, Michał Rózański, and Grzegorz Stachowiak. Subquadratic non-adaptive threshold group testing. Journal of Computer and System Sciences, 111(??): 42–56, August 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300131.

Tang:2020:DPB

[3547] Qiyi Tang and Franck van Breugel.

Deciding probabilistic bisimilarity distance one for probabilistic automata. Journal of Computer and System Sciences, 111(??):57– 84, August 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300209.

Anonymous:2020:PS

[3548] Anonymous. Pages 1–96 (September 2020). Journal of Computer and System Sciences, 112(??):1–96, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2020:EBf

[3549] Anonymous. Editorial Board. Journal of Computer and System Sciences, 112(??):ii, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300428.

Downey:2020:GUO

[3550] Rodney Downey, Matthew Harrison-Trainor, Iskander Kalimullin, Alexander Melnikov, and Daniel Turetsky. Graphs are not universal for online computability. Journal of Computer and System Sciences, 112(??): 1–12, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300210.

Baumeister:2020:CCJ

[3551] Dorothea Baumeister, Gábor Erdélyi, Olivia J. Erdélyi, Jörg Rothe, and Ann-Kathrin Selker. Complexity of control in judgment aggregation for uniform premise-based quota rules. Journal of Computer and System Sciences, 112 (??):13–33, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300222.

Balogh:2020:OPC

[3552] János Balogh, József Békési, György Dósa, Leah Epstein, and Asaf Levin. Online bin packing with cardinality constraints resolved. Journal of Computer and System Sciences, 112(??): 34–49, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018300722.

Dragan:2020:ETH

[3553] Feodor F. Dragan and Heather M.

Guarnera. Eccentricity terrain of δhyperbolic graphs. Journal of Computer and System Sciences, 112(??): 50–65, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300350.

Hsieh:2020:MLB

[3554] Sun-Yuan Hsieh, Sin-Jie Huang, and Tai-Hsuan Ho. Multilevel lengthbased classified index table for IP lookups and updates. Journal of Computer and System Sciences, 112 (??):66–84, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300362.

Censor-Hillel:2020:DRM

[3555] Keren Censor-Hillel and Mikaël Rabie. Distributed reconfiguration of maximal independent sets. Journal of Computer and System Sciences, 112 (??):85–96, September 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300349.

Anonymous:2020:PN

[3556] Anonymous. Pages 1–178 (November 2020). Journal of Computer and System Sciences, 113(??):1–178, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2020:EBg

[3557] Anonymous. Editorial Board. Journal of Computer and System Sciences, 113(??):ii, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300581.

Montealegre:2020:GRC

[3558] P. Montealegre, S. Perez-Salazar, I. Rapaport, and I. Todinca. Graph reconstruction in the congested clique. Journal of Computer and System Sciences, 113(??):1–17, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000018306329.

Reghizzi:2020:BOP

[3559] Stefano Crespi Reghizzi and Matteo Pradella. Beyond operator-precedence grammars and languages. Journal of Computer and System Sciences, 113 (??):18–41, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001830182X.

Martin:2020:DCC

[3560] Barnaby Martin, Daniël Paulusma, and Erik Jan van Leeuwen. Disconnected cuts in claw-free graphs. Journal of Computer and System Sciences, 113 (??):60–75, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018306056.

Chalopin:2020:CTC

[3561] Jérémie Chalopin and Victor Chepoi.

A counterexample to Thiagarajan’s conjecture on regular event structures. Journal of Computer and System Sciences, 113(??):76–100, November 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018301041.

Golovach:2020:FCS

[3562] Petr A. Golovach, Pinar Heggernes, Paloma T. Lima, and Pedro Montealegre. Finding connected secluded subgraphs. Journal of Computer and System Sciences, 113(??): 101–124, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300519.

Misra:2020:SAC

[3563] Neeldhara Misra, Fahad Panolan, and Saket Saurabh. Subexponential algorithm for d-cluster edge deletion: Exception or rule? Journal of Computer and System Sciences, 113(??): 150–162, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300532.

Pu:2020:AAU

[3564] Lianrong Pu, Daming Zhu, and Haitao Jiang. A 1.375-approximation algorithm for unsigned translocation sorting. Journal of Computer and System Sciences, 113(??):163– 178, November 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300490.

Korman:2020:MRC

[3565] Amos Korman and Yoav Rodeh. Multiround cooperative search games with multiple players. Journal of Computer and System Sciences, 113(??): 125–149, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300489.

Arvind:2020:WLI

[3566] V. Arvind, Frank Fuhlbrück, Johannes Köbler, and Oleg Verbitsky. On Weisfeiler–Leman invariance: Subgraph counts and related graph properties. Journal of Computer and System Sciences, 113(??): 42–59, November 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300386.

Anonymous:2020:PD

[3567] Anonymous. Pages 1–146 (December 2020). Journal of Computer and System Sciences, 114(??):1–146, December 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2020:EBh

[3568] Anonymous. Editorial Board. Journal of Computer and System Sciences, 114(??):ii, December 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300763.

Fearnley:2020:UEP

[3569] John Fearnley, Spencer Gordon, Ruta Mehta, and Rahul Savani. Unique end of potential line. Journal of Computer and System Sciences, 114 (??):1–35, December 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300520.

Wang:2020:IAM

[3570] Biing-Feng Wang, Jhih-Hong Ye, and Chih-Yu Li. An improved algorithm for the minmax regret path center problem on trees. Journal of Computer and System Sciences, 114(??): 36–47, December 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018305531.

Chlebus:2020:USM

[3571] Bogdan S. Chlebus, Vicent Cholvi, and Dariusz R. Kowalski. Universal stability in multi-hop radio networks. Journal of Computer and System Sciences, 114(??):48– 64, December 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300544.

Cardona:2020:CET

[3572] Gabriel Cardona and Louxin Zhang. Counting and enumerating tree-child networks and their subclasses. Journal of Computer and System Sciences, 114 (??):84–104, December 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300611.

Emura:2020:SCF

[3573] Keita Emura, Katsuhiko Ito, and Toshihiro Ohigashi. Secure-channel free searchable encryption with multiple keywords: a generic construction, an instantiation, and its implementation. Journal of Computer and System Sciences, 114(??):107– 125, December 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000018304598.

Duris:2020:THD

[3574] Pavol Duris, Rastislav Královic, and Dana Pardubská. Tight hierarchy of data-independent multi-head automata. Journal of Computer and System Sciences, 114(??):126– 136, December 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300647.

Akrida:2020:HFC

[3575] Eleni C. Akrida, George B. Mertzios, Sotiris Nikoletseas, Christoforos Raptopoulos, Paul G. Spirakis, and Viktor Zamaraev. How fast can we reach a target vertex in stochastic temporal graphs? Journal of Computer and System Sciences, 114(??): 65–83, December 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300507.

Klasing:2020:SPI

[3576] Ralf Klasing and Marc Zeitoun. Selected papers of the 21st International Symposium on Fundamentals of Computation Theory, FCT 2017. Journal of Computer and System Sciences, 114 (??):105–106, December 2020. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300635.

Das:2020:SER

[3577] Bireswar Das, Shivdutt Sharma, and P. R. Vaidyanathan. Space efficient representations of finite groups. Journal of Computer and System Sciences, 114(??):137–146, December 2020. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300660.

Anonymous:2021:PF

[3578] Anonymous. Pages 1–264 (February 2021). Journal of Computer and System Sciences, 115(??):1–264, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBa

[3579] Anonymous. Editorial Board. Journal of Computer and System Sciences, 115(??):ii, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300891.

Dereniowski:2021:SHA

[3580] Dariusz Dereniowski, Lukasz Kuszner, and Robert Ostrowski. Searching by heterogeneous agents. Journal of Computer and System Sciences, 115 (??):1–21, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300672.

Mukdasanit:2021:ADL

[3581] Supachai Mukdasanit and Sanpawat Kantabutra. Attack and defense in the layered cyber-security model and their (1±ϵ)-approximation schemes. Journal of Computer and System Sciences, 115 (??):54–63, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300684.

Zhao:2021:EFT

[3582] Shuang Zhao, Zongqing Chen, Weihua Yang, and Jixiang Meng. Edge fault-tolerance analysis of maximally edge-connected graphs and super edgeconnected graphs. Journal of Computer and System Sciences, 115(??): 64–72, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300696.

Jancar:2021:EPA

[3583] Petr Jancar. Equivalence of pushdown automata via first-order grammars. Journal of Computer and System Sciences, 115(??):86–112, February 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300714.

Kuijpers:2021:DPF

[3584] Bart Kuijpers. Deciding the pointto-fixed-point problem for skew tent maps on an interval. Journal of Computer and System Sciences, 115(??): 113–120, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200001830237X.

Moses:2021:DPS

[3585] William K. Moses, Jr. and Shailesh Vaya. Deterministic protocols in the SINR model without knowledge of coordinates. Journal of Computer and System Sciences, 115(??): 121–145, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300623.

Deng:2021:UPC

[3586] Xiaotie Deng, Jack R. Edmonds, Zhe Feng, Zhengyang Liu, Qi Qi, and Zeying Xu. Understanding PPAcompleteness. Journal of Computer and System Sciences, 115(??):146– 168, February 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300726.

Enright:2021:ATM

[3587] Jessica Enright, Kitty Meeks, and Fiona Skerman. Assigning times to minimise reachability in temporal graphs. Journal of Computer and System Sciences, 115(??):169– 186, February 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300799.

Galanis:2021:APF

[3588] Andreas Galanis, Leslie Ann Goldberg, and Kuan Yang. Approximating partition functions of boundeddegree Boolean counting constraint satisfaction problems. Journal of Computer and System Sciences, 115(??): 187–213, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300817.

Berlinkov:2021:PPD

[3589] Mikhail V. Berlinkov, Robert Ferens, and Marek Szykuła. Preimage problems for deterministic finite automata. Journal of Computer and System Sciences, 115(??):

214–234, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300805.

Charalampopoulos:2021:CPM

[3590] Panagiotis Charalampopoulos, Tomasz Kociumaka, Solon P. Pissis, Jakub Radoszewski, Wojciech Rytter, Juliusz Straszyński, Tomasz Waleń, and Wiktor Zuba. Circular pattern matching with k mismatches. Journal of Computer and System Sciences, 115 (??):73–85, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300702.

Lohrey:2021:LCP

[3591] Markus Lohrey and Sebastian Maneth.

Largest common prefix of a regular tree language. Journal of Computer and System Sciences, 115(??): 235–245, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300830.

Carvalho:2021:FDP

[3592] Iago A. Carvalho, Thomas Erlebach, and Kleitos Papadopoulos. On the fast delivery problem with one or two packages. Journal of Computer and System Sciences, 115(??): 246–263, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300842.

Bollig:2021:CFS

[3593] Benedikt Bollig, Marie Fortin, and Paul Gastin. Communicating finitestate machines, first-order logic, and star-free propositional dynamic logic. Journal of Computer and System Sciences, 115(??):22–53, February 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000020300659.

Anonymous:2021:PMa

[3594] Anonymous. Pages 1–54 (March 2021). Journal of Computer and System Sciences, 116(??):1–54, March 2021. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBb

[3595] Anonymous. Editorial Board. Journal of Computer and System Sciences, 116(??):ii, March 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020301070.

Wang:2021:ECQ

[3596] Qisheng Wang, Junyi Liu, and Mingsheng Ying. Equivalence checking of quantum finite-state machines. Journal of Computer and System Sciences, 116(??):1–21, March 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300829.

Taleb:2021:SVD

[3597] Abdul Rahman Taleb and Damien Vergnaud. Speeding-up verification of digital signatures. Journal of Computer and System Sciences, 116 (??):22–39, March 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300854.

Durand:2021:DCP

[3598] Arnaud Durand, Anselm Haak, Juha Kontinen, and Heribert Vollmer. Descriptive complexity of #P functions: a new perspective. Journal of Computer and System Sciences, 116(??): 40–54, March 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300374.

Anonymous:2021:PMb

[3599] Anonymous. Pages 1–242 (May 2021).

Journal of Computer and System Sciences, 117(??):1–242, May 2021. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBc

[3600] Anonymous. Editorial Board. Journal of Computer and System Sciences, 117(??):ii, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301173.

Belazzougui:2021:BT

[3601] Djamal Belazzougui, Manuel Cáceres, Travis Gagie, Paweł Gawrychowski, Juha KÃrkkÃinen, Gonzalo Navarro, Alberto Ordóñez, Simon J. Puglisi, and Yasuo Tabei. Block trees. Journal of Computer and System Sciences, 117(??):1–22, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020301033.

Lagerkvist:2021:CIC

[3602] Victor Lagerkvist and Biman Roy. Complexity of inverse constraint problems and a dichotomy for the inverse satisfiability problem. Journal of Computer and System Sciences, 117(??):23–39, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300970.

Fomin:2021:PCT

[3603] Fedor V. Fomin, Petr A. Golovach, and Kirill Simonov. Parameterized kclustering: Tractability island. Journal of Computer and System Sciences, 117(??):50–74, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300982.

Li:2021:PCC

[3604] Keqin Li. On the profits of competing cloud service providers: a game theoretic approach. Journal of Computer and System Sciences, 117 (??):130–153, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200002030101X.

Livshits:2021:CSR

[3605] Ester Livshits, Benny Kimelfeld, and Jef Wijsen. Counting subset repairs with functional dependencies. Journal of Computer and System Sciences, 117(??):154–164, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300933.

Jones:2021:MPD

[3606] Mark Jones, Steven Kelk, and Leen Stougie. Maximum parsimony distance on phylogenetic trees: a linear kernel and constant factor approximation algorithm. Journal of Computer and System Sciences, 117 (??):165–181, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300969.

Kawase:2021:AAA

[3607] Yasushi Kawase, Tomomi Matsui, and Atsushi Miyauchi. Additive approximation algorithms for modularity maximization. Journal of Computer and System Sciences, 117(??): 182–201, May 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301124.

Clemente:2021:RRT

[3608] Lorenzo Clemente and Sławomir Lasota. Reachability relations of timed pushdown automata. Journal of Computer and System Sciences, 117 (??):202–241, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301100.

Caragiannis:2021:APN

[3609] Ioannis Caragiannis and Angelo Fanelli. On approximate pure Nash equilibria in weighted congestion games with polynomial latencies. Journal of Computer and System Sciences, 117(??):40–48, May 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020301008.

Flocchini:2021:ICA

[3610] Paola Flocchini. 46th International Colloquium on Automata, Languages and Programming (ICALP 2019) — Track C: Foundations of networks and multi-agent systems. Journal of Computer and System Sciences, 117 (??):49, May 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300957.

Deligkas:2021:CES

[3611] Argyrios Deligkas, John Fearnley, Themistoklis Melissourgos, and Paul G. Spirakis. Computing exact solutions of consensus halving and the Borsuk–Ulam theorem. Journal of Computer and System Sciences, 117 (??):75–98, May 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000020300994.

Grigore:2021:SM

[3612] Radu Grigore and Stefan Kiefer. Selective monitoring. Journal of Computer and System Sciences, 117(??):

99–129, May 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301021.

Anonymous:2021:PJ

[3613] Anonymous. Pages 1–140 (June 2021).

Journal of Computer and System Sciences, 118(??):1–140, June 2021. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBd

[3614] Anonymous. Editorial Board. Journal of Computer and System Sciences, 118(??):ii, June 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000064.

Bjorklund:2021:UPH

[3615] Henrik Björklund, Frank Drewes, Petter Ericson, and Florian Starke. Uniform parsing for hyperedge replacement grammars. Journal of Computer and System Sciences, 118(??): 1–27, June 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020300945.

He:2021:MOC

[3616] Huaiwen He and Hong Shen. Minimizing the operation cost of distributed green data centers with energy storage under carbon capping. Journal of Computer and System Sciences, 118(??):28–52, June 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301112.

Claude:2021:GCI

[3617] Francisco Claude, Gonzalo Navarro, and Alejandro Pacheco. Grammarcompressed indexes with logarithmic search time. Journal of Computer and System Sciences, 118(??): 53–74, June 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301136.

Gasieniec:2021:POB

[3618] Leszek Gasieniec, Jesper Jansson, Christos Levcopoulos, Andrzej Lingas, and Mia Persson. Pushing the online Boolean matrix-vector multiplication conjecture off-line and identifying its easy cases. Journal of Computer and System Sciences, 118 (??):108–118, June 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000027.

Tsur:2021:ROT

[3619] Dekel Tsur. Representation of ordered trees with a given degree distribution. Journal of Computer and System Sciences, 118(??):119– 130, June 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301203.

Li:2021:AAM

[3620] Xingfu Li, Daming Zhu, and Lusheng Wang. A 43-approximation algorithm for the maximum internal spanning tree problem. Journal of Computer and System Sciences, 118(??): 131–140, June 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000015.

Kozachinskiy:2021:AKC

[3621] Alexander Kozachinskiy and Alexander Shen. Automatic Kolmogorov complexity, normality, and finite-state dimension revisited. Journal of Computer and System Sciences, 118(??): 75–107, June 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000020301215.

Anonymous:2021:PA

[3622] Anonymous. Pages 1–230 (August 2021). Journal of Computer and System Sciences, 119(??):1–230, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBe

[3623] Anonymous. Editorial Board. Journal of Computer and System Sciences, 119(??):ii, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000337.

Enright:2021:DER

[3624] Jessica Enright, Kitty Meeks, George B.

Mertzios, and Viktor Zamaraev. Deleting edges to restrict the size of an epidemic in temporal networks. Journal of Computer and System Sciences, 119(??):60–77, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000155.

Daviaud:2021:WEC

[3625] Laure Daviaud, Marcin Jurdziński, Ranko Lazić, Filip Mazowiecki, Guillermo A. Pérez, and James Worrell. When are emptiness and containment decidable for probabilistic automata? Journal of Computer and System Sciences, 119(??): 78–96, August 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000143.

Barcelo:2021:RCF

[3626] Pablo Barceló, Alexander Baumgartner, Victor Dalmau, and Benny Kimelfeld. Regularizing conjunctive features for classification. Journal of Computer and System Sciences, 119 (??):97–124, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000118.

Eiben:2021:BBT

[3627] Eduard Eiben, Daniel Lokshtanov, and Amer E. Mouawad. Bisection of bounded treewidth graphs by convolutions. Journal of Computer and System Sciences, 119(??): 125–132, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000179.

Ferrarotti:2021:DCD

[3628] Flavio Ferrarotti, Senén González, José María Turull Torres, Jan Van den Bussche, and Jonni Virtema. Descriptive complexity of deterministic polylogarithmic time and space. Journal of Computer and System Sciences, 119 (??):145–163, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000180.

Donkers:2021:TKD

[3629] Huib Donkers and Bart M. P. Jansen.

A Turing kernelization dichotomy for structural parameterizations of F - minor-free deletion. Journal of Computer and System Sciences, 119(??): 164–182, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000209.

DeMarco:2021:OCU

[3630] Gianluca De Marco, Tomasz Jurdziński, and Dariusz R. Kowalski. Optimal channel utilization with limited feedback. Journal of Computer and System Sciences, 119(??): 21–33, August 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002100012X.

Dominguez:2021:NSA

[3631] Jesús Domínguez and Maribel Fernández. Nominal syntax with atom substitutions. Journal of Computer and System Sciences, 119(??):34– 59, August 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000106.

Avni:2021:BMG

[3632] Guy Avni, Thomas A. Henzinger, and Dorde Zikelić. Bidding mechanisms in graph games. Journal of Computer and System Sciences, 119 (??):133–144, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000234.

Junges:2021:CRP

[3633] Sebastian Junges, Joost-Pieter Katoen, Guillermo A. Pérez, and Tobias Winkler. The complexity of reachability in parametric Markov decision processes. Journal of Computer and System Sciences, 119(??): 183–210, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000210.

Brihaye:2021:RER

[3634] Thomas Brihaye, Véronique Bruyère, Aline Goeminne, and Nathan Thomasset. On relevant equilibria in reachability games. Journal of Computer and System Sciences, 119(??): 211–230, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000246.

Erlebach:2021:TGE

[3635] Thomas Erlebach, Michael Hoffmann, and Frank Kammer. On temporal graph exploration. Journal of Computer and System Sciences, 119 (??):1–18, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000131.

Schewe:2021:ESI

[3636] Sven Schewe and Lijun Zhang. Editorial — special issue on Concurrency Theory (CONCUR 2018). Journal of Computer and System Sciences, 119 (??):19–20, August 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000167.

Anonymous:2021:PS

[3637] Anonymous. Pages 1–194 (September 2021). Journal of Computer and System Sciences, 120(??):1–194, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBf

[3638] Anonymous. Editorial Board. Journal of Computer and System Sciences, 120(??):ii, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000507.

Ishizuka:2021:CPA

[3639] Takashi Ishizuka. The complexity of the parity argument with potential. Journal of Computer and System Sciences, 120(??):14– 41, September 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000283.

Lima:2021:RGT

[3640] Paloma T. Lima, Vinicius F. dos Santos, Ignasi Sau, and Uéverton S. Souza. Reducing graph transversals via edge contractions. Journal of Computer and System Sciences, 120 (??):62–74, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000271.

Bulteau:2021:YRM

[3641] Laurent Bulteau, Niels Grüttemeier, Christian Komusiewicz, and Manuel Sorge. Your rugby mates don’t need to know your colleagues: Triadic closure with edge colors. Journal of Computer and System Sciences, 120 (??):75–96, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000258.

Mertzios:2021:SWT

[3642] George B. Mertzios, Hendrik Molter, and Viktor Zamaraev. Sliding window temporal graph coloring. Journal of Computer and System Sciences, 120 (??):97–115, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000295.

Bonifaci:2021:AHS

[3643] Vincenzo Bonifaci, Gianlorenzo D’Angelo, and Alberto Marchetti-Spaccamela. Algorithms for hierarchical and semipartitioned parallel scheduling. Journal of Computer and System Sciences, 120(??):116–136, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000021000362.

Chen:2021:LBU

[3644] Wenbin Chen, Jin Li, Zhengan Huang, Chongzhi Gao, Siuming Yiu, and Zoe L. Jiang. Lattice-based unidirectional infinite-use proxy resignatures with private re-signature key. Journal of Computer and System Sciences, 120(??):137–148, September 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000398.

Abascal:2021:CNP

[3645] Jackson Abascal, Lane A. Hemaspaandra, Shir Maimon, and Daniel Rubery. Closure and nonclosure properties of the classes of compressible and rankable sets. Journal of Computer and System Sciences, 120(??): 162–176, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000192.

Schmitt:2021:BBS

[3646] Dominique Schmitt. Bivariate Bsplines from convex configurations. Journal of Computer and System Sciences, 120(??):42–61, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S002200002100026X.

Gasieniec:2021:FSP

[3647] Leszek Gasieniec, Jesper Jansson, and Christos Levcopoulos. Foreword: Selected papers from the 22nd International Symposium on Fundamentals of Computation Theory (FCT 2019). Journal of Computer and System Sciences, 120(??): 177–178, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000404.

Protasov:2021:AMR

[3648] Vladimir Yu. Protasov. Analytic methods for reachability problems. Journal of Computer and System Sciences, 120 (??):1–13, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000222.

Hampson:2021:TST

[3649] Christopher Hampson. On the termination and structural termination problems for counter machines with incrementing errors. Journal of Computer and System Sciences, 120(??): 149–161, September 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000374.

Akrida:2021:TEW

[3650] Eleni C. Akrida, George B. Mertzios, Paul G. Spirakis, and Christoforos Raptopoulos. The temporal explorer who returns to the base. Journal of Computer and System Sciences, 120(??):179–193, September 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000386.

Anonymous:2021:PN

[3651] Anonymous. Pages 1–76 (November 2021). Journal of Computer and System Sciences, 121(??):1–76, November 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBg

[3652] Anonymous. Editorial Board. Journal of Computer and System Sciences, 121(??):ii, November 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000593.

Zheng:2021:LBW

[3653] Yunhai Zheng, Chengliang Tian, Hanlin Zhang, Jia Yu, and Fengjun Li. Lattice-based weak-key analysis on single-server outsourcing protocols of modular exponentiations and basic countermeasures. Journal of Computer and System Sciences, 121(??): 18–33, November 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000441.

Torres-Aviles:2021:USP

[3654] Rodrigo Torres-Avilés. Undecidability of the speed positiveness problem in reversible and complete Turing machines. Journal of Computer and System Sciences, 121(??): 49–56, November 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000465.

Eiben:2021:MWM

[3655] Eduard Eiben, Robert Ganian, Thekla Hamm, and O joung Kwon. Measuring what matters: a hybrid approach to dynamic programming with treewidth. Journal of Computer and System Sciences, 121(??):57– 75, November 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200002100043X.

Casteigts:2021:TCA

[3656] Arnaud Casteigts, Joseph G. Peters, and Jason Schoeters. Temporal cliques admit sparse spanners. Journal of Computer and System Sciences, 121 (??):1–17, November 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000428.

Aspnes:2021:CPP

[3657] James Aspnes. Clocked population protocols. Journal of Computer and System Sciences, 121(??): 34–48, November 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000453.

Anonymous:2021:PD

[3658] Anonymous. Pages 1–126 (December 2021). Journal of Computer and System Sciences, 122(??):1–126, December 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2021:EBh

[3659] Anonymous. Editorial Board. Journal of Computer and System Sciences, 122(??):ii, December 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002100074X.

Bell:2021:IQF

[3660] Paul C. Bell and Mika Hirvensalo. On injectivity of quantum finite automata. Journal of Computer and System Sciences, 122(??): 19–33, December 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000532.

Goldberg:2021:HBP

[3661] Paul W. Goldberg and Alexandros Hollender. The Hairy Ball problem is PPAD-complete. Journal of Computer and System Sciences, 122 (??):34–62, December 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000556.

Dennunzio:2021:ECC

[3662] Alberto Dennunzio, Enrico Formenti, Darij Grinberg, and Luciano Margara. An efficiently computable characterization of stability and instability for linear cellular automata. Journal of Computer and System Sciences, 122 (??):63–71, December 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000635.

Gotoh:2021:EDN

[3663] Tsuyoshi Gotoh, Paola Flocchini, Toshimitsu Masuzawa, and Nicola Santoro. Exploration of dynamic networks: Tight bounds on the number of agents. Journal of Computer and System Sciences, 122(??): 1–18, December 2021. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000416.

Luthra:2021:TTO

[3664] Manisha Luthra, Boris Koldehofe, Niels Danger, Pascal Weisenberger, Guido Salvaneschi, and Ioannis Stavrakakis. TCEP: Transitions in operator placement to adapt to dynamic network environments. Journal of Computer and System Sciences, 122 (??):94–125, December 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000544.

Feldmann:2021:PTA

[3665] Andreas Emil Feldmann and David Saulpic. Polynomial time approximation schemes for clustering in low highway dimension graphs. Journal of Computer and System Sciences, 122 (??):72–93, December 2021. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000647.

Anonymous:2022:PF

[3666] Anonymous. Pages 1–202 (February 2022). Journal of Computer and System Sciences, 123(??):1–202, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EBa

[3667] Anonymous. Editorial Board. Journal of Computer and System Sciences, 123(??):ii, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000957.

Chakraborty:2022:FDP

[3668] Sankardeep Chakraborty, Anish Mukherjee, Venkatesh Raman, and Srinivasa Rao Satti. Frameworks for designing in-place graph algorithms. Journal of Computer and System Sciences, 123(??):1–19, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000684.

Chen:2022:IAS

[3669] Chi-Yeh Chen and Sun-Yuan Hsieh.

An improved algorithm for the Steiner tree problem with bounded edge-length. Journal of Computer and System Sciences, 123(??):20– 36, February 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000672.

Nutov:2022:ACS

[3670] Zeev Nutov. A 4 + ϵ approximation for k-connected subgraphs. Journal of Computer and System Sciences, 123 (??):64–75, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000702.

Golovach:2022:RNP

[3671] Petr A. Golovach, Christian Komusiewicz, Dieter Kratsch, and Van Bang Le. Refined notions of parameterized enumeration kernels with applications to matching cut enumeration. Journal of Computer and System Sciences, 123(??): 76–102, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000696.

Frei:2022:CS

[3672] Fabian Frei, Edith Hemaspaandra, and Jörg Rothe. Complexity of stability. Journal of Computer and System Sciences, 123(??):103– 121, February 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000659.

Jez:2022:WEN

[3673] Artur Jez. Word equations in nondeterministic linear space. Journal of Computer and System Sciences, 123 (??):122–142, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000775.

Marino:2022:CTG

[3674] Andrea Marino and Ana Silva. Coloring temporal graphs. Journal of Computer and System Sciences, 123 (??):171–185, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000805.

Mukhopadhyay:2022:PGC

[3675] Priyanka Mukhopadhyay. The Projection Games Conjecture and the hardness of approximation of super-SAT and related problems. Journal of Computer and System Sciences, 123(??): 186–201, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000830.

Kowalski:2022:PAD

[3676] Dariusz R. Kowalski and Miguel A.

Mosteiro. Polynomial anonymous dynamic distributed computing without a unique leader. Journal of Computer and System Sciences, 123 (??):37–63, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000660.

Bonomi:2022:SIA

[3677] Silvia Bonomi, Giuseppe A. Di Luna, Othon Michail, and Leonardo Querzoni. Special issue on algorithmic theory of dynamic networks and its applications — preface. Journal of Computer and System Sciences, 123(??): 157–158, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000787.

Zhang:2022:IPI

[3678] Hanrui Zhang. Improved prophet inequalities for combinatorial welfare maximization with (approximately) subadditive agents. Journal of Computer and System Sciences, 123(??): 143–156, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000799.

Ren:2022:IDS

[3679] Hanlin Ren. Improved distance sensitivity oracles with subcubic preprocessing time. Journal of Computer and System Sciences, 123(??): 159–170, February 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000817.

Anonymous:2022:PMa

[3680] Anonymous. Pages 1–234 (March 2022). Journal of Computer and System Sciences, 124(??):1–234, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EBb

[3681] Anonymous. Editorial Board. Journal of Computer and System Sciences, 124(??):ii, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021001082.

Drange:2022:TI

[3682] Pål Grønås Drange, Markus Fanebust Dregi, Daniel Lokshtanov, and Blair D. Sullivan. On the threshold of intractability. Journal of Computer and System Sciences, 124 (??):1–25, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000842.

Dublois:2022:AMM

[3683] Louis Dublois, Tesshu Hanaka, Mehdi Khosravian Ghadikolaei, Michael Lampis, and Nikolaos Melissinos. (In)approximability of maximum minimal FVS. Journal of Computer and System Sciences, 124(??):26–40, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000829.

Tao:2022:IJU

[3684] Yufei Tao and Ke Yi. Intersection joins under updates. Journal of Computer and System Sciences, 124 (??):41–64, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000854.

Ambos-Spies:2022:NID

[3685] Klaus Ambos-Spies, Wolfgang Merkle, and Sebastiaan A. Terwijn. Normalized information distance and the oscillation hierarchy. Journal of Computer and System Sciences, 124(??): 65–76, March 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000866.

Melissourgos:2022:EMP

[3686] Themistoklis Melissourgos, Sotiris E. Nikoletseas, Christoforos L. Raptopoulos, and Paul G. Spirakis. An extension of the Moran process using type-specific connection graphs. Journal of Computer and System Sciences, 124(??):77–96, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000878.

Agrawal:2022:FGC

[3687] Akanksha Agrawal. Fine-grained complexity of rainbow coloring and its variants. Journal of Computer and System Sciences, 124(??):140– 158, March 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000982.

Golovach:2022:IDP

[3688] Petr A. Golovach, Daniël Paulusma, and Erik Jan van Leeuwen. Induced Disjoint Paths in AT-free graphs. Journal of Computer and System Sciences, 124(??):170–191, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021001008.

Blasius:2022:EEH

[3689] Thomas Bläsius, Tobias Friedrich, Julius Lischeid, Kitty Meeks, and Martin Schirneck. Efficiently enumerating hitting sets of hypergraphs arising in data profiling. Journal of Computer and System Sciences, 124 (??):192–213, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021000994.

Kuhlmann:2022:TGC

[3690] Marco Kuhlmann, Andreas Maletti, and Lena Katharina Schiffer. The tree-generative capacity of combinatory categorial grammars. Journal of Computer and System Sciences, 124 (??):214–233, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021001021.

Charalampopoulos:2022:SSS

[3691] Panagiotis Charalampopoulos and Adam Karczmarz. Single-source shortest paths and strong connectivity in dynamic planar graphs. Journal of Computer and System Sciences, 124 (??):97–111, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002100091X.

Kowalik:2022:MVT

[3692] Lukasz Kowalik, Shaohua Li, Wojciech Nadara, Marcin Smulewicz, and Magnus Wahlström. Manyvisits TSP revisited. Journal of Computer and System Sciences, 124 (??):112–128, March 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000908.

Eisenbrand:2022:ACT

[3693] Friedrich Eisenbrand and Moritz Venzin. Approximate CVP_{p} in time 2^{0.802n}. Journal of Computer and System Sciences, 124(??):129– 139, March 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021000891.

Chan:2022:MCM

[3694] Timothy M. Chan and Qizheng He. More on change-making and related problems. Journal of Computer and System Sciences, 124(??):159– 169, March 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200002100088X.

Anonymous:2022:PMb

[3695] Anonymous. Pages 1–166 (May 2022). Journal of Computer and System Sciences, 125(??):1–166, May 2022. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EBc

[3696] Anonymous. Editorial Board. Journal of Computer and System Sciences, 125(??):ii, May 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021001203.

Cseresnyes:2022:REL

[3697] Ehud Cseresnyes and Hannes Seiwert. Regular expression length via arithmetic formula complexity. Journal of Computer and System Sciences, 125(??):1–24, May 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002100101X.

Bringmann:2022:GRA

[3698] Karl Bringmann, Ralph Keusch, Johannes Lengler, Yannic Maus, and Anisur R. Molla. Greedy routing and the algorithmic small-world phenomenon. Journal of Computer and System Sciences, 125(??): 59–105, May 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021001112.

Deligkas:2022:AET

[3699] Argyrios Deligkas, John Fearnley, Themistoklis Melissourgos, and Paul G. Spirakis. Approximating the existential theory of the reals. Journal of Computer and System Sciences, 125 (??):106–128, May 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021001045.

DiGiacomo:2022:OPT

[3700] Emilio Di Giacomo, Giuseppe Liotta, and Fabrizio Montecchiani. Orthogonal planarity testing of bounded treewidth graphs. Journal of Computer and System Sciences, 125(??): 129–148, May 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021001124.

Marx:2022:IFE

[3701] Dániel Marx and R. B. Sandeep. Incompressibility of H-free edge modification problems: Towards a dichotomy. Journal of Computer and System Sciences, 125(??):25– 58, May 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021001033.

Fairstein:2022:AOA

[3702] Yaron Fairstein, Ariel Kulik, Joseph (Seffi) Naor, Danny Raz, and Hadas Shachnai. An almost optimal approximation algorithm for monotone submodular multiple knapsack. Journal of Computer and System Sciences, 125(??): 149–165, May 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000021001136.

Anonymous:2022:PJ

[3703] Anonymous. Pages 1–164 (June 2022). Journal of Computer and System Sciences, 126(??):1–164, June 2022. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EBd

[3704] Anonymous. Editorial Board. Journal of Computer and System Sciences, 126(??):ii, June 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000113.

Augustine:2022:LCD

[3705] John Augustine, Seth Gilbert, Fabian Kuhn, Peter Robinson, and Suman Sourav. Latency, capacity, and distributed minimum spanning trees. Journal of Computer and System Sciences, 126(??):1–20, June 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000021001148.

Bentert:2022:LBC

[3706] Matthias Bentert, Klaus Heeger, and Dusan Knop. Length-bounded cuts: Proper interval graphs and structural parameters. Journal of Computer and System Sciences, 126(??): 21–43, June 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021001161.

Jansen:2022:PVD

[3707] Bart M. P. Jansen and Jari J. H.

de Kroon. Preprocessing vertexdeletion problems: Characterizing graph properties by low-rank adjacencies. Journal of Computer and System Sciences, 126(??):59– 79, June 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000021001239.

Beneteau:2022:MMG

[3708] Laurine Bénéteau, Jérémie Chalopin, Victor Chepoi, and Yann Vaxès. Medians in median graphs and their cube complexes in linear time. Journal of Computer and System Sciences, 126(??):80–105, June 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000034.

Bampis:2022:MK

[3709] Evripidis Bampis, Bruno Escoffier, and Alexandre Teiller. Multistage knapsack. Journal of Computer and System Sciences, 126(??):106– 118, June 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000022.

Andola:2022:SSE

[3710] Nitish Andola, Sourabh Prakash, Vijay Kumar Yadav, Raghav, S. Venkatesan, and Shekhar Verma. A secure searchable encryption scheme for cloud using hash-based indexing. Journal of Computer and System Sciences, 126 (??):119–137, June 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000010.

Bridoux:2022:CFP

[3711] Florian Bridoux, Amélia Durbec, Kevin Perrot, and Adrien Richard. Complexity of fixed point counting problems in Boolean networks. Journal of Computer and System Sciences, 126(??):138–164, June 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000046.

Friggstad:2022:CFA

[3712] Zachary Friggstad and Chaitanya Swamy. A constant-factor approximation for directed latency in quasipolynomial time. Journal of Computer and System Sciences, 126(??): 44–58, June 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002100115X.

Anonymous:2022:PA

[3713] Anonymous. Pages 1–146 (August 2022). Journal of Computer and System Sciences, 127(??):1–146, August 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EBe

[3714] Anonymous. Editorial Board. Journal of Computer and System Sciences, 127(??):ii, August 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000241.

Chalopin:2022:USC

[3715] Jérémie Chalopin, Victor Chepoi, Shay Moran, and Manfred K. Warmuth. Unlabeled sample compression schemes and corner peelings for ample and maximum classes. Journal of Computer and System Sciences, 127(??):1–28, August 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002200006X.

Abboud:2022:SLB

[3716] Amir Abboud, Karl Bringmann, Danny Hermelin, and Dvir Shabtay. Scheduling lower bounds via AND subset sum. Journal of Computer and System Sciences, 127(??): 29–40, August 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000058.

Gebhardt:2022:M

[3717] Kilian Gebhardt, Frédéric Meunier, and Sylvain Salvati. O_{n} is an n-MCFL. Journal of Computer and System Sciences, 127(??):41– 52, August 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000174.

Bell:2022:PAP

[3718] Paul C. Bell. Polynomially ambiguous probabilistic automata on restricted languages. Journal of Computer and System Sciences, 127(??): 53–65, August 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000162.

Hemaspaandra:2022:COB

[3719] Edith Hemaspaandra, Lane A. Hemaspaandra, and Jörg Rothe. The complexity of online bribery in sequential elections. Journal of Computer and System Sciences, 127(??): 66–90, August 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000071.

Canini:2022:RSS

[3720] Marco Canini, Iosif Salem, Liron Schiff, Elad M. Schiller, and Stefan Schmid. Renaissance: a self-stabilizing distributed SDN control plane using inband communications. Journal of Computer and System Sciences, 127 (??):91–121, August 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000150.

Nabli:2022:CMC

[3721] Adel Nabli, Margarida Carvalho, and Pierre Hosteins. Complexity of the multilevel critical node problem. Journal of Computer and System Sciences, 127(??):122– 145, August 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000186.

Anonymous:2022:PS

[3722] Anonymous. Pages 1–134 (September 2022). Journal of Computer and System Sciences, 128(??):1–134, September 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EBf

[3723] Anonymous. Editorial Board. Journal of Computer and System Sciences, 128(??):ii, September 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200002200037X.

Linz:2022:NEA

[3724] Simone Linz and Charles Semple. Non-essential arcs in phylogenetic networks. Journal of Computer and System Sciences, 128(??): 1–17, September 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000198.

Cordasco:2022:DDP

[3725] Gennaro Cordasco, Luisa Gargano, and Adele A. Rescigno. Dual domination problems in graphs. Journal of Computer and System Sciences, 128 (??):18–34, September 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000277.

Salo:2022:CC

[3726] Ville Salo. Cutting corners. Journal of Computer and System Sciences, 128 (??):35–70, September 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000204.

Brettell:2022:CWS

[3727] Nick Brettell, Matthew Johnson, and Daniël Paulusma. Computing weighted subset odd cycle transversals in H-free graphs. Journal of Computer and System Sciences, 128 (??):71–85, September 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000289.

Borba:2022:MAE

[3728] Eric Borba, Eduardo Tavares, and Paulo Maciel. A modeling approach for estimating performance and energy consumption of storage systems. Journal of Computer and System Sciences, 128(??): 86–106, September 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000290.

Destombes:2022:RBK

[3729] Julien Destombes and Andrei Romashchenko. Resource-bounded Kolmogorov complexity provides an obstacle to soficness of multidimensional shifts. Journal of Computer and System Sciences, 128(??):107– 134, September 2022. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000307.

Anonymous:2022:PN

[3730] Anonymous. Pages 1–106 (November 2022). Journal of Computer and System Sciences, 129(??):1–106, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EBg

[3731] Anonymous. Editorial Board. Journal of Computer and System Sciences, 129(??):ii, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000460.

Chatterjee:2022:GPE

[3732] Krishnendu Chatterjee and Laurent Doyen. Graph planning with expected finite horizon. Journal of Computer and System Sciences, 129 (??):1–21, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000319.

Saurabh:2022:PCG

[3733] Saket Saurabh, Uéverton dos Santos Souza, and Prafullkumar Tale. On the parameterized complexity of Grid Contraction. Journal of Computer and System Sciences, 129(??): 22–38, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000332.

Ferro:2022:ACE

[3734] Manuel Vilares Ferro, Víctor M. Darriba Bilbao, and Jesús Vilares Ferro. Absolute convergence and error thresholds in non-active adaptive sampling. Journal of Computer and System Sciences, 129(??): 39–61, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002200040X.

Brand:2022:DSC

[3735] Cornelius Brand. Discriminantal subset convolution: Refining exterioralgebraic methods for parameterized algorithms. Journal of Computer and System Sciences, 129(??): 62–71, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000423.

Kadivar:2022:MCB

[3736] Mehdi Kadivar and Neda Mohammadi.

A maximum clique based approximation algorithm for wireless link scheduling under SINR model. Journal of Computer and System Sciences, 129 (??):72–89, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000320.

Jukna:2022:LBB

[3737] Stasys Jukna and Andrzej Lingas.

Lower bounds for Boolean circuits of bounded negation width. Journal of Computer and System Sciences, 129 (??):90–105, November 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000411.

Anonymous:2022:PD

[3738] Anonymous. Pages 1–70 (December 2022). Journal of Computer and System Sciences, 130(??):1–70, December 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2022:EB

[3739] Anonymous. Editorial Board. Journal of Computer and System Sciences, 130(??):ii, December 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000563.

Czyzowicz:2022:CTP

[3740] Jurek Czyzowicz, Leszek Gasieniec, Adrian Kosowski, Evangelos Kranakis, Paul G. Spirakis, and Przemysław Uznański. On convergence and threshold properties of discrete Lotka– Volterra population protocols. Journal of Computer and System Sciences, 130(??):1–25, December 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000502.

Bartal:2022:CMS

[3741] Yair Bartal, Ora Nova Fandina, and Ofer Neiman. Covering metric spaces by few trees. Journal of Computer and System Sciences, 130(??): 26–42, December 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000496.

Deng:2022:AAC

[3742] Shichuan Deng, Jian Li, and Yuval Rabani. Approximation algorithms for clustering with dynamic points. Journal of Computer and System Sciences, 130(??): 43–70, December 2022. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000514.

Anonymous:2023:PF

[3743] Anonymous. Pages 1–104 (February 2023). Journal of Computer and System Sciences, 131(??):1–104, February 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2023:EBa

[3744] Anonymous. Editorial Board. Journal of Computer and System Sciences, 131(??):ii, February 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000678.

Chakraborty:2023:SRN

[3745] Sankardeep Chakraborty, Roberto Grossi, Kunihiko Sadakane, and Srinivasa Rao Satti. Succinct representation for (non)deterministic finite automata. Journal of Computer and System Sciences, 131(??): 1–12, February 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000526.

Wang:2023:QRA

[3746] Qisheng Wang and Mingsheng Ying.

Quantum random access storedprogram machines. Journal of Computer and System Sciences, 131(??): 13–63, February 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000599.

Saurabh:2023:PCM

[3747] Saket Saurabh and Meirav Zehavi. Parameterized complexity of multinode hubs. Journal of Computer and System Sciences, 131(??):64– 85, February 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000605.

deBernardo:2023:FCQ

[3748] Guillermo de Bernardo, Travis Gagie, Susana Ladra, Gonzalo Navarro, and Diego Seco. Faster compressed quadtrees. Journal of Computer and System Sciences, 131(??):86– 104, February 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000629.

Anonymous:2023:PMa

[3749] Anonymous. Pages 1–88 (March 2023). Journal of Computer and System Sciences, 132(??):1–88, March 2023. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2023:EBb

[3750] Anonymous. Editorial Board. Journal of Computer and System Sciences, 132(??):ii, March 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000782.

Govorov:2023:DBD

[3751] Artem Govorov, Jin-Yi Cai, and Martin Dyer. A dichotomy for bounded degree graph homomorphisms with nonnegative weights. Journal of Computer and System Sciences, 132(??): 1–15, March 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000630.

Salo:2023:PUT

[3752] Ville Salo and Ilkka Törmä. A physically universal Turing machine. Journal of Computer and System Sciences, 132(??):16–44, March 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000617.

Koana:2023:CBM

[3753] Tomohiro Koana, Vincent Froese, and Rolf Niedermeier. The complexity of binary matrix completion under diameter constraints. Journal of Computer and System Sciences, 132 (??):45–67, March 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002200071X.

Fomin:2023:BLC

[3754] Fedor V. Fomin, Danil Sagunov, and Kirill Simonov. Building large kcores from sparse graphs. Journal of Computer and System Sciences, 132 (??):68–88, March 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000721.

Anonymous:2023:PMb

[3755] Anonymous. Pages 1–56 (May 2023). Journal of Computer and System Sciences, 133(??):1–56, May 2023. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2023:EBc

[3756] Anonymous. Editorial Board. Journal of Computer and System Sciences, 133(??):ii, May 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000041.

DeMarco:2023:DNA

[3757] Gianluca De Marco, Dariusz R. Kowalski, and Grzegorz Stachowiak. Deterministic non-adaptive contention resolution on a shared channel. Journal of Computer and System Sciences, 133(??):1–22, May 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000022000745.

Einarson:2023:EVC

[3758] Carl Einarson, Gregory Gutin, Bart M. P. Jansen, Diptapriyo Majumdar, and Magnus Wahlström. pedge/vertex-connected vertex cover: Parameterized and approximation algorithms. Journal of Computer and System Sciences, 133(??):23– 40, May 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000824.

Lin:2023:IIL

[3759] Yen-Heng Lin and Sun-Yuan Hsieh.

Improved IP lookup technology for trie-based data structures. Journal of Computer and System Sciences, 133(??):41–55, May 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000733.

Anonymous:2023:PJ

[3760] Anonymous. Pages 1–86 (June 2023).

Journal of Computer and System Sciences, 134(??):1–86, June 2023. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2023:EBd

[3761] Anonymous. Editorial Board. Journal of Computer and System Sciences, 134(??):ii, June 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002300020X.

Eiben:2023:PCC

[3762] Eduard Eiben, Robert Ganian, Iyad Kanj, Sebastian Ordyniak, and Stefan Szeider. On the parameterized complexity of clustering problems for incomplete data. Journal of Computer and System Sciences, 134(??):1–19, June 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000022000836.

Esnay:2023:ACL

[3763] Solène J. Esnay, Alonso Núñez, and Ilkka Törmä. Arithmetical complexity of the language of generic limit sets of cellular automata. Journal of Computer and System Sciences, 134(??):20–41, June 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000090.

Berndt:2023:OCL

[3764] Sebastian Berndt, Leah Epstein, Klaus Jansen, Asaf Levin, Marten Maack, and Lars Rohwedder. Online bin covering with limited migration. Journal of Computer and System Sciences, 134(??):42–72, June 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000089.

Haviv:2023:BBR

[3765] Ishay Haviv and Michal Parnas. On the binary and Boolean rank of regular matrices. Journal of Computer and System Sciences, 134(??): 73–86, June 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000120.

Anonymous:2023:PA

[3766] Anonymous. Pages 1–174 (August 2023). Journal of Computer and System Sciences, 135(??):1–174, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2023:EBe

[3767] Anonymous. Editorial Board. Journal of Computer and System Sciences, 135(??):ii, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002300034X.

Dey:2023:EAM

[3768] Palash Dey and Neeldhara Misra.

On the exact amount of missing information that makes finding possible winners hard. Journal of Computer and System Sciences, 135(??): 32–54, August 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000168.

Gupta:2023:BSD

[3769] Diksha Gupta, Jared Saia, and Maxwell Young. Bankrupting Sybil despite churn. Journal of Computer and System Sciences, 135(??): 89–124, August 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000235.

Liotta:2023:PCG

[3770] Giuseppe Liotta, Ignaz Rutter, and Alessandra Tappini. Parameterized complexity of graph planarity with restricted cyclic orders. Journal of Computer and System Sciences, 135 (??):125–144, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000284.

Fernandes:2023:HHI

[3771] Cristina G. Fernandes and Carla N. Lintzmayer. How heavy independent sets help to find arborescences with many leaves in DAGs. Journal of Computer and System Sciences, 135 (??):158–174, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000260.

Maack:2023:FST

[3772] Nicolas Maack, Hendrik Molter, Rolf Niedermeier, and Malte Renken. On finding separators in temporal split and permutation graphs. Journal of Computer and System Sciences, 135 (??):1–14, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000119.

Cai:2023:BRC

[3773] Jin-Yi Cai, Austen Z. Fan, and Yin Liu. Bipartite 3-regular counting problems with mixed signs. Journal of Computer and System Sciences, 135 (??):15–31, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000144.

Lohrey:2023:CWP

[3774] Markus Lohrey. Complexity of word problems for HNN-extensions. Journal of Computer and System Sciences, 135 (??):145–157, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000272.

Andrews:2023:EET

[3775] John Andrews, Michele Ciampi, and Vassilis Zikas. Etherless Ethereum tokens: Simulating native tokens in Ethereum. Journal of Computer and System Sciences, 135(??): 55–72, August 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000132.

Erlebach:2023:PTE

[3776] Thomas Erlebach and Jakob T.

Spooner. Parameterised temporal exploration problems. Journal of Computer and System Sciences, 135 (??):73–88, August 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000107.

Anonymous:2023:PS

[3777] Anonymous. Pages 1–326 (September 2023). Journal of Computer and System Sciences, 136(??):1–326, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2023:EBf

[3778] Anonymous. Editorial Board. Journal of Computer and System Sciences, 136(??):ii, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000521.

Gupta:2023:GRC

[3779] Siddharth Gupta, Guy Sa’ar, and Meirav Zehavi. Grid recognition: Classical and parameterized computational perspectives. Journal of Computer and System Sciences, 136(??): 17–62, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000259.

Lima:2023:RVC

[3780] Paloma T. Lima, Vinicius F. dos Santos, Ignasi Sau, Uéverton S. Souza, and Prafullkumar Tale. Reducing the vertex cover number via edge contractions. Journal of Computer and System Sciences, 136(??): 63–87, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000302.

Yamakami:2023:BFS

[3781] Tomoyuki Yamakami. The 2CNF Boolean formula satisfiability problem and the linear space hypothesis. Journal of Computer and System Sciences, 136(??):88–112, September 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000247.

Baier:2023:MCU

[3782] Christel Baier, Stefan Kiefer, Joachim Klein, David Müller, and James Worrell. Markov chains and unambiguous automata. Journal of Computer and System Sciences, 136(??):

113–134, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000442.

Jain:2023:AMA

[3783] Sanjay Jain, Xiaodong Jia, Ammar Fathin Sabili, and Frank Stephan. Addition machines, automatic functions and open problems of Floyd and Knuth. Journal of Computer and System Sciences, 136(??):135– 156, September 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000387.

Fomin:2023:PCC

[3784] Fedor V. Fomin, Petr A. Golovach, and Nidhi Purohit. Parameterized complexity of categorical clustering with size constraints. Journal of Computer and System Sciences, 136(??): 171–194, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000430.

Hannula:2023:CEI

[3785] Miika Hannula, Xinyi Li, and Sebastian Link. Controlling entity integrity with key sets. Journal of Computer and System Sciences, 136(??): 195–219, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000454.

Bartier:2023:GTS

[3786] Valentin Bartier, Nicolas Bousquet, and Amer E. Mouawad. Galactic token sliding. Journal of Computer and System Sciences, 136(??): 220–248, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000375.

Aracena:2023:SBN

[3787] Julio Aracena, Adrien Richard, and Lilian Salinas. Synchronizing Boolean networks asynchronously. Journal of Computer and System Sciences, 136(??):249–279, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000023000399.

Behrendt:2023:SAG

[3788] Lukas Behrendt, Katrin Casel, Tobias Friedrich, J. A. Gregor Lagodzinski, Alexander Löser, and Marcus Wilhelm. From symmetry to asymmetry: Generalizing TSP approximations by parametrization. Journal of Computer and System Sciences, 136(??): 157–170, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000429.

Jacob:2023:DSGa

[3789] Ashwin Jacob, Diptapriyo Majumdar, and Venkatesh Raman. Deletion to scattered graph classes II — improved FPT algorithms for deletion to pairs of graph classes. Journal of Computer and System Sciences, 136(??):

280–301, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000417.

Adsul:2023:ACB

[3790] Bharat Adsul, Saptarshi Sarkar, and A. V. Sreejith. Algebraic characterizations and block product decompositions for first order logic and its infinitary quantifier extensions over countable words. Journal of Computer and System Sciences, 136(??): 302–326, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000405.

Caballero:2023:CVS

[3791] David Caballero, Timothy Gomez, Robert Schweller, and Tim Wylie. Complexity of verification in selfassembly with prebuilt assemblies. Journal of Computer and System Sciences, 136(??):1–16, September 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000023000296.

Anonymous:2023:PN

[3792] Anonymous. Pages 1–86 (November 2023). Journal of Computer and System Sciences, 137(??):1–86, November 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Anonymous:2023:EBg

[3793] Anonymous. Editorial Board. Journal of Computer and System Sciences, 137(??):ii, November 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000648.

Mertzios:2023:CMM

[3794] George B. Mertzios, Hendrik Molter, Rolf Niedermeier, Viktor Zamaraev, and Philipp Zschoche. Computing maximum matchings in temporal graphs. Journal of Computer and System Sciences, 137(??): 1–19, November 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000466.

Papamakarios:2023:SCC

[3795] Theodoros Papamakarios and Alexander Razborov. Space characterizations of complexity measures and size-space trade-offs in propositional proof systems. Journal of Computer and System Sciences, 137(??): 20–36, November 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000478.

Bishnu:2023:AOQ

[3796] Arijit Bishnu, Arijit Ghosh, Sudeshna Kolay, Gopinath Mishra, and Saket Saurabh. Almost optimal query algorithm for hitting set using a subset query. Journal of Computer and System Sciences, 137(??):50– 65, November 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000156.

Fomin:2023:DDG

[3797] Fedor V. Fomin, Petr A. Golovach, William Lochet, Danil Sagunov, Saket Saurabh, and Kirill Simonov. Detours in directed graphs. Journal of Computer and System Sciences, 137 (??):66–86, November 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000569.

Falk:2023:LTP

[3798] Brett Hemenway Falk, Rohit Nema, and Rafail Ostrovsky. Linear-time 2party secure merge from additively homomorphic encryption. Journal of Computer and System Sciences, 137 (??):37–49, November 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002300048X.

Angelopoulos:2023:WOS

[3799] Spyros Angelopoulos and Konstantinos Panagiotou. Weighted online search. Journal of Computer and System Sciences, 138(??): ??, December 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000557.

Bridoux:2023:IGI

[3800] Florian Bridoux, Kévin Perrot, Aymeric Picard Marchetto, and Adrien Richard. Interaction graphs of isomorphic automata networks I: Complete digraph and minimum in-degree. Journal of Computer and System Sciences, 138 (??):??, December 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000570.

Greenhill:2023:BAH

[3801] Catherine Greenhill, Bernard Mans, and Ali Pourmiri. Balanced allocation on hypergraphs. Journal of Computer and System Sciences, 138 (??):??, December 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000582.

Jacob:2023:DSGb

[3802] Ashwin Jacob, Jari J. H. de Kroon, Diptapriyo Majumdar, and Venkatesh Raman. Deletion to scattered graph classes I — case of finite number of graph classes. Journal of Computer and System Sciences, 138(??): ??, December 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000600.

Donkers:2023:FST

[3803] Huib Donkers, Bart M. P. Jansen, and Jari J. H. de Kroon. Finding k-secluded trees faster. Journal of Computer and System Sciences, 138 (??):??, December 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000594.

Alaniz:2023:BSO

[3804] Robert M. Alaniz, David Caballero, Sonya C. Cirlos, Timothy Gomez, Elise Grizzell, Andrew Rodriguez, Robert Schweller, Armando Tenorio, and Tim Wylie. Building squares with optimal state complexity in restricted active self-assembly. Journal of Computer and System Sciences, 138(??):??, December 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000673.

Hradovich:2023:RMA

[3805] Elijah Hradovich, Marek Klonowski, and Dariusz R. Kowalski. Restrained medium access control on adversarial shared channels. Journal of Computer and System Sciences, 138(??): ??, December 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000685.

Baumeister:2023:PWU

[3806] Dorothea Baumeister, Marc Neveling, Magnus Roos, Jörg Rothe, Lena Schend, Robin Weishaupt, and Lirong Xia. The possible winner with uncertain weights problem. Journal of Computer and System Sciences, 138 (??):??, December 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000697.

Anonymous:2023:EBh

[3807] Anonymous. Editorial Board. Journal of Computer and System Sciences, 138(??):??, December 2023. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000764.

Anonymous:2023:D

[3808] Anonymous. December 2023. Journal of Computer and System Sciences, 138 (??):??, December 2023. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Binucci:2024:CSP

[3809] Carla Binucci, Emilio Di Giacomo, William J. Lenhart, Giuseppe Liotta, Fabrizio Montecchiani, Martin Nöllenburg, and Antonios Symvonis. On the complexity of the storyplan problem. Journal of Computer and System Sciences, 139(??): ??, February 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000715.

Ibiapina:2024:MGC

[3810] Allen Ibiapina and Ana Silva. Mengerian graphs: Characterization and recognition. Journal of Computer and System Sciences, 139(??): ??, February 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000727.

Sorbelli:2024:WIS

[3811] Francesco Betti Sorbelli, Alfredo Navarra, Lorenzo Palazzetti, Cristina M. Pinotti, and Giuseppe Prencipe.

Wireless IoT sensors data collection reward maximization by leveraging multiple energy- and storageconstrained UAVs. Journal of Computer and System Sciences, 139(??): ??, February 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000806.

Gasieniec:2024:PMM

[3812] Leszek Gasieniec, Tomasz Jurdziński, Ralf Klasing, Christos Levcopoulos, Andrzej Lingas, Jie Min, and Tomasz Radzik. Perpetual maintenance of machines with different urgency requirements. Journal of Computer and System Sciences, 139 (??):??, February 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000818.

Chodil:2024:SPQ

[3813] Miroslav Chodil and Antonín Kucera.

The satisfiability problem for a quantitative fragment of PCTL. Journal of Computer and System Sciences, 139(??):??, February 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000831.

vanBevern:2024:SPK

[3814] René van Bevern, Artem M. Kirilin, Daniel A. Skachkov, Pavel V. Smirnov, and Oxana Yu. Tsidulko. Serial and parallel kernelization of Multiple Hitting Set parameterized by the Dilworth number, implemented on the GPU. Journal of Computer and System Sciences, 139(??): ??, February 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000843.

Anonymous:2024:EBa

[3815] Anonymous. Editorial Board. Journal of Computer and System Sciences, 139(??):??, February 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000909.

Anonymous:2024:F

[3816] Anonymous. February 2024. Journal of Computer and System Sciences, 139 (??):??, February 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Bartlett:2024:CPL

[3817] Peter L. Bartlett and Philip M. Long. Corrigendum to “Prediction, learning, uniform convergence, and scalesensitive dimensions” [J. Comput. Syst. Sci. 56(2) (1998) 174–190]. Journal of Computer and System Sciences, 140(??):??, March 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000703.

Deen:2024:NLK

[3818] Elise Deen, Leo van Iersel, Remie Janssen, Mark Jones, Yukihiro Murakami, and Norbert Zeh. A nearlinear kernel for bounded-state parsimony distance. Journal of Computer and System Sciences, 140(??): ??, March 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002300082X.

Huber:2024:OUP

[3819] Katharina T. Huber, Leo van Iersel, Remie Janssen, Mark Jones, Vincent Moulton, Yukihiro Murakami, and Charles Semple. Orienting undirected phylogenetic networks. Journal of Computer and System Sciences, 140(??):??, March 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000855.

Czerner:2024:FSP

[3820] Philipp Czerner, Roland Guttenberg, Martin Helfrich, and Javier Esparza. Fast and succinct population protocols for Presburger arithmetic. Journal of Computer and System Sciences, 140(??):??, March 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000867.

Huang:2024:BIM

[3821] Haoqiang Huang, Zihe Wang, Zhide Wei, and Jie Zhang. Bounded incentives in manipulating the probabilistic serial rule. Journal of Computer and System Sciences, 140 (??):??, March 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002300096X.

Li:2024:PMA

[3822] Keqin Li. Performance modeling and analysis for randomly walking mobile users with Markov chains. Journal of Computer and System Sciences, 140(??):??, March 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023000971.

Bonomo-Braberman:2024:SPG

[3823] Flavia Bonomo-Braberman, Nick Brettell, Andrea Munaro, and Daniël Paulusma. Solving problems on generalized convex graphs via mimwidth. Journal of Computer and System Sciences, 140(??):??, March 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000023000983.

Anonymous:2024:EBb

[3824] Anonymous. Editorial Board. Journal of Computer and System Sciences, 140(??):??, March 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023001058.

Anonymous:2024:Ma

[3825] Anonymous. March 2024. Journal of Computer and System Sciences, 140 (??):??, March 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Zheng:2024:LQC

[3826] Shenggen Zheng, Yaqiao Li, Minghua Pan, Jozef Gruska, and Lvzhou Li.

Lifting query complexity to timespace complexity for two-way finite automata. Journal of Computer and System Sciences, 141 (??):??, May 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023000995.

Bilo:2024:BTT

[3827] Davide Bilò, Gianlorenzo D’Angelo, Luciano Gualà, Stefano Leucci, and Mirko Rossi. Blackout-tolerant temporal spanners. Journal of Computer and System Sciences, 141 (??):??, May 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023001009.

Su:2024:PLH

[3828] Jinyan Su, Jinhui Xu, and Di Wang.

PAC learning halfspaces in noninteractive local differential privacy model with public unlabeled data. Journal of Computer and System Sciences, 141(??):??, May 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000023001010.

Aspnes:2024:SIS

[3829] James Aspnes and Othon Michail.

Special issue on the 1st Symposium on Algorithmic Foundations of Dynamic Networks (SAND 2022). Journal of Computer and System Sciences, 141(??):??, May 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000023001083.

Anonymous:2024:EBc

[3830] Anonymous. Editorial Board. Journal of Computer and System Sciences, 141(??):??, May 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000059.

Anonymous:2024:Mb

[3831] Anonymous. May 2024. Journal of Computer and System Sciences, 141 (??):??, May 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Bandyapadhyay:2024:CFC

[3832] Sayan Bandyapadhyay, Fedor V. Fomin, and Kirill Simonov. On coresets for fair clustering in metric and Euclidean spaces and their applications. Journal of Computer and System Sciences, 142(??):??, June 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000011.

Poindron:2024:MCP

[3833] Alexis Poindron. The maximal coordination principle in regulatory Boolean networks. Journal of Computer and System Sciences, 142 (??):??, June 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000138.

Kelk:2024:DKT

[3834] Steven Kelk, Simone Linz, and Ruben Meuwese. Deep kernelization for the Tree Bisection and Reconnection (TBR) distance in phylogenetics. Journal of Computer and System Sciences, 142(??):??, June 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002400014X.

Dreier:2024:SBD

[3835] Jan Dreier, Sebastian Ordyniak, and Stefan Szeider. SAT backdoors: Depth beats size. Journal of Computer and System Sciences, 142(??):??, June 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000151.

Anonymous:2024:EBd

[3836] Anonymous. Editorial Board. Journal of Computer and System Sciences, 142(??):??, June 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000199.

Anonymous:2024:J

[3837] Anonymous. June 2024. Journal of Computer and System Sciences, 142 (??):??, June 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Yang:2024:MEG

[3838] Chenxu Yang, Gang Yang, Sun-Yuan Hsieh, Yaping Mao, and Ralf Klasing.

Monitoring the edges of a graph using distances with given girth. Journal of Computer and System Sciences, 143(??):??, August 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000230.

Beal:2024:DPS

[3839] Marie-Pierre Béal, Dominique Perrin, and Antonio Restivo. Decidable problems in substitution shifts. Journal of Computer and System Sciences, 143(??):??, August 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000242.

Babu:2024:PAD

[3840] Jasine Babu, Ajay Saju Jacob, R. Krithika, and Deepak Rajendraprasad. Packing arc-disjoint cycles in oriented graphs. Journal of Computer and System Sciences, 143(??):??, August 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000254.

Anonymous:2024:EBe

[3841] Anonymous. Editorial Board. Journal of Computer and System Sciences, 143(??):??, August 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000333.

Anonymous:2024:A

[3842] Anonymous. August 2024. Journal of Computer and System Sciences, 143 (??):??, August 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Cerveny:2024:KPV

[3843] Radovan Cervený, Pratibha Choudhary, and Ondrej Suchý. On kernels for d-path vertex cover. Journal of Computer and System Sciences, 144(??):??, September 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000266.

Donkers:2024:PRS

[3844] Huib Donkers and Bart M. P. Jansen.

Preprocessing to reduce the search space: Antler structures for feedback vertex set. Journal of Computer and System Sciences, 144(??): ??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000278.

Hermelin:2024:PCI

[3845] Danny Hermelin, Yuval Itzhaki, Hendrik Molter, and Dvir Shabtay. On the parameterized complexity of interval scheduling with eligible machine sets. Journal of Computer and System Sciences, 144(??): ??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002400028X.

Erlebach:2024:CRG

[3846] Thomas Erlebach, Nils Morawietz, Jakob T. Spooner, and Petra Wolf. A cop and robber game on edgeperiodic temporal graphs. Journal of Computer and System Sciences, 144 (??):??, September 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000291.

Chepoi:2024:LSC

[3847] Victor Chepoi, Kolja Knauer, and Manon Philibert. Labeled sample compression schemes for complexes of oriented matroids. Journal of Computer and System Sciences, 144 (??):??, September 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000382.

Samadi:2024:DDF

[3848] Mahboubeh Samadi, Fatemeh Ghassemi, and Ramtin Khosravi. Decentralized deadlock-free enforcement of message orderings in messagebased systems. Journal of Computer and System Sciences, 144(??): ??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000394.

Angelopoulos:2024:OCU

[3849] Spyros Angelopoulos, Christoph Dürr, Shendan Jin, Shahin Kamali, and Marc Renault. Online computation with untrusted advice. Journal of Computer and System Sciences, 144 (??):??, September 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000400.

Bulteau:2024:TCN

[3850] Laurent Bulteau and Louxin Zhang.

The tree-child network inference problem for line trees and the shortest common supersequence problem for permutation strings. Journal of Computer and System Sciences, 144 (??):??, September 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000412.

Ordyniak:2024:BD

[3851] Sebastian Ordyniak, Andre Schidler, and Stefan Szeider. Backdoor DNFs. Journal of Computer and System Sciences, 144(??): ??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000424.

Costa:2024:CLT

[3852] Isnard Lopes Costa, Raul Lopes, Andrea Marino, and Ana Silva. On computing large temporal (unilateral) connected components. Journal of Computer and System Sciences, 144 (??):??, September 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000436.

Molter:2024:TRM

[3853] Hendrik Molter, Malte Renken, and Philipp Zschoche. Temporal reachability minimization: Delaying vs. deleting. Journal of Computer and System Sciences, 144(??):??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000448.

Diaz:2024:MVB

[3854] Josep Díaz, Öznur Yasar Diner, Maria Serna, and Oriol Serra. On minimum vertex bisection of random dregular graphs. Journal of Computer and System Sciences, 144(??): ??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200002400045X.

Bocci:2024:NTI

[3855] Cristiano Bocci, Chiara Capresi, Kitty Meeks, and John Sylvester. A new temporal interpretation of cluster editing. Journal of Computer and System Sciences, 144(??): ??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000461.

Anonymous:2024:EBf

[3856] Anonymous. Editorial Board. Journal of Computer and System Sciences, 144(??):??, September 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000539.

Anonymous:2024:S

[3857] Anonymous. September 2024. Journal of Computer and System Sciences, 144 (??):??, September 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Wang:2024:ECG

[3858] Zhao Wang, Yaping Mao, Sun-Yuan Hsieh, Ralf Klasing, and Yuzhi Xiao. The g-extra connectivity of graph products. Journal of Computer and System Sciences, 145(??): ??, November 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000473.

Dawar:2024:CHP

[3859] Anuj Dawar. Corrigendum to “Homomorphism preservation on quasiwide classes” [J. Comput. Syst. Sci. 76 (5) (2010) 324-332]. Journal of Computer and System Sciences, 145 (??):??, November 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000485. See

[2507].

Anonymous:2024:EBg

[3860] Anonymous. Editorial Board. Journal of Computer and System Sciences, 145(??):??, November 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000631.

Anonymous:2024:N

[3861] Anonymous. November 2024. Journal of Computer and System Sciences, 145 (??):??, November 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Bjorklund:2024:ISM

[3862] Johanna Björklund. The impact of state merging on predictive accuracy in probabilistic tree automata: Dietze’s conjecture revisited. Journal of Computer and System Sciences, 146 (??):??, December 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000588.

Klobas:2024:CCO

[3863] Nina Klobas, George B. Mertzios, Hendrik Molter, and Paul G. Spirakis. The complexity of computing optimum labelings for temporal connectivity. Journal of Computer and System Sciences, 146(??):??, December 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S002200002400059X.

Gomes:2024:MSR

[3864] Guilherme C. M. Gomes, Clément Legrand-Duchesne, Reem Mahmoud, Amer E. Mouawad, Yoshio Okamoto, Vinicius F. dos Santos, and Tom C. van der Zanden. Minimum separator reconfiguration. Journal of Computer and System Sciences, 146 (??):??, December 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000692.

Durand:2024:MSB

[3865] Arnaud Durand, Juha Kontinen, and Jouko Väänänen. Modular SATbased techniques for reasoning tasks in team semantics. Journal of Computer and System Sciences, 146(??): ??, December 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000709.

Nourine:2024:TDC

[3866] Lhouari Nourine, Jean-Marc Petit, and Simon Vilmin. Towards declarative comparabilities: Application to functional dependencies. Journal of Computer and System Sciences, 146 (??):??, December 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000710.

Anonymous:2024:EBh

[3867] Anonymous. Editorial Board. Journal of Computer and System Sciences, 146(??):??, December 2024. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000795.

Anonymous:2024:D

[3868] Anonymous. December 2024. Journal of Computer and System Sciences, 146 (??):??, December 2024. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Chakraborty:2025:LAR

[3869] Dipayan Chakraborty, Soumen Nandi, Sagnik Sen, and D. K. Supraja. A linear algorithm for radio kcoloring of powers of paths having small diameters. Journal of Computer and System Sciences, 147(??): ??, February 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000722.

Agrawal:2025:FSR

[3870] Akanksha Agrawal, Soumita Hait, and Amer E. Mouawad. On finding short reconfiguration sequences between independent sets. Journal of Computer and System Sciences, 147(??):??, February 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000734.

Boehmer:2025:ASM

[3871] Niclas Boehmer and Klaus Heeger. Adapting stable matchings to forced and forbidden pairs. Journal of Computer and System Sciences, 147 (??):??, February 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000746.

Benerecetti:2025:PPP

[3872] Massimo Benerecetti, Daniele Dell’Erba, Fabio Mogavero, Sven Schewe, and Dominik Wojtczak. Priority promotion with Parysian flair. Journal of Computer and System Sciences, 147(??):??, February 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000758.

Mertzios:2025:CGG

[3873] George Mertzios, Othon Michail, George Skretas, Paul G. Spirakis, and Michail Theofilatos. The complexity of growing a graph. Journal of Computer and System Sciences, 147(??):??, February 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000825.

Anonymous:2025:EBa

[3874] Anonymous. Editorial Board. Journal of Computer and System Sciences, 147(??):??, February 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000874.

Anonymous:2025:F

[3875] Anonymous. February 2025. Journal of Computer and System Sciences, 147 (??):??, February 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Bar-Noy:2025:ARO

[3876] Amotz Bar-Noy, Toni Böhnlein, David Peleg, Yingli Ran, and Dror Rawitz. Approximate realizations for outerplanaric degree sequences. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000837.

Bubboloni:2025:COT

[3877] Daniela Bubboloni, Costanza Catalano, Andrea Marino, and Ana Silva. On computing optimal temporal branchings and spanning subgraphs. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0022000024000916.

Jansen:2025:SEF

[3878] Bart M. P. Jansen, Jari J. H. de Kroon, and Michał Włodarczyk. Single-exponential FPT algorithms for enumerating secluded F-free subgraphs and deleting to scattered graph classes. Journal of Computer and System Sciences, 148(??): ??, March 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000928.

Rauch:2025:EPA

[3879] Johannes Rauch, Dieter Rautenbach, and Uéverton S. Souza. Exact and parameterized algorithms for the independent cutset problem. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002400093X.

Chaudhary:2025:PRA

[3880] Juhi Chaudhary and Meirav Zehavi. Parameterized results on acyclic matchings with implications for related problems. Journal of Computer and System Sciences, 148(??): ??, March 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000941.

Nederlof:2025:ATK

[3881] Jesper Nederlof and Krisztina Szilágyi.

Algorithms and Turing kernels for detecting and counting small patterns in unit disk graphs. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000953.

Kocher:2025:BRC

[3882] Chris Köcher and Dietrich Kuske.

Backwards-reachability for cooperating multi-pushdown systems. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024000965.

Li:2025:MEP

[3883] Wen Li, Ralf Klasing, Yaping Mao, and Bo Ning. Monitoring the edges of product networks using distances. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000977.

Tang:2025:EHT

[3884] Zhiyi Tang. Embedding hypercubes into torus and Cartesian product of paths and/or cycles for minimizing wirelength. Journal of Computer and System Sciences, 148(??): ??, March 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024000989.

Feldmann:2025:PCS

[3885] Andreas Emil Feldmann, Anish Mukherjee, and Erik Jan van Leeuwen. The parameterized complexity of the survivable network design problem. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0022000024000990.

Karakostas:2025:TSS

[3886] George Karakostas and Stavros G. Kolliopoulos. Time-sharing scheduling with tolerance capacities. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000024001004.

Anonymous:2025:EBb

[3887] Anonymous. Editorial Board. Journal of Computer and System Sciences, 148(??):??, March 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024001053.

Anonymous:2025:Ma

[3888] Anonymous. March 2025. Journal of Computer and System Sciences, 148 (??):??, March 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Dragan:2025:FDA

[3889] Feodor F. Dragan, Guillaume Ducoffe, and Heather M. Guarnera. Fast deterministic algorithms for computing all eccentricities in (hyperbolic) Helly graphs. Journal of Computer and System Sciences, 149 (??):??, May 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024001016.

Jaakkola:2025:RDC

[3890] Reijo Jaakkola, Antti Kuusisto, and Miikka Vilander. Relating description complexity to entropy. Journal of Computer and System Sciences, 149(??):??, May 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024001107.

Nasz:2025:WHP

[3891] Andreea-Teodora Nász. The weighted HOM-problem over fields. Journal of Computer and System Sciences, 149(??):??, May 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024001119.

Kim:2025:CFA

[3892] Eun Jung Kim, Euiwoong Lee, and Dimitrios M. Thilikos. A constantfactor approximation for weighted bond cover. Journal of Computer and System Sciences, 149 (??):??, May 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024001120.

Anonymous:2025:EBc

[3893] Anonymous. Editorial Board. Journal of Computer and System Sciences, 149(??):??, May 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000066.

Anonymous:2025:Mb

[3894] Anonymous. May 2025. Journal of Computer and System Sciences, 149 (??):??, May 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Connor:2025:TMR

[3895] Matthew Connor and Othon Michail. Transformation of modular robots by rotation: 3 + 1 musketeers for all orthogonally convex shapes. Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000024001132.

Bazgan:2025:DGP

[3896] Cristina Bazgan, Katrin Casel, and Pierre Cazals. Dense graph partitioning on sparse and dense graphs. Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000017.

Csikos:2025:FPT

[3897] Mónika Csikós, Michel Habib, Minh-Hang Nguyen, Mikaël Rabie, and Laurent Viennot. Forbidden patterns in temporal graphs resulting from encounters in a corridor. Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000029.

Mertzios:2025:CTO

[3898] George B. Mertzios, Hendrik Molter, Malte Renken, Paul G. Spirakis, and Philipp Zschoche. The complexity of transitively orienting temporal graphs. Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000121.

Goyal:2025:PCD

[3899] Dishant Goyal, Ashwin Jacob, Kaushtubh Kumar, Diptapriyo Majumdar, and Venkatesh Raman. Parameterized complexity of dominating set variants in almost cluster and split graphs.

Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000133.

Faliszewski:2025:HST

[3900] Piotr Faliszewski, Piotr Skowron, Arkadii Slinko, Krzysztof Sornat, Stanisław Szufa, and Nimrod Talmon. How similar are two elections? Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000145.

Banik:2025:DCC

[3901] Aritra Banik, Prahlad Narasimhan Kasthurirangan, and Venkatesh Raman. Dominator coloring and CD coloring in almost cluster graphs. Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000157.

Anonymous:2025:EBd

[3902] Anonymous. Editorial Board. Journal of Computer and System Sciences, 150(??):??, June 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000248.

Czerwinski:2025:LGF

[3903] Wojciech Czerwiński, Maciej Debski, Tomasz Gogasz, Gordon Hoi, Sanjay Jain, Michał Skrzypczak, Frank Stephan, and Christopher Tan. Languages given by finite automata over the unary alphabet. Journal of Computer and System Sciences, 151 (??):??, August 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000169.

Bazgan:2025:DDS

[3904] Cristina Bazgan, André Nichterlein, and Sofia Vazquez Alferez. Destroying densest subgraphs is hard. Journal of Computer and System Sciences, 151(??):??, August 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000170.

Cunha:2025:ITC

[3905] Luís Cunha, Gabriel Duarte, Fábio Protti, Loana Nogueira, and Uéverton Souza. Induced tree covering and the generalized Yutsis property. Journal of Computer and System Sciences, 151(??):??, August 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000182.

Anonymous:2025:EBe

[3906] Anonymous. Editorial Board. Journal of Computer and System Sciences, 151(??):??, August 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000303.

Anonymous:2025:A

[3907] Anonymous. August 2025. Journal of Computer and System Sciences, 151 (??):??, August 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Tada:2025:LDA

[3908] Takumi Tada and Kazuya Haraguchi. A linear delay algorithm in SD set system and its application to subgraph enumeration. Journal of Computer and System Sciences, 152(??): ??, September 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000194.

Goldberg:2025:IBP

[3909] Leslie Ann Goldberg and John Lapinskas. Instability of backoff protocols with arbitrary arrival rates. Journal of Computer and System Sciences, 152(??):??, September 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000200.

Frohn:2025:RSD

[3910] Martin Frohn, Niels Holtgrefe, Leo van Iersel, Mark Jones, and Steven Kelk. Reconstructing semi-directed level-1 networks using few quarnets. Journal of Computer and System Sciences, 152(??):??, September 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000376.

Kshemkalyani:2025:NOD

[3911] Ajay D. Kshemkalyani and Gokarna Sharma. Near-optimal dispersion on arbitrary anonymous graphs. Journal of Computer and System Sciences, 152(??):??, September 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000388.

Wu:2025:BNR

[3912] Yufeng Wu and Louxin Zhang. Bounding the number of reticulation events for displaying multiple trees in a phylogenetic network. Journal of Computer and System Sciences, 152 (??):??, September 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002500039X.

Anonymous:2025:EBf

[3913] Anonymous. Editorial Board. Journal of Computer and System Sciences, 152(??):??, September 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000480.

Anonymous:2025:S

[3914] Anonymous. September 2025. Journal of Computer and System Sciences, 152 (??):??, September 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Bilo:2025:FDR

[3915] Davide Bilò, Luciano Gualà, Stefano Leucci, and Luca Pepè Sciarria. Finding diameter-reducing shortcuts in trees. Journal of Computer and System Sciences, 153(??): ??, November 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000406.

Anonymous:2025:EBg

[3916] Anonymous. Editorial Board. Journal of Computer and System Sciences, 153(??):??, November 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000546.

Anonymous:2025:N

[3917] Anonymous. November 2025. Journal of Computer and System Sciences, 153 (??):??, November 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Xiao:2025:SCO

[3918] Ligang Xiao and Daowen Qiu. State complexity of one-way quantum finite automata together with classical states. Journal of Computer and System Sciences, 154(??): ??, December 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000418.

Jauregui:2025:CDC

[3919] Benjamin Jauregui, Pedro Montealegre, Diego Ramirez-Romero, and Ivan Rapaport. Compact distributed certification of geometric graph classes. Journal of Computer and System Sciences, 154(??):??, December 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000431.

Bok:2025:ASI

[3920] Jan Bok, Nikola Jedličková, Barnaby Martin, Pascal Ochem, Daniël Paulusma, and Siani Smith. Acyclic, star and injective colouring: a complexity picture for H-free graphs. Journal of Computer and System Sciences, 154(??):??, December 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000443.

Georgiou:2025:WGS

[3921] Konstantinos Georgiou and Xin Wang.

Weighted group search on the disk and improved lower bounds for priority evacuation. Journal of Computer and System Sciences, 154(??): ??, December 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000601.

Dondi:2025:FAT

[3922] Riccardo Dondi and Manuel Lafond. An FPT algorithm for timeline cover. Journal of Computer and System Sciences, 154(??): ??, December 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000613.

Bergougnoux:2025:PCL

[3923] Benjamin Bergougnoux, Nello Blaser, Michael Fellows, Petr Golovach, Frances Rosamond, and Emmanuel Sam. On the parameterized complexity of lineal topologies (depthfirst spanning trees) with many or few leaves. Journal of Computer and System Sciences, 154(??): ??, December 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000625.

Day:2025:EDS

[3924] Joel D. Day, Pamela Fleischmann, Maria Kosche, Tore Koß, Florin Manea, and Stefan Siemer. The edit distance to k-subsequence universality. Journal of Computer and System Sciences, 154(??):??, December 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000637.

Bodlaender:2025:CFF

[3925] Hans L. Bodlaender, Matthew Johnson, Barnaby Martin, Jelle J. Oostveen, Sukanya Pandey, Daniël Paulusma, Siani Smith, and Erik Jan van Leeuwen. Complexity framework for forbidden subgraphs IV: the Steiner Forest problem. Journal of Computer and System Sciences, 154(??): ??, December 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000649.

Anonymous:2025:EBh

[3926] Anonymous. Editorial Board. Journal of Computer and System Sciences, 154(??):??, December 2025. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000704.

Anonymous:2025:D

[3927] Anonymous. December 2025. Journal of Computer and System Sciences, 154 (??):??, December 2025. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Dyrseth:2026:CRV

[3928] Jakob Dyrseth and Paloma T. de Lima.

On the complexity of rainbow vertex colouring diametral path graphs. Journal of Computer and System Sciences, 155(??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000650.

Christodoulou:2026:MIM

[3929] Filippos Christodoulou, Pierluigi Crescenzi, Andrea Marino, Ana Silva, and Dimitrios M. Thilikos. Making the interval membership width of temporal graphs connected and bidirectional. Journal of Computer and System Sciences, 155(??): ??, February 2026. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000662.

Kiya:2026:SST

[3930] Hironori Kiya, Yuto Okada, Hirotaka Ono, and Yota Otachi. Sequentially swapping tokens: Further on graph classes. Journal of Computer and System Sciences, 155(??): ??, February 2026. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002500073X.

Didimo:2026:PAO

[3931] Walter Didimo, Siddharth Gupta, Philipp Kindermann, Giuseppe Liotta, Alexander Wolff, and Meirav Zehavi. Parameterized approaches to orthogonal compaction. Journal of Computer and System Sciences, 155(??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000741.

Wang:2026:JSS

[3932] Ru Wang and Yufei Tao. Join and subgraph sampling under degree constraints. Journal of Computer and System Sciences, 155(??): ??, February 2026. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000753.

Barcelo:2026:CDC

[3933] Pablo Barceló, Gaëlle Fontaine, Sylvain Salvati, and Sophie Tison. Corrections to “On the data complexity of consistent query answering over graph databases [Journal of Computer and System Sciences 88 (2017) 164–194]”. Journal of Computer and System Sciences, 155(??): ??, February 2026. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000765. See

[3310].

Georgiou:2026:TEA

[3934] Konstantinos Georgiou and Woojin Jang. Triangle evacuation of 2 agents in the wireless model and the power of choosing a starting point. Journal of Computer and System Sciences, 155(??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000777.

Bhagat:2026:EWO

[3935] Subhash Bhagat and Andrzej Pelc. Exploring wedges of an oriented grid by an automaton with pebbles. Journal of Computer and System Sciences, 155(??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000789.

Bockenhauer:2026:OKR

[3936] Hans-Joachim Böckenhauer, Ralf Klasing, Tobias Mömke, Peter Rossmanith, Moritz Stocker, and David Wehner. Online knapsack with removal and recourse. Journal of Computer and System Sciences, 155 (??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000790.

Wang:2026:GRD

[3937] Mu-Jiang-Shan Wang, Shu-Hao Xu, Jin-Cheng Jiang, Dong Xiang, and Sun-Yuan Hsieh. Global reliable diagnosis of networks based on selfcomparative diagnosis model and ggood-neighbor property. Journal of Computer and System Sciences, 155 (??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000807.

Huang:2026:ASU

[3938] Shengyu Huang, Chih-Hung Liu, and Daniel Rutschmann. Approximate selection with unreliable comparisons in sublinear time. Journal of Computer and System Sciences, 155(??): ??, February 2026. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000819.

Bredereck:2026:HTP

[3939] Robert Bredereck, Piotr Faliszewski, Michał Furdyna, Andrzej Kaczmarczyk, Joanna Kaczmarek, Martin Lackner, Christian Laußmann, Jörg Rothe, and Tessa Seeger. How to tamper with a Parliament: Strategic campaigns in apportionment elections. Journal of Computer and System Sciences, 155(??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0022000025000820.

Kutner:2026:TRD

[3940] David C. Kutner and Laura Larios-Jones. Temporal reachability dominating sets: Contagion in temporal graphs. Journal of Computer and System Sciences, 155(??): ??, February 2026. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0022000025000832.

Anonymous:2026:EBa

[3941] Anonymous. Editorial Board. Journal of Computer and System Sciences, 155(??):??, February 2026. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S002200002500090X.

Anonymous:2026:F

[3942] Anonymous. February 2026. Journal of Computer and System Sciences, 155 (??):??, February 2026. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Baeza-Yates:1989:TAW

[3943] Ricardo A. Baeza-Yates. A trivial algorithm whose analysis is not: a continuation. BIT, 29(3):378–394, September 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&
volume=29&issue=3&spage=378. See